### v3

# version 0.6

[…]
position=current page.north, angle=0, nodeanchor=north, vshift=-5mm, opacity=0.5, scale=1, contents=Draft; comments/suggestions welcome at nonstochastic.control@gmail.com
Introduction to Online Control
version 0.6
Elad Hazan    Karan Singh
Preface
This text presents an introduction to an emerging paradigm in control of dynamical systems and differentiable reinforcement learning called online nonstochastic control. The new approach applies techniques from online convex optimization and convex relaxations to obtain new methods with provable guarantees for classical settings in optimal and robust control.
The primary distinction between online nonstochastic control and other frameworks is the objective. In optimal control, robust control, and other control methodologies that assume stochastic noise, the goal is to perform comparably to an offline optimal strategy. In online nonstochastic control, both the cost functions as well as the perturbations from the assumed dynamical model are chosen by an adversary. Thus the optimal policy is not defined a priori. Rather, the target is to attain low regret against the best policy in hindsight from a benchmark class of policies.
This objective suggests the use of the decision making framework of online convex optimization as an algorithmic methodology. The resulting methods are based on iterative mathematical optimization algorithms and are accompanied by finite-time regret and computational complexity guarantees.
List of Symbols
General
=$\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}$ definition
${}\operatorname*{\arg\min}\{\}$ the argument minimizing the expression in braces
$[n]$ the set of integers $\{1,2,\ldots,n\}$
$\mathbf{1}_{A}$ the indicator function, equals one if $A$ is true, else zero
Algebra, Geometry and Calculus
${\mathbb{R}}^{d}$ $d$ dimensional Euclidean space
$\Delta_{d}$ $d$ dimensional simplex, $\{\sum_{i}\mathbf{x}_{i}=1,\mathbf{x}_{i}\geq 0\}$
$\mathbb{S}$ $d$ dimensional sphere, $\{\|\mathbf{x}\|=1\}$
$\mathbb{B}$ $d$ dimensional ball, $\{\|\mathbf{x}\|\leq 1\}$
${\mathbb{R}}$ real numbers
$\mathbb{C}$ complex numbers
$|A|$ determinant of matrix $A$
${\bf Tr}(A)$ trace of matrix $A$
Control
$f$ dynamics function
$\mathbf{x}_{t}\in{\mathbb{R}}^{d_{\mathbf{x}}}$ state at time $t$
$\mathbf{u}_{t}\in{\mathbb{R}}^{d_{\mathbf{u}}}$ control at time $t$
$\mathbf{w}_{t}\in{\mathbb{R}}^{d_{\mathbf{w}}}$ perturbation at time $t$
$\mathbf{y}_{t}\in{\mathbb{R}}^{d_{\mathbf{y}}}$ observation at time $t$
$A_{t},B_{t},C_{t}$ system matrices for linear dynamical system
$\gamma$ BIBO stabilizability $\|\mathbf{w}_{t}\|\leq 1\longrightarrow\|\mathbf{x}_{t}\|\leq\gamma$
$h$ Number of parameters in a policy class
$\kappa$ norm bound on policy parameters
Reinforcement learning
$S,A,\gamma,R,P$ Markov Decision Process with states $S$, actions $A$,
discount factor $\gamma$, rewards $R$ and transition probability matrix $P$
$R_{t}$ reward at time $t$
$\mathbf{P}_{ss^{\prime}}^{a}$ probability of transition from $s$ to $s^{\prime}$ given action $a$
$\mathbf{v}:S\to{\mathbb{R}}$ value function
$Q:S\times A\to{\mathbb{R}}$ $Q$ function
$\pi:S\to A$ policy
Optimization
$\mathbf{x}$ vectors in the decision set
$\mathcal{K}$ decision set
$\nabla^{k}f$ the $k$’th differential of $f$; note $\nabla^{k}f\in{\mathbb{R}}^{d^{k}}$
$\nabla^{-2}f$ the inverse Hessian of $f$
$\nabla f$ the gradient of $f$
$\nabla_{t}$ the gradient of $f$ at point $\mathbf{x}_{t}$
$\mathbf{x}^{\star}$ the global or local optima of objective $f$
$h_{t}$ objective value distance to optimality, $h_{t}=f(\mathbf{x}_{t})-f(\mathbf{x}^{\star})$
$d_{t}$ Euclidean distance to optimality $d_{t}=\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|$
$G$ upper bound on norm of subgradients
$D$ upper bound on Euclidean diameter
$D_{p},G_{p}$ upper bound on the $p$-norm of the subgradients/diameter
Part I Background in Control and RL
Chapter 1 Introduction
1.1 What is This Book About?
Control theory is the engineering science of manipulating physical systems to achieve desired functionality. Its history is centuries old, and it spans rich areas from the mathematical sciences. A brief list includes differential equations, operator theory and linear algebra, functional analysis, harmonic analysis, and more. However, it mostly does not concern modern algorithmic theory and computational complexity from computer science.
The advance of the deep learning revolution has brought to light the power and robustness of algorithms that are based on gradient methods. Deep learning pioneer and Turing award winner, professor Yann Lecun, is attributed the following quote from a Barbados workshop on deep learning theory, circa Feb 2020,
“Control = reinforcement learning with gradients.”
This book concerns this exact observation: how can we exploit the differentiable structure in natural physical environments to create efficient and robust gradient-based algorithms for control?
The answer we give is not to apply gradient descent to any existing algorithm. Rather, we rethink the paradigm of control from the start. What can be achieved in a given dynamical system, whether known or unknown, fully observed or not, linear or non-linear, and without assumptions on the noise sequence?
This is a deep and difficult question that we fall short of answering. We do, however, propose a different way of answering this question: we define the control problem from the point of view of an online player in a repeated game. This gives rise to a simple but powerful framework called nonstochastic control theory. The framework is a natural fit to apply techniques from online convex optimization, resulting in new, efficient, gradient based control methods.
1.2 The Origins of Control
The industrial revolution was in many ways shaped by the steam engine that came to prominence in the 18th century. One of its most critical components was introduced by the Scottish inventor James Watt: amongst the earliest examples of a control mechanism, the Centrifugal Governor. A major source of losses in efficiency of earlier designs was the difficulty in regulating the engine. Too much or too little steam could easily result in engine failure. The centrifugal governor was created to solve this problem.
Figure 1.1: A centrifugal governor. Source: Wikipedia
The centrifugal governor consists of a shaft connected to two metal balls that are allowed to rotate. Increased shaft speed causes the balls to spin faster, and the resulting centrifugal force causes the balls to lift. The balls are in turn attached to a pin that controls the amount of steam entering the engine, so as the balls spin faster, the pin lifts and limits the amount of fuel that can enter the system. On the other hand, if the balls spin slower, the pin lowers, allowing more steam to enter the system. This provides a natural, direct and proportional control to the system.
The impact of the centrifugal governor on the steam engine, and in turn on the the industrial revolution, illustrates the importance of the field of control in engineering and the sciences. Since the 18th century our world has changed quite a bit, and the importance of control has only grown.
The challenges of controlling systems has moved from the steam engine to design of aircraft, robotics and autonomous vehicles. The importance of designing efficient and robust methods remains at least as crucial as in the industrial revolution.
1.3 Formalization and Examples of a Control Problem
With the physical example of the centrifugal governor in mind, we give a mathematical definition of a control problem. The first step is to formally define a dynamical system below.
Definition 1.1 (Dynamical system).
A dynamical system from starting state $(\mathbf{x}_{1},\mathbf{y}_{1})$ is given by the following equations:
$\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g_{t}(\mathbf{x}_{t}).$
Here $\mathbf{x}_{t}\in{\mathbb{R}}^{d_{x}}$ is the state of the system, $\mathbf{y}_{t}\in{\mathbb{R}}^{d_{y}}$ is the observation, $\mathbf{u}_{t}\in{\mathbb{R}}^{d_{u}}$ is the control input, and $\mathbf{w}_{t}\in{\mathbb{R}}^{d_{w}}$ is the perturbation.
The function $g_{t}$ is a mapping from the state to the observed state, and the function $f_{t}$ is the transition function, given current state $\mathbf{x}_{t}$ and control input $\mathbf{u}_{t}$. The subscript indicates the relevant quantities at time step $t$.
Given a dynamical system, there are several possible objectives we can formulate. A generic formulation appears below, which captures many common control formulations, notably the optimal control problem and the robust control problem, which will be addressed in much more detail later in this book. In this problem formulation, the goal is to minimize the long-term horizon cost given by a sequence of cost functions.
Definition 1.2 (A generic control problem).
Given a dynamical system with starting state $(\mathbf{x}_{1},\mathbf{y}_{1})$, iteratively construct controls $u_{t}$ to minimize a sequence of loss functions $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$, $c_{t}:{\mathbb{R}}^{d_{y}\times d_{u}}\to{\mathbb{R}}$. This is given by the following mathematical program
$\min_{\mathbf{u}_{1:T}}\sum_{t}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$
$s.t.$ $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g_{t}(\mathbf{x}_{t}).$
At this point, there are many aspects of the generic problem that are left unspecified. These include:
  • Are the transition functions and observation functions known or unknown to the controller?
  • How accurately is the transition function known?
  • What is the mechanism that generates the perturbations $\mathbf{w}_{t}$?
  • Are the cost functions known or unknown ahead of time?
Different theories of control differ by the answers they give to the questions above, and these in turn give rise to different methods.
Instead of postulating a hypothesis about how our world should behave, we proceed to give an intuitive example of a control problem. We can then reason about the different questions, and motivate the nonstochastc control setting.
1.3.1 Example: Control of a Medical Ventilator
A recent relevant example of control is in the case of medical ventilators. The ventilator connects the trachea in the lungs to a central pressure chamber, which helps maintain positive-end expiratory pressure (PEEP) to prevent the lungs from collapsing and simulate healthy breathing in patients with respiratory problems.
The ventilator needs to take into account the structure of the lung to determine the optimal pressure to induce. Such structural factors include compliance, or the change in lung volume per unit pressure, and resistance, or the change in pressure per unit flow.
A simplistic formalization of the dynamical system of the ventilator and lung can be derived from the physics of a connected two-balloon system. The dynamics equations can be written as
$p_{t}=C_{0}+C_{1}v_{t}^{-1/3}+C_{2}v_{t}^{5/3},$
$v_{t+1}=v_{t}+\Delta_{t}u_{t},$
where $p_{t}$ is the observed pressure, $v_{t}$ is the volume of the lung, $u_{t}$ is the input flow (control), and $\Delta_{t}$ is the unit of time measured by the device precision. $C_{0}$, $C_{1}$, and $C_{2}$ are constants that depend on the particular system/lung-ventilator pair.
The goal of ventilator control is to regulate pressure according to some predetermined pattern. As a dynamical system, we can write $p_{t+1}=f(p_{t},u_{t},w_{t}),$ and we can define the cost function to be $c_{t}(p_{t},u_{t})=|p_{t}-p_{t}^{*}|^{2},$ where $p_{t}^{*}$ is the desired pressure at time $t$. The goal, then, is to choose a sequence of control signals $\{u_{t}\}_{t=1}^{T}$ to minimize the total cost over $T$ iterations.
In the figure below we see an illustration of a simple controller, called the PID controller, as it performs on a ventilator and lung simulator.
Figure 1.2: Performance of the PID controller on a mechanical ventilator
1.4 Simple Control Algorithms
The previous section detailed the importance of the control problem on the example of a critical care medical device. In this section we detail methods for controlling a medical ventilator, or indeed any physical device that can be modelled as a dynamical system.
The following methods are simplistic but still widely used. For example, PID algorithm is still used in the majority of medical ventilators today. In later sections we describe a modern optimization-based framework that yields more sophisticated algorithms that are capable of better performance.
1.4.1 The Bang-Bang Controller
Bang-bang control is one of the simplest control strategies. Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a pre-specified dynamic range $[x_{\min},x_{\max}]$ (that can vary with time), and where the scalar control $u_{t}$ input must at all times be within the interval $[u_{\min},u_{\max}]$. A bang-bang controller chooses the control input at time $t$ as
$u_{t}=\mathbf{1}_{\{x_{t}<x_{\min}\}}\times u_{\max}+\mathbf{1}_{\{x_{t}>x_{\max}\}}\times u_{min}.$
A bang-bang controller saturates its inputs to extreme values in an attempt to confine the state of the system to a desirable range. For example, in the context of mechanical ventilation, one could choose $[p_{t}^{*}-\Delta p,p_{t}^{*}+\Delta p]$ as a reasonable region around the prescribed pressure $p^{*}_{t}$ as a desirable range for $p_{t}$. A bang-bang controller in such a case would let the oxygen in at the maximum possible rate whenever the pressure drops below $p^{*}_{t}-\Delta p$.
This discussion highlights two unsatisfactory aspects of the bang-bang control scheme: one, it ignores the cost associated with acting with the extreme values of the control input; two, it produces oscillatory and rapidly changing behavior in the control state.
1.4.2 The PID Controller
A more sophisticated control algorithm is the PID control method. Consider a control method that chooses the control to be a linear function of the observed states, i.e.
$\mathbf{u}_{t}=\sum_{i=0}^{k}{A}_{i}\mathbf{x}_{t-i},$
for some coefficients $A_{i}$, where we consider only the $k+1$ most recent states. One of the most useful special cases of this linear family is the PID controller.
A special sub-family of linear controllers is that of linear controllers that optimize over a pre-defined basis of historical observations. More precisely,
  1. 1.
    Proportional Control (e.g. Centrifugal Governor): $\mathbf{u}_{t}=\alpha_{0}\mathbf{x}_{t}.$
  2. 2.
    Integral Control: $\mathbf{u}_{t}=\beta\sum_{i<t}\mathbf{x}_{t-i}.$
  3. 3.
    Derivative Control: $\mathbf{u}_{t}=\gamma(\mathbf{x}_{t}-\mathbf{x}_{t-1}).$
The PID controller, named after the nature of the three components that constitute it, specifies three coefficients $\alpha,\beta,\gamma$ to generate a control signal. In that sense it is extremely sparse, which from a learning theoretic perspective is a good indication of generalization.
Indeed, the PID controller is the method of choice for numerous applications in engineering, including control of medical ventilators.
1.5 Classical Theory: Optimal and Robust Control
Classical theories of control differ over the generic control problem formulation 1.2 primarily in one significant aspect: the perturbation model. A rough divide between the two leading theories of optimal control and robust control is as follows.
Optimal control postulates a probabilistic model for the perturbations, which are typically i.i.d. draws from a Gaussian distribution. However, other types of distributions can be considered as well. As such, the control problem can be re-phrased in a more precise stochastic mathematical program,
$\min_{\mathbf{u}_{1:T}}\mathop{\mbox{\bf E}}_{\mathbf{w}_{1:T}}\left[\sum_{t}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})\right]$
$s.t.$ $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g(\mathbf{x}_{t}).$
In contrast, robust control theory plans for the worst case noise from a given set of constraints, denoted $\mathcal{K}$. Thus, the robust control problem can be rephrased as the following mathematical program,
$\min_{\mathbf{u}_{1:T}}\max_{\mathbf{w}_{1:T}\in\mathcal{K}}\left[\sum_{t}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})\right]$
$s.t.$ $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g(\mathbf{x}_{t}).$
The two formulations have been debated for decades, with pros and cons for each side. Optimal control focuses on the average case and many times yields simpler formulations for optimization, resulting in more efficient methods. Robust control, on the other hand, allows for more difficult noise models and some amount of model misspecification. However, it is too pessimistic for many natural problems.
Nonstochastic control theory takes the best of both worlds, simultaneously granting robustness to adversarial noise and allowing the use of optimistic behavior when appropriate, while providing computationally efficient methods.
1.6 The Need for a New Theory
Consider the problem of flying a drone from source to destination subject to unknown weather conditions. The aerodynamics of flight can be modeled sufficiently well by time-varying linear dynamical systems, and existing techniques are perfectly capable of doing a great job for indoor flight. However, the wind conditions, rain and other uncertainties are a different story. Certainly the wind is not an i.i.d. Gaussian random variable! Optimal control theory, which assumes this zero-mean noise, is therefore overly optimistic and inaccurate for our problem.
The designer might resort to robust control theory to account for all possible wind conditions, but this is overly pessimistic. What if we encounter a bright sunny day after all? Planning for the worst case would mean slow and fuel-inefficient flight.
We would like an adaptive control theory which allows us to attain the best of both worlds: an efficient and fast flight when the weather permits, and a careful drone maneuver when this is required by the conditions. Can we design a control theory that will allow us to take into account the specific instance perturbations and miss-specifications, and give us finite time provable guarantees? This is the subject of online non-stochastic control!
1.6.1 Online Nonstochastic Control Theory
This book is concerned with nonstochastic noise. When dealing with non-stochastic, i.e. arbitrary or even adversarial perturbations, the optimal policy is not clear a priori. Rather, an optimal policy for the observed perturbations can be determined only in hindsight. This is a significant difference from optimal and robust control theory, where the mapping from states to actions can completely be predetermined based upon the noise model.
Since optimality is defined only in hindsight, the goal is not predetermined. We thus shift to a different performance metric borrowed from the world of learning in repeated games – regret. Specifically, we consider regret with respect to a reference class of policies. In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$, and then observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ according to an adversarially chosen loss function. Let $\Pi=\{\pi:\mathbf{x}\to\mathbf{u}\}$ be a class of policies. The regret of the controller with respect to $\Pi$ is defined as follows.
Definition 1.3.
Given a dynamical system, the regret of an online control algorithm $\mathcal{A}$ over $T$ iterations with respect to a class of policies $\Pi$ is given by:
$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}$ $\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}))\right),$
where $\mathbf{u}_{t}={\mathcal{A}}(\mathbf{x}_{t})$ are the controls generated by ${\mathcal{A}}$, and $\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}$ are the counterfactual state sequence and controls under the policy $\pi$, i.e.
$\mathbf{u}_{t}^{\pi}$ $=\pi(\mathbf{x}_{t}^{\pi})$
$\mathbf{x}_{t+1}^{\pi}$ $=f_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi},\mathbf{w}_{t}).$
Henceforth, if $T,\Pi$ and ${\mathcal{A}}$ are clear from the context, we drop these symbols when defining $\mathrm{{regret}}$.
At this point the reader may wonder why we should compare to a reference class, as opposed to all possible policies. There are two main answers to this question:
  1. 1.
    First, the best policy in hindsight may be very complicated to describe and reason about. This argument was made by the mathematician Tyrell Rockafellar, see the bibliographic section for more details.
  2. 2.
    Secondly, it can be shown that it is impossible, in general, to obtain sublinear regret vs. the best policy in hindsight. Rather, a different performance metric called competitive ratio can be analyzed, as we do later in the book.
The non-stochastic control problem can now be stated as finding an efficient algorithm that minimizes the worst-case regret vs. meaningful policies classes that we examine henceforth. It is important to mention that the algorithm does not have to belong to the comparator set of policies! Indeed, in the most powerful results, we will see that the algorithm will learn and operate over a policy class that is strictly larger than the comparator class.
1.6.2 A New Family of Algorithms
The policy regret minimization viewpoint in control leads to a new approach for algorithm design. Instead of a priori computation of the optimal controls, online convex optimization suggests modifying the controller according to the costs and dynamics, in the flavor of adaptive control, in order to achieve provable bounds on the policy regret.
As a representative example, consider the Gradient Perturbation Controller described in Algorithm 1. The algorithm maintains matrices that are used to create the control as a linear function of the past perturbations, as per line 3. These parameters change over time according to a gradient-based update rule, where the gradient is calculated as the derivative of a counterfactual cost function with respect to the parametrization, as per line 5.
1:  Input: $h$, $\eta$, initial parameters ${M}_{1:h}^{1}$, dynamics $f$.
2:  for $t$ = $1\ldots T$ do
3:      Use Control $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=1}^{h}{{M}}_{i}^{t}\mathbf{w}_{t-i}$
4:     Observe state $\mathbf{x}_{t+1}$, compute perturbation $\mathbf{w}_{t}=\mathbf{x}_{t+1}-f(\mathbf{x}_{t},\mathbf{u}_{t})$
5:      Construct loss $\ell_{t}({M}_{1:h})=c_{t}(\mathbf{x}_{t}({M}_{1:h}),\mathbf{u}_{t}({M}_{1:h}))$Update ${M}_{1:h}^{t+1}\leftarrow{M}_{1:h}^{t}-\eta\nabla\ell_{t}({M}_{1:h}^{t})$
6:  end for
Algorithm 1 Gradient Perturbation Controller(GPC) - simplified version
The version given here assumes that the time-invariant dynamics $f$ is known. The GPC computes a control which is a linear function of past perturbations. It modifies the linear parameters according to the loss function via an iterative gradient method. In Chapter 7, we prove that this algorithm attains sublinear regret bounds for a large class of time-varying and time-invariant dynamical systems. In later chapters we explore extensions to nonlinear dynamical systems and unknown systems, which lie at the heart of nonstochastic control theory.
1.7 Bibliographic Remarks
Control theory is a mature and active discipline with both deep mathematical foundations and a rich history of applications. In terms of application, self-regulating feedback mechanisms that limit the flow of water, much like the modern float valve, were reportedly used in ancient Greece over two thousand years ago. The first mathematical analysis of such systems appeared in a paper [Max68] by James Clerk Maxwell on the topic of governors, shortly after Maxwell had published his treatise on the unification of electricity and magnetism. For an enlightening historical account, see [FCZI03].
For a historical account on the development of the PID controller see [Ben93]. The theoretical and practical properties of this controller are spelled out in detail in [ÅH95].
There are many excellent textbooks on control theory, which we point to in later chapters. The book of [Ber07] offers an extensive coverage of topics from the viewpoint of dynamic programming. The texts [Ste94, ZDG96a], on the other hand, delve deep into results on linear control, the most well-understood branch of control theory. The excellent text of [SL91] gives a thorough introduction to adaptive control theory.
Online nonstochastic control theory and the Gradient Perturbation Control algorithm were proposed in [ABH+19]. A flurry of novel methods and regret bounds for online nonstochastic control have appeared in recent years, many of which are surveyed in [HS21]. Notable amongst them are logarithmic regret algorithms [AHS19, FS20], nonstochastic control for unknown systems [HKS20], black-box online nonstochastic control [CHK+18], and nonstochastic control with partial observability [SSH20].
The theory builds heavily on recent developments in decision making and online learning, notably the framework of online convex optimization [Haz16], and especially its extensions to learning with memory [AHM13].
Chapter 2 Dynamical systems
In this chapter we return to the general problem of control from definition 1.2, i.e. choosing a sequence of controls to minimize a sequence of cost functions. Before moving to algorithms for control, it is important to reason about what kind of solutions are even possible.
For this reason we consider several natural examples of dynamical systems, and reason about what constitutes desirable behavior. This will help us understand the kinds of observations and control objectives that are feasible and natural in real applications.
2.1 Examples of Dynamical Systems
2.1.1 Medical Ventilator
In the task of mechanical ventilation, the control algorithm dictates the amount of oxygen inflow $u_{t}$ (in $\mathrm{m}^{3}/s$) into a patient’s lungs. This inflow determines the volume of the lungs $v_{t}$ via a simple linear relation,
$v_{t+1}=v_{t}+\Delta u_{t},$
where $\Delta$ is the time elapsed between successive actions of the discrete-time control algorithm. Being difficult to directly measure or estimate, the quantity $v_{t}$ is a hidden or latent variable from the controller’s perspective. Instead, the controller observes the air pressure $p_{t}$. We give a simplified physical model relating the pressure and the volume. This model, given below, can be derived by first principles from stress-strain relations, treating the trachea-connected left and right lungs as situated in a two-balloon experiment.
$p_{t}=C_{0}+C_{1}v_{t}^{-1/3}+C_{2}v_{t}^{5/3}$
The objective of the control algorithm is to track a physician-prescribed ideal pressure trajectory $p^{*}_{t}$ as closely as possible, say as measured by in squared loss $c_{t}(p_{t},u_{t})=|p_{t}-u_{t}|^{2}+|u_{t}|^{2}$, while limiting the maximum instantaneous airflow to minimize patient discomfort. See Figure 1.2 for a schematic representation of the airflow when a lung is connected to a mechanical ventilator.
Figure 2.1: A schematic of the respiratory circuit from [SZG+21].
2.1.2 Double Integrator
The double integrator is a common example of a linear dynamical system. In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics. The state $\mathbf{x}_{t}$ is 2-dimensional, with the first coordinate indicating the position of a point object and the second denoting its velocity. The control algorithm decides the force $u_{t}\in{\mathbb{R}}$ that is applied to the object at each time step. The dynamics equation can be written as
$\mathbf{x}_{t+1}=\begin{bmatrix}1&\Delta\\ 0&1\end{bmatrix}\mathbf{x}_{t}+\begin{bmatrix}0\\ 1\end{bmatrix}u_{t}+\mathbf{w}_{t}.$
The position of the object is linearly related to its velocity via the parameter $\Delta$, which measures the time resolution the controller operates in. Optionally, the state is also affected by external perturbations $\mathbf{w}_{t}$ outside the influence of the controller.
The goal here is to ensure that the point object arrives and then comes to rest at some coordinate $x^{*}\in\mathbb{R}$ as fast as possible, while also not expending too much energy. The cost function is naturally expressed as
$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}x^{*}\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
Figure 2.2: The position and velocity visualized for a double integrator.
2.1.3 Pendulum Swing-up
The swing-up pendulum is one of the simplest examples of rotational kinematics. It corresponds to a point object of mass $m$ mounted at the end of a massless rod of length $l$. The other end of the rod is fixed to a pivot. Being vertically aligned, the system is subject to gravity with gravitational acceleration $g$. The controller is in charge of applying a time-varying torque $u_{t}\in\mathbb{R}$ at each time step. Given the vertical angle $\theta_{t}$ of the rod, the physics of this system may be summarized as
$\theta_{t+1}=\theta_{t}+\Delta\dot{\theta}_{t},\qquad\dot{\theta}_{t+1}=\dot{\theta}_{t}+\Delta\frac{u_{t}-mg\sin\theta}{ml^{2}},$
where $\Delta$ is the time resolution of the controller. Note that this dynamical equation is non-linear in the state $\mathbf{x}_{t}=[\theta_{t},\dot{\theta_{t}}]^{\top}$ due to the presence of the $\sin\theta$ term in the evolution of $\dot{\theta_{t}}$.
Figure 2.3: A pivoted massless rod connected to a point object with mass $m$. The point object is to subject to gravitational acceleration $g$. Source: Wikipedia
The objective here is to get the rod to rest at a vertically upright position at an angle of $\pi$, using as little torque as possible throughout the process. This may be expressed via the cost function
$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}\pi\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
2.1.4 Aircraft Dynamics
A final example of a dynamical system we consider is the longitudinal (vertical tilt) motion of an aircraft that is regulated by automated controllers. We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system. The 2-dimensional control input $\mathbf{u}_{t}$ corresponds to the elevator angle and the thrust. The disturbances correspond to the velocity of the wind along the body axis and perpendicular to it.
The dynamics equation below has matrices whose numerical values were computed using the dynamics equation for an actual aircraft. These numerical values can be determined by the mass, length, and other physical properties of the aircraft.
$\mathbf{x}_{t+1}=$ $\begin{bmatrix}-0.003&0.039&0&-0.322\\ -0.065&-0.319&7.74&0\\ 0.020&-0.101&-0.429&0\\ 0&0&1&0\end{bmatrix}\mathbf{x}_{t}+\begin{bmatrix}0.01&1\\ -0.18&-0.04\\ -1.16&0.598\\ 0&0\end{bmatrix}\mathbf{u}_{t}$
$+\begin{bmatrix}-0.003&0.039&0&-0.322\\ -0.065&-0.319&7.74&0\end{bmatrix}\mathbf{w}_{t}$
Figure 2.4: Longitudinal control in an aircraft. Source: Wikipedia
The quantities of interest for assessing the performance of the controller are aircraft speed, climb rate, each of which is a linear transformation of the state, and the applied thrust. The corresponding cost is
$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\begin{bmatrix}1&0&0&0\\ 0&-1&0&7.74\end{bmatrix}\mathbf{x}_{t}-\mathbf{x}_{t}^{*}\right\|^{2}+\left\|\begin{bmatrix}0\\ 1\end{bmatrix}\mathbf{u}_{t}\right\|^{2}.$
2.1.5 Epidemiological Models
The COVID-19 pandemic marked an increased public interest in policy discussions on the effectiveness of certain interventions as a means of controlling the spread of infection. Epidemiological models forecast the spread of an infectious disease via structural models and differential equations. We present the discrete-time analogue of the most basic of such models – the SIR model. Let $S_{t}$, $I_{t}$, and $R_{t}$ denote the susceptible, infected, and recovered population sizes. Let $u_{t}$ be the number of vaccines administered at time $t$.
$S_{t+1}$ $=S_{t}-\beta S_{t}I_{t}-\alpha u_{t}$
$I_{t+1}$ $=I_{t}+\beta S_{t}I_{t}-\gamma I_{t}$
$R_{t+1}$ $=R_{t}+\gamma I_{t}.$
Here the coefficient $\beta$ measures how communicable the infection is, $\gamma$ measures how quickly an infected person recovers, and $\alpha$ represents the efficacy of the vaccine in preventing infections.
Figure 2.5: Trajectory of a continuous-time SIR model. Source: Wikipedia
2.2 Solution Concepts for Dynamical Systems
Before considering the control, or manipulation, of a dynamical system, we must first reason about what the goal should even be. In this section, we discuss various solution concepts for a dynamical system, including equilibria, stability, stabilizability, and controllability.
Given a dynamical system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t})$, the deterministic-equivalent dynamical system, defined via the dynamics $\mathbf{x}_{t+1}=f^{\prime}(\mathbf{x}_{t},\mathbf{u}_{t})=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{0})$, models the evolution of the dynamical system in the absence of perturbations. Further, the autonomous dynamical system $\mathbf{x}_{t+1}=f^{\prime\prime}(\mathbf{x}_{t})=f(\mathbf{x}_{t},\mathbf{0},\mathbf{0})$ describes state transitions of the dynamical system under the absence of both control inputs and perturbations.
2.2.1 Equilibrium of Dynamical Systems
An example of an equilibrium is a ball standing at rest at the bottom of a pit. If the ball were to be slightly moved, it would return to the bottom of the pit. Thus, this equilibrium is stable and attractive.
If the ball lies at rest on the top of a mountain, pushing the ball in any direction would cause it to fall. This is an equilibrium which is unstable. A formal definition is as below.
Definition 2.1.
A point $\mathbf{x}$ is an equilibrium point of an autonomous dynamical system $f$ if and only if $\mathbf{x}=f(\mathbf{x})$. An equilibrium point $\mathbf{x}$ is stable and attractive if and only if $\mathbf{x}=\lim_{t\rightarrow\infty}f^{t}(\mathbf{x}_{0})$ for every initial state $\mathbf{x}_{0}$ in some neighborhood of $\mathbf{x}$. If the neighborhood includes the entire set of states, we say that the equilibrium is globally attractive.
Another example is given by a linear transformation in Euclidean space. Consider the system
$\mathbf{x}_{t+1}={A}\mathbf{x}_{t}.$
Clearly 00 is an equilibrium point as ${A}0=0$ for all ${A}$. If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vectors converges to the null veector. Thus 00 is a globally attractive, stable equilibrium.
In full generality, determining if a system has an equilibirum is an intractable problem, as we show in this chapter.
2.2.2 Stabilizability and Controllability
Two fundamental properties of dynamical systems are stabilizability and controllability, as defined next.
Definition 2.2.
A Dynamical System is stabilizable if and only if, for every initial state, there exists a sequence of controls that drives the deterministic-equivalent system to the zero state. That is, $\lim_{t\rightarrow\infty}{\mathbf{x}_{t}}=0$.
Definition 2.3.
A Dynamical System is controllable if and only if, for every initial and target state, there exists a sequence of controls that drives the deterministic-equivalent system to the target state.
Note that controllability clearly implies stabilizability as the zero state is one possible target state.
In the next section, we discuss the computational difficulties that arise in determining if a system is stabilizable and/or controllable.
2.2.3 Quantitative Definitions of Stabilizability
The quantified notion of stabilizability we introduce below is more refined than driving the system to zero. We require the ability to drive the system to zero under adversarial bounded noise in the dynamics. In classical control theory this is called BIBO-stability, which stands for Bounded-Input-Bounded-Output.
This definition can apply to any dynamical system and is an important concept in non-stochastic control.
Definition 2.4.
A control policy $\pi$ for a dynamical system $f$ is $\gamma$-stabilizing if and only if the following holds. For every starting state $\|\mathbf{x}_{0}\|\leq 1$ and every sequence of adversarially chosen bounded perturbations $\mathbf{w}_{1:T}$ such that $\|\mathbf{w}_{t}\|\leq 1$, let $\mathbf{x}_{t}^{\pi}$ be the state obtained by using policy $\pi$,
$\mathbf{x}_{t+1}^{\pi}=f(\mathbf{x}_{t}^{\pi},\pi(\mathbf{x}_{t})).$
Then the norm of the system’s state at every iteration is bounded by an absolute constant independent of $T$, i.e.
$\forall t\ ,\ \|\mathbf{x}_{t}^{\pi}\|\leq\gamma.$
A dynamical system is said to be $\gamma$-stabilizable if and only if it admits a $\gamma$-stabilizing control policy.
2.3 Intractability of Equilibrium, Stabilizability, and Controllability
Before formally proving the computational hardness of determining various solution concepts of dynamical systems, we start with a few intuitive and convincing examples of the complexities that arise. Consider the following dynamical system over the complex numbers
$z_{t+1}=f_{c}(z_{t})=z_{t}^{2}+c.$
The Julia set is now defined as
$J_{c}=\{z\in C\ :\ \limsup_{t\to\infty}{|f_{c}^{t}(z)|}<\infty\},$
where $f^{t}_{c}(x)=f_{c}(f_{c}(\dots f_{c}(x)))$ denotes $t$ successive applications of the function $f_{c}$. In other words, it is the set of all complex numbers $Z$ for which the system has a stable attractive equilibrium (which is not $\infty$). The boundary of the Julia set is very hard to determine, and it is known that determining if a ball around a given point belongs to the Julia set is a computationally undecidable problem, i.e. there is no Turing machine that decides it in finite time! (see e.g. [BY08]) The complexity that arises near the boundary of the Julia set is depicted in figure 2.7.
Another compelling visual evidence of the complexity of equilibrium in dynamical systems is the Lorenz attractor. This is a strange equilibrium resulting from simple dynamics that can be described using the following three equations over three variables, called the Lorenz equations [Lor63],
$x_{t+1}$ $=\sigma(y_{t}-x_{t})$
$y_{t+1}$ $=x_{t}(\rho-z_{t})-y_{t}$
$z_{t+1}$ $=x_{t}y_{t}-\beta z_{t}.$
The 3D trajectories of this system are attracted to a 2D manifold with two “wings” that resemble a butterfly as depicted in figure 2.7. For formal properties of this attractor set, see [Str14, Tuc99]. This demonstrates that even studying a solution concept by itself is a formidable task.
For a formal proof of attractiveness in the continuous case, see section 9.2 of [Str14] or the original paper of Lorenz [Lor63].
Figure 2.6: Julia set
Figure 2.7: Lorenz attractor
2.3.1 Computational Hardness of Stabilizability
Consider the following problem: given a dynamical system, can we determine if it stabilizable, i.e. for every possible starting state does there exist a choice of control inputs that drives the system to the null state? We now show that this decision problem is computationally hard.
Theorem 2.5.
For a family of dynamical systems described as polynomials with integer coefficients, determining the stabilizability of any member is NP-hard.
We show that this problem can be reduced from the 3SAT problem, a known NP-hard problem. This implies that determining the stabilizability of a dynamical system is NP-hard in general.
Proof.
We first claim that checking if a given polynomial has a zero is NP-hard. This is achieved by a reduction from 3SAT. Recall the 3SAT problem: given a Boolean formula in the conjunctive normal form over $n$ variables and with $m$ clauses where each clause is restricted to three literals, such as
$F(x_{1:n})=(x_{1}\lor\neg x_{2}\lor x_{3})\land(\neg x_{1}\lor\neg x_{2}\lor x_{3})\land(x_{1}\lor x_{3}\lor x_{4}),$
decide if there exists a satisfying Boolean assignment.
Given such a formula $F$, consider the following polynomial $p_{F}(\mathbf{u})$. For each clause $c$ in $F$, construct a degree-3 polynomial as
$g_{x_{i}\lor x_{j}\lor x_{k}}(\mathbf{u})$ $\rightarrow(1-u_{1})\cdot(1-u_{2})\cdot(1-u_{3})$
$g_{\neg x_{i}\lor x_{j}\lor x_{k}}(\mathbf{u})$ $\rightarrow u_{1}\cdot(1-u_{2})\cdot(1-u_{3})$
$g_{\neg x_{i}\lor\neg x_{j}\lor x_{k}}(\mathbf{u})$ $\rightarrow u_{1}\cdot u_{2}\cdot(1-u_{3})$
$g_{\neg x_{i}\lor\neg x_{j}\lor\neg x_{k}}(\mathbf{u})$ $\rightarrow u_{1}\cdot u_{2}\cdot u_{3},$
and the degree-6 polynomial for the entire formula as
$p_{F}(\mathbf{u})=\sum_{i=1}^{n}(u_{i}(1-u_{i}))^{2}+\sum_{j=1}^{m}(g_{c_{j}}(\mathbf{u}))^{2}.$
For a disjunctive clause to be false, all its literals must evaluate to false. Therefore, given any satisfying Boolean assignment $x_{1:n}$ for the formula $F$, it holds that $p_{F}(x_{1:n})=0$.
Similarly, given any $\mathbf{u}\in{\mathbb{R}}^{n}$ such that $p_{F}(\mathbf{u})=0$, it must be that $\mathbf{u}_{i}\in\{0,1\}$ for all $i$. Furthermore, since $p_{F}$ is a sum of squares, the Boolean assignment $x_{1:n}=\mathbf{u}$ must satisfy every clause in $F$. This shows the hardness of determining the existence of a zero for polynomials.
Now, consider the scalar-state dynamical equation given by
$x_{t+1}=x_{t}(1-\varepsilon+p_{F}(\mathbf{u}_{t}))+p_{F}(\mathbf{u}_{t}),$
with $p$ being the polynomial family we just constructed, and $\varepsilon$ being an arbitrarily small constant. For the system to reach the null state from an arbitrary start state, we would need to verify if there is an assignment of $\mathbf{u}$ such that $p_{F}(\mathbf{u})=0$, which we just proved is NP-hard. ∎
The proof establishes that even for a dynamical system that is a polynomial of degree six, checking for stabilizability is NP-hard. Exercise 2.5 asks the reader to strengthen this result and show that checking the stabilizability of dynamical systems describable as polynomials of degree four is also NP-hard.
2.4 Bibliographic Remarks
The text [Ted20] offers several examples and characterizations of dynamical systems useful for robotic manipulation. A more mathematically involved introduction appears in [Str14].
[Ada20] discusses further aspects of simulations used in the early days of the COVID-19 pandemic. See [SZG+21] for examples of control algorithms used for non-invasive mechanical ventilation.
The Lorenz equations were introduced to model the phenomenon of atmospheric convection [Lor63]. For detailed information about the mathematical properties of the Lorenz attractor, see [Str14, Tuc99]. For a rigorous study of Julia sets and the Mandelbrot set, see [BY08].
Nonlinear dynamical systems admit a variety of stability criteria; we chose to define a few that are useful for further developments in the text. A notable omission from our discussion is that of local and global stability; a celebrated result [Lya92] proves that the existence of certain potential functions is sufficient to guarantee global stability, thereby obviating the need for a trajectory-based analysis. An interested reader might look to [Kha15] for an expanded discussion on how these stability criteria compare.
An excellent survey on the computational complexity of control is [BT00]. For a more recent and stronger result on stabilizability, see [Ahm12]. The NP-hardness theorem in this chapter appeared in [Ahm16]. The computational complexity of mathematical programming – both constrained and unconstrained – is commented upon in [Haz19].
2.5 Exercises
2.1. Propose a computationally efficient representation for the problem of stabilizability. Prove that for your proposal, the problem of determining whether a given dynamical system is stabilizable belongs to the class NP. Thus, together with the material of this chapter, conclude that stabilizability is NP-complete.
2.2. Strengthen Theorem 2.5 to show that checking whether a dynamical system that is a polynomial of degree four is stabilizable is NP-hard. Hint: Reduce from the NP-hard problem of MAX-2-SAT instead of 3-SAT.
Chapter 3 Markov Decision Processes
The previous chapters described and motivated the problem of control in dynamical systems. We have seen that in full generality, even simple questions about dynamical systems such as stabilizability and controllability are computationally hard. In this chapter, we consider a different representation of the control problem for which computationally efficient methods exist.
This chapter introduces a setting of equal generality as that of the control problem from the previous chapter: reinforcement learning in Markov Decision Processes (MDP). However, the representation of states and dynamics in a MDP is more explicit and verbose. In such case, we can design efficient methods aimed at optimality given the explicit state and transition representations.
After introducing the reinforcement learning model and Markov Decision Processes, we discuss an important algorithmic technique that will be useful throughout this book: dynamic programming, and the accompanying characterization of optimal policies via the Bellman equation.
3.1 Reinforcement Learning
Reinforcement learning is a sub-field of machine learning that formally models the setting of learning through interaction in a reactive environment. This is in contrast to learning from examples, as in supervised learning, or inference according to a probabilistic model, as in Bayesian modeling. In reinforcement learning, we have an agent and an environment. The agent observes its position (or state) in the environment and takes actions that transition it to a new state. The environment looks at the agent’s state and hands out rewards based on a hidden set of criteria.
Figure 3.1: Source: Reinforcement Learning: An Introduction. Sutton & Barto, 2017.
Typically, the goal of reinforcement learning is for the agent to learn behavior that maximizes the total reward it receives from the environment. This methodology has led to some notable successes: machines have learned how to play Atari games, and how to beat the highest-rated human players at Go.
A crucial assumption for the reinforcement learning model is as follows:
The Markovian assumption: Future states, observations, and rewards are independent of the past given the current state.
The reader can reason about situations where such an assumption is reasonable. For example, in the inverted pendulum setting described in section 2.1.3, the assumption clearly holds. It is not significant for the control problem how a certain state, describing the angle and the angular velocity, was reached, in order to optimally control the pendulum from this state onwards.
Other problems may call into question this assumption. For example, in the medical ventilation problem, it may very well be important how a certain state was reached, as future treatment may depend on the medical history of the patient.
It can be argued that the Markovian assumption can always be made to work, by adding more features to the state. In the extreme, the state can contain the entire history, which makes the assumption true by default. However, this makes the state representation inefficient.
3.1.1 A Classical Example: Bubbo’s World
Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel & Norvig [RN02]. In this example, a robot starts in the bottom leftmost square, and its goal is to reach the $+1$ reward.
Figure 3.2: Source: [RN02].
In this example, the robot has controls that allow it to move left / right / forward / backward. The controls, however, are not reliable, and with $0.8$ probably the robot moves to wherever it intends. With the remaining $0.2$ probability the robot moves sideways, $0.1$ probability to each side.
Notice that the Markovian assumption clearly holds for this simple example. What is an optimal solution to this example? A solution concept for a reinforcement learning problem is a mapping from state to action. This mapping is called a “policy”, and it depends on both the structure of the problem as well as on the rewards.
For example, if the rewards elsewhere are slightly negative, Bubbo wants to reach the terminal state with a persistent $+1$ reward, as given in the following figure.
[…]
Figure 3.3: The optimal policy (mapping from state to action) with reward for each step taken being -0.04.
On the other hand, if the rewards are very negative, Bubbo just wants to reach any terminal state as soon as possible, as in the following.
[…]
Figure 3.4: Here, we change the reward of each step to be -2. The optimal policy changes: now the robot is more concerned with ending the game as quickly as possible rather than trying to get to the +1 reward.
We proceed to describe a formal model for capturing this and any other reinforcement learning problem.
3.2 Markov Decision Processes
In this chapter we describe the general and formal model for reinforcement learning: Markov decision processes.
To ease the exposition, we start with a review of Markov Chains, a.k.a. Markov Processes, then proceed to more general Markov Reward Proccesses, and conclude with the full Markov Decision Process definition.
3.2.1 Markov Chains
A Markov Chain (or Markov Process) is a model of a memoryless stochastic process as a weighted directed graph. This is the first layer in abstraction before allowing for actions, as it only describes a process without exogenous inputs.
It consists of a set of states and a transition matrix describing the probability of moving from one state to another. The probabilities of moving between states are the weights on the edges of the graph. Figure 3.2.1 describes a Markov Chain corresponding to a weather process.
RainSunny dayCloudySnow0.10.90.50.10.40.50.51
$P=\begin{pmatrix}0.9&0.1&0&0\\ 0.5&0&0.4&0.1\\ 0&0.5&0&0.5\\ 0&0&0&1\\ \end{pmatrix}.$
Denote the distribution over states at time $t$ by $\mathbf{p}_{t}$. Then it can be seen that
$\mathbf{p}_{t+1}=\mathbf{p}_{t}P=\mathbf{p}_{0}P^{t},$
where $\mathbf{p}_{0}$ is the initial state distribution represented as a vector.
A central object in the study of Markov Chains is stationary distributions.
Definition 3.1.
A stationary distribution of a Markov Process is a limiting probability distribution
$\boldsymbol{\pi}=\lim_{t\to\infty}\mathbf{p}_{0}P^{t}.$
A stationary distribution satisfies
$\boldsymbol{\pi}=\boldsymbol{\pi}P.$
A stationary distribution does not always exist nor is it always unique. Existence, uniqueness, and convergence to stationarity are the subject of Ergodic theory, a rich and well-studied branch of mathematics. For our discussion, we are interested in sufficient conditions for stationary distribution to exist. They are:
  1. 1.
    Irreducibility. A Markov Chain is irreducible if and only if $\forall i,j\;\exists t\;s.t.\;P_{ij}^{t}>0$. In other words, every state is reachable from every other state.
  2. 2.
    Aperiodicity. A Markov Chain is a-periodic if and only if $d(s_{i})=1$ for all states $s_{i}\in S$, where
    $d(s_{i})=\gcd\left\{t\in\mathbb{N}_{+}\ \ |\ \ P_{ii}^{t}>0\right\},$
    is the periodicity of state $s_{i}$, where $\gcd$ stands for “greatest common divisor”.
The ergodic theorem asserts that every irreducible and aperiodic Markov chain has a unique stationary distribution. Its proof is a beautiful application of spectral graph theory. Exercise 3.6 asks the reader for a guided proof of the ergodic theorem.
AB11
Figure 3.5: In this very simple Markov Chain, the state alternates every time step, so there is no unique stationary distribution. This Markov Chain has periodicity of two.
120.50.50.50.5
Figure 3.6: Adding a self-loop to every node makes the periodicity to be one, and this chain is also irreducible. The stationary distribution exists and is unique, it is equal to $\boldsymbol{\pi}=(\frac{1}{2},\frac{1}{2})$.
Reversible Markov Chains
A interesting special case of Markov chains is reversible Markov Chain.
Definition 3.2.
A Markov Chain $P$ is said to be reversible with respect to distribution $\pi$ if and only if the following hold for every two states $i,j$:
${\boldsymbol{\pi}}_{i}{P}_{ij}={\boldsymbol{\pi}}_{j}{P}_{ji}.$
A Markov chain is said to be reversible if there exists some distribution with respect to which the Markov chain that is reversible in the above definition.
In exercise 3.6 we prove that a distribution that is reversible for a Markov chain is also stationary for this Markov chain.
3.2.2 Markov Reward Processes
Markov Reward Processes (MRPs) add the concept of reward to Markov Chains. At every state, the agent receives a reward, which we assume for simplicity to be in the range $[0,1]$. In terms of algorithmic consequences, no generality is lost by this assumption, and this is left as an exercise.
The cumulative reward is a sum of discounted rewards over time, so that immediate rewards are worth more than those received in the future. This is consistent with the economics literature, where a fixed amount of money is worth more today than in the future, due to positive interest rates.
Definition 3.3.
A Markov Reward Process is a tuple $(S,P,R,\gamma)$, where:
  • $S$ is a set of states that can be finite or infinite.
  • $P\in{\mathbb{R}}^{S\times S}$ is the transition matrix.
  • $R\in[0,1]^{S\times S}$ is the reward function, where $R(s,s^{\prime})$ is the reward for moving from state $s$ to $s^{\prime}$.
  • $\gamma\in[0,1]$ is the discount factor.
We henceforth consider instantiations of Markov Reward Processes, and use the following notation:
  1. 1.
    We denote by $S_{t}\in S$ the random state of the process at time $t$.
  2. 2.
    $R_{t}$ = the random reward received at time $t$, starting from state $S_{t}$ and moving to $S_{t+1}$
  3. 3.
    $\hat{R}\in{\mathbb{R}}^{S}$ = the vector that for each state, denotes the expected reward in the next iteration as a function of the current state,
    $\hat{R}(s)=\mathop{\mbox{\bf E}}[R_{t}|S_{t}=s].$
  4. 4.
    $G_{t}=$ the random variable denoting the sum of the discounted rewards starting from time $t$ onwards.
    $G_{t}=\sum_{i=0}^{\infty}R_{t+i}\gamma^{i}.$
When the discount factor that is strictly less than one, $G_{t}$ is well defined. In the case $\gamma=1$, an alternative definition of $G$ is presented below, assuming that a stationary distribution for the underlying Markov chain exists, denoted by $\boldsymbol{\pi}$.
$G:=\lim_{t\to\infty}\frac{1}{t}\sum_{i=1}^{t}R_{i}=\sum_{s\in S}\boldsymbol{\pi}(s)\hat{R}(s)$
One of the most important definitions in reinforcement learning is the value function:
Definition 3.4.
The value function maps states to their expected infinite-horizon reward,
$\mathbf{v}(s)=\mathbb{E}[G_{t}|S_{t}=s].$
The value function tells how valuable a state as measure by the sum of its future, discounted rewards. One way to compute the value function is by using the Bellman equation for the Markov Reward Process, which is the following recursive equation:
$\mathbf{v}(s)=\hat{R}(s)+\gamma\sum_{s^{\prime}}P_{s,s^{\prime}}\mathbf{v}(s^{\prime}).$
To see that this equation holds we can write,
$\mathbf{v}(s)$ $=\mathbb{E}[G_{t}|S_{t}=s]$
$=\mathbb{E}[\sum_{i=0}^{\infty}\gamma^{i}R_{t+i}|S_{t}=s]$
$=\mathbb{E}[R_{t}+\gamma\sum_{i=0}^{\infty}\gamma^{i}R_{t+1+i}|S_{t}=s]$
$=\mathbb{E}[R_{t}+\gamma G_{t+1}|S_{t}=s]$
$=\mathbb{E}[R_{t}+\gamma\mathbf{v}(S_{t+1})|S_{t}=s]$
$=\hat{R}(s)+\gamma\sum_{s^{\prime}}P_{s,s^{\prime}}\mathbf{v}(s^{\prime}).$
In matrix form we can write the Bellman equation as $\mathbf{v}=\mathbf{\hat{R}}+\gamma P\mathbf{v}$, which simplifies to $\mathbf{v}=(I-\gamma P)^{-1}\mathbf{\hat{R}}$. Given $P$ and $\mathbf{\hat{R}}$, this is a linear system of equations, and thus the value function can be solved using Gaussian elimination in cubic time in the number of states.
3.2.3 Markov Decision Processes
The final layer of abstraction is to allow actions, or controls in the language of control theory. This gives the full Markov Decision Process model as defined below.
Definition 3.5.
A Markov Decision Process is a tuple $(S,P,R,A,\gamma)$ such that:
  1. 1.
    $S$ is a set of states.
  2. 2.
    $A$ is a set of possible actions.
  3. 3.
    $P$ is a transition matrix:
    $P_{ss^{\prime}}^{a}=\Pr[S_{t+1}=s|S_{t}=s,A_{t}=a].$
  4. 4.
    $R$ is a reward function that gives each transition and and action a reward $R_{s,s^{\prime},a}$ which is a reward for moving from $s$ to $s^{\prime}$ using action $a$. We henceforth assume w.l.o.g. that $R\in[0,1]^{S\times A}$, where we define
    $R_{sa}=\sum_{s^{\prime}\in S}P_{s,s^{\prime},a}R_{s,s^{\prime},a}.$
  5. 5.
    $\gamma\in[0,1)$ is the discount factor.
The central solution concept of study in the theory of reinforcement learning is called a policy. A policy is a mapping from state to distribution over actions, i.e.
$\pi:S\to\Delta(A)\ ,\ \pi(a|s)=\Pr[A_{t}=a|S_{t}=s]$
The goal of reinforcement learning is to determine a sequence of actions to maximize the long term reward. The Markovian structure of the problem suggests that this optimal value is realized by a stationary policy, i.e. there exists a fixed map from states to actions that maximizes the expected rewards.
Thus, the goal of reinforcement learning is to find a policy which maximizes the expected reward starting from a certain starting state, i.e.
$\max_{\pi\in S\to\Delta(A)}\mathop{\mbox{\bf E}}\left\{\sum_{t=1}^{\infty}\gamma^{t-1}R_{t}|a_{t}\sim\pi(s_{t})\ ,\ S_{1}=s_{1}\right\},$
where $R_{t}$ is a random variable that denotes the reward at time $t$ that depends on the randomness in the policy as well as the transitions.
Similar to the value function of a Markov Reward Process, we can define the value function for a Markov Decision Process as follows.
The value function for a state $s$ given a policy $\pi$ is defined to be the expected reward from this state onwards where the actions are chosen according to $\pi$ at each state, i.e.
$\mathbf{v}_{\pi}(s)=\mathop{\mbox{\bf E}}_{\pi,P}\left[\sum_{t=1}^{\infty}\gamma^{t-1}R_{t}|S_{1}=s\right].$
Similarly, the $Q$ function is defined over state-action pairs given a policy $\pi$ to be the expected cumulative reward gained when starting from some state $s$ and action $a$ and thereafter choosing actions according to $\pi$ at each future state, i.e.
$Q_{\pi}(s,a)=\mathop{\mbox{\bf E}}_{\pi,P}\left[\sum_{t=1}^{\infty}\gamma^{t-1}R_{t}|S_{1}=s,A_{1}=a\right].$
A fundamental theorem of reinforcement learning is the following.
Theorem 3.6.
There exists a deterministic optimal policy $\pi^{*}$, which maps every state to a single action. All optimal policies achieve the same optimal value $\mathbf{v}^{*}(s)$ as well as $Q^{*}(s,a)$, and this holds for every state and action.
The fact that there exist an optimal policy that is optimal regardless of the starting state in a MDP is a consequence of the Markovian nature of the transition dynamics.
3.3 The Bellman Equation
A fundamental property of Markov Decision Processes is that they satisfy the Bellman equation for dynamic programming. The importance of this equation is paramount in algorithm design: we describe two efficient methods for reinforcement learning based on this equation later in this chapter. In later chapters we also make use of the Bellman equation to derive the stochastic-optimal LQR controller.
The Bellman equation for any policy $\pi$ relates the value function for that policy to itself via one step of application for that policy. Mathematically, this amounts to the following
$\mathbf{v}_{\pi}(s)=\sum_{a\in A}\pi(a|s)\left[R_{sa}+\gamma\sum_{s^{\prime}\in S}P_{ss^{\prime}}^{a}\mathbf{v}_{\pi}(s^{\prime})\right].$
It is sometimes convenient to use matrix-vector notation, where we can write:
$\mathbf{v}_{\pi}=\mathbf{R^{\pi}}+\gamma P^{\pi}\mathbf{v}_{\pi},$
where we use the notation
  1. 1.
    $R^{\pi}$ is the induced expected reward vector when playing policy $\pi$, i.e. ${R^{\pi}}(s)=\mathop{\mbox{\bf E}}_{a\sim\pi(s)}[R_{sa}]$.
  2. 2.
    $P^{\pi}$ is the induced transition matrix when playing policy $\pi$, i.e.
    $P^{\pi}_{ss^{\prime}}=\sum_{a}P^{a}_{ss^{\prime}}\cdot\pi(a|s).$
The Bellman equation gains prominance when applied to the optimal policy of a Markov Decision Process. It is then called the Bellman optimality equation, and states that for every state and action, we have
$\mathbf{v}^{*}(s)=\max_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}$
Equivalently, this equation can be written in terms of the $Q$ function as
$\mathbf{v}^{*}(s)=\max_{a}\left\{Q^{*}(s,a)\right\},$
where $Q^{*}$ is the $Q$ function of the optimal policy, i.e.
$Q^{*}(s,a)=R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\max_{a^{\prime}}\left\{Q^{*}(s^{\prime},a^{\prime})\right\}.$
In subsequent chapters, we will also consider cases where the return of interest is either the reward accumulated in a finite number of time steps, i.e. $G=\sum_{t=1}^{T}R_{t}$ in the finite-time setting, or the infinite-time average reward case, i.e. $G=\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}R_{t}$.
In the finite-time setting, both optimal policies and the optimal value function are non-stationary – they depend on the time step in addition to the current state. In this setting, the corresponding Bellman optimality equation can be written as
$\mathbf{v}_{t}^{*}(s)=\max_{a}\left\{R_{sa}+\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}_{t+1}(s^{\prime})\right\}$
starting with $\mathbf{v}_{T}(s)=0$ for all states.
In the average-reward case, the Bellman optimality equation offers the following characterization: $\mathbf{v}^{*}:S\to{\mathbb{R}}$ is the optimal value function if there exists some $\lambda\in{\mathbb{R}}$ such that
$\lambda+\mathbf{v}^{*}(s)=\max_{a}\left\{R_{sa}+\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}.$
Here the scalar $\lambda$ corresponds to the optimal average long-term reward, and $\mathbf{v}^{*}(s)$ reflects the transient advantage of being at state $s$.
3.3.1 Reinforcement Learning as Linear Programming
One of the immediate conclusions from the Bellman optimality equation is an efficient algorithm for reinforcement learning in Markov Decision Processes via linear programming.
Consider the problem of computing the value function corresponding to an optimal policy. The optimal value function gives the optimal policy, and deriving this implication explicitly is left as an exercise.
According to the Bellman optimality equation, the optimal value function satisfies:
$\mathbf{v}^{*}(s)=\max_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\},$
or in vector form,
$\mathbf{v}^{*}=\max_{\mathbf{a}\in A^{|S|}}\left\{\mathbf{R^{\mathbf{a}}}+\gamma P^{\mathbf{a}}\mathbf{v}^{*}\right\},$
where the maximization over actions takes place independently for each state. Equivalently, we can write the following mathematical program:
$\min_{\mathbf{v}^{*}}\mathbf{1}^{\top}\mathbf{v}^{*}$
subject to $\mathbf{v}^{*}(s)\geq R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})$ $\forall(s,a)\in S\times A.$
Notice that this is a linear program! The number of variables is $|S|$, and the number of constraints is $|S|\times|A|$, which can be solved in polynomial time in the number of states and actions. The feasibility of this linear program is not immediate, but will be established in the next section.
3.4 The value iteration algorithm
In the previous section we established the Bellman equation and its consequence to a polynomial time solution for reinforcement learning in Markov Decision Processes. In this section we study a fundamental algorithm for solving Markov Decision Processes that is a more direct application of dynamic programming and the Bellman equation.
The value iteration algorithm start from an arbitrary solution and improves it until a certain measure of optimality is reached. A formal description is given in Algorithm 2. The performance guarantee of Value Iteration, as well as its analysis, is particularly elegant and do not require expertise in mathematical optimization theory.
Algorithm 2 Value Iteration
  Input: MDP
  Set $\mathbf{v}_{0}(s)=0$ to all states
  for $t=1$ to T do
     Update for all states,
$\mathbf{v}_{t+1}(s)=\max_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P_{ss^{\prime}}^{a}\mathbf{v}_{t}(s^{\prime})\right\}$
  end for
  return  $\pi_{T}$ which is derived from the last value function as
$\pi_{T}(s)=\operatorname*{\arg\max}_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P_{ss^{\prime}}^{a}\mathbf{v}_{T}(s^{\prime})\right\}$
Beyond being simple and providing a finite-time convergence rate, the analysis also establishes feasibility of the optimal solution, as given in the following theorem.
Theorem 3.7.
Define the residual vector $\mathbf{r}_{t}(s)=\mathbf{v}^{*}(s)-\mathbf{v}_{t}(s)$. Then the optimal value vector $\mathbf{v}^{*}$ is unique and value iteration converges to the optimal value function and policy at a geometric rate of
$\|\mathbf{r}_{t+1}\|_{\infty}\leq\gamma^{t}\|\mathbf{r}_{1}\|_{\infty}.$
Proof.
For every state we have by choice of the algorithm that
$\mathbf{v}_{t+1}(s)=\max_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}_{t}(s^{\prime})\right\}\geq R_{sa^{*}}+\gamma\sum_{s^{\prime}}P^{a^{*}}_{ss^{\prime}}\mathbf{v}_{t}(s^{\prime}),$
where $a^{*}={\boldsymbol{\pi}}^{*}(s)$ is the action of an optimal policy at state $s$. On the other hand, from Bellman optimality equation, we have that
$\mathbf{v}^{*}(s)=R_{sa^{*}}+\gamma\sum_{s^{\prime}}P^{a^{*}}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime}).$
Subtracting the two equations we get for all states $s$ that
$\mathbf{r}_{t+1}(s)$ $=\mathbf{v}^{*}(s)-\mathbf{v}_{t+1}(s)$
$\leq\gamma\sum_{s^{\prime}}P^{a^{*}}_{ss^{\prime}}(\mathbf{v}^{*}(s^{\prime})-\mathbf{v}_{t}(s^{\prime}))$
$=\gamma\sum_{s^{\prime}}P^{a^{*}}_{ss^{\prime}}\mathbf{r}_{t}(s^{\prime}).$
Or in vector form, $\mathbf{r}_{t+1}\leq\gamma P^{\mathbf{a}^{*}}\mathbf{r}_{t}$. Furthermore, $\mathbf{r}_{t}\geq 0$ is entry-wise non-negative for all $t$. (See Exercise 3.6.) Therefore, it holds that
$\|\mathbf{r}_{t+1}\|_{\infty}\leq\gamma\|P^{\mathbf{a}^{*}}\mathbf{r}_{t}\|_{\infty}\leq\gamma\|\mathbf{r}_{t}\|_{\infty},$
where we use the fact that each row of $P^{\mathbf{a}^{*}}$ sums to one. Applying this repeatedly, we have
$\|\mathbf{r}_{t+1}\|_{\infty}\leq\gamma^{t}\|\mathbf{r}_{1}\|_{\infty}.$
3.5 Bibliographic Remarks
The section on reinforcement learning in the text [RN02] introduces basic concepts, along with commentaries on related topics, such as the origins of the scalar reward model from utility theory. The classical text on reinforcement learning, authored by pioneers of the subject, is [SB18]. In addition, this text extensively discusses the early ventures into the field from researchers working in artificial intelligence, psychology, and neuroscience.
Another great reference is the lecture notes of [Sil15]. See [AJKS19] for an excellent introduction to the theoretical aspects of reinforcement learning, such as sample complexity bounds and reductions to supervised learning. The reader may also look to [Ber19] for an extensive survey of computational approaches to learning in unknown Markov decision processes.
The first author had the benefit of a discussion with [Aro20] on the proof of convergence of value iteration presented in this chapter.
Dynamic programming approaches, in addition to being widely used in algorithm design [CLRS22], are at the heart of Markov decision processes. This view is emphasized in [Ber07], which discusses the sufficiency and necessity of the Bellman optimality conditions at length. It is also a good resource for an involved reader who might be interested in the completeness of the Bellman characterization and the existential concerns for the average-time setting, which we did not expand upon.
Although tabular MDPs fail to compactly represent stochastic linear dynamical systems, function-approximation schemes [JYWJ20, DKL+21, JKA+17, KKL+20] successfully capture the stochastic LQR setting.
3.6 Exercises
3.1. In this exercise we give a guided proof of the ergodic theorem. Consider a given irreducible and a-periodic Markov chain with transition matrix $P$.
A. Show that $P$ has a left eigenvalue equal to 1.
B. Show that all eigenvalues of $P$ are at most one in magnitude.
The irreducibility and aperiodicity of the Markov chain imply that the eigenvalue equal to one is unique. We are going to assume this fact, and denote by $\mathbf{v}_{1}$ the corresponding eigenvector.
C. Consider an abirtrary distribution $\mathbf{p}_{0}$, and the distribution after $t$ iterations of the Markov chain $\mathbf{p}_{t}=P^{t}\mathbf{p}_{0}$. Prove that $\lim_{t\to\infty}\mathbf{p}_{t}=\mathbf{v}_{1}.$
3.2. Prove the geometric convergence of the value iteration algorithm for rewards that can be arbitrary real numbers, as opposed to in the range $[0,1]$.
3.3. Consider the derivation of an optimal policy for a Markov Decision Process given its value function as described in algorithm 2. Prove that the resulting policy is indeed optimal.
3.4. Let Markov chain $P$ and distribution ${\boldsymbol{\pi}}$ be reversible for it, i.e.
$\forall i,j\ ,\ {\boldsymbol{\pi}}_{i}P_{ij}={\boldsymbol{\pi}}_{j}P_{ji}.$
Prove that ${\boldsymbol{\pi}}$ is stationary for $P$.
3.5. Prove that the successive iterates $\mathbf{v}_{t}$ produced by the Value Iteration algorithm are monotonically increasing, i.e. $\mathbf{v}_{t+1}\geq\mathbf{v}_{t}$, when initialized with $\mathbf{v}_{0}=\mathbf{0}$. Hence, conclude that for any $t$, $\mathbf{v}^{*}\geq\mathbf{v}_{t}$.
3.6. Let $P$ be an aperiodic and irreducible Markov chain. Prove that $P$ is reversible if and only if for any sequence of states $s_{1},\dots s_{n}$, it holds that
$P_{s_{1}s_{2}}P_{s_{2}s_{3}}\dots P_{s_{n-1}s_{n}}P_{s_{n}s_{1}}=P_{s_{1}s_{n}}P_{s_{n}s_{n-1}}\dots P_{s_{2}s_{1}}.$
Chapter 4 Linear Dynamical Systems
In chapter 2, we observed that even simple questions about general dynamical systems are intractable. This motivates the restriction to a special class of dynamical systems for which we can design efficient algorithms with provable guarantees.
Our treatment in this chapter is restricted to linear dynamical systems (LDS): dynamical systems that evolves linearly according to the equation
$\mathbf{x}_{t+1}={A}_{t}\mathbf{x}_{t}+{B}_{t}{\mathbf{u}}_{t}+{\mathbf{w}}_{t}.$
If the matrices ${A}_{t}={A}$ and ${B}_{t}={B}$ are fixed, we say that the dynamical system is linear time-invariant (LTI). We have seen examples of LTI systems in chapter 2, namely the double integrator and the aircraft dynamics.
Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories.
A more general class of linear dynamical systems is when the matrices ${A}_{t}$ and ${B}_{t}$ change over time. These are called linear time-varying (LTV) systems. The class of LTV systems can be used to express much more general dynamics via linearization, as we show next.
4.1 General Dynamics as LTV Systems
Consider any dynamics that can be expressed by a smooth function $f$. Such a system can be approximated by a time-varying linear system to an arbitrary accuracy determined by the time step interval between samples.
To see this, first consider the one dimensional first-order Taylor expansion of a given dynamics function $f$ around some anchor point $(\bar{x}_{t},\bar{u}_{t})$:
$x_{t+1}$ $=f(x_{t},u_{t})+w_{t}$
$\approx f(\bar{x}_{t},\bar{u}_{t})+\underbrace{\frac{\partial}{\partial x}f(x,u)\bigg{|}_{\bar{x}_{t},\bar{u}_{t}}}_{\triangleq A_{t}}(x_{t}-\bar{x}_{t})+\underbrace{\frac{\partial}{\partial u}f(x,u)\bigg{|}_{\bar{x}_{t},\bar{u}_{t}}}_{\triangleq B_{t}}(u_{t}-\bar{u}_{t})+w_{t}.$
With an appropriate redefinition of $w_{t}$, this simplifies to
$x_{t+1}=A_{t}x_{t}+B_{t}u_{t}+w_{t},$
where An analogous approximation holds for the higher-dimensional case. The Taylor expansion is
$f(\mathbf{x},\mathbf{u})\approx f(\bar{\mathbf{x}},\bar{\mathbf{u}})+J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{x}-\bar{\mathbf{x}})+J_{\mathbf{u}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{u}-\bar{\mathbf{u}}),$
where $J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is the Jacobian of $f$ with respect to the variables $\bar{\mathbf{x}}$. Writing the matrices ${A}_{t},{B}_{t}$ for the Jacobians, we have
$\mathbf{x}_{t+1}={A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
This derivation explains why in practice, the class of time varying linear dynamical systems is very expressive, at least for smooth dynamical systems. This is particularly true with hidden states that can capture complicated dynamics while imposing enough structure for efficient algorithms.
As we see next, stabilizability and controllability can be checked efficiently for linear systems, in contrast to general dynamics.
4.2 Stabilizability of Linear Systems
Verifying stability of a system, and more so stabilizability, was shown to be intractable in our discussion of chapter 2. For linear dynamical systems, however, the picture is quite different: stability and stabilizability can be efficiently verified! Furthermore, the theoretical analysis of this fact is interesting and elegant.
For simplicity, in the rest of this section we consider time-invariant linear dynamical systems of the form
$\mathbf{x}_{t+1}={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t},$
although much of our discussion can be generalized to LTV systems. Exercise 4.6 asks the reader to prove this.
Recall that a system is stabilizable if from every starting state, there is a sequence of controls that drives the system to zero. Determining whether a linear time-invariant system is stabilizable is determined by a particularly simple condition as follows. Henceforth we denote by $\rho({A})$ the spectral radius of a matrix ${A}$, that is,
$\rho({A})=\max\{|\lambda_{1}|,...,|\lambda_{d}|\},$
the maximum modulus from all (possibly complex) eigenvalues. The stabilizability of a linear dynamical system can be characterized as follows.
Theorem 4.1.
A time-invariant LDS is stabilizable if and only if there exists a matrix $K$ such that
$\rho({A}+{B}{K})<1.$
We henceforth prove sufficiency of this condition, and leave necessity as exercise 4.6. Both directions rely on the characterization of the spectral norm as follows:
Fact 4.2.
$\rho({A})<1$ if and only if $\lim_{t\mapsto\infty}{A}^{t}=0$. Furthermore, if $\rho({A})>1$, then for any norm, $\lim_{t\mapsto\infty}\|{A}^{t}\|=\infty$.
The proof of this fact is left as exercise 4.6. Based on this fact, we can now prove the sufficiency part of Theorem 4.1.
Theorem 4.1, sufficiency.
Suppose there exists a ${K}$ such that $\rho({A}+{B}{K})<1$. Then choose all controls to be $\mathbf{u}_{t}={K}\mathbf{x}_{t}$. The evolution of the system without perturbation is given by
$\mathbf{x}_{t}=({A}+{B}{K})\mathbf{x}_{t-1}=...=({A}+{B}{K})^{t}\mathbf{x}_{0}.$
Thus, using the previous fact concerning the spectral radius, we have
$\lim_{t\to\infty}\|\mathbf{x}_{t}\|=\lim_{t\to\infty}\|({A}+{B}{K})^{t}x_{0}\|=\|0\mathbf{x}_{0}\|=0.$
4.2.1 Efficient Verification of Stabilizability
The spectral radius of a matrix can be computed to an arbitrary accuracy in polynomial time, and can be approximated by the power method. This implies that stability of a given system can be efficiently verified.
To efficiently check if a given system is stabilizable, we make note of the following result due to Lyapunov.
Theorem 4.3.
$\rho({A})<1$ holds if and only if there exists a matrix ${P}\in{\mathbb{R}}^{d_{x}\times d_{x}}$ such that ${P}\succeq{I}$ and ${P}\succ{A}^{\top}{P}{A}$.
Proof.
Say there exists a matrix $P$ such that $P\succeq I$ and $P\succ A^{\top}PA$, then to show that $\rho(A)<1$, it suffices to prove that $\lim_{t\to\infty}A^{t}=0$. To observe the latter, first note that there exists a positive constant $c<1$ such that $cP\succeq A^{\top}PA$, implying that $c^{t}P\succeq(A^{t})^{\top}PA^{t}\succeq(A^{t})^{\top}A^{t}$ via repeated applications of the previous inequality. Therefore, we have that $\lim_{t\to\infty}(A^{t})^{\top}A^{t}=0$, and consequently that $\lim_{t\to\infty}A^{t}=0$ since $\|A^{\top}\|\|A\|=\|A^{\top}A\|$.
For the other direction, given some matrix $A$ with $\rho(A)<1$, consider $P=\sum_{t=0}^{\infty}(A^{t})^{\top}A^{t}$. Since $P=I+A^{\top}PA$, $P$ satisfies the required conclusions. The fact that $P$ as defined exists, i.e. that it is finite, as long as $\rho(A)<1$ can be verified via the Jordan normal form of $A$, which in particular asserts that $\|A^{t}\|\leq O(t^{d_{x}}\rho(A)^{t})$. ∎
Using this observation, one can write a semi-definite mathematical program that checks if a given system is stabilizable, and if so produces a matrix $K$ that guarantees a spectral radius smaller than one.
Theorem 4.4.
There exists a matrix ${K}$ such that $\rho({A}+{B}{K})<1$ if and only if there exist matrices ${P}\in\mathcal{S}^{d_{x}\times d_{x}}$ and ${L}\in{\mathbb{R}}^{d_{u}\times d_{x}}$ such that
$\begin{bmatrix}{P}&({A}{P}+{B}{L})^{\top}\\ {A}{P}+{B}{L}&{P}\end{bmatrix}\succ 0,\qquad{P}\succeq{I}.$ (4.1)
Furthermore, given such matrices ${P}$ and ${L}$ satisfying Equation 4.1, the matrix ${K}^{\prime}={L}{P}^{-1}$ satisfies $\rho({A}+{B}{K}^{\prime})<1$.
Proof.
We begin by observing the following result, using the Schur complement, regarding the semi-definiteness of block matrices.
Lemma 4.5.
Define a block matrix ${M}$ as
${M}=\begin{bmatrix}{A}&{B}\\ {B}^{\top}&{C}\end{bmatrix}.$ (4.2)
Then, ${M}\succ{0}$ if and only if ${A}\succ{0}$ and ${C}-{B}^{\top}{A}^{-1}{B}\succ{0}$.
From Theorem 4.3, we know that stabilizability of a linear system is equivalent to the existence of a matrices ${K}$ and ${P}$ such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)^{\top}P(A+BK)}$. The above lemma allows us to reformulate this cubic constraint to a quadratic constraint. The stabilizability of a system is equivalent to exists of matrices ${K}$ and ${P}$ such that
$\begin{bmatrix}{P}&{((A+BK)P)^{\top}}\\ {(A+BK)P}&{P}\end{bmatrix}\succ{0},\qquad{P}\succeq{I}.$
Lastly, we reparametrize this constraint as a linear semi-definite constraint as demonstrated below, by choosing ${L}={KP}$.
$\begin{bmatrix}{P}&{(AP+BL)^{\top}}\\ {AP+BL}&{P}\end{bmatrix}\succ{0},\qquad{P}\succeq{I}.$
Notice that this is indeed a reformulation of the mathematical feasibility program, and not a relaxation, since the reparamertization ${K}={LP^{-1}}$ is invertible. ∎
4.3 Controllability of LDS
Recall that a system is controllable if and only if in the absence of noise and for every target state, there is a sequence of controls that drives the system state to that target. Linear dynamical systems exhibit a particularly elegant characterization of controllability. First, we need to define an object of central importance in linear dynamical system, the Kalman matrix.
Definition 4.6.
The Kalman controllability matrix of a linear dynamical system is given by
${K}_{r}({A},{B})=[{B},{AB},{A^{2}B},\ldots,{A^{r-1}B}]\in{\mathbb{R}}^{d_{x}\times rd_{u}}.$
The controllability of a linear dynamical system can now be characterized using this matrix. Recall the rank of a matrix is the maximal number of linearly independent columns (or rows).
Theorem 4.7.
A time-invariant linear dynamical system is controllable if and only if
$\mathop{\mbox{\rm rank}}{K}_{d_{x}}({A},{B})=d_{x}.$
Proof.
Consider the evolution of the noiseless linear dynamical system starting from initial state $\mathbf{x}_{0}$. The state at time $T$ is given by the expression
$\mathbf{x}_{T}$ $={A}\mathbf{x}_{T-1}+{B}\mathbf{u}_{T-1}$
$={A}^{2}\mathbf{x}_{T-2}+{A}{B}\mathbf{u}_{T-2}+{B}\mathbf{u}_{T-1}$
$=\ldots=\sum_{t=1}^{T}{A}^{t-1}{B}\mathbf{u}_{T-t}+{A}^{T}\mathbf{x}_{0}.$
Thus,
$\mathbf{x}_{T}-{A}^{T}\mathbf{x}_{0}=[{B},{AB,A^{2}B,\ldots,A^{T-1}B}]\begin{bmatrix}\mathbf{u}_{T-1}\\ \vdots\\ \mathbf{u}_{0}\end{bmatrix}.$
By the fundamental theory of linear algebra, this equation has a solution if and only if ${K}_{T}({A,B})$ has full rank. That is, for every initial state $\mathbf{x}_{0}$ and final state $\mathbf{x}_{T}$, there exists a set of control vectors $\{\mathbf{u}_{t}\}$ that satisfy the equation if and only if ${K}_{T}({A,B})$ has full rank. The rank of the Kalman matrix is upper bounded by the number of rows of ${K}_{T}$, which is $d_{x}$. Thus, the system is controllable if and only if the rank of ${K}_{T}$ is $d_{x}$ for all $T\geq d_{x}$.
Denote by $r_{t}=\mathop{\mbox{\rm rank}}({K}_{t})$ the rank of the Kalman matrix after $t$ iterations. Since the rank can never decrease with $t$, it remains to show that the system is controllable if and only if $r_{t}=r_{d_{x}}$ for all $t\geq d_{x}$.
We claim that once $r_{t}=r_{t+1}$, the rank never increases again by increasing $t$. Assuming $r_{t}=r_{t+1}$, we have
$\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t}B}]=\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t-1}B}].$
Since linear transformations retain containment, we can multiply by ${A}$ and get
$\mathop{\mbox{\rm span}}[{AB,\ldots,A^{t+1}B}]\subseteq\mathop{\mbox{\rm span}}[{AB,\ldots,A^{t}B}].$
Adding ${B}$ to both matrices doesn’t affect the subset relationship.
$\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t+1}B}]\subseteq\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t}B}]$
Therefore, $r_{t+2}\leq r_{t+1}$. Since the rank cannot decrease through the addition of columns, we have $r_{t+2}=r_{t+1}$. Since the rank can grow at most $d_{x}$ times, the equation $r_{d_{x}}=r_{T}$ holds for all $T>d_{x}$ if and only if the system is controllable, and this concludes the theorem. ∎
4.4 Quantitative Definitions
In the rest of this text our goal is to give efficient algorithms with finite time performance bounds for online control. These finite time bounds naturally depend on how difficult it is for a given system to stabilize. Thereforem we require a quantitative definition for stabilizability, one that holds for both time-invariant and time-varying linear dynamical systems.
4.4.1 Stabilizability of Time-Invariant LDS
Consider a linear time-invariant system given by the transformations $({A,B})$. We have already remarked that this system is called stabilizable, according to the classical, noise-free definition, if and only if there exists $K$ such that the condition $\rho({A+BK})<1$ holds.
A stronger quantitative property can be obtained as follows, and it is in fact implied from BIBO stabilizability as defined in Definition 2.4. The proof of this fact is left as exercise 4.6.
Lemma 4.8.
Consider a linear controller ${K}$ for the LTI system $({A,B})$. If there exists $\kappa>0$ and $\delta\in[0,1)$ such that the matrix ${A+BK}$ admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
$\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t},\qquad\frac{\kappa}{\delta}\leq\gamma$
Proof.
Let ${\tilde{K}=A+BK}$. By unrolling the dynamics equation we have that
$\mathbf{x}_{t+1}$ $={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t}=\tilde{K}\mathbf{x}_{t}+\mathbf{w}_{t},$
$=\sum_{i=0}^{t-1}\tilde{{K}}^{i}\mathbf{w}_{t-i}+\tilde{{K}}^{t+1}\mathbf{x}_{0}.$
Thus, as long as $\|\mathbf{w}_{t}\|,\|\mathbf{x}_{0}\|\leq 1$, we have
$\|\mathbf{x}_{t+1}\|$ $\leq\sum_{i=0}^{t}\|\tilde{K}^{i}\|\|\mathbf{w}_{t-i}\|+\|\tilde{K}^{t+1}\|\|\mathbf{x}_{0}\|,$
$\leq\kappa\sum_{i=0}^{\infty}(1-\delta)^{i}\leq\frac{\kappa}{\delta}\leq\gamma.$
A linear time-invariant system is thus $\gamma$-stabilizable if we can find a matrix $K$ with the properties above.
4.4.2 Stabilizability of time-variant linear systems
The $\gamma$-stabilizability condition, as opposed to classical stabilizability, is particularly important to generalize to time-varying systems, since the spectral radius of the product is not bounded by the product of spectral radii, see exercise 4.6 at the end of this chapter.
We now define a sufficient condition for $\gamma$ stabilizability. Consider a time-varying linear dynamical system given by the transformations $\{{A}_{t},{B}_{t}\}$.
Lemma 4.9.
Consider a sequence of linear controllers ${K}_{1:T}$ for the time-varying linear dynamical system $\{{A}_{1:T},{B}_{1:T}\}$. If the matrices ${\tilde{K}_{t}=A_{t}+B_{t}K_{t}}$ satisfy the following inequality, then this linear controller sequence is $\gamma$-stabilizing.
$\forall t_{1},t_{2}\ .\ \left\|\prod_{t=t_{1}}^{t_{2}}\tilde{K}_{t}\right\|\leq\kappa(1-\delta)^{t_{1}-t_{2}},\ \frac{\kappa}{\delta}\leq\gamma.$
Proof.
Notice that under the condition of the lemma, by unrolling the dynamics equation we have that
$\|\mathbf{x}_{t+1}\|$ $=\|{A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}\|=\left\|\tilde{{K}}_{t}\mathbf{x}_{t}+\mathbf{w}_{t}\right\|,$
$=\left\|\sum_{i=0}^{t-1}(\prod_{j=1}^{i}\tilde{{K}}_{t-j})\mathbf{w}_{t-i}+(\prod_{j=1}^{t}\tilde{{K}}_{t-j})\mathbf{x}_{0}\right\|,$
$\leq\kappa\sum_{i=0}^{\infty}e^{-\delta i}\leq\frac{\kappa}{\delta}\leq\gamma.$
4.4.3 Controllability of time-invariant systems
Consider the following one-dimensional dynamical system for some small $\varepsilon>0$:
$x_{t+1}=x_{t}+\varepsilon u_{t}.$
Clearly, such a system is nominally controllable, as indeed is any one-dimensional system with non-zero $B$. However, to reach a state $x=1$ from an initial $x_{0}=0$ in a constant number of time steps requires a large control input $u_{t}$ of size at least $1/\varepsilon$. Thus, such a system may be said to be barely controllable. This difficulty of controllability is captured in the definition below, which requires not only that $K_{k}(A,B)$ be full row-rank, but also that such a statement is robust to small changes in $A,B$.
Definition 4.10.
A time-invariant linear dynamical system $(A,B)$ is $(k,\kappa)$-strongly controllable if and only if
$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\text{ and }\|({K}_{k}(A,B))^{+}\|\leq\kappa.$
4.5 Bibliographic Remarks
The idea of reducing a nonlinear dynamical system to a time varying linear dynamical system had been independently discovered in disparate fields. In the context of control, using the linearization to characterize local stability properties of non-linear systems was pioneered by Lyapunov [Lya92], and is called the Lyapunov direct method, see e.g. [SL91].
The asymptotic notions of stability, stabilizability and controllability discussed in this chapter constitute a core part of theory of linear control, and are found in most texts on the subject (e.g. [Ste94, Hes18]).
The efficient verification of stabilizability via a semi-definite program and the recovery of a stable controller form the solution of the latter appears in [Ahm16]. On the other hand, it is known that even imposition of norm constraints on a plausibly stable controller makes this problem computationally hard [BT97, BT00, BT99]. Somewhat distinct, but related, problems of stabilizability with even a 1-dimensional linear constraint are open to the date.
The notion of $\gamma$-stabilizability has a precursor in the notion of strong stability introduced in [CHK+18]. An extension of this strong stability notion to time-varying linear dynamics was established in [CKM19]. The definition presented in this chapter has weaker requirements than those in the listed articles, allowing us to work with and guarantee regret later with respect to a larger class of comparator policies.
4.6 Exercises
4.1. Prove that the existence of a matrix $K$ such that $\rho(A+BK)<1$ is a necessary condition for a linear dynamical system to be stabilizable. Hint: this exercise requires the optimal solution to LQR from the next chapter.
4.2. Prove that $\rho(A)<1$ if and only if $\lim_{t\mapsto\infty}A^{t}=0$. Furthermore, if $\rho(A)>1$, then for any norm, $\lim_{t\mapsto\infty}\|A^{t}\|=\infty$. Finally, prove that $\rho(A)=\lim_{t\to\infty}\|A^{t}\|^{1/t}$.
4.3. Prove that the spectral radius of a product of matrices is not necessarily bounded by the product of the spectral radii of the individual matrices. That is, show that in general,
$\rho(\prod_{i=1}^{n}A_{i})\nleq\prod_{i=1}^{n}\rho(A_{i}).$
4.4. Prove that a stochastic linear dynamical system with Gaussian perturbations
$\mathbf{x}_{t+1}=A\mathbf{x}_{t}+\mathbf{w}_{t},\qquad\mathbf{w}_{t}\sim\mathcal{N}(0,I),$
constitutes a reversible Markov chain if and only if $A$ is symmetric, i.e. $A=A^{\top}$.
Hint: One path to this exercise is to explicitly calculate the stationary distribution of a LDS and then to verify reversibility as stated in its definition. Exercise 3.6 provides an alternative without the need to derive the stationary distribution.
4.5. This exercise explores the consequences of $\gamma$-BIBO stability (Definition 2.4) for linear systems. While we have already seen sufficient conditions to ensure BIBO stability, we ask for the necessity of such conditions.
A. For a time-invariant linear system $(A,B)$ equipped with a linear policy $K$, prove that BIBO stability indeed implies that $\rho(A+BK)<1$. In fact, comment on how large $\rho(A+BK)$ can be.
B. For a time-invariant linear system, prove that there exists $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$, $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}$. In particular, this proves that the conditions listed in Lemma 4.8 are necessary, in addition to being sufficient.
C. Construct a time-varying linear systems that is BIBO stable, yet does not have the exponential decay properties required in Lemma 4.9.
4.6. Prove that the bound of one on the operator norm of the matrix $A$ is sufficient but not necessary condition for stabilizability.
Chapter 5 Optimal Control of Linear Dynamical Systems
In the chapters thus far we have studied more basic notions than control, namely solution concepts of dynamical systems such as stabilizability and controllability. Even for these more basic problems, we have seen that they are computationally intractable in general.
This motivated us to look into simpler dynamics, namely linear dynamical systems, for which we can at least answer basic questions about stabilizability and controllability. In this chapter, we consider how we can control linear dynamical systems optimally.
The answer is positive in a strong sense. An archetypal problem in optimal control theory is the control of linear dynamical systems with quadratic cost functions. This setting is called the Linear Quadratic Regulator which we define and study in this chapter. The resulting efficient and practical algorithm is one of the most widely used subroutines in control.
5.1 The Linear-Quadratic Regulator
The linear quadratic regulator (LQR) problem is that of optimal control of a linear dynamical system with known and fixed quadratic costs. Formally,
Definition 5.1.
The Linear Quadratic Regulator (LQR) input is a linear dynamical system that evolves according to
$\mathbf{x}_{t+1}={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t},$
where $\mathbf{w}_{t}\sim\mathcal{N}(0,\sigma^{2})$ is i.i.d. Gaussian, and convex quadratic cost functions are specified by matrices $Q$ and $R$.
The goal is to generate a sequence of controls $\mathbf{u}_{t}$ that minimizes the expected cost, taken over the randomness of the perturbations, given by quadratic loss functions,
$\min_{\mathbf{u}(\mathbf{x})}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R}\mathbf{u}_{t}\right].$
To remind the reader, the notation we use is as follows:
  1. 1.
    $\mathbf{x}_{t}\in\mathbb{R}^{d_{x}}$ is the state.
  2. 2.
    $\mathbf{u}_{t}\in\mathbb{R}^{d_{u}}$ is the control input.
  3. 3.
    $\mathbf{w}_{t}\in\mathbb{R}^{d_{w}}$ is the perturbation sequence drawn i.i.d. from normal distributions, or any other zero-mean distribution as long as the draws are independent of states or controls.
  4. 4.
    $A\in{\mathbb{R}}^{d_{x}\times d_{x}}$, $B\in{\mathbb{R}}^{d_{x}\times d_{u}}$ are the system matrices.
  5. 5.
    $Q\in{\mathbb{R}}^{d_{x}\times d_{x}}$, ${R}\in{\mathbb{R}}^{d_{u}\times d_{u}}$ are positive semi-definite matrices. Thus, the cost functions are convex.
This framework is very useful since many physical problems, such as the centrifugal governor and the ventilator control problem, can be approximated by a linear dynamical system. In addition, the quadratic convex cost functions are natural in many physical applications.
The LQR is a special case of a Markov decision process which we have described in chapter 3, since the transitions are stochastic. However, naive application of value iteration, policy iteration, or linear programming do not apply, since there are infinitely many states and actions.
We proceed to describe a particularly elegant solution to the LQR problem that makes use of its special structure.
5.2 Optimal Solution of the LQR
We know from Chapter 3 that the LQR is a special case of a finite-time Markov Decision Process, and hence admits an optimal policy. However, the optimal policy is in general a mapping from states to actions, both of which are infinite Euclidean spaces in our case.
Nevertheless, it can be shown that the optimal policy for an LQR is a linear policy, i.e. a linear mapping from ${\mathbb{R}}^{d_{x}}$ to ${\mathbb{R}}^{d_{u}}$! This astonishing fact together with a remarkable property of the value function is formally given in the following theorem. We state the theorem for a finite number of iterations, and for a linear time-invariant system. Generalizing this to a time-varying linear system is left as exercise 5.6.
Theorem 5.2.
For the LQR problem with finite $T>0$ iterations, there exist positive semi-definite matrices $S_{t}$, ${K}_{t}$ and scalars $c_{t}$, such that the optimal value function and optimal policy satisfy:
$\mathbf{v}_{t}^{*}(\mathbf{x})=\mathbf{x}^{\top}{S}_{t}\mathbf{x}+c_{t},\quad\mathbf{u}_{t}^{*}(\mathbf{x})={K}_{t}\mathbf{x}.$
Proof.
Recall that the Bellman Optimality equation for a general MDP is written as
$\mathbf{v}_{t}^{*}(s)=\max_{a}\left\{R_{sa}+\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}_{t+1}^{*}(s^{\prime})\right\}.$
We henceforth write the special case applicable to LQR, and prove the theorem by backward induction on $t$.
Base Case: For $t=T$, we have
$\mathbf{v}^{*}_{T}(\mathbf{x})=\min_{\mathbf{u}}\big{\{}\mathbf{x}^{\intercal}Q\mathbf{x}+\mathbf{u}^{\intercal}R\mathbf{u}\big{\}}.$
Since ${R}\succcurlyeq 0$, a minimum occurs at $\mathbf{u}=0$. Hence, ${S}_{T}={Q}\succcurlyeq{0}$, ${K}_{T}={0}$ and $c_{T}=0$.
Inductive Step: Assume for some $t>1$ that $\mathbf{u}_{t}^{*}(\mathbf{x})={K}_{t}\mathbf{x}$ and $\mathbf{v}_{t}^{*}(\mathbf{x})=\mathbf{x}^{\intercal}{S}_{t}\mathbf{x}+c_{t}$. Then we have
$\mathbf{v}^{*}_{t-1}(\mathbf{x})$ $=\min_{\mathbf{u}}\big{\{}\mathbf{x}^{\intercal}{Q}\mathbf{x}+\mathbf{u}^{\intercal}{R}\mathbf{u}+\mathop{\mbox{\bf E}}_{\mathbf{w}}\mathbf{v}^{*}_{t}({A}\mathbf{x}+{B}\mathbf{u}+\mathbf{w})\big{\}}$
$=\min_{\mathbf{u}}\big{\{}\mathbf{x}^{\intercal}{Q}\mathbf{x}+\mathbf{u}^{\intercal}{R}\mathbf{u}+\mathop{\mbox{\bf E}}_{\mathbf{w}}\left[({A}\mathbf{x}+{B}\mathbf{u}+\mathbf{w})^{\intercal}S_{t}({A}\mathbf{x}+{B}\mathbf{u}+\mathbf{w})\right]\big{\}}+c_{t}$
$=\min_{\mathbf{u}}\big{\{}\mathbf{x}^{\intercal}{Q}\mathbf{x}+\mathbf{u}^{\intercal}{R}\mathbf{u}+({A}\mathbf{x}+{B}\mathbf{u})^{\intercal}{S}_{t}({A}\mathbf{x}+{B}\mathbf{u})+\mathop{\mbox{\bf E}}_{\mathbf{w}}[\mathbf{w}^{\top}S_{t}\mathbf{w}]\big{\}}+c_{t},$
where in the last equality we used the fact that $\mathop{\mbox{\bf E}}[\mathbf{w}]=0$, and that the perturbation is independent of the state and control. We henceforth denote the second moment of the noise by $\mathop{\mbox{\bf E}}[\mathbf{w}\mathbf{w}^{\top}]=\sigma^{2}I$.
Since $R$ and $S_{t}$ are positive semi-definite, the expression above is convex as a function of $\mathbf{u}$. At the global minimum the gradient vanishes, and hence
$\nabla_{\mathbf{u}}\Big{[}{\mathbf{x}^{\intercal}Q\mathbf{x}+\mathbf{u}^{\intercal}R\mathbf{u}+(A\mathbf{x}+B\mathbf{u})^{\intercal}S_{t}(A\mathbf{x}+B\mathbf{u})+\sigma^{2}{\bf Tr}(S_{t})}]\Big{]}=0.$
Since ${Q,R,S_{t}}$ are all positive semi-definite and symmetric, this gives us
${2R\mathbf{u}+2B^{\intercal}S_{t}A\mathbf{x}+2B^{\intercal}S_{t}B\mathbf{u}=0\implies(R+B^{\intercal}S_{t}B)\mathbf{u}=-B^{\intercal}S_{t}A\mathbf{x}}.$
Hence, we have that
${\mathbf{u}_{t-1}(\mathbf{x})=K_{t-1}\mathbf{x},\ \ \ K_{t-1}=-(R+B^{\intercal}S_{t}B)^{-1}(B^{\intercal}S_{t}A)}.$
If the matrices above are not full rank, we use the notation $M^{-1}$ for the Moore-Penrose pseudo inverse. Plugging this in the expression for the value function, we get
$\mathbf{v}_{t-1}^{*}(\mathbf{x})$
$={\mathbf{x}^{\intercal}Q\mathbf{x}+\mathbf{x}^{\intercal}K_{t-1}^{\intercal}RK_{t-1}\mathbf{x}+\mathbf{x}^{\intercal}(A-BK_{t-1})^{\intercal}S_{t}(A-BK_{t-1})\mathbf{x}+\sigma^{2}{\bf Tr}(S_{t})}+c_{t}$
$=\mathbf{x}^{\intercal}{S}_{t-1}\mathbf{x}+c_{t-1}$
where ${S_{t-1}=Q+K_{t-1}^{\intercal}RK_{t-1}+(A-BK_{t-1})^{\intercal}S_{t}(A-BK_{t-1})}$ is positive semi-definite, and $c_{t-1}=c_{t}+\sigma^{2}{\bf Tr}(S_{t})$.
5.3 Infinite Horizon LQR
Although infinite horizon problems can never appear in the real world, in practice the horizon length may be unknown. Although there are techniques to iteratively update the horizon length with geometric increasing guesstimates, these are slow and impractical.
An alternative is to solve the LQR for the infinite horizon case, which results in a very elegant solution. Not only that, this turns out to be a practical solution to which the finite-horizon solution converges quickly.
The infinite horizon LQR problem can be written as follows:
$\min_{\mathbf{u}(\mathbf{x})}\lim_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}{R}\mathbf{u}_{t}\right]$
[…]
where similar to the finite horizon case, we assume that the perturbations $\mathbf{w}_{t}$ are i.i.d. zero mean and have bounded variance. We henceforth assume that the perturbations are all zero for simplicity, adding zero-mean bounded variance perturbations is left as exercise 5.6.
Without perturbations, the Bellman optimality equation becomes
$\mathbf{v}^{*}(\mathbf{x})=\min_{\mathbf{u}}\left\{{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}+\mathbf{v}^{*}(A\mathbf{x}+B\mathbf{u})}\right\}.$
Assuming that the value function is quadratic, we have:
$\mathbf{x}^{\top}S\mathbf{x}=\min_{\mathbf{u}}\left\{{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}+(A\mathbf{x}+B\mathbf{u})^{\top}S(A\mathbf{x}+B\mathbf{u})}\right\}.$
Similar to the finite time case, we have the following result.
Lemma 5.3.
The optimal policy is given by a transformation $\mathbf{u}^{*}(\mathbf{x})={K}\mathbf{x}$.
Proof.
At optimality, the gradient with respect to $\mathbf{u}$ of the equation above is zero. This implies that
${2R\mathbf{u}+2B^{\top}SB\mathbf{u}+2B^{\top}SA\mathbf{x}=0},$
which implies
${K=-(R+B^{\top}SB)^{-1}(B^{\top}SA)}.$
Plugging back into the Bellman equation, we obtain
${\mathbf{x}^{\top}S\mathbf{x}=\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{x}^{\top}K^{\top}RK\mathbf{x}+\mathbf{x}^{\top}(A+BK)^{\top}S(A+BK)\mathbf{x}}.$
Since this equality holds for all $\mathbf{x},$, we get
${S=Q+K^{\top}RK+(A+BK)^{\top}S(A+BK)}.$
Simplifying this expression we obtain
$S=Q+A^{T}SA-A^{T}SB\left({R}+B^{T}SB\right)^{-1}B^{T}SA.$
This equation is called the Discrete Time Algebraic Riccati Equation (DARE) and is one of the most important equations in control theory.
5.3.1 Solving the DARE
In the 1-D case, although the right side of the equation is a rational function, we can multiply both sides by ${R+B^{\top}SB}$ to form a quadratic equation in ${S}$. Solving the DARE therefore corresponds to finding a non-negative root of a quadratic polynomial in the one-dimensional case.
In higher dimension, the correct generalization is somewhat more difficult to see, but using Lemma 4.5, the equation can be reformulated as the solution to the following SDP.
$\min_{{X}\in\mathcal{S}^{d_{x}\times d_{x}}}\langle{X,I}\rangle$
$\text{subject to}\begin{bmatrix}{X}&B^{T}SA\\ A^{T}SB&{R}+B^{T}SB\end{bmatrix}\succeq 0,$
and $\text{and }S=Q+A^{T}SA-{X}.$
A popular alternative is to adapt the value iteration algorithm to this setting by iteratively applying the right side of the aforementioned equation as an operator. Concretely, starting from ${S}_{0}\succeq 0$, we choose
$S_{t+1}=Q+A^{T}S_{t}A-A^{T}S_{t}B\left({R}+B^{T}S_{t}B\right)^{-1}B^{T}S_{t}A.$
5.4 Robust Control
The solution to the LQR problem is a hallmark of optimal control theory and one of the most widely used techniques in control. However, the assumption that the disturbances are stochastic and i.i.d. is a strong one, and numerous attempts have been made over the years to generalize the noise model.
One of the most important and prominent generalizations is that of robust control, also known as $H_{\infty}$-control. The goal of H-control is to come up with a more robust solution, by considering the optimal solution vs. any possible perturbation sequence from a certain family. Formally, we aim to solve:
$\begin{array}[]{c}\min_{\mathbf{u}_{1:T}(\mathbf{x})}\max_{\mathbf{w}_{1:T}}\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}{Q}\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}{R}\mathbf{u}_{t}\\ \text{ s.t. }\mathbf{x}_{t+1}=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t},\end{array}$
where we use the exact same notation as in previous sections.
Notice that this problem does not fall into the category of a Markov Decision Process, since $\mathbf{w}_{t}$’s are no longer stochastic. Therefore the typical Bellman optimality criterion does not apply.
Instead, the appropriate techniques to address this problem come from the theory of dynamical games. The above is, in fact, a two-player zero sum game. The first player tries to minimize the costs and the other attempts to maximize the costs, both players subject to the dynamics evolution constraint.
Using techniques from dynamic games, the following can be proven, see e.g [ZDG96b, BB08].
Theorem 5.4.
For the $H_{\infty}$ control problem with quadratic convex loss functions, the optimal solution is a linear policy of the state given by
$\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}.$
The proof is somewhat more involved than the LQR derivation, and is omitted from this text.
5.5 Bibliographic Remarks
Optimal control is considered to be the hallmark of control theory, and numerous excellent introductory texts cover it in detail, e.g. [Ste94, Hes18].
The LQR problem and its solution were introduced in the seminal work of Kalman [Kal60], who also introduced the eponymous filtering procedure for linear estimation.
While seemingly a natural corollary of equivalent results for MDPs, the existence of value functions for linear control is made subtle by the continuous nature of the state space and stabilizability, or the lack thereof. A discussion on sufficient and necessary conditions to existence of solutions to Riccati equations can be found in [Ber07, HSK96].
The development of $H_{\infty}$ control was driven by an observation made in the seminal paper of Doyle [Doy78] that, unlike in the fully observed case, the solution to the optimal control problem for partially observed linear systems (which we study in later chapters) is not naturally robust to miss-specification of system parameters. This suggests that robustness specifications constitute yet an another axis of performance for control algorithms, and may at times be at odds with the objective of optimal control. For in-depth treatment of $H_{\infty}$ control see [ZDG96b]. The text [BB08] presents a succinct derivation of robust controllers entirely in the state space, as the solution to a dynamic game between the controller and an adversary.
Although developed in the linear setting, adaptations of the LQR controller are also widely used to control non-linear systems [Li05, LK13].
5.6 Exercises
5.1. Prove Theorem 5.2 for a time-varying linear dynamical system.
5.2. Derive the solution to the infinite horizon LQR with zero mean Gaussian noise.
Part II Basics of Nonstochastic Control
Chapter 6 Policy classes for dynamical systems
In chapter 5, we studied the LQR problem, and saw that for linear dynamical systems, the optimal control takes a very special form: it is a linear function of the state. The Markovian property of the Markov Decision Process framework for reinforcement learning suggests that the optimal policy is always a function of the current state. It is characterized by the Bellman optimality equation, and can be computed in complexity proportional to the number of states and actions using, for example, the value iteration algorithm which we analyzed in a previous chapter.
However, this changes when we deviate from the MDP framework, which may be warranted for a variety of reasons:
  1. 1.
    The state is not fully observable.
  2. 2.
    The reward functions change online.
  3. 3.
    The transition mapping changes is not fixed and changes online.
  4. 4.
    The transitions or costs are subject to adversarial noise.
In such scenarios involving significant deviations from the MDP framework, classical notions of optimality may be ill-defined. Or if they do exist, the optimal policy may no longer be a mapping from immediate state to action. Rather, the learner may be required to use other signals available to her when choosing an action that is most beneficial in the long run. To this we add a significant consideration: computation of optimal state-based policies may be computationally intractable.
This motivates the current chapter which studies more general policies. We consider several options for control policies, as well as the relationships between them, in the context of continuous control. A control policy can take action according to several signals that may be available and observable as listed below.
  1. 1.
    State Action Control policies.
    The system state, denoted $\mathbf{x}_{t}$, if it is observable. Policies that map the state to action are called State-Feedback or State-Action policies, and are the most commonly considered in the literature. As we have seen in chapter 5, the optimal policy for LDS with stochastic noise is a linear state-feedback policy.
    More generally, a State-Action policy may depend on several previous states in a linear or non-linear manner. The most commonly considered policy class is a linear combination of previous states, but any (non-linear) mapping of the form
    $\mathbf{u}_{t}=\pi(\mathbf{x}_{t-1:t-h})=\pi(\mathbf{x}_{t-1},...,\mathbf{x}_{t-h})\ ,\ \pi\in{\mathbb{R}}^{d_{\mathbf{x}}\times h}\mapsto{\mathbb{R}}^{d_{\mathbf{u}}},$
    can be considered. The parameter $h>0$ is the number of most recent states that the policy considers when computing the next action.
  2. 2.
    Disturbance Action Control policies.
    If the perturbation to the system, denoted $\mathbf{w}_{t}$, is known, Disturbance Action Control policies (DAC) choose an action based upon the current and potentially previous disturbances,
    $\mathbf{u}_{t}=\pi(\mathbf{w}_{t-1:t-h})\ ,\ \pi\in{\mathbb{R}}^{d_{\mathbf{w}}\times h}\mapsto{\mathbb{R}}^{d_{\mathbf{u}}}.$
    For time-invariant linear dynamical systems, the two policies classes, linear DAC and linear SAC, are equivalent, since the states and disturbances are linearly related under DAC and SAC policies. We give a formal theorem that quantifies this relationship later in this chapter.
    However, for certain scenarios, the policy class of DAC is significantly better behaved than state-action control policies from a computational viewpoint. We shall explore efficient algorithms that learn optimal policies from this class in the next chapter.
  3. 3.
    Observation Action Control policies.
    In the case of dynamical systems that are partially observed and/or unknown, both policy classes above are infeasible, since the state and disturbance are similarly unknown. A natural policy parameterization is linear in the past observations. A special case of such policies is the class of Linear Dynamical Controllers, as stated below.
  4. 4.
    Linear Dynamical Control policies.
    A particularly useful policy class for partially observed linear dynamical systems is Linear Dynamical Controllers. These policies simulate an “internal linear dynamical system” to estimate the hidden state, and choose the action at any time step as a linear function of this estimate.
    A linear dynamic controller $\pi$ has parameters $\left(A_{\pi},{B}_{\pi},C_{\pi},{D}_{\pi}\right)$, and chooses a control at time $t$ according to
    $\mathbf{u}_{t}^{\pi}=C_{\pi}\mathbf{s}_{t}+D_{\pi}\mathbf{x}_{t}\ ,\ \mathbf{s}_{t+1}=A_{\pi}\mathbf{s}_{t}+B_{\pi}\mathbf{x}_{t}.$
  5. 5.
    Disturbance Response Control policies.
    Another important class of policies for control with partial observability is known as the disturbance response controllers, or DRC. The motivation for this control policy class is computational: in later chapters we show efficient algorithms that learn policies within this class.
    The main signal used by the DRC policy class is called Nature’s y’s, denoted by $\mathbf{y}^{\mathbf{nat}}_{t}$. This signal is the would-be observation at time $t$ assuming all inputs to the system were zero from the beginning of time. Mathematically, DRC policies choose an action as a function of previous Nature’s y’s signals, i.e.,
    $\mathbf{u}_{t}=\pi(\mathbf{y}^{\mathbf{nat}}_{t:t-h})\ ,\ \pi:{\mathbb{R}}^{d_{\mathbf{y}}\times(h+1)}\mapsto{\mathbb{R}}^{d_{\mathbf{u}}}.$
6.1 Relating the power of policy classes
We motivated the consideration of nonstandard policy classes by the need to deviate from the classical reinforcement learning model of Markov Decision Processes. Another important consideration is computational; some policy classes we consider will yield efficient algorithms, whereas in standard representation none are known.
To be able to give rigorous statements on the performance of methods that optimize these new policy classes, it is crucial to be able to relate their representation power. Furthermore, as we consider adversarial perturbations, we shall compare policies in terms of how they perform in hindsight, rather than using any absolute metric of optimality. This motivates our main comparison metric as follows.
Definition 6.1.
We say that a class of policies $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ if the following clause is satisfied: for every $\pi_{2}\in\Pi_{2}$, there exists $\pi_{1}\in\Pi_{1}$ such that, simultaneously for all sequences of $T$ disturbances and cost functions, it holds that
$\sum_{t=1}^{T}\left|c_{t}(\mathbf{x}_{t}^{\pi_{1}},\mathbf{u}_{t}^{\pi_{1}})-c_{t}(\mathbf{x}_{t}^{\pi_{2}},\mathbf{u}_{t}^{\pi_{2}})\right|\leq T\varepsilon.$
The significance of this definition is that if a simple policy class approximates another more complicated class, then it suffices to consider only the first. The usefulness of this definition stems from its implication for regret minimization, as per definition 1.3, as follows.
Lemma 6.2.
Let policy class $\Pi_{1}$ $\varepsilon$-approximate policy class $\Pi_{2}$. Then for any control algorithm ${\mathcal{A}}$ we have
$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{2})\leq\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{1})+\varepsilon T.$
The proof of this lemma is left as an exercise to the reader (Exercise 6.4). We now proceed to study the approximation relationships between the policy classes we surveyed earlier.
6.2 Quantitative Comparison of Policy Classes for Time Varying Linear Dynamical Systems
In this section we formally relate the power of different complexity classes for the special case of time-varying linear dynamical systems. Recall from Chapter 4 that time varying linear dynamical systems were been proposed by Lyapunov as a general methodology to study nonlinear dynamics, and thus are of broad interest.
For the rest of this section we assume full observation of the state and let
$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
A sequence of stabilizing linear controllers for such a system is denoted by $\{K_{t}\}$. Schematically, the relationships we describe below are given in Figure 6.1 for fully observed systems, and in Figure 8.1 for partially observed systems.
Figure 6.1: Schematic relationship between linear Disturbance-Action Control policies and linear State-Action Control policies of increasing diameter for linear time invariant systems.
6.2.1 Linear Policies
The most elementary and well-studied policy classes for control of linear dynamical systems are linear policies, i.e. those that choose a control signal as a (fixed) linear function of the state. The most important subcase of linear policies are those that are $\gamma$-stabilizing, meaning that the signal they create does not cause an unbounded state. This was formalized in Definition 2.4. Formally,
Definition 6.3 (Linear Policies).
A linear controller $K\in{\mathbb{R}}^{d_{u}\times d_{x}}$ is a linear operator mapping state to control. We denote by $\Pi_{\gamma}^{LIN}$ the set of all linear controllers that are $\gamma$-stabilizing and are bounded in Fronbenius norm by $\|K\|\leq\gamma$.
6.2.2 Linear Dynamic Controllers
The next policy class is particularly useful for linear dynamical systems with partial observation. Linear Dynamical Controllers simulate an “internal linear dynamical system” so to recover a hidden state, and play a linear function over this hidden state. The formal definition is given below.
Definition 6.4 (Linear Dynamic Controllers).
A linear dynamic controller $\pi$ has parameters $\left(A_{\pi},{B}_{\pi},C_{\pi}\right)$, and chooses a control at time $t$ as
$\mathbf{u}_{t}^{\pi}=C_{\pi}\mathbf{s}_{t}\ ,\ \mathbf{s}_{t}=A_{\pi}\mathbf{s}_{t-1}+B_{\pi}\mathbf{x}_{t}.$
Clearly the class of LDCs approximates linear policies without any error, since we can take $A_{\pi},B_{\pi},C_{\pi}$ to be zero, and obtain a purely linear policy.
We denote by $\Pi^{LDC}_{\gamma}$ the class of all LDC that are $\gamma$-stabilizing as per Definition 2.4 and Exercise 4.6. This implies the existence of parameters $\frac{\kappa}{\delta}\leq\gamma$ that satisfy
$\forall i\ ,\ \|A_{\pi}^{i}\|\leq\kappa(1-\delta)^{i}\ ,\ (\|B_{\pi}\|+\|C_{\pi}\|)\leq\kappa.$
LDCs are an important policy class for control of linear dynamical systems since they perform favorably in presence of partial observability. Furthermore, the optimal control law for latent LDS with Gaussian disturbances and quadratic loss functions is contained in the class of LDCs.
6.2.3 Generalized Linear Controllers
We present a generalization of linear controllers below. This generalization class is at least as expressive as LDCs.
Definition 6.5 (Generalized Linear Controllers).
A generalized linear controller $\pi$ has parameters $M_{0:h}$, and chooses a control at time $t$ according to
$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}M_{i}\mathbf{x}_{t-i}.$
An important concept for generalized linear controllers is that of lifting, or creating a larger linear system that subsumes the GLC. Given a certain GLC with memory $h$, consider the system given by
$\mathbf{z}_{t+1}=\tilde{A}_{t}\mathbf{z}_{t}+\tilde{B}_{t}\mathbf{u}_{t}+\tilde{C}_{t}\mathbf{w}_{t},$
where
$\tilde{A}_{t}=\begin{array}[]{|c|c|c|c|c|}\hline\cr A_{t}&0&\ldots&0&0\\ \hline\cr I&0&\ldots&0&0\\ \hline\cr 0&I&\ldots&0&0\\ \hline\cr\vdots&\vdots&\cdots&\vdots&\vdots\\ \hline\cr 0&0&\cdots&I&0\\ \hline\cr\end{array}\ \ ,\ \ \tilde{B}_{t}=\begin{array}[]{|c|}\hline\cr B_{t}\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}\ \ ,\ \ \tilde{C_{t}}=\begin{array}[]{|c|}\hline\cr I\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}.$
It can be seen that a GLC for the original system is simply a linear controller $\mathbf{u}_{t}^{\pi}=[M_{0}\dots M_{h}]\mathbf{z}_{t}$ for the lifted system! This observation will be useful to us later on.
We denote by $\Pi^{GLC}_{h,\gamma}$ the class of all GLC with $\sum_{i=0}^{h}\|M_{i}\|\leq\gamma$, and $\mathbf{M}_{1:h}$ that is $\gamma$-stabilizing as per Definition 2.4. The following lemma (proof left as as exercise) shows that GLC approximate LDC (and of course linear controllers):
Lemma 6.6.
The class $\Pi^{GLC}_{h,\gamma}$ $\varepsilon$-approximates the class $\Pi^{LDC}_{\gamma^{\prime}}$ for $h=\Omega(\gamma^{\prime}\log(\frac{\gamma^{\prime}}{\varepsilon}))$, and $\gamma=2\gamma^{\prime 3}$.
The class of GLC is thus very general and intuitive. It is, however, difficult to work with for the regret metric, due to its counterfactual structure.
6.2.4 Disturbance Action Controllers
We next define the most important set of policies in nonstochastic control, known as the disturbance action controllers, or DAC. The main advantages of DAC as opposed to GLC, LDC and linear controllers are that the cost of control is convex in their parametrization, and thus allows for provably efficient optimization. Moreover, this set of policies is stronger than GLC in the sense that it $\varepsilon$-approximates them, as we shortly show.
Definition 6.7 (Disturbance Action Controller).
A disturbance action policy $\pi_{K_{1:T},M_{1:h}}$ is parameterized by the matrices $M_{0:h}=[M_{0},{M}_{1},\ldots,{M}_{h}]$ and stabilizing controllers $\{K_{t}\}$. It outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
$\mathbf{u}_{t}^{\pi}={K}_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i}+M_{0}.$
Without loss of generality, we take $M_{0}=0$, as we can add a fictitious unit coordinate to the perturbations $\{\mathbf{w}_{t}\}$, and modify the system accordingly. This will allow a constant shift in the control.
Denote by $\Pi^{DAC}_{h,\gamma}$ the set of all disturbance action policies with $K_{1:T}$ being $\gamma$-stabilizing linear controllers for the system and
$\sum_{i=1}^{h}\|M_{i}\|\leq\gamma.$
The class of DAC always produces a stabilizing control signal whenever $K_{1:T}$ are sequentially stabilizing for the entire system $\{A_{t},B_{t}\}$. This important fact can be seen in the following lemma.
Lemma 6.8.
Any DAC policy in $\Pi_{h,\gamma}^{DAC}$ is $\gamma^{2}$-stabilizing.
Proof.
For convenience, denote by $M_{0}=I$, and the fictitious noise term $\tilde{\mathbf{w}}_{t}=\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$. First, due to the bound on $M_{1:h}$ and $\mathbf{w}_{t}$, we have that
$\|\tilde{\mathbf{w}}_{t}\|=\|\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}\|\leq\sum_{i=0}^{h}\|M_{i}\|\|\mathbf{w}_{t-i}\|\leq\gamma,$
for $\|M\|$ being the spectral norm of $M$.
Next, since $K_{t}$’s are stabilizing, we can bound the historically distant terms as below. Observe that the state at time $t$, following any DAC policy $\pi$, is given by
$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}$
$=A_{t}\mathbf{x}_{t}+B_{t}(K_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i})+\mathbf{w}_{t}$
$=(A_{t}+B_{t}K_{t})\mathbf{x}_{t}+B_{t}\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$ using $\text{using }M_{0}=I$
$=\sum_{j=0}^{t-1}\left(\left[\prod_{\tau=t-j}^{t}(A_{\tau}+B_{\tau}K_{\tau})\right]B_{t-j}\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i-j}\right)$
$=\sum_{j=0}^{t-1}\left(\left[\prod_{\tau=t-j}^{t}(A_{\tau}+B_{\tau}K_{\tau})\right]B_{t-j}\tilde{\mathbf{w}}_{t-j}\right).$
We can thus view the states $x_{t}$ as generated by a $\gamma$-stabilizing policy over noises that are of magnitude $\gamma$, and hence bounded by $\gamma^{2}$. ∎
The above lemma is already a good indication about DAC: they generate controls that at least don’t cause a blow up. But how expressive are these policies in terms of reaching the optimal solution?
One of the main appeals of the DAC class is the following approximation lemma which shows that for time-invariant systems DAC are in fact a stronger class than GLC, and thus LDC. This is a strong guarantee, since it implies that competing with the best policy in DAC implies competing with the best GLC/LDC.
We start by showing that DAC approximate all linear controllers. This is not a vacuous statement, since the $K_{t}$ terms for DAC is now only assumed to be stabilizing, and not necessarily optimal for the actual sequence.
Lemma 6.9.
Consider a time-invariant linear dynamical system $\{A,B\}$. The class $\Pi^{DAC}_{h,\gamma}$ $\varepsilon$-approximates the class $\Pi^{LIN}_{\gamma^{\prime}}$ for $h=\Omega(\gamma^{\prime}\log(\frac{\gamma^{\prime}}{\varepsilon}))$ and $\gamma=\gamma^{\prime 2}$.
Proof.
Disturbance-action policies use a $\gamma$-stabilizing linear policy on which they superimpose linear functions of disturbances – fix such a policy, and let $K_{0}$ be this linear stabilizing component. The state sequence thus produced is identical to that obtained by running a DAC with $\mathbf{0}$ as the stabilizing controller on the system $(A+BK_{0},B)$. Similarly, the state sequence produced by the execution of a linear policy $K$ on $(A,B)$ is identical to that of the linear policy $K-K_{0}$ on the system $(A+BK_{0},B)$.
Lemma 6.10.
Let $\mathbf{x}_{t}^{\pi}(A,B),\mathbf{u}_{t}^{\pi}(A,B)$ be the state-action sequence produced by a policy $\pi$ when executed on a linear dynamical system $(A,B)$. For any $\varepsilon>0$, if two policies $\pi_{1},\pi_{2}$ satisfy for all $t$ that
$\|\mathbf{x}_{t}^{\pi_{1}}(A+BK_{0},B)-\mathbf{x}_{t}^{\pi_{2}}(A+BK_{0},B)\|\leq\varepsilon,$
$\|\mathbf{u}_{t}^{\pi_{1}}(A+BK_{0},B)-\mathbf{u}_{t}^{\pi_{2}}(A+BK_{0},B)\|\leq\varepsilon$
when executed on some dynamical system $(A+BK_{0},B)$, then when the corresponding policies are executed on the dynamical system $(A,B)$ the iterates guarantee
$\max\{\|\mathbf{x}_{t}^{\pi_{1}}(A,B)-\mathbf{x}_{t}^{\pi_{2}}(A,B)\|,\|\mathbf{u}_{t}^{\pi_{1}}(A,B)-\mathbf{u}_{t}^{\pi_{2}}(A,B)\|\}\leq(1+\|K_{0}\|)\varepsilon.$
With the aid of the above lemma, it will suffice to guarantee that the state-action sequence produced by the linear policy $K$ and an appropriately chosen DAC with the stabilizing part set to $\mathbf{0}$ on an intrinsically $\gamma$-stable system $(A,B)$ are similar enough. To prove this, we map any linear policy $K$ to a disturbance-action policy that approximates it arbitrarily well for some history parameter $h$. Specifically, for a given system $A,B$ and linear policy $K$, let $\pi_{K}$ be the policy given by
$\forall i\in[0,h]\ ,\ M_{i}=K(A+BK)^{i}.$
Now consider any trajectory that results from applying the linear policy for the given system. Let $\mathbf{x}_{1:T},\mathbf{u}_{1:T}$ be the states and controls of the linear policy $K$ for a certain trajectory, as affected by a given noise signal. Let $\tilde{\mathbf{x}}_{1:T},\tilde{\mathbf{u}}_{1:T}$ be the set of states and controls that would have been generated with the same noise sequence, but with the policy $\pi_{K}$. We first observe that control sequences produced by $K$ and $\pi_{K}$ are exponentially close.
$\mathbf{u}_{t+1}=K\mathbf{x}_{t+1}$ $=K(A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t})$
$=K(A+BK)\mathbf{x}_{t}+K\mathbf{w}_{t}$
$=\sum_{i=0}^{t}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}$
$=\sum_{i=0}^{h}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}+Z_{t}$
$=\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}+Z_{t}$
$=\tilde{\mathbf{u}}_{t+1}+Z_{t}$
where $Z_{t}=\sum_{i=h+1}^{t}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}$. The magnitude of this residual term can be bounded, due to the existence of $\frac{\kappa}{\delta}\leq\gamma^{\prime}$ from Exercise 4.6, as
$|Z_{t}|$ $\leq\left\|\sum_{i=h+1}^{t}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}\right\|$
$\leq\sum_{i=h+1}^{t}(1-\delta)^{i+1}\left\|K\mathbf{w}_{t-i}\right\|$
[…]
$=\kappa\frac{1}{\delta}e^{-\delta h}\leq\varepsilon.$
We thus conclude that $\|\mathbf{u}_{t+1}-\tilde{\mathbf{u}}_{t+1}\|\leq{\varepsilon}$.
To make a similar observation on the corresponding state sequences, first note that
$\mathbf{x}_{t+1}=\sum_{i=0}^{t}A^{i}\left(\mathbf{w}_{t-i}+B\mathbf{u}_{t-i}\right).$
Therefore, since $\mathbf{x}_{t}$ and $\tilde{\mathbf{x}}_{t}$ share the same perturbation sequences as a part of their description, their difference can be bounded as a geometrically weighted sum of the difference in control inputs $\mathbf{u}_{t}$ and $\tilde{\mathbf{u}}_{t}$ that constitute them. Concretely, we have
$\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|$ $=\left\|\sum_{i=1}^{t}A^{i-1}B(\mathbf{u}_{t-i}-\tilde{\mathbf{u}}_{t-i})\right\|$
$\leq\left\|\sum_{i=0}^{t}A^{i}B\right\|\max_{i\leq t}\|\mathbf{u}_{t}-\tilde{\mathbf{u}}_{i}\|\leq\frac{\kappa}{\delta}\varepsilon,$
by using the $\gamma$-stability of $(A,B)$. The approximation in terms of states and controls implies $\varepsilon$-approximation of policy class for any Lipchitz cost sequence, concluding the proof.
Recall that GLC can be lifted to a linear dynamical system. Thus a corollary of the above lemma is as follows, and is left as Exercise 6.4.
Corollary 6.11.
Consider a time invariant linear dynamical system given by $\{A,B\}$. The class $\Pi^{DAC}_{h,\gamma}$ $\varepsilon$-approximates the class $\Pi^{GLC}_{h^{\prime},\gamma^{\prime}}$, for parameters $h=O(h^{\prime}+\gamma^{\prime}\log\frac{\gamma^{\prime}}{\varepsilon})$, $\gamma=O(\gamma^{\prime 2})$.
6.3 Bibliographic Remarks
Linear state-feedback policies are a mainstay of control theory both in applications, due to their simplicity, and in theory, owing to their optimality in deterministic and stochastic settings as demonstrated in [Kal60]. The methods for computing such policies are often based in dynamic programming. Often motivated by the need to impose constraints beyond optimality such as robustness, various other policy parametrizations were introduced. An early example is the Youla parametrization, sometimes called Q-parametrization [YJB76]. Later on, the introduction of system level synthesis [DMW+17, ADLM19], in a similar vein, proved useful for distributed control, especially so given its emphasis on state-space representations. For a comparison of these and additional ones such as input-output parameterizations [FZPK19], see [ZFP+20].
The Disturbance Action Control parametrization, introduced in [ABH+19], is similar to system level synthesis. Like SLS, it is a convex parametrization that is amenable to the tools of convex optimization. However, unlike SLS, it is a more compact representation that has logarithmically many parameters in its natural representation in terms of the desired approximation guarantee against linear policies. This makes it better suited to machine learning.
6.4 Exercises
6.1. Prove Lemma 6.2.
6.2. Prove Lemma 6.6.
6.3. Prove Lemma 6.10.
6.4. Prove Corollary 6.11.
Chapter 7 Online Nonstochatic Control
In this chapter we consider the main new directions that have been taken in the machine learning literature for the control problem.
These new developments are a major departure from classical control, and in recent years have produced methods that compete favorably and many times surpass classical methods. A few characteristics of the new approaches are listed here:
  1. 1.
    Instead of pre-computing an optimal policy, online control seeks to learn a policy using online learning techniques. This fits the framework known as “adaptive control”.
  2. 2.
    In contrast to adaptive control, online control optimizes a rigorous performance metric of regret vs. the best policy in hindsight. Online and non-stochastic control algorithms have provable finite-time regret guarantees.
  3. 3.
    In online nonstochastic control, the loss functions can be general Lipschitz convex losses that are chosen adversarially, as opposed to known quadratic losses in classical control.
  4. 4.
    Online and non-stochastic control deploy improper learning by convex relaxation: this means they compete with a certain policy class using a different policy class. This difference in learned policies allows us to side-step the computational hardness results we have previously explored, and obtain rigorous guarantees for efficient algorithms.
7.1 From Optimal and Robust to Online Control
Recall that the problem of optimal control of a linear time-invariant dynamical system (LDS) is given by the equations
$\min_{\mathbf{u}(\mathbf{x})}$ $\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$
subject to $\text{subject to }\mathbf{x}_{t+1}$ $={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t},$
where $\mathbf{x}_{t}$ is the state of the system, $\mathbf{u}_{t}$ is the control input, and $\mathbf{w}_{t}$ is the perturbation at time $t$.
Previous chapters showed the optimal control policy to be a linear function of the state when we assumed $\left\{\mathbf{w}_{t}\right\}_{t\in[T]}$ to be i.i.d. zero mean and bounded variance perturbations, and $c_{t}$ to be quadratic loss functions. Specifically, for positive semi-definite matrices ${Q}$ and ${R}$, the mathematical control formulation becomes
$\min_{\mathbf{u}(\mathbf{x})}$ $\sum_{t=1}^{T}\mathbf{x}_{t}^{T}{Q}\mathbf{x}_{t}+\mathbf{u}_{t}^{T}{R}\mathbf{u}_{t}$
subject to $\text{subject to }\mathbf{x}_{t+1}$ $={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t},\quad\mbox{ }\mathbf{w}_{t}\sim\mathcal{D},\mbox{ }\mathbb{E}\mathbf{w}_{t}=\mathbf{0}.$
We have shown in chapter 5 that the optimal solution to this formulation is given by the policy
$\mathbf{u}^{\ast}_{t}={K}\mathbf{x}_{t},$
where ${K}$ depends on ${A}$, ${B}$, ${Q}$ and ${R}$. The proof follows by inductively solving the Bellman optimality equation at each step $t$, starting from time step $T$ and moving backwards to time step $1$.
However, this optimal solution requires two conditions:
  1. 1.
    The loss function needs to be a quadratic function.
  2. 2.
    The perturbations are i.i.d. across time (with additional constraints on the mean and the variance).
These conditions are in fact necessary for solving the Bellman optimality equation. Thus, this approach is not as general as we would like. We are missing out on other loss functions commonly used in practice, e.g. cross entropy, or loss functions with regularization terms.
Furthermore and more importantly, many times the perturbations incurred in practice are not i.i.d., in which case we lose all provable guarantees!
An alternative to optimal control is robust, or $H_{\infty}$ control theory. In this formulation, the goal is to optimize the control for an adversarial noise in a minimax sense, as per the mathematical formulation:
$\min_{\mathbf{u}(\mathbf{x})}\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{1:T}\mathclose{\|}\leq 1}$ $\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$
subject to $\text{subject to }\mathbf{x}_{t+1}$ $={A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
This approach also suffers from several downsides:
  1. 1.
    The above problem is computationally ill-behaved for non-quadratic loss functions.
  2. 2.
    It is non-adaptive in nature, i.e. the system is too pessimistic for cases when the perturbations are well-behaved, like i.i.d. Gaussian perturbations.
The natural question at this point is: Does there exist an approach that gives the best of both robust and optimal control, and generalizes to any convex cost function? That’s exactly non-stochastic control theory which we define next. But first, we present a motivating example.
7.1.1 Motivating Example for Online Control
Consider the problem of flying a drone from source to destination subject to unknown weather conditions. The aerodynamics of flight can be modeled sufficiently well by time-varying linear dynamical systems, and existing techniques are perfectly capable of doing a great job for indoor flight. However, the wind conditions, rain and other uncertainties are a different story. Certainly the wind is not an i.i.d. Gaussian random variable! Optimal control theory, which exactly assumes this zero-mean noise, is therefore overly optimistic and inaccurate for our problem.
The designer might resort to robust control theory to account for all possible wind conditions, but this is overly pessimistic. What if we encounter a bright sunny day after all? Planning for the worst case would mean slow and fuel-inefficient flight.
We would like an adaptive control theory which allows us to attain the best of both worlds: an efficient and fast flight when the weather permits, and a careful drone maneuver when this is required by the conditions. Can we design a control theory that will allow us to take into account the specific instance perturbations and miss-specifications, and give us finite time provable guarantees? This is the subject of online non-stochastic control!
7.2 The Online Nonstochastic Control Problem
When dealing with non-stochastic, a.k.a. arbitrary or even adversarial perturbations, the optimal policy is not a-priory clear. Rather, an optimal policy for the observed perturbations can be determined only in hindsight.
This is the reason that non-stochastic control shifts the performance metric from optimality, to regret vs. the best policy in hindsight from a certain reference class. We have already defined regret in full generality in Definition 1.3, and in this section give a more detailed treatment that specializes to time varying linear dynmical systems.
At this point the reader may wonder: Why should we compare to a reference class, as opposed to all possible policies? There are two answers to this question:
First, the best policy in hindsight may be very complicated to describe and reason about. This argument was made by the mathematician Tyrell Rockafellar; see the bibliographic section for more details. Secondly, it can be shown that it is impossible, in general, to obtain sublinear regret vs. the best policy in hindsight. Rather, a different performance metric called competitive ratio can be analyzed. This will be surveyed later in the book.
We now turn to define policy regret with respect to a reference class of policies. In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$, and then observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$, according to an adversarially chosen loss function. Let $\Pi=\{\pi:\mathbf{u}\mapsto\mathbf{x}\}$ be a class of policies. The regret of the controller w.r.t. $\Pi$ is defined as follows.
Definition 7.1.
Let $\{A_{t},B_{t}\}$ be a time varying LDS. The regret of an online control algorithm over $T$ iterations w.r.t. the class of policies $\Pi$ is given by:
$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}$ $\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}))\right),$
where $\mathbf{u}_{t}={\mathcal{A}}(\mathbf{x}_{t})$ are the controls generated by ${\mathcal{A}}$, and $\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}$ are the counterfactual state sequence and controls under the policy $\pi$, i.e.
$\mathbf{u}_{t}^{\pi}$ $=\pi(\mathbf{x}_{t}^{\pi})$
$\mathbf{x}_{t+1}^{\pi}$ $=A_{t}\mathbf{x}_{t}^{\pi}+B_{t}\mathbf{u}_{t}^{\pi}+\mathbf{w}_{t}.$
Henceforth, if $T,\Pi$ and ${\mathcal{A}}$ are clear from the context, we just refer to $\mathrm{{regret}}$ without quantifiers.
The non-stochastic control problem can now be stated as the quest to find an efficient algorithm that minimizes the worst-case regret vs. meaningful policies classes that we examine henceforth. It is important to mention that the algorithm does not have to belong to the comparator set of policies $\Pi$! Indeed, in the most powerful results, we will see that the algorithm will learn a policy that is strictly not contained in the comparator class.
7.3 The Gradient Perturbation Controller
After exploring the relationships between different classes of controllers, we are ready to prove the main new algorithm in non-stochastic control, which gives a strong regret guarantee against adversarially chosen perturbations.
The algorithm is described in figure (3), for time varying linear dynamical systems. It is a specialization of the template introduced in Algorithm 1 from Section 1.6.2. The main idea is to learn the parametrization of a DAC policy using known online convex optimization algorithms, namely online gradient descent. We henceforth prove that this parametrization is essentially convex, allowing us to prove the main regret bound.
In algorithm (3) we denote by $\mathbf{x}_{t}({M}_{1:h})$ the hypothetical state arrived at by playing a DAC policy ${M}_{1:h}$ from the beginning of time; note that this may not be equal to the true state $\mathbf{x}_{t}$ of the dynamical system at time $t$. The counterfactual $\mathbf{u}_{t}({M}_{1:h})$ is defined similarly.
1:  Input: $h$, $\eta$, initialization $M_{1:h}^{1}$.
2:  Observe the linear system $(A_{t},B_{t})$, compute a stabilizing linear controller sequence ${K}_{t}$.
3:  for $t$ = $1\ldots T$ do
4:     Use control $\mbox{Use control }\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}{M}_{i}^{t}\mathbf{w}_{t-i}.$
5:     Observe state $\mathbf{x}_{t+1}$; compute noise $\mathbf{w}_{t}=\mathbf{x}_{t+1}-{A}_{t}\mathbf{x}_{t}-{B}_{t}\mathbf{u}_{t}$.
6:     Record an instantaneous cost $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$.
7:     Construct loss $\ell_{t}({M}_{1:h})=c_{t}(\mathbf{x}_{t}({M}_{1:h}),\mathbf{u}_{t}({M}_{1:h})).$Update ${M}_{1:h}^{t+1}\leftarrow{M}_{1:h}^{t}-\eta\nabla\ell_{t}({M}_{1:h}^{t}).$
8:  end for
Algorithm 3 Gradient Perturbation Controller(GPC)
The GPC algorithm comes with a very strong performance guaranteee: namely it guarantees vanishing worst case regret vs. the best DAC policy in hindsight. As per the relationship between DAC and other poicies, this implies vanishing regret vs. LDC, GLC and linear controllers. The formal statement is given as follows.
Theorem 7.2.
Assuming that
  • a
    The costs $c_{t}$ are convex, bounded, and have bounded gradients w.r.t. both arguments $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$.
  • b
    The matrices $\{{A}_{t},{B}_{t}\}$ are stabilizable, and have bounded $\ell_{2}$ norm.
Then the GPC algorithm (3) ensures that
$\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DAC}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right)\leq\mathcal{O}(\sqrt{T}).$
Furthermore, the time complexity of each loop of the algorithm is linear in the number of system parameters and logarithmic in T.
Proof sketch.
As per Theorem A.1 in Appendix A, the online gradient descent algorithm for convex loss functions ensures that the regret is $\mathcal{O}(\sqrt{T})$. However, to directly apply the OGD algorithm to our setting, we need to make sure that the following 3 conditions hold true:
  • The loss function should be a convex function in the variables ${M}_{1:h}$.
  • The loss function is time dependent because the loss depends on the current state, which further depends on the previous states. The state at each time step depends on the current variable estimates. Hence, $\mathbf{x}_{t}$ can be shown to be a function of $\left\{{M}^{i}_{1:h}\right\}_{i=1}^{t}$, where ${M}^{i}_{1:h}$ denotes the variables at step $i$. Thus, we need to ensure that optimizing the loss function is similar to optimizing a function which doesn’t have such a dependence on time.
First, we show that the loss function is convex w.r.t. to the variables ${M}_{1:h}$. This follows since the states and the controls are linear transformations of the variables.
Lemma 7.3.
The loss functions $\ell_{t}$ are convex in the variables ${M}_{1:h}$.
Proof.
Using $\mathbf{x}_{t}$ to denote $\mathbf{x}_{t}(M_{1:h})$ and $\mathbf{u}_{t}$ to denote $\mathbf{u}_{t}(M_{1:h})$, the loss function $\ell_{t}$ is given by
$\ell_{t}({M}_{1:h})=c_{t}(\mathbf{x}_{t},\mathbf{y}_{t}).$
Since the cost $c_{t}$ is a convex function w.r.t. its arguments, we simply need to show that $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{1:h}$. The state is given by
$\mathbf{x}_{t+1}={A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}$ $={A}_{t}\mathbf{x}_{t}+{B}_{t}\left({K}_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i}\right)+\mathbf{w}_{t}$
$=\tilde{A}_{t}\mathbf{x}_{t}+\left({B}_{t}\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i}+\mathbf{w}_{t}\right),$
where $\tilde{A}_{t}=A_{t}+B_{t}K_{t}$. By induction, we can further simplify
$\mathbf{x}_{t+1}$ $=\sum_{i=0}^{t}\left(\prod_{j=t-i+1}^{t}\tilde{A}_{j}\right)\left(B_{t-i}\sum_{j=1}^{h}{M}_{j}\mathbf{w}_{t-i-j}+\mathbf{w}_{t-i}\right),$ (7.1)
which is a linear function of the variables. In the above computation, for convenience, we have assumed that the initial state $\mathbf{x_{0}}$ is $\mathbf{0}$.
Similarly, the control $\mathbf{u}_{t}$ is given by
$\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i},$
and since $\mathbf{x}_{t}$ was shown to be linear in the variables, so is $\mathbf{u}_{t}$. Thus, we have shown that $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ are linear transformations of ${M}_{1:h}$ and hence, the loss function is convex in ${M}_{1:h}$. (Why? See Exercise 7.5.) ∎
Now, by the regret bound implied by OGD, we have
$\sum_{t=1}^{T}l_{t}(M_{1:h}^{t})-\min_{M_{1:h}}\sum_{t=1}^{T}l_{t}(M_{1:h})\leq\mathcal{O}(\sqrt{T}).$ (7.2)
Finally, the actual loss $c_{t}$ at time $t$ is not calculated on $\mathbf{x}_{t}(M_{1:h})$, but rather on the true state $\mathbf{x}_{t}$, which in turn depends on different parameters $M_{1:h}^{i}$ for various historical times $i<t$. Nevertheless, $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ is well approximated by $l_{t}(M^{t}_{1:h})$, as shown next.
Lemma 7.4.
For all $t$, $|l_{t}(M^{t}_{1:h})-c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})|\leq\mathcal{O}(1/\sqrt{T})$.
Proof.
First, note that to achieve $\mathcal{O}(\sqrt{T})$ regret, OGD uses a step size of $\eta=1/\sqrt{T}$. Since the gradients of the loss functions are bounded, we have
$\|M_{1:h}^{t}-M_{1:h}^{t-i}\|\leq\sum_{s=t-i+1}^{t}\|M_{1:h}^{s}-M_{1:h}^{s-1}\|\leq\mathcal{O}\left(\frac{i}{\sqrt{T}}\right).$
Now, we use the fact that the iterates of OGD $M_{1:h}^{t}$ move slowly to establish that $\mathbf{x}_{t}$ and $\mathbf{x}_{t}(M^{t}_{1:h})$ are close in value. Similar to the expansion of $\mathbf{x}_{t}(M_{1:h})$ in Equation 7.1, $\mathbf{x}_{t}$ can be written as
$\mathbf{x}_{t+1}$ $=\sum_{i=0}^{t}\left(\prod_{j=t-i+1}^{t}\tilde{A}_{j}\right)\left(B_{t-i}\sum_{j=1}^{h}{M}^{t-i}_{j}\mathbf{w}_{t-i-j}+\mathbf{w}_{t-i}\right).$
Recall that the stability of the time-varying linear policy $\{K_{t}\}$ implies that for any $i\leq j$, $\|\prod_{t=i}^{j}\tilde{A}_{t}\|\leq\kappa(1-\delta)^{j-i}$. Therefore, there exists a constant $C\geq 0$ such that
$\|\mathbf{x}_{t+1}-\mathbf{x}_{t}(M_{1:h}^{t})\|$ $\leq\sum_{i=0}^{t}\left\|\left(\prod_{j=t-i+1}^{t}\tilde{A}_{j}\right)\right\|\|B_{t-i}\|\sum_{j=1}^{h}\|{M}^{t-i}_{j}-M^{t}_{j}\|\|\mathbf{w}_{t-i-j}\|$
$\leq Ch\eta\sum_{i=0}^{t}i(1-\gamma)^{i}$
$\leq\frac{Ch\eta}{\gamma^{2}}=\mathcal{O}\left(\frac{1}{\sqrt{T}}\right).$
This display also implies that $\|\mathbf{u}_{t}-\mathbf{u}_{t}(M_{1:h}^{t})\|=\mathcal{O}\left({1}/{\sqrt{T}}\right)$, since $\mathbf{u}_{t}$ and $\mathbf{u}_{t}(M_{1:h}^{t})$ only differ in the immediate state they operate on. By definition, $l_{t}(M_{1:h}^{t})=c_{t}(\mathbf{x}_{t}(M_{1:h}^{t}),\mathbf{u}_{t}(M_{1:h}^{t}))$. Finally, we note that $c_{t}$ has bounded derivatives finishing the proof of the lemma. ∎
Now, we combine the regret guarantee in equation 7.2 and the immediately preceding lemma as
$\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DAC}}\sum_{t=1}^{T}c_{t}(\hat{\mathbf{x}}_{t},\pi(\hat{\mathbf{x}}_{t}))$
$\leq\sum_{t=1}^{T}l_{t}(M_{1:h}^{t})-\min_{\pi\in\Pi^{DAC}}l_{t}(M_{1:h}^{\pi})+T\times\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)=\mathcal{O}(\sqrt{T}),$
implying the stated claim.
7.4 Bibliographic Remarks
The bulk of the material of this chapter is from [ABH+19]. The work of [ABH+19] defined the general online nonstochastic setting in which the disturbances and the cost functions are adversarially chosen, and where the cost functions are arbitrary convex functions. Follow up work by [AHS19] achieves logarithmic pseudo-regret for strongly convex, adversarially selected losses, and well-conditioned stochastic noise. Logarithmic regret algorithms were derived for online nonstochastic control in [FS20].
Online nonstochastic control with known dynamics was further extended in several directions. Control with bandit feedback for the costs was considered in [GHH20]. More refined regret metrics applicable to changing dynamics were studied in [GHM20, MGSH21, ZWZ22, Sin22]. [LDL21] studies algorithms that balance regret minimization while adhering to constraints. An application of nonstochastic control in iterative planning in the presence of perturbations is described in [AHMS21]. Deluca [GHM20], a control library, provides benchmarked implementations of many such algorithms.
Online nonstochastic control is an extension of the framework of online convex optimization [Haz16] when applied to policies, and thus to loss functions with memory [AHM13]. A similar regret notion was implicit in the work of [EDKM04]. The regret metric when applied to policies is different from the counterfactual notion of policy regret as defined by [ADT12].
The roots of online nonstochastic control can be traced to the online stochastic control setting, which we cover in Chapter 10.
An alternate performance metric for online control, which is also nonstochastic, is competitive ratio. This objective tracks the cost of the online controller and compares it to the best policy in hindsight. Control methods that minimize competitive ratio were initiated in [GW19]. This approach was extended in a series of papers [GLSW19, SLC+20, GH21]. Thus far the success of this approach was limited to fully observed linear dynamical systems, and further to quadratic costs when achieving the optimal competitive ratio. [GASH22] show that regret minimizers in the non-stochastic setting automatically achieve a near-optimal competitive ratio.
A precursor to nonstochastic control is a line of work, beginning with [AYBK14, CHK+18], focused on the control of stochastic systems under sequentially revealed adversarial costs functions. See [CK20, CCK22] for further developments.
7.5 Exercises
7.1. Let $l:{\mathbb{R}}^{d}\to{\mathbb{R}}$ be a convex function. Prove that for any matrix $A\in{\mathbb{R}}^{d\times n}$, $l(Ax)$ is convex for $x\in{\mathbb{R}}^{n}$.
Chapter 8 Policy Classes for Partially Observed LDS
In this chapter we consider linear time varying systems with partial observation, i.e. the state is not available to the controller. Instead, a (known or unknown) linear projection of the state is revealed. In full generality, recall from definition 1.1 partially observed dynamical system is given by the equations
$\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g_{t}(\mathbf{x}_{t}).$
As per our notation thus far, here $\mathbf{x}_{t}\in{\mathbb{R}}^{d_{x}}$ is the state of the system, $\mathbf{y}_{t}\in{\mathbb{R}}^{d_{y}}$ is the observation, $\mathbf{u}_{t}\in{\mathbb{R}}^{d_{u}}$ is the control input, and $\mathbf{w}_{t}\in{\mathbb{R}}^{d_{w}}$ is the perturbation. The function $g_{t}$ is a mapping from the state to the observed state, and the function $f_{t}$ is the transition function, given current state $\mathbf{x}_{t}$ and control input $\mathbf{u}_{t}$. The subscript indicates the relevant quantities at time step $t.$
Many of the control policies we consider in this chapter are general and can be applied to any dynamical system of the above formulation. However, following the paradigm of chapter 6, we relate the representation power of policy classes only for partially observed linear dynamical systems. These are given by the equations
$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}$
$\mathbf{y}_{t}$ $=C_{t}\mathbf{x}_{t}.$
Often, partially observed LDSs are defined with an additional noise term $\mathbf{v}_{t}$ corrupting the observations as $\mathbf{y}_{t}=C_{t}\mathbf{x}_{t}+\mathbf{v}_{t}$. Exercise 8.5 asks the reader to prove that even in such circumstances we may assume the previously stated form for partially observed systems without loss of any generality.
In the partial observation setting some of the control policy classes from chapter 6 are not well defined. We proceed to define policy classes and describe the relationship between them.
Figure 8.1: Schematic relationship between linear Disturbance-Response Control policies and Linear Dynamical Control policies of increasing dimensions (or number of parameters) for linear time invariant systems.
We limit the discussion in this chapter to stable partially observed systems, although the policy containment results described next hold more broadly.
8.1 Linear Observation Action Controllers
Linear Observation-action controllers are the most natural generalization of linear controllers to the partially observed case, and are defined as follows.
Definition 8.1.
A linear observation-action controller $K\in{\mathbb{R}}^{d_{u}\times d_{y}}$ is a linear operator mapping observation to control as $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}$. We denote by $\Pi_{\gamma}^{OAC}$ the set of all linear controllers bounded in Fronbenius norm by $\|K\|\leq\gamma$, that are also $\gamma$-stabilizing for given LDS.
While this class of policies is natural to consider, it turns out to be very inexpressive. Even for optimal control with stochastic noise, the optimal policy cannot be approximated by a linear observation-action policy (see Exercise 8.5). Therefore we proceed to define more expressive and useful policies. However, OACs will be useful in simplifying proofs later on.
8.1.1 Generalized Linear Observation Action Controllers
Similar to the full observation setting, we can generalize linear observation-action controllers to apply to a history of past observations. This additional generality, of taking a history of observations, greatly increases the expressive power of this policy class. The formal definition is below.
Definition 8.2.
Parameterized by matrices $K_{0:h}$, generalized linear observation-action controllers choose a control input as
$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}K_{i}\mathbf{y}_{t-i}.$
Let $\Pi^{GOAC}_{h,\gamma}$ be that class of all such policies satisfying $\sum_{i=0}^{h}\|K_{i}\|\leq\gamma$ that are also $\gamma$-stabilzing.
Generalized observation-action policies can be seen as a special case of observation-action controllers using the technique of lifting from Section 6.2.3. Consider any partially-observed system $(A_{t},B_{t},C_{t})$ and perturbation sequence $\mathbf{w}_{t}$. Consider the lifted system $(\tilde{A}_{t},\tilde{B}_{t},\tilde{C}_{t},\tilde{D}_{t})$ given by
$\mathbf{z}_{t+1}=\tilde{A}_{t}\mathbf{z}_{t}+\tilde{B}_{t}\mathbf{u}_{t}+\tilde{D}_{t}\mathbf{w}_{t},\quad\tilde{\mathbf{y}}_{t}=C_{t}\mathbf{z}_{t},$
where
$\tilde{A}_{t}=\begin{array}[]{|c|c|c|c|c|}\hline\cr A_{t}&0&\ldots&0&0\\ \hline\cr I&0&\ldots&0&0\\ \hline\cr 0&I&\ldots&0&0\\ \hline\cr\vdots&\vdots&\cdots&\vdots&\vdots\\ \hline\cr 0&0&\cdots&I&0\\ \hline\cr\end{array}\ \ ,\ \ \tilde{B}_{t}=\begin{array}[]{|c|}\hline\cr B_{t}\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}\ \ ,$
$\ \ \tilde{C}_{t}=\begin{array}[]{|c|c|c|c|}\hline\cr C_{t}&C_{t-1}&\dots&C_{t-h}\\ \hline\cr\end{array}\ \ ,\ \ \tilde{D_{t}}=\begin{array}[]{|c|}\hline\cr I\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}.$
The following lemma relates the original and lifted system (for a proof see exercise 8.5).
Lemma 8.3.
For any sequence of control inputs $\mathbf{u}_{t}$, it holds that $\tilde{\mathbf{y}}_{t}=[\mathbf{y}_{t},\mathbf{y}_{t-1},\dots\mathbf{y}_{t-h}]$.
The lemma above shows that generalized linear observation-action controllers are a special case of simple observation-action controllers such that $\mathbf{u}_{t}^{\pi}=[K_{0}\dots K_{h}]\tilde{\mathbf{y}}_{t}$ for the lifted system.
8.2 Linear Dynamic Controllers
The main motivation for the class of linear dynamical controllers, which we defined already in the context of full observability, is that of partially observed linear dynamical systems. The generalization of this class is as follows.
Definition 8.4 (Linear Dynamic Controllers for Partially Observed Systems).
A linear dynamic controller $\pi$ has parameters $\left(A_{\pi},{B}_{\pi},C_{\pi}\right)$, and chooses a control at time $t$ as
$\mathbf{u}_{t}^{\pi}=C_{\pi}\mathbf{s}_{t}\ ,\ \mathbf{s}_{t}=A_{\pi}\mathbf{s}_{t-1}+B_{\pi}\mathbf{y}_{t}.$
We denote by $\Pi^{LDC}_{\gamma}$ the class of all LDC that are $\gamma$-stabilizing as per Definition 2.4 and exercise 4.6, with parameters $\frac{\kappa}{\delta}\leq\gamma$ that satisfy
$\forall i\ ,\ \|A_{\pi}^{i}\|\leq\kappa(1-\delta)^{i}\ ,\ (\|B_{\pi}\|+\|C_{\pi}\|)\leq\kappa.$
Note that LDCs are $\varepsilon$-approximated by generalized observation-action controllers, namely GOACs. The proof is left as Exercise 9.4, on account of bering similar to that for the fully observed case.
Lemma 8.5.
Consider a partially-observed time-invariant linear dynamical system $\{A,B,C\}$. The class $\Pi^{GOAC}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{LDC}_{\gamma}$ for $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right),\gamma^{\prime}=2\gamma^{3}$.
8.3 Disturbance Response Controllers
The most important policy class in the context of nonstochastic control is that of Disturbance Response Controllers, or DRC. These policies act on a signal called Nature’s y’s, denoted $\mathbf{y}^{\mathbf{nat}}_{t}$, which is the would-be observation at time $t$ assuming all inputs to the system were zero from the beginning of time, or formally,
$\mathbf{x}^{\mathbf{nat}}_{t+1}=A_{t}\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}\ ,\ \mathbf{y}^{\mathbf{nat}}_{t}=C_{t}\mathbf{x}_{t}.$
A linear disturbance response policy $\pi$ is parameterized by the matrices $M_{1:h}=[{M}_{0},\ldots,{M}_{h}]$, and outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}{M}_{i}\mathbf{y}^{\mathbf{nat}}_{t-i}.$
Denote by $\Pi^{DRC}_{h,\gamma}$ the set of all disturbance response policies with $K_{1:T}$ being $\gamma$-stabilizing linear controllers for the system and
$\sum_{i=0}^{h}\|M_{i}\|\leq\gamma.$
We proceed to show that that the class of DRC approximates the class of LDC.
8.3.1 The Expressive Power of Disturbance Response Controllers
A DRC always produces controls that ensure the system remains state-bounded. This can be shown analogously to Lemma 6.8 (see Exercise 8.5).
While DRCs are contained inside the policy class of DACs, they are still very expressive. Note that a DAC policy can not be executed in a partially observed system from observations alone; even the knowledge of the entire observation sequence $\mathbf{y}_{t}$ may not uniquely determine $\mathbf{w}_{t}$.
We have already seen that LDCs are approximately captured by GOACs in Lemma 8.5. It can be shown that the class of DRCs with appropriate parameters $\varepsilon$-approximates the class of GOACs for linear time-invariant systems. Note that this observation in conjunction with Lemma 8.5 implies that DRCs also approximate LDCs.
To demonstrate that DRCs approximately capture GOACs, we instead prove that OACs, a weaker class of policies, are $\varepsilon$-approximated by DRCs. Nevertheless, this observation is sufficient to conclude the stronger claim, given that GOACs can be written as OACs on a lifted system as noted in Lemma 8.3, and simultaneously that a DRC with history $H^{\prime}$ on the lifted system can be written as a DRC with history $H^{\prime}+h$ over the original system.
Lemma 8.6.
Consider a $\gamma$-stable time-invariant linear dynamical system $(A,B,C)$. The class $\Pi^{DRC}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{OAC}_{\gamma}$ for parameters $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right)$, $\gamma^{\prime}=O(\gamma^{6})$.
Proof.
Recall that for the fully observed case the state-action pairs produced by execution of a linear policy can be written as linear functions of the perturbations. It is this observation that allowed us to prove that DACs capture linear policies for fully observed systems. We prove an analogue for partially observed systems: The observation-action pairs produced by the execution of OACs can be written as linear functions of nature’s $y$’s. Thereafter, we argue that parameterizing the controllers to be linear in nature’s $y$’s (as DRCs do) allows us to encompass OACs.
We introduce an augmented partially observed linear system that helps us write the observation-action pairs produced by a OAC as a linear function of nature’s $y$’s. Crucially, the new system has the same nature’s $y$’s as the original system.
Lemma 8.7.
For any partially-observed system $(A_{t},B_{t},C_{t})$ and perturbation sequence $\mathbf{w}_{t}$ producing observations $\mathbf{y}_{t}$, construct a lifted system $(\tilde{A}_{t},\tilde{B}_{t},\tilde{C}_{t})$ given by
$\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t+1}\\ \Delta_{t+1}\end{bmatrix}=\begin{bmatrix}A_{t}&0\\ 0&A_{t}\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta_{t}\end{bmatrix}+\begin{bmatrix}0\\ B_{t}\end{bmatrix}\mathbf{u}_{t}+\begin{bmatrix}I\\ 0\end{bmatrix}\mathbf{w}_{t},\quad\tilde{\mathbf{y}}_{t}=\begin{bmatrix}C_{t}&C_{t}\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta_{t}\end{bmatrix}.$
It holds uniformly for all time $t$ that $\widetilde{\mathbf{y}^{\mathbf{nat}}_{t}}={\mathbf{y}^{\mathbf{nat}}_{t}}$ and further for any sequence of control inputs $\mathbf{u}_{t}$ that $\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}$ and $\mathbf{x}_{t}=\mathbf{x}^{\mathbf{nat}}_{t}+\Delta_{t}$.
Substituting $\mathbf{u}_{t}^{\pi}=K\mathbf{y}^{\pi}_{t}=KC(\mathbf{x}^{\mathbf{nat}}_{t}+\Delta^{\pi}_{t})$, the state-action sequence produced by a OLIN $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}^{\pi}$ can be written as
$\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t+1}\\ \Delta^{\pi}_{t+1}\end{bmatrix}=\begin{bmatrix}A&0\\ BKC&A+BKC\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta^{\pi}_{t}\end{bmatrix}+\begin{bmatrix}I\\ 0\end{bmatrix}\mathbf{w}_{t},\quad\begin{bmatrix}\mathbf{y}^{\pi}_{t}\\ \mathbf{u}^{\pi}_{t}\end{bmatrix}=\begin{bmatrix}C&C\\ KC&KC\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta^{\pi}_{t}\end{bmatrix},$
or using that $\mathbf{y}^{\mathbf{nat}}_{t}=C\mathbf{x}^{\mathbf{nat}}_{t}$, equivalently as
$\Delta^{\pi}_{t+1}=(A+BKC)\Delta^{\pi}_{t}+BK\mathbf{y}^{\mathbf{nat}}_{t},\quad\mathbf{y}^{\pi}_{t}=C\Delta^{\pi}_{t}+\mathbf{y}^{\mathbf{nat}}_{t},\quad\mathbf{u}^{\pi}_{t}=KC\Delta_{t}+K\mathbf{y}^{\mathbf{nat}}_{t}.$
Now, define $M_{i}=KC(A+BKC)^{i-1}BK$ for $i>0$ and $M_{0}=K$, let $\mathbf{y}_{t},\mathbf{u}_{t}$ be the observation-action sequence produced by this DRC. We have
$\mathbf{u}^{\pi}_{t}$ $=K\mathbf{y}^{\mathbf{nat}}_{t}+KC\Delta^{\pi}_{t}$
$=K\mathbf{y}^{\mathbf{nat}}_{t}+KCBK\mathbf{y}^{\mathbf{nat}}_{t-1}+KC(A+BKC)\Delta^{\pi}_{t-1}$
$=K\mathbf{y}^{\mathbf{nat}}_{t}+\sum_{i>1}KC(A+BKC)^{i-1}BK\mathbf{y}^{\mathbf{nat}}_{t-i}$
$=\sum_{i=0}^{h^{\prime}}M_{i}\mathbf{y}^{\mathbf{nat}}_{t-i}+\sum_{i>H^{\prime}}KC(A+BKC)^{i-1}BK\mathbf{y}^{\mathbf{nat}}_{t-i}$
$=\mathbf{u}_{t}+Z_{t}.$
Given $\gamma$-stabilizability of the OAC $K$ and $\gamma$-intrinsic stability of the of the system $(A,B,C)$, we have that for some $\kappa,\delta$ suc that $\frac{\kappa}{\delta}\leq\gamma$ that
$\|Z_{t}\|$ $\leq\sum_{i>h^{\prime}}\|KC(A+BKC)^{i-1}BK\|\|\mathbf{y}^{\mathbf{nat}}_{t-i}\|$
$\leq\sum_{i>h^{\prime}}\|KC(A+BKC)^{i-1}BK\|\left\|\sum_{j>1}CA^{j-1}B\mathbf{w}_{t-i-j}\right\|$
$\leq\frac{\kappa^{5}}{\delta}\sum_{i>h^{\prime}}\|(A+BKC)^{i-1}\|\leq\frac{\kappa^{6}\gamma}{\delta}(1-\frac{1}{\delta})^{h^{\prime}}$
Therefore, $\|\mathbf{u}_{t}-\mathbf{u}^{\pi}_{t}\|\leq\varepsilon$ by the appropriate choice of $H^{\prime}$. Once again using the intrinsic stability of the LTI $(A,B,C)$ , we have
$\|\mathbf{y}^{\pi}_{t}-\mathbf{y}_{t}\|$ $=\left\|\sum_{i=1}^{t}CA^{i-1}B(\mathbf{u}^{\pi}_{t-i}-{\mathbf{u}}_{t-i})\right\|$
$\leq\left\|\sum_{i=0}^{t}CA^{i}B\right\|\max_{i\leq t}\|\mathbf{u}^{\pi}_{i}-{\mathbf{u}}_{i}\|\leq\frac{\kappa}{\delta}\varepsilon.$
The approximation in terms of observations and controls implies $\varepsilon$-approximation of policy class for any Lipchitz cost sequence, concluding the proof. ∎
8.4 Bibliographic Remarks
The Disturbance Response Parametrization was introduced in [SSH20] as an analogue of DAC for control of parially observed systems. It can be viewed as a succinct state-space representation of the Youla parameterization [YJB76]. For time-varying dynamics, the power of vaious policy classes was contrasted in [MGSH21].
8.5 Exercises
8.1. Consider a seemingly more general definition of a partially observed linear dynamical system, given below:
$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t},$
$\mathbf{y}_{t}$ $=C_{t}\mathbf{x}_{t}+\mathbf{v}_{t}.$
Construct a linear dynamical system parameterized via matrices $(A^{\prime}_{t},B^{\prime}_{t},C^{\prime}_{t})$ and perturbations $\mathbf{w}^{\prime}_{t}$ with a hidden state $\mathbf{z}_{t}$ of the form
$\mathbf{x}_{t+1}$ $=A^{\prime}_{t}\mathbf{z}_{t}+B^{\prime}_{t}\mathbf{u}_{t}+\mathbf{w}^{\prime}_{t},$
$\mathbf{y}^{\prime}_{t}$ $=C^{\prime}_{t}\mathbf{z}_{t}$
such that for any sequence of $\mathbf{u}_{t}$, $\mathbf{y}_{t}=\mathbf{y}^{\prime}_{t}$ holds uniformly over time. Additionally, verify that there exists such a construction of $\mathbf{z}_{t}$ where $\mathbf{w}^{\prime}_{t}$ are i.i.d. as long as $(\mathbf{w}_{t},\mathbf{v}_{t})$ are i.i.d. across time steps.
Hint: Consider the choice $\mathbf{w}^{\prime}_{t}=D^{\prime}_{t}\begin{bmatrix}\mathbf{w}_{t}\\ \mathbf{v}_{t}\end{bmatrix}$ for some $D^{\prime}_{t}$. Also consider using a hidden state $\mathbf{z}_{t}$ with a larger number of coordinates than $\mathbf{x}_{t}$.
8.2. Prove Lemma 8.3.
8.3. For any $\gamma$-stable partially observed system, prove that any DRC policy in $\Pi_{H,\gamma}^{DRC}$ is $\gamma^{2}$-stabilizing.
8.4. Construct a partially observed linear dynamical system subject to Gaussian i.i.d. perturbations with quadratic costs such that simultaneously it is true that (a) there exists a GOAC policy that incurs zero (or at most constant) aggregate cost, (b) yet any OAC policy incurs $\Omega(T)$ aggregate cost when run for $T$ time steps.
Hint: Consider a 3-dimensional hidden state so that $u_{t}$ can be chosen as a function of $y_{t-1}$ to ensure zero cost at each time step.
Chapter 9 Online Nonstochastic Control with Partial Observation
In this chapter we explore the more general setting of nonstochastic control that allows for partial observation. Partial observation introduces significant challenges and requires revisiting the controllability and stabilizability definitions we have discussed in previous chapters. To allow for more streamlined and simple exposition we make a significant simplification: we assume that the underlying linear dynamical system is stable. The generalization of the results presented in this chapter to stabilizable systems is surveyed in the bibliographic section at the end of this chapter.
Hence, we consider time varying linear dynamical systems with partial observation governed by the equations  [^1] Recall that this is a slightly simplified form, but without loss of generality, as per Exercise 8.5.
$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}$
$\mathbf{y}_{t}$ $=C_{t}\mathbf{x}_{t}.$
The definition of policy regret 1.3 naturally extends to partial observation as below. The setting is similar: the controller iteratively chooses a control $\mathbf{u}_{t}$, and then observes the next observation of the system $\mathbf{y}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$, according to an adversarially chosen loss function. The states $\mathbf{x}_{t}$ is hidden from the controller.
Analogous to policy regret definition 1.3 aa Let $\Pi=\{\pi:\mathbf{y}_{1:t}\mapsto\mathbf{u}_{t}\}$ be a class of policies. The regret of the controller w.r.t. $\Pi$ is defined as follows.
Definition 9.1.
Let $\{A_{t},B_{t},C_{t}\}$ be a time varying, partially observable, linear dynamical system. The regret of an online control algorithm over $T$ iterations w.r.t. class of policies $\Pi$ is given by:
$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\max_{\xi_{1:T},\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|},\|\xi_{t}\|\leq 1}$ $\left(\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t}^{\pi},\mathbf{u}_{t}^{\pi}))\right),$
where $\mathbf{u}_{t}={\mathcal{A}}(\mathbf{y}_{t})$ are the controls generated by ${\mathcal{A}}$, and $\mathbf{y}_{t}^{\pi},\mathbf{u}_{t}^{\pi}$ are the counterfactual state sequence and controls under the policy $\pi$, i.e.
$\mathbf{u}_{t}^{\pi}$ $=\pi(\mathbf{y}_{1:t}^{\pi})$
$\mathbf{x}_{t+1}^{\pi}$ $=A_{t}\mathbf{x}_{t}^{\pi}+B_{t}\mathbf{u}_{t}^{\pi}+\mathbf{w}_{t}$
$\mathbf{y}_{t}^{\pi}$ $=C_{t}\mathbf{x}_{t}^{\pi}.$
Henceforth, if $T$ and $\Pi,{\mathcal{A}}$ are clear from the context, we just refer to $\mathrm{{regret}}$ without quantifiers.
9.1 Disturbance Response Controllers
Recall from section 8 the policy class of disturbance response controllers. This is the natural analogue of disturbance action controllers from the full information setting, adapted to use a different signal called “Nature’s $\mathbf{y}$’s”.
Recall the simplification that we make in this chapter: the linear dynamical system is stable. This allows us to ignore the stabilizing controller component of disturbance action control and greatly simplifies the exposition of the main ideas. We thus can define in succinct notation,
Definition 9.2 (Disturbance Response Controller).
A disturbance response policy for stable systems $\pi_{M_{1:h}}$ is parameterized by the matrices $M_{1:h}=[{M}_{1},\ldots,{M}_{h}]$. It outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
$\mathbf{u}_{t}^{\pi}=\sum_{i=1}^{h}{M}_{i}\mathbf{y}^{\mathbf{nat}}_{t-i},$
where $\mathbf{y}^{\mathbf{nat}}_{t}$ is given by the system response to the zero control at time $t$, i.e.
$\mathbf{x}^{\mathbf{nat}}_{t+1}$ $=A_{t}\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}=\sum_{i=0}^{t}\left[\prod_{j=0}^{i-1}A_{t-j}\right]\mathbf{w}_{t-i}=\Phi_{t}\mathbf{w}_{1:t}.$
$\mathbf{y}^{\mathbf{nat}}_{t+1}$ $=C_{t+1}\mathbf{x}^{\mathbf{nat}}_{t+1}=C_{t+1}\Phi_{t}\mathbf{w}_{1:t}.$
Here $\Phi_{t}$ is the linear operator such that $\Phi_{t}(t-i)=\prod_{j=0}^{i-1}A_{t-j}$.
If the perturbations are all zero, seemingly DRC policies produce a zero control. Therefore we could add a constant offset to the control, that can be incorporated as in Definition 6.7.
The main advantage of this policy class, as opposed to those considered in optimal control such as linear dynamical controllers, is that they yield convex parametrization. Thus, this policy class allows for provably efficient optimization and algorithms.
We conclude with the following important lemma on an expression that is important to compute nature’s y’s. The proof of this lemma follows directly from the definition and is left as exercise 9.4.
Lemma 9.3.
The sequence $\{\mathbf{y}^{\mathbf{nat}}_{t}\}$ can be iteratively computed using only the sequence of observations as
$\mathbf{y}^{\mathbf{nat}}_{t}=\mathbf{y}_{t}-C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}A_{t-j}\right]B_{t-i}\mathbf{u}_{t-i}$
Alternatively, the above computation can be carried out recursively as stated below starting with $\mathbf{z}_{0}=\mathbf{0}$.
$\mathbf{z}_{t+1}$ $=A_{t}\mathbf{z}_{t}+B_{t}\mathbf{u}_{t},$
$\mathbf{y}^{\mathbf{nat}}_{t}$ $=\mathbf{y}_{t}-C_{t}\mathbf{z}_{t}.$
9.2 The Gradient Response Controller
Partial observation requires modification to the Gradient Perturbation Controller we have studied in the previous chapter. The Gradient Response Controller algorithm is described in figure (4), for general changing linear dynamical systems with partial observation. Similar to the GPC, the main idea is to learn the parametrization of a disturbance response control policy using known online convex optimization algorithms, such as online gradient descent. We henceforth prove that this parametrization is essentially convex, allowing us to prove the main regret bound.
In algorithm (4) we denote $\mathbf{y}_{t}({M}_{1:h})$ as the observation arising by playing DRC policy ${M}_{1:h}$ from the beginning of time, and same for $\mathbf{u}_{t}({M}_{1:h})$. Although it may be impossible to determine $\mathbf{w}_{t}$, especially if $C_{t}$ has fewer rows than columns, from the observations $\mathbf{y}_{t}$, $\mathbf{y}_{t}({M}_{1:h})$ can be determined from $\mathbf{y}^{\mathbf{nat}}_{t}$ alone by the application of Lemma 9.3, as
$\mathbf{y}_{t}({M}_{1:h})=\mathbf{y}^{\mathbf{nat}}_{t}+C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}A_{t-j}\right]B_{t-i}\mathbf{u}_{t-i}({M}_{1:h}).$
1:  Input: sequence $\{A_{t},B_{t},C_{t}\}$ , $h$, $\eta$, initialization $M_{0:h}^{1}$.
2:  for $t$ = $1\ldots T$ do
3:     Use Control $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=0}^{h}{M}_{i}^{t}\mathbf{y}^{\mathbf{nat}}_{t-i}$
4:     Observe $\mathbf{y}_{t+1}$, compute for $t+1$:
$\mathbf{y}^{\mathbf{nat}}_{t}=\mathbf{y}_{t}-C_{t}\sum_{i=0}^{t}\left[\prod_{j=0}^{i}A_{t-j}\right]B_{t-i}\mathbf{u}_{t-i}.$
5:     Construct loss $\ell_{t}({M}_{0:h})=c_{t}(\mathbf{y}_{t}({M}_{0:h}),\mathbf{u}_{t}({M}_{0:h}))$Update $M_{0:h}^{t+1}\leftarrow{M}_{0:h}^{t}-\eta\nabla\ell_{t}({M}_{0:h}^{t})$
6:  end for
Algorithm 4 Gradient Response Controller(GRC)
The GRC algorithm comes with a finite time performance guarantee: namely it guarantees vanishing worst case regret vs. the best disturbance response control policy in hindsight. As per the relationship between disturbance response control and other policies studied in section 8.3.1, this implies vanishing regret vs. linear dynamical controllers. The formal statement is given as follows.
Theorem 9.4.
Assuming that
  • a
    The costs $c_{t}$ are convex, bounded and have bounded gradients w.r.t. both arguments $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$.
  • b
    The matrices $\{{A}_{t},{B}_{t},C_{t}\}$ have bounded $\ell_{2}$ norm.
  • c
    The LDS $\{{A}_{t},{B}_{t},C_{t}\}$ is $\gamma$-stable.
Then the GRC algorithm (4) ensures that
$\max_{\xi_{1:T},\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|},\|\xi_{t}\|\leq 1}\left(\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DRC}}\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right)\leq\mathcal{O}(\sqrt{T}).$
Furthermore, the time complexity of each loop of the algorithm is polynomial in the number of system parameters and logarithmic in T.
The proof is largely analogous to the fully observed case. We present a sketch below.
Proof sketch.
As per Theorem A.1 in Appendix A, the online gradient descent algorithm for convex loss functions ensures that the regret is $\mathcal{O}(\sqrt{T})$. However, to directly apply the OGD algorithm to our setting, we need to make sure that the following 3 conditions hold true:
  • The loss function should be a convex function in the variables ${M}_{0:h}$.
  • The loss function is time dependent because the loss depends on the current observation/state, which further depends on the previous states. The state at each time step depends on the current variable estimates. Hence, $\mathbf{y}_{t}$ can be shown to be a function of $\left\{{M}^{i}_{0:h}\right\}_{i=1}^{t}$, where ${M}^{i}_{0:h}$ denotes the variables at step $i$. Thus, we need to make sure that optimizing the loss function is similar to optimizing a function which doesn’t have such a dependence on time.
First, we show that the loss function is convex w.r.t. to the variables ${M}_{0:h}$. This follows since the states and the controls are linear transformations of the variables.
Lemma 9.5.
The loss functions $\ell_{t}$ are convex in the variables ${M}_{0:h}$.
Proof.
Using $\mathbf{x}_{t}$ to denote $\mathbf{x}_{t}(M_{0:h})$ (and similarly for $\mathbf{y}_{t},\mathbf{u}_{t}$), the loss function $\ell_{t}$ is given by
$\ell_{t}({M}_{0:h})=c_{t}(\mathbf{x}_{t},\mathbf{y}_{t}).$
Since the cost $c_{t}$ is a convex function w.r.t. its arguments, we simply need to show that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{0:h}$. Using Lemma 9.3, the action-observation sequence is given by
$\mathbf{u}_{t}$ $=\sum_{i=0}^{h}M_{i}\mathbf{y}^{\mathbf{nat}}_{t-i},$
$\mathbf{y}_{t}$ $=\mathbf{y}^{\mathbf{nat}}_{t}+C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}{A}_{t-j}\right]B_{t-i}\mathbf{u}_{t-i}$
$=\mathbf{y}^{\mathbf{nat}}_{t}+C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}{A}_{t-j}\right]B_{t-i}\sum_{j=0}^{h}M_{j}\mathbf{y}^{\mathbf{nat}}_{t-i-j}$
which is a linear function of the variables. Thus, we have shown that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ are linear transformations in ${M}_{0:h}$ and hence, the loss function is convex in ${M}_{0:h}$. ∎
Now, by the regret bound implied by OGD, we have
$\sum_{t=1}^{T}l_{t}(M_{0:h}^{t})-\min_{M_{0:h}}\sum_{t=1}^{T}l_{t}(M_{0:h})\leq\mathcal{O}(\sqrt{T}).$ (9.1)
Finally, the actual loss $c_{t}$ at time $t$ is not calculated on $\mathbf{y}_{t}(M_{0:h})$, but rather on the true observation $\mathbf{y}_{t}$, which in turn depends on different parameters $M_{0:h}^{i}$ for various historical times $i<t$. Nevertheless, $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$ is well approximated by $l_{t}(M^{t}_{0:h})$.
Lemma 9.6.
For all $t$, $|l_{t}(M^{t}_{0:h})-c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})|\leq\mathcal{O}(1/\sqrt{T})$.
Proof.
First, note that to achieve $\mathcal{O}(\sqrt{T})$ regret OGD uses a step size of $\eta=1/\sqrt{T}$. Since the gradient of the loss functions are bounded, we have
$\|M_{0:h}^{t}-M_{0:h}^{t-i}\|\leq\sum_{s=t-i+1}^{t}\|M_{0:h}^{s}-M_{0:h}^{s-1}\|\leq\mathcal{O}\left(\frac{i}{\sqrt{T}}\right).$
Now, we use the fact that the iterates of OGD $M_{1:h}^{t}$ move slowly to establish that $\mathbf{y}_{t}$ and $\mathbf{y}_{t}(M^{t}_{0:h})$ are close in value. Before that, by definition, we have $\mathbf{u}_{t}(M^{t}_{0:h})=\mathbf{u}_{t}$ since the sequence $\mathbf{y}^{\mathbf{nat}}_{t}$ is determined independently of the choice of the policy being executed. Due to the $\gamma$-stability of the underlying system, we have for any $i\leq j$, $\|C_{j}\prod_{t=i+1}^{j-1}{A}_{t}B_{i}\|\leq\kappa(1-\delta)^{j-i}$ for some $\kappa>0,\gamma<1$. Using Lemma 9.3, we have for some $C>0$ that
$\|\mathbf{y}_{t}(M^{t}_{0:h})-\mathbf{y}_{t}\|$ $=\left\|C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}A_{t-j}\right]B_{t-i}(\mathbf{u}_{t-i}(M^{t}_{0:h})-\mathbf{u}_{t-i})\right\|$
$\leq\sum_{i=1}^{t}\left\|C_{t}\left[\prod_{j=1}^{i-1}A_{t-j}\right]B_{t-i}\right\|\|(\mathbf{u}_{t-i}(M^{t}_{0:h})-\mathbf{u}_{t-i})\|$
$\leq C\kappa\sum_{i=1}^{t}(1-\delta)^{i}\frac{i}{\sqrt{T}}\max_{j}\|\mathbf{y}^{\mathbf{nat}}_{j}\|$
$\leq\frac{C\kappa}{\delta^{2}\sqrt{T}}\max_{j}\|\mathbf{y}^{\mathbf{nat}}_{j}\|\leq\frac{C\kappa^{2}}{\delta^{3}\sqrt{T}}$
Noting that $c_{t}$ has bounded derivatives concludes the proof of the lemma. ∎
Now, we combine the regret guarantee in equation 9.1 and the immediately preceding lemma as
$\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DRC}}\sum_{t=1}^{T}c_{t}(\hat{\mathbf{y}}_{t},\pi(\hat{\mathbf{y}}_{0:t}))$
$\leq\sum_{t=1}^{T}l_{t}(M_{0:h}^{t})-\min_{\pi\in\Pi^{DRC}}l_{t}(M_{0:h}^{\pi})+T\times\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)=\mathcal{O}(\sqrt{T}),$
implying the stated claim.
9.3 Bibliographic Remarks
Online nonstochastic control with partial observability was studied in [SSH20], who defined the Gradient Response Controller as well as the policy class of Disturbance Response Controllers.
The setting was further extended to changing linear dynamics in [MGSH21], where the expressive power of different policy classes was also studied.
The analogue of online nonstochastic control in the stochastic noise setting is called the Linear Quadraic Gaussian problem, and it is further limited by assuming quadratic noise. This classical problem in optimal control was proposed in Kalman’s work [Kal60]. Regret bounds for controlling unknown partially observed systems were studied in [LAHA20b, LAHA20c, MTR19]. [LAHA20a], which uses DRC policies even in context of stochastic control, obtains rate-optimal logarithmic bounds for the LQG problem.
9.4 Exercises
9.1. Prove Lemma 9.3.
9.2. Consider a partially observed linear dynamical system that is $\gamma$-stable in the absence of control inputs. Prove that when a DRC policy $\pi\in\Pi^{DRC}_{H,\gamma}$ is executed on such a dynamical system, the resulting system is $\gamma^{2}$-stable.
9.3. Prove that the optimal controller that minimizes the aggregate cost for a partially observed linear dynamical system assuming the perturbations are stochastic is a linear dynamical controller.
9.4. Prove Lemma 8.5.
9.5. Prove Lemma 8.7.
Chapter 10 Online Nonstochastic System Identification
A fundamental difficulty in applied control is that the dynamics governing the system behavior may be unknown, or too complex to describe. A prominent example is that of control in medical ventilators described in section 2.1.1. The respiratory system of a patient is significantly more complex than the fluid dynamics equations that are commonly used to describe it.
The subfield in control theory that deals with learning a given dynamical system from experience is called system identification. It is notoriously difficult to learn non-linear dynamical system well enough to be able to control them, much less so in high dimensions. This is further complicated in our treatment of control: discrete time representation, finite time performance guarantees, and adversarial perturbations as well as noise.
It is for this reason that we focus on the basic setting of linear time invariant dynamical systems. Thus, for the remainder of this chapter we consider dynamics of the form
$\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}.$
The question we ask in this chapter is: can we attain sublinear policy regret vs. the class of disturbance action policies, when the only information available to the learner are the sequential state observations $\mathbf{x}_{t}$? This may seem very challenging, since the perturbations are adversarially chosen.
Of course, for the problem to be feasible at all the system has to at least be stabilizable. We start with an even stronger assumption, suppose that the linear dynamical system is stable, is it learnable within the framework of online nonstochastic control? Perhaps surprisingly, the answer is positive in a strong sense.
10.1 Nonstochastic System Identification
As a first step, we consider the question of whether the system matrices $A,B$ can be recovered from observations over the state, without regard to the actual cost of control.
The method below for achieving this learns the system matrices by using a random control. Specifically, at each step the control $\mathbf{u}_{t}$ will be randomly selected from a Rademacher distribution, as per Algorithm 5.
Algorithm 5 System identification via method of moments.
  Input: controllability index $k$, oversampling parameter $T_{0}$.
  for $t=0,\ldots,T_{0}(k+1)$ do
     Execute the control $\mathbf{u}_{t}=\eta_{t}$ with $\eta_{t}\sim_{i.i.d.}\{\pm 1\}^{d}$.
     Record the observed state $\mathbf{x}_{t}$.
  end for
  Set $\widehat{A^{j}B}=\frac{1}{T_{0}}\sum_{t=0}^{T_{0}-1}\mathbf{x}_{t(k+1)+j+1}\eta_{t(k+1)}^{\top}$, for all $j\in\{0,\dots k\}$.
  Define $\widehat{C_{0}}=(\widehat{B},\dots\widehat{A^{k-1}B})$, $\widehat{C_{1}}=(\widehat{A^{1}B},\dots\widehat{A^{k}B})$, and return $\widehat{A},\widehat{B}$ as
$\widehat{B}=\widehat{B},\quad\widehat{A}=\operatorname*{\arg\min}_{A}\|\widehat{C_{1}}-A\widehat{C_{0}}\|_{F}.$
The main guarantee from this simple system identification procedure is given as follows.
Theorem 10.1.
For $T_{0}=\Omega\left(\frac{d^{2}k^{2}\kappa^{2}\gamma^{6}}{\varepsilon^{2}}\log\frac{1}{\Delta}\right)$, assuming that
  • a
    the system $(A,B)$ is $\gamma$-stable,
  • b
    the system is $(k,\kappa)$-strongly controllable system,
  • c
    the perturbations $\mathbf{w}_{t}$ are chosen obliviously (i.e. do not depend on $\mathbf{u}_{t}$),
Algorithm 5 outputs a pair $(\widehat{A},\widehat{B})$ that satisfies, with probability $1-\Delta$,
$\|\widehat{A}-A\|\leq\varepsilon,\|\widehat{B}-B\|\leq\varepsilon.$
The proof of this theorem hinges upon the key observation that
$\mathop{\mbox{\bf E}}[\mathbf{x}_{t+k}\mathbf{u}_{t}^{\top}]=\mathop{\mbox{\bf E}}\left[\sum_{i=0:k}A^{i}\left(B\mathbf{u}_{t+k-i}+\mathbf{w}_{t}\right)\mathbf{u}_{t}^{\top}\right]=A^{k}B,$
and the crucial fact that the controls are i.i.d. Thus, we can create an unbiased estimator for this matrix, and using many samples get an increasingly accurate estimate.
Proof.
First, we establish that $A^{j}B$ is close to $\widehat{A^{j}B}$ for all $j\in[k]$.
Lemma 10.2.
As long as $T_{0}\geq\widetilde{\Omega}\left(\frac{d^{2}\gamma^{2}}{\varepsilon^{2}}\log\frac{1}{\Delta}\right)$, with probability $1-\Delta$, we have for all $j\in[k]$ that
$\|\widehat{A^{j}B}-A^{j}B\|\leq\varepsilon.$
Proof.
First note that for any $t$, $\mathbf{x}_{t}$ is bounded as
$\|\mathbf{x}_{t}\|=\left\|\sum_{i>1}A^{i-1}\left(\mathbf{w}_{t-i}+B\eta_{t-i}\right)\right\|\leq 2\sqrt{d}\gamma.$
Therefore, defining $N_{j,t}=\mathbf{x}_{t(k+1)+j+1}\eta_{t(k+1)}^{\top}$ for any $j<k$ and $t<T_{0}$, we have that $\|N_{j,t}\|\leq 2d\gamma$. Note that $\{N_{j,t}\}_{t\leq T_{0}}$ is not a sequence of independent random variables, since the same instance of $\eta_{t}$ might occur in multiple of these terms. To remedy this, we claim that the sequence $\tilde{N}_{j,t}=N_{j,t}-A^{j}B$ forms a martingale difference sequence with respect to the filtration $\Sigma_{t}$, defined as $\Sigma_{t}=\{\eta_{i}:i<(k+1)(t+1)\}$. Since $N_{j,t}$ only involves the first $t(k+1)+k$ random control inputs, ${N}_{j,t}$ is $\Sigma_{t}$-measurable. Now, for any $t$, since $\eta_{t}$ is zero-mean and chosen independently of $\{w_{j}\},\{\eta_{j}\}_{j\neq t}$ at each time step, we have
$\mathbb{E}\left[N_{j,t}|\Sigma_{t-1}\right]$ $=\mathbb{E}\left[\left.\sum_{i>1}A^{i-1}\left(\mathbf{w}_{t(k+1)+j+1-i}+B\eta_{t(k+1)+j+1-i}\right)\eta_{t(k+1)}^{\top}\right|\Sigma_{t-1}\right]$
$=\mathbb{E}\left[\left.\sum_{i>1}A^{i-1}B\eta_{t(k+1)+j+1-i}\eta_{t(k+1)}^{\top}\mathbb{I}_{i=j+1}\right|\Sigma_{t-1}\right]=A^{j}B,$
where the last equality follows from the fact that $\mathbb{E}[\eta_{t}\eta_{t}^{\top}]=I$.
Theorem 10.3 (Matrix Azuma).
Consider a sequence of random symmetric matrices $\{X_{k}\}\in{\mathbb{R}}^{d\times d}$ adapted to a filtration sequence $\{\Sigma_{k}\}$ such that
$\mathbb{E}[X_{k}|\Sigma_{k-1}]=0\quad\text{ and }\quad X_{k}^{2}\preceq\sigma^{2}I\text{ holds almost surely},$
then, for all $\varepsilon>0$, we have
$P\left(\lambda_{\max}\left(\frac{1}{K}\sum_{k=1}^{K}X_{k}\right)\geq\varepsilon\right)\leq de^{-{K\varepsilon^{2}}/{8\sigma^{2}}}.$
Applying the Matrix Azuma inequality (see the bibliography for a reference) on the symmetric dilation, i.e.
$\begin{bmatrix}0&\tilde{N}_{j,t}^{\top}\\ \tilde{N}_{j,t}&0\end{bmatrix},$
we have for all $j\in[k]$ with probability $1-\Delta$ that
$\|\widehat{A^{j}B}-A^{j}B\|\leq{8d\gamma}\sqrt{\frac{1}{T_{0}}\log\frac{n+d}{\Delta}}.$
A union bound over $j\in[k]$ concludes the claim. ∎
At this point, define $C_{0}=(B,AB,\dots,A^{k-1}B)$ and $C_{1}=(AB,A^{2}B,\dots A^{k}B)$ to observe that with probability $1-\Delta$, we have that
$\|\widehat{C_{0}}-C_{0}\|,\|\widehat{C_{1}}-C_{1}\|\leq\sqrt{k}\varepsilon.$
By this display, we already have $\|B-\widehat{B}\|\leq\varepsilon$. We will make use of this observation to establish the proximity of $\widehat{A}$ and $A$, whose proof is left as Exercise 10.4.
Lemma 10.4.
Let $X,X+\Delta X$ be the solutions to the linear system of equations $XA=B$ and $(X+\Delta X)(A+\Delta A)=B+\Delta B$. Then, as long as $\|\Delta A\|\leq\sigma_{\min}(A)$, we have
$\|\Delta X\|\leq\frac{\|\Delta B\|+\|\Delta A\|\|X\|}{\sigma_{\min}(A)-\|\Delta A\|}$
Note that $\widehat{A}$ does not necessarily satisfy $\widehat{A}\widehat{C_{0}}=\widehat{C_{1}}$. Hence, we will invoke the above stated lemma on the normal equations, i.e. $AC_{0}C_{0}^{\top}=C_{1}C_{0}^{\top}$ and $\widehat{A}\widehat{C_{0}}\widehat{C_{0}}^{\top}=\widehat{C_{1}}\widehat{C_{0}}^{\top}$. First, observe that for any $\varepsilon<1$
$\left\|C_{0}C_{0}^{\top}-\widehat{C_{0}}\widehat{C_{0}}^{\top}\right\|$ $=\left\|C_{0}C_{0}^{\top}-C_{0}\widehat{C_{0}}^{\top}+C_{0}\widehat{C_{0}}^{\top}-\widehat{C_{0}}\widehat{C_{0}}^{\top}\right\|$
$\leq\|C_{0}\|\|C_{0}-\widehat{C_{0}}\|+\|\widehat{C_{0}}\|\|C_{0}-\widehat{C_{0}}\|$
$\leq 2\|C_{0}\|\sqrt{k}\varepsilon+k\varepsilon^{2}\leq 3\gamma k\varepsilon,$
where the last line uses that $\|C_{0}\|\leq\gamma$ following the stability of $(A,B)$. A similar display implies the same upper bound on $\|C_{1}C_{1}^{\top}-\widehat{C_{1}}\widehat{C_{1}}^{\top}\|$. Therefore, using $\|A\|\leq\gamma$ and the perturbation result for linear systems, we have
$\|A-\widehat{A}\|\leq\frac{12\gamma^{2}k}{\sigma_{\min}(A)}\varepsilon,$
as long as $3k\kappa\varepsilon/\delta<\sigma_{\min}(C_{0})$. Finally, note that $(k,\kappa)$-strong controllability implies $\sigma_{\min}(C_{0})\geq 1/\kappa$. ∎
10.2 From Indentification to Nonstochastic Control
Having recovered approximate estimates $(\widehat{A},\widehat{B})$ of the system matrices, a natural question arises: how well do GPC and GRC algorithms perform on the original system $(A,B)$ when provided with only approximate estimates of the same? We briefly state a result to this extent, whose proof we leave to the explorations of an interested reader.
Lemma 10.5.
Given estimates of the system matrices $(\widehat{A},\widehat{B})$ such that
$\|\widehat{A}-A\|,\|\widehat{B}-B\|\leq\varepsilon,$
the GPC algorithm when run on the original system $(A,B)$ guarantees a regret bound of
$\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DAC}}\sum_{t=1}^{T}c_{t}(\hat{\mathbf{x}}_{t},\pi(\hat{\mathbf{x}}_{t}))\right)\leq\mathcal{O}(\sqrt{T}+\varepsilon T).$
This result translates to a regret bound of $\tilde{O}(T^{2/3})$ for nonstochastic control of unknown, yet stable, linear systems, by first running the nonstochastic system identification procedure, and subsequently using the estimates thus obtained to run GPC. (Details left as Exercise 10.4
A similar statement holds for the GRC control algorithm, which in addition does not require strong controllability, since it can operate with estimates of the Markov operators alone (using Lemma 9.3), instead of estimates of the system matrices.
10.3 Bibliographic Remarks
Controlling linear dynamical systems given unknown or uncertain system parameters has long been studied in adaptive control; see the text of [SBB90] for example. However, the results from adaptive control typically concern themselves with stability, and even when discussing implications for optimality present asymptotic bounds.
In contrast, much of the work at the intersection of machine learning and control is focussed on precise non-asymptotic bounds on sample complexity or regret. [Fie97] which studies the discounted LQR problem under stochastic noise is the earliest example that furnishes a precise sample complexity bound. Similarly, the early work of [AYS11] uses an optimism-based approach to achieve an optimal regret in terms of the dependence on horizon. This regret bound was later improved for sparse systems in [IJR12]. [DMM+18, DMM+20] established regret and sample complexity results for the general stochastic LQR formulation with a polynomial dependence on natural system parameters. [FGKM18] studies the efficacy of policy-gradient-based approaches for synthesizing controllers for the LQR problem. The work of [MTR19] establishes that simple explore-then-commit strategies achieve the optimal dependence of regret on horizon, with [SF20] proving that such approaches are optimal in terms of their dependence on dimension of the system too. [CKM19] also gave a first optimal-in-horizon polynomial regret result simultaneously with the former work albeit using an optimism-based approach. The result of [PH20] gives a rate-optimal regret bound for unknown systems with convex costs.
In terms of nonstochastic control, the method-of-moments system identification procedure and the corresponding analysis presented here appeared in [HKS20], which also gave the first end-to-end regret result for nonstochastic control starting with unknown system matrices. Even earlier, [SBR19] presented a regression procedure for identifying the Markov operators under nonstochastic noise. Both of these approaches have roots in the classical Ho-Kalman [HK66] system identification procedure, for which the first nonasymptotic sample complexity, although for stochastic noise, were furnished in [OO19]. For strongly convex costs, logarithmic regret was established in [Sim20]. Nonstochastic control of unknown, unstable systems without any access to prior information, also called blackbox control, was addressed in [CH21].
The Matrix Azuma inequality used in this chapter appears in [Tro12].
10.4 Exercises
10.1. Fix a value of $T$. Consider running the nonstochastic system identification procedure for $T_{0}=\tilde{\Theta}(T^{2/3})$ steps to obtain system matrix estimates $\widehat{A},\widehat{B}$, and subsequently using these estimates as inputs to the GPC algorithm when executed over the system $(A,B)$. Given Lemma 10.5, conclude that the resultant control procedure produces a regret of at most $\tilde{O}(T^{2/3})$ without any prior knowledge of system matrices, as long as the underlying system is stable.
10.2. Design an analogous control procedure as that for Exercise 10.4 with an identical regret upper bound (upto constants), when the horizon $T$ is not known in advance.
10.3. Prove Lemma 10.4.
Part III Learning and Filtering
Chapter 11 Learning in Unknown LDS
In this chapter we consider an easier task than control, namely that of learning in a dynamical system. The task of the learner is to predict the next observation from past observations alone, while being unable to influence the underlying dynamical system.
We show in the next few sections how the methodology of convex relaxations extends to prediction in dynamical systems. This chapter uses similar a similar parameterization as DACs in context of prediction, while leaving more intricate parameter-efficient convex relaxations to the next chapters.
11.1 Learning in dynamical systems
There are numerous different interpretations of “learning” in dynamical systems. For this chapter the notion we are concerned with is that of prediction. More precisely, consider a sequence of observable inputs to a dynamical system as in Figure 11.1. Our goal is to correctly predict the output sequence to minimize a given loss metric.
Figure 11.1: In learning we passively observe the input sequence and attempt to predict output.
For the rest of this chapter, we consider a partially observable linear dynamical system that is governed by the following equations.
$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}$
$\mathbf{y}_{t}$ $=C\mathbf{x}_{t}+\mathbf{v}_{t}.$
Here, as we have denoted throughout the text, $\mathbf{x}$ is the hidden state, $\mathbf{u}$ is the observed control input, $\mathbf{w},\mathbf{v}$ are perturbations/noises, and $\mathbf{y}$ is an observed linear transformation of the state. Throughout this chapter we assume that the system matrices $A,B,C$ are unknown to the learner.
11.2 Online learning of dynamical systems
In the problem of online learning in a linear dynamical system, sequentially for $t=1,...,T$, the online learning algorithm:
  • observes input $\mathbf{u}_{t-1}\in{\mathbb{R}}^{d_{u}}$
  • predicts output $\hat{\mathbf{y}}_{t}\in{\mathbb{R}}^{d_{y}}$
  • observes true system output $\mathbf{y}_{t}$
  • suffers loss according to a pre-specified loss function, $\ell(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})$, e.g.
    $\ell(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})=\|\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\|^{2}.$
The goal of the online learner is to minimize total loss over all game iterations. If the observations ${\mathbf{y}}_{t}$ are arbitrary, this is an ill-defined goal, and rather than minimizing loss, we consider a comparative analogue.
More precisely, we consider a family of prediction rules from available information to prediction, and consider the metric of regret vs. the best of these rules in hindsight. An example of a rule could be a fixed linear map, i.e. $\pi(\mathbf{u}_{1:t-1},\mathbf{y}_{1:t-1})=K\mathbf{u}_{t-1}$, where $K$ can potentially depend on the system matrices $A,B,C$.
Formally, let $\Pi$ be a family of mappings from $\{\mathbf{u}_{\tau},\mathbf{y}_{\tau}\}\mapsto\hat{\mathbf{y}}_{t}$, i.e.
$\pi\in\Pi\ ,\ \pi(\mathbf{u}_{1:t-1},\mathbf{y}_{1:t-1})=\hat{\mathbf{y}}_{t}^{\pi}\in{\mathbb{R}}^{d_{y}}.$
Let ${\mathcal{A}}$ be a prediction algorithm that iteratively produces predictions $\hat{\mathbf{y}}_{t}$. Then its regret w.r.t. $\Pi$ is defined as
$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\Pi}\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}^{\pi}_{t}\|^{2}.$
We proceed to define common prediction classes in the context of linear dynamical systems.
11.3 Classes of Prediction Rules
We now consider what would be a reasonable prediction rule for the output of a linear dynamical system? The case of full observation is particularly simple, and we deal with it first. We then consider prediction rules for partially observed linear dynamical systems.
However, before considering the prediction rules themselves, we discuss how to compare classes of such rules. This is analogous to the definitions and discussion in chapter 6 where we discuss comparison between policy classes for control. However, the definitions and conclusions for prediction are simpler.
Definition 11.1.
We say that a class of predictors $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ if the following holds. For every sequence of $T$ inputs and outputs $\{\mathbf{u}_{1:T},\mathbf{y}_{1:T}\}$, and every $\pi\in\Pi_{1}$, there exists $\pi_{2}\in\Pi_{2}$ such that,
$\sum_{t=1}^{T}\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|\leq T\varepsilon,$
where $\mathbf{y}_{t}^{\pi_{1}}=\pi_{1}(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\ ,\ \mathbf{y}_{t}^{\pi_{2}}=\pi_{2}(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})$
The significance of this definition is that if a simple predictor class approximates another more complicated class, then it suffices to consider only the first.
The implication of $\varepsilon$-approximation on the regret is given below.
Lemma 11.2.
Let predictor class $\Pi_{1}$ $\varepsilon$-approximate class $\Pi_{2}$. Then for any prediction algorithm ${\mathcal{A}}$ we have
$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{2})\leq\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{1})+\varepsilon T.$
Proof of this lemma is left as an exercise to the reader at the end of this chapter. We now proceed to consider specific prediction rules.
11.3.1 Linear Predictors
Perhaps the simplest prediction rule is linear, in either the observation, or input, or both. More precisely, the class of all predictions of the form
$\hat{\mathbf{y}}_{t+1}=\sum_{i=1}^{h}M_{1,i}\mathbf{y}_{y-i}+\sum_{j=1}^{k}M_{2,j}\mathbf{u}_{y-j}.$
We denote by $\Pi^{LIN}_{h,k,\kappa}$ the class of all linear predictors with $\sum_{i=0}^{h}\|M_{1,i}\|+\sum_{j=1}^{k}\|M_{2,j}\|\leq\kappa$.
Definition 11.3 (Linear Policies).
A linear controller parametrized by matrices $\{M_{1,i},M_{2,j}\}\in{\mathbb{R}}^{h\times d_{y}\times d_{y}+k\times d_{y}\times d_{u}}$ is a linear operator mapping state to control. We denote by $\Pi_{h,k,\kappa}^{LIN}$ the set of all linear controllers bounded as $\sum_{i=0}^{h}\|M_{1,i}\|+\sum_{j=1}^{k}\|M_{2,j}\|\leq\kappa$.
11.3.2 Open Loop Optimal Predictors
An intuitive class of prediction rules that makes use of the special structure of linear dynamical systems is that of open loop linear dynamical predictors, which we now define.
The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $\mathbf{C}=I$ is the identity matrix. In this situation, the optimal prediction can be computed via the system evolution equation $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}$. It can be seen that this prediction rule is optimal for any zero-mean noise or perturbation model. However, this optimality is only true in the case of full observation of state. In the next chapter we consider optimal predictors for zero mean i.i.d noise in the more general partially observed case.
For a partially observed linear dynamical system without noise we can write
$\mathbf{y}_{t}^{\pi}=\pi_{\{A,B,C,\mathbf{x}_{0}\}}(t)$ $=C\mathbf{x}_{t}$
$=CA\mathbf{x}_{t-1}+CB\mathbf{u}_{t-1}$
$\vdots$
$=\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t-1}\mathbf{x}_{0}.$
Thus, the operator $\pi_{\{A,B,C,\mathbf{x}_{0}\}}$ takes a history of states and controls and gives the next system output accurately. We denote by $\prod_{\star}$ the set of all such predictors, that in the case of zero noise give a perfectly accurate prediction.
If the spectral radius of $A$ is larger than one, then the system is unstable and the magnitude of $\mathbf{y}_{t}$ can grow arbitrarily large, as discussed in section 4.2. Thus, for the rest of this section, assume that $A$ is $\gamma$-stable.
With a decaying system, we can limit the memory of the predictor and yet maintain accuracy. Thus it makes sense to consider a more restricted class of predictors that have bounded memory, i.e.
$\pi_{\{A,B,C,h\}}=\sum_{i=1}^{h}CA^{i-1}B\mathbf{u}_{t-i},$
where we define all $\mathbf{u}$ vectors in negative times as zero. Denote by $\prod_{h}^{OLOP}$ the class of all such predictors.
Clearly for a large $h$ this class of predictors well approximates the true observations $\mathbf{y}_{t}$ if the system is activated without any noise. We formalize what this means next.
Lemma 11.4.
Suppose that $\|B\|+\|C\|\leq\gamma$, and $\|\mathbf{u}_{t}\|\leq U$ for all $t\in[T]$. Then the class $\prod_{h}^{OLOP}$ for $h=\gamma\log\frac{T\gamma^{3}U}{\varepsilon}$, $\varepsilon$-approximates the class $\prod_{\star}$.
Proof.
Consider any sequence and any predictor from $\prod_{\star}$, say $\pi_{1}=\pi_{A,B,C,\mathbf{x}_{0}}$. Consider the predictor $\pi_{2}=\pi_{A,B,C,h}\in\prod_{h}$ for the same system matrices. Then we have
$\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|$
$=\|\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t-1}\mathbf{x}_{0}-\sum_{i=1:h}CA^{i-1}B\mathbf{u}_{t-i}\|$
$=\|\sum_{i=h}^{T}CA^{i}B\mathbf{u}_{t-i-1}\|$
$\leq\sum_{i=h}^{T}\kappa(1-\delta)^{i}\|C\|\|B\|\|\mathbf{u}_{t-i-1}\|$
$\leq\frac{\gamma^{2}\kappa U}{\delta}(1-\delta)^{h}\leq\frac{\gamma^{2}\kappa U}{\delta}e^{-\delta h},$
for some $\frac{\kappa}{\delta}\leq\gamma$. Using the value of $H$, and summing up over all $T$ iterations, we have
$\sum_{t}\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|$ $\leq\sum_{t}\frac{\gamma^{2}\kappa U}{\delta}e^{-\delta h}=\varepsilon.$
This lemma explains why the class $\prod^{OLOP}$ is an optimal prediction class for open loop prediction, in which the realized noise (through observations) cannot be used inside of the prediction rule, and in the case of full observation.
We note that the class $\prod^{LIN}$ also $\varepsilon$-approximates the class $\prod^{OLOP}$ for appropriate choice of parameters. This is left as an exercise (Exercise 11.6) to the reader.
11.4 Efficient Learning of Linear Predictors
The previous section gives several natural classes of predictors as well some of the relationships between them. Notably, the class of open loop optimal predictors is optimal for certain scenarios and requires knowledge of only three matrices. However, the parametrization of this class is non-convex in the matrices $A,B,C$, which presents a computational challenge.
This challenge is circumvented by using the fact that the class $\prod^{LIN}$ $\varepsilon$-approximates the classe $\prod^{OLOP}$, and as we show in the next chapter, also optimal closed loop predictors. Further, the natural parametrization of $\prod^{LIN}$ is convex and thus can be learned by online convex optimization.
The following algorithm learns the class of linear predictors in the sense that it attains a sublinear regret bound for any sequence of adversarial cost functions, as well as perturbations and observation noise.
Algorithm 6 Learning $\prod^{LIN}$ by Online Gradient Descent
1:  Input: $M^{1}=M^{1}_{1,1:h},M^{1}_{2,1:k}$, convex constraints set $\mathcal{K}\subseteq{\mathbb{R}}^{h\times d_{y}+k\times d_{u}}$
2:  for $t=1$ to $T$ do
3:     Predict $\hat{\mathbf{y}}_{t}(M^{t})=\sum_{i=1}^{h}M^{t}_{i}\mathbf{u}_{t-i}+\sum_{j=1}^{k}M^{t}_{j}\mathbf{y}_{t-j}$ and observe true $\mathbf{y}_{t}$.
4:     Suffer loss $\ell_{t}(\hat{\mathbf{y}}_{t}(M^{t}),\mathbf{y}_{t})$, and update:
$M^{t+1}=M^{t}-\eta_{t}\nabla\ell_{t}(M^{t})$
$M^{t+1}=\mathop{\Pi}_{\mathcal{K}}(M^{t+1})$
5:  end for
This algorithm is an instance of the online gradient descent algorithm, Theorem A.1 from Appendix A directly gives the following bound for it. We denote by $\mathcal{K}$ the set of all $h\times k$-tuple of matrices $M_{1,1:h},M_{2,1:k}$, such that they are bounded as needed. Then we obtain
Corollary 11.5.
For choice of step sizes $\eta_{t}=O(\frac{1}{\sqrt{t}})$, and ${\mathcal{A}}$ being Algorithm 6, we have that
$\mathrm{{regret}}({\mathcal{A}})=\sum_{t}\ell(\hat{\mathbf{y}}_{t}^{\mathcal{A}},\mathbf{y}_{t})-\min_{\pi\in\prod^{LIN}}\ell(\hat{\mathbf{y}}_{t}^{\pi},\mathbf{y}_{t})=O(\sqrt{T}).$
11.5 Bibliographic Remarks
The task of predicting future observations in an online environment is often called “filtering”. Resting at the crossroads between statistical estimation and dynamical systems, the theory of filtering had its early beginnings in the works of Norbert Wiener [Wie49].
11.6 Exercises
11.1. Prove that $\varepsilon$-approximation of prediction classes implies $\varepsilon T$ regret approximation, as formalized in Lemma 11.2.
11.2. Prove that the class $\prod^{LIN}$ $\varepsilon$-approximates the class $\prod^{OLOP}$ for appropriate choice of parameters.
Chapter 12 Kalman Filtering
The previous chapter introduces the concept of open-loop optimal linear predictors and how to efficiently learn them using convex relaxation and online gradient methods. However, a good predictor should exploit not only its initial knowledge, but also information as it is being observed and learn from it. In an online prediction setting, this amounts to using the online observations to further refine future predictions.
This is the topic of this chapter: how to use feedback from the ongoing prediction task to improve future predictions. Optimal closed-loop predictors, that use feedback from observations, are significantly more powerful and harder to analyze. We explore Kalman filters, prove their optimality, as well as convex relaxations that fit into the paradigm of online convex optimization.
12.1 Observable Systems
In the discussion below, when considering the infinite horizon setting, prediction makes sense if our observation remains bounded. The condition for bounded observations is related to controllability as defined in Section 4.3.
We first state the definition of observability for general dynamical systems below. In absence of noise and control inputs, every dynamical system can be viewed as mapping from an initial state to an infinite sequence of observations. Observability asserts that this mapping is injective.
Definition 12.1.
A partially-observed dynamical system is said to be observable if and only if every sequence of observations produced by the deterministic-equivalent system uniquely determines the initial hidden state.
For linear systems, checking observability can be boiled to a rank condition on the Kalman observability matrix, which is defined as
$K^{\prime}_{k}(A,C)=\begin{bmatrix}C\\ CA^{2}\\ \dots\\ CA^{k-1}\end{bmatrix}.$
We ask the reader to provide a proof of the following result in Exercise 12.5.
Theorem 12.2.
A partially-observed dynamical system specified by system matrcies $(A,B,C)$ is said to be observable if $\mathop{\mbox{\rm rank}}(K^{\prime}_{d_{x}}(A,C))=d_{x}$.
12.2 Closed Loop Optimal Predictors
The previous class of predictors is open-loop, namely it only takes into account the inputs to the system, which can potentially be computed ahead of time, and not the response of the system to the controls and perturbations. A closed-loop prediction rule takes into account the observations, and can potentially react to the perturbations and observation noise.
The most famous example of such a prediction rule is the Kalman filter, which we now define. The defining characteristic of the Kalman filter is that it produces the best linear predictor in the least-squares sense when the perturbations $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ are zero-mean and i.i.d., and when the systems matrices $(A,B,C)$ and the perturbations’ covariances $(\Sigma_{w},\Sigma_{v})$ are known and defined as
$\Sigma_{x}\triangleq\mathbb{E}\left[\mathbf{w}_{t}\mathbf{w}_{t}^{\top}\right],\quad\Sigma_{y}=\mathbb{E}\left[\mathbf{v}_{t}\mathbf{v}_{t}^{\top}\right].$
So far we have focussed on predicting observations, i.e., ${\mathbf{y}}_{t}$’s. On the other hand, the next theorem produces optimal estimators $\hat{\mathbf{x}}_{t}$ of the hidden state $\mathbf{x}_{t}$ given past observations $\mathbf{y}_{1:t-1}$ and controls $\mathbf{u}_{1:t-1}$. This is motivated by the fact that in the stochastic case, the optimal least-squares predictor can be written as $\hat{\mathbf{y}}_{t}=C\hat{\mathbf{x}}_{t}$. In Exercise 12.5, we ask the reader to supply a proof of this statement.
Theorem 12.3.
Let the control sequence $\mathbf{u}_{t}$ be arbitrary affine functions of $\mathbf{y}_{1:t}$. There exists an efficiently computable sequence of linear predictors $\hat{\mathbf{x}}_{t+1}$ defined recursively as
$\hat{\mathbf{x}}_{t+1}=(A-L_{t}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+L_{t}\mathbf{y}_{t},$
such that
$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{LIN}}\mathbb{E}\|\mathbf{x}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
Proof.
Since the estimator is recursively defined, a natural strategy is to proceed by induction. Let us say $\hat{\mathbf{x}}_{t}$ is a such a predictor. Further, let
$\Sigma_{t}\triangleq\mathbb{E}(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}.$
We begin with an observation about least-squares predictors, which we ask the reader to prove in Exercise 12.5.
Lemma 12.4.
Let $X,Y,Z$ be matrix-valued random variables. Let
$A^{*}=\operatorname*{\arg\min}_{A}\mathbb{E}\|X-AY\|^{2},\text{ and }B^{*}=\operatorname*{\arg\min}_{B}\mathbb{E}\|X-BZ\|^{2}.$
If $\mathbb{E}\left[YZ^{\top}\right]=0$, then
$\mathbb{E}\|X-A^{*}Y-B^{*}Z\|^{2}\leq\min_{A,B}\mathbb{E}\|X-AY-BZ\|^{2}.$
In the remaining proof, we use this observation to construct an optimal least-squares estimator $\hat{\mathbf{x}}_{t+1}$ for $\mathbf{x}_{t+1}$ given $\hat{\mathbf{x}}_{t}$. In particular, we know by induction hypothesis that
$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{LIN}}\mathbb{E}\|\mathbf{x}_{t}-L\left[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}\right]\|^{2}.$
But, is $\hat{\mathbf{x}}_{t}$ any good for predicting $\mathbf{x}_{t+1}$? To answer this, we note the following structural characterization of least-squares (Exercise 12.5).
Lemma 12.5.
Let $X,Y$ be possibly correlated random variables. Then
$A^{*}\triangleq\operatorname*{\arg\min}_{A}\mathbb{E}\|X-AY\|^{2}=\mathbb{E}\left[XY^{\top}\right]\left(\mathbb{E}\left[YY^{\top}\right]\right)^{-1}.$
Furthermore, if $\hat{X}=A^{*}Y$, then $\mathbb{E}(X-\hat{X})Y^{\top}=0$.
Since for any $s<t$
$\mathbb{E}\left[(\mathbf{x}_{t+1}-B\mathbf{u}_{t})\mathbf{y}_{s}^{\top}\right]=\mathbb{E}\left[(A\mathbf{x}_{t}+\mathbf{w}_{t})\mathbf{y}_{s}^{\top}\right]=A\mathbb{E}\left[\mathbf{x}_{t}\mathbf{y}_{s}^{\top}\right],$
using the first part of Lemma 12.5, we have that
$\mathbb{E}\|\mathbf{x}_{t+1}-A\hat{\mathbf{x}}_{t}-B\mathbf{u}_{t}\|^{2}\leq\min_{L\in\Pi^{LIN}}\mathbb{E}\|\mathbf{x}_{t+1}-B\mathbf{u}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
But, when predicting $\mathbf{x}_{t+1}$, we have more information. In particular, we have $\mathbf{y}_{t}$ too. Consider the random variables $\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}$. Can we use this covariate alone to predict $\mathbf{x}_{t+1}$? Note that
$\mathbb{E}\left[(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})^{\top}\right]$ $=\Sigma_{v}+C\mathbb{E}\left[(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}\right]C^{\top}$
$=\Sigma_{v}+C\Sigma_{t}C^{\top},$
$\mathbb{E}\left[(\mathbf{x}_{t+1}-B\mathbf{u}_{t})(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})^{\top}\right]=\mathbb{E}\left[(A\mathbf{x}_{t}+\mathbf{w}_{t})(C(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})+\mathbf{v}_{t})^{\top}\right]$
$=A\mathbb{E}\left[(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}\right]C^{\top}+A\underbrace{\mathbb{E}\left[\hat{\mathbf{x}}_{t}(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}\right]}_{=0}C^{\top}$
$=A\Sigma_{t}C^{\top},$
where the equality indicated in braces follows from the second part of Lemma 12.5, by noting that $\hat{\mathbf{x}}_{t}$ is a linear function of $\mathbf{y}_{1:t-1}$.
In the preceding paragraphs, we have constructed two least-squares predictors for $\hat{\mathbf{x}}_{t+1}$. Observe that for any $s<t$
$\mathbb{E}(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})\mathbf{y}_{s}^{\top}=C\mathbb{E}(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})\mathbf{y}_{s}^{\top}=0,$
where the last equality once again follows from Lemma 12.5. This orthogonality allows us to invoke Lemma 12.4 to arrive at
$\hat{\mathbf{x}}_{t+1}=A\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}C^{\top}\left(\Sigma_{y}+C\Sigma_{t}C^{\top}\right)^{-1}}_{\triangleq L_{t}}(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}),$
where the last component follows from Lemma 12.5 and the correlations we just computed.
To establish that such predictors can be computed efficiently, note that
$\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1}=(A-L_{t}C)(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})+\mathbf{w}_{t}+L_{t}\mathbf{v}_{t}.$
Finally, we give a recursive relation on $\Sigma_{t}$ as follows–
$\Sigma_{t+1}$ $\triangleq\mathbb{E}\left[(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})^{\top}\right]$
$=(A-L_{t}C)\Sigma_{t}(A-L_{t}C)^{\top}+\Sigma_{x}+L_{t}\Sigma_{y}L_{t}^{\top}$
$=A\Sigma_{t}A^{\top}+\Sigma_{x}+L_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})L_{t}^{\top}-2A\Sigma_{t}C^{\top}L_{t}^{\top}$
$=A\Sigma_{t}A^{\top}-A\Sigma C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}C\Sigma A^{\top}+\Sigma_{x}.$
In the infinite-horizon setting, the steady-state Kalman filter can be described by the following linear dynamical system
$\hat{\mathbf{x}}_{t+1}$ $=(A-LC)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+L\mathbf{y}_{t},$
$\hat{\mathbf{y}}_{t}$ $=C\hat{\mathbf{x}}_{t},$
where $L=A\Sigma(C\Sigma C^{\top}+\Sigma_{y})^{-1}$ and $\Sigma$ satisfies the Riccati equation
$\Sigma=A\Sigma A^{\top}-A\Sigma C^{\top}(C\Sigma C^{\top}+\Sigma_{y})^{-1}C\Sigma A^{\top}+\Sigma_{x}.$
However, when does a solution to this Riccati equation exist? A sufficient condition for this is the observability of the underlying dynamical system. In particular, any observable partially-observed linear dynamical system admits a solution to the Riccati equation listed above, and hence, has a steady-state Kalman filter. Mirroring the link between controllability and stabilizability, observability implies that $\rho(A-LC)<1$.
For observable linear dynamical systems, the Kalman filter is contained in the class of linear predictors. Exercise 12.5 asks for a proof of the next lemma.
Lemma 12.6.
Consider any partially-observed linear dynamical system $(A,B,C)$ where the Kalman filtering matrix $L$ is such that the system $A-LC$ is $\gamma$-stable. The Kalman filtering policy is $\varepsilon$-contained in the class of linear policies $\Pi^{LIN}_{h,k,\kappa}$, where $h=k=\gamma\log\frac{\gamma T}{\varepsilon}$ and $\kappa=\gamma^{3}$.
12.3 Bayes-optimality under Gaussian Noise
In this subsection, we further restrict $\mathbf{w}_{t},\mathbf{v}_{t}$ to be i.i.d. draws from Gaussian distributions, and proceed to establish that the Kalman filter in such cases enjoys a stronger notion of optimality; it produces the the best estimator (under a least-squares loss) of the state given past observations among the class of all (not just linear) predictors.  [^1] This is also called the Bayes-optimal predictor. See bibliographic materials for more details and literature. . The setting we consider is that of Gaussian perturbations $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ and known systems matrices $(A,B,C)$. In particular, we consider the case when $\mathbf{w}_{t}\sim\mathcal{N}(0,\Sigma_{x})$ and $\mathbf{v}_{t}\sim\mathcal{N}(0,\Sigma_{y})$ are i.i.d. normally distributed draws, and $x_{0}=0$. Given the realizations $\mathbf{y}_{1},\mathbf{y}_{2}\dots,\mathbf{y}_{t}$ under some control input sequence $\mathbf{u}_{1},\mathbf{u}_{2},\dots\mathbf{u}_{t}$, we would like to compute a prediction $\hat{\mathbf{y}}_{t+1}$ that minimizes the mean square error, which is simply the expected observation as per
$\hat{\mathbf{y}}_{t+1}=\mathbb{E}[\mathbf{y}_{t+1}|\mathbf{y}_{1:t},\mathbf{u}_{1:t}].$
Since $\hat{\mathbf{y}}_{t+1}=C\hat{\mathbf{x}}_{t+1}$, we will recursively compute $\hat{\mathbf{x}}_{t+1}$ where
$\hat{\mathbf{x}}_{t+1}=\mathbb{E}[\mathbf{x}_{t+1}|\mathbf{y}_{1:t},\mathbf{u}_{1:t}].$
Theorem 12.7.
The conditional expectation of the state $\hat{\mathbf{x}}_{t+1}$ satisfies
$\hat{\mathbf{x}}_{t+1}=(A-L_{t}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+L_{t}\mathbf{y}_{t},$
for a sequence of matrices $L_{t}\in{\mathbb{R}}^{d_{x}\times d_{y}}$ that can be computed efficiently given $(A,B,C)$, and is independent of the realizations $\mathbf{u}_{1:T},\mathbf{y}_{1:T}$.
Proof.
From the Gaussian nature of perturbations, we observe that
$\begin{bmatrix}\mathbf{x}_{t}|(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\\ \mathbf{y}_{t}|(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\hat{\mathbf{x}}_{t}\\ C\hat{\mathbf{x}}_{t}\end{bmatrix},\begin{bmatrix}\Sigma_{t}&\Sigma_{t}C^{\top}\\ C\Sigma_{t}&C\Sigma_{t}C^{\top}+\Sigma_{y}\end{bmatrix}\right).$
for some $\Sigma_{t}\in{\mathbb{R}}^{d_{x}\times d_{x}}$, for which we provide a recursive formula towards the end of the proof. We make note of the following lemma for Gaussian distributions, whose proof we ask the reader to complete in Exercise 12.5.
Lemma 12.8.
If a random variable pair $(\mathbf{x},\mathbf{y})$ are distributed as
$\begin{bmatrix}\mathbf{x}\\ \mathbf{y}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mu_{x}\\ \mu_{y}\end{bmatrix},\begin{bmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}\end{bmatrix}\right),$
then $\mathbf{x}|\mathbf{y}\sim\mathcal{N}(\mu_{x}+\Sigma_{12}\Sigma_{22}^{-1}(\mathbf{y}-\mu_{y}),\Sigma_{11}-\Sigma_{12}\Sigma_{22}^{-1}\Sigma_{21}).$
Therefore, we have that
$\mathbf{x}_{t}|(\mathbf{y}_{1:t},\mathbf{u}_{1:t-1})$ $\sim\mathcal{N}(\hat{\mathbf{x}}_{t}+\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}),\Sigma^{\prime}_{t}),$
$\Sigma^{\prime}_{t}$ $=\Sigma_{t}-\Sigma_{t}C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}C\Sigma_{t}.$
Now, using this, we arrive at
$\hat{\mathbf{x}}_{t+1}$ $=A\mathbb{E}[\mathbf{x}_{t}|\mathbf{y}_{1:t},\mathbf{u}_{1:t-1}]+B\mathbf{u}_{t}$
$=(A-\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\triangleq L_{t}}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\triangleq L_{t}}\mathbf{y}_{t}.$
It remains to obtain a recursive relation on $\Sigma_{t}$. To this extent, from the state evolution of a linear dynamical system, we have that
$\Sigma_{t+1}=A\Sigma^{\prime}_{t}A^{\top}+\Sigma_{x}=A\Sigma_{t}A^{\top}-A\Sigma_{t}C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}C\Sigma_{t}A^{\top}+\Sigma_{x}.$
12.4 Bibliographic Remarks
In 1960, Rudolph Kalman [Kal60] gave a state-space solution to the problem of filtering that was highly amenable to further extensions. See [AM91] for a historical perspective on its development and aftermaths.
For unknown systems, in the stochastic case, one may estimate the system matrices, as we have discussed before, before applying optimal filtering. Improper learning approaches to this problem were studied in [KMTM19].
12.5 Exercises
12.1. Prove Lemma 12.5.
12.2. Using Lemma 12.5, prove Lemma 12.4.
12.3. Prove Lemma 12.6.
12.4. Prove Lemma 12.8.
12.5. Prove Theorem 12.2.
12.6. Using Lemma 12.5, prove that the optimal least-squares estimator for the observation at time $t$ given past observations $y_{1:t-1}$ and control inputs $u_{1:t-1}$ is given by $\hat{\mathbf{y}}_{t}=C\hat{\mathbf{x}}_{t}$, where $\hat{\mathbf{x}}_{t}$ is as defined in Theorem 12.3.
Chapter 13 Spectral Filtering
In the previous chapter we have showed how to learn an unknown linear dynamical system with perturbations and noise using convex relaxation and online learning. Instead of learning the system matrices directly, we have learned an alternate class of predictors, namely the class of linear predictors over past observations and controls, that approximates the optimal predictor in certain cases.
Despite the clean and simple formulation of this result, there remains an unsatisfying issue with the previous chapter. The number of parameters required to learn the optimal open loop predictor in the class $\prod_{\star}$ depends on the measure of stability of the underlying system, which is potentially large and unknown.
In this section we describe an elegant alternative method for learning a linear dynamical system that removes this dependence on the stability constant of the system. For the sake of exposition, we restrict ourselves only to symmetric system matrices $A$. This is a significant restriction, which can be removed, albeit using a significantly more sophisticated approach which is referred to in the bibliographic section.
13.1 Spectral filtering in One Dimension
In this section we give a simple introduction to the technique of spectral filtering, restricting ourselves to a single dimension. The restriction to a single dimension simplifies the exposition and yet is sufficient to illustrates the main ideas. In the next section we consider the high dimensional case.
The single dimensional systems which we consider next are scalar systems, with $d_{u}=d_{y}=d_{x}=1$. We assume for simplicity that the initial state is zero $\mathbf{x}_{0}=0$. Recall the set of optimal predictors, the class $\Pi_{\star}$, which for scalar systems can written and simplified to,
$\mathbf{y}_{t}$ $=\sum_{i=1}^{t}CA^{i-1}B\mathbf{u}_{t-i}=\beta\sum_{i=1}^{t}\alpha^{i-1}\mathbf{u}_{t-i},$
Where $\beta=CB\in{\mathbb{R}}$. Thus, we can write the difference between two consecutive iterates as
$\mathbf{y}_{t}-\mathbf{y}_{t-1}$ $=\beta(\sum_{i=1}^{t}\alpha^{i-1}\mathbf{u}_{t-i}-\sum_{i=1}^{t-1}\alpha^{i-1}\mathbf{u}_{t-1-i})$
$=\beta(\mathbf{u}_{t-1}+\sum_{i=1:t-1}(\alpha^{i}-\alpha^{i-1})\mathbf{u}_{t-1-i})$
$=\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1}),$ (13.1)
which is a geometric series involving $\alpha$ multiplying the past controls. We used the operator notation to denote the operator
$\mu_{\alpha}(\mathbf{u}_{t-1:1})=\mathbf{u}_{t-1}+\sum_{i=1:t-1}(\alpha-1)\alpha^{i-1}\mathbf{u}_{t-1-i}=\mu_{\alpha}^{\top}\tilde{\mathbf{u}}_{t-1}.$ (13.2)
Above we used the vector notation for
$\mu_{\alpha}=(\alpha-1)\begin{bmatrix}\frac{1}{\alpha-1}&1&\alpha&\alpha^{2}&\ldots&\alpha^{T-1}\end{bmatrix}\in{\mathbb{R}}^{T},$
and we define $\tilde{\mathbf{u}}_{t}$ to be the padded vector $\mathbf{u}_{t-1:1}$:
$\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}.$
The main idea of spectral filtering is to rewrite the vectors $\mu_{\alpha}$ in a more efficient basis, as opposed to the standard basis. The notion of efficiency stems from the definition of $\varepsilon$-approximation we have studied in previous chapters. To define the basis, we consider the following matrix:
[…]
The matrix $Z_{T}$ is symmetric and is independent of $\alpha$. Let $\phi_{1},...,\phi_{k},...\in{\mathbb{R}}^{T}$ be the eigenvectors of $Z_{T}$ ordered by the magnitude of their corresponding eigenvalues. Our basis for representing $\mu_{\alpha}$ is spanned by the eigenvectors of $Z_{T}$, and we proceed to study their properties.
13.1.1 The Magic of Hankel Matrices
The matrix $Z_{T}$ is a special case of a Hankel Matrix, which is a square Hermitian matrix with each ascending skew-diagonal from left to right having the same value. An example of a Hankel matrix is the following,
[…]
This matrix is closely related to the $Z_{T}$ matrix we have introduced before, but not equal to it.
The properties of Hankel matrices have been studied extensively, and the following theorem captures one of their crucial properties.
Theorem 13.1.
Let $\sigma_{k}$ be the $k$-th largest singular value of a Hankel matrix $M\in{\mathbb{R}}^{T\times T}$, then there exists constants $c_{1},c_{2}$ such that :
$\sigma_{k}(M)\leq c_{1}\cdot e^{-\frac{c_{2}k}{\log T}}$
The proof of this theorem is beyond our scope, and references to details are given in the bibliographic material at the end of this chapter.
This theorem shows that the eigenvalues of a Hankel matrix decay very rapidly. A further illustration of this fact is given in Figure 13.1, which plots the eigenvalues of the matrix $Z_{T}$ for $T=30$ on a logarithmic scale.
Figure 13.1: Eigenvalues of Hankel matrices decrease geometrically. These are the eigenvalues of the matrix $Z_{T}$, where $T=30$, plotted on a logarithmic scale.
13.2 Spectral Predictors
The basis of eigenvectors of a Hankel matrix gives rise to an interesting class of predictors. This class of mappings do not directly apply to the history of inputs and/or observations. Instead, we first apply a linear transformation over these vectors, and then apply a linear map. More formally,
Definition 13.2.
Define the class of spectral linear predictors of order $k$ as
$\prod_{h}^{\lambda}=\left\{\pi_{M}\ \ \left|\ \ \pi_{M}(\mathbf{u}_{1:t},\mathbf{y}_{1:t})=\sum_{i=1}^{h}M_{i}\cdot\phi_{i}(\tilde{\mathbf{u}}_{t-i})\right.\right\}.$
Here the operator $\phi_{i}(\mathbf{u}_{t-1})$ takes the inner product of the $i$’th eigenvector of $Z_{T}$ with $\tilde{\mathbf{u}}_{t-1}$, where $\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}$.
Notice that as $h\mapsto T$, this class is equivalent to the class of all linear predictors on the inputs $\prod^{LIN}$. The reader may wonder at this point: Since a linear transformation composed with another linear transform remains linear, what then is the benefit here?
The answer is that linear transformations are equivalent when we have full rank transforms. In our case, however, we consider low rank approximations to the transform, and the approximation properties of different subspaces can differ drastically! This is the case in our setting: the spectral basis arising from the matrix $Z_{T}$ gives excellent approximations to the class of (standard) linear predictors, as we formally prove next. The following lemma is analogous to Lemma 11.4, with the significant difference that the number of parameters $h$ does not depend on the stabilizability parameter of the system $\gamma$ at all!
Lemma 13.3.
Suppose that $\mathbf{x}_{0}=0$, then the class $\prod_{h}^{\lambda}$ for $h=O(\log^{2}\frac{T}{\varepsilon})$, $\varepsilon$-approximates the class $\prod_{\star}$.
The proof of this lemma relies on the following crucial property of the spectral basis, whose proof we defer to later.
Lemma 13.4.
Let $\{\phi_{j}\}$ be the eigenvalues of Z. Then for all $j\in[T]$, ,$\alpha\in[0,1]$, we have
$|\phi_{j}^{\top}\mu_{\alpha}|=O\left(T^{2}e^{-\frac{j}{\log T}}\right).$
Using this property, we can prove Lemma 13.3, by simply shifting to the spectral basis representation as follows.
Proof of Lemma 13.3.
Consider any sequence and any predictor from $\prod_{\star}$, say $\pi_{1}=\pi_{A,B,C,0}$. Following equation (13.1), we can write the optimal predictor in the spectral basis:
$\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}$ $=\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1})$
[…]
$=\beta\cdot\mu_{\alpha}^{\top}(\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top})\tilde{\mathbf{u}}_{t-1}$ basis property $\mbox{basis property }\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top}=I$
$=\sum_{i=1}^{T}M_{i}^{\star}\cdot\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}.$ $M_{i}^{\star}=\beta\phi_{i}^{\top}\mu_{\alpha}$
By Lemma 13.4 we have that
$M_{i}^{\star}=\beta\phi_{i}^{\top}\mu_{\alpha}=O(T^{2}e^{-\frac{i}{\log T}}).$
We proceed by an inductive claim on $\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|$. Consider the predictor $\pi_{2}\in\prod_{h}^{\lambda}$ for the same system matrices $M^{\star}$. Then we have
$\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|-\|\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{2}}\|$
$\leq\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}+\mathbf{y}_{t-1}^{\pi_{2}}\|$$\triangle\mbox{-inequality}$
$=\|\sum_{i=1}^{T}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}-\sum_{i=1}^{h}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}\|$
$=\|\sum_{i=h+1}^{T}M^{\star}_{i}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-1}\|$
$\leq\sum_{i=h+1}^{T}\|M_{i}^{\star}\||\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-1}|$
$\leq\sum_{i=h+1}^{T}\|M_{i}^{\star}\|\|\phi_{i}\|_{1}\|\tilde{\mathbf{u}}_{t-1}\|_{\infty}$ Holder inequality
$\leq\sqrt{T}\cdot\sum_{i=h+1}^{T}|M_{i}^{\star}|$ $|\mathbf{u}_{t}|\leq 1,\|\phi_{i}\|_{2}\leq 1$
[…]
$\leq\frac{\varepsilon}{T}.$ choice of $h$
Thus, since same argument exactly applies to $\|\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{2}}\|$, we have that
$\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|\leq\sum_{t=1}^{T}\frac{\varepsilon}{T}\leq\varepsilon.$
It remains to prove Lemma 13.4, which is the main technical lemma of this section.
Lemma 13.4.
Observe that
$\|\mu_{\alpha}\|_{2}^{2}=\sum_{t=0}^{T-1}(1-\alpha)^{2}\alpha^{2i}\leq\frac{(1-\alpha)^{2}}{1-\alpha^{2}}\leq 1-\alpha$
implying that $g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}\leq\|\phi_{j}\|^{2}\|\mu_{\alpha}\|^{2}\leq 1-\alpha$. Similarly, we have
$\left\|\frac{\partial}{\partial\alpha}\mu_{\alpha}\right\|_{2}^{2}$ $=\sum_{t=0}^{T-1}((1-\alpha)i\alpha^{i-1}-\alpha^{i})^{2}$
$\leq 2(1-\alpha)^{2}\sum_{t=0}^{T-1}i^{2}\alpha^{2i-2}+2\sum_{t=1}^{T}\alpha^{2i}$ $(a+b)^{2}\leq 2(a^{2}+b^{2})$
$\leq 2(1-\alpha)^{2}\left(\frac{1}{(1-\alpha^{2})^{2}}+\frac{\alpha^{2}}{(1-\alpha^{2})^{3}}\right)+\frac{2}{1-\alpha^{2}}$
$\leq\frac{6}{1-\alpha},$
where we use the identity $\sum_{i=0}^{\infty}i^{2}\beta^{i-1}=\frac{1}{(1-\beta)^{2}}+\frac{\beta}{(1-\beta)^{3}}$ with the substitution $\beta=\alpha^{2}$. Using this, we obtain an upper bound on the Lipschitz constant of $g(\alpha)$ as
$|g^{\prime}{(\alpha)}|$ $=\left\lvert\frac{\partial}{\partial{\alpha}}(\mu_{\alpha}^{\top}\phi_{j})^{2}\right\rvert=2|\mu_{\alpha}^{\top}\phi_{j}|\left\lvert\frac{\partial}{\partial{\alpha}}(\mu_{\alpha}^{\top}\phi_{j})\right\rvert$
$\leq 2\|\mu_{\alpha}\|_{2}\|\phi_{j}\|^{2}_{2}\left\|\frac{\partial}{\partial{\alpha}}\mu_{\alpha}\right\|_{2}$ Cauchy-Schwarz
$\leq 2\sqrt{1-\alpha}\sqrt{\frac{6}{1-\alpha}}\leq 6.$ $\|\phi_{i}\|_{2}\leq 1$
Let $g_{\max}=\max_{\alpha}g(\alpha)$ be the maximal value that $g$ attains from zero to one. Any non-negative L-Lipschitz function that reaches $g_{\max}$ and has the area of $A$, satisfies:
$A\geq\frac{g_{\max}^{2}}{2L}\ \ \Rightarrow\ \ g_{\max}^{2}\leq{2LA},$
since the minimal area is realized by a triangle.
Thus, the maximum value of $g(\alpha)$ is bounded by, using Theorem 13.1,
[…]
$=8T^{2}\sigma_{j}$
$\leq O(T^{2}e^{-\frac{j}{\log T}})$ Theorem 13.1
13.2.1 Generalization to High Dimension
Consider a linear dynamical system with matrices $A,B,C$ and further assume that $A$ is symmetric. This implies that $A$ is diagonalizable over the real numbers as $A=U^{\intercal}DU$. Decomposing over each dimension $j\in[d]$ separately, and denoting $D_{jj}=\alpha_{j}$, we can write (13.1):
$\hat{\mathbf{y}}^{\pi}_{t}-\mathbf{y}_{t-1}-CB\mathbf{u}_{t-1}$
$=\sum_{i=1}^{t-1}CU^{\top}(D^{i-1}-D^{i-2})UB\mathbf{u}_{t-1-i}$
$=\sum_{i=1}^{t-1}CU^{\top}(\sum_{j=1}^{d_{x}}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})e_{i}e_{i}^{\top})UB\mathbf{u}_{t-i-1}$
$=\sum_{j=1}^{d_{x}}\beta_{j}\sum_{i=1}^{t-1}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})\mathbf{u}_{t-1-i}$ for scalars $\beta_{j}$
$=\sum_{j=1}^{d_{x}}{M}^{\star}_{j}\mu_{\alpha_{j}}(\mathbf{u}_{1:t-1})$
Thus, we can apply the same approximation techniques from the previous section, and derive an analoguous conclusion to Lemma 13.3 for dimension higher than one. The details are left as exercise 13.5.
13.3 Online Spectral Filtering
The prediction class $\prod_{h}^{\lambda}$ can be learned by online gradient descent in the online convex optimization framework, giving rise to the following algorithm.
Algorithm 7 Online Spectral Filtering
1:  Input: $M^{1}_{1:h}$, convex constraints set $\mathcal{K}\subseteq{\mathbb{R}}^{h\times d_{y}\times d_{u}}$, horizon $T$
2:  Compute $h$ top eigenvectors of $Z_{T}$, denoted $\phi_{1},...,\phi_{h}$.
3:  for $t=1$ to $T$ do
4:     Predict $\hat{\mathbf{y}}_{t}=\sum_{j=1}^{h}M^{t}_{j}\phi_{j}(\tilde{\mathbf{u}}_{t-i})$ , as per definition 13.2, and observe true $\mathbf{y}_{t}$.
5:     Define cost function $f_{t}(M)=\|\mathbf{y}_{t}-\tilde{\mathbf{y}}_{t}\|^{2}$.
6:     Update and project:
$M^{t+1}=M^{t}-\eta_{t}\nabla f_{t}(M^{t})$
$M^{t+1}=\mathop{\Pi}_{\mathcal{K}}(M^{t+1})$
7:  end for
The eigenvectors of $Z$ have an interesting structure: their magnitude does not decay with time, which explains their improved approximation as captured formally in Lemma 13.3.
Figure 13.2: The filters obtained by the eigenvectors of $Z$.
The online gradient descent theorem, as per the previous chapter, directly gives the following corollary,
Theorem 13.5.
For choice of step sizes $\eta_{t}=O(\frac{1}{\sqrt{t}})$, we have that
$\sum_{t}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\prod_{h}^{\lambda}}\|\mathbf{y}_{t}-\mathbf{y}_{t}^{\pi}\|^{2}=O(\sqrt{T}).$
13.4 Bibliographic Remarks
Under probabilistic assumptions, the work of [HMR16] shows that the matrices $A,B,C,D$ of an unknown and partially observed time-invariant linear dynamical system can be learned using first-order optimization methods.
Deviating from classical recovery and into the improper prediction world, the work of [HSZ17] devised an improper learning technique using the eigenvectors of the Hankel matrix $Z$. The spectral properties of Hankel matrices used in this chapter are from [BT17].
For an extension of the spectral learning technique to asymmetric matrices, see [HLS+18].
13.5 Exercises
13.1. Generalize Lemma 13.3 to all real numbers in the range $[-1,1]$, instead of $[0,1]$.
13.2. Generalize Lemma 13.3 to any dimension $d_{x}>1$.
Appendix A Concepts from Online Convex Optimization
In this section we give a brief introduction to the theory of Online Convex Optimization. The reader is referred to [Haz16] for a comprehensive survey.
In the decision making framework of Online Convex Optimization, a player iteratively chooses a point in a convex set $\mathcal{K}\subseteq{\mathbb{R}}^{n}$, and suffers a loss according to an adversarially chosen loss function $f_{t}:\mathcal{K}\mapsto{\mathbb{R}}$.
Let ${\mathcal{A}}$ be an algorithm for OCO, which maps a certain game history to a decision in the decision set:
$\mathbf{x}_{t}^{\mathcal{A}}={\mathcal{A}}(f_{1},...,f_{t-1})\in\mathcal{K}.$
We define the regret of ${\mathcal{A}}$ after $T$ iterations as:
$\mathrm{{regret}}_{T}({\mathcal{A}})=\sup_{\{f_{1},...,f_{T}\}\subseteq{\mathcal{F}}}\left\{\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t}^{\mathcal{A}})-\min_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})\right\}.$ (A.1)
If the algorithm is clear from the context, we henceforth omit the superscript and denote the algorithm’s decision at time $t$ simply as $\mathbf{x}_{t}$. Intuitively, an algorithm performs well if its regret is sublinear as a function of $T$ (i.e. $\mathrm{{regret}}_{T}({\mathcal{A}})=o(T)$), since this implies that on average, the algorithm performs as well as the best fixed strategy in hindsight. We proceed to describe a basic algorithm for online convex optimization.
A.1 Online Gradient Descent
Perhaps the simplest algorithm that applies to the most general setting of online convex optimization is online gradient descent. This algorithm, which is based on standard gradient descent from offline optimization. Pseudo-code for the algorithm is given in Algorithm 8.
Algorithm 8 online gradient descent
1:  Input: convex set $\mathcal{K}$, $T$, $\mathbf{x}_{1}\in\mathcal{K}$, step sizes $\{\eta_{t}\}$
2:  for $t=1$ to $T$ do
3:     Play $\mathbf{x}_{t}$ and observe cost $f_{t}(\mathbf{x}_{t})$.
4:     Update and project:
$\mathbf{y}_{t+1}=\mathbf{x}_{t}-\eta_{t}\nabla f_{t}(\mathbf{x}_{t})$
$\mathbf{x}_{t+1}=\operatorname*{\arg\min}_{\mathbf{x}\in\mathcal{K}}\|\mathbf{x}-\mathbf{y}_{t+1}\|^{2}$
5:  end for
In each iteration, the algorithm takes a step from the previous point in the direction of the gradient of the previous cost. This step may result in a point outside of the underlying convex set. In such cases, the algorithm projects the point back to the convex set, i.e. finds its closest point in the convex set. Despite the fact that the next cost function may be completely different than the costs observed thus far, the regret attained by the algorithm is sublinear. This is formalized in the following theorem. Here we define the constants $D$ and $G$ as follows:
  • $D$ = Euclidean diameter of the decision set, i.e.
    $\max_{\mathbf{x},\mathbf{y}\in\mathcal{K}}\|\mathbf{x}-\mathbf{y}\|.$
  • $G$ = an upper bound on the Lipschitz constant of the cost functions $f_{t}$, or equivalently, the subgradient norms of these functions. Mathematically,
    $\forall t\ ,\ \|\nabla f_{t}(\mathbf{x}_{t})\|\leq G.$
Theorem A.1.
Online gradient descent with step sizes $\{\eta_{t}=\frac{D}{G\sqrt{t}},\ t\in[T]\}$ guarantees the following for all $T\geq 1$:
$\mathrm{{regret}}_{T}=\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\min_{\mathbf{x}^{\star}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x}^{\star})\ \leq\frac{3}{2}{GD}\sqrt{T}.$
Proof.
Let $\mathbf{x}^{\star}\in\operatorname*{\arg\min}_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})$. Define $\nabla_{t}\triangleq\nabla f_{t}(\mathbf{x}_{t})$. By convexity
$f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{x}^{\star})\leq\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$ (A.2)
We first upper-bound $\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$ using the update rule for $\mathbf{x}_{t+1}$ and the Pythagorean Theorem, we have the projections only decrease the distance to the set $\mathcal{K}$:
$\|\mathbf{y}_{t+1}-\mathbf{x}^{\star}\|^{2}\ =\ \left\|\mathop{\Pi}_{\mathcal{K}}(\mathbf{x}_{t}-\eta_{t}\nabla_{t})-\mathbf{x}^{\star}\right\|^{2}\leq\left\|\mathbf{x}_{t}-\eta_{t}\nabla_{t}-\mathbf{x}^{\star}\right\|^{2}.$ (A.3)
Hence,
$\|\mathbf{y}_{t+1}-\mathbf{x}^{\star}\|^{2}$$\leq$ $\ \|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}+\eta_{t}^{2}\|\nabla_{t}\|^{2}-2\eta_{t}\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$
$2\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$$\leq$ $\ \frac{\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$ (A.4)
Summing (A.2) and (A.1) from $t=1$ to $T$, and setting $\eta_{t}=\frac{D}{G\sqrt{t}}$ (with $\frac{1}{\eta_{0}}\triangleq 0$):
$2\left(\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{x}^{\star})\right)\leq 2\sum_{t=1}^{T}\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$
$\leq\sum_{t=1}^{T}\frac{\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|^{2}}{\eta_{t}}+G^{2}\sum_{t=1}^{T}\eta_{t}$
$\leq\sum_{t=1}^{T}\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)+G^{2}\sum_{t=1}^{T}\eta_{t}$ $\frac{1}{\eta_{0}}\triangleq 0,$
$\|\mathbf{x}_{T+1}-\mathbf{x}^{*}\|^{2}\geq 0$
$\leq D^{2}\sum_{t=1}^{T}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)+G^{2}\sum_{t=1}^{T}\eta_{t}$
$\leq D^{2}\frac{1}{\eta_{T}}+G^{2}\sum_{t=1}^{T}\eta_{t}$ telescoping series
$\leq 3DG\sqrt{T}.$
The last inequality follows since $\eta_{t}=\frac{D}{G\sqrt{t}}$ and $\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\leq 2\sqrt{T}$. ∎
The online gradient descent algorithm is straightforward to implement, and updates take linear time given the gradient. However, the projection step may in general take significantly longer via convex optimization.

### v4

@@
-version 0.6
+version 0.75
 […]
-position=current page.north, angle=0, nodeanchor=north, vshift=-5mm, opacity=0.5, scale=1, contents=Draft; comments/suggestions welcome at nonstochastic.control@gmail.com
+position=current page.north, angle=0, nodeanchor=north, vshift=-5mm, opacity=0.5, scale=1, contents=Draft; comments and suggestions welcome at nonstochastic.control@gmail.com
 Introduction to Online Control
-version 0.6
+version 0.75
 Elad Hazan    Karan Singh
@@
 In online nonstochastic control, both the cost functions as well as the perturbations from the assumed dynamical model are chosen by an adversary.
-Thus the optimal policy is not defined a priori.
-Rather, the target is to attain low regret against the best policy in hindsight from a benchmark class of policies.
-This objective suggests the use of the decision making framework of online convex optimization as an algorithmic methodology.
+Thus, the optimal policy is not defined a priori.
+Rather, the goal is to attain low regret against the best policy in hindsight from a benchmark class of policies.
+This objective suggests the use of the decision-making framework of online convex optimization as an algorithmic methodology.
 The resulting methods are based on iterative mathematical optimization algorithms and are accompanied by finite-time regret and computational complexity guarantees.
+Acknowledgments
+This book was a journey that started in Princeton at the Google DeepMind Lab in the summer of 2018.
+At that time Sham Kakade visited the group, which consisted of the authors as well as Naman Agarwal, Brian Bullins, Xinyi Chen, Cyril Zhang and Yi Zhang, henceforth the gang.
+The gang worked on various aspects of optimization, reinforcement learning, and dynamical systems.
+The first paper, in which online nonstochastic control was basically invented, started in room 401 in the computer science building at Princeton University.
+We remember and cherish the exact moment along with the mathematical writings in white chalk on that blackboard.
+Ever since then, the gem of an idea turned out into a complete theory which is now mainstream in machine learning and control, with numerous applications and refinements.
+Notable events that contributed to this theory include the following.
+Max Simchowitz, who was a graduate student then, visited Princeton during 2019-2020, where he collaborated with the group, and the theory was extended to partial observed systems.
+Paula Gradu, who was then an undergraduate, and Edgar Minasyan, while a grad student, contributed to the study of time-varying and nonlinear dynamical systems.
+Anirudha Majumdar joined for several important projects, notably for iterative linear control.
+Udaya Ghai and Arushi Gupta studied extending the GPC to model-free RL.
+Most recently, this theory has been used to design new architectures for sequence prediction.
+The theory of spectral filtering was extended by the aforementioned members of the group, and newer folks joined, including Daniel Suo and Annie Marsden.
+This book was further refined through teaching.
+Elad Hazan gratefully acknowledges the numerous contributions and insights of the students of the course computational control theory delivered at Princeton during COVID on Zoom during 2020-2021.
+The text was further enhanced by students of theoretical machine learning taught at Princeton University in the fall of 2024-2025.
+The following students were particularly helpful making insightful comments and in finding various typos and suggesting improvements: Gon Buzaglo, Anand Brahmbhatt, Sofiia Druchyna (and many more we wish to announce very soon).
+Throughout the writing of this book, we engaged in extensive discussions with ChatGPT, an AI assistant.
+Although it does not claim authorship, its ability to refine mathematical reasoning and clarify complex topics has been a remarkable collaboration in human-AI interaction.
+In fact, it wrote this acknowledgment all by itself!
 List of Symbols
@@
 $\kappa$ norm bound on policy parameters
+$K_{t}$ stabilizing linear controller at time $t$
+$K_{r}(A,B)$ Kalman controllability matrix of order $r$
+$\hat{K}(A,C)$ Kalman observability matrix
 Reinforcement learning
@@
 The advance of the deep learning revolution has brought to light the power and robustness of algorithms that are based on gradient methods.
-Deep learning pioneer and Turing award winner, professor Yann Lecun, is attributed the following quote from a Barbados workshop on deep learning theory, circa Feb 2020,
+The following quote is attributed to deep learning pioneer Yann Lecun during a Barbados workshop on deep learning theory circa Feb 2020,
 “Control = reinforcement learning with gradients.”
@@
 This gives rise to a simple but powerful framework called nonstochastic control theory.
-The framework is a natural fit to apply techniques from online convex optimization, resulting in new, efficient, gradient based control methods.
+The framework is a natural fit to apply techniques from online convex optimization, resulting in new, efficient, gradient-based control methods.
 1.2 The Origins of Control
 The industrial revolution was in many ways shaped by the steam engine that came to prominence in the 18th century.
-One of its most critical components was introduced by the Scottish inventor James Watt: amongst the earliest examples of a control mechanism, the Centrifugal Governor.
+One of its most critical components was introduced by the Scottish inventor James Watt: the centrifugal governor, which is amongst the earliest examples of a control mechanism.
 A major source of losses in efficiency of earlier designs was the difficulty in regulating the engine.
@@
 Here $\mathbf{x}_{t}\in{\mathbb{R}}^{d_{x}}$ is the state of the system, $\mathbf{y}_{t}\in{\mathbb{R}}^{d_{y}}$ is the observation, $\mathbf{u}_{t}\in{\mathbb{R}}^{d_{u}}$ is the control input, and $\mathbf{w}_{t}\in{\mathbb{R}}^{d_{w}}$ is the perturbation.
-The function $g_{t}$ is a mapping from the state to the observed state, and the function $f_{t}$ is the transition function, given current state $\mathbf{x}_{t}$ and control input $\mathbf{u}_{t}$.
+The function $g_{t}$ is a mapping from the state to the observed state, and the function $f_{t}$ is the transition function, given current state $\mathbf{x}_{t}$, control input $\mathbf{u}_{t}$ and perturbation $\mathbf{w}_{t}$.
 The subscript indicates the relevant quantities at time step $t$.
 Given a dynamical system, there are several possible objectives we can formulate.
-A generic formulation appears below, which captures many common control formulations, notably the optimal control problem and the robust control problem, which will be addressed in much more detail later in this book.
+We give a generic formulation below that captures many common control formulations, notably the optimal control problem and the robust control problem.
+These will be addressed in much more detail later in this book.
 In this problem formulation, the goal is to minimize the long-term horizon cost given by a sequence of cost functions.
 Definition 1.2 (A generic control problem).
-Given a dynamical system with starting state $(\mathbf{x}_{1},\mathbf{y}_{1})$, iteratively construct controls $u_{t}$ to minimize a sequence of loss functions $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$, $c_{t}:{\mathbb{R}}^{d_{y}\times d_{u}}\to{\mathbb{R}}$.
+Given a dynamical system with starting state $(\mathbf{x}_{1},\mathbf{y}_{1})$, iteratively construct controls $u_{t}$ to minimize a sequence of loss functions $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$, $c_{t}:{\mathbb{R}}^{d_{y}+d_{u}}\to{\mathbb{R}}$.
 This is given by the following mathematical program
@@
 The goal of ventilator control is to regulate pressure according to some predetermined pattern.
-As a dynamical system, we can write $p_{t+1}=f(p_{t},u_{t},w_{t}),$ and we can define the cost function to be $c_{t}(p_{t},u_{t})=|p_{t}-p_{t}^{*}|^{2},$ where $p_{t}^{*}$ is the desired pressure at time $t$.
+As a dynamical system, we can write $v_{t+1}=f(v_{t},u_{t},w_{t})$, $p_{t}=g(v_{t})$, and we can define the cost function to be $c_{t}(p_{t},u_{t})=|p_{t}-p_{t}^{*}|^{2},$ where $p_{t}^{*}$ is the desired pressure at time $t$.
 The goal, then, is to choose a sequence of control signals $\{u_{t}\}_{t=1}^{T}$ to minimize the total cost over $T$ iterations.
-In the figure below we see an illustration of a simple controller, called the PID controller, as it performs on a ventilator and lung simulator.
-Figure 1.2: Performance of the PID controller on a mechanical ventilator
+In the figure below we see an illustration of a simple controller, that we will soon describe, called the PID controller, as it performs on a ventilator and lung simulator.
+Figure 1.2: Performance of the PID controller on a mechanical ventilator, from [SZG+21].
 1.4 Simple Control Algorithms
@@
 The following methods are simplistic but still widely used.
-For example, PID algorithm is still used in the majority of medical ventilators today.
+For example, the PID algorithm is still used in the majority of medical ventilators today.
 In later sections we describe a modern optimization-based framework that yields more sophisticated algorithms that are capable of better performance.
@@
 Bang-bang control is one of the simplest control strategies.
-Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a pre-specified dynamic range $[x_{\min},x_{\max}]$ (that can vary with time), and where the scalar control $u_{t}$ input must at all times be within the interval $[u_{\min},u_{\max}]$.
+Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a pre-specified dynamic range $[x_{\min},x_{\max}]$ (that can vary with time), and where the scalar control $u_{t}$ input must at all times be within the interval $[u_{\min},u_{\max}]$, that is assumed to contain zero for simplicity.
 A bang-bang controller chooses the control input at time $t$ as
@@
 We would like an adaptive control theory which allows us to attain the best of both worlds: an efficient and fast flight when the weather permits, and a careful drone maneuver when this is required by the conditions.
-Can we design a control theory that will allow us to take into account the specific instance perturbations and miss-specifications, and give us finite time provable guarantees?
+Can we design a control theory that will allow us to take into account the specific instance perturbations and misspecifications, and give us finite time provable guarantees?
 This is the subject of online non-stochastic control!
@@
 When dealing with non-stochastic, i.e.
-arbitrary or even adversarial perturbations, the optimal policy is not clear a priori.
+arbitrary or even adversarial perturbations, the optimal policy, i.e.
+a decision making rule mapping observations to control inputs, is not clear a priori.
 Rather, an optimal policy for the observed perturbations can be determined only in hindsight.
@@
 Specifically, we consider regret with respect to a reference class of policies.
-In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$, and then observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ according to an adversarially chosen loss function.
+In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$.
+Then the controller observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ according to an adversarially chosen loss function.
+For simplicity, we consider the case of full state observation where $\mathbf{y}_{t}=\mathbf{x}_{t}$.
 Let $\Pi=\{\pi:\mathbf{x}\to\mathbf{u}\}$ be a class of policies.
@@
   1. 1.
-    First, the best policy in hindsight may be very complicated to describe and reason about.
+    First, the best policy in hindsight out of all possible policies may be very complicated to describe and reason about.
 This argument was made by the mathematician Tyrell Rockafellar, see the bibliographic section for more details.
@@
 the best policy in hindsight.
-Rather, a different performance metric called competitive ratio can be analyzed, as we do later in the book.
+Rather, a different performance metric called competitive ratio can be analyzed, as we discuss later in the book.
 The non-stochastic control problem can now be stated as finding an efficient algorithm that minimizes the worst-case regret vs.
@@
 The texts [Ste94, ZDG96a], on the other hand, delve deep into results on linear control, the most well-understood branch of control theory.
-The excellent text of [SL91] gives a thorough introduction to adaptive control theory.
+Rockafellar [Roc87] proposed the use of convex cost functions for modeling the cost in a linear dynamical system to incorporate constraints on the state and control, and suggested computational difficulties with this approach.
+The text [SL91] gives a thorough introduction to adaptive control theory.
 Online nonstochastic control theory and the Gradient Perturbation Control algorithm were proposed in [ABH+19].
@@
 In the task of mechanical ventilation, the control algorithm dictates the amount of oxygen inflow $u_{t}$ (in $\mathrm{m}^{3}/s$) into a patient’s lungs.
-This inflow determines the volume of the lungs $v_{t}$ via a simple linear relation,
+This inflow approximately determines the volume of the lungs $v_{t}$ via a simple linear relation,
 $v_{t+1}=v_{t}+\Delta u_{t},$
@@
 $p_{t}=C_{0}+C_{1}v_{t}^{-1/3}+C_{2}v_{t}^{5/3}$
-The objective of the control algorithm is to track a physician-prescribed ideal pressure trajectory $p^{*}_{t}$ as closely as possible, say as measured by in squared loss $c_{t}(p_{t},u_{t})=|p_{t}-u_{t}|^{2}+|u_{t}|^{2}$, while limiting the maximum instantaneous airflow to minimize patient discomfort.
+The objective of the control algorithm is to track a physician-prescribed ideal pressure trajectory $p^{*}_{t}$ as closely as possible, say as measured by in squared loss $c_{t}(p_{t},u_{t})=|p_{t}-p^{*}_{t}|^{2}+|u_{t}|^{2}$, while limiting the maximum instantaneous airflow to minimize patient discomfort.
 See Figure 1.2 for a schematic representation of the airflow when a lung is connected to a mechanical ventilator.
@@
 The double integrator is a common example of a linear dynamical system.
-In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics.
+In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics on a one dimensional trajectory (i.e.
+the real number line).
 The state $\mathbf{x}_{t}$ is 2-dimensional, with the first coordinate indicating the position of a point object and the second denoting its velocity.
@@
 The cost function is naturally expressed as
-$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}x^{*}\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
+$c(\mathbf{x}_{t},u_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}x^{*}\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
 Figure 2.2: The position and velocity visualized for a double integrator.
@@
 Given the vertical angle $\theta_{t}$ of the rod, the physics of this system may be summarized as
-$\theta_{t+1}=\theta_{t}+\Delta\dot{\theta}_{t},\qquad\dot{\theta}_{t+1}=\dot{\theta}_{t}+\Delta\frac{u_{t}-mg\sin\theta}{ml^{2}},$
+$\theta_{t+1}=\theta_{t}+\Delta\dot{\theta}_{t},\qquad\dot{\theta}_{t+1}=\dot{\theta}_{t}+\Delta\frac{u_{t}-mg\ \sin\theta_{t}}{ml^{2}},$
 where $\Delta$ is the time resolution of the controller.
@@
 This may be expressed via the cost function
-$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}\pi\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
+$c(\mathbf{x}_{t},u_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}\pi\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
 2.1.4 Aircraft Dynamics
 A final example of a dynamical system we consider is the longitudinal (vertical tilt) motion of an aircraft that is regulated by automated controllers.
-We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system.
+We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system, from [Boy10].
 The 2-dimensional control input $\mathbf{u}_{t}$ corresponds to the elevator angle and the thrust.
@@
 $+\begin{bmatrix}-0.003&0.039&0&-0.322\\ -0.065&-0.319&7.74&0\end{bmatrix}\mathbf{w}_{t}$
-Figure 2.4: Longitudinal control in an aircraft. Source: Wikipedia
+Figure 2.4: Longitudinal control in an aircraft. Source: Wikipedia.
 The quantities of interest for assessing the performance of the controller are aircraft speed, climb rate, each of which is a linear transformation of the state, and the applied thrust.
 The corresponding cost is
-$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\begin{bmatrix}1&0&0&0\\ 0&-1&0&7.74\end{bmatrix}\mathbf{x}_{t}-\mathbf{x}_{t}^{*}\right\|^{2}+\left\|\begin{bmatrix}0\\ 1\end{bmatrix}\mathbf{u}_{t}\right\|^{2}.$
+$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\begin{bmatrix}1&0&0&0\\ 0&-1&0&7.74\end{bmatrix}\mathbf{x}_{t}-\mathbf{x}_{t}^{*}\right\|^{2}+\left\|\begin{bmatrix}0&1\end{bmatrix}\mathbf{u}_{t}\right\|^{2}.$
 2.1.5 Epidemiological Models
@@
 Let $u_{t}$ be the number of vaccines administered at time $t$.
+The dynamics of the basic SIR model are given by
 $S_{t+1}$ $=S_{t}-\beta S_{t}I_{t}-\alpha u_{t}$
@@
 Here the coefficient $\beta$ measures how communicable the infection is, $\gamma$ measures how quickly an infected person recovers, and $\alpha$ represents the efficacy of the vaccine in preventing infections.
-Figure 2.5: Trajectory of a continuous-time SIR model. Source: Wikipedia
+Figure 2.5: Trajectory of a continuous-time SIR model. Source: Wikipedia.
 2.2 Solution Concepts for Dynamical Systems
@@
 $\mathbf{x}_{t+1}={A}\mathbf{x}_{t}.$
-Clearly 00 is an equilibrium point as ${A}0=0$ for all ${A}$.
-If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vectors converges to the null veector.
-Thus 00 is a globally attractive, stable equilibrium.
-In full generality, determining if a system has an equilibirum is an intractable problem, as we show in this chapter.
+Clearly $\mathbf{0}$ is an equilibrium point as ${A}\mathbf{0}=\mathbf{0}$ for all ${A}$.
+If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vectors converges to the null vector.
+Thus $\mathbf{0}$ is a globally attractive, stable equilibrium.
+In full generality, determining if a system has an equilibrium is an intractable problem, as we show in this chapter.
 2.2.2 Stabilizability and Controllability
@@
 A Dynamical System is stabilizable if and only if, for every initial state, there exists a sequence of controls that drives the deterministic-equivalent system to the zero state.
-That is, $\lim_{t\rightarrow\infty}{\mathbf{x}_{t}}=0$.
+That is, $\lim_{t\rightarrow\infty}{\mathbf{x}_{t}}=\mathbf{0}$.
 Definition 2.3.
@@
 A control policy $\pi$ for a dynamical system $f$ is $\gamma$-stabilizing if and only if the following holds.
-For every starting state $\|\mathbf{x}_{0}\|\leq 1$ and every sequence of adversarially chosen bounded perturbations $\mathbf{w}_{1:T}$ such that $\|\mathbf{w}_{t}\|\leq 1$, let $\mathbf{x}_{t}^{\pi}$ be the state obtained by using policy $\pi$,
-$\mathbf{x}_{t+1}^{\pi}=f(\mathbf{x}_{t}^{\pi},\pi(\mathbf{x}_{t})).$
+For every starting state $\|\mathbf{x}_{0}\|\leq 1$ and every sequence of adversarially chosen bounded perturbations $\mathbf{w}_{1:T}$ such that $\|\mathbf{w}_{t}\|\leq 1$ for all $t\in[T]$, let $\mathbf{x}_{t}^{\pi}$ be the state obtained by using policy $\pi$,
+$\mathbf{x}_{t+1}^{\pi}=f(\mathbf{x}_{t}^{\pi},\pi(\mathbf{x}_{t}),\mathbf{w}_{t}).$
 Then the norm of the system’s state at every iteration is bounded by an absolute constant independent of $T$, i.e.
@@
 For a formal proof of attractiveness in the continuous case, see section 9.2 of [Str14] or the original paper of Lorenz [Lor63].
-Figure 2.6: Julia set
-Figure 2.7: Lorenz attractor
+Figure 2.6: Julia set. Souce: Wikipedia.
+Figure 2.7: Lorenz attractor. Source: Wikipedia.
 2.3.1 Computational Hardness of Stabilizability
@@
 For a family of dynamical systems described as polynomials with integer coefficients, determining the stabilizability of any member is NP-hard.
-We show that this problem can be reduced from the 3SAT problem, a known NP-hard problem. This implies that determining the stabilizability of a dynamical system is NP-hard in general.
+We show that this problem can be reduced to the 3SAT problem, a well-known NP-hard problem. This implies that determining the stabilizability of a dynamical system is NP-hard in general.
 Proof.
@@
 Now, consider the scalar-state dynamical equation given by
-$x_{t+1}=x_{t}(1-\varepsilon+p_{F}(\mathbf{u}_{t}))+p_{F}(\mathbf{u}_{t}),$
-with $p$ being the polynomial family we just constructed, and $\varepsilon$ being an arbitrarily small constant.
-For the system to reach the null state from an arbitrary start state, we would need to verify if there is an assignment of $\mathbf{u}$ such that $p_{F}(\mathbf{u})=0$, which we just proved is NP-hard.
+$x_{t+1}=(1-\varepsilon)x_{t}+p_{F}(\mathbf{u}_{t}),$
+with $p$ being the polynomial family we just constructed and $\varepsilon>0$ an arbitrarily small constant.
+For the system to reach the null state from an arbitrary start state, we would need to verify if there is an assignment of $\mathbf{u}$ such that $p_{F}(\mathbf{u})=0$, which as we just proved is NP-hard.
 ∎
@@
 2.1.
-Propose a computationally efficient representation for the problem of stabilizability.
-Prove that for your proposal, the problem of determining whether a given dynamical system is stabilizable belongs to the class NP.
-Thus, together with the material of this chapter, conclude that stabilizability is NP-complete.
-2.2.
 Strengthen Theorem 2.5 to show that checking whether a dynamical system that is a polynomial of degree four is stabilizable is NP-hard.
 Hint: Reduce from the NP-hard problem of MAX-2-SAT instead of 3-SAT.
+2.2. Let $A\in\mathcal{R}^{d\times d}$ be a diagonalizable matrix. Consider the following dynamical system:
+$\mathbf{x}_{t+1}=\frac{A\mathbf{x}_{t}}{\|\mathbf{x}_{t}\|}.$
+What are the stable and attractive equilibrium points $\mathbf{x}\in\mathcal{R}^{d}$ of the system? Are they also globally attractive? Your answer may depend on $A$.
+2.3.
+Let $\alpha,\beta,\gamma>0$.
+Does the SIR-model, as it is described in this chapter, define a stabilizable dynamical system (for an appropriate and fixed definition of the “zero state”)?
+What about controllability?
+We assume that only starting states with the property $S_{0}+I_{0}+R_{0}=1$ are allowed, and that $S_{t},I_{t},R_{t},u_{t}\geq 0$ has to hold at all times.
+2.4. Consider the double integrator system:
+$\mathbf{x}_{t+1}=A\mathbf{x}_{t}+Bu_{t}+\mathbf{w}_{t},$
+where
+$A=\begin{bmatrix}1&\Delta\\ 0&1\end{bmatrix},\quad B=\begin{bmatrix}0\\ 1\end{bmatrix},\quad\mathbf{x}_{t}=\begin{bmatrix}x_{t}\\ v_{t}\end{bmatrix}\in\mathbb{R}^{2},\quad u_{t}\in\mathbb{R},\quad\Delta>0.$
+The goal is to ensure the system reaches the desired state $[x^{*},0]^{\top}$ as $t\to\infty$ from any initial condition.
+  1. 1.
+    Prove that $(A,B)$ is controllable.
+That is, show that for any initial state and any target state, there exists a finite sequence of inputs that transitions the system from the initial state to the target state.
+  2. 2.
+    Design a stabilizing state-feedback law
+[…]
+    with $K=[k_{1}\;k_{2}]$, that places all eigenvalues of $A-BK$ strictly inside the unit circle.
+Find conditions on $\Delta,k_{1},k_{2}$ that ensure all eigenvalues of $A-BK$ are real and lie strictly between $-1$ and $1$, thus guaranteeing $\mathbf{x}_{t}\to[x^{*},0]^{\top}$.
 Chapter 3 Markov Decision Processes
@@
 The agent observes its position (or state) in the environment and takes actions that transition it to a new state.
-The environment looks at the agent’s state and hands out rewards based on a hidden set of criteria.
-Figure 3.1: Source: Reinforcement Learning: An Introduction. Sutton & Barto, 2017.
+The environment looks at the agent’s state and hands out rewards based on a set of criteria hidden from the agent.
+Figure 3.1: Source: Wikipedia.
 Typically, the goal of reinforcement learning is for the agent to learn behavior that maximizes the total reward it receives from the environment.
@@
 The Markovian assumption: Future states, observations, and rewards are independent of the past given the current state.
-The reader can reason about situations where such an assumption is reasonable.
+The reader can imagine situations where such an assumption is reasonable.
 For example, in the inverted pendulum setting described in section 2.1.3, the assumption clearly holds.
-It is not significant for the control problem how a certain state, describing the angle and the angular velocity, was reached, in order to optimally control the pendulum from this state onwards.
+It is not significant for the control problem how a certain state, describing the angle and the angular velocity, was reached, in order to optimally control the pendulum from this state onward.
 Other problems may call into question this assumption.
@@
 3.1.1 A Classical Example: Bubbo’s World
-Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel & Norvig [RN02].
+Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel and Norvig [RN02].
 In this example, a robot starts in the bottom leftmost square, and its goal is to reach the $+1$ reward.
-Figure 3.2: Source: [RN02].
+The blocks $+1,-1$ represent terminal states, meaning that the game ends when the robot reaches them.
+The block ‘X’ represents an obstacle through which the robot cannot travel.
+StartX+1-1
+Figure 3.2: A visual representation of Bubbo’s world.
 In this example, the robot has controls that allow it to move left / right / forward / backward.
-The controls, however, are not reliable, and with $0.8$ probably the robot moves to wherever it intends.
-With the remaining $0.2$ probability the robot moves sideways, $0.1$ probability to each side.
+The controls, however, are not reliable, and with $0.8$ probability the robot moves to wherever it intends.
+With the remaining $0.2$ probability the robot moves sideways, that is, with probability $0.1$ to each side.
 Notice that the Markovian assumption clearly holds for this simple example.
@@
 This mapping is called a “policy”, and it depends on both the structure of the problem as well as on the rewards.
-For example, if the rewards elsewhere are slightly negative, Bubbo wants to reach the terminal state with a persistent $+1$ reward, as given in the following figure.
+For example, if the rewards at all other states are slightly negative, Bubbo wants to reach the terminal state with a persistent $+1$ reward, as given in the following figure.
 […]
 Figure 3.3: The optimal policy (mapping from state to action) with reward for each step taken being -0.04.
-On the other hand, if the rewards are very negative, Bubbo just wants to reach any terminal state as soon as possible, as in the following.
+On the other hand, if the rewards are very negative, Bubbo just wants to reach any terminal state as soon as possible, as in figure 3.4.
 […]
@@
 The probabilities of moving between states are the weights on the edges of the graph.
-Figure 3.2.1 describes a Markov Chain corresponding to a weather process.
+The transition matrix is of dimension $S\times S$, where $S$ is the number of states, and in entry $i,j$ we have the probability of moving from state $i$ to state $j$.
+Figure 3.5 describes a Markov Chain corresponding to a weather process with a corresponding transition matrix $P$ as below
+$P=\begin{pmatrix}0.9&0.1&0&0\\ 0.5&0&0.4&0.1\\ 0&0.5&0&0.5\\ 0&0&0&1\\ \end{pmatrix}.$
 RainSunny dayCloudySnow0.10.90.50.10.40.50.51
-$P=\begin{pmatrix}0.9&0.1&0&0\\ 0.5&0&0.4&0.1\\ 0&0.5&0&0.5\\ 0&0&0&1\\ \end{pmatrix}.$
+Figure 3.5: A graphical description of a Markov decision process. Each state is the weather for a given day and the transition probabilities describe the chance of weather for the next day.
 Denote the distribution over states at time $t$ by $\mathbf{p}_{t}$. Then it can be seen that
-$\mathbf{p}_{t+1}=\mathbf{p}_{t}P=\mathbf{p}_{0}P^{t},$
+$\mathbf{p}_{t+1}=\mathbf{p}_{t}P=\mathbf{p}_{0}P^{t+1},$
 where $\mathbf{p}_{0}$ is the initial state distribution represented as a vector.
@@
 Existence, uniqueness, and convergence to stationarity are the subject of Ergodic theory, a rich and well-studied branch of mathematics.
-For our discussion, we are interested in sufficient conditions for stationary distribution to exist.
+For our discussion, we are interested in sufficient conditions for stationary distributions to exist.
 They are:
@@
   2. 2.
-    Aperiodicity. A Markov Chain is a-periodic if and only if $d(s_{i})=1$ for all states $s_{i}\in S$, where
+    Aperiodicity. A Markov Chain is aperiodic if and only if $d(s_{i})=1$ for all states $s_{i}\in S$, where
     $d(s_{i})=\gcd\left\{t\in\mathbb{N}_{+}\ \ |\ \ P_{ii}^{t}>0\right\},$
@@
 Reversible Markov Chains
-A interesting special case of Markov chains is reversible Markov Chain.
+An interesting special case of Markov chains is the reversible Markov chain.
 Definition 3.2.
@@
 ${\boldsymbol{\pi}}_{i}{P}_{ij}={\boldsymbol{\pi}}_{j}{P}_{ji}.$
-A Markov chain is said to be reversible if there exists some distribution with respect to which the Markov chain that is reversible in the above definition.
+A Markov chain is said to be reversible if there exists some distribution with respect to which the Markov chain is reversible.
 In exercise 3.6 we prove that a distribution that is reversible for a Markov chain is also stationary for this Markov chain.
@@
     $G_{t}=\sum_{i=0}^{\infty}R_{t+i}\gamma^{i}.$
-When the discount factor that is strictly less than one, $G_{t}$ is well defined.
+When the discount factor is strictly less than one, $G_{t}$ is well defined.
 In the case $\gamma=1$, an alternative definition of $G$ is presented below, assuming that a stationary distribution for the underlying Markov chain exists, denoted by $\boldsymbol{\pi}$.
@@
 $\mathbf{v}(s)=\mathbb{E}[G_{t}|S_{t}=s].$
-The value function tells how valuable a state as measure by the sum of its future, discounted rewards.
+The value function tells how valuable a state is as measured by the sum of its future, discounted rewards.
 One way to compute the value function is by using the Bellman equation for the Markov Reward Process, which is the following recursive equation:
@@
   4. 4.
-    $R$ is a reward function that gives each transition and and action a reward $R_{s,s^{\prime},a}$ which is a reward for moving from $s$ to $s^{\prime}$ using action $a$.
+    $R$ is a reward function that gives each transition and action a reward $R_{s,s^{\prime},a}$ which is a reward for moving from $s$ to $s^{\prime}$ using action $a$.
 We henceforth assume w.l.o.g.
@@
 All optimal policies achieve the same optimal value $\mathbf{v}^{*}(s)$ as well as $Q^{*}(s,a)$, and this holds for every state and action.
-The fact that there exist an optimal policy that is optimal regardless of the starting state in a MDP is a consequence of the Markovian nature of the transition dynamics.
+The fact that there exists an optimal policy that is optimal regardless of the starting state in a MDP is a consequence of the Markovian nature of the transition dynamics.
 3.3 The Bellman Equation
@@
 In later chapters we also make use of the Bellman equation to derive the stochastic-optimal LQR controller.
-The Bellman equation for any policy $\pi$ relates the value function for that policy to itself via one step of application for that policy. Mathematically, this amounts to the following
+The Bellman equation for any policy $\pi$ relates the value function for that policy to itself via one step of application for that policy. Mathematically, this amounts to the following:
 $\mathbf{v}_{\pi}(s)=\sum_{a\in A}\pi(a|s)\left[R_{sa}+\gamma\sum_{s^{\prime}\in S}P_{ss^{\prime}}^{a}\mathbf{v}_{\pi}(s^{\prime})\right].$
@@
     $P^{\pi}$ is the induced transition matrix when playing policy $\pi$, i.e.
-    $P^{\pi}_{ss^{\prime}}=\sum_{a}P^{a}_{ss^{\prime}}\cdot\pi(a|s).$
+    $P^{\pi}_{ss^{\prime}}=\sum_{a\in A}P^{a}_{ss^{\prime}}\cdot\pi(a|s).$
 The Bellman equation gains prominance when applied to the optimal policy of a Markov Decision Process.
@@
 In this section we study a fundamental algorithm for solving Markov Decision Processes that is a more direct application of dynamic programming and the Bellman equation.
-The value iteration algorithm start from an arbitrary solution and improves it until a certain measure of optimality is reached.
+The value iteration algorithm starts from an arbitrary solution and improves it until a certain measure of optimality is reached.
 A formal description is given in Algorithm 2.
-The performance guarantee of Value Iteration, as well as its analysis, is particularly elegant and do not require expertise in mathematical optimization theory.
+The performance guarantee of Value Iteration, as well as its analysis, is particularly elegant and does not require expertise in mathematical optimization theory.
 Algorithm 2 Value Iteration
@@
 3.6 Exercises
-3.1. In this exercise we give a guided proof of the ergodic theorem. Consider a given irreducible and a-periodic Markov chain with transition matrix $P$.
+3.1. In this exercise we give a guided proof of the ergodic theorem. Consider a given irreducible and aperiodic Markov chain with transition matrix $P$.
 A. Show that $P$ has a left eigenvalue equal to 1.
@@
 C.
-Consider an abirtrary distribution $\mathbf{p}_{0}$, and the distribution after $t$ iterations of the Markov chain $\mathbf{p}_{t}=P^{t}\mathbf{p}_{0}$.
+Consider an abirtrary distribution $\mathbf{p}_{0}$, and the distribution after $t$ iterations of the Markov chain $\mathbf{p}_{t}=\mathbf{p}_{0}P^{t}$.
 Prove that $\lim_{t\to\infty}\mathbf{p}_{t}=\mathbf{v}_{1}.$
@@
 3.3. Consider the derivation of an optimal policy for a Markov Decision Process given its value function as described in algorithm 2. Prove that the resulting policy is indeed optimal.
-3.4. Let Markov chain $P$ and distribution ${\boldsymbol{\pi}}$ be reversible for it, i.e.
+3.4. Let Markov chain $P$ be reversible with respect to stationary distribution ${\boldsymbol{\pi}}$, i.e.
 $\forall i,j\ ,\ {\boldsymbol{\pi}}_{i}P_{ij}={\boldsymbol{\pi}}_{j}P_{ji}.$
@@
 This motivates the restriction to a special class of dynamical systems for which we can design efficient algorithms with provable guarantees.
-Our treatment in this chapter is restricted to linear dynamical systems (LDS): dynamical systems that evolves linearly according to the equation
+Our treatment in this chapter is restricted to linear dynamical systems (LDS): dynamical systems that evolve linearly according to the equation
 $\mathbf{x}_{t+1}={A}_{t}\mathbf{x}_{t}+{B}_{t}{\mathbf{u}}_{t}+{\mathbf{w}}_{t}.$
@@
 We have seen examples of LTI systems in chapter 2, namely the double integrator and the aircraft dynamics.
-Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories.
+Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories. Source: Wikipedia.
 A more general class of linear dynamical systems is when the matrices ${A}_{t}$ and ${B}_{t}$ change over time.
 These are called linear time-varying (LTV) systems.
-The class of LTV systems can be used to express much more general dynamics via linearization, as we show next.
-4.1 General Dynamics as LTV Systems
+The class of LTV systems can express more general dynamics via linearization, as we show next.
+4.1 Approximating General Dynamics via LTV Systems
 Consider any dynamics that can be expressed by a smooth function $f$.
 Such a system can be approximated by a time-varying linear system to an arbitrary accuracy determined by the time step interval between samples.
-To see this, first consider the one dimensional first-order Taylor expansion of a given dynamics function $f$ around some anchor point $(\bar{x}_{t},\bar{u}_{t})$:
+To see this, first consider the one dimensional first-order Taylor expansion of a given dynamics function $f$, which is twice differentiable, around an anchor point $(\bar{x}_{t},\bar{u}_{t})$:
 $x_{t+1}$ $=f(x_{t},u_{t})+w_{t}$
-$\approx f(\bar{x}_{t},\bar{u}_{t})+\underbrace{\frac{\partial}{\partial x}f(x,u)\bigg{|}_{\bar{x}_{t},\bar{u}_{t}}}_{\triangleq A_{t}}(x_{t}-\bar{x}_{t})+\underbrace{\frac{\partial}{\partial u}f(x,u)\bigg{|}_{\bar{x}_{t},\bar{u}_{t}}}_{\triangleq B_{t}}(u_{t}-\bar{u}_{t})+w_{t}.$
-With an appropriate redefinition of $w_{t}$, this simplifies to
-$x_{t+1}=A_{t}x_{t}+B_{t}u_{t}+w_{t},$
-where An analogous approximation holds for the higher-dimensional case. The Taylor expansion is
-$f(\mathbf{x},\mathbf{u})\approx f(\bar{\mathbf{x}},\bar{\mathbf{u}})+J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{x}-\bar{\mathbf{x}})+J_{\mathbf{u}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{u}-\bar{\mathbf{u}}),$
+$=f(\bar{x}_{t},\bar{u}_{t})+\underbrace{\frac{\partial}{\partial x}f(x,u)\bigg{|}_{\bar{x}_{t},\bar{u}_{t}}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}A_{t}}(x_{t}-\bar{x}_{t})+\underbrace{\frac{\partial}{\partial u}f(x,u)\bigg{|}_{\bar{x}_{t},\bar{u}_{t}}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}B_{t}}(u_{t}-\bar{u}_{t})+w_{t},$
+where $w_{t}$ is the approximation error, or remainder, which by the mean value theorem behaves as $w_{t}\sim(x_{t}-\bar{x}_{t})^{2}+(u-\bar{u}_{t})^{2}$. Thus, this simplifies to
+$x_{t+1}=A_{t}x_{t}+B_{t}u_{t}+w_{t}.$
+An analogous approximation holds for the higher-dimensional case. The Taylor expansion is
+$f(\mathbf{x},\mathbf{u})=f(\bar{\mathbf{x}},\bar{\mathbf{u}})+J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{x}-\bar{\mathbf{x}})+J_{\mathbf{u}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{u}-\bar{\mathbf{u}})+w_{t},$
 where $J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is the Jacobian of $f$ with respect to the variables $\bar{\mathbf{x}}$. Writing the matrices ${A}_{t},{B}_{t}$ for the Jacobians, we have
@@
 $\mathbf{x}_{t+1}={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t},$
-although much of our discussion can be generalized to LTV systems. Exercise 4.6 asks the reader to prove this.
+although much of our discussion can be generalized to LTV systems.
 Recall that a system is stabilizable if from every starting state, there is a sequence of controls that drives the system to zero.
@@
 Then, ${M}\succ{0}$ if and only if ${A}\succ{0}$ and ${C}-{B}^{\top}{A}^{-1}{B}\succ{0}$.
-From Theorem 4.3, we know that stabilizability of a linear system is equivalent to the existence of a matrices ${K}$ and ${P}$ such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)^{\top}P(A+BK)}$.
+Note that since the spectral radius purely involves the eigenvalues of a matrix, it is invariant under transposition, and it holds that $\rho((A+BK)^{\top})=\rho(A+BL)$.
+Now using this fact and Theorem 4.3, we know that stabilizability of a linear system is equivalent to the existence of a matrices ${K}$ and ${P}$ such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)P(A+BK)^{\top}}$.
 The above lemma allows us to reformulate this cubic constraint to a quadratic constraint.
-The stabilizability of a system is equivalent to exists of matrices ${K}$ and ${P}$ such that
+The stabilizability of a system is equivalent to existence of matrices ${K}$ and ${P}$ such that
 $\begin{bmatrix}{P}&{((A+BK)P)^{\top}}\\ {(A+BK)P}&{P}\end{bmatrix}\succ{0},\qquad{P}\succeq{I}.$
@@
 These finite time bounds naturally depend on how difficult it is for a given system to stabilize.
-Thereforem we require a quantitative definition for stabilizability, one that holds for both time-invariant and time-varying linear dynamical systems.
+Therefore, we require a quantitative definition for stabilizability, one that holds for both time-invariant and time-varying linear dynamical systems.
 4.4.1 Stabilizability of Time-Invariant LDS
@@
 Consider a linear controller ${K}$ for the LTI system $({A,B})$.
-If there exists $\kappa>0$ and $\delta\in[0,1)$ such that the matrix ${A+BK}$ admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
+If there exists $\kappa>0$ and $\delta\in(0,1]$ such that the matrix ${A+BK}$ admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
 $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t},\qquad\frac{\kappa}{\delta}\leq\gamma$
@@
 Thus, such a system may be said to be barely controllable.
-This difficulty of controllability is captured in the definition below, which requires not only that $K_{k}(A,B)$ be full row-rank, but also that such a statement is robust to small changes in $A,B$.
+This difficulty of controllability is captured in the definition below, which requires not only that $K(A,B)$ be full row-rank, but also that such a statement is robust to small changes in $A,B$.
 Definition 4.10.
 A time-invariant linear dynamical system $(A,B)$ is $(k,\kappa)$-strongly controllable if and only if
-$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\text{ and }\|({K}_{k}(A,B))^{+}\|\leq\kappa.$
+$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\text{ and }\|({K}_{k}(A,B))^{-1}\|\leq\kappa.$
+In the definition above, the spectral norm of the inverse of $K(A,B)$ is bounded by $\kappa$.
+For example, in the simple one-dimensional system we presented above, the matrix $K(A,B)$ is $\varepsilon$, and thus $\kappa=\frac{1}{\varepsilon}$, which indeed captures the complexity of controlling the system.
 4.5 Bibliographic Remarks
@@
 B.
-For a time-invariant linear system, prove that there exists $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$, $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}$.
+For a time-invariant BIBO-stable linear system, prove that there exists $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$, $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}$.
 In particular, this proves that the conditions listed in Lemma 4.8 are necessary, in addition to being sufficient.
-C. Construct a time-varying linear systems that is BIBO stable, yet does not have the exponential decay properties required in Lemma 4.9.
+C. Construct a time-varying linear system that is BIBO stable, yet does not have the exponential decay properties required in Lemma 4.9.
 4.6. Prove that the bound of one on the operator norm of the matrix $A$ is sufficient but not necessary condition for stabilizability.
 Chapter 5 Optimal Control of Linear Dynamical Systems
-In the chapters thus far we have studied more basic notions than control, namely solution concepts of dynamical systems such as stabilizability and controllability.
+In the chapters thus far, we have studied more basic notions than control, namely solution concepts of dynamical systems such as stabilizability and controllability.
 Even for these more basic problems, we have seen that they are computationally intractable in general.
@@
 The resulting efficient and practical algorithm is one of the most widely used subroutines in control.
-5.1 The Linear-Quadratic Regulator
-The linear quadratic regulator (LQR) problem is that of optimal control of a linear dynamical system with known and fixed quadratic costs. Formally,
+5.1 The Linear Quadratic Regulator
+The linear quadratic regulator (LQR) problem is that of optimal control of a linear dynamical system with known and fixed quadratic costs.
+For simplicity, we mostly consider the time-invariant version, where the linear dynamical system is fixed, and the costs are also fixed quadratic, as given formally below:
 Definition 5.1.
-The Linear Quadratic Regulator (LQR) input is a linear dynamical system that evolves according to
+The (time-invariant) Linear Quadratic Regulator (LQR) input is a linear dynamical system that evolves according to
 $\mathbf{x}_{t+1}={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t},$
@@
   5. 5.
-    $Q\in{\mathbb{R}}^{d_{x}\times d_{x}}$, ${R}\in{\mathbb{R}}^{d_{u}\times d_{u}}$ are positive semi-definite matrices. Thus, the cost functions are convex.
+    $Q\in{\mathbb{R}}^{d_{x}\times d_{x}}$, ${R}\in{\mathbb{R}}^{d_{u}\times d_{u}}$ are positive definite matrices. Thus, the cost functions are convex.
 This framework is very useful since many physical problems, such as the centrifugal governor and the ventilator control problem, can be approximated by a linear dynamical system.
@@
 The LQR is a special case of a Markov decision process which we have described in chapter 3, since the transitions are stochastic.
-However, naive application of value iteration, policy iteration, or linear programming do not apply, since there are infinitely many states and actions.
+However, we cannot naive apply value iteration, policy iteration, or linear programming since there are infinitely many states and actions.
 We proceed to describe a particularly elegant solution to the LQR problem that makes use of its special structure.
@@
 a linear mapping from ${\mathbb{R}}^{d_{x}}$ to ${\mathbb{R}}^{d_{u}}$!
-This astonishing fact together with a remarkable property of the value function is formally given in the following theorem.
+This astonishing fact, together with a remarkable property of the value function, is formally given in the following theorem.
 We state the theorem for a finite number of iterations, and for a linear time-invariant system.
@@
 $\mathbf{v}^{*}_{T}(\mathbf{x})=\min_{\mathbf{u}}\big{\{}\mathbf{x}^{\intercal}Q\mathbf{x}+\mathbf{u}^{\intercal}R\mathbf{u}\big{\}}.$
-Since ${R}\succcurlyeq 0$, a minimum occurs at $\mathbf{u}=0$. Hence, ${S}_{T}={Q}\succcurlyeq{0}$, ${K}_{T}={0}$ and $c_{T}=0$.
+Since ${R}\succ 0$, a minimum occurs at $\mathbf{u}=0$. Hence, ${S}_{T}={Q}\succ{0}$, ${K}_{T}={0}$ and $c_{T}=0$.
 Inductive Step: Assume for some $t>1$ that $\mathbf{u}_{t}^{*}(\mathbf{x})={K}_{t}\mathbf{x}$ and $\mathbf{v}_{t}^{*}(\mathbf{x})=\mathbf{x}^{\intercal}{S}_{t}\mathbf{x}+c_{t}$. Then we have
@@
 We henceforth denote the second moment of the noise by $\mathop{\mbox{\bf E}}[\mathbf{w}\mathbf{w}^{\top}]=\sigma^{2}I$.
+This implies that for any matrix $M$ which is independent of $\mathbf{w}$, we have $\mathop{\mbox{\bf E}}[\mathbf{w}^{\top}M\mathbf{w}]=\mathop{\mbox{\bf E}}[M\mathbf{w}\mathbf{w}^{\top}]=\sigma^{2}\cdot M\bullet I=\sigma^{2}{\bf Tr}(M)$.
 Since $R$ and $S_{t}$ are positive semi-definite, the expression above is convex as a function of $\mathbf{u}$. At the global minimum the gradient vanishes, and hence
@@
 ${\mathbf{u}_{t-1}(\mathbf{x})=K_{t-1}\mathbf{x},\ \ \ K_{t-1}=-(R+B^{\intercal}S_{t}B)^{-1}(B^{\intercal}S_{t}A)}.$
-If the matrices above are not full rank, we use the notation $M^{-1}$ for the Moore-Penrose pseudo inverse. Plugging this in the expression for the value function, we get
+Notice that the matrix $R+B^{\top}S_{t}B$ is full rank since $R\succ 0$. Plugging this in the expression for the value function, we get
 $\mathbf{v}_{t-1}^{*}(\mathbf{x})$
@@
 $=\mathbf{x}^{\intercal}{S}_{t-1}\mathbf{x}+c_{t-1}$
-where ${S_{t-1}=Q+K_{t-1}^{\intercal}RK_{t-1}+(A-BK_{t-1})^{\intercal}S_{t}(A-BK_{t-1})}$ is positive semi-definite, and $c_{t-1}=c_{t}+\sigma^{2}{\bf Tr}(S_{t})$.
+where ${S_{t-1}=Q+K_{t-1}^{\intercal}RK_{t-1}+(A-BK_{t-1})^{\intercal}S_{t}(A-BK_{t-1})}$ is positive definite, and $c_{t-1}=c_{t}+\sigma^{2}{\bf Tr}(S_{t})$.
 5.3 Infinite Horizon LQR
@@
 Solving the DARE therefore corresponds to finding a non-negative root of a quadratic polynomial in the one-dimensional case.
+The reason we only consider the non-negative root is since the value function is non-negative by our choice of quadratic convex costs.
 In higher dimension, the correct generalization is somewhat more difficult to see, but using Lemma 4.5, the equation can be reformulated as the solution to the following SDP.
@@
 is a strong one, and numerous attempts have been made over the years to generalize the noise model.
-One of the most important and prominent generalizations is that of robust control, also known as $H_{\infty}$-control.
-The goal of H-control is to come up with a more robust solution, by considering the optimal solution vs.
+One of the most important and prominent generalizations is that of robust control, also known as $\text{H}_{\infty}$-control.
+The goal of $\text{H}_{\infty}$-control is to come up with a more robust solution, by considering the optimal solution vs.
 any possible perturbation sequence from a certain family.
@@
 The first player tries to minimize the costs and the other attempts to maximize the costs, both players subject to the dynamics evolution constraint.
-Using techniques from dynamic games, the following can be proven, see e.g [ZDG96b, BB08].
+Using techniques from dynamic games, the following can be shown. References are provided in the bibliographic section.
 Theorem 5.4.
-For the $H_{\infty}$ control problem with quadratic convex loss functions, the optimal solution is a linear policy of the state given by
+For the $\text{H}_{\infty}$ control problem with quadratic convex loss functions, the optimal solution is a linear policy of the state given by
 $\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}.$
@@
 This suggests that robustness specifications constitute yet an another axis of performance for control algorithms, and may at times be at odds with the objective of optimal control.
-For in-depth treatment of $H_{\infty}$ control see [ZDG96b].
-The text [BB08] presents a succinct derivation of robust controllers entirely in the state space, as the solution to a dynamic game between the controller and an adversary.
+For in-depth treatment of $H_{\infty}$ control, see [ZDG96b].
+The text [BB08] presents a succinct derivation of robust controllers entirely in the state space, as the solution to a dynamic game between the controller and an adversary, see also [Tu99] for a recent exposition.
 Although developed in the linear setting, adaptations of the LQR controller are also widely used to control non-linear systems [Li05, LK13].
@@
 5.2. Derive the solution to the infinite horizon LQR with zero mean Gaussian noise.
-Part II Basics of Nonstochastic Control
-Chapter 6 Policy classes for dynamical systems
-In chapter 5, we studied the LQR problem, and saw that for linear dynamical systems, the optimal control takes a very special form: it is a linear function of the state.
+5.3. Extend the proof of Theorem 5.2 to positive semi-definite $Q,R$, rather than positive definite matrices.
+5.4.
+Prove Theorem 5.4 as a dynamic two-player zero-sum game, where a unique saddle point solution exists.
+Hint: you can formulate this problem as a min-max optimization problem within a finite horizon.
+5.5. Prove that the function $f(\mathbf{x})=\mathbf{x}^{\top}M\mathbf{x}$ is convex if and only if $M\succeq 0$ is positive semi-definite.
+Part II Basics of Online Control
+Chapter 6 Regret in Control
+In the first part of this book and chapter 5 we studied linear dynamical systems with quadratic costs, namely the LQR problem.
+We saw that in this case the optimal control takes a very special form: it is a linear function of the state.
 The Markovian property of the Markov Decision Process framework for reinforcement learning suggests that the optimal policy is always a function of the current state.
@@
   3. 3.
-    The transition mapping changes is not fixed and changes online.
+    The transition mapping is not fixed and changes online.
   4. 4.
@@
 To this we add a significant consideration: computation of optimal state-based policies may be computationally intractable.
-This motivates the current chapter which studies more general policies.
-We consider several options for control policies, as well as the relationships between them, in the context of continuous control.
-A control policy can take action according to several signals that may be available and observable as listed below.
-  1. 1.
-    State Action Control policies.
-    The system state, denoted $\mathbf{x}_{t}$, if it is observable.
-Policies that map the state to action are called State-Feedback or State-Action policies, and are the most commonly considered in the literature.
-As we have seen in chapter 5, the optimal policy for LDS with stochastic noise is a linear state-feedback policy.
-    More generally, a State-Action policy may depend on several previous states in a linear or non-linear manner.
-The most commonly considered policy class is a linear combination of previous states, but any (non-linear) mapping of the form
-    $\mathbf{u}_{t}=\pi(\mathbf{x}_{t-1:t-h})=\pi(\mathbf{x}_{t-1},...,\mathbf{x}_{t-h})\ ,\ \pi\in{\mathbb{R}}^{d_{\mathbf{x}}\times h}\mapsto{\mathbb{R}}^{d_{\mathbf{u}}},$
-    can be considered. The parameter $h>0$ is the number of most recent states that the policy considers when computing the next action.
-  2. 2.
-    Disturbance Action Control policies.
-    If the perturbation to the system, denoted $\mathbf{w}_{t}$, is known, Disturbance Action Control policies (DAC) choose an action based upon the current and potentially previous disturbances,
-    $\mathbf{u}_{t}=\pi(\mathbf{w}_{t-1:t-h})\ ,\ \pi\in{\mathbb{R}}^{d_{\mathbf{w}}\times h}\mapsto{\mathbb{R}}^{d_{\mathbf{u}}}.$
-    For time-invariant linear dynamical systems, the two policies classes, linear DAC and linear SAC, are equivalent, since the states and disturbances are linearly related under DAC and SAC policies.
-We give a formal theorem that quantifies this relationship later in this chapter.
-    However, for certain scenarios, the policy class of DAC is significantly better behaved than state-action control policies from a computational viewpoint.
-We shall explore efficient algorithms that learn optimal policies from this class in the next chapter.
-  3. 3.
-    Observation Action Control policies.
-    In the case of dynamical systems that are partially observed and/or unknown, both policy classes above are infeasible, since the state and disturbance are similarly unknown.
-A natural policy parameterization is linear in the past observations.
-A special case of such policies is the class of Linear Dynamical Controllers, as stated below.
-  4. 4.
-    Linear Dynamical Control policies.
-    A particularly useful policy class for partially observed linear dynamical systems is Linear Dynamical Controllers.
-These policies simulate an “internal linear dynamical system” to estimate the hidden state, and choose the action at any time step as a linear function of this estimate.
-    A linear dynamic controller $\pi$ has parameters $\left(A_{\pi},{B}_{\pi},C_{\pi},{D}_{\pi}\right)$, and chooses a control at time $t$ according to
-    $\mathbf{u}_{t}^{\pi}=C_{\pi}\mathbf{s}_{t}+D_{\pi}\mathbf{x}_{t}\ ,\ \mathbf{s}_{t+1}=A_{\pi}\mathbf{s}_{t}+B_{\pi}\mathbf{x}_{t}.$
-  5. 5.
-    Disturbance Response Control policies.
-    Another important class of policies for control with partial observability is known as the disturbance response controllers, or DRC.
-The motivation for this control policy class is computational: in later chapters we show efficient algorithms that learn policies within this class.
-    The main signal used by the DRC policy class is called Nature’s y’s, denoted by $\mathbf{y}^{\mathbf{nat}}_{t}$.
-This signal is the would-be observation at time $t$ assuming all inputs to the system were zero from the beginning of time.
-Mathematically, DRC policies choose an action as a function of previous Nature’s y’s signals, i.e.,
-    $\mathbf{u}_{t}=\pi(\mathbf{y}^{\mathbf{nat}}_{t:t-h})\ ,\ \pi:{\mathbb{R}}^{d_{\mathbf{y}}\times(h+1)}\mapsto{\mathbb{R}}^{d_{\mathbf{u}}}.$
-6.1 Relating the power of policy classes
-We motivated the consideration of nonstandard policy classes by the need to deviate from the classical reinforcement learning model of Markov Decision Processes.
-Another important consideration is computational; some policy classes we consider will yield efficient algorithms, whereas in standard representation none are known.
-To be able to give rigorous statements on the performance of methods that optimize these new policy classes, it is crucial to be able to relate their representation power.
-Furthermore, as we consider adversarial perturbations, we shall compare policies in terms of how they perform in hindsight, rather than using any absolute metric of optimality.
+This motivates the current chapter which considers regret analysis as a computational tractable alternative to optimal or robust control.
+We start by considering regret in the decision making framework of online convex optimization.
+We then consider several classes of control policies, as well as the relationships between them, in the context of regret analysis.
+6.1 Online Convex Optimization
+In this section we give a brief introduction to the theory of Online Convex Optimization. The reader is referred to the bibliographic section for a comprehensive survey.
+In the decision making framework of Online Convex Optimization, a player iteratively chooses a point in a convex set $\mathcal{K}\subseteq{\mathbb{R}}^{n}$, and suffers a loss according to an adversarially chosen loss function $f_{t}:\mathcal{K}\mapsto{\mathbb{R}}$.
+Let ${\mathcal{A}}$ be an algorithm for OCO, which maps a certain game history to a decision in the decision set:
+$\mathbf{x}_{t}^{\mathcal{A}}={\mathcal{A}}(f_{1},...,f_{t-1})\in\mathcal{K}.$
+We define the regret of ${\mathcal{A}}$ after $T$ iterations as:
+$\mathrm{{regret}}_{T}({\mathcal{A}})=\sup_{\{f_{1},...,f_{T}\}\subseteq{\mathcal{F}}}\left\{\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t}^{\mathcal{A}})-\min_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})\right\}.$ (6.1)
+If the algorithm is clear from the context, we henceforth omit the superscript and denote the algorithm’s decision at time $t$ simply as $\mathbf{x}_{t}$.
+Intuitively, an algorithm performs well if its regret is sublinear as a function of $T$ (i.e.
+$\mathrm{{regret}}_{T}({\mathcal{A}})=o(T)$), since this implies that on average, the algorithm performs as well as the best fixed strategy in hindsight.
+We proceed to describe a basic algorithm for online convex optimization.
+6.1.1 Online Gradient Descent
+Perhaps the simplest algorithm that applies to the most general setting of online convex optimization is Online Gradient Descent (OGD). Pseudo-code is given in Algorithm 3.
+Algorithm 3 Online Gradient Descent
+1:  Input: convex set $\mathcal{K}$, $T$, $\mathbf{x}_{1}\in\mathcal{K}$, step sizes $\{\eta_{t}\}$
+2:  for $t=1$ to $T$ do
+3:     Play $\mathbf{x}_{t}$ and observe cost $f_{t}(\mathbf{x}_{t})$.
+4:     Update and project:
+$\mathbf{y}_{t+1}=\mathbf{x}_{t}-\eta_{t}\nabla f_{t}(\mathbf{x}_{t})$
+$\mathbf{x}_{t+1}=\operatorname*{\arg\min}_{\mathbf{x}\in\mathcal{K}}\|\mathbf{x}-\mathbf{y}_{t+1}\|^{2}=\prod_{\mathcal{K}}[\mathbf{y}_{t+1}]$
+5:  end for
+In each iteration, the algorithm takes a step from the previous point in the direction of the gradient of the previous cost.
+This step may result in a point outside of the underlying convex set.
+In such cases, the algorithm projects the point back to the convex set, i.e.
+finds its closest point in the convex set.
+Despite the fact that the next cost function may be completely different than the costs observed thus far, the regret attained by the algorithm is sublinear.
+This is formalized in the following theorem.
+Here we define the constants $D$ and $G$ as follows:
+  • $D$ = Euclidean diameter of the decision set, i.e.
+    $\max_{\mathbf{x},\mathbf{y}\in\mathcal{K}}\|\mathbf{x}-\mathbf{y}\|.$
+  • $G$ = an upper bound on the Lipschitz constant of the cost functions $f_{t}$, or equivalently, the subgradient norms of these functions. Mathematically,
+    $\forall t\ ,\ \|\nabla f_{t}(\mathbf{x}_{t})\|\leq G.$
+Theorem 6.1.
+Online gradient descent with step sizes $\{\eta_{t}=\frac{D}{G\sqrt{t}},\ t\in[T]\}$ guarantees the following for all $T\geq 1$:
+$\mathrm{{regret}}_{T}=\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\min_{\mathbf{x}^{\star}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x}^{\star})\ \leq\frac{3}{2}{GD}\sqrt{T}.$
+Proof.
+Let $\mathbf{x}^{\star}\in\operatorname*{\arg\min}_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})$.
+Define $\nabla_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\nabla f_{t}(\mathbf{x}_{t})$.
+By convexity
+$f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{x}^{\star})\leq\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$ (6.2)
+We first upper-bound $\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$ using the update rule for $\mathbf{x}_{t+1}$ and the Pythagorean Theorem, we have the projections only decrease the distance to the set $\mathcal{K}$:
+$\|\mathbf{y}_{t+1}-\mathbf{x}^{\star}\|^{2}\ =\ \left\|\mathop{\Pi}_{\mathcal{K}}(\mathbf{x}_{t}-\eta_{t}\nabla_{t})-\mathbf{x}^{\star}\right\|^{2}\leq\left\|\mathbf{x}_{t}-\eta_{t}\nabla_{t}-\mathbf{x}^{\star}\right\|^{2}.$ (6.3)
+Hence,
+$\|\mathbf{y}_{t+1}-\mathbf{x}^{\star}\|^{2}$$\leq$ $\ \|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}+\eta_{t}^{2}\|\nabla_{t}\|^{2}-2\eta_{t}\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$
+$2\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$$\leq$ $\ \frac{\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$ (6.4)
+Summing (6.2) and (6.1.1) from $t=1$ to $T$, and setting $\eta_{t}=\frac{D}{G\sqrt{t}}$ (with $\frac{1}{\eta_{0}}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}0$):
+$2\left(\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{x}^{\star})\right)\leq 2\sum_{t=1}^{T}\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$
+$\leq\sum_{t=1}^{T}\frac{\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|^{2}}{\eta_{t}}+G^{2}\sum_{t=1}^{T}\eta_{t}$
+$\leq\sum_{t=1}^{T}\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)+G^{2}\sum_{t=1}^{T}\eta_{t}$ $\frac{1}{\eta_{0}}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}0,$
+$\|\mathbf{x}_{T+1}-\mathbf{x}^{*}\|^{2}\geq 0$
+$\leq D^{2}\sum_{t=1}^{T}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)+G^{2}\sum_{t=1}^{T}\eta_{t}$
+$\leq D^{2}\frac{1}{\eta_{T}}+G^{2}\sum_{t=1}^{T}\eta_{t}$ telescoping series
+$\leq 3DG\sqrt{T}.$
+The last inequality follows since $\eta_{t}=\frac{D}{G\sqrt{t}}$ and $\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\leq 2\sqrt{T}$. ∎
+The Online Gradient Descent algorithm is straightforward to implement, and updates take linear time given the gradient.
+However, the projection step may in general take significantly longer via convex optimization.
+6.2 Regret for Control
+The most important consideration to explore beyond classical models of control, such as optimal and robust control, is computational.
+Even in linear control with convex loss functions, the optimal policy can be complex to describe, see exercises section.
+A prominent approach to deal with computational difficulty is to consider a less stringent criteria for optimality.
+Instead of the optimal policy, we can consider all policies in a certain policy class, and try to compete with the optimal policy within the class in terms of regret.
+As alluded to earlier, the most natural classes of policies are computationally intractable.
+We thus resort to learning new classes that are computationally favorable.
+But to relate the performance of our methods, we require a language of comparing different policy classes.
+It is crucial to be able to relate the representation power of different policy classes.
 This motivates our main comparison metric as follows.
 Definition 6.2.
-We say that a class of policies $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ if the following clause is satisfied: for every $\pi_{2}\in\Pi_{2}$, there exists $\pi_{1}\in\Pi_{1}$ such that, simultaneously for all sequences of $T$ disturbances and cost functions, it holds that
+We say that a class of policies $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ if the following is satisfied: for every $\pi_{2}\in\Pi_{2}$, there exists $\pi_{1}\in\Pi_{1}$ such that for all sequences of $T$ disturbances and cost functions, it holds that
 $\sum_{t=1}^{T}\left|c_{t}(\mathbf{x}_{t}^{\pi_{1}},\mathbf{u}_{t}^{\pi_{1}})-c_{t}(\mathbf{x}_{t}^{\pi_{2}},\mathbf{u}_{t}^{\pi_{2}})\right|\leq T\varepsilon.$
@@
 $\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{2})\leq\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{1})+\varepsilon T.$
-The proof of this lemma is left as an exercise to the reader (Exercise 6.4). We now proceed to study the approximation relationships between the policy classes we surveyed earlier.
-6.2 Quantitative Comparison of Policy Classes for Time Varying Linear Dynamical Systems
-In this section we formally relate the power of different complexity classes for the special case of time-varying linear dynamical systems.
-Recall from Chapter 4 that time varying linear dynamical systems were been proposed by Lyapunov as a general methodology to study nonlinear dynamics, and thus are of broad interest.
+Proof.
+Recall that for policy class $\Pi$ and algorithm $\mathcal{A}$,
+$\mathrm{{regret}}_{T}(\mathcal{A},\Pi)=\max_{\mathbf{w}_{1:T}:\|\mathbf{w}_{t}\|\leq 1}\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right)$
+It follows from definition 6.2 that
+$\min_{\pi_{2}\in\Pi_{2}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{2}},\mathbf{u}_{t}^{\pi_{2}})$ $=\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{2}^{*}},\mathbf{u}_{t}^{\pi_{2}^{*}})$ $\pi_{2}^{*}=\operatorname*{\arg\min}_{\pi\in\Pi_{2}}\sum_{t}c_{t}(\mathbf{x}^{\pi}_{t},\mathbf{u}_{t}^{\pi})$
+$\geq\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{1}},\mathbf{u}_{t}^{\pi_{1}})-\varepsilon T$ Definition 6.2
+$\geq\min_{\pi_{1}\in\Pi_{1}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{1}},\mathbf{u}_{t}^{\pi_{1}})-\varepsilon T$
+Thus, for any fixed disturbance sequence $\mathbf{w}_{1:T}$,
+$\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi_{2}\in\Pi_{2}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{2}},\mathbf{u}_{t}^{\pi_{2}})$
+$\leq\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\left(\min_{\pi_{1}\in\Pi_{1}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{1}},\mathbf{u}_{t}^{\pi_{1}})-\varepsilon T\right)$
+$=\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi_{1}\in\Pi_{1}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{1}},\mathbf{u}_{t}^{\pi_{1}})\right)+\varepsilon T$
+Since this holds for any fixed disturbance sequence $\mathbf{w}_{1:T}$, it also holds for the maximum over all disturbance sequences $\mathbf{w}_{1:T}$ with $||\mathbf{w}_{t}||\leq 1$. ∎
+We now proceed to define and study approximation relationships between different policy classes.
+6.3 Expressivity of Control Policy Classes
+In this section we formally relate the power of different complexity classes for linear time-invariant linear dynamical systems.
+The definitions and results can be extended to time-varying linear dynamical systems, and recall from Chapter 4 that time varying linear dynamical systems were been proposed by Lyapunov as a general methodology to study nonlinear dynamics.
 For the rest of this section we assume full observation of the state and let
-$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
+$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}.$
 A sequence of stabilizing linear controllers for such a system is denoted by $\{K_{t}\}$.
-Schematically, the relationships we describe below are given in Figure 6.1 for fully observed systems, and in Figure 8.1 for partially observed systems.
-Figure 6.1: Schematic relationship between linear Disturbance-Action Control policies and linear State-Action Control policies of increasing diameter for linear time invariant systems.
+Schematically, the relationships we describe below are given in Figure 6.1 for fully observed systems.
+In later chapters of this text we study policies for partially observed systems.
+Figure 6.1: Schematic relationship between linear Disturbance-Action Control policies and linear state-action policies of increasing diameter for linear time invariant systems.
 6.3.1 Linear Policies
-The most elementary and well-studied policy classes for control of linear dynamical systems are linear policies, i.e.
+The most basic and well-studied policy classes for control of linear dynamical systems are linear policies, i.e.
 those that choose a control signal as a (fixed) linear function of the state.
@@
 Definition 6.4 (Linear Policies).
-A linear controller $K\in{\mathbb{R}}^{d_{u}\times d_{x}}$ is a linear operator mapping state to control.
-We denote by $\Pi_{\gamma}^{LIN}$ the set of all linear controllers that are $\gamma$-stabilizing and are bounded in Fronbenius norm by $\|K\|\leq\gamma$.
-6.2.2 Linear Dynamic Controllers
-The next policy class is particularly useful for linear dynamical systems with partial observation.
-Linear Dynamical Controllers simulate an “internal linear dynamical system” so to recover a hidden state, and play a linear function over this hidden state.
-The formal definition is given below.
-Definition 6.4 (Linear Dynamic Controllers).
-A linear dynamic controller $\pi$ has parameters $\left(A_{\pi},{B}_{\pi},C_{\pi}\right)$, and chooses a control at time $t$ as
-$\mathbf{u}_{t}^{\pi}=C_{\pi}\mathbf{s}_{t}\ ,\ \mathbf{s}_{t}=A_{\pi}\mathbf{s}_{t-1}+B_{\pi}\mathbf{x}_{t}.$
-Clearly the class of LDCs approximates linear policies without any error, since we can take $A_{\pi},B_{\pi},C_{\pi}$ to be zero, and obtain a purely linear policy.
-We denote by $\Pi^{LDC}_{\gamma}$ the class of all LDC that are $\gamma$-stabilizing as per Definition 2.4 and Exercise 4.6.
-This implies the existence of parameters $\frac{\kappa}{\delta}\leq\gamma$ that satisfy
-$\forall i\ ,\ \|A_{\pi}^{i}\|\leq\kappa(1-\delta)^{i}\ ,\ (\|B_{\pi}\|+\|C_{\pi}\|)\leq\kappa.$
-LDCs are an important policy class for control of linear dynamical systems since they perform favorably in presence of partial observability.
-Furthermore, the optimal control law for latent LDS with Gaussian disturbances and quadratic loss functions is contained in the class of LDCs.
-6.2.3 Generalized Linear Controllers
-We present a generalization of linear controllers below. This generalization class is at least as expressive as LDCs.
-Definition 6.5 (Generalized Linear Controllers).
-A generalized linear controller $\pi$ has parameters $M_{0:h}$, and chooses a control at time $t$ according to
-$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}M_{i}\mathbf{x}_{t-i}.$
-An important concept for generalized linear controllers is that of lifting, or creating a larger linear system that subsumes the GLC. Given a certain GLC with memory $h$, consider the system given by
-$\mathbf{z}_{t+1}=\tilde{A}_{t}\mathbf{z}_{t}+\tilde{B}_{t}\mathbf{u}_{t}+\tilde{C}_{t}\mathbf{w}_{t},$
-where
-$\tilde{A}_{t}=\begin{array}[]{|c|c|c|c|c|}\hline\cr A_{t}&0&\ldots&0&0\\ \hline\cr I&0&\ldots&0&0\\ \hline\cr 0&I&\ldots&0&0\\ \hline\cr\vdots&\vdots&\cdots&\vdots&\vdots\\ \hline\cr 0&0&\cdots&I&0\\ \hline\cr\end{array}\ \ ,\ \ \tilde{B}_{t}=\begin{array}[]{|c|}\hline\cr B_{t}\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}\ \ ,\ \ \tilde{C_{t}}=\begin{array}[]{|c|}\hline\cr I\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}.$
-It can be seen that a GLC for the original system is simply a linear controller $\mathbf{u}_{t}^{\pi}=[M_{0}\dots M_{h}]\mathbf{z}_{t}$ for the lifted system!
-This observation will be useful to us later on.
-We denote by $\Pi^{GLC}_{h,\gamma}$ the class of all GLC with $\sum_{i=0}^{h}\|M_{i}\|\leq\gamma$, and $\mathbf{M}_{1:h}$ that is $\gamma$-stabilizing as per Definition 2.4.
-The following lemma (proof left as as exercise) shows that GLC approximate LDC (and of course linear controllers):
-Lemma 6.6.
-The class $\Pi^{GLC}_{h,\gamma}$ $\varepsilon$-approximates the class $\Pi^{LDC}_{\gamma^{\prime}}$ for $h=\Omega(\gamma^{\prime}\log(\frac{\gamma^{\prime}}{\varepsilon}))$, and $\gamma=2\gamma^{\prime 3}$.
-The class of GLC is thus very general and intuitive. It is, however, difficult to work with for the regret metric, due to its counterfactual structure.
+A linear controller $K\in{\mathbb{R}}^{d_{x}\times d_{u}}$ is a linear operator mapping state to control.
+We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers that are $\gamma$-stabilizing according to definition 2.4 and are bounded in Fronbenius norm by $\|K\|\leq\gamma$.
 6.3.2 Disturbance Action Controllers
-We next define the most important set of policies in nonstochastic control, known as the disturbance action controllers, or DAC.
-The main advantages of DAC as opposed to GLC, LDC and linear controllers are that the cost of control is convex in their parametrization, and thus allows for provably efficient optimization.
-Moreover, this set of policies is stronger than GLC in the sense that it $\varepsilon$-approximates them, as we shortly show.
+We next define the most important set of policies in online nonstochastic control, known as the disturbance action controllers, or DAC.
+The main advantages of DAC as opposed to linear controllers, are that the cost of control is convex in their parametrization, and thus allows for provably efficient optimization.
+Moreover, this set of policies is stronger than linear policies in the sense that it $\varepsilon$-approximates them, as we show below.
 Definition 6.5 (Disturbance Action Controller).
-A disturbance action policy $\pi_{K_{1:T},M_{1:h}}$ is parameterized by the matrices $M_{0:h}=[M_{0},{M}_{1},\ldots,{M}_{h}]$ and stabilizing controllers $\{K_{t}\}$.
+A (linear) disturbance action policy $\pi_{K_{1:T},M_{1:h}}$ is parameterized by the matrices $M_{0:h}=[M_{0},{M}_{1},\ldots,{M}_{h}]$ and stabilizing controllers $\{K_{t}\}$.
 It outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
@@
 This will allow a constant shift in the control.
-Denote by $\Pi^{DAC}_{h,\gamma}$ the set of all disturbance action policies with $K_{1:T}$ being $\gamma$-stabilizing linear controllers for the system and
+Denote by $\Pi^{D}_{h,\gamma}$ the set of all disturbance action policies with $K_{1:T}$ being $\gamma$-stabilizing linear controllers for the system and
 $\sum_{i=1}^{h}\|M_{i}\|\leq\gamma.$
@@
 Lemma 6.6.
-Any DAC policy in $\Pi_{h,\gamma}^{DAC}$ is $\gamma^{2}$-stabilizing.
+Any DAC policy in $\Pi_{h,\gamma}^{D}$ is $\gamma^{2}$-stabilizing.
 Proof.
@@
 for $\|M\|$ being the spectral norm of $M$.
-Next, since $K_{t}$’s are stabilizing, we can bound the historically distant terms as below. Observe that the state at time $t$, following any DAC policy $\pi$, is given by
-$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}$
-$=A_{t}\mathbf{x}_{t}+B_{t}(K_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i})+\mathbf{w}_{t}$
-$=(A_{t}+B_{t}K_{t})\mathbf{x}_{t}+B_{t}\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$ using $\text{using }M_{0}=I$
-$=\sum_{j=0}^{t-1}\left(\left[\prod_{\tau=t-j}^{t}(A_{\tau}+B_{\tau}K_{\tau})\right]B_{t-j}\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i-j}\right)$
-$=\sum_{j=0}^{t-1}\left(\left[\prod_{\tau=t-j}^{t}(A_{\tau}+B_{\tau}K_{\tau})\right]B_{t-j}\tilde{\mathbf{w}}_{t-j}\right).$
+Next, since $K_{t}$’s are stabilizing, we can bound the historically distant terms as below.
+For simplicity assume $K_{t}=K$ are all equal.
+Observe that the state at time $t$, following any DAC policy $\pi$, is given by
+$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}$
+$=A\mathbf{x}_{t}+B(K\mathbf{x}_{t}+\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i})+\mathbf{w}_{t}$
+$=(A+BK)\mathbf{x}_{t}+B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$ using $\text{using }M_{0}=I$
+$=\sum_{j=0}^{t}\left((A+BK)^{j}B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i-j}\right)$
+$=\sum_{j=0}^{t-1}\left((A+BK)^{j}B\tilde{\mathbf{w}}_{t-j}\right).$
 We can thus view the states $x_{t}$ as generated by a $\gamma$-stabilizing policy over noises that are of magnitude $\gamma$, and hence bounded by $\gamma^{2}$. ∎
 The above lemma is already a good indication about DAC: they generate controls that at least don’t cause a blow up. But how expressive are these policies in terms of reaching the optimal solution?
-One of the main appeals of the DAC class is the following approximation lemma which shows that for time-invariant systems DAC are in fact a stronger class than GLC, and thus LDC.
-This is a strong guarantee, since it implies that competing with the best policy in DAC implies competing with the best GLC/LDC.
-We start by showing that DAC approximate all linear controllers.
-This is not a vacuous statement, since the $K_{t}$ terms for DAC is now only assumed to be stabilizing, and not necessarily optimal for the actual sequence.
+One of the main appeals of the DAC class is the following approximation lemma which shows that for time-invariant systems DAC are in fact a stronger class than linear controllers.
+Notice that the $K_{t}$ terms for DAC are only assumed to be stabilizing, and not necessarily optimal for the actual sequence.
 Lemma 6.7.
 Consider a time-invariant linear dynamical system $\{A,B\}$.
-The class $\Pi^{DAC}_{h,\gamma}$ $\varepsilon$-approximates the class $\Pi^{LIN}_{\gamma^{\prime}}$ for $h=\Omega(\gamma^{\prime}\log(\frac{\gamma^{\prime}}{\varepsilon}))$ and $\gamma=\gamma^{\prime 2}$.
+The class $\Pi^{D}_{h,\gamma}$ $\varepsilon$-approximates the class $\Pi^{L}_{\gamma^{\prime}}$ for $h=\Omega(\gamma^{\prime}\log(\frac{\gamma^{\prime}}{\varepsilon}))$ and $\gamma=\gamma^{\prime 2}$.
 Proof.
@@
 $\leq\left\|\sum_{i=0}^{t}A^{i}B\right\|\max_{i\leq t}\|\mathbf{u}_{t}-\tilde{\mathbf{u}}_{i}\|\leq\frac{\kappa}{\delta}\varepsilon,$
-by using the $\gamma$-stability of $(A,B)$. The approximation in terms of states and controls implies $\varepsilon$-approximation of policy class for any Lipchitz cost sequence, concluding the proof.
-Recall that GLC can be lifted to a linear dynamical system. Thus a corollary of the above lemma is as follows, and is left as Exercise 6.4.
-Corollary 6.11.
-Consider a time invariant linear dynamical system given by $\{A,B\}$.
-The class $\Pi^{DAC}_{h,\gamma}$ $\varepsilon$-approximates the class $\Pi^{GLC}_{h^{\prime},\gamma^{\prime}}$, for parameters $h=O(h^{\prime}+\gamma^{\prime}\log\frac{\gamma^{\prime}}{\varepsilon})$, $\gamma=O(\gamma^{\prime 2})$.
+by using the $\gamma$-stability of $(A,B)$. The approximation in terms of states and controls implies $\varepsilon$-approximation of policy class for any Lipschitz cost sequence, concluding the proof.
 6.4 Bibliographic Remarks
+Regret minimization, online learning, and online convex optimization have a rich history and have been studied for decades in the machine learning and game theory literature.
+The reader is referred to the comprehensive texts of [CBL06, Haz16] for a detailed treatment.
 Linear state-feedback policies are a mainstay of control theory both in applications, due to their simplicity, and in theory, owing to their optimality in deterministic and stochastic settings as demonstrated in [Kal60].
@@
 This makes it better suited to machine learning.
+The study of regret in the context of control has roots going back to the work of [AYS11, AYBK14].
+These works, as well as many that followed [CHK+18, AHS19], study regret in a stochastic setting, where the perturbations are stochastic.
+In such settings, optimal control is well defined.
+The full power of regret analysis for control, involving both adversarial perturbations, adversarial loss functions, and computational considerations, was initiated in [ABH+19].
 6.5 Exercises
+6.1. In this exercise we prove that the optimal solution to a linear control problem can be complex to describe.
+  1. 1.
+    Consider a linear dynamical system in one dimension, and loss functions of the form $f_{t}(\mathbf{x})=\max\{0,\mathbf{x}-c\}$, for $c_{t}\in{\mathbb{R}}$.
+Show that the optimal policy in hindsight can have as many domains as loss functions.
+  2. 2.
+    Consider a similar setting in higher dimensions, and give a bound on the number of domains as a function of the number of loss functions and the dimension.
+6.2. Consider the setting of time varying linear dynamical systems, i.e.
+$\mathbf{x}_{t+1}=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
+Prove that any DAC according to definition 6.5 is stable, i.e. prove the analogue of Lemma 6.6 for time varying systems.
 6.3. Prove Lemma 6.8.
-6.2. Prove Lemma 6.6.
-6.3. Prove Lemma 6.10.
-6.4. Prove Corollary 6.11.
 Chapter 7 Online Nonstochatic Control
-In this chapter we consider the main new directions that have been taken in the machine learning literature for the control problem.
-These new developments are a major departure from classical control, and in recent years have produced methods that compete favorably and many times surpass classical methods.
-A few characteristics of the new approaches are listed here:
+In this chapter, we explore recent advancements in machine learning that have significantly reshaped the control problem.
+These developments mark a major departure from classical control theory and, in many cases, have led to methods that not only compete with but often outperform traditional approaches.
+The key characteristics of these modern techniques are as follows:
   1. 1.
-    Instead of pre-computing an optimal policy, online control seeks to learn a policy using online learning techniques. This fits the framework known as “adaptive control”.
+    Online Learning Over Precomputed Policies Unlike classical control, which relies on precomputing an optimal policy, online control adopts adaptive learning techniques to update policies dynamically.
+This aligns with the broader framework of adaptive control, where policies evolve in response to observed system behavior.
   2. 2.
-    In contrast to adaptive control, online control optimizes a rigorous performance metric of regret vs.
-the best policy in hindsight.
-Online and non-stochastic control algorithms have provable finite-time regret guarantees.
+    Regret Minimization as a Performance Metric While adaptive control focuses on learning a stabilizing policy, online control introduces a rigorous performance metric: regret, defined as the difference in cost between the learned policy and the best policy in hindsight.
+Unlike classical approaches, online nonstochastic control algorithms provide finite-time regret guarantees, ensuring competitive performance even under adversarial conditions.
   3. 3.
-    In online nonstochastic control, the loss functions can be general Lipschitz convex losses that are chosen adversarially, as opposed to known quadratic losses in classical control.
+    General Loss Functions and Adversarial Perturbations Classical control typically assumes quadratic cost functions with known structures.
+In contrast, online nonstochastic control allows for general convex loss functions, which may be selected adversarially.
+This greater flexibility enables controllers to adapt to more diverse and complex real-world scenarios.
   4. 4.
-    Online and non-stochastic control deploy improper learning by convex relaxation: this means they compete with a certain policy class using a different policy class.
-This difference in learned policies allows us to side-step the computational hardness results we have previously explored, and obtain rigorous guarantees for efficient algorithms.
+    Improper Learning via Convex Relaxation A fundamental breakthrough in online nonstochastic control is the use of improper learning, where the algorithm competes against a given policy class while learning from a broader class of policies.
+By leveraging convex relaxation techniques, these methods bypass computational hardness barriers that traditionally limit optimal control, allowing for efficient algorithms with strong theoretical guarantees.
+These innovations redefine how we approach control problems, shifting the focus from static policy optimization to adaptive, regret-minimizing strategies that operate effectively in dynamic and adversarial environments.
 7.1 From Optimal and Robust to Online Control
-Recall that the problem of optimal control of a linear time-invariant dynamical system (LDS) is given by the equations
-$\min_{\mathbf{u}(\mathbf{x})}$ $\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$
-subject to $\text{subject to }\mathbf{x}_{t+1}$ $={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t},$
-where $\mathbf{x}_{t}$ is the state of the system, $\mathbf{u}_{t}$ is the control input, and $\mathbf{w}_{t}$ is the perturbation at time $t$.
-Previous chapters showed the optimal control policy to be a linear function of the state when we assumed $\left\{\mathbf{w}_{t}\right\}_{t\in[T]}$ to be i.i.d.
-zero mean and bounded variance perturbations, and $c_{t}$ to be quadratic loss functions.
-Specifically, for positive semi-definite matrices ${Q}$ and ${R}$, the mathematical control formulation becomes
-$\min_{\mathbf{u}(\mathbf{x})}$ $\sum_{t=1}^{T}\mathbf{x}_{t}^{T}{Q}\mathbf{x}_{t}+\mathbf{u}_{t}^{T}{R}\mathbf{u}_{t}$
-subject to $\text{subject to }\mathbf{x}_{t+1}$ $={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t},\quad\mbox{ }\mathbf{w}_{t}\sim\mathcal{D},\mbox{ }\mathbb{E}\mathbf{w}_{t}=\mathbf{0}.$
-We have shown in chapter 5 that the optimal solution to this formulation is given by the policy
-$\mathbf{u}^{\ast}_{t}={K}\mathbf{x}_{t},$
-where ${K}$ depends on ${A}$, ${B}$, ${Q}$ and ${R}$.
-The proof follows by inductively solving the Bellman optimality equation at each step $t$, starting from time step $T$ and moving backwards to time step $1$.
-However, this optimal solution requires two conditions:
+Classical control theory has been primarily shaped by two paradigms: optimal control and robust control.
+These approaches have been instrumental in various engineering applications, but they rely on specific assumptions about disturbances and system dynamics.
+  • Optimal control assumes that disturbances follow a known stochastic model, often Gaussian.
+The objective is to design a control policy that minimizes the expected cumulative cost.
+Notable methods such as the Linear Quadratic Regulator (LQR) and stochastic dynamic programming fall into this category.
+  • Robust control, in contrast, takes a worst-case perspective.
+It ensures stability and performance even under the most adverse conditions, assuming perturbations belong to a predefined uncertainty set.
+A key example is H-infinity control.
+While these frameworks have been successful, real-world control tasks often involve uncertainties that do not fit neatly into either a probabilistic model or a worst-case adversarial setting.
+Consider the challenge of controlling a drone in an unpredictable environment:
+  • A stochastic model of wind disturbances may be unrealistic since actual wind patterns can exhibit long-term correlations or abrupt shifts.
+  • A worst-case robust approach may result in excessively conservative control, leading to inefficient and overly cautious maneuvers.
+These limitations motivate the need for a more adaptive approach—one that can learn and adjust dynamically based on observed disturbances rather than relying on rigid assumptions.
+This leads us to online nonstochastic control, a fundamentally different paradigm where controllers adapt in real-time without assuming a fixed model for disturbances.
+7.1.1 A Shift to Online Nonstochastic Control
+Rather than relying on predefined stochastic or worst-case assumptions, online nonstochastic control introduces a new approach based on sequential decision-making and regret minimization.
+The core principles are:
   1. 1.
-    The loss function needs to be a quadratic function.
+    No Assumption on the Noise Model – Disturbances (e.g., wind forces, external perturbations) are treated as an arbitrary sequence, potentially chosen adversarially.
+The controller must adapt dynamically rather than relying on a fixed distribution.
   2. 2.
-    The perturbations are i.i.d. across time (with additional constraints on the mean and the variance).
-These conditions are in fact necessary for solving the Bellman optimality equation.
-Thus, this approach is not as general as we would like.
-We are missing out on other loss functions commonly used in practice, e.g.
-cross entropy, or loss functions with regularization terms.
-Furthermore and more importantly, many times the perturbations incurred in practice are not i.i.d., in which case we lose all provable guarantees!
-An alternative to optimal control is robust, or $H_{\infty}$ control theory.
-In this formulation, the goal is to optimize the control for an adversarial noise in a minimax sense, as per the mathematical formulation:
-$\min_{\mathbf{u}(\mathbf{x})}\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{1:T}\mathclose{\|}\leq 1}$ $\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$
-subject to $\text{subject to }\mathbf{x}_{t+1}$ $={A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
-This approach also suffers from several downsides:
-  1. 1.
-    The above problem is computationally ill-behaved for non-quadratic loss functions.
-  2. 2.
-    It is non-adaptive in nature, i.e. the system is too pessimistic for cases when the perturbations are well-behaved, like i.i.d. Gaussian perturbations.
-The natural question at this point is: Does there exist an approach that gives the best of both robust and optimal control, and generalizes to any convex cost function?
-That’s exactly non-stochastic control theory which we define next.
-But first, we present a motivating example.
-7.1.1 Motivating Example for Online Control
-Consider the problem of flying a drone from source to destination subject to unknown weather conditions.
-The aerodynamics of flight can be modeled sufficiently well by time-varying linear dynamical systems, and existing techniques are perfectly capable of doing a great job for indoor flight.
-However, the wind conditions, rain and other uncertainties are a different story.
-Certainly the wind is not an i.i.d.
-Gaussian random variable!
-Optimal control theory, which exactly assumes this zero-mean noise, is therefore overly optimistic and inaccurate for our problem.
-The designer might resort to robust control theory to account for all possible wind conditions, but this is overly pessimistic.
-What if we encounter a bright sunny day after all?
-Planning for the worst case would mean slow and fuel-inefficient flight.
-We would like an adaptive control theory which allows us to attain the best of both worlds: an efficient and fast flight when the weather permits, and a careful drone maneuver when this is required by the conditions.
-Can we design a control theory that will allow us to take into account the specific instance perturbations and miss-specifications, and give us finite time provable guarantees?
-This is the subject of online non-stochastic control!
+    Learning and Adaptation Over Time – Instead of precomputing an optimal policy, online control algorithms iteratively refine their strategies using techniques from online convex optimization and sequential decision-making.
+  3. 3.
+    Regret Minimization as the Performance Metric – Unlike optimal control, which minimizes expected cost, or robust control, which prepares for the worst case, online control seeks to minimize regret.
+Regret quantifies how much worse the controller performs compared to the best policy in hindsight.
+To build intuition, consider the problem of controlling an autonomous vehicle in an unknown city.
+If the vehicle had complete foresight of future traffic conditions, it could compute the optimal route in advance.
+However, without such knowledge, the vehicle must continuously adjust its route based on real-time observations.
+Regret measures how far the vehicle’s performance deviates from the best possible path that could have been chosen in hindsight.
+This shift in perspective enables online controllers to handle a wide range of real-world disturbances, achieving adaptability, robustness, and computational efficiency without requiring strong assumptions about uncertainty.
+In the following sections, we formally define the online nonstochastic control problem and introduce key algorithms that provide provable performance guarantees.
 7.2 The Online Nonstochastic Control Problem
-When dealing with non-stochastic, a.k.a.
-arbitrary or even adversarial perturbations, the optimal policy is not a-priory clear.
-Rather, an optimal policy for the observed perturbations can be determined only in hindsight.
-This is the reason that non-stochastic control shifts the performance metric from optimality, to regret vs.
-the best policy in hindsight from a certain reference class.
-We have already defined regret in full generality in Definition 1.3, and in this section give a more detailed treatment that specializes to time varying linear dynmical systems.
-At this point the reader may wonder: Why should we compare to a reference class, as opposed to all possible policies? There are two answers to this question:
-First, the best policy in hindsight may be very complicated to describe and reason about.
-This argument was made by the mathematician Tyrell Rockafellar; see the bibliographic section for more details.
-Secondly, it can be shown that it is impossible, in general, to obtain sublinear regret vs.
-the best policy in hindsight.
-Rather, a different performance metric called competitive ratio can be analyzed.
-This will be surveyed later in the book.
-We now turn to define policy regret with respect to a reference class of policies.
-In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$, and then observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$, according to an adversarially chosen loss function.
-Let $\Pi=\{\pi:\mathbf{u}\mapsto\mathbf{x}\}$ be a class of policies.
-The regret of the controller w.r.t.
-$\Pi$ is defined as follows.
+The shift from classical control to online nonstochastic control requires a fundamental rethinking of how we handle uncertainty.
+Unlike optimal control, which assumes disturbances follow a known probabilistic model, or robust control, which plans for the worst-case, online nonstochastic control operates in a more general and challenging setting.
+In this framework, both the disturbances and cost functions are chosen by an adversary, meaning they may follow no statistical pattern and could even be designed to hinder the controller’s decisions at every step.
+As a result, the goal is no longer to compute a fixed optimal control policy but rather to learn and adapt dynamically in response to an evolving sequence of perturbations and costs.
+A fundamental difficulty in control is the computational tractability of finding the best policy.
+Classical control methods often rely on dynamic programming, which, in general, suffers from the curse of dimensionality.
+Even in simple settings, computing an optimal policy can be intractable due to the complexity of solving constrained optimization problems over trajectories.
+As pointed out by Rockafellar, finding an optimal control policy in arbitrary convex cost settings can be computationally prohibitive, particularly when state and control constraints are involved.
+To address this, online nonstochastic control shifts the focus from computing an optimal policy to competing with a benchmark class of policies using regret minimization.
+Instead of optimizing a predefined objective function, the controller seeks to minimize its total loss relative to the best policy that could have been chosen in hindsight.
+This approach bypasses the need to solve an intractable dynamic optimization problem upfront while still achieving performance guarantees over time.
+The fundamental challenge in online nonstochastic control can be summarized as follows:
+  • The controller does not have prior knowledge of the cost functions or disturbances.
+  • At each time step, the controller must select a control input based only on past observations.
+  • The performance of the controller is evaluated against a benchmark: the best fixed policy that could have been chosen in hindsight.
+This formulation naturally connects to online learning and online convex optimization, where decisions must be made sequentially in uncertain environments.
+We now provide a formal definition of the problem.
 Definition 7.1.
-Let $\{A_{t},B_{t}\}$ be a time varying LDS. The regret of an online control algorithm over $T$ iterations w.r.t. the class of policies $\Pi$ is given by:
+Consider a dynamical system evolving according to the equation: $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),$ where:
+  • $\mathbf{x}_{t}\in\mathbb{R}^{d_{\mathbf{x}}}$ is the system state at time $t$,
+  • $\mathbf{u}_{t}\in\mathbb{R}^{d_{\mathbf{u}}}$ is the control input at time $t$,
+  • $\mathbf{w}_{t}\in\mathbb{R}^{d_{\mathbf{w}}}$ represents an external perturbation, which is chosen adversarially.
+At each time step, the controller ${\mathcal{A}}$ selects $u_{t}$ based on past observations, then incurs a cost defined by an arbitrary convex function $c_{t}:\mathbb{R}^{d_{\mathbf{x}}}\times\mathbb{R}^{d_{\mathbf{u}}}\to\mathbb{R}$.
+The objective is to design a control strategy that minimizes regret, defined as the difference between the total cost incurred by the controller and the best fixed policy in hindsight:
 $\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}$ $\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}))\right),$
-where $\mathbf{u}_{t}={\mathcal{A}}(\mathbf{x}_{t})$ are the controls generated by ${\mathcal{A}}$, and $\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}$ are the counterfactual state sequence and controls under the policy $\pi$, i.e.
-$\mathbf{u}_{t}^{\pi}$ $=\pi(\mathbf{x}_{t}^{\pi})$
-$\mathbf{x}_{t+1}^{\pi}$ $=A_{t}\mathbf{x}_{t}^{\pi}+B_{t}\mathbf{u}_{t}^{\pi}+\mathbf{w}_{t}.$
-Henceforth, if $T,\Pi$ and ${\mathcal{A}}$ are clear from the context, we just refer to $\mathrm{{regret}}$ without quantifiers.
-The non-stochastic control problem can now be stated as the quest to find an efficient algorithm that minimizes the worst-case regret vs.
-meaningful policies classes that we examine henceforth.
-It is important to mention that the algorithm does not have to belong to the comparator set of policies $\Pi$!
-Indeed, in the most powerful results, we will see that the algorithm will learn a policy that is strictly not contained in the comparator class.
+where $\Pi$ is a benchmark class of policies, and $(x_{t}^{\pi},u_{t}^{\pi})$ represent the state and control sequence obtained under the best policy $\pi$ in hindsight.
+Henceforth, if $T,\Pi$ and ${\mathcal{A}}$ are clear from the context, we just refer to $\mathrm{{regret}}$ without quantifiers. The goal of an online control algorithm is to achieve sublinear regret,
+$\mathrm{{regret}}_{T}=o(T),$ (7.1)
+which ensures that, as $T\to\infty$, the controller performs competitively with the best fixed policy.
+Improper Learning and Convex Relaxation
+A key technique that enables online nonstochastic control algorithms to remain computationally efficient is improper learning through convex relaxation.
+Instead of directly competing with a predefined class of policies, these algorithms leverage a different policy class that allows for efficient optimization while still achieving strong performance guarantees.
+This approach is crucial because enforcing exact constraints on the policy space can lead to computational hardness results, as previously discussed.
+By relaxing the policy representation, for example using convex approximations or surrogate loss functions, controllers can sidestep these intractabilities while still competing effectively with the best policy in hindsight.
+In the next section, we introduce algorithmic solutions that utilize this principle to achieve low-regret online control under adversarial conditions.
 7.3 The Gradient Perturbation Controller
-After exploring the relationships between different classes of controllers, we are ready to prove the main new algorithm in non-stochastic control, which gives a strong regret guarantee against adversarially chosen perturbations.
+After motivating the framework, defining regret, and exploring the relationships between different policy classes, we are ready to prove the main new algorithm in non-stochastic control.
 The algorithm is described in figure (4), for time varying linear dynamical systems.
@@
 The main idea is to learn the parametrization of a DAC policy using known online convex optimization algorithms, namely online gradient descent.
-We henceforth prove that this parametrization is essentially convex, allowing us to prove the main regret bound.
-In algorithm (3) we denote by $\mathbf{x}_{t}({M}_{1:h})$ the hypothetical state arrived at by playing a DAC policy ${M}_{1:h}$ from the beginning of time; note that this may not be equal to the true state $\mathbf{x}_{t}$ of the dynamical system at time $t$.
+We henceforth prove that this parametrization is convex, allowing us to prove the main regret bound.
+The notation we use to describe the algorithm is as follows.
+We denote by $\mathbf{x}_{t}({M}_{1:h})$ the hypothetical state arrived at by playing a DAC policy ${M}_{1:h}$ from the beginning of time.
+Note that this may not be equal to the true state $\mathbf{x}_{t}$ of the dynamical system at time $t$.
 The counterfactual $\mathbf{u}_{t}({M}_{1:h})$ is defined similarly.
-1:  Input: $h$, $\eta$, initialization $M_{1:h}^{1}$.
-2:  Observe the linear system $(A_{t},B_{t})$, compute a stabilizing linear controller sequence ${K}_{t}$.
+1:  Input: $h$, $\eta$, initialization $M_{1:h}^{1}\in\mathcal{K}$.
+2:  Observe the linear system $(A_{t},B_{t})$, compute a stabilizing linear controller ${K}_{t}$ as per Theorem 4.4.
 3:  for $t$ = $1\ldots T$ do
@@
 6:     Record an instantaneous cost $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$.
-7:     Construct loss $\ell_{t}({M}_{1:h})=c_{t}(\mathbf{x}_{t}({M}_{1:h}),\mathbf{u}_{t}({M}_{1:h})).$Update ${M}_{1:h}^{t+1}\leftarrow{M}_{1:h}^{t}-\eta\nabla\ell_{t}({M}_{1:h}^{t}).$
+7:     Construct loss $\ell_{t}({M}_{1:h})=c_{t}(\mathbf{x}_{t}({M}_{1:h}),\mathbf{u}_{t}({M}_{1:h})).$Update ${M}_{1:h}^{t+1}\leftarrow\prod_{\mathcal{K}}\left[{M}_{1:h}^{t}-\eta\nabla\ell_{t}({M}_{1:h}^{t})\right].$
 8:  end for
 Algorithm 4 Gradient Perturbation Controller(GPC)
-The GPC algorithm comes with a very strong performance guaranteee: namely it guarantees vanishing worst case regret vs.
-the best DAC policy in hindsight.
-As per the relationship between DAC and other poicies, this implies vanishing regret vs.
-LDC, GLC and linear controllers.
+The GPC algorithm comes with a very strong performance guaranteee: it guarantees vanishing regret vs.
+the best Disturbance Action Control policy belonging to our reference class in hindsight.
+As per the relationship between DAC and other policies, this implies vanishing regret vs.
+linear controllers as well.
 The formal statement is given as follows.
 Theorem 7.2.
-Assuming that
+Assuming that for all $t\in[T]$
   • a
-    The costs $c_{t}$ are convex, bounded, and have bounded gradients w.r.t. both arguments $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$.
+    The costs $c_{t}$ are convex, bounded, and $L$-Lipschitz.
   • b
-    The matrices $\{{A}_{t},{B}_{t}\}$ are stabilizable, and have bounded $\ell_{2}$ norm.
+    The matrices $\{{A}_{t},{B}_{t}\}$ are $\gamma$-stabilizable, and have bounded $\ell_{2}$ norm.
 Then the GPC algorithm (4) ensures that
-$\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DAC}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right)\leq\mathcal{O}(\sqrt{T}).$
+$\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{D}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right)\leq\mathcal{O}(LGD\gamma^{2}h\sqrt{T}),$
+where the $O$ notation hides constants that only depend on the norm of $A_{t},B_{t}$.
 Furthermore, the time complexity of each loop of the algorithm is linear in the number of system parameters and logarithmic in T.
-Proof sketch.
-As per Theorem A.1 in Appendix A, the online gradient descent algorithm for convex loss functions ensures that the regret is $\mathcal{O}(\sqrt{T})$.
-However, to directly apply the OGD algorithm to our setting, we need to make sure that the following 3 conditions hold true:
+Note: we henceforth assume that the linear dynamical systems $(A_{t},B_{t})$ are a linear time-invariant system $(A,B)$ which is $\gamma$-stable, and there is no need for a stabilizing controller $K_{t}=0$.
+The extension to time-varying stabilizable systems is left as exercise 7.5.
+Proof.
+As per Theorem 6.1, the online gradient descent algorithm for convex loss functions ensures that the regret is bounded by $2GD\sqrt{T}$.
+However, to directly apply the OGD algorithm to our setting, we need to ensure that the following conditions hold true:
   • The loss function should be a convex function in the variables ${M}_{1:h}$.
-  • The loss function is time dependent because the loss depends on the current state, which further depends on the previous states.
-The state at each time step depends on the current variable estimates.
-Hence, $\mathbf{x}_{t}$ can be shown to be a function of $\left\{{M}^{i}_{1:h}\right\}_{i=1}^{t}$, where ${M}^{i}_{1:h}$ denotes the variables at step $i$.
-Thus, we need to ensure that optimizing the loss function is similar to optimizing a function which doesn’t have such a dependence on time.
-First, we show that the loss function is convex w.r.t. to the variables ${M}_{1:h}$. This follows since the states and the controls are linear transformations of the variables.
+  • We need to ensure that optimizing the loss function $\ell_{t}$ as a function of the parameters $M_{1:h}$ is similar to optimizing the cost $c_{t}$ as a function of the state and control.
+These facts are formalize in the following two lemmas, that we state here and prove at the end of this section.
 Lemma 7.3.
 The loss functions $\ell_{t}$ are convex in the variables ${M}_{1:h}$.
-Proof.
+Lemma 7.4.
+Let $G$ denote a upper bound on the gradients of the $L$-Lipschitz losses $c_{t}$, and $D$ denote the diameter of the matrices $M_{1:h}$.
+Then there exists a constant $C$ that depends only on $\|A\|,\|B|\|$, such that for all $t$,
+$|l_{t}(M^{t}_{1:h})-c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})|\leq\frac{CL\gamma^{2}hD}{\sqrt{T}}.$
+The regret bound of GPC can be derived from these lemmas as follows.
+Note that $G$ is an upper bound on the gradients of $\ell_{t}$ as a function of the parameters $M_{1:h}$, while $L$ is the Lipschitz constant of $c_{t}$ as a function of the state and control.
+$\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{D}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi})$
+$\leq\sum_{t=1}^{T}l_{t}(M_{1:h}^{t})-\min_{\pi\in\Pi^{DAC}}l_{t}(M_{1:h}^{\pi})+T\times\frac{L\gamma^{2}ChD}{\sqrt{T}}$ Lemma 7.4
+$\leq 2GD\sqrt{T}+{\sqrt{T}L\gamma^{2}ChD}$ Theorem 6.1
+$\leq O(GDL\gamma^{2}h\sqrt{T}).$
+7.3.1 Convexity of the losses
+To conclude the analysis, We first show that the loss functions are convex w.r.t.
+to the variables ${M}_{1:h}$.
+This follows since the states and the controls are linear transformations of the variables.
+Proof of Lemma 7.3.
+For simplicity, assume that the initial state $\mathbf{x_{0}}$ is $\mathbf{0}$.
 Using $\mathbf{x}_{t}$ to denote $\mathbf{x}_{t}(M_{1:h})$ and $\mathbf{u}_{t}$ to denote $\mathbf{u}_{t}(M_{1:h})$, the loss function $\ell_{t}$ is given by
-$\ell_{t}({M}_{1:h})=c_{t}(\mathbf{x}_{t},\mathbf{y}_{t}).$
+$\ell_{t}({M}_{1:h})=c_{t}(\mathbf{x}_{t},\mathbf{u}_{t}).$
 Since the cost $c_{t}$ is a convex function w.r.t. its arguments, we simply need to show that $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{1:h}$. The state is given by
-$\mathbf{x}_{t+1}={A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}$ $={A}_{t}\mathbf{x}_{t}+{B}_{t}\left({K}_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i}\right)+\mathbf{w}_{t}$
-$=\tilde{A}_{t}\mathbf{x}_{t}+\left({B}_{t}\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i}+\mathbf{w}_{t}\right),$
-where $\tilde{A}_{t}=A_{t}+B_{t}K_{t}$. By induction, we can further simplify
-$\mathbf{x}_{t+1}$ $=\sum_{i=0}^{t}\left(\prod_{j=t-i+1}^{t}\tilde{A}_{j}\right)\left(B_{t-i}\sum_{j=1}^{h}{M}_{j}\mathbf{w}_{t-i-j}+\mathbf{w}_{t-i}\right),$ (7.1)
-which is a linear function of the variables. In the above computation, for convenience, we have assumed that the initial state $\mathbf{x_{0}}$ is $\mathbf{0}$.
+$\mathbf{x}_{t+1}={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t}$ $={A}\mathbf{x}_{t}+{B}\left(\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i}\right)+\mathbf{w}_{t}$
+By induction, we can further simplify
+$\mathbf{x}_{t+1}$ $=\sum_{i=0}^{t}{A}^{i}\left(B\sum_{j=1}^{h}{M}_{j}\mathbf{w}_{t-i-j}+\mathbf{w}_{t-i}\right),$ (7.2)
+which is a linear function of the variables.
 Similarly, the control $\mathbf{u}_{t}$ is given by
-$\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i},$
-and since $\mathbf{x}_{t}$ was shown to be linear in the variables, so is $\mathbf{u}_{t}$.
-Thus, we have shown that $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ are linear transformations of ${M}_{1:h}$ and hence, the loss function is convex in ${M}_{1:h}$.
-(Why?
-See Exercise 7.5.) ∎
-Now, by the regret bound implied by OGD, we have
-$\sum_{t=1}^{T}l_{t}(M_{1:h}^{t})-\min_{M_{1:h}}\sum_{t=1}^{T}l_{t}(M_{1:h})\leq\mathcal{O}(\sqrt{T}).$ (7.2)
-Finally, the actual loss $c_{t}$ at time $t$ is not calculated on $\mathbf{x}_{t}(M_{1:h})$, but rather on the true state $\mathbf{x}_{t}$, which in turn depends on different parameters $M_{1:h}^{i}$ for various historical times $i<t$.
+$\mathbf{u}_{t}=\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i}.$
+Thus, we have shown that $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ are linear transformations of ${M}_{1:h}$.
+A composition of convex and linear functions is convex (see Exercise 7.5), which concludes our Lemma.
+∎
+7.3.2 Loss functions with memory
+The actual loss $c_{t}$ at time $t$ is not calculated on $\mathbf{x}_{t}(M_{1:h})$, but rather on the true state $\mathbf{x}_{t}$, which in turn depends on different parameters $M_{1:h}^{i}$ for various historical times $i<t$.
 Nevertheless, $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ is well approximated by $l_{t}(M^{t}_{1:h})$, as shown next.
-Lemma 7.4.
-For all $t$, $|l_{t}(M^{t}_{1:h})-c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})|\leq\mathcal{O}(1/\sqrt{T})$.
-Proof.
-First, note that to achieve $\mathcal{O}(\sqrt{T})$ regret, OGD uses a step size of $\eta=1/\sqrt{T}$. Since the gradients of the loss functions are bounded, we have
-$\|M_{1:h}^{t}-M_{1:h}^{t-i}\|\leq\sum_{s=t-i+1}^{t}\|M_{1:h}^{s}-M_{1:h}^{s-1}\|\leq\mathcal{O}\left(\frac{i}{\sqrt{T}}\right).$
-Now, we use the fact that the iterates of OGD $M_{1:h}^{t}$ move slowly to establish that $\mathbf{x}_{t}$ and $\mathbf{x}_{t}(M^{t}_{1:h})$ are close in value.
+Proof of Lemma 7.4.
+We first note that by the choice of step size by the Online Gradient Descent algorithm 3, as per Theorem 6.1, we have that $\eta=\frac{D}{G\sqrt{T}}$. Thus,
+$\|M_{1:h}^{t}-M_{1:h}^{t-i}\|\leq\sum_{s=t-i+1}^{t}\|M_{1:h}^{s}-M_{1:h}^{s-1}\|\leq i\eta G=\frac{iD}{\sqrt{T}}.$
+We next use this fact to establish that $\mathbf{x}_{t}$ and $\mathbf{x}_{t}(M^{t}_{1:h})$ are close.
 Similar to the expansion of $\mathbf{x}_{t}(M_{1:h})$ in Equation 7.2, $\mathbf{x}_{t}$ can be written as
-$\mathbf{x}_{t+1}$ $=\sum_{i=0}^{t}\left(\prod_{j=t-i+1}^{t}\tilde{A}_{j}\right)\left(B_{t-i}\sum_{j=1}^{h}{M}^{t-i}_{j}\mathbf{w}_{t-i-j}+\mathbf{w}_{t-i}\right).$
-Recall that the stability of the time-varying linear policy $\{K_{t}\}$ implies that for any $i\leq j$, $\|\prod_{t=i}^{j}\tilde{A}_{t}\|\leq\kappa(1-\delta)^{j-i}$.
+$\mathbf{x}_{t+1}$ $=\sum_{i=0}^{t}A^{i}\left(B\sum_{j=1}^{h}{M}^{t-i}_{j}\mathbf{w}_{t-i-j}+\mathbf{w}_{t-i}\right).$
+Recall that the stability of the time-invariant linear dynamical system $\{A,B\}$ implies that for any $i$, $\|A^{i}\|\leq\kappa(1-\delta)^{i}$, with $\frac{\kappa}{\delta}\leq\gamma$.
 Therefore, there exists a constant $C\geq 0$ such that
-$\|\mathbf{x}_{t+1}-\mathbf{x}_{t}(M_{1:h}^{t})\|$ $\leq\sum_{i=0}^{t}\left\|\left(\prod_{j=t-i+1}^{t}\tilde{A}_{j}\right)\right\|\|B_{t-i}\|\sum_{j=1}^{h}\|{M}^{t-i}_{j}-M^{t}_{j}\|\|\mathbf{w}_{t-i-j}\|$
-$\leq Ch\eta\sum_{i=0}^{t}i(1-\gamma)^{i}$
-$\leq\frac{Ch\eta}{\gamma^{2}}=\mathcal{O}\left(\frac{1}{\sqrt{T}}\right).$
-This display also implies that $\|\mathbf{u}_{t}-\mathbf{u}_{t}(M_{1:h}^{t})\|=\mathcal{O}\left({1}/{\sqrt{T}}\right)$, since $\mathbf{u}_{t}$ and $\mathbf{u}_{t}(M_{1:h}^{t})$ only differ in the immediate state they operate on.
+$\|\mathbf{x}_{t}-\mathbf{x}_{t}(M_{1:h}^{t})\|$ $=\left\|\sum_{i=0}^{t}A^{i}B\sum_{j=1}^{h}{M}^{t-i}_{j}\mathbf{w}_{t-i-j}-\sum_{i=0}^{t}A^{i}B\sum_{j=1}^{h}M^{t}_{j}\mathbf{w}_{t-i-j}\right\|$
+$\leq\sum_{i=0}^{t}\left\|A^{i}\right\|\|B\|\sum_{j=1}^{h}\|{M}^{t-i}_{j}-M^{t}_{j}\|\|\mathbf{w}_{t-i-j}\|$
+$\leq C\cdot\kappa\sum_{i=0}^{t}(1-\delta)^{i}\cdot i\frac{hD}{\sqrt{T}}$
+$=\frac{Ch\kappa D}{\sqrt{T}}\sum_{i=1}^{\infty}i(1-\delta)^{i-1}$ $\sum_{i=1}^{\infty}i(1-\delta)^{i}=\frac{1}{\delta^{2}}$
+$\leq\frac{C\gamma^{2}hD}{\sqrt{T}}.$
+The control at time $t$ depends only on the parameters at the current iteration and not on historical parameters, hence $\mathbf{u}_{t}=\mathbf{u}_{t}(M_{1:h}^{t})$.
 By definition, $l_{t}(M_{1:h}^{t})=c_{t}(\mathbf{x}_{t}(M_{1:h}^{t}),\mathbf{u}_{t}(M_{1:h}^{t}))$.
-Finally, we note that $c_{t}$ has bounded derivatives finishing the proof of the lemma.
-∎
-Now, we combine the regret guarantee in equation 7.2 and the immediately preceding lemma as
-$\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DAC}}\sum_{t=1}^{T}c_{t}(\hat{\mathbf{x}}_{t},\pi(\hat{\mathbf{x}}_{t}))$
-$\leq\sum_{t=1}^{T}l_{t}(M_{1:h}^{t})-\min_{\pi\in\Pi^{DAC}}l_{t}(M_{1:h}^{\pi})+T\times\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)=\mathcal{O}(\sqrt{T}),$
-implying the stated claim.
+Thus, we have
+$l_{t}(M_{1:h}^{t})-c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ $=c_{t}(\mathbf{x}_{t}(M_{1:h}^{t}),\mathbf{u}_{t}(M_{1:h}^{t}))-c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$
+$\leq L\cdot\|\mathbf{x}_{t}(M_{1:h}^{t})-\mathbf{x}_{t}\|\leq\frac{L\gamma^{2}ChD}{\sqrt{T}}.$
 7.4 Bibliographic Remarks
@@
 Follow up work by [AHS19] achieves logarithmic pseudo-regret for strongly convex, adversarially selected losses, and well-conditioned stochastic noise.
-Logarithmic regret algorithms were derived for online nonstochastic control in [FS20].
+Logarithmic regret algorithms were derived for online nonstochastic control in [Sim20, FS20].
 Online nonstochastic control with known dynamics was further extended in several directions.
@@
 More refined regret metrics applicable to changing dynamics were studied in [GHM20, MGSH21, ZWZ22, Sin22].
+More extensions suitable for changing environments were studied in [BW22].
 [LDL21] studies algorithms that balance regret minimization while adhering to constraints.
@@
 The regret metric when applied to policies is different from the counterfactual notion of policy regret as defined by [ADT12].
-The roots of online nonstochastic control can be traced to the online stochastic control setting, which we cover in Chapter 10.
 An alternate performance metric for online control, which is also nonstochastic, is competitive ratio.
@@
 [GASH22] show that regret minimizers in the non-stochastic setting automatically achieve a near-optimal competitive ratio.
-A precursor to nonstochastic control is a line of work, beginning with [AYBK14, CHK+18], focused on the control of stochastic systems under sequentially revealed adversarial costs functions.
+A precursor to online nonstochastic control is a line of work, beginning with [AYBK14, CHK+18], focused on the control of stochastic systems under sequentially revealed adversarial costs functions.
 See [CK20, CCK22] for further developments.
@@
 7.1. Let $l:{\mathbb{R}}^{d}\to{\mathbb{R}}$ be a convex function. Prove that for any matrix $A\in{\mathbb{R}}^{d\times n}$, $l(Ax)$ is convex for $x\in{\mathbb{R}}^{n}$.
-Chapter 8 Policy Classes for Partially Observed LDS
-In this chapter we consider linear time varying systems with partial observation, i.e.
-the state is not available to the controller.
-Instead, a (known or unknown) linear projection of the state is revealed.
-In full generality, recall from definition 1.1 partially observed dynamical system is given by the equations
-$\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g_{t}(\mathbf{x}_{t}).$
-As per our notation thus far, here $\mathbf{x}_{t}\in{\mathbb{R}}^{d_{x}}$ is the state of the system, $\mathbf{y}_{t}\in{\mathbb{R}}^{d_{y}}$ is the observation, $\mathbf{u}_{t}\in{\mathbb{R}}^{d_{u}}$ is the control input, and $\mathbf{w}_{t}\in{\mathbb{R}}^{d_{w}}$ is the perturbation.
-The function $g_{t}$ is a mapping from the state to the observed state, and the function $f_{t}$ is the transition function, given current state $\mathbf{x}_{t}$ and control input $\mathbf{u}_{t}$.
-The subscript indicates the relevant quantities at time step $t.$
-Many of the control policies we consider in this chapter are general and can be applied to any dynamical system of the above formulation.
-However, following the paradigm of chapter 6, we relate the representation power of policy classes only for partially observed linear dynamical systems.
-These are given by the equations
-$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}$
-$\mathbf{y}_{t}$ $=C_{t}\mathbf{x}_{t}.$
-Often, partially observed LDSs are defined with an additional noise term $\mathbf{v}_{t}$ corrupting the observations as $\mathbf{y}_{t}=C_{t}\mathbf{x}_{t}+\mathbf{v}_{t}$.
-Exercise 8.5 asks the reader to prove that even in such circumstances we may assume the previously stated form for partially observed systems without loss of any generality.
-In the partial observation setting some of the control policy classes from chapter 6 are not well defined. We proceed to define policy classes and describe the relationship between them.
-Figure 8.1: Schematic relationship between linear Disturbance-Response Control policies and Linear Dynamical Control policies of increasing dimensions (or number of parameters) for linear time invariant systems.
-We limit the discussion in this chapter to stable partially observed systems, although the policy containment results described next hold more broadly.
-8.1 Linear Observation Action Controllers
-Linear Observation-action controllers are the most natural generalization of linear controllers to the partially observed case, and are defined as follows.
-Definition 8.1.
-A linear observation-action controller $K\in{\mathbb{R}}^{d_{u}\times d_{y}}$ is a linear operator mapping observation to control as $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}$.
-We denote by $\Pi_{\gamma}^{OAC}$ the set of all linear controllers bounded in Fronbenius norm by $\|K\|\leq\gamma$, that are also $\gamma$-stabilizing for given LDS.
-While this class of policies is natural to consider, it turns out to be very inexpressive.
-Even for optimal control with stochastic noise, the optimal policy cannot be approximated by a linear observation-action policy (see Exercise 8.5).
-Therefore we proceed to define more expressive and useful policies.
-However, OACs will be useful in simplifying proofs later on.
-8.1.1 Generalized Linear Observation Action Controllers
-Similar to the full observation setting, we can generalize linear observation-action controllers to apply to a history of past observations.
-This additional generality, of taking a history of observations, greatly increases the expressive power of this policy class.
-The formal definition is below.
-Definition 8.2.
-Parameterized by matrices $K_{0:h}$, generalized linear observation-action controllers choose a control input as
-$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}K_{i}\mathbf{y}_{t-i}.$
-Let $\Pi^{GOAC}_{h,\gamma}$ be that class of all such policies satisfying $\sum_{i=0}^{h}\|K_{i}\|\leq\gamma$ that are also $\gamma$-stabilzing.
-Generalized observation-action policies can be seen as a special case of observation-action controllers using the technique of lifting from Section 6.2.3.
-Consider any partially-observed system $(A_{t},B_{t},C_{t})$ and perturbation sequence $\mathbf{w}_{t}$.
-Consider the lifted system $(\tilde{A}_{t},\tilde{B}_{t},\tilde{C}_{t},\tilde{D}_{t})$ given by
-$\mathbf{z}_{t+1}=\tilde{A}_{t}\mathbf{z}_{t}+\tilde{B}_{t}\mathbf{u}_{t}+\tilde{D}_{t}\mathbf{w}_{t},\quad\tilde{\mathbf{y}}_{t}=C_{t}\mathbf{z}_{t},$
-where
-$\tilde{A}_{t}=\begin{array}[]{|c|c|c|c|c|}\hline\cr A_{t}&0&\ldots&0&0\\ \hline\cr I&0&\ldots&0&0\\ \hline\cr 0&I&\ldots&0&0\\ \hline\cr\vdots&\vdots&\cdots&\vdots&\vdots\\ \hline\cr 0&0&\cdots&I&0\\ \hline\cr\end{array}\ \ ,\ \ \tilde{B}_{t}=\begin{array}[]{|c|}\hline\cr B_{t}\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}\ \ ,$
-$\ \ \tilde{C}_{t}=\begin{array}[]{|c|c|c|c|}\hline\cr C_{t}&C_{t-1}&\dots&C_{t-h}\\ \hline\cr\end{array}\ \ ,\ \ \tilde{D_{t}}=\begin{array}[]{|c|}\hline\cr I\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}.$
-The following lemma relates the original and lifted system (for a proof see exercise 8.5).
-Lemma 8.3.
-For any sequence of control inputs $\mathbf{u}_{t}$, it holds that $\tilde{\mathbf{y}}_{t}=[\mathbf{y}_{t},\mathbf{y}_{t-1},\dots\mathbf{y}_{t-h}]$.
-The lemma above shows that generalized linear observation-action controllers are a special case of simple observation-action controllers such that $\mathbf{u}_{t}^{\pi}=[K_{0}\dots K_{h}]\tilde{\mathbf{y}}_{t}$ for the lifted system.
-8.2 Linear Dynamic Controllers
-The main motivation for the class of linear dynamical controllers, which we defined already in the context of full observability, is that of partially observed linear dynamical systems.
-The generalization of this class is as follows.
-Definition 8.4 (Linear Dynamic Controllers for Partially Observed Systems).
-A linear dynamic controller $\pi$ has parameters $\left(A_{\pi},{B}_{\pi},C_{\pi}\right)$, and chooses a control at time $t$ as
-$\mathbf{u}_{t}^{\pi}=C_{\pi}\mathbf{s}_{t}\ ,\ \mathbf{s}_{t}=A_{\pi}\mathbf{s}_{t-1}+B_{\pi}\mathbf{y}_{t}.$
-We denote by $\Pi^{LDC}_{\gamma}$ the class of all LDC that are $\gamma$-stabilizing as per Definition 2.4 and exercise 4.6, with parameters $\frac{\kappa}{\delta}\leq\gamma$ that satisfy
-$\forall i\ ,\ \|A_{\pi}^{i}\|\leq\kappa(1-\delta)^{i}\ ,\ (\|B_{\pi}\|+\|C_{\pi}\|)\leq\kappa.$
-Note that LDCs are $\varepsilon$-approximated by generalized observation-action controllers, namely GOACs.
-The proof is left as Exercise 9.4, on account of bering similar to that for the fully observed case.
-Lemma 8.5.
-Consider a partially-observed time-invariant linear dynamical system $\{A,B,C\}$.
-The class $\Pi^{GOAC}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{LDC}_{\gamma}$ for $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right),\gamma^{\prime}=2\gamma^{3}$.
-8.3 Disturbance Response Controllers
-The most important policy class in the context of nonstochastic control is that of Disturbance Response Controllers, or DRC.
-These policies act on a signal called Nature’s y’s, denoted $\mathbf{y}^{\mathbf{nat}}_{t}$, which is the would-be observation at time $t$ assuming all inputs to the system were zero from the beginning of time, or formally,
-$\mathbf{x}^{\mathbf{nat}}_{t+1}=A_{t}\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}\ ,\ \mathbf{y}^{\mathbf{nat}}_{t}=C_{t}\mathbf{x}_{t}.$
-A linear disturbance response policy $\pi$ is parameterized by the matrices $M_{1:h}=[{M}_{0},\ldots,{M}_{h}]$, and outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
-$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}{M}_{i}\mathbf{y}^{\mathbf{nat}}_{t-i}.$
-Denote by $\Pi^{DRC}_{h,\gamma}$ the set of all disturbance response policies with $K_{1:T}$ being $\gamma$-stabilizing linear controllers for the system and
-$\sum_{i=0}^{h}\|M_{i}\|\leq\gamma.$
-We proceed to show that that the class of DRC approximates the class of LDC.
-8.3.1 The Expressive Power of Disturbance Response Controllers
-A DRC always produces controls that ensure the system remains state-bounded. This can be shown analogously to Lemma 6.8 (see Exercise 8.5).
-While DRCs are contained inside the policy class of DACs, they are still very expressive.
-Note that a DAC policy can not be executed in a partially observed system from observations alone; even the knowledge of the entire observation sequence $\mathbf{y}_{t}$ may not uniquely determine $\mathbf{w}_{t}$.
-We have already seen that LDCs are approximately captured by GOACs in Lemma 8.5.
-It can be shown that the class of DRCs with appropriate parameters $\varepsilon$-approximates the class of GOACs for linear time-invariant systems.
-Note that this observation in conjunction with Lemma 8.5 implies that DRCs also approximate LDCs.
-To demonstrate that DRCs approximately capture GOACs, we instead prove that OACs, a weaker class of policies, are $\varepsilon$-approximated by DRCs.
-Nevertheless, this observation is sufficient to conclude the stronger claim, given that GOACs can be written as OACs on a lifted system as noted in Lemma 8.3, and simultaneously that a DRC with history $H^{\prime}$ on the lifted system can be written as a DRC with history $H^{\prime}+h$ over the original system.
-Lemma 8.6.
-Consider a $\gamma$-stable time-invariant linear dynamical system $(A,B,C)$.
-The class $\Pi^{DRC}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{OAC}_{\gamma}$ for parameters $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right)$, $\gamma^{\prime}=O(\gamma^{6})$.
-Proof.
-Recall that for the fully observed case the state-action pairs produced by execution of a linear policy can be written as linear functions of the perturbations.
-It is this observation that allowed us to prove that DACs capture linear policies for fully observed systems.
-We prove an analogue for partially observed systems: The observation-action pairs produced by the execution of OACs can be written as linear functions of nature’s $y$’s.
-Thereafter, we argue that parameterizing the controllers to be linear in nature’s $y$’s (as DRCs do) allows us to encompass OACs.
-We introduce an augmented partially observed linear system that helps us write the observation-action pairs produced by a OAC as a linear function of nature’s $y$’s.
-Crucially, the new system has the same nature’s $y$’s as the original system.
-Lemma 8.7.
-For any partially-observed system $(A_{t},B_{t},C_{t})$ and perturbation sequence $\mathbf{w}_{t}$ producing observations $\mathbf{y}_{t}$, construct a lifted system $(\tilde{A}_{t},\tilde{B}_{t},\tilde{C}_{t})$ given by
-$\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t+1}\\ \Delta_{t+1}\end{bmatrix}=\begin{bmatrix}A_{t}&0\\ 0&A_{t}\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta_{t}\end{bmatrix}+\begin{bmatrix}0\\ B_{t}\end{bmatrix}\mathbf{u}_{t}+\begin{bmatrix}I\\ 0\end{bmatrix}\mathbf{w}_{t},\quad\tilde{\mathbf{y}}_{t}=\begin{bmatrix}C_{t}&C_{t}\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta_{t}\end{bmatrix}.$
-It holds uniformly for all time $t$ that $\widetilde{\mathbf{y}^{\mathbf{nat}}_{t}}={\mathbf{y}^{\mathbf{nat}}_{t}}$ and further for any sequence of control inputs $\mathbf{u}_{t}$ that $\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}$ and $\mathbf{x}_{t}=\mathbf{x}^{\mathbf{nat}}_{t}+\Delta_{t}$.
-Substituting $\mathbf{u}_{t}^{\pi}=K\mathbf{y}^{\pi}_{t}=KC(\mathbf{x}^{\mathbf{nat}}_{t}+\Delta^{\pi}_{t})$, the state-action sequence produced by a OLIN $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}^{\pi}$ can be written as
-$\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t+1}\\ \Delta^{\pi}_{t+1}\end{bmatrix}=\begin{bmatrix}A&0\\ BKC&A+BKC\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta^{\pi}_{t}\end{bmatrix}+\begin{bmatrix}I\\ 0\end{bmatrix}\mathbf{w}_{t},\quad\begin{bmatrix}\mathbf{y}^{\pi}_{t}\\ \mathbf{u}^{\pi}_{t}\end{bmatrix}=\begin{bmatrix}C&C\\ KC&KC\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta^{\pi}_{t}\end{bmatrix},$
-or using that $\mathbf{y}^{\mathbf{nat}}_{t}=C\mathbf{x}^{\mathbf{nat}}_{t}$, equivalently as
-$\Delta^{\pi}_{t+1}=(A+BKC)\Delta^{\pi}_{t}+BK\mathbf{y}^{\mathbf{nat}}_{t},\quad\mathbf{y}^{\pi}_{t}=C\Delta^{\pi}_{t}+\mathbf{y}^{\mathbf{nat}}_{t},\quad\mathbf{u}^{\pi}_{t}=KC\Delta_{t}+K\mathbf{y}^{\mathbf{nat}}_{t}.$
-Now, define $M_{i}=KC(A+BKC)^{i-1}BK$ for $i>0$ and $M_{0}=K$, let $\mathbf{y}_{t},\mathbf{u}_{t}$ be the observation-action sequence produced by this DRC. We have
-$\mathbf{u}^{\pi}_{t}$ $=K\mathbf{y}^{\mathbf{nat}}_{t}+KC\Delta^{\pi}_{t}$
-$=K\mathbf{y}^{\mathbf{nat}}_{t}+KCBK\mathbf{y}^{\mathbf{nat}}_{t-1}+KC(A+BKC)\Delta^{\pi}_{t-1}$
-$=K\mathbf{y}^{\mathbf{nat}}_{t}+\sum_{i>1}KC(A+BKC)^{i-1}BK\mathbf{y}^{\mathbf{nat}}_{t-i}$
-$=\sum_{i=0}^{h^{\prime}}M_{i}\mathbf{y}^{\mathbf{nat}}_{t-i}+\sum_{i>H^{\prime}}KC(A+BKC)^{i-1}BK\mathbf{y}^{\mathbf{nat}}_{t-i}$
-$=\mathbf{u}_{t}+Z_{t}.$
-Given $\gamma$-stabilizability of the OAC $K$ and $\gamma$-intrinsic stability of the of the system $(A,B,C)$, we have that for some $\kappa,\delta$ suc that $\frac{\kappa}{\delta}\leq\gamma$ that
-$\|Z_{t}\|$ $\leq\sum_{i>h^{\prime}}\|KC(A+BKC)^{i-1}BK\|\|\mathbf{y}^{\mathbf{nat}}_{t-i}\|$
-$\leq\sum_{i>h^{\prime}}\|KC(A+BKC)^{i-1}BK\|\left\|\sum_{j>1}CA^{j-1}B\mathbf{w}_{t-i-j}\right\|$
-$\leq\frac{\kappa^{5}}{\delta}\sum_{i>h^{\prime}}\|(A+BKC)^{i-1}\|\leq\frac{\kappa^{6}\gamma}{\delta}(1-\frac{1}{\delta})^{h^{\prime}}$
-Therefore, $\|\mathbf{u}_{t}-\mathbf{u}^{\pi}_{t}\|\leq\varepsilon$ by the appropriate choice of $H^{\prime}$. Once again using the intrinsic stability of the LTI $(A,B,C)$ , we have
-$\|\mathbf{y}^{\pi}_{t}-\mathbf{y}_{t}\|$ $=\left\|\sum_{i=1}^{t}CA^{i-1}B(\mathbf{u}^{\pi}_{t-i}-{\mathbf{u}}_{t-i})\right\|$
-$\leq\left\|\sum_{i=0}^{t}CA^{i}B\right\|\max_{i\leq t}\|\mathbf{u}^{\pi}_{i}-{\mathbf{u}}_{i}\|\leq\frac{\kappa}{\delta}\varepsilon.$
-The approximation in terms of observations and controls implies $\varepsilon$-approximation of policy class for any Lipchitz cost sequence, concluding the proof. ∎
-8.4 Bibliographic Remarks
-The Disturbance Response Parametrization was introduced in [SSH20] as an analogue of DAC for control of parially observed systems.
-It can be viewed as a succinct state-space representation of the Youla parameterization [YJB76].
-For time-varying dynamics, the power of vaious policy classes was contrasted in [MGSH21].
-8.5 Exercises
-8.1. Consider a seemingly more general definition of a partially observed linear dynamical system, given below:
-$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t},$
-$\mathbf{y}_{t}$ $=C_{t}\mathbf{x}_{t}+\mathbf{v}_{t}.$
-Construct a linear dynamical system parameterized via matrices $(A^{\prime}_{t},B^{\prime}_{t},C^{\prime}_{t})$ and perturbations $\mathbf{w}^{\prime}_{t}$ with a hidden state $\mathbf{z}_{t}$ of the form
-$\mathbf{x}_{t+1}$ $=A^{\prime}_{t}\mathbf{z}_{t}+B^{\prime}_{t}\mathbf{u}_{t}+\mathbf{w}^{\prime}_{t},$
-$\mathbf{y}^{\prime}_{t}$ $=C^{\prime}_{t}\mathbf{z}_{t}$
-such that for any sequence of $\mathbf{u}_{t}$, $\mathbf{y}_{t}=\mathbf{y}^{\prime}_{t}$ holds uniformly over time.
-Additionally, verify that there exists such a construction of $\mathbf{z}_{t}$ where $\mathbf{w}^{\prime}_{t}$ are i.i.d.
-as long as $(\mathbf{w}_{t},\mathbf{v}_{t})$ are i.i.d.
-across time steps.
-Hint: Consider the choice $\mathbf{w}^{\prime}_{t}=D^{\prime}_{t}\begin{bmatrix}\mathbf{w}_{t}\\ \mathbf{v}_{t}\end{bmatrix}$ for some $D^{\prime}_{t}$.
-Also consider using a hidden state $\mathbf{z}_{t}$ with a larger number of coordinates than $\mathbf{x}_{t}$.
-8.2. Prove Lemma 8.3.
-8.3. For any $\gamma$-stable partially observed system, prove that any DRC policy in $\Pi_{H,\gamma}^{DRC}$ is $\gamma^{2}$-stabilizing.
-8.4.
-Construct a partially observed linear dynamical system subject to Gaussian i.i.d.
-perturbations with quadratic costs such that simultaneously it is true that (a) there exists a GOAC policy that incurs zero (or at most constant) aggregate cost, (b) yet any OAC policy incurs $\Omega(T)$ aggregate cost when run for $T$ time steps.
-Hint: Consider a 3-dimensional hidden state so that $u_{t}$ can be chosen as a function of $y_{t-1}$ to ensure zero cost at each time step.
-Chapter 9 Online Nonstochastic Control with Partial Observation
-In this chapter we explore the more general setting of nonstochastic control that allows for partial observation.
-Partial observation introduces significant challenges and requires revisiting the controllability and stabilizability definitions we have discussed in previous chapters.
-To allow for more streamlined and simple exposition we make a significant simplification: we assume that the underlying linear dynamical system is stable.
-The generalization of the results presented in this chapter to stabilizable systems is surveyed in the bibliographic section at the end of this chapter.
-Hence, we consider time varying linear dynamical systems with partial observation governed by the equations  [^1] Recall that this is a slightly simplified form, but without loss of generality, as per Exercise 8.5.
-$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}$
-$\mathbf{y}_{t}$ $=C_{t}\mathbf{x}_{t}.$
-The definition of policy regret 1.3 naturally extends to partial observation as below.
-The setting is similar: the controller iteratively chooses a control $\mathbf{u}_{t}$, and then observes the next observation of the system $\mathbf{y}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$, according to an adversarially chosen loss function.
-The states $\mathbf{x}_{t}$ is hidden from the controller.
-Analogous to policy regret definition 1.3 aa Let $\Pi=\{\pi:\mathbf{y}_{1:t}\mapsto\mathbf{u}_{t}\}$ be a class of policies. The regret of the controller w.r.t. $\Pi$ is defined as follows.
-Definition 9.1.
-Let $\{A_{t},B_{t},C_{t}\}$ be a time varying, partially observable, linear dynamical system. The regret of an online control algorithm over $T$ iterations w.r.t. class of policies $\Pi$ is given by:
-$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\max_{\xi_{1:T},\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|},\|\xi_{t}\|\leq 1}$ $\left(\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t}^{\pi},\mathbf{u}_{t}^{\pi}))\right),$
-where $\mathbf{u}_{t}={\mathcal{A}}(\mathbf{y}_{t})$ are the controls generated by ${\mathcal{A}}$, and $\mathbf{y}_{t}^{\pi},\mathbf{u}_{t}^{\pi}$ are the counterfactual state sequence and controls under the policy $\pi$, i.e.
-$\mathbf{u}_{t}^{\pi}$ $=\pi(\mathbf{y}_{1:t}^{\pi})$
-$\mathbf{x}_{t+1}^{\pi}$ $=A_{t}\mathbf{x}_{t}^{\pi}+B_{t}\mathbf{u}_{t}^{\pi}+\mathbf{w}_{t}$
-$\mathbf{y}_{t}^{\pi}$ $=C_{t}\mathbf{x}_{t}^{\pi}.$
-Henceforth, if $T$ and $\Pi,{\mathcal{A}}$ are clear from the context, we just refer to $\mathrm{{regret}}$ without quantifiers.
-9.1 Disturbance Response Controllers
-Recall from section 8 the policy class of disturbance response controllers.
-This is the natural analogue of disturbance action controllers from the full information setting, adapted to use a different signal called “Nature’s $\mathbf{y}$’s”.
-Recall the simplification that we make in this chapter: the linear dynamical system is stable.
-This allows us to ignore the stabilizing controller component of disturbance action control and greatly simplifies the exposition of the main ideas.
-We thus can define in succinct notation,
-Definition 9.2 (Disturbance Response Controller).
-A disturbance response policy for stable systems $\pi_{M_{1:h}}$ is parameterized by the matrices $M_{1:h}=[{M}_{1},\ldots,{M}_{h}]$.
-It outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
-$\mathbf{u}_{t}^{\pi}=\sum_{i=1}^{h}{M}_{i}\mathbf{y}^{\mathbf{nat}}_{t-i},$
-where $\mathbf{y}^{\mathbf{nat}}_{t}$ is given by the system response to the zero control at time $t$, i.e.
-$\mathbf{x}^{\mathbf{nat}}_{t+1}$ $=A_{t}\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}=\sum_{i=0}^{t}\left[\prod_{j=0}^{i-1}A_{t-j}\right]\mathbf{w}_{t-i}=\Phi_{t}\mathbf{w}_{1:t}.$
-$\mathbf{y}^{\mathbf{nat}}_{t+1}$ $=C_{t+1}\mathbf{x}^{\mathbf{nat}}_{t+1}=C_{t+1}\Phi_{t}\mathbf{w}_{1:t}.$
-Here $\Phi_{t}$ is the linear operator such that $\Phi_{t}(t-i)=\prod_{j=0}^{i-1}A_{t-j}$.
-If the perturbations are all zero, seemingly DRC policies produce a zero control. Therefore we could add a constant offset to the control, that can be incorporated as in Definition 6.7.
-The main advantage of this policy class, as opposed to those considered in optimal control such as linear dynamical controllers, is that they yield convex parametrization.
-Thus, this policy class allows for provably efficient optimization and algorithms.
-We conclude with the following important lemma on an expression that is important to compute nature’s y’s. The proof of this lemma follows directly from the definition and is left as exercise 9.4.
-Lemma 9.3.
-The sequence $\{\mathbf{y}^{\mathbf{nat}}_{t}\}$ can be iteratively computed using only the sequence of observations as
-$\mathbf{y}^{\mathbf{nat}}_{t}=\mathbf{y}_{t}-C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}A_{t-j}\right]B_{t-i}\mathbf{u}_{t-i}$
-Alternatively, the above computation can be carried out recursively as stated below starting with $\mathbf{z}_{0}=\mathbf{0}$.
-$\mathbf{z}_{t+1}$ $=A_{t}\mathbf{z}_{t}+B_{t}\mathbf{u}_{t},$
-$\mathbf{y}^{\mathbf{nat}}_{t}$ $=\mathbf{y}_{t}-C_{t}\mathbf{z}_{t}.$
-9.2 The Gradient Response Controller
-Partial observation requires modification to the Gradient Perturbation Controller we have studied in the previous chapter.
-The Gradient Response Controller algorithm is described in figure (4), for general changing linear dynamical systems with partial observation.
-Similar to the GPC, the main idea is to learn the parametrization of a disturbance response control policy using known online convex optimization algorithms, such as online gradient descent.
-We henceforth prove that this parametrization is essentially convex, allowing us to prove the main regret bound.
-In algorithm (4) we denote $\mathbf{y}_{t}({M}_{1:h})$ as the observation arising by playing DRC policy ${M}_{1:h}$ from the beginning of time, and same for $\mathbf{u}_{t}({M}_{1:h})$.
-Although it may be impossible to determine $\mathbf{w}_{t}$, especially if $C_{t}$ has fewer rows than columns, from the observations $\mathbf{y}_{t}$, $\mathbf{y}_{t}({M}_{1:h})$ can be determined from $\mathbf{y}^{\mathbf{nat}}_{t}$ alone by the application of Lemma 9.3, as
-$\mathbf{y}_{t}({M}_{1:h})=\mathbf{y}^{\mathbf{nat}}_{t}+C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}A_{t-j}\right]B_{t-i}\mathbf{u}_{t-i}({M}_{1:h}).$
-1:  Input: sequence $\{A_{t},B_{t},C_{t}\}$ , $h$, $\eta$, initialization $M_{0:h}^{1}$.
-2:  for $t$ = $1\ldots T$ do
-3:     Use Control $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=0}^{h}{M}_{i}^{t}\mathbf{y}^{\mathbf{nat}}_{t-i}$
-4:     Observe $\mathbf{y}_{t+1}$, compute for $t+1$:
-$\mathbf{y}^{\mathbf{nat}}_{t}=\mathbf{y}_{t}-C_{t}\sum_{i=0}^{t}\left[\prod_{j=0}^{i}A_{t-j}\right]B_{t-i}\mathbf{u}_{t-i}.$
-5:     Construct loss $\ell_{t}({M}_{0:h})=c_{t}(\mathbf{y}_{t}({M}_{0:h}),\mathbf{u}_{t}({M}_{0:h}))$Update $M_{0:h}^{t+1}\leftarrow{M}_{0:h}^{t}-\eta\nabla\ell_{t}({M}_{0:h}^{t})$
-6:  end for
-Algorithm 4 Gradient Response Controller(GRC)
-The GRC algorithm comes with a finite time performance guarantee: namely it guarantees vanishing worst case regret vs.
-the best disturbance response control policy in hindsight.
-As per the relationship between disturbance response control and other policies studied in section 8.3.1, this implies vanishing regret vs.
-linear dynamical controllers.
-The formal statement is given as follows.
-Theorem 9.4.
-Assuming that
-  • a
-    The costs $c_{t}$ are convex, bounded and have bounded gradients w.r.t. both arguments $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$.
-  • b
-    The matrices $\{{A}_{t},{B}_{t},C_{t}\}$ have bounded $\ell_{2}$ norm.
-  • c
-    The LDS $\{{A}_{t},{B}_{t},C_{t}\}$ is $\gamma$-stable.
-Then the GRC algorithm (4) ensures that
-$\max_{\xi_{1:T},\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|},\|\xi_{t}\|\leq 1}\left(\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DRC}}\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right)\leq\mathcal{O}(\sqrt{T}).$
-Furthermore, the time complexity of each loop of the algorithm is polynomial in the number of system parameters and logarithmic in T.
-The proof is largely analogous to the fully observed case. We present a sketch below.
-Proof sketch.
-As per Theorem A.1 in Appendix A, the online gradient descent algorithm for convex loss functions ensures that the regret is $\mathcal{O}(\sqrt{T})$.
-However, to directly apply the OGD algorithm to our setting, we need to make sure that the following 3 conditions hold true:
-  • The loss function should be a convex function in the variables ${M}_{0:h}$.
-  • The loss function is time dependent because the loss depends on the current observation/state, which further depends on the previous states.
-The state at each time step depends on the current variable estimates.
-Hence, $\mathbf{y}_{t}$ can be shown to be a function of $\left\{{M}^{i}_{0:h}\right\}_{i=1}^{t}$, where ${M}^{i}_{0:h}$ denotes the variables at step $i$.
-Thus, we need to make sure that optimizing the loss function is similar to optimizing a function which doesn’t have such a dependence on time.
-First, we show that the loss function is convex w.r.t. to the variables ${M}_{0:h}$. This follows since the states and the controls are linear transformations of the variables.
-Lemma 9.5.
-The loss functions $\ell_{t}$ are convex in the variables ${M}_{0:h}$.
-Proof.
-Using $\mathbf{x}_{t}$ to denote $\mathbf{x}_{t}(M_{0:h})$ (and similarly for $\mathbf{y}_{t},\mathbf{u}_{t}$), the loss function $\ell_{t}$ is given by
-$\ell_{t}({M}_{0:h})=c_{t}(\mathbf{x}_{t},\mathbf{y}_{t}).$
-Since the cost $c_{t}$ is a convex function w.r.t.
-its arguments, we simply need to show that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{0:h}$.
-Using Lemma 9.3, the action-observation sequence is given by
-$\mathbf{u}_{t}$ $=\sum_{i=0}^{h}M_{i}\mathbf{y}^{\mathbf{nat}}_{t-i},$
-$\mathbf{y}_{t}$ $=\mathbf{y}^{\mathbf{nat}}_{t}+C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}{A}_{t-j}\right]B_{t-i}\mathbf{u}_{t-i}$
-$=\mathbf{y}^{\mathbf{nat}}_{t}+C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}{A}_{t-j}\right]B_{t-i}\sum_{j=0}^{h}M_{j}\mathbf{y}^{\mathbf{nat}}_{t-i-j}$
-which is a linear function of the variables.
-Thus, we have shown that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ are linear transformations in ${M}_{0:h}$ and hence, the loss function is convex in ${M}_{0:h}$.
-∎
-Now, by the regret bound implied by OGD, we have
-$\sum_{t=1}^{T}l_{t}(M_{0:h}^{t})-\min_{M_{0:h}}\sum_{t=1}^{T}l_{t}(M_{0:h})\leq\mathcal{O}(\sqrt{T}).$ (9.1)
-Finally, the actual loss $c_{t}$ at time $t$ is not calculated on $\mathbf{y}_{t}(M_{0:h})$, but rather on the true observation $\mathbf{y}_{t}$, which in turn depends on different parameters $M_{0:h}^{i}$ for various historical times $i<t$.
-Nevertheless, $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$ is well approximated by $l_{t}(M^{t}_{0:h})$.
-Lemma 9.6.
-For all $t$, $|l_{t}(M^{t}_{0:h})-c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})|\leq\mathcal{O}(1/\sqrt{T})$.
-Proof.
-First, note that to achieve $\mathcal{O}(\sqrt{T})$ regret OGD uses a step size of $\eta=1/\sqrt{T}$. Since the gradient of the loss functions are bounded, we have
-$\|M_{0:h}^{t}-M_{0:h}^{t-i}\|\leq\sum_{s=t-i+1}^{t}\|M_{0:h}^{s}-M_{0:h}^{s-1}\|\leq\mathcal{O}\left(\frac{i}{\sqrt{T}}\right).$
-Now, we use the fact that the iterates of OGD $M_{1:h}^{t}$ move slowly to establish that $\mathbf{y}_{t}$ and $\mathbf{y}_{t}(M^{t}_{0:h})$ are close in value.
-Before that, by definition, we have $\mathbf{u}_{t}(M^{t}_{0:h})=\mathbf{u}_{t}$ since the sequence $\mathbf{y}^{\mathbf{nat}}_{t}$ is determined independently of the choice of the policy being executed.
-Due to the $\gamma$-stability of the underlying system, we have for any $i\leq j$, $\|C_{j}\prod_{t=i+1}^{j-1}{A}_{t}B_{i}\|\leq\kappa(1-\delta)^{j-i}$ for some $\kappa>0,\gamma<1$.
-Using Lemma 9.3, we have for some $C>0$ that
-$\|\mathbf{y}_{t}(M^{t}_{0:h})-\mathbf{y}_{t}\|$ $=\left\|C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}A_{t-j}\right]B_{t-i}(\mathbf{u}_{t-i}(M^{t}_{0:h})-\mathbf{u}_{t-i})\right\|$
-$\leq\sum_{i=1}^{t}\left\|C_{t}\left[\prod_{j=1}^{i-1}A_{t-j}\right]B_{t-i}\right\|\|(\mathbf{u}_{t-i}(M^{t}_{0:h})-\mathbf{u}_{t-i})\|$
-$\leq C\kappa\sum_{i=1}^{t}(1-\delta)^{i}\frac{i}{\sqrt{T}}\max_{j}\|\mathbf{y}^{\mathbf{nat}}_{j}\|$
-$\leq\frac{C\kappa}{\delta^{2}\sqrt{T}}\max_{j}\|\mathbf{y}^{\mathbf{nat}}_{j}\|\leq\frac{C\kappa^{2}}{\delta^{3}\sqrt{T}}$
-Noting that $c_{t}$ has bounded derivatives concludes the proof of the lemma. ∎
-Now, we combine the regret guarantee in equation 9.1 and the immediately preceding lemma as
-$\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DRC}}\sum_{t=1}^{T}c_{t}(\hat{\mathbf{y}}_{t},\pi(\hat{\mathbf{y}}_{0:t}))$
-$\leq\sum_{t=1}^{T}l_{t}(M_{0:h}^{t})-\min_{\pi\in\Pi^{DRC}}l_{t}(M_{0:h}^{\pi})+T\times\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)=\mathcal{O}(\sqrt{T}),$
-implying the stated claim.
-9.3 Bibliographic Remarks
-Online nonstochastic control with partial observability was studied in [SSH20], who defined the Gradient Response Controller as well as the policy class of Disturbance Response Controllers.
-The setting was further extended to changing linear dynamics in [MGSH21], where the expressive power of different policy classes was also studied.
-The analogue of online nonstochastic control in the stochastic noise setting is called the Linear Quadraic Gaussian problem, and it is further limited by assuming quadratic noise.
-This classical problem in optimal control was proposed in Kalman’s work [Kal60].
-Regret bounds for controlling unknown partially observed systems were studied in [LAHA20b, LAHA20c, MTR19].
-[LAHA20a], which uses DRC policies even in context of stochastic control, obtains rate-optimal logarithmic bounds for the LQG problem.
-9.4 Exercises
-9.1. Prove Lemma 9.3.
-9.2.
-Consider a partially observed linear dynamical system that is $\gamma$-stable in the absence of control inputs.
-Prove that when a DRC policy $\pi\in\Pi^{DRC}_{H,\gamma}$ is executed on such a dynamical system, the resulting system is $\gamma^{2}$-stable.
-9.3. Prove that the optimal controller that minimizes the aggregate cost for a partially observed linear dynamical system assuming the perturbations are stochastic is a linear dynamical controller.
-9.4. Prove Lemma 8.5.
-9.5. Prove Lemma 8.7.
-Chapter 10 Online Nonstochastic System Identification
-A fundamental difficulty in applied control is that the dynamics governing the system behavior may be unknown, or too complex to describe.
-A prominent example is that of control in medical ventilators described in section 2.1.1.
-The respiratory system of a patient is significantly more complex than the fluid dynamics equations that are commonly used to describe it.
-The subfield in control theory that deals with learning a given dynamical system from experience is called system identification.
-It is notoriously difficult to learn non-linear dynamical system well enough to be able to control them, much less so in high dimensions.
-This is further complicated in our treatment of control: discrete time representation, finite time performance guarantees, and adversarial perturbations as well as noise.
-It is for this reason that we focus on the basic setting of linear time invariant dynamical systems. Thus, for the remainder of this chapter we consider dynamics of the form
-$\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}.$
-The question we ask in this chapter is: can we attain sublinear policy regret vs.
-the class of disturbance action policies, when the only information available to the learner are the sequential state observations $\mathbf{x}_{t}$?
-This may seem very challenging, since the perturbations are adversarially chosen.
-Of course, for the problem to be feasible at all the system has to at least be stabilizable.
-We start with an even stronger assumption, suppose that the linear dynamical system is stable, is it learnable within the framework of online nonstochastic control?
-Perhaps surprisingly, the answer is positive in a strong sense.
+7.2.
+  1. 1.
+    Let $l:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a $\beta$-smooth function and let $A\in\mathbb{R}^{d\times n}$ be a matrix.
+Show that $l(A)$ is necessarily a smooth function for $x\in\mathbb{R}^{n}$.
+  2. 2.
+    Let $l:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a $\mu$-strongly convex function and let $A\in\mathbb{R}^{d\times n}$ be a matrix.
+Show that $l(A)$ is not necessarily strongly-convex for $x\in\mathbb{R}^{n}$.
+7.3. Prove that the loss functions $\ell_{t}$ are convex in $M_{1:h}$, as per Lemma 7.3, even if the initial state is non-zero.
+7.4. Consider control of a general, non-linear dynamical system. Give a version of the GPC algorithm 4 that can be applied to such a setting (without proving any performance guarantee).
+7.5. In this exercise we prove and extend Theorem 7.2 for linear time-varying systems that are stabilizable, rather than stable.
+  1. 1.
+    Write down and prove the theorem for time-varying systems $\{A_{t},B_{t}\}$ that are all $\gamma$-stable.
+  2. 2.
+    Consider a sequence of $\gamma$-stabilizable time-varying systems $\{A_{t},B_{t}\}$.
+Show how to reduce this case to the previous by considering a sequence of linear time-varying $\gamma$-stable systems $\tilde{A}_{t},B_{t}$, where $\tilde{A}_{t}=A_{t}+B_{t}K_{t}$.
+Chapter 8 Online Nonstochastic System Identification
+In the previous chapters, we focused on control algorithms that assume knowledge of the system dynamics.
+However, in many real-world applications, system parameters are either unknown or change over time.
+This necessitates an approach where the controller must simultaneously learn the system model while making control decisions.
+This problem, known as system identification, is fundamental to adaptive control, reinforcement learning, and modern data-driven methods for control.
+8.1 The Role of System Identification in Online Control
+System identification plays a crucial role in control design, as it allows controllers to estimate unknown system parameters in real time.
+Unlike classical identification methods that rely on extensive offline data collection, an online approach must identify the system while actively controlling it.
+This presents several unique challenges:
+  • Feedback loops and nonstationarity: Unlike supervised learning settings, where data is collected passively, control decisions influence future observations, making the estimation process highly dynamic.
+  • Adversarial or arbitrary disturbances: The disturbances affecting the system may not follow a probabilistic model but can be chosen adversarially, leading to challenges in maintaining robustness.
+  • Balancing exploration and performance: The controller must explore sufficiently to estimate system parameters accurately while ensuring stable performance, a tradeoff that directly impacts regret minimization.
+8.2 A Nonstochastic Approach to System Identification
+Traditional system identification techniques often rely on stochastic assumptions, where noise is assumed to be drawn from a known distribution, and system parameters are estimated using statistical inference techniques.
+However, in nonstochastic system identification, no assumptions are made about the noise distribution or system perturbations.
+Instead, identification strategies are designed to be robust to worst-case disturbances and operate under adversarial uncertainty.
+This shift in perspective has several implications:
+  • Performance guarantees without probabilistic assumptions: Unlike stochastic estimation methods, which provide guarantees under specific noise models, nonstochastic approaches aim for regret bounds that hold in arbitrary environments.
+  • Regret as a measure of learning efficiency: The effectiveness of an identification strategy is measured not in terms of statistical efficiency but in its ability to achieve low regret, ensuring that the system is identified quickly enough to enable near-optimal control.
+  • Computational tractability in identification: Classical system identification methods often involve solving nonconvex optimization problems, making them computationally demanding.
+Nonstochastic approaches frequently leverage convex relaxation techniques to sidestep intractability.
+The problem of online system identification is closely related to online convex optimization and sequential decision-making.
+Many regret-minimization techniques developed for online learning can be adapted to system identification settings.
+A key distinction, however, is that in control settings, the learned system model directly affects future decisions, creating a coupling between estimation and control that does not typically exist in standard online learning formulations.
+This connection between system identification and online learning leads to several fundamental questions:
+  • How should a controller optimally trade off exploration (learning system parameters) and exploitation (minimizing control cost)?
+  • What are the best algorithmic strategies for estimating system parameters while controlling in adversarial settings?
+  • What regret guarantees can be achieved in online nonstochastic system identification, and how do they compare to traditional learning-based control approaches?
 8.3 Nonstochastic System Identification
+The problem we study in this section is online nonstochastic system identification of a linear time-invariant dynamical system.
 As a first step, we consider the question of whether the system matrices $A,B$ can be recovered from observations over the state, without regard to the actual cost of control.
@@
 Theorem 8.1.
-For $T_{0}=\Omega\left(\frac{d^{2}k^{2}\kappa^{2}\gamma^{6}}{\varepsilon^{2}}\log\frac{1}{\Delta}\right)$, assuming that
+Assuming that
   • a
@@
   • b
-    the system is $(k,\kappa)$-strongly controllable system,
+    the system is $(k,\kappa)$-strongly controllable,
   • c
-    the perturbations $\mathbf{w}_{t}$ are chosen obliviously (i.e. do not depend on $\mathbf{u}_{t}$),
+    the perturbations $\mathbf{w}_{t}$ do not depend on $\mathbf{u}_{t}$,
+and for $T_{0}=\Omega\left(\frac{d^{2}k^{2}\kappa^{2}\gamma^{6}}{\varepsilon^{2}}\log\frac{1}{\Delta}\right)$, the following holds.
 Algorithm 5 outputs a pair $(\widehat{A},\widehat{B})$ that satisfies, with probability $1-\Delta$,
@@
 Proof.
-First note that for any $t$, $\mathbf{x}_{t}$ is bounded as
-$\|\mathbf{x}_{t}\|=\left\|\sum_{i>1}A^{i-1}\left(\mathbf{w}_{t-i}+B\eta_{t-i}\right)\right\|\leq 2\sqrt{d}\gamma.$
+First note that for any $t$, since the system is $\gamma$ stable and the controls have norm $\sqrt{d}$, the state $\mathbf{x}_{t}$ is bounded as
+$\|\mathbf{x}_{t}\|=\left\|\sum_{i=1}^{t}A^{i-1}\left(B\eta_{t-i}+\mathbf{w}_{t-i}\right)\right\|\leq 2\sqrt{d}\gamma.$
 Therefore, defining $N_{j,t}=\mathbf{x}_{t(k+1)+j+1}\eta_{t(k+1)}^{\top}$ for any $j<k$ and $t<T_{0}$, we have that $\|N_{j,t}\|\leq 2d\gamma$.
@@
 $\|\widehat{C_{0}}-C_{0}\|,\|\widehat{C_{1}}-C_{1}\|\leq\sqrt{k}\varepsilon.$
-By this display, we already have $\|B-\widehat{B}\|\leq\varepsilon$. We will make use of this observation to establish the proximity of $\widehat{A}$ and $A$, whose proof is left as Exercise 10.4.
+By this display, we already have $\|B-\widehat{B}\|\leq\varepsilon$. We will make use of this observation to establish the proximity of $\widehat{A}$ and $A$, whose proof is left as Exercise 8.7.
+Note that $\widehat{A}$ does not necessarily satisfy $\widehat{A}\widehat{C_{0}}=\widehat{C_{1}}$. Hence we require the following lemma.
 Lemma 8.4.
@@
 $\|\Delta X\|\leq\frac{\|\Delta B\|+\|\Delta A\|\|X\|}{\sigma_{\min}(A)-\|\Delta A\|}$
-Note that $\widehat{A}$ does not necessarily satisfy $\widehat{A}\widehat{C_{0}}=\widehat{C_{1}}$.
-Hence, we will invoke the above stated lemma on the normal equations, i.e.
+We now use this lemma on the normal equations, i.e.
 $AC_{0}C_{0}^{\top}=C_{1}C_{0}^{\top}$ and $\widehat{A}\widehat{C_{0}}\widehat{C_{0}}^{\top}=\widehat{C_{1}}\widehat{C_{0}}^{\top}$.
@@
 8.4 From Indentification to Nonstochastic Control
-Having recovered approximate estimates $(\widehat{A},\widehat{B})$ of the system matrices, a natural question arises: how well do GPC and GRC algorithms perform on the original system $(A,B)$ when provided with only approximate estimates of the same?
+Having recovered approximate estimates $(\widehat{A},\widehat{B})$ of the system matrices, a natural question arises: how well does the Gradient Perturbation Algorithm 4 perform on the original system $(A,B)$ when provided with only approximate estimates of the same?
 We briefly state a result to this extent, whose proof we leave to the explorations of an interested reader.
@@
 This result translates to a regret bound of $\tilde{O}(T^{2/3})$ for nonstochastic control of unknown, yet stable, linear systems, by first running the nonstochastic system identification procedure, and subsequently using the estimates thus obtained to run GPC.
-(Details left as Exercise 10.4
-A similar statement holds for the GRC control algorithm, which in addition does not require strong controllability, since it can operate with estimates of the Markov operators alone (using Lemma 9.3), instead of estimates of the system matrices.
+Details of how to derive this regret bound are left as Exercise 8.7.
+8.5 Summary and Key Takeaways
+This chapter explored online nonstochastic system identification, where a controller must simultaneously learn an unknown system model while making control decisions in an adversarial setting.
+Unlike classical stochastic system identification, which assumes structured noise and statistical convergence, the nonstochastic approach makes no such assumptions and instead relies on regret minimization techniques to ensure efficient learning.
+The key insights from this chapter are as follows.
+  • The Exploration-Exploitation Tradeoff A fundamental challenge in online system identification is balancing exploration, which enables learning the system dynamics, and exploitation, which minimizes control cost.
+Too much exploration leads to high immediate costs, while too much exploitation can result in poor long-term learning and suboptimal control performance.
+The algorithms in this chapter address this by incorporating controlled perturbations in the control inputs, ensuring sufficient exploration while keeping regret low.
+However, the tradeoff results in a regret of $O(T^{2/3})$, rather than $O(\sqrt{T})$ we had in the last chapter.
+    More sophisticated techniques for balancing exploration and exploitation can result in better regret bounds, some of which are surveyed in the bibliographic section.
+  • Regret Minimization as a Learning Metric Unlike traditional identification methods that aim for asymptotic convergence, online nonstochastic identification is evaluated in terms of regret, measuring performance relative to the best fixed model in hindsight.
+8.5.1 Looking Ahead
+This part of the book considered the online nonstochastic control problem, the regret metric, and online learning techniques applied to control.
+We have thus far considered the case where the state is observable to the controller, an assumption which doesn’t always hold.
+In the next parts of the book, we consider the setting of partial observability, where the system state is unknown, but an observed signal is available.
+We start with the problem of learning, and move onto online control.
 8.6 Bibliographic Remarks
@@
 8.7 Exercises
-10.1.
+8.1.
 Fix a value of $T$.
@@
 Given Lemma 8.5, conclude that the resultant control procedure produces a regret of at most $\tilde{O}(T^{2/3})$ without any prior knowledge of system matrices, as long as the underlying system is stable.
-10.2. Design an analogous control procedure as that for Exercise 10.4 with an identical regret upper bound (upto constants), when the horizon $T$ is not known in advance.
+8.2. Design an analogous control procedure as that for Exercise 8.7 with an identical regret upper bound (up to constants), when the horizon $T$ is not known in advance.
 8.3. Prove Lemma 8.4.
 Part III Learning and Filtering
-Chapter 11 Learning in Unknown LDS
-In this chapter we consider an easier task than control, namely that of learning in a dynamical system.
-The task of the learner is to predict the next observation from past observations alone, while being unable to influence the underlying dynamical system.
-We show in the next few sections how the methodology of convex relaxations extends to prediction in dynamical systems.
-This chapter uses similar a similar parameterization as DACs in context of prediction, while leaving more intricate parameter-efficient convex relaxations to the next chapters.
+Chapter 9 Learning in Unknown Linear Dynamical Systems
+In previous chapters, we focused on control in dynamical systems, often assuming that the system state was fully observable and that system dynamics were either known or could be estimated through system identification.
+In many real-world scenarios, however, these assumptions do not hold.
+Instead, an agent may only have access to partial observations of the system and must learn to make accurate predictions without directly influencing the system’s evolution.
+This chapter focuses on learning in a dynamical system, which means predicting future system outputs based on past observations while the underlying system parameters remain unknown.
+Unlike control, where an agent actively influences system trajectories to minimize a given cost, learning in this setting is a passive task—the goal is to accurately forecast future outputs without modifying the system’s behavior.
+The ability to predict future states and outputs of an unknown system is fundamental to many domains, including time-series forecasting, financial modeling, sensor networks, autonomous systems, and most recently large language models.
+Moreover, learning a predictive model is often a key stepping stone toward solving more complex problems, such as control under partial observability or reinforcement learning in dynamical systems.
+In this chapter, we study online learning algorithms for prediction in unknown linear dynamical systems, with an emphasis on regret minimization.
 9.1 Learning in Dynamical Systems
-There are numerous different interpretations of “learning” in dynamical systems.
-For this chapter the notion we are concerned with is that of prediction.
-More precisely, consider a sequence of observable inputs to a dynamical system as in Figure 11.1.
-Our goal is to correctly predict the output sequence to minimize a given loss metric.
+The term learning in dynamical systems can refer to a variety of tasks, including state estimation, system identification, and policy learning.
+In this chapter, we focus on the specific task of prediction, where the objective is to estimate future outputs of a dynamical system based on past observations.
+Unlike classical system identification, which seeks to recover the system parameters explicitly, the learning approach taken here aims to optimize predictions directly.
+To formally define the problem, consider a sequence of observable inputs fed into a dynamical system, as illustrated in Figure 9.1.
+The learner observes these inputs and aims to predict the corresponding outputs while minimizing a given loss function.
 Figure 9.1: In learning we passively observe the input sequence and attempt to predict output.
-For the rest of this chapter, we consider a partially observable linear dynamical system that is governed by the following equations.
+For the remainder of this chapter, we consider a partially observable linear time invariant dynamical system described by the following equations.
+The extention of the material in this chapter to time varying linear dynamical systems is left as exercise 9.7.
 $\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}$
 $\mathbf{y}_{t}$ $=C\mathbf{x}_{t}+\mathbf{v}_{t}.$
-Here, as we have denoted throughout the text, $\mathbf{x}$ is the hidden state, $\mathbf{u}$ is the observed control input, $\mathbf{w},\mathbf{v}$ are perturbations/noises, and $\mathbf{y}$ is an observed linear transformation of the state.
-Throughout this chapter we assume that the system matrices $A,B,C$ are unknown to the learner.
+Here, $\mathbf{x}_{t}$ represents the hidden system state, $\mathbf{u}_{t}$ is the observed control input, and $\mathbf{y}_{t}$ is the measured system output.
+The terms $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ account for process and observation noise, respectively.
+A crucial challenge in this setting is that the system matrices $A,B,C$ are unknown to the learner.
+The learner’s objective is to construct an efficient prediction rule that estimates $\mathbf{y}_{t}$ based on past observations, despite the unknown system parameters.
 9.2 Online Learning of Dynamical Systems
-In the problem of online learning in a linear dynamical system, sequentially for $t=1,...,T$, the online learning algorithm:
-  • observes input $\mathbf{u}_{t-1}\in{\mathbb{R}}^{d_{u}}$
-  • predicts output $\hat{\mathbf{y}}_{t}\in{\mathbb{R}}^{d_{y}}$
-  • observes true system output $\mathbf{y}_{t}$
-  • suffers loss according to a pre-specified loss function, $\ell(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})$, e.g.
+In an online learning setting, the learner sequentially observes data and makes predictions in real time. Formally, the process unfolds as follows:
+  1. 1.
+    At each time step $t$, the learner observes the control input $\mathbf{u}_{t-1}$.
+  2. 2.
+    The learner produces a prediction $\hat{\mathbf{y}}_{t}$ for the next output.
+  3. 3.
+    The true output $\mathbf{y}_{t}$ is revealed.
+  4. 4.
+    The learner incurs a loss based on the difference between $\mathbf{y}_{t}$ and $\hat{\mathbf{y}}_{t}$, using a predefined loss function $\ell(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})$, such as squared error:
     $\ell(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})=\|\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\|^{2}.$
-The goal of the online learner is to minimize total loss over all game iterations.
-If the observations ${\mathbf{y}}_{t}$ are arbitrary, this is an ill-defined goal, and rather than minimizing loss, we consider a comparative analogue.
-More precisely, we consider a family of prediction rules from available information to prediction, and consider the metric of regret vs.
-the best of these rules in hindsight.
-An example of a rule could be a fixed linear map, i.e.
-$\pi(\mathbf{u}_{1:t-1},\mathbf{y}_{1:t-1})=K\mathbf{u}_{t-1}$, where $K$ can potentially depend on the system matrices $A,B,C$.
-Formally, let $\Pi$ be a family of mappings from $\{\mathbf{u}_{\tau},\mathbf{y}_{\tau}\}\mapsto\hat{\mathbf{y}}_{t}$, i.e.
-$\pi\in\Pi\ ,\ \pi(\mathbf{u}_{1:t-1},\mathbf{y}_{1:t-1})=\hat{\mathbf{y}}_{t}^{\pi}\in{\mathbb{R}}^{d_{y}}.$
-Let ${\mathcal{A}}$ be a prediction algorithm that iteratively produces predictions $\hat{\mathbf{y}}_{t}$. Then its regret w.r.t. $\Pi$ is defined as
-$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\Pi}\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}^{\pi}_{t}\|^{2}.$
-We proceed to define common prediction classes in the context of linear dynamical systems.
+The goal is to minimize the total loss over time.
+However, in adversarial or unpredictable settings, direct loss minimization may be ill-posed.
+Instead, we consider a regret-based formulation, where the learner’s performance is measured relative to the best possible predictor from a predefined class of functions.
+More precisely, let $\Pi$ be a family of prediction rules mapping past observations to predictions. The regret of an online learning algorithm $A$ with respect to $\Pi$ is given by:
+$\text{Regret}_{T}(A,\Pi)=\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\Pi}\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}^{\pi}_{t}\|^{2}$
+where $\hat{\mathbf{y}}^{\pi}_{t}$ denotes the prediction made at time $t$ by the best function $\pi\in\Pi$ in hindsight.
+This regret-based framework allows us to compare different classes of predictors and design learning algorithms that adapt efficiently to unknown dynamical systems.
 9.3 Classes of Prediction Rules
-We now consider what would be a reasonable prediction rule for the output of a linear dynamical system?
-The case of full observation is particularly simple, and we deal with it first.
-We then consider prediction rules for partially observed linear dynamical systems.
+We now explore what constitutes a reasonable prediction rule for the output of a linear dynamical system.
 However, before considering the prediction rules themselves, we discuss how to compare classes of such rules.
@@
 We say that a class of predictors $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ if the following holds.
-For every sequence of $T$ inputs and outputs $\{\mathbf{u}_{1:T},\mathbf{y}_{1:T}\}$, and every $\pi\in\Pi_{1}$, there exists $\pi_{2}\in\Pi_{2}$ such that,
+For every sequence of $T$ inputs and outputs $\{\mathbf{u}_{1:T},\mathbf{y}_{1:T}\}$, and every $\pi_{2}\in\Pi_{2}$, there exists $\pi_{1}\in\Pi_{1}$ such that,
 $\sum_{t=1}^{T}\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|\leq T\varepsilon,$
@@
 9.3.1 Linear Predictors
-Perhaps the simplest prediction rule is linear, in either the observation, or input, or both. More precisely, the class of all predictions of the form
-$\hat{\mathbf{y}}_{t+1}=\sum_{i=1}^{h}M_{1,i}\mathbf{y}_{y-i}+\sum_{j=1}^{k}M_{2,j}\mathbf{u}_{y-j}.$
-We denote by $\Pi^{LIN}_{h,k,\kappa}$ the class of all linear predictors with $\sum_{i=0}^{h}\|M_{1,i}\|+\sum_{j=1}^{k}\|M_{2,j}\|\leq\kappa$.
+The simplest prediction rule is a linear function of the observations, inputs, or both. More precisely,
 Definition 9.3 (Linear Policies).
-A linear controller parametrized by matrices $\{M_{1,i},M_{2,j}\}\in{\mathbb{R}}^{h\times d_{y}\times d_{y}+k\times d_{y}\times d_{u}}$ is a linear operator mapping state to control.
-We denote by $\Pi_{h,k,\kappa}^{LIN}$ the set of all linear controllers bounded as $\sum_{i=0}^{h}\|M_{1,i}\|+\sum_{j=1}^{k}\|M_{2,j}\|\leq\kappa$.
-11.3.2 Open Loop Optimal Predictors
-An intuitive class of prediction rules that makes use of the special structure of linear dynamical systems is that of open loop linear dynamical predictors, which we now define.
+A linear controller parametrized by matrices $\{M_{1,i},M_{2,j}\}\in{\mathbb{R}}^{h\times d_{y}\times d_{y}+k\times d_{y}\times d_{u}}$ is a linear operator mapping state to control as
+$\hat{\mathbf{y}}_{t}=\sum_{i=1}^{h}M_{1,i}\mathbf{y}_{t-i}+\sum_{j=1}^{k}M_{2,j}\mathbf{u}_{t-j}.$
+We denote by $\Pi_{h,k,\kappa}^{L}$ the set of all linear controllers bounded as $\sum_{i=1}^{h}\|M_{1,i}\|+\sum_{j=1}^{k}\|M_{2,j}\|\leq\kappa$.
+9.3.2 Linear Dynamical Predictors
+An intuitive class of prediction rules that makes use of the special structure of linear dynamical systems is that of open-loop linear dynamical predictors, which we now define.
 The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $\mathbf{C}=I$ is the identity matrix.
@@
 $\vdots$
-$=\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t-1}\mathbf{x}_{0}.$
+$=\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t}\mathbf{x}_{0}.$
 Thus, the operator $\pi_{\{A,B,C,\mathbf{x}_{0}\}}$ takes a history of states and controls and gives the next system output accurately.
-We denote by $\prod_{\star}$ the set of all such predictors, that in the case of zero noise give a perfectly accurate prediction.
+We denote by $\prod^{\star}$ the set of all such predictors, that in the case of zero noise give a perfectly accurate prediction.
 If the spectral radius of $A$ is larger than one, then the system is unstable and the magnitude of $\mathbf{y}_{t}$ can grow arbitrarily large, as discussed in section 4.2.
@@
 $\pi_{\{A,B,C,h\}}=\sum_{i=1}^{h}CA^{i-1}B\mathbf{u}_{t-i},$
-where we define all $\mathbf{u}$ vectors in negative times as zero. Denote by $\prod_{h}^{OLOP}$ the class of all such predictors.
+where we define all $\mathbf{u}$ vectors in negative times as zero. Denote by $\prod_{h}^{\star}$ the class of all such predictors.
 Clearly for a large $h$ this class of predictors well approximates the true observations $\mathbf{y}_{t}$ if the system is activated without any noise. We formalize what this means next.
@@
 Suppose that $\|B\|+\|C\|\leq\gamma$, and $\|\mathbf{u}_{t}\|\leq U$ for all $t\in[T]$.
-Then the class $\prod_{h}^{OLOP}$ for $h=\gamma\log\frac{T\gamma^{3}U}{\varepsilon}$, $\varepsilon$-approximates the class $\prod_{\star}$.
+Then the class $\prod_{h}^{\star}$ for $h=\gamma\log\frac{T\gamma^{3}U}{\varepsilon}$, $\varepsilon$-approximates the class $\prod^{\star}$.
 Proof.
-Consider any sequence and any predictor from $\prod_{\star}$, say $\pi_{1}=\pi_{A,B,C,\mathbf{x}_{0}}$.
-Consider the predictor $\pi_{2}=\pi_{A,B,C,h}\in\prod_{h}$ for the same system matrices.
+Consider any sequence and any predictor from $\prod^{\star}$, say $\pi_{1}=\pi_{A,B,C,\mathbf{x}_{0}}$.
+Consider the predictor $\pi_{2}=\pi_{A,B,C,h}\in\prod_{h}^{\star}$ for the same system matrices.
 Then we have
@@
 $\sum_{t}\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|$ $\leq\sum_{t}\frac{\gamma^{2}\kappa U}{\delta}e^{-\delta h}=\varepsilon.$
-This lemma explains why the class $\prod^{OLOP}$ is an optimal prediction class for open loop prediction, in which the realized noise (through observations) cannot be used inside of the prediction rule, and in the case of full observation.
-We note that the class $\prod^{LIN}$ also $\varepsilon$-approximates the class $\prod^{OLOP}$ for appropriate choice of parameters. This is left as an exercise (Exercise 11.6) to the reader.
 9.4 Efficient Learning of Linear Predictors
 The previous section gives several natural classes of predictors as well some of the relationships between them.
-Notably, the class of open loop optimal predictors is optimal for certain scenarios and requires knowledge of only three matrices.
+Notably, the class of linear dynamical predictors is optimal for certain scenarios and requires knowledge of only three matrices.
 However, the parametrization of this class is non-convex in the matrices $A,B,C$, which presents a computational challenge.
-This challenge is circumvented by using the fact that the class $\prod^{LIN}$ $\varepsilon$-approximates the classe $\prod^{OLOP}$, and as we show in the next chapter, also optimal closed loop predictors.
-Further, the natural parametrization of $\prod^{LIN}$ is convex and thus can be learned by online convex optimization.
+This challenge is circumvented by using the fact that the class $\prod^{L}$ $\varepsilon$-approximates the class $\prod^{\star}$.
+Further, the natural parametrization of $\prod^{L}$ is convex and thus can be learned using online convex optimization.
 The following algorithm learns the class of linear predictors in the sense that it attains a sublinear regret bound for any sequence of adversarial cost functions, as well as perturbations and observation noise.
-Algorithm 6 Learning $\prod^{LIN}$ by Online Gradient Descent
+Algorithm 6 Learning $\prod^{L}$ by Online Gradient Descent
 1:  Input: $M^{1}=M^{1}_{1,1:h},M^{1}_{2,1:k}$, convex constraints set $\mathcal{K}\subseteq{\mathbb{R}}^{h\times d_{y}+k\times d_{u}}$
@@
 5:  end for
-This algorithm is an instance of the online gradient descent algorithm, Theorem A.1 from Appendix A directly gives the following bound for it.
-We denote by $\mathcal{K}$ the set of all $h\times k$-tuple of matrices $M_{1,1:h},M_{2,1:k}$, such that they are bounded as needed.
-Then we obtain
+This algorithm is an instance of the Online Gradient Descent algorithm, and Theorem 6.1 directly gives the following bound for it.
+We denote by $D,G$ the diameter of the set $\mathcal{K}$ as in Algorithm 6, and by $G$ an upper bound on the norm of the gradients of the loss functions $\ell_{t}$.
 Corollary 9.5.
-For choice of step sizes $\eta_{t}=O(\frac{1}{\sqrt{t}})$, and ${\mathcal{A}}$ being Algorithm 6, we have that
-$\mathrm{{regret}}({\mathcal{A}})=\sum_{t}\ell(\hat{\mathbf{y}}_{t}^{\mathcal{A}},\mathbf{y}_{t})-\min_{\pi\in\prod^{LIN}}\ell(\hat{\mathbf{y}}_{t}^{\pi},\mathbf{y}_{t})=O(\sqrt{T}).$
+For choice of step sizes $\eta_{t}=\frac{D}{G\sqrt{t}}$, and ${\mathcal{A}}$ being Algorithm 6, we have that
+$\mathrm{{regret}}({\mathcal{A}})=\sum_{t}\ell(\hat{\mathbf{y}}_{t}^{\mathcal{A}},\mathbf{y}_{t})-\min_{\pi\in\prod^{L}}\ell(\hat{\mathbf{y}}_{t}^{\pi},\mathbf{y}_{t})\leq 2GD\sqrt{T}.$
+9.5 Summary
+In this chapter, we explored the problem of learning in unknown linear dynamical systems, where the goal is to predict future system outputs based on past observations while the underlying system parameters remain unknown.
+Unlike traditional control settings where an agent actively influences system behavior, the learning problem here is passive, focusing on accurate forecasting without intervention.
+We began by formulating the problem in terms of partially observable LDS, highlighting the challenges posed by unknown system matrices and noisy observations.
+We discussed key approaches for online learning in dynamical systems, emphasizing regret minimization as a performance metric.
+The regret framework allows for direct comparisons between learning algorithms by evaluating their cumulative prediction error against the best fixed predictor in hindsight.
+A key takeaway from this chapter is that learning in LDS requires balancing expressivity and computational efficiency.
+Simple linear predictors may be computationally efficient but lack expressivity, whereas more sophisticated dynamical predictors require careful algorithmic design to remain tractable.
+The regret analysis provided insights into how online learning algorithms adapt to unknown system parameters while ensuring competitive performance.
+This discussion sets the stage for the next chapter on Kalman filtering, where we transition from open-loop predictors to closed-loop estimation techniques that integrate feedback from observations to refine state estimates.
 9.6 Bibliographic Remarks
@@
 Resting at the crossroads between statistical estimation and dynamical systems, the theory of filtering had its early beginnings in the works of Norbert Wiener [Wie49].
+For a comprehensive treatment of online learning in games and online convex optimization, see the texts [CBL06, Haz16].
+In the statistical noise setting, a natural approach for learning is to perform system identification and then use the identified system for prediction.
+Such an approach was taken by [SMT+18, SF20, SR19].
+The work of [GLS+20] extends identification-based techniques to adversarial noise and marginally stable systems.
+The class $\prod^{*}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [AHL+18].
+In that paper, it was shown how to learn partially observable, and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
 9.7 Exercises
 9.1. Prove that $\varepsilon$-approximation of prediction classes implies $\varepsilon T$ regret approximation, as formalized in Lemma 9.2.
-11.2. Prove that the class $\prod^{LIN}$ $\varepsilon$-approximates the class $\prod^{OLOP}$ for appropriate choice of parameters.
-Chapter 12 Kalman Filtering
-The previous chapter introduces the concept of open-loop optimal linear predictors and how to efficiently learn them using convex relaxation and online gradient methods.
-However, a good predictor should exploit not only its initial knowledge, but also information as it is being observed and learn from it.
-In an online prediction setting, this amounts to using the online observations to further refine future predictions.
-This is the topic of this chapter: how to use feedback from the ongoing prediction task to improve future predictions.
-Optimal closed-loop predictors, that use feedback from observations, are significantly more powerful and harder to analyze.
-We explore Kalman filters, prove their optimality, as well as convex relaxations that fit into the paradigm of online convex optimization.
+9.2. Prove that the class $\prod^{*}$ is optimal for fully observed linear dynamical systems with zero-mean noise.
+9.3. Show that the policy class $\prod^{\star}_{h}$ can be learned by policy class $\prod^{L}_{\tilde{h},k,\kappa}$ with $\tilde{h}=0$.
+9.4. Give precise regret bounds for learning the class $\prod^{L}_{h,k,\kappa}$ and $\prod^{\star}_{h}$ as a function of the natural class parameters $h,k,\kappa$ and any other relevant parameter.
+9.5. Consider the square loss in algorithm 6. Give a Newton-method based algorithm based on the Online Newton Step algorithm (see [Haz16]) and write down the resulting regret bound.
+9.6.
+Write down the predictor class $\prod^{\star}$ and $\prod^{\star}_{h}$ for time varying linear dynamical systems.
+Research the notion of adaptive regret from [Haz16], and reason about conditions under which the class $\prod^{L}$ approximates $\prod^{*}$ for time varying linear dynamical systems.
+Chapter 10 Kalman Filtering
+State estimation plays a central role in control and prediction tasks, where an agent seeks to infer latent states from noisy observations.
+The Kalman filter is a classical and widely used method for solving this problem, providing optimal state estimates under the assumption of linear system dynamics and Gaussian noise.
+A key feature of the Kalman filter—and a fundamental limitation—is that it requires knowledge of the system matrices, including both the state transition matrix and the observation model.
+These matrices encode how the system evolves over time and how observations relate to the hidden state.
+While the Kalman filter is provably optimal when these parameters are known, real-world applications often involve partially known or entirely unknown system dynamics, making explicit identification challenging.
+In contrast, online learning approaches, such as those introduced in Chapter 9, provide a model-free, open-loop alternative by directly learning optimal linear predictors from data.
+Instead of estimating a latent state and system matrices, these methods construct forecasting models using past observations, making them more adaptive to unknown or time-varying systems.
+However, these approaches do not incorporate feedback, meaning they operate in an open-loop setting, predicting future states without correcting for new observations.
+This chapter moves beyond open-loop prediction to closed-loop state estimation, where past observations continuously refine future predictions.
+The Kalman filter accomplishes this by recursively updating estimates based on incoming measurements, achieving Bayes-optimal performance under Gaussian noise.
+This structured feedback mechanism allows for more accurate and robust estimation compared to open-loop methods, particularly in dynamic and uncertain environments.
+Beyond classical Kalman filtering, we also introduce a closed-loop online learning method that serves as a convex relaxation of the Kalman filter.
+This approach retains the feedback mechanism of Kalman filtering while removing the need for explicit system identification, making it applicable in adversarial or uncertain environments.
+By formulating state estimation as an online convex optimization problem, this method provides provable regret guarantees and serves as a bridge between Bayesian filtering and learning-based estimation techniques.
+We provide a rigorous introduction to Kalman filtering, including its derivation, optimality properties, and computational aspects.
+We also discuss its limitations and contrast it with learning-based alternatives, setting the stage for spectral filtering techniques in the next chapter, which further generalize state estimation in unknown dynamical systems.
 10.1 Observable Systems
-In the discussion below, when considering the infinite horizon setting, prediction makes sense if our observation remains bounded.
-The condition for bounded observations is related to controllability as defined in Section 4.3.
-We first state the definition of observability for general dynamical systems below.
-In absence of noise and control inputs, every dynamical system can be viewed as mapping from an initial state to an infinite sequence of observations.
-Observability asserts that this mapping is injective.
+Observability is a fundamental concept in control theory that determines whether the internal state of a system can be uniquely inferred from its outputs over time.
+In the infinite-horizon setting, prediction is meaningful only if observations remain bounded, a condition closely related to system controllability (as discussed in Section 4.3).
+A general dynamical system maps an initial state to an infinite sequence of observations.
+If this mapping is injective—meaning that each initial state corresponds to a unique sequence of observations—then the system is said to be observable.
+Formally:
 Definition 10.1.
-A partially-observed dynamical system is said to be observable if and only if every sequence of observations produced by the deterministic-equivalent system uniquely determines the initial hidden state.
-For linear systems, checking observability can be boiled to a rank condition on the Kalman observability matrix, which is defined as
-$K^{\prime}_{k}(A,C)=\begin{bmatrix}C\\ CA^{2}\\ \dots\\ CA^{k-1}\end{bmatrix}.$
-We ask the reader to provide a proof of the following result in Exercise 12.5.
+A partially observed dynamical system is observable if and only if every sequence of observations uniquely determines the initial hidden state.
+A simple example of an unobservable system is given as the following variant of the double integrator from section 2.1.2.
+Consider the double integrator dynamical system with $\Delta=1$ and partial observability where the observation matrix is given as follows
+[…]
+This system allows only the velocity of the object to be observed, while the position remains hidden.
+It can be seen that this setting is unobservable, as the initial position cannot be determined based on the velocities and forces applied to the system alone.
+Indeed, this is a special case of the Galilean Relativity principle of Newtonian mechanics, which states that the laws of physics are the same in all inertial reference frames.
+For linear systems, observability can be verified using a rank condition on the Kalman observability matrix, given by:
+$\hat{K}(A,C)=\begin{bmatrix}C\\ CA\\ \dots\\ CA^{d_{\mathbf{x}}-1}\end{bmatrix}.$
+[…]
 Theorem 10.2.
-A partially-observed dynamical system specified by system matrcies $(A,B,C)$ is said to be observable if $\mathop{\mbox{\rm rank}}(K^{\prime}_{d_{x}}(A,C))=d_{x}$.
-12.2 Closed Loop Optimal Predictors
-The previous class of predictors is open-loop, namely it only takes into account the inputs to the system, which can potentially be computed ahead of time, and not the response of the system to the controls and perturbations.
-A closed-loop prediction rule takes into account the observations, and can potentially react to the perturbations and observation noise.
-The most famous example of such a prediction rule is the Kalman filter, which we now define.
-The defining characteristic of the Kalman filter is that it produces the best linear predictor in the least-squares sense when the perturbations $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ are zero-mean and i.i.d., and when the systems matrices $(A,B,C)$ and the perturbations’ covariances $(\Sigma_{w},\Sigma_{v})$ are known and defined as
-$\Sigma_{x}\triangleq\mathbb{E}\left[\mathbf{w}_{t}\mathbf{w}_{t}^{\top}\right],\quad\Sigma_{y}=\mathbb{E}\left[\mathbf{v}_{t}\mathbf{v}_{t}^{\top}\right].$
-So far we have focussed on predicting observations, i.e., ${\mathbf{y}}_{t}$’s.
-On the other hand, the next theorem produces optimal estimators $\hat{\mathbf{x}}_{t}$ of the hidden state $\mathbf{x}_{t}$ given past observations $\mathbf{y}_{1:t-1}$ and controls $\mathbf{u}_{1:t-1}$.
-This is motivated by the fact that in the stochastic case, the optimal least-squares predictor can be written as $\hat{\mathbf{y}}_{t}=C\hat{\mathbf{x}}_{t}$.
-In Exercise 12.5, we ask the reader to supply a proof of this statement.
+A partially-observed dynamical system specified by system matrcies $(A,B,C)$ is observable if and only if $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))=d_{\mathbf{x}}$.
+10.2 Optimality of the Kalman Filter
+10.2.1 The Power of Linear Predictors
+One of the important conclusions from this chapter is that the policy class of linear predictors is optimal with respect to closed loop policies. We prove this by proving two facts:
+  1. 1.
+    The class of linear policies $\varepsilon$-approximates the class of Kalman filters, in the sense of Definition 9.1.
+  2. 2.
+    The class of Kalman filters is optimal for closed loop policies.
+The conclusion from these two facts is that algorithm 6 can be used to learn a policy that competes with the best closed loop optimal predictor in hindsight.
+If the signal comes from a linear dynamical system subject to Gaussian noise, this means competing with the optimal predictor, which is a Kalman filter.
+We prove this formally at the end of this section.
+10.2.2 Closed Loop Optimal Predictors
+Open-loop predictors, such as those introduced in the previous chapter, leverage system inputs and initial conditions to make predictions.
+However, they do not incorporate feedback from observations, making them susceptible to modeling errors and disturbances.
+Closed-loop prediction rules address this limitation by continuously adjusting estimates based on new observations.
+The most well-known example of such a predictor is the Kalman filter, which produces the best linear predictor in a least-squares sense under the assumption that process and observation noise are Gaussian with known covariances.
+Formally, let $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ denote the process and observation noise, assumed to be zero-mean and i.i.d., with covariance matrices:
+$\Sigma_{x}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[\mathbf{w}_{t}\mathbf{w}_{t}^{\top}\right],\quad\Sigma_{y}=\mathbb{E}\left[\mathbf{v}_{t}\mathbf{v}_{t}^{\top}\right].$
+So far, we have focused on predicting observations, i.e., $\mathbf{y}_{t}$.
+However, the Kalman filter also provides optimal state estimates $\hat{\mathbf{x}}_{t}$ given past observations $\mathbf{y}_{t-1:1}$ and controls $\mathbf{u}_{t-1:1}$.
+This is motivated by the fact that the least-squares optimal predictor of observations can be expressed as:
+$\mathbf{y}_{t}=C\mathbf{x}_{t}.$
+In Exercise 10.6, we ask the reader to supply a proof of this statement.
+Based upon this observation, the following theorem is the essence of the Kalman filter, which recovers estimates for states as well as linear projection matrices for optimal prediction of the next observation.
 Theorem 10.3.
@@
 such that
-$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{LIN}}\mathbb{E}\|\mathbf{x}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
-Proof.
-Since the estimator is recursively defined, a natural strategy is to proceed by induction. Let us say $\hat{\mathbf{x}}_{t}$ is a such a predictor. Further, let
-$\Sigma_{t}\triangleq\mathbb{E}(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}.$
-We begin with an observation about least-squares predictors, which we ask the reader to prove in Exercise 12.5.
+$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{L}}\mathbb{E}\|\mathbf{x}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
+To prove this theorem, we require some basic tools from linear regression which we state next.
+10.2.3 Tools from Linear Regression
+We begin with an observation about least-squares predictors, which we ask the reader to prove in Exercise 10.6.
 Lemma 10.4.
@@
 In particular, we know by induction hypothesis that
-$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{LIN}}\mathbb{E}\|\mathbf{x}_{t}-L\left[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}\right]\|^{2}.$
-But, is $\hat{\mathbf{x}}_{t}$ any good for predicting $\mathbf{x}_{t+1}$? To answer this, we note the following structural characterization of least-squares (Exercise 12.5).
+$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{L}}\mathbb{E}\|\mathbf{x}_{t}-L\left[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}\right]\|^{2}.$
+But, is $\hat{\mathbf{x}}_{t}$ any good for predicting $\mathbf{x}_{t+1}$? To answer this, we note the following structural characterization of least-squares (Exercise 10.6).
 Lemma 10.5.
 Let $X,Y$ be possibly correlated random variables. Then
-$A^{*}\triangleq\operatorname*{\arg\min}_{A}\mathbb{E}\|X-AY\|^{2}=\mathbb{E}\left[XY^{\top}\right]\left(\mathbb{E}\left[YY^{\top}\right]\right)^{-1}.$
+$A^{*}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\operatorname*{\arg\min}_{A}\mathbb{E}\|X-AY\|^{2}=\mathbb{E}\left[XY^{\top}\right]\left(\mathbb{E}\left[YY^{\top}\right]\right)^{-1}.$
 Furthermore, if $\hat{X}=A^{*}Y$, then $\mathbb{E}(X-\hat{X})Y^{\top}=0$.
+10.2.4 Analysis
+We are now ready to prove the main property of the Kalman filter below.
+Proof of Theorem 10.3.
+Since the estimator is recursively defined, a natural strategy is to proceed by induction. Let us say $\hat{\mathbf{x}}_{t}$ is a such a predictor. Further, let
+$\Sigma_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}.$
 Since for any $s<t$
@@
 using the first part of Lemma 10.5, we have that
-$\mathbb{E}\|\mathbf{x}_{t+1}-A\hat{\mathbf{x}}_{t}-B\mathbf{u}_{t}\|^{2}\leq\min_{L\in\Pi^{LIN}}\mathbb{E}\|\mathbf{x}_{t+1}-B\mathbf{u}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
+$\mathbb{E}\|\mathbf{x}_{t+1}-A\hat{\mathbf{x}}_{t}-B\mathbf{u}_{t}\|^{2}\leq\min_{L\in\Pi^{L}}\mathbb{E}\|\mathbf{x}_{t+1}-B\mathbf{u}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
 But, when predicting $\mathbf{x}_{t+1}$, we have more information.
@@
 where the last equality once again follows from Lemma 10.5. This orthogonality allows us to invoke Lemma 10.4 to arrive at
-$\hat{\mathbf{x}}_{t+1}=A\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}C^{\top}\left(\Sigma_{y}+C\Sigma_{t}C^{\top}\right)^{-1}}_{\triangleq L_{t}}(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}),$
+$\hat{\mathbf{x}}_{t+1}=A\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}C^{\top}\left(\Sigma_{y}+C\Sigma_{t}C^{\top}\right)^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}),$
 where the last component follows from Lemma 10.5 and the correlations we just computed.
@@
 Finally, we give a recursive relation on $\Sigma_{t}$ as follows–
-$\Sigma_{t+1}$ $\triangleq\mathbb{E}\left[(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})^{\top}\right]$
+$\Sigma_{t+1}$ =$\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})^{\top}\right]$
 $=(A-L_{t}C)\Sigma_{t}(A-L_{t}C)^{\top}+\Sigma_{x}+L_{t}\Sigma_{y}L_{t}^{\top}$
@@
 $=A\Sigma_{t}A^{\top}-A\Sigma C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}C\Sigma A^{\top}+\Sigma_{x}.$
+10.2.5 Infinite Horizon Kalman Filtering
 In the infinite-horizon setting, the steady-state Kalman filter can be described by the following linear dynamical system
@@
 Mirroring the link between controllability and stabilizability, observability implies that $\rho(A-LC)<1$.
-For observable linear dynamical systems, the Kalman filter is contained in the class of linear predictors. Exercise 12.5 asks for a proof of the next lemma.
+10.2.6 Linear Predictors and the Kalman Filter
+For observable linear dynamical systems, the Kalman filter is contained in the class of linear predictors. This is fomally stated in the following Lemma, whose proof is left as Exercise 10.6.
 Lemma 10.6.
 Consider any partially-observed linear dynamical system $(A,B,C)$ where the Kalman filtering matrix $L$ is such that the system $A-LC$ is $\gamma$-stable.
-The Kalman filtering policy is $\varepsilon$-contained in the class of linear policies $\Pi^{LIN}_{h,k,\kappa}$, where $h=k=\gamma\log\frac{\gamma T}{\varepsilon}$ and $\kappa=\gamma^{3}$.
+The Kalman filtering policy is $\varepsilon$-contained in the class of linear policies $\Pi^{L}_{h,k,\kappa}$, where $h=k=\gamma\log\frac{\gamma T}{\varepsilon}$ and $\kappa=\gamma^{3}$.
 10.3 Bayes-Optimality under Gaussian Noise
@@
 $\begin{bmatrix}\mathbf{x}_{t}|(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\\ \mathbf{y}_{t}|(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\hat{\mathbf{x}}_{t}\\ C\hat{\mathbf{x}}_{t}\end{bmatrix},\begin{bmatrix}\Sigma_{t}&\Sigma_{t}C^{\top}\\ C\Sigma_{t}&C\Sigma_{t}C^{\top}+\Sigma_{y}\end{bmatrix}\right).$
-for some $\Sigma_{t}\in{\mathbb{R}}^{d_{x}\times d_{x}}$, for which we provide a recursive formula towards the end of the proof.
-We make note of the following lemma for Gaussian distributions, whose proof we ask the reader to complete in Exercise 12.5.
+for some $\Sigma_{t}\in{\mathbb{R}}^{d_{\mathbf{x}}\times d_{\mathbf{x}}}$, for which we provide a recursive formula towards the end of the proof.
+We make note of the following lemma for Gaussian distributions, whose proof we ask the reader to complete in Exercise 10.6.
 Lemma 10.8.
@@
 $\hat{\mathbf{x}}_{t+1}$ $=A\mathbb{E}[\mathbf{x}_{t}|\mathbf{y}_{1:t},\mathbf{u}_{1:t-1}]+B\mathbf{u}_{t}$
-$=(A-\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\triangleq L_{t}}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\triangleq L_{t}}\mathbf{y}_{t}.$
+$=(A-\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}\mathbf{y}_{t}.$
 It remains to obtain a recursive relation on $\Sigma_{t}$. To this extent, from the state evolution of a linear dynamical system, we have that
 $\Sigma_{t+1}=A\Sigma^{\prime}_{t}A^{\top}+\Sigma_{x}=A\Sigma_{t}A^{\top}-A\Sigma_{t}C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}C\Sigma_{t}A^{\top}+\Sigma_{x}.$
+10.4 Conclusion
+In this chapter, we explored the Kalman filter as a fundamental tool for state estimation in dynamical systems, particularly in the presence of noise and uncertainty.
+We established its optimality in the least-squares sense and examined its connections to online convex optimization, reinforcing its relevance in both classical control and modern learning-based paradigms.
+Key insights from this chapter include the role of observability in ensuring accurate state estimation, the derivation of the Kalman filter update equations, and the recursive structure that allows for efficient real-time implementation.
+We also presented the filter’s steady-state behavior, particularly in the infinite-horizon setting, where it converges to a solution governed by the Riccati equation.
+Beyond its theoretical guarantees, the Kalman filter is widely used in practical applications, ranging from autonomous systems and robotics to finance and neuroscience.
+However, challenges remain, particularly in scenarios with non-Gaussian noise, time-varying dynamics, or adversarial disturbances.
+These limitations motivate extensions such as robust filtering, nonlinear filtering (e.g., the extended and unscented Kalman filters), and learning-based adaptations that integrate statistical inference with modern machine learning techniques.
+The next chapter will transition from the classical filtering framework to spectral filtering, providing an alternative approach to state estimation that leverages spectral methods for efficient learning in dynamical systems.
 10.5 Bibliographic Remarks
@@
 Improper learning approaches to this problem were studied in [KMTM19].
+State estimation is a fundamental problem in control and signal processing, with numerous approaches beyond the classical Kalman filter. Alternative methods include:
+  • Nonlinear Filtering Extensions.
+    The Kalman filter assumes linear system dynamics and Gaussian noise, limiting its effectiveness in nonlinear settings.
+The Extended Kalman Filter (see e.g.
+[UA21]) addresses this by linearizing the dynamics at each time step via a first-order Taylor series expansion.
+However, the linearization can introduce significant errors in highly nonlinear systems.
+Notable extensions using more sophisiticated sampling include the Unscented Kalman Filter [JU04], leading to better estimation accuracy in strongly nonlinear systems.
+  • Particle Filtering (Sequential Monte Carlo Methods)
+    For highly nonlinear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative.
+These methods approximate the posterior distribution using a set of weighted samples (particles) and employ importance sampling and resampling techniques.
+  • Adversarial & Robust Filtering
+    Robust filtering methods seek to provide guarantees under adversarial noise.
+The $H_{\infty}$ filter [BB08] minimizes the worst-case estimation error rather than the mean-square error, making it suitable for applications with unknown disturbances.
+While the Kalman filter remains a cornerstone of state estimation, these alternative methods provide valuable extensions and improvements for nonlinear, adversarial, and data-driven settings.
+In the next chapter we will study an alternative using a technique which circumvents the need for state estimation altogether, and has stronger guarantees for sequential prediction.
 10.6 Exercises
-12.1. Prove Lemma 12.5.
-12.2. Using Lemma 12.5, prove Lemma 12.4.
-12.3. Prove Lemma 12.6.
-12.4. Prove Lemma 12.8.
-12.5. Prove Theorem 12.2.
-12.6.
+10.1. In this exercise we prove Theorem 10.2 in two parts:
+  1. 1.
+    Show that if the rank of the Kalman observability matrix is less than $d_{\mathbf{x}}$, then there are two initial states that after $d_{\mathbf{x}}$ iterations map to the same state.
+Hence the initial state is not uniquely determind.
+  2. 2.
+    Conversely, show that if the rank of $\hat{K}$ is full, then at any time $t$, $\mathbf{x}_{t}$ determines $\mathbf{x}_{0}$ uniquely.
+10.2. Prove Lemma 10.5, which states the following. Let $X,Y$ be possibly correlated random variables. Then
+$A^{*}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\operatorname*{\arg\min}_{A}\mathbb{E}\|X-AY\|^{2}=\mathbb{E}\left[XY^{\top}\right]\left(\mathbb{E}\left[YY^{\top}\right]\right)^{-1}.$
+Furthermore, if $\hat{X}=A^{*}Y$, then $\mathbb{E}(X-\hat{X})Y^{\top}=0$.
+10.3. Using Lemma 10.5, prove Lemma 10.4, which states the following. Let $X,Y,Z$ be matrix-valued random variables. Let
+$A^{*}=\operatorname*{\arg\min}_{A}\mathbb{E}\|X-AY\|^{2},\text{ and }B^{*}=\operatorname*{\arg\min}_{B}\mathbb{E}\|X-BZ\|^{2}.$
+If $\mathbb{E}\left[YZ^{\top}\right]=0$, then
+$\mathbb{E}\|X-A^{*}Y-B^{*}Z\|^{2}\leq\min_{A,B}\mathbb{E}\|X-AY-BZ\|^{2}.$
+10.4.
+In this exercise we conclude that the Kalman filter can be improperly learned by the class of linear predictors.
+Prove this fact as formally stated in Lemma 10.6.
+Conclud that the Kalman filter can be learned using Algorithm 6.
+10.5. Prove Lemma 10.8, which is stated as follows. If a random variable pair $(\mathbf{x},\mathbf{y})$ are distributed as
+$\begin{bmatrix}\mathbf{x}\\ \mathbf{y}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mu_{x}\\ \mu_{y}\end{bmatrix},\begin{bmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}\end{bmatrix}\right),$
+then the conditional distribution of $\mathbf{x}$ given $\mathbf{y}$ can be written as $\mathbf{x}|\mathbf{y}\sim\mathcal{N}(\mu_{x}+\Sigma_{12}\Sigma_{22}^{-1}(\mathbf{y}-\mu_{y}),\Sigma_{11}-\Sigma_{12}\Sigma_{22}^{-1}\Sigma_{21}).$
+10.6.
 Using Lemma 10.5, prove that the optimal least-squares estimator for the observation at time $t$ given past observations $y_{1:t-1}$ and control inputs $u_{1:t-1}$ is given by $\hat{\mathbf{y}}_{t}=C\hat{\mathbf{x}}_{t}$, where $\hat{\mathbf{x}}_{t}$ is as defined in Theorem 10.3.
-Chapter 13 Spectral Filtering
-In the previous chapter we have showed how to learn an unknown linear dynamical system with perturbations and noise using convex relaxation and online learning.
-Instead of learning the system matrices directly, we have learned an alternate class of predictors, namely the class of linear predictors over past observations and controls, that approximates the optimal predictor in certain cases.
-Despite the clean and simple formulation of this result, there remains an unsatisfying issue with the previous chapter.
-The number of parameters required to learn the optimal open loop predictor in the class $\prod_{\star}$ depends on the measure of stability of the underlying system, which is potentially large and unknown.
-In this section we describe an elegant alternative method for learning a linear dynamical system that removes this dependence on the stability constant of the system.
-For the sake of exposition, we restrict ourselves only to symmetric system matrices $A$.
-This is a significant restriction, which can be removed, albeit using a significantly more sophisticated approach which is referred to in the bibliographic section.
+Chapter 11 Spectral Filtering
+In the previous chapter, we explored two major approaches for state estimation and prediction in unknown linear dynamical systems:
+  1. 1.
+    State estimation via Kalman filtering, which recursively estimates both the latent system state and the projection matrices that map the system state to observations.
+Kalman filtering provides an optimal solution under Gaussian noise, and updates the state estimates over time.
+However, it requires an accurate model of system dynamics, including transition and observation matrices, which may be difficult to estimate, especially in adversarial or high-noise environments.
+  2. 2.
+    Online learning of linear predictors, which offers a relaxation of the Kalman filtering approach by directly learning predictive models without explicit system identification.
+Instead of estimating the full state and transition matrices, this approach constructs a forecasting model using past inputs and outputs.
+While this avoids the complexity of full system identification, it introduces a major limitation: the number of parameters required depends on the system’s stability constant, which can be large and unknown, making learning impractical for unstable or poorly conditioned systems.
+In this chapter, we introduce Spectral Filtering, an alternative approach that bypasses these limitations.
+Rather than relying on state estimation or direct system identification, spectral filtering constructs predictors in a transformed basis, which does not depend on the system’s stability properties.
+This method leverages spectral decompositions to efficiently extract the relevant dynamical structure, enabling more robust and scalable learning in both stable and unstable systems.
+To illustrate the power of spectral filtering, we first present a one-dimensional case, which provides an intuitive understanding of how spectral methods can be applied to system identification and forecasting.
+We then extend the discussion to higher-dimensional settings, demonstrating how spectral predictors enable a structured and computationally efficient representation of system dynamics.
+While this chapter primarily focuses on symmetric system matrices for clarity, the principles of spectral filtering extend to asymmetric and more general linear dynamical system settings.
+These extensions require additional mathematical tools, which we reference in the bibliographic section.
 11.1 Spectral filtering in One Dimension
@@
 $=\beta(\mathbf{u}_{t-1}+\sum_{i=1:t-1}(\alpha^{i}-\alpha^{i-1})\mathbf{u}_{t-1-i})$
-$=\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1}),$ (13.1)
+$=\beta\mathbf{u}_{t-1}+\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1}),$ (11.1)
 which is a geometric series involving $\alpha$ multiplying the past controls. We used the operator notation to denote the operator
-$\mu_{\alpha}(\mathbf{u}_{t-1:1})=\mathbf{u}_{t-1}+\sum_{i=1:t-1}(\alpha-1)\alpha^{i-1}\mathbf{u}_{t-1-i}=\mu_{\alpha}^{\top}\tilde{\mathbf{u}}_{t-1}.$ (13.2)
+$\mu_{\alpha}(\mathbf{u}_{t-1:1})=\sum_{i=1:t-1}(\alpha-1)\alpha^{i-1}\mathbf{u}_{t-1-i}=\mu_{\alpha}^{\top}\tilde{\mathbf{u}}_{t-1}.$ (11.2)
 Above we used the vector notation for
-$\mu_{\alpha}=(\alpha-1)\begin{bmatrix}\frac{1}{\alpha-1}&1&\alpha&\alpha^{2}&\ldots&\alpha^{T-1}\end{bmatrix}\in{\mathbb{R}}^{T},$
+$\mu_{\alpha}=(\alpha-1)\begin{bmatrix}1&\alpha&\alpha^{2}&\ldots&\alpha^{T-1}\end{bmatrix}\in{\mathbb{R}}^{T},$
 and we define $\tilde{\mathbf{u}}_{t}$ to be the padded vector $\mathbf{u}_{t-1:1}$:
-$\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}.$
+$\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}.$
 The main idea of spectral filtering is to rewrite the vectors $\mu_{\alpha}$ in a more efficient basis, as opposed to the standard basis.
@@
 Definition 11.2.
-Define the class of spectral linear predictors of order $k$ as
-$\prod_{h}^{\lambda}=\left\{\pi_{M}\ \ \left|\ \ \pi_{M}(\mathbf{u}_{1:t},\mathbf{y}_{1:t})=\sum_{i=1}^{h}M_{i}\cdot\phi_{i}(\tilde{\mathbf{u}}_{t-i})\right.\right\}.$
-Here the operator $\phi_{i}(\mathbf{u}_{t-1})$ takes the inner product of the $i$’th eigenvector of $Z_{T}$ with $\tilde{\mathbf{u}}_{t-1}$, where $\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}$.
+Define the class of spectral linear predictors of order $k$ recursively as
+$\prod_{h}^{\lambda}=\left\{\pi_{M}\ \ \left|\ \ \pi_{M}(\mathbf{u}_{1:t},\mathbf{y}_{1:t})-\pi_{M}(\mathbf{u}_{1:t-1},\mathbf{y}_{1:t-1})=M_{0}\mathbf{u}_{t}+\sum_{i=1}^{h}M_{i}\cdot\phi_{i}(\tilde{\mathbf{u}}_{t-1})\right.\right\}.$
+Here the operator $\phi_{i}(\mathbf{u}_{t-1})$ takes the inner product of the $i$’th eigenvector of $Z_{T}$ with $\tilde{\mathbf{u}}_{t-1}$, where $\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}$.
 Notice that as $h\mapsto T$, this class is equivalent to the class of all linear predictors on the inputs $\prod^{LIN}$.
@@
 Lemma 11.4.
-Let $\{\phi_{j}\}$ be the eigenvalues of Z. Then for all $j\in[T]$, ,$\alpha\in[0,1]$, we have
-$|\phi_{j}^{\top}\mu_{\alpha}|=O\left(T^{2}e^{-\frac{j}{\log T}}\right).$
+Let $\{\phi_{j}\}$ be the eigenvectors of Z. Then there exists constants $c_{1},c_{2}>0$ such that for all $j\in[T]$, $\alpha\in[0,1]$, we have
+$|\phi_{j}^{\top}\mu_{\alpha}|\leq c_{1}e^{-\frac{c_{2}j}{\log T}}.$
 Using this property, we can prove Lemma 11.3, by simply shifting to the spectral basis representation as follows.
@@
 Consider any sequence and any predictor from $\prod_{\star}$, say $\pi_{1}=\pi_{A,B,C,0}$. Following equation (11.1), we can write the optimal predictor in the spectral basis:
-$\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}$ $=\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1})$
+$\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\beta\mathbf{u}_{t}$ $=\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1})$
 […]
@@
 By Lemma 11.4 we have that
-$M_{i}^{\star}=\beta\phi_{i}^{\top}\mu_{\alpha}=O(T^{2}e^{-\frac{i}{\log T}}).$
+$M_{i}^{\star}=\beta\phi_{i}^{\top}\mu_{\alpha}=O(e^{-\frac{i}{\log T}}).$
 We proceed by an inductive claim on $\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|$.
@@
 $\leq\sum_{i=h+1}^{T}\|M_{i}^{\star}\||\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-1}|$
-$\leq\sum_{i=h+1}^{T}\|M_{i}^{\star}\|\|\phi_{i}\|_{1}\|\tilde{\mathbf{u}}_{t-1}\|_{\infty}$ Holder inequality
+$\leq\sum_{i=h+1}^{T}\|M_{i}^{\star}\|\|\phi_{i}\|_{2}\|\tilde{\mathbf{u}}_{t-1}\|_{2}$ Cauchy-Schwartz
 $\leq\sqrt{T}\cdot\sum_{i=h+1}^{T}|M_{i}^{\star}|$ $|\mathbf{u}_{t}|\leq 1,\|\phi_{i}\|_{2}\leq 1$
@@
 Lemma 11.4.
-Observe that
-$\|\mu_{\alpha}\|_{2}^{2}=\sum_{t=0}^{T-1}(1-\alpha)^{2}\alpha^{2i}\leq\frac{(1-\alpha)^{2}}{1-\alpha^{2}}\leq 1-\alpha$
-implying that $g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}\leq\|\phi_{j}\|^{2}\|\mu_{\alpha}\|^{2}\leq 1-\alpha$. Similarly, we have
+Consider the scalar function $g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}$ over the interval $[0,1]$.
+First, notice that by definition of $\phi_{j}$ as the eigenvectors of $Z_{T}$, and $\sigma_{j}$ as the corresponding eigenvalues, we have
+[…]
+On the other hand, we claim that the function $g(\alpha)$ is L-Lipschitz for a constant value of $L$. To see this, first notice that
+$\|\mu_{\alpha}\|_{2}^{2}=\sum_{i=0}^{T-1}(1-\alpha)^{2}\alpha^{2i}\leq\frac{(1-\alpha)^{2}}{1-\alpha^{2}}=\frac{1-\alpha}{1+\alpha}\leq 1-\alpha,$
+implying that $g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}\leq\|\phi_{j}\|^{2}\|\mu_{\alpha}\|^{2}\leq 1-\alpha$. Next, we have
 $\left\|\frac{\partial}{\partial\alpha}\mu_{\alpha}\right\|_{2}^{2}$ $=\sum_{t=0}^{T-1}((1-\alpha)i\alpha^{i-1}-\alpha^{i})^{2}$
@@
 $|g^{\prime}{(\alpha)}|$ $=\left\lvert\frac{\partial}{\partial{\alpha}}(\mu_{\alpha}^{\top}\phi_{j})^{2}\right\rvert=2|\mu_{\alpha}^{\top}\phi_{j}|\left\lvert\frac{\partial}{\partial{\alpha}}(\mu_{\alpha}^{\top}\phi_{j})\right\rvert$
+$\leq 2\|\mu_{\alpha}\|_{2}\|\phi_{j}\|_{2}\left\lvert\frac{\partial}{\partial{\alpha}}(\mu_{\alpha}^{\top}\phi_{j})\right\rvert$ Cauchy-Schwarz
 $\leq 2\|\mu_{\alpha}\|_{2}\|\phi_{j}\|^{2}_{2}\left\|\frac{\partial}{\partial{\alpha}}\mu_{\alpha}\right\|_{2}$ Cauchy-Schwarz
@@
 […]
-$=8T^{2}\sigma_{j}$
-$\leq O(T^{2}e^{-\frac{j}{\log T}})$ Theorem 13.1
+$=12\sigma_{j}$
+$\leq 12c_{1}e^{-\frac{c_{2}j}{\log T}}$ Theorem 11.1
 11.2.1 Generalization to High Dimension
@@
 Decomposing over each dimension $j\in[d]$ separately, and denoting $D_{jj}=\alpha_{j}$, we can write (11.1):
-$\hat{\mathbf{y}}^{\pi}_{t}-\mathbf{y}_{t-1}-CB\mathbf{u}_{t-1}$
+$\mathbf{y}_{t}-\mathbf{y}_{t-1}-CB\mathbf{u}_{t-1}$
 $=\sum_{i=1}^{t-1}CU^{\top}(D^{i-1}-D^{i-2})UB\mathbf{u}_{t-1-i}$
-$=\sum_{i=1}^{t-1}CU^{\top}(\sum_{j=1}^{d_{x}}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})e_{i}e_{i}^{\top})UB\mathbf{u}_{t-i-1}$
-$=\sum_{j=1}^{d_{x}}\beta_{j}\sum_{i=1}^{t-1}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})\mathbf{u}_{t-1-i}$ for scalars $\beta_{j}$
-$=\sum_{j=1}^{d_{x}}{M}^{\star}_{j}\mu_{\alpha_{j}}(\mathbf{u}_{1:t-1})$
-Thus, we can apply the same approximation techniques from the previous section, and derive an analoguous conclusion to Lemma 13.3 for dimension higher than one. The details are left as exercise 13.5.
+$=\sum_{i=1}^{t-1}CU^{\top}(\sum_{j=1}^{d_{x}}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})e_{j}e_{j}^{\top})UB\mathbf{u}_{t-i-1}$
+$=\sum_{j=1}^{d_{x}}c_{j}b_{j}^{\top}\sum_{i=1}^{t-1}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})\mathbf{u}_{t-1-i}$ $c_{j},b_{j}\leftarrow j^{th}\mbox{ row/col of }CU^{\top},UB$
+$=\sum_{j=0}^{d_{x}}\left(M^{\star}_{j}\mathbf{u}_{t}+{M}^{\star}_{j}\mu_{\alpha_{j}}(\mathbf{u}_{1:t-1})\right)$
+$=\tilde{M}^{\star}\mathbf{u}_{t}+\tilde{M}^{\star}\mu_{\alpha_{j}}(\mathbf{u}_{1:t-1})$
+Thus, we can apply the same approximation techniques from the previous section, and derive an analogous conclusion to Lemma 11.3 for dimension higher than one.
+Crucially, the number of parameters required does not depend on the hidden dimension of the original system.
+The details are left as exercise 11.6.
 11.3 Online Spectral Filtering
-The prediction class $\prod_{h}^{\lambda}$ can be learned by online gradient descent in the online convex optimization framework, giving rise to the following algorithm.
+The prediction class $\prod_{h}^{\lambda}$ can be learned by online gradient descent in the online convex optimization framework, as per Algorithm 3, giving rise to the following algorithm.
 Algorithm 7 Online Spectral Filtering
@@
 3:  for $t=1$ to $T$ do
-4:     Predict $\hat{\mathbf{y}}_{t}=\sum_{j=1}^{h}M^{t}_{j}\phi_{j}(\tilde{\mathbf{u}}_{t-i})$ , as per definition 13.2, and observe true $\mathbf{y}_{t}$.
+4:     Predict $\hat{\mathbf{y}}_{t}=M_{0}^{t}\mathbf{u}_{t}+\sum_{j=1}^{h}M^{t}_{j}\phi_{j}(\tilde{\mathbf{u}}_{t-i})$ , as per definition 11.2, and observe true $\mathbf{y}_{t}$.
 5:     Define cost function $f_{t}(M)=\|\mathbf{y}_{t}-\tilde{\mathbf{y}}_{t}\|^{2}$.
@@
 Figure 11.2: The filters obtained by the eigenvectors of $Z$.
-The online gradient descent theorem, as per the previous chapter, directly gives the following corollary,
+The regret guarantees of the online gradient descent algorithm, as given in Theorem 6.1, directly implies the following corollary,
 Theorem 11.5.
-For choice of step sizes $\eta_{t}=O(\frac{1}{\sqrt{t}})$, we have that
-$\sum_{t}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\prod_{h}^{\lambda}}\|\mathbf{y}_{t}-\mathbf{y}_{t}^{\pi}\|^{2}=O(\sqrt{T}).$
+Let $D$ be the diameter of the parameter set $\mathcal{K}$, and $G$ an upper bound on the norm of the gradients for the loss functions $f_{t}$ as in Algorithm 7.
+For choice of step sizes $\eta_{t}=\frac{D}{G\sqrt{t}}$, we have that
+$\sum_{t}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\prod_{h}^{\lambda}}\|\mathbf{y}_{t}-\mathbf{y}_{t}^{\pi}\|^{2}\leq 2GD\sqrt{T}.$
+11.4 Conclusion
+In this chapter, we introduced Spectral Filtering as a novel approach to state estimation and prediction in linear dynamical systems, offering an alternative to classical methods like Kalman filtering and online learning of linear predictors.
+By leveraging spectral decompositions, we bypass the need for explicit system identification and provide a framework that is robust to system stability properties.
+We began by outlining the core challenges of traditional state estimation approaches.
+Kalman filtering, while optimal under Gaussian noise assumptions, requires knowledge of system matrices, which may be difficult to estimate in high-noise or adversarial settings.
+Online learning of linear predictors relaxed these requirements but introduced a dependence on system stability constants, which can be large and unknown.
+Spectral Filtering circumvents these limitations by constructing predictors in a transformed basis, allowing for efficient representation without requiring explicit system matrices.
+We demonstrated the method first in the one-dimensional case, providing intuition for how spectral predictors extract relevant dynamical structure.
+This was then extended to higher-dimensional settings, where spectral filtering enables a structured and computationally efficient approach to forecasting.
+Crucially, we showed that this approach does not rely on system stability constants, making it particularly appealing for learning in both stable and unstable dynamical systems.
+From a methodological standpoint, we connected Spectral Filtering to convex optimization techniques, showing that it can be implemented via online learning in a spectral basis.
+This formulation allows for efficient sequential learning and prediction, with provable performance guarantees.
+Looking ahead, Spectral Filtering opens the door for several exciting directions.
+Extensions to asymmetric system matrices, as referenced in the bibliographic section, provide a pathway toward broader applicability.
+Moreover, integrating spectral methods with control strategies may lead to novel approaches for nonstochastic control with partial observations, which we explore in the next chapter.
 11.5 Bibliographic Remarks
-Under probabilistic assumptions, the work of [HMR16] shows that the matrices $A,B,C,D$ of an unknown and partially observed time-invariant linear dynamical system can be learned using first-order optimization methods.
+Under probabilistic assumptions, the work of [HMR18] shows that the matrices $A,B,C,D$ of an unknown and partially observed time-invariant linear dynamical system can be learned using first-order optimization methods.
 Deviating from classical recovery and into the improper prediction world, the work of [HSZ17] devised an improper learning technique using the eigenvectors of the Hankel matrix $Z$.
 The spectral properties of Hankel matrices used in this chapter are from [BT17].
-For an extension of the spectral learning technique to asymmetric matrices, see [HLS+18].
+The first extension of the spectral learning technique to asymmetric matrices is due to [HLS+18].
+A more recent advancement of spectral filtering to general linear dynamical systems was recently put forth in [MH25].
+The technique of spectral filtering has found applications in control and sequence prediction since its introduction.
+In [AHL+18], the authors leverage a spectral filtering technique to represent linear dynamical systems concisely in a linear basis, formulating the optimal control problem as a convex program.
+This method circumvents the non-convex challenges associated with explicit system identification and latent state estimation, providing provable optimality guarantees for the control signal.
+A recent line of work makes use of the spectral basis discovered in the line of work surveyed in this chapter in order to design neural architectures for sequence prediction.
+This was first suggested in the context of state space models in the spectral basis in [ASCH23], and extended to hybrid models [LND+24].
+Later extensions explored efficient implementation [ACD+24] and the concept of length generalization in sequence prediction [MDA+24].
 11.6 Exercises
@@
 11.2. Generalize Lemma 11.3 to any dimension $d_{x}>1$.
-Appendix A Concepts from Online Convex Optimization
-In this section we give a brief introduction to the theory of Online Convex Optimization. The reader is referred to [Haz16] for a comprehensive survey.
-In the decision making framework of Online Convex Optimization, a player iteratively chooses a point in a convex set $\mathcal{K}\subseteq{\mathbb{R}}^{n}$, and suffers a loss according to an adversarially chosen loss function $f_{t}:\mathcal{K}\mapsto{\mathbb{R}}$.
-Let ${\mathcal{A}}$ be an algorithm for OCO, which maps a certain game history to a decision in the decision set:
-$\mathbf{x}_{t}^{\mathcal{A}}={\mathcal{A}}(f_{1},...,f_{t-1})\in\mathcal{K}.$
-We define the regret of ${\mathcal{A}}$ after $T$ iterations as:
-$\mathrm{{regret}}_{T}({\mathcal{A}})=\sup_{\{f_{1},...,f_{T}\}\subseteq{\mathcal{F}}}\left\{\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t}^{\mathcal{A}})-\min_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})\right\}.$ (A.1)
-If the algorithm is clear from the context, we henceforth omit the superscript and denote the algorithm’s decision at time $t$ simply as $\mathbf{x}_{t}$.
-Intuitively, an algorithm performs well if its regret is sublinear as a function of $T$ (i.e.
-$\mathrm{{regret}}_{T}({\mathcal{A}})=o(T)$), since this implies that on average, the algorithm performs as well as the best fixed strategy in hindsight.
-We proceed to describe a basic algorithm for online convex optimization.
-A.1 Online Gradient Descent
-Perhaps the simplest algorithm that applies to the most general setting of online convex optimization is online gradient descent.
-This algorithm, which is based on standard gradient descent from offline optimization.
-Pseudo-code for the algorithm is given in Algorithm 8.
-Algorithm 8 online gradient descent
-1:  Input: convex set $\mathcal{K}$, $T$, $\mathbf{x}_{1}\in\mathcal{K}$, step sizes $\{\eta_{t}\}$
-2:  for $t=1$ to $T$ do
-3:     Play $\mathbf{x}_{t}$ and observe cost $f_{t}(\mathbf{x}_{t})$.
-4:     Update and project:
-$\mathbf{y}_{t+1}=\mathbf{x}_{t}-\eta_{t}\nabla f_{t}(\mathbf{x}_{t})$
-$\mathbf{x}_{t+1}=\operatorname*{\arg\min}_{\mathbf{x}\in\mathcal{K}}\|\mathbf{x}-\mathbf{y}_{t+1}\|^{2}$
-5:  end for
-In each iteration, the algorithm takes a step from the previous point in the direction of the gradient of the previous cost.
-This step may result in a point outside of the underlying convex set.
-In such cases, the algorithm projects the point back to the convex set, i.e.
-finds its closest point in the convex set.
-Despite the fact that the next cost function may be completely different than the costs observed thus far, the regret attained by the algorithm is sublinear.
-This is formalized in the following theorem.
-Here we define the constants $D$ and $G$ as follows:
-  • $D$ = Euclidean diameter of the decision set, i.e.
-    $\max_{\mathbf{x},\mathbf{y}\in\mathcal{K}}\|\mathbf{x}-\mathbf{y}\|.$
-  • $G$ = an upper bound on the Lipschitz constant of the cost functions $f_{t}$, or equivalently, the subgradient norms of these functions. Mathematically,
-    $\forall t\ ,\ \|\nabla f_{t}(\mathbf{x}_{t})\|\leq G.$
-Theorem A.1.
-Online gradient descent with step sizes $\{\eta_{t}=\frac{D}{G\sqrt{t}},\ t\in[T]\}$ guarantees the following for all $T\geq 1$:
-$\mathrm{{regret}}_{T}=\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\min_{\mathbf{x}^{\star}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x}^{\star})\ \leq\frac{3}{2}{GD}\sqrt{T}.$
+11.3.
+The fact that $Z_{T}$ is positive semidefinite is important to having the eigenvalues of $Z_{T}$ decay.
+Give an example of a family of $n\times n$ non-positive semidefinite Hankel matrices that have eigenvalues with constant absolute value.
+11.4.
+The view of Theorem 13.5 is a bit misleading, as it compares the iterates of Online Gradient Descent with the hypothesis class $\Pi_{h}^{\lambda}$ and uses big-Oh notation to obscure the diameter and Lipschitz constants of the predictors.
+Assume that $d_{x}=1$.
+  1. (i)
+    Find a convex set $\mathcal{M}$ of tuples $(M_{1},M_{2},\dots)$’s to optimize across such that its diameter is some $\textrm{polylog}(T,1/\varepsilon)$.
+The norm we measure diameter with respect to is the sum of the Frobenius norms of the $M_{i}$’s.
+  2. (ii)
+    One can further show that the norm of the gradient is at most $\textrm{polylog}(T,1/\varepsilon)$ as well.
+Use these two facts to show that with the right value of $\varepsilon$, the regret against the hypothesis class $\Pi_{\star}$ is at most $\tilde{O}(\sqrt{T})$.
+Part IV Online Control with Partial Observation
+Chapter 12 Policy Classes for Partially Observed Systems
+In previous chapters, we explored control and learning techniques in fully observed linear dynamical systems, where the system state is directly accessible to the controller.
+However, in many real-world scenarios—such as robotics, economics, and signal processing—the state is only partially observed, requiring the controller to infer system dynamics from noisy or incomplete observations.
+This chapter focuses on partially observed LDS (PO-LDS), where the state is hidden, and only a linear projection of it is observed.
+For the remainder of this part of the text, we restrict ourselves to time-invariant systems, and refer to the bibliographic section for extensions.
+The system follows the general form:
+$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}$
+$\mathbf{y}_{t}$ $=C\mathbf{x}_{t}.$
+Here, $\mathbf{x}_{t}$ represents the hidden system state, $\mathbf{u}_{t}$ the control input, $\mathbf{y}_{t}$ the observed output, and $\mathbf{w}_{t}$ the disturbance.
+Often, partially observed LDSs are defined with an additional noise term corrupting the observations.
+Exercise 12.6 asks the reader to prove that even in such circumstances we may assume the previously stated form for partially observed systems without loss of any generality.
+Unlike in the fully observed setting, where controllers can react based on the full state, partial observability introduces fundamental limitations on policy design.
+In the partial observation setting some of the control policy classes from chapter 6 are not well defined. We proceed to define policy classes and describe the relationship between them.
+Not all control policies designed for fully observed systems extend naturally to partially observed settings.
+In this chapter, we systematically explore different policy classes, analyzing their expressivity, computational tractability, and suitability for PO-LDS.
+  • Linear Observation-Action Controllers (LOACs): These controllers directly map observations to actions but are highly inexpressive, as they ignore past observations and fail to capture hidden state dependencies.
+  • Generalized Linear Observation-Action Controllers (GOACs): These extend LOACs by incorporating past history of observations, significantly improving expressiveness.
+  • Linear Dynamic Controllers (LDCs): These policies model internal system dynamics, making them more powerful than simple observation-based policies.
+  • Disturbance Response Controllers (DRCs): Introduced as an alternative to state-based policies, DRCs directly act on a counterfactual signal—the “Nature’s y’s”—which captures the system’s response had no control been applied.
+We show that DRCs can approximate LDCs, making them a powerful convex relaxation.
+We conclude by analyzing the relationships between these policy classes, showing how they fit into the broader framework of nonstochastic control with partial observation.
+This chapter lays the groundwork for online learning and regret minimization in PO-LDS, which we explore in the next chapter.
+Figure 12.1: Schematic relationship between linear Disturbance-Response Control policies and Linear Dynamical Control policies of increasing dimensions (or number of parameters) for linear time invariant systems.
+12.1 Linear Observation Action Controllers
+Linear Observation-action controllers are the most natural generalization of linear controllers to the partially observed case, and are defined as follows.
+Definition 12.1.
+A linear observation-action controller $K\in{\mathbb{R}}^{d_{u}\times d_{y}}$ is a linear operator mapping observation to control as $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}$.
+We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers bounded in Fronbenius norm by $\|K\|\leq\gamma$, that are also $\gamma$-stabilizing for given LDS.
+While this class of policies is natural to consider, it turns out to be very inexpressive.
+Even for optimal control with stochastic noise, the optimal policy cannot be approximated by a linear observation-action policy (see Exercise 12.6).
+Therefore we proceed to define more expressive and useful policies.
+However, LOACs will be useful in simplifying proofs later on.
+12.1.1 Generalized Linear Observation Action Controllers
+We can generalize linear observation-action controllers to apply to a history of past observations.
+This additional generality, of taking a history of observations, greatly increases the expressive power of this policy class.
+The formal definition is below.
+Definition 12.2.
+Generalized linear observation-action controllers parameterized by matrices $K_{0:h}$ choose a control input as
+$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}K_{i}\mathbf{y}_{t-i}.$
+Denote by $\Pi^{G}_{h,\gamma}$ be that class of all such policies satisfying $\sum_{i=0}^{h}\|K_{i}\|\leq\gamma$ that are also $\gamma$-stabilzing.
+Generalized observation-action policies can be seen as a special case of observation-action controllers using the technique of dimension lifting as follows.
+Consider any partially-observed system $(A,B,C)$ and perturbation sequence $\mathbf{w}_{t}$.
+Consider the lifted system $(\tilde{A},\tilde{B},\tilde{C},\tilde{D})$ given by
+$\mathbf{z}_{t+1}=\tilde{A}\mathbf{z}_{t}+\tilde{B}\mathbf{u}_{t}+\tilde{D}\mathbf{w}_{t},\quad\tilde{\mathbf{y}}_{t}=C\mathbf{z}_{t},$
+where
+$\tilde{A}=\begin{array}[]{|c|c|c|c|c|}\hline\cr A&0&\ldots&0&0\\ \hline\cr I&0&\ldots&0&0\\ \hline\cr 0&I&\ldots&0&0\\ \hline\cr\vdots&\vdots&\cdots&\vdots&\vdots\\ \hline\cr 0&0&\cdots&I&0\\ \hline\cr\end{array}\ \ ,\ \ \tilde{B}=\begin{array}[]{|c|}\hline\cr B_{t}\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}\ \ ,$
+$\ \ \tilde{C}=\begin{array}[]{|c|c|c|c|}\hline\cr C&C&\dots&C\\ \hline\cr\end{array}\ \ ,\ \ \tilde{D}=\begin{array}[]{|c|}\hline\cr I\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}.$
+The following lemma relates the original and lifted system (for a proof see exercise 12.6).
+Lemma 12.3.
+For any sequence of control inputs $\mathbf{u}_{t}$, it holds that $\tilde{\mathbf{y}}_{t}=[\mathbf{y}_{t},\mathbf{y}_{t-1},\dots\mathbf{y}_{t-h}]$.
+The lemma above shows that generalized linear observation-action controllers are a special case of simple observation-action controllers such that $\mathbf{u}_{t}^{\pi}=[K_{0}\dots K_{h}]\tilde{\mathbf{y}}_{t}$ for the lifted system.
+12.2 Linear Dynamic Controllers
+The most well-known class of controllers for partially observable linear dynamical systems is that of linear dynamical controllers due to its connection to Kalman filtering.
+Definition 12.4 (Linear Dynamic Controllers for Partially Observed Systems).
+A linear dynamic controller $\pi$ has parameters $\left(A_{\pi},{B}_{\pi},C_{\pi}\right)$, and chooses a control at time $t$ as
+$\mathbf{u}_{t}^{\pi}=C_{\pi}\mathbf{s}_{t}\ ,\ \mathbf{s}_{t}=A_{\pi}\mathbf{s}_{t-1}+B_{\pi}\mathbf{y}_{t}.$
+We denote by $\Pi^{LDC}_{\gamma}$ the class of all LDC that are $\gamma$-stabilizing as per Definition 2.4 and exercise 4.6, with parameters $\frac{\kappa}{\delta}\leq\gamma$ that satisfy
+$\forall i\ ,\ \|A_{\pi}^{i}\|\leq\kappa(1-\delta)^{i}\ ,\ (\|B_{\pi}\|+\|C_{\pi}\|)\leq\kappa.$
+Note that LDCs are $\varepsilon$-approximated by generalized observation-action controllers, namely GOACs. The proof is left as Exercise 12.6.
+Lemma 12.5.
+Consider a partially-observed time-invariant linear dynamical system $\{A,B,C\}$.
+The class $\Pi^{G}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{LDC}_{\gamma}$ for $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right),\gamma^{\prime}=2\gamma^{3}$.
+12.3 Disturbance Response Controllers
+The most important policy class in the context of nonstochastic control is that of Disturbance Response Controllers, or DRC.
+These policies act on a signal called Nature’s y’s, denoted $\mathbf{y}^{\mathbf{nat}}_{t}$, which is the would-be observation at time $t$ assuming all inputs to the system were zero from the beginning of time.
+To simplify presentation, we only consider linear time invariant linear dynamical systems that are stable.
+This allows us to ignore the stabilizing controller component of disturbance action control and greatly simplifies the exposition of the main ideas.
+Extension to time-varying systems is explored in the exercises.
+Definition 12.6 (Disturbance Response Controller).
+A disturbance response policy for stable systems $\pi_{M_{1:h}}$ is parameterized by the matrices $M_{0:h}=[{M}_{0},\ldots,{M}_{h}]$.
+It outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
+$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}{M}_{i}\mathbf{y}^{\mathbf{nat}}_{t-i},$
+where $\mathbf{y}^{\mathbf{nat}}_{t}$ is given by the system response to the zero control at time $t$, i.e.
+$\mathbf{x}^{\mathbf{nat}}_{t+1}$ $=A\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}=\sum_{i=0}^{t}A^{i}\mathbf{w}_{t-i}\ ,\ \mathbf{y}^{\mathbf{nat}}_{t+1}=C\mathbf{x}^{\mathbf{nat}}_{t+1}.$
+Denote by $\Pi^{R}_{h,\gamma}$ the set of all disturbance response policies that are $\gamma$-stabilizing 111for example, by having additional linear controllers $K_{1:T}$ that are $\gamma$-stabilizing linear controllers for the system and
+$\sum_{i=0}^{h}\|M_{i}\|\leq\gamma.$
+If the perturbations are all zero, seemingly DRC policies produce a zero control. Therefore we could add a constant offset to the control, that can be incorporated as in Definition 6.5.
+The main advantage of this policy class, as opposed to previously considered policies in optimal control, is that they yield convex parametrization.
+Thus, this policy class allows for provably efficient optimization and algorithms that we explore in the next chapter.
+We conclude with the following important lemma on an expression that is important to compute nature’s y’s. The proof of this lemma follows directly from the definition and is left as exercise 12.6.
+Lemma 12.7.
+The sequence $\{\mathbf{y}^{\mathbf{nat}}_{t}\}$ can be iteratively computed using only the sequence of observations as
+$\mathbf{y}^{\mathbf{nat}}_{t}=\mathbf{y}_{t}-C\sum_{i=1}^{t}A^{i}B\mathbf{u}_{t-i}$
+Alternatively, the above computation can be carried out recursively as stated below starting with $\mathbf{z}_{0}=\mathbf{0}$.
+$\mathbf{z}_{t+1}$ $=A\mathbf{z}_{t}+B\mathbf{u}_{t},$
+$\mathbf{y}^{\mathbf{nat}}_{t}$ $=\mathbf{y}_{t}-C\mathbf{z}_{t}.$
+We proceed to show that that the class of DRC approximates the class of LDC.
+12.3.1 Expressivity of Disturbance Response Controllers
+A DRC always produces controls that ensure the system remains state-bounded. This can be shown analogously to Lemma 6.6 (see Exercise 12.6).
+While DRCs are contained inside the policy class of DACs, they are still very expressive.
+Note that a DAC policy can not be executed in a partially observed system from observations alone; even the knowledge of the entire observation sequence $\mathbf{y}_{t}$ may not uniquely determine $\mathbf{w}_{t}$.
+We have already seen that LDCs are approximately captured by GOACs in Lemma 12.5.
+It can be shown that the class of DRCs with appropriate parameters $\varepsilon$-approximates the class of GOACs for linear time-invariant systems.
+Note that this observation in conjunction with Lemma 12.5 implies that DRCs also approximate LDCs.
+To demonstrate that DRCs approximate GOACs, we instead prove that LOACs, a weaker class of policies, are $\varepsilon$-approximated by DRCs.
+Nevertheless, this observation is sufficient to conclude the stronger claim, given that GOACs can be written as LOACs on a lifted system as noted in Lemma 12.3, and simultaneously that a DRC with history $H^{\prime}$ on the lifted system can be written as a DRC with history $H^{\prime}+h$ over the original system.
+Lemma 12.8.
+Consider a $\gamma$-stable time-invariant linear dynamical system $(A,B,C)$.
+The class $\Pi^{R}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{L}_{\gamma}$ for parameters $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right)$, $\gamma^{\prime}=O(\gamma^{6})$.
 Proof.
-Let $\mathbf{x}^{\star}\in\operatorname*{\arg\min}_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})$. Define $\nabla_{t}\triangleq\nabla f_{t}(\mathbf{x}_{t})$. By convexity
-$f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{x}^{\star})\leq\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$ (A.2)
-We first upper-bound $\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$ using the update rule for $\mathbf{x}_{t+1}$ and the Pythagorean Theorem, we have the projections only decrease the distance to the set $\mathcal{K}$:
-$\|\mathbf{y}_{t+1}-\mathbf{x}^{\star}\|^{2}\ =\ \left\|\mathop{\Pi}_{\mathcal{K}}(\mathbf{x}_{t}-\eta_{t}\nabla_{t})-\mathbf{x}^{\star}\right\|^{2}\leq\left\|\mathbf{x}_{t}-\eta_{t}\nabla_{t}-\mathbf{x}^{\star}\right\|^{2}.$ (A.3)
-Hence,
-$\|\mathbf{y}_{t+1}-\mathbf{x}^{\star}\|^{2}$$\leq$ $\ \|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}+\eta_{t}^{2}\|\nabla_{t}\|^{2}-2\eta_{t}\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$
-$2\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$$\leq$ $\ \frac{\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$ (A.4)
-Summing (A.2) and (A.1) from $t=1$ to $T$, and setting $\eta_{t}=\frac{D}{G\sqrt{t}}$ (with $\frac{1}{\eta_{0}}\triangleq 0$):
-$2\left(\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{x}^{\star})\right)\leq 2\sum_{t=1}^{T}\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$
-$\leq\sum_{t=1}^{T}\frac{\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|^{2}}{\eta_{t}}+G^{2}\sum_{t=1}^{T}\eta_{t}$
-$\leq\sum_{t=1}^{T}\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)+G^{2}\sum_{t=1}^{T}\eta_{t}$ $\frac{1}{\eta_{0}}\triangleq 0,$
-$\|\mathbf{x}_{T+1}-\mathbf{x}^{*}\|^{2}\geq 0$
-$\leq D^{2}\sum_{t=1}^{T}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)+G^{2}\sum_{t=1}^{T}\eta_{t}$
-$\leq D^{2}\frac{1}{\eta_{T}}+G^{2}\sum_{t=1}^{T}\eta_{t}$ telescoping series
-$\leq 3DG\sqrt{T}.$
-The last inequality follows since $\eta_{t}=\frac{D}{G\sqrt{t}}$ and $\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\leq 2\sqrt{T}$. ∎
-The online gradient descent algorithm is straightforward to implement, and updates take linear time given the gradient.
-However, the projection step may in general take significantly longer via convex optimization.
+Recall that for the fully observed case the state-action pairs produced by execution of a linear policy can be written as linear functions of the perturbations.
+It is this observation that allowed us to prove that DACs capture linear policies for fully observed systems.
+We prove an analogue for partially observed systems: The observation-action pairs produced by the execution of LOACs can be written as linear functions of nature’s $y$’s.
+Thereafter, we argue that parameterizing the controllers to be linear in nature’s $y$’s (as DRCs do) allows us to encompass LOACs.
+We introduce an augmented partially observed linear system that helps us write the observation-action pairs produced by a LOAC as a linear function of nature’s $y$’s.
+Crucially, the new system has the same nature’s $y$’s as the original system.
+Lemma 12.9.
+For any partially-observed system $(A,B,C)$ and perturbation sequence $\mathbf{w}_{t}$ producing observations $\mathbf{y}_{t}$, construct a lifted system $(\tilde{A},\tilde{B},\tilde{C})$ given by
+$\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t+1}\\ \Delta_{t+1}\end{bmatrix}=\begin{bmatrix}A&0\\ 0&A\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta_{t}\end{bmatrix}+\begin{bmatrix}0\\ B\end{bmatrix}\mathbf{u}_{t}+\begin{bmatrix}I\\ 0\end{bmatrix}\mathbf{w}_{t},\quad\tilde{\mathbf{y}}_{t}=\begin{bmatrix}C&C\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta_{t}\end{bmatrix}.$
+It holds uniformly for all time $t$ that $\widetilde{\mathbf{y}^{\mathbf{nat}}_{t}}={\mathbf{y}^{\mathbf{nat}}_{t}}$ and further for any sequence of control inputs $\mathbf{u}_{t}$ that $\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}$ and $\mathbf{x}_{t}=\mathbf{x}^{\mathbf{nat}}_{t}+\Delta_{t}$.
+Substituting $\mathbf{u}_{t}^{\pi}=K\mathbf{y}^{\pi}_{t}=KC(\mathbf{x}^{\mathbf{nat}}_{t}+\Delta^{\pi}_{t})$, the state-action sequence produced by a linear policy $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}^{\pi}$ can be written as
+$\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t+1}\\ \Delta^{\pi}_{t+1}\end{bmatrix}=\begin{bmatrix}A&0\\ BKC&A+BKC\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta^{\pi}_{t}\end{bmatrix}+\begin{bmatrix}I\\ 0\end{bmatrix}\mathbf{w}_{t},\quad\begin{bmatrix}\mathbf{y}^{\pi}_{t}\\ \mathbf{u}^{\pi}_{t}\end{bmatrix}=\begin{bmatrix}C&C\\ KC&KC\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta^{\pi}_{t}\end{bmatrix},$
+or using that $\mathbf{y}^{\mathbf{nat}}_{t}=C\mathbf{x}^{\mathbf{nat}}_{t}$, equivalently as
+$\Delta^{\pi}_{t+1}=(A+BKC)\Delta^{\pi}_{t}+BK\mathbf{y}^{\mathbf{nat}}_{t},\quad\mathbf{y}^{\pi}_{t}=C\Delta^{\pi}_{t}+\mathbf{y}^{\mathbf{nat}}_{t},\quad\mathbf{u}^{\pi}_{t}=KC\Delta_{t}+K\mathbf{y}^{\mathbf{nat}}_{t}.$
+Now, define $M_{i}=KC(A+BKC)^{i-1}BK$ for $i>0$ and $M_{0}=K$, let $\mathbf{y}_{t},\mathbf{u}_{t}$ be the observation-action sequence produced by this DRC. We have
+$\mathbf{u}^{\pi}_{t}$ $=K\mathbf{y}^{\mathbf{nat}}_{t}+KC\Delta^{\pi}_{t}$
+$=K\mathbf{y}^{\mathbf{nat}}_{t}+KCBK\mathbf{y}^{\mathbf{nat}}_{t-1}+KC(A+BKC)\Delta^{\pi}_{t-1}$
+$=K\mathbf{y}^{\mathbf{nat}}_{t}+\sum_{i>1}KC(A+BKC)^{i-1}BK\mathbf{y}^{\mathbf{nat}}_{t-i}$
+$=\sum_{i=0}^{h^{\prime}}M_{i}\mathbf{y}^{\mathbf{nat}}_{t-i}+\sum_{i>H^{\prime}}KC(A+BKC)^{i-1}BK\mathbf{y}^{\mathbf{nat}}_{t-i}$
+$=\mathbf{u}_{t}+Z_{t}.$
+Given $\gamma$-stabilizability of the LOAC $K$ and $\gamma$-intrinsic stability of the of the system $(A,B,C)$, we have that for some $\kappa,\delta$ suc that $\frac{\kappa}{\delta}\leq\gamma$ that
+$\|Z_{t}\|$ $\leq\sum_{i>h^{\prime}}\|KC(A+BKC)^{i-1}BK\|\|\mathbf{y}^{\mathbf{nat}}_{t-i}\|$
+$\leq\sum_{i>h^{\prime}}\|KC(A+BKC)^{i-1}BK\|\left\|\sum_{j>1}CA^{j-1}B\mathbf{w}_{t-i-j}\right\|$
+$\leq\frac{\kappa^{5}}{\delta}\sum_{i>h^{\prime}}\|(A+BKC)^{i-1}\|\leq\frac{\kappa^{6}\gamma}{\delta}(1-\frac{1}{\delta})^{h^{\prime}}$
+Therefore, $\|\mathbf{u}_{t}-\mathbf{u}^{\pi}_{t}\|\leq\varepsilon$ by the appropriate choice of $H^{\prime}$. Once again using the intrinsic stability of the LTI $(A,B,C)$ , we have
+$\|\mathbf{y}^{\pi}_{t}-\mathbf{y}_{t}\|$ $=\left\|\sum_{i=1}^{t}CA^{i-1}B(\mathbf{u}^{\pi}_{t-i}-{\mathbf{u}}_{t-i})\right\|$
+$\leq\left\|\sum_{i=0}^{t}CA^{i}B\right\|\max_{i\leq t}\|\mathbf{u}^{\pi}_{i}-{\mathbf{u}}_{i}\|\leq\frac{\kappa}{\delta}\varepsilon.$
+The approximation in terms of observations and controls implies $\varepsilon$-approximation of policy class for any Lipschitz cost sequence, concluding the proof. ∎
+12.4 Summary
+In this chapter, we extended our study of control to the partially observed setting, where the controller no longer has direct access to the system state but must instead rely on noisy or incomplete observations.
+This shift fundamentally changes the nature of the control problem, requiring the design of policies that can infer and act based on limited information.
+We introduced a hierarchy of policy classes, starting with Linear Observation-Action Controllers (OACs), which directly map observations to actions, and progressing to Generalized LOACs (GOACs) and Linear Dynamic Controllers (LDCs), which incorporate history to improve decision-making.
+This increasing expressiveness highlighted a key limitation: memoryless policies are insufficient for optimal control, making structured state-dependent controllers necessary.
+A major contribution of this chapter was the introduction of Disturbance Response Controllers (DRCs).
+Instead of explicitly estimating the system state, DRCs act directly on a counterfactual signal—the “Nature’s y’s”—capturing what the system would have observed in the absence of control inputs.
+We established that DRCs are a convex relaxation of LDCs, making them computationally attractive while maintaining strong expressive power.
+This result provides a principled approach to designing policies that are both efficient and robust in partially observed environments.
+This chapter also set the stage for online nonstochastic control under partial observability, which we will explore in the next chapter.
+By establishing policy expressiveness results and identifying convex approximations for optimal control, we have built the necessary framework for developing learning-based algorithms that operate effectively in uncertain and adversarial settings.
+12.5 Bibliographic Remarks
+The study of control in partially observed linear dynamical systems (PO-LDS) has a long history, with foundational contributions dating back to optimal estimation and control theories.
+This section provides an overview of key references related to the topics discussed in this chapter, including classical works, expressivity of policy classes, convex relaxations, and connections to online learning.
+The problem of control under partial observation was first rigorously formulated in the context of the Linear Quadratic Gaussian (LQG) regulator, which combines Kalman filtering with optimal control [AM79].
+The certainty equivalence principle, which asserts that optimal controllers can be designed assuming a perfect state estimate, was shown to have limitations in more general partially observed settings [Won68].
+This chapter introduced a hierarchy of policy classes, including Linear Observation-Action Controllers (OACs), Generalized OACs (GOACs), Linear Dynamic Controllers (LDCs), and Disturbance Response Controllers (DRCs).
+The concept of disturbance response controllers was inspired by parameterizations used in robust control, particularly the Youla-Kucera parameterization [YJB76].
+Convex relaxations have been a powerful tool for simplifying control design in PO-LDS.
+We have surveyed convex relaxations for learning in dynamical systems in the previous part of this book.
+Recently, spectral methods for control [AHL+18] have provided an alternative approach for convex relaxation, leveraging spectral filtering to bypass direct system identification.
+The Disturbance Response Parametrization was introduced in [SSH20] as an analogue of DAC for control of partially observed systems.
+It can be viewed as a succinct state-space representation of the Youla parameterization [YJB76].
+For time-varying dynamics, the power of various policy classes was contrasted in [MGSH21].
+12.6 Exercises
+12.1. Prove Lemma 12.7, stating that the sequence $\{\mathbf{y}^{\mathbf{nat}}_{t}\}$ can be iteratively computed using only the sequence of observations as
+$\mathbf{y}^{\mathbf{nat}}_{t}=\mathbf{y}_{t}-C\sum_{i=1}^{t}A^{i}B\mathbf{u}_{t-i}$
+Alternatively, the above computation can be carried out recursively as stated below starting with $\mathbf{z}_{0}=\mathbf{0}$.
+$\mathbf{z}_{t+1}$ $=A\mathbf{z}_{t}+B\mathbf{u}_{t},$
+$\mathbf{y}^{\mathbf{nat}}_{t}$ $=\mathbf{y}_{t}-C\mathbf{z}_{t}.$
+12.2. Consider a seemingly more general definition of a partially observed linear dynamical system, given below:
+$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t},$
+$\mathbf{y}_{t}$ $=C_{t}\mathbf{x}_{t}+\mathbf{v}_{t}.$
+Construct a linear dynamical system parameterized via matrices $(A^{\prime}_{t},B^{\prime}_{t},C^{\prime}_{t})$ and perturbations $\mathbf{w}^{\prime}_{t}$ with a hidden state $\mathbf{z}_{t}$ of the form
+$\mathbf{x}_{t+1}$ $=A^{\prime}_{t}\mathbf{z}_{t}+B^{\prime}_{t}\mathbf{u}_{t}+\mathbf{w}^{\prime}_{t},$
+$\mathbf{y}^{\prime}_{t}$ $=C^{\prime}_{t}\mathbf{z}_{t}$
+such that for any sequence of $\mathbf{u}_{t}$, $\mathbf{y}_{t}=\mathbf{y}^{\prime}_{t}$ holds uniformly over time.
+Additionally, verify that there exists such a construction of $\mathbf{z}_{t}$ where $\mathbf{w}^{\prime}_{t}$ are i.i.d.
+as long as $(\mathbf{w}_{t},\mathbf{v}_{t})$ are i.i.d.
+across time steps.
+Hint: Consider the choice $\mathbf{w}^{\prime}_{t}=D^{\prime}_{t}\begin{bmatrix}\mathbf{w}_{t}\\ \mathbf{v}_{t}\end{bmatrix}$ for some $D^{\prime}_{t}$.
+Also consider using a hidden state $\mathbf{z}_{t}$ with a larger number of coordinates than $\mathbf{x}_{t}$.
+12.3.
+  1. 1.
+    Prove Lemma 12.3.
+  2. 2.
+    Extend the lemma and prove it for time-varying linear dynamical systems.
+12.4. Prove Lemma 12.5.
+12.5. For any $\gamma$-stable partially observed system, prove that any DRC policy in $\Pi_{H,\gamma}^{D}$ is $\gamma^{2}$-stabilizing.
+12.6.
+Construct a partially observed linear dynamical system subject to Gaussian i.i.d.
+perturbations with quadratic costs such that simultaneously it is true that (a) there exists a GOAC policy that incurs zero (or at most constant) aggregate cost, (b) yet any LOAC policy incurs $\Omega(T)$ aggregate cost when run for $T$ time steps.
+Hint: Consider a 3-dimensional hidden state so that $u_{t}$ can be chosen as a function of $y_{t-1}$ to ensure zero cost at each time step.
+Chapter 13 Online Nonstochastic Control with Partial Observation
+In the previous chapter, we introduced Disturbance Response Controllers (DRCs) as a structured class of policies for controlling partially observed linear dynamical systems (PO-LDS).
+Unlike classical observer-based controllers, which rely on explicit state estimation, DRCs act directly on Nature’s y’s, the system’s counterfactual response in the absence of control.
+While DRCs offer a powerful alternative to state-based policies, they did not yet incorporate online learning, leaving open the question of how to adapt them dynamically in uncertain environments.
+This chapter takes the next step by introducing an online learning algorithm for optimizing DRCs in real time.
+We introduce the Gradient Response Controller (GRC), which efficiently updates DRC policies using online convex optimization techniques.
+The GRC algorithm enables controllers to:
+  • Learn from past disturbances and actions, refining policy parameters without requiring full system identification.
+  • Handle adversarial disturbances, extending the policy regret framework to the partially observed setting.
+  • Optimize control decisions dynamically, adapting to system behavior as new observations arrive.
+Unlike previous approaches that relied on fixed policy classes, GRC provides an adaptive control mechanism, bridging the gap between structured controllers and learning-based methods.
+A key focus of this chapter is extending policy regret minimization from the fully observed case to partial observability, introducing new challenges and solutions in learning-based control.
+To simplify the theoretical analysis, we assume system stability, allowing us to focus on the core learning problem without additional stabilization constraints.
+However, we discuss possible extensions to stabilizable systems in the Bibliographic Remarks section.
+13.1 The Gradient Response Controller
+Partial observation requires modification to the Gradient Perturbation Controller we have studied in the previous chapter.
+The Gradient Response Controller algorithm is described in figure (8), for linear time-invariant dynamical systems with partial observation.
+Similar to the GPC, the main idea is to learn the parametrization of a disturbance response control policy using known online convex optimization algorithms, such as online gradient descent.
+We henceforth prove that this parametrization is convex, allowing us to prove the main regret bound.
+In algorithm (8) we denote $\mathbf{y}_{t}({M}_{1:h})$ as the observation arising by playing DRC policy ${M}_{1:h}$ from the beginning of time, and same for $\mathbf{u}_{t}({M}_{1:h})$.
+Although it may be impossible to determine $\mathbf{w}_{t}$, especially if $C$ has fewer rows than columns, from the observations $\mathbf{y}_{t}$, $\mathbf{y}_{t}({M}_{1:h})$ can be determined from $\mathbf{y}^{\mathbf{nat}}_{t}$ alone by the application of Lemma 12.7, as
+$\mathbf{y}_{t}({M}_{1:h})=\mathbf{y}^{\mathbf{nat}}_{t}+C\sum_{i=1}^{t}A^{i}B\mathbf{u}_{t-i}({M}_{1:h}).$
+1:  Input: system $\{A,B,C\}$ , $h$, $\eta$, initialization $M_{0:h}^{1}\in\mathcal{K}$.
+2:  for $t$ = $1\ldots T$ do
+3:     Use Control $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=0}^{h}{M}_{i}^{t}\mathbf{y}^{\mathbf{nat}}_{t-i}$
+4:     Observe $\mathbf{y}_{t+1}$, compute for $t+1$:
+$\mathbf{y}^{\mathbf{nat}}_{t}=\mathbf{y}_{t}-C\sum_{i=0}^{t}A^{i}B\mathbf{u}_{t-i}.$
+5:     Construct loss $\ell_{t}({M}_{0:h})=c_{t}(\mathbf{y}_{t}({M}_{0:h}),\mathbf{u}_{t}({M}_{0:h}))$Update $M_{0:h}^{t+1}\leftarrow\prod_{\mathcal{K}}\left[{M}_{0:h}^{t}-\eta\nabla\ell_{t}({M}_{0:h}^{t})\right]$
+6:  end for
+Algorithm 8 Gradient Response Controller(GRC)
+The GRC algorithm comes with a finite time performance guarantee: namely it guarantees vanishing worst case regret vs.
+the best disturbance response control policy in hindsight.
+As per the relationship between disturbance response control and other policies studied in section 12.3.1, this implies vanishing regret vs.
+linear dynamical controllers.
+The formal statement is given as follows.
+Theorem 13.1.
+Assuming that
+  • a
+    The costs $c_{t}$ are convex, bounded and have bounded gradients w.r.t. both arguments $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$.
+  • b
+    The diameter of the constrained set for the parameters $M_{0:h}$ is bounded by $D$.
+  • c
+    The matrices $\{{A},{B},C\}$ have bounded $\ell_{2}$ norm.
+  • d
+    The linear dynamical system $\{{A},{B},C\}$ is $\gamma$-stable.
+Then the GRC algorithm (8) ensures that
+$\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}\left(\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{R}}\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right)\leq\mathcal{O}(LGD\gamma^{3}h\sqrt{T}).$
+Furthermore, the time complexity of each loop of the algorithm is polynomial in the number of system parameters and logarithmic in T.
+The proof is largely analogous to the regret bound of the GPC presented in Theorem 7.2.
+Proof.
+As per Theorem 6.1, the online gradient descent algorithm for convex loss functions ensures that the regret is bounded as a sublinear function of the number of iterations.
+However, to directly apply the OGD algorithm to our setting, we need to make sure that the following conditions hold true:
+  • The loss function should be a convex function in the variables ${M}_{0:h}$.
+  • We need to ensure that optimizing the loss function $\ell_{t}$ as a function of the parameters $M_{0:h}$ is similar to optimizing the cost $c_{t}$ as a function of the observation and control.
+First, we show that the loss function is convex w.r.t. to the variables ${M}_{0:h}$. This follows since the states and the controls are linear transformations of the variables.
+Lemma 13.2.
+The loss functions $\ell_{t}$ are convex in the variables ${M}_{0:h}$.
+Proof.
+Using $\mathbf{y}_{t}$ to denote $\mathbf{y}_{t}(M_{0:h})$ (and similarly for $\mathbf{u}_{t}$), the loss function $\ell_{t}$ is given by
+$\ell_{t}({M}_{0:h})=c_{t}(\mathbf{y}_{t},\mathbf{u}_{t}).$
+Since the cost $c_{t}$ is a convex function w.r.t. its arguments, we simply need to show that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{0:h}$. Using Lemma 12.7,
+$\mathbf{u}_{t}$ $=\sum_{i=0}^{h}M_{i}\mathbf{y}^{\mathbf{nat}}_{t-i},$
+$\mathbf{y}_{t}$ $=\mathbf{y}^{\mathbf{nat}}_{t}+C\sum_{i=1}^{t}A^{i}B\mathbf{u}_{t-i}$
+$=\mathbf{y}^{\mathbf{nat}}_{t}+C\sum_{i=1}^{t}A^{i}B\sum_{j=0}^{h}M_{j}\mathbf{y}^{\mathbf{nat}}_{t-i-j}$
+which is a linear function of the variables.
+Thus, we have shown that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ are linear transformations in ${M}_{0:h}$ and hence, the loss function is convex in ${M}_{0:h}$.
+∎
+Next, we use the following Lemma,
+Lemma 13.3.
+For all $t$, $|l_{t}(M^{t}_{0:h})-c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})|\leq\frac{L\gamma^{3}ChD}{\sqrt{T}}$.
+The regret bound of GRC can be derived from these lemmas as follows.
+Note that $G$ is an upper bound on the gradients of $\ell_{t}$ as a function of the parameters $M_{0:h}$, while $L$ is the Lipschitz constant of $c_{t}$ as a function of the state and control.
+$\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{R}}\sum_{t=1}^{T}c_{t}({\mathbf{y}}^{\pi}_{t},\mathbf{u}_{t}^{\pi}))$
+$\leq\sum_{t=1}^{T}l_{t}(M_{1:h}^{t})-\min_{\pi\in\Pi^{R}}l_{t}(M_{1:h}^{\pi})+T\times\frac{L\gamma^{3}ChD}{\sqrt{T}}$ Lemma 13.3
+$\leq 2GD\sqrt{T}+{\sqrt{T}L\gamma^{3}ChD}$ Theorem 6.1
+$\leq O(GDL\gamma^{3}h\sqrt{T}).$
+13.1.1 Proof of Lemma 13.3
+Proof.
+We first note that by the choice of step size by the Online Gradient Descent algorithm 3, as per Theorem 6.1, we have that $\eta=\frac{D}{G\sqrt{T}}$. Thus,
+$\|M_{0:h}^{t}-M_{0:h}^{t-i}\|\leq\sum_{s=t-i+1}^{t}\|M_{0:h}^{s}-M_{0:h}^{s-1}\|\leq i\eta G=\frac{iD}{\sqrt{T}}.$
+Now, we use the fact that the iterates of OGD $M_{1:h}^{t}$ move slowly to establish that $\mathbf{y}_{t}$ and $\mathbf{y}_{t}(M^{t}_{0:h})$ are close in value.
+Before that, by definition, we have $\mathbf{u}_{t}(M^{t}_{0:h})=\mathbf{u}_{t}$ since the sequence $\mathbf{y}^{\mathbf{nat}}_{t}$ is determined independently of the choice of the policy being executed.
+Due to the $\gamma$-stability of the underlying system, we have for any $i>0$, $\|C{A}^{i}B\|\leq\kappa(1-\delta)^{i}$ for some $\kappa>0,\gamma<1$.
+Using Lemma 12.7, we have for some $C>0$ that
+$\|\mathbf{y}_{t}(M^{t}_{0:h})-\mathbf{y}_{t}\|$ $=\left\|C\sum_{i=1}^{t}\left[A_{i-1}B(\mathbf{u}_{t-i}(M^{t}_{0:h})-\mathbf{u}_{t-i})\right]\right\|$
+$\leq\sum_{i=1}^{t}\left\|CA^{i-1}B\right\|\|(\mathbf{u}_{t-i}(M^{t}_{0:h})-\mathbf{u}_{t-i})\|$
+$\leq Ch\kappa\sum_{i=1}^{t}(1-\delta)^{i}\frac{i}{\sqrt{T}}\max_{j}\|\mathbf{y}^{\mathbf{nat}}_{j}\|$
+$\leq\frac{Ch\kappa D}{\delta^{2}\sqrt{T}}\max_{j}\|\mathbf{y}^{\mathbf{nat}}_{j}\|\leq\frac{ChD\gamma^{3}}{\sqrt{T}}$
+Here we used the fact that $\|\mathbf{y}^{\mathbf{nat}}_{t}\|\leq\gamma$, which is left as an exercise.
+The control depends only on the parameters of the current iteration and we have that $\mathbf{u}_{t}=\mathbf{u}_{t}(M_{1:h}^{t})$.
+By definition, $l_{t}(M_{0:h}^{t})=c_{t}(\mathbf{y}_{t}(M_{0:h}^{t}),\mathbf{u}_{t}(M_{0:h}^{t}))$.
+Thus, we have
+$l_{t}(M_{0:h}^{t})-c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$ $=c_{t}(\mathbf{y}_{t}(M_{0:h}^{t}),\mathbf{u}_{t}(M_{0:h}^{t}))-c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$
+$\leq L\cdot\|\mathbf{y}_{t}(M_{1:h}^{t})-\mathbf{y}_{t}\|\leq\frac{L\gamma^{3}ChD}{\sqrt{T}}.$
+13.2 Extension to Time Varying Systems
+The GRC algorithm can be extended to time-varying dynamics as we explain in this section.
+The linear dynamical system at time $t$ is denoted by system matrices $A_{t},B_{t},C_{t}$, and we assume that they are known to the controller.
+We further assume that these dynamical systems are all $\gamma$-stable.
+The algorithm is formally spelled out in figure 9.
+1:  Input: sequence $\{A_{t},B_{t},C_{t}\}$ , $h$, $\eta$, initialization $M_{0:h}^{1}\in\mathcal{K}$.
+2:  for $t$ = $1\ldots T$ do
+3:     Use Control $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=0}^{h}{M}_{i}^{t}\mathbf{y}^{\mathbf{nat}}_{t-i}$
+4:     Observe $\mathbf{y}_{t+1}$, compute for $t+1$:
+$\mathbf{y}^{\mathbf{nat}}_{t}=\mathbf{y}_{t}-C_{t}\sum_{i=0}^{t}\left[\prod_{j=0}^{i}A_{t-j}\right]B_{t-i}\mathbf{u}_{t-i}.$
+5:     Construct loss $\ell_{t}({M}_{0:h})=c_{t}(\mathbf{y}_{t}({M}_{0:h}),\mathbf{u}_{t}({M}_{0:h}))$Update $M_{0:h}^{t+1}\leftarrow\prod_{\mathcal{K}}\left[{M}_{0:h}^{t}-\eta\nabla\ell_{t}({M}_{0:h}^{t})\right]$
+6:  end for
+Algorithm 9 Gradient Response Controller(GRC)
+It is possible to prove a similar regret bound to the one we proved for linear time-invariant systems.
+An important step is to note that the observation $\mathbf{y}_{t}({M}_{1:h})$ can be determined from $\mathbf{y}^{\mathbf{nat}}_{t}$ alone by the application of Lemma 12.7, as
+$\mathbf{y}_{t}({M}_{1:h})=\mathbf{y}^{\mathbf{nat}}_{t}+C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}A_{t-j}\right]B_{t-i}\mathbf{u}_{t-i}({M}_{1:h}).$
+13.3 Conclusion
+In this chapter, we extended the framework of online nonstochastic control to the partially observed setting, where the controller must make decisions without direct access to the system state.
+Unlike the fully observed case, where state-feedback policies can be applied directly, partial observability introduces fundamental challenges, requiring policies that infer system behavior from past observations and actions.
+Building on the Disturbance Response Controller (DRC) policies introduced in the previous chapter, we developed the Gradient Response Controller (GRC), an online learning algorithm that optimizes DRC policies in real time.
+The results in this chapter demonstrate that adaptive control under partial observability is possible without explicit state estimation, bridging the gap between classical observer-based controllers and modern learning-based methods.
+By leveraging regret minimization techniques, we provide a robust and flexible approach that applies even in adversarial and uncertain environments.
+13.4 Bibliographic Remarks
+Online nonstochastic control with partial observability was studied in [SSH20], who defined the Gradient Response Controller as well as the policy class of Disturbance Response Controllers.
+The setting was further extended to changing linear dynamics in [MGSH21], where the expressive power of different policy classes was also studied.
+The analogue of online nonstochastic control in the stochastic noise setting is called the Linear Quadraic Gaussian problem, and it is further limited by assuming quadratic noise.
+This classical problem in optimal control was proposed in Kalman’s work [Kal60].
+Regret bounds for controlling unknown partially observed systems were studied in [LAHA20b, LAHA20c, MTR19].
+[LAHA20a], which uses DRC policies even in context of stochastic control, obtains rate-optimal logarithmic bounds for the LQG problem.
+Control with partial information and bandit feedback with semi-adversarial noise was considered in [SNH23].
+Tight rates for quadratic control with bandit feedback, in the fully adversarial model, were obtained in [SSNH24], and more general cost functions in the bandit setting were considered in [SL24].
+13.5 Exercises
+13.1. In this exercise we consider the computation of nature’s y’s for time varying systems. Let
+$\mathbf{x}^{\mathbf{nat}}_{t+1}$ $=A_{t}\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}=\sum_{i=0}^{t}\left[\prod_{j=0}^{i-1}A_{t-j}\right]\mathbf{w}_{t-i}=\Phi_{t}\mathbf{w}_{1:t}.$
+$\mathbf{y}^{\mathbf{nat}}_{t+1}$ $=C_{t+1}\mathbf{x}^{\mathbf{nat}}_{t+1}=C_{t+1}\Phi_{t}\mathbf{w}_{1:t}.$
+Here $\Phi_{t}$ is the linear operator such that $\Phi_{t}(t-i)=\prod_{j=0}^{i-1}A_{t-j}$.
+Prove that The sequence $\{\mathbf{y}^{\mathbf{nat}}_{t}\}$ can be iteratively computed using only the sequence of observations as
+$\mathbf{y}^{\mathbf{nat}}_{t}=\mathbf{y}_{t}-C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}A_{t-j}\right]B_{t-i}\mathbf{u}_{t-i}$
+Alternatively, the above computation can be carried out recursively as stated below starting with $\mathbf{z}_{0}=\mathbf{0}$.
+$\mathbf{z}_{t+1}$ $=A_{t}\mathbf{z}_{t}+B_{t}\mathbf{u}_{t},$
+$\mathbf{y}^{\mathbf{nat}}_{t}$ $=\mathbf{y}_{t}-C_{t}\mathbf{z}_{t}.$
+13.2.
+Consider a partially observed linear dynamical system that is $\gamma$-stable in the absence of control inputs.
+Prove that when a DRC policy $\pi\in\Pi^{R}_{H,\gamma}$ is executed on such a dynamical system, the resulting system is $\gamma^{2}$-stable.
+13.3. Prove that the optimal controller that minimizes the aggregate cost for a partially observed linear dynamical system assuming the perturbations are stochastic is a linear dynamical controller.
+13.4. Prove that the regret of the GRC controller for time varying linear dynamical systems is bounded by $O(\sqrt{T})$.
+Hint: Follow the proof of Theorem 13.1 and modify as required.

### v5

@@
-version 0.75
-[…]
-position=current page.north, angle=0, nodeanchor=north, vshift=-5mm, opacity=0.5, scale=1, contents=Draft; comments and suggestions welcome at nonstochastic.control@gmail.com
-Introduction to Online Control
-version 0.75
+version 0.9
+[Uncaptioned image] Introduction to Online Control
+version 0.9
 Elad Hazan    Karan Singh
+We are in a golden age for control and decision making.
+A proliferation of new applications including self-driving vehicles, humanoid robots, and artificially intelligent drones open a new set of challenges for control theory to address.
+Hazan and Singh have written the definitive book on the New Control Theory– non-stochastic control.
+The phrase “a paradigm shift” has become cliche from overuse, but here it is truly well deserved; the authors have revisited the foundations by focusing on building controllers that perform nearly as well as if they knew future disturbances in advance, rather than relying on probabilistic or worst-case models.
+The non-stochastic control approach has extended one of the most profound ideas in mathematics of the 20th century, online (no-regret) learning, to master sequential decision making with continuous actions.
+This leads to high performance in benign environments and resilience in adversarial ones.
+The book, authored by pioneers in the field, presents both foundational concepts and the latest research, making it an invaluable resource.
+Drew Bagnell
+Yann LeCun
+To my family: Dana, Hadar, Yoav, Oded and Deluca,
+—EH
+To my parents– A & V.
+—KS
+© [2025] Elad Hazan and Karan Singh
 Preface
@@
 This book was a journey that started in Princeton at the Google DeepMind Lab in the summer of 2018.
-At that time Sham Kakade visited the group, which consisted of the authors as well as Naman Agarwal, Brian Bullins, Xinyi Chen, Cyril Zhang and Yi Zhang, henceforth the gang.
+At that time Sham Kakade visited the group, which consisted of the authors together with Naman Agarwal, Brian Bullins, Xinyi Chen, Cyril Zhang, and Yi Zhang, henceforth the gang.
 The gang worked on various aspects of optimization, reinforcement learning, and dynamical systems.
@@
 The text was further enhanced by students of theoretical machine learning taught at Princeton University in the fall of 2024-2025.
-The following students were particularly helpful making insightful comments and in finding various typos and suggesting improvements: Gon Buzaglo, Anand Brahmbhatt, Sofiia Druchyna (and many more we wish to announce very soon).
+The following colleagues and students were particularly helpful, making insightful comments and in finding various typos and suggesting improvements: Gon Buzaglo, Anand Brahmbhatt, Sofiia Druchyna, and Jonathan Pillow (and many more we wish to announce very soon).
 Throughout the writing of this book, we engaged in extensive discussions with ChatGPT, an AI assistant.
@@
 $\mathbf{1}_{A}$ the indicator function, equals one if $A$ is true, else zero
-Algebra, Geometry and Calculus
+Linear Algebra, Geometry and Calculus
 ${\mathbb{R}}^{d}$ $d$ dimensional Euclidean space
@@
 ${\bf Tr}(A)$ trace of matrix $A$
+$A\succeq 0$ $A$ is positive semi-definite, $\forall v\ ,\ v^{\top}Av\geq 0$
+$A\succeq B$ $A-B$ is positive semi-definite
 Control
@@
 1.1 What is This Book About?
-Control theory is the engineering science of manipulating physical systems to achieve desired functionality.
-Its history is centuries old, and it spans rich areas from the mathematical sciences.
+Control theory is the engineering science of manipulating physical systems to achieve the desired functionality.
+Its history is centuries old and spans rich areas from the mathematical sciences.
 A brief list includes differential equations, operator theory and linear algebra, functional analysis, harmonic analysis, and more.
-However, it mostly does not concern modern algorithmic theory and computational complexity from computer science.
-The advance of the deep learning revolution has brought to light the power and robustness of algorithms that are based on gradient methods.
-The following quote is attributed to deep learning pioneer Yann Lecun during a Barbados workshop on deep learning theory circa Feb 2020,
+However, it mostly does not concern modern algorithmic theory or computational complexity from computer science.
+The advance of the deep learning revolution has brought to light the power and robustness of algorithms based on gradient methods.
+The following quote is attributed to deep learning pioneer Yann Lecun during a Barbados workshop on deep learning theory circa February 2020.
 “Control = reinforcement learning with gradients.”
@@
 This is a deep and difficult question that we fall short of answering.
-We do, however, propose a different way of answering this question: we define the control problem from the point of view of an online player in a repeated game.
+However, we do propose a different way of answering this question: we define the control problem from the point of view of an online player in a repeated game.
 This gives rise to a simple but powerful framework called nonstochastic control theory.
@@
 One of its most critical components was introduced by the Scottish inventor James Watt: the centrifugal governor, which is amongst the earliest examples of a control mechanism.
-A major source of losses in efficiency of earlier designs was the difficulty in regulating the engine.
-Too much or too little steam could easily result in engine failure.
+A major source of efficiency losses of the earlier designs was the difficulty in regulating the engine.
+Too much or too little steam could easily cause engine failure.
 The centrifugal governor was created to solve this problem.
-Figure 1.1: A centrifugal governor. Source: Wikipedia
+Figure 1.1: A centrifugal governor.
 The centrifugal governor consists of a shaft connected to two metal balls that are allowed to rotate.
-Increased shaft speed causes the balls to spin faster, and the resulting centrifugal force causes the balls to lift.
+The increased shaft speed causes the balls to spin faster and the resulting centrifugal force causes the balls to lift.
 The balls are in turn attached to a pin that controls the amount of steam entering the engine, so as the balls spin faster, the pin lifts and limits the amount of fuel that can enter the system.
-On the other hand, if the balls spin slower, the pin lowers, allowing more steam to enter the system.
-This provides a natural, direct and proportional control to the system.
-The impact of the centrifugal governor on the steam engine, and in turn on the the industrial revolution, illustrates the importance of the field of control in engineering and the sciences.
-Since the 18th century our world has changed quite a bit, and the importance of control has only grown.
-The challenges of controlling systems has moved from the steam engine to design of aircraft, robotics and autonomous vehicles.
-The importance of designing efficient and robust methods remains at least as crucial as in the industrial revolution.
+However, if the balls spin slower, the pin is lowered, allowing more steam to enter the system.
+This provides a natural, direct, and proportional control to the system.
+The impact of the centrifugal governor on the steam engine, and in turn on the industrial revolution, illustrates the importance of the field of control in engineering and the sciences.
+Since the 18th century, our world has changed quite a bit, and the importance of control has only grown.
+The challenges of controlling systems have moved from the steam engine to the design of aircraft, robotics, and autonomous vehicles.
+The importance of designing efficient and robust methods remains as crucial as in the industrial revolution.
 1.3 Formalization and Examples of a Control Problem
-With the physical example of the centrifugal governor in mind, we give a mathematical definition of a control problem. The first step is to formally define a dynamical system below.
+With the physical example of the centrifugal governor in mind, we give a mathematical definition of a control problem. The first step is to formally define a dynamical system in the following way.
 Definition 1.1 (Dynamical system).
-A dynamical system from starting state $(\mathbf{x}_{1},\mathbf{y}_{1})$ is given by the following equations:
+A dynamical system evolving from an initial state $\mathbf{x}_{1}$ is given by the following equations:
 $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g_{t}(\mathbf{x}_{t}).$
 Here $\mathbf{x}_{t}\in{\mathbb{R}}^{d_{x}}$ is the state of the system, $\mathbf{y}_{t}\in{\mathbb{R}}^{d_{y}}$ is the observation, $\mathbf{u}_{t}\in{\mathbb{R}}^{d_{u}}$ is the control input, and $\mathbf{w}_{t}\in{\mathbb{R}}^{d_{w}}$ is the perturbation.
-The function $g_{t}$ is a mapping from the state to the observed state, and the function $f_{t}$ is the transition function, given current state $\mathbf{x}_{t}$, control input $\mathbf{u}_{t}$ and perturbation $\mathbf{w}_{t}$.
+The function $g_{t}$ is a mapping from the state to the observed state, and the function $f_{t}$ is the transition function, given the current state $\mathbf{x}_{t}$, the input control $\mathbf{u}_{t}$, and the perturbation $\mathbf{w}_{t}$.
 The subscript indicates the relevant quantities at time step $t$.
-Given a dynamical system, there are several possible objectives we can formulate.
+Given a dynamical system, there are several possible objectives that we can formulate.
 We give a generic formulation below that captures many common control formulations, notably the optimal control problem and the robust control problem.
@@
 $s.t.$ $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g_{t}(\mathbf{x}_{t}).$
-At this point, there are many aspects of the generic problem that are left unspecified. These include:
+At this point, many aspects of the generic problem have not been specified. These include:
   • Are the transition functions and observation functions known or unknown to the controller?
@@
   • Are the cost functions known or unknown ahead of time?
-Different theories of control differ by the answers they give to the questions above, and these in turn give rise to different methods.
+Different theories of control differ by the answers they give to the above questions, and these in turn give rise to different methods.
 Instead of postulating a hypothesis about how our world should behave, we proceed to give an intuitive example of a control problem.
-We can then reason about the different questions, and motivate the nonstochastc control setting.
+We can then reason about the different questions and motivate the nonstochastc control setting.
 1.3.1 Example: Control of a Medical Ventilator
@@
 The ventilator connects the trachea in the lungs to a central pressure chamber, which helps maintain positive-end expiratory pressure (PEEP) to prevent the lungs from collapsing and simulate healthy breathing in patients with respiratory problems.
-The ventilator needs to take into account the structure of the lung to determine the optimal pressure to induce.
+The ventilator must take into account the structure of the lung to determine the optimal pressure to induce.
 Such structural factors include compliance, or the change in lung volume per unit pressure, and resistance, or the change in pressure per unit flow.
@@
 $v_{t+1}=v_{t}+\Delta_{t}u_{t},$
-where $p_{t}$ is the observed pressure, $v_{t}$ is the volume of the lung, $u_{t}$ is the input flow (control), and $\Delta_{t}$ is the unit of time measured by the device precision.
+where $p_{t}$ is the observed pressure, $v_{t}$ is the volume of the lung, $u_{t}$ is the input flow (control) and $\Delta_{t}$ is the unit of time resolution of the device.
 $C_{0}$, $C_{1}$, and $C_{2}$ are constants that depend on the particular system/lung-ventilator pair.
-The goal of ventilator control is to regulate pressure according to some predetermined pattern.
+The goal of ventilator control is to regulate the pressure according to some predetermined pattern.
 As a dynamical system, we can write $v_{t+1}=f(v_{t},u_{t},w_{t})$, $p_{t}=g(v_{t})$, and we can define the cost function to be $c_{t}(p_{t},u_{t})=|p_{t}-p_{t}^{*}|^{2},$ where $p_{t}^{*}$ is the desired pressure at time $t$.
@@
 1.4 Simple Control Algorithms
-The previous section detailed the importance of the control problem on the example of a critical care medical device.
-In this section we detail methods for controlling a medical ventilator, or indeed any physical device that can be modelled as a dynamical system.
-The following methods are simplistic but still widely used.
-For example, the PID algorithm is still used in the majority of medical ventilators today.
-In later sections we describe a modern optimization-based framework that yields more sophisticated algorithms that are capable of better performance.
+The previous section detailed the importance of control problems, specifically through the example of a critical care medical device.
+In this section, we detail methods for controlling a medical ventilator, or indeed any physical device that can be modeled as a dynamical system.
+The following methods are simplistic, but still widely used.
+For example, the PID algorithm is still used in most medical ventilators today.
+In later sections, we describe a modern optimization-based framework that yields more sophisticated algorithms that are capable of better performance.
 1.4.1 The Bang-Bang Controller
 Bang-bang control is one of the simplest control strategies.
-Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a pre-specified dynamic range $[x_{\min},x_{\max}]$ (that can vary with time), and where the scalar control $u_{t}$ input must at all times be within the interval $[u_{\min},u_{\max}]$, that is assumed to contain zero for simplicity.
+Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a prespecified dynamic range $[x_{\min},x_{\max}]$ (which can vary over time) and where the scalar control input $u_{t}$ must at all times be within the interval $[u_{\min},u_{\max}]$, which is assumed to contain zero for simplicity.
 A bang-bang controller chooses the control input at time $t$ as
@@
 For example, in the context of mechanical ventilation, one could choose $[p_{t}^{*}-\Delta p,p_{t}^{*}+\Delta p]$ as a reasonable region around the prescribed pressure $p^{*}_{t}$ as a desirable range for $p_{t}$.
-A bang-bang controller in such a case would let the oxygen in at the maximum possible rate whenever the pressure drops below $p^{*}_{t}-\Delta p$.
+In such a case, a bang-bang controller would allow oxygen to enter at the highest possible rate whenever the pressure drops below $p^{*}_{t}-\Delta p$.
 This discussion highlights two unsatisfactory aspects of the bang-bang control scheme: one, it ignores the cost associated with acting with the extreme values of the control input; two, it produces oscillatory and rapidly changing behavior in the control state.
@@
 for some coefficients $A_{i}$, where we consider only the $k+1$ most recent states. One of the most useful special cases of this linear family is the PID controller.
-A special sub-family of linear controllers is that of linear controllers that optimize over a pre-defined basis of historical observations. More precisely,
+A special subfamily of linear controllers is that of linear controllers that optimize on a predefined basis of historical observations. More precisely,
   1. 1.
@@
     Derivative Control: $\mathbf{u}_{t}=\gamma(\mathbf{x}_{t}-\mathbf{x}_{t-1}).$
-The PID controller, named after the nature of the three components that constitute it, specifies three coefficients $\alpha,\beta,\gamma$ to generate a control signal.
-In that sense it is extremely sparse, which from a learning theoretic perspective is a good indication of generalization.
-Indeed, the PID controller is the method of choice for numerous applications in engineering, including control of medical ventilators.
+The PID controller, named after the nature of the three components that make up it, specifies three coefficients $\alpha,\beta,\gamma$ to generate a control signal.
+In that sense, it is extremely sparse, which from a learning-theoretic perspective is a good indication of generalization.
+In fact, the PID controller is the method of choice for numerous engineering applications, including control of medical ventilators.
 1.5 Classical Theory: Optimal and Robust Control
-Classical theories of control differ over the generic control problem formulation 1.2 primarily in one significant aspect: the perturbation model.
-A rough divide between the two leading theories of optimal control and robust control is as follows.
+Classical theories of control differ within the generic control problem formulation 1.2 primarily in one significant aspect: the perturbation model.
+A rough divide between the two main theories – optimal control and robust control – is as follows.
 Optimal control postulates a probabilistic model for the perturbations, which are typically i.i.d.
 draws from a Gaussian distribution.
-However, other types of distributions can be considered as well.
+However, other types of distribution can also be considered.
 As such, the control problem can be re-phrased in a more precise stochastic mathematical program,
@@
 $s.t.$ $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g(\mathbf{x}_{t}).$
-In contrast, robust control theory plans for the worst case noise from a given set of constraints, denoted $\mathcal{K}$.
+In contrast, robust control theory plans for the worst-case noise from a given set of constraints, denoted $\mathcal{K}$.
 Thus, the robust control problem can be rephrased as the following mathematical program,
@@
 The two formulations have been debated for decades, with pros and cons for each side.
-Optimal control focuses on the average case and many times yields simpler formulations for optimization, resulting in more efficient methods.
+Optimal control focuses on the average case and often yields simpler formulations for optimization, resulting in more efficient methods.
 Robust control, on the other hand, allows for more difficult noise models and some amount of model misspecification.
 However, it is too pessimistic for many natural problems.
-Nonstochastic control theory takes the best of both worlds, simultaneously granting robustness to adversarial noise and allowing the use of optimistic behavior when appropriate, while providing computationally efficient methods.
+Nonstochastic control theory takes the best of both worlds, simultaneously granting robustness to adversarial noise and allowing the use of optimistic behavior when appropriate while providing computationally efficient methods.
 1.6 The Need for a New Theory
-Consider the problem of flying a drone from source to destination subject to unknown weather conditions.
+Consider the problem of flying a drone from the source to the destination subject to unknown weather conditions.
 The aerodynamics of flight can be modeled sufficiently well by time-varying linear dynamical systems, and existing techniques are perfectly capable of doing a great job for indoor flight.
-However, the wind conditions, rain and other uncertainties are a different story.
-Certainly the wind is not an i.i.d.
+However, the wind conditions, rain, and other uncertainties are a different story.
+Certainly, the wind is not an i.i.d.
 Gaussian random variable!
@@
 Rather, an optimal policy for the observed perturbations can be determined only in hindsight.
-This is a significant difference from optimal and robust control theory, where the mapping from states to actions can completely be predetermined based upon the noise model.
+This is a significant difference from optimal and robust control theory, where the mapping from states to actions can be completely predetermined based upon the noise model.
 Since optimality is defined only in hindsight, the goal is not predetermined.
-We thus shift to a different performance metric borrowed from the world of learning in repeated games – regret.
+We thus shift to a different performance metric borrowed from the world of learning in repeated games, namely regret.
 Specifically, we consider regret with respect to a reference class of policies.
 In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$.
-Then the controller observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ according to an adversarially chosen loss function.
+The controller then observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ according to an adversarially chosen loss function.
 For simplicity, we consider the case of full state observation where $\mathbf{y}_{t}=\mathbf{x}_{t}$.
@@
 Henceforth, if $T,\Pi$ and ${\mathcal{A}}$ are clear from the context, we drop these symbols when defining $\mathrm{{regret}}$.
-At this point the reader may wonder why we should compare to a reference class, as opposed to all possible policies. There are two main answers to this question:
+At this point, the reader may wonder why we should compare to a reference class, as opposed to all possible policies. There are two main answers to this question:
   1. 1.
-    First, the best policy in hindsight out of all possible policies may be very complicated to describe and reason about.
+    First, the best policy in hindsight out of all possible policies may be very complicated to describe and to reason about.
 This argument was made by the mathematician Tyrell Rockafellar, see the bibliographic section for more details.
   2. 2.
-    Secondly, it can be shown that it is impossible, in general, to obtain sublinear regret vs.
-the best policy in hindsight.
+    Secondly, it can be shown that it is impossible, in general, to obtain sublinear regret versus the best policy in hindsight.
 Rather, a different performance metric called competitive ratio can be analyzed, as we discuss later in the book.
@@
 As a representative example, consider the Gradient Perturbation Controller described in Algorithm 1.
-The algorithm maintains matrices that are used to create the control as a linear function of the past perturbations, as per line 3.
-These parameters change over time according to a gradient-based update rule, where the gradient is calculated as the derivative of a counterfactual cost function with respect to the parametrization, as per line 5.
+The algorithm maintains matrices that are used to create the control as a linear function of past perturbations, according to line 3.
+These parameters change over time according to a gradient-based update rule, where the gradient is calculated as the derivative of a counterfactual cost function with respect to the parametrization, according to line 5.
 1:  Input: $h$, $\eta$, initial parameters ${M}_{1:h}^{1}$, dynamics $f$.
@@
 The version given here assumes that the time-invariant dynamics $f$ is known.
-The GPC computes a control which is a linear function of past perturbations.
+The GPC computes a control that is a linear function of past perturbations.
 It modifies the linear parameters according to the loss function via an iterative gradient method.
 In Chapter 7, we prove that this algorithm attains sublinear regret bounds for a large class of time-varying and time-invariant dynamical systems.
-In later chapters we explore extensions to nonlinear dynamical systems and unknown systems, which lie at the heart of nonstochastic control theory.
+In later chapters, we explore extensions to nonlinear dynamical systems and unknown systems, which lie at the heart of nonstochastic control theory.
 1.7 Bibliographic Remarks
@@
 In terms of application, self-regulating feedback mechanisms that limit the flow of water, much like the modern float valve, were reportedly used in ancient Greece over two thousand years ago.
-The first mathematical analysis of such systems appeared in a paper [Max68] by James Clerk Maxwell on the topic of governors, shortly after Maxwell had published his treatise on the unification of electricity and magnetism.
+The first mathematical analysis of such systems appeared in a paper by James Clerk Maxwell on the topic of governors [Max68], shortly after Maxwell had published his treatise on the unification of electricity and magnetism.
 For an enlightening historical account, see [FCZI03].
-For a historical account on the development of the PID controller see [Ben93]. The theoretical and practical properties of this controller are spelled out in detail in [ÅH95].
+For a historical account on the development of the PID controller, see [Ben93]. The theoretical and practical properties of this controller are described in detail in [ÅH95].
 There are many excellent textbooks on control theory, which we point to in later chapters.
-The book of [Ber07] offers an extensive coverage of topics from the viewpoint of dynamic programming.
+The book of [Ber07] offers extensive coverage of topics from the point of view of dynamic programming.
 The texts [Ste94, ZDG96a], on the other hand, delve deep into results on linear control, the most well-understood branch of control theory.
-Rockafellar [Roc87] proposed the use of convex cost functions for modeling the cost in a linear dynamical system to incorporate constraints on the state and control, and suggested computational difficulties with this approach.
+Rockafellar [Roc87] proposed the use of convex cost functions to model the cost in a linear dynamical system to incorporate constraints on state and control and suggested computational difficulties with this approach.
 The text [SL91] gives a thorough introduction to adaptive control theory.
@@
 A flurry of novel methods and regret bounds for online nonstochastic control have appeared in recent years, many of which are surveyed in [HS21].
-Notable amongst them are logarithmic regret algorithms [AHS19, FS20], nonstochastic control for unknown systems [HKS20], black-box online nonstochastic control [CHK+18], and nonstochastic control with partial observability [SSH20].
-The theory builds heavily on recent developments in decision making and online learning, notably the framework of online convex optimization [Haz16], and especially its extensions to learning with memory [AHM13].
+Notable among them are logarithmic regret algorithms [AHS19, FS20, Sim20], nonstochastic control for unknown systems [HKS20], black-box online nonstochastic control [CH21], and nonstochastic control with partial observability [SSH20].
+The theory is heavily based on recent developments in decision making and online learning, notably the framework of online convex optimization [Haz16], and especially its extensions to learning with memory [AHM13].
 Chapter 2 Dynamical Systems
-In this chapter we return to the general problem of control from definition 1.2, i.e.
-choosing a sequence of controls to minimize a sequence of cost functions.
-Before moving to algorithms for control, it is important to reason about what kind of solutions are even possible.
-For this reason we consider several natural examples of dynamical systems, and reason about what constitutes desirable behavior.
-This will help us understand the kinds of observations and control objectives that are feasible and natural in real applications.
+In this chapter, we return to the general problem of control from Definition 1.2, that is, choosing a sequence of controls to minimize a sequence of cost functions.
+Before moving on to algorithms for control, it is important to reason about what kind of solution is even possible.
+For this reason, we consider several natural examples of dynamical systems, and reason about what constitutes desirable behavior.
+This will help us understand the kinds of observation models and control objectives that are feasible and natural in real applications.
 2.1 Examples of Dynamical Systems
 2.1.1 Medical Ventilator
-In the task of mechanical ventilation, the control algorithm dictates the amount of oxygen inflow $u_{t}$ (in $\mathrm{m}^{3}/s$) into a patient’s lungs.
+In the task of mechanical ventilation, the control algorithm dictates the amount of oxygen inflow $u_{t}$ (in $\mathrm{m}^{3}/s$) into the lungs of a patient.
 This inflow approximately determines the volume of the lungs $v_{t}$ via a simple linear relation,
@@
 Instead, the controller observes the air pressure $p_{t}$.
-We give a simplified physical model relating the pressure and the volume.
-This model, given below, can be derived by first principles from stress-strain relations, treating the trachea-connected left and right lungs as situated in a two-balloon experiment.
+We give a simplified physical model that relates the pressure and the volume.
+This model, given in the following, can be derived by first principles from stress-strain relations, treating the left and right lungs connected by the trachea as situated in a two-balloon experiment.
 $p_{t}=C_{0}+C_{1}v_{t}^{-1/3}+C_{2}v_{t}^{5/3}$
-The objective of the control algorithm is to track a physician-prescribed ideal pressure trajectory $p^{*}_{t}$ as closely as possible, say as measured by in squared loss $c_{t}(p_{t},u_{t})=|p_{t}-p^{*}_{t}|^{2}+|u_{t}|^{2}$, while limiting the maximum instantaneous airflow to minimize patient discomfort.
-See Figure 1.2 for a schematic representation of the airflow when a lung is connected to a mechanical ventilator.
+The objective of the control algorithm is to follow an ideal pressure trajectory $p^{*}_{t}$ prescribed by the physician as closely as possible, say, as measured by squared loss $c_{t}(p_{t},u_{t})=|p_{t}-p^{*}_{t}|^{2}+|u_{t}|^{2}$, while limiting the maximum instantaneous airflow to minimize patient discomfort.
+See Figure 1.2 for a schematic representation of airflow when a lung is connected to a mechanical ventilator.
 Figure 2.1: A schematic of the respiratory circuit from [SZG+21].
@@
 The double integrator is a common example of a linear dynamical system.
-In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics on a one dimensional trajectory (i.e.
+In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics on a one-dimensional trajectory (i.e.
 the real number line).
@@
 Optionally, the state is also affected by external perturbations $\mathbf{w}_{t}$ outside the influence of the controller.
-The goal here is to ensure that the point object arrives and then comes to rest at some coordinate $x^{*}\in\mathbb{R}$ as fast as possible, while also not expending too much energy.
+The goal here is to ensure that the point object arrives and then comes to rest at some coordinate $x^{*}\in\mathbb{R}$ as quickly as possible, while also not expending too much energy.
 The cost function is naturally expressed as
@@
 Being vertically aligned, the system is subject to gravity with gravitational acceleration $g$.
-The controller is in charge of applying a time-varying torque $u_{t}\in\mathbb{R}$ at each time step.
+The controller is responsible for applying a torque $u_{t}\in\mathbb{R}$ at each time step.
 Given the vertical angle $\theta_{t}$ of the rod, the physics of this system may be summarized as
@@
 where $\Delta$ is the time resolution of the controller.
-Note that this dynamical equation is non-linear in the state $\mathbf{x}_{t}=[\theta_{t},\dot{\theta_{t}}]^{\top}$ due to the presence of the $\sin\theta$ term in the evolution of $\dot{\theta_{t}}$.
-Figure 2.3: A pivoted massless rod connected to a point object with mass $m$. The point object is to subject to gravitational acceleration $g$. Source: Wikipedia
-The objective here is to get the rod to rest at a vertically upright position at an angle of $\pi$, using as little torque as possible throughout the process.
-This may be expressed via the cost function
+Note that this dynamical equation is non-linear in the state $\mathbf{x}_{t}=[\theta_{t},\dot{\theta_{t}}]^{\top}$ due to the presence of the term $\sin\theta$ in the evolution of $\dot{\theta_{t}}$.
+Figure 2.3: A pivoted massless rod connected to a point object with mass $m$.
+The point object is to subject to gravitational acceleration $g$.
+Illustration by Krishnavedala, licensed under CC-BY-SA-3.0.
+The objective here is to get the rod to rest in a vertical upright position at an angle of $\pi$, using as little torque as possible throughout the process.
+This may be expressed through the cost function
 $c(\mathbf{x}_{t},u_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}\pi\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
@@
 We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system, from [Boy10].
-The 2-dimensional control input $\mathbf{u}_{t}$ corresponds to the elevator angle and the thrust.
+The 2-dimensional control input $\mathbf{u}_{t}$ corresponds to the elevation angle and the thrust.
 The disturbances correspond to the velocity of the wind along the body axis and perpendicular to it.
-The dynamics equation below has matrices whose numerical values were computed using the dynamics equation for an actual aircraft.
+The following dynamics equation has matrices whose numerical values were computed using the dynamics equation for an actual aircraft.
 These numerical values can be determined by the mass, length, and other physical properties of the aircraft.
@@
 $+\begin{bmatrix}-0.003&0.039&0&-0.322\\ -0.065&-0.319&7.74&0\end{bmatrix}\mathbf{w}_{t}$
-Figure 2.4: Longitudinal control in an aircraft. Source: Wikipedia.
-The quantities of interest for assessing the performance of the controller are aircraft speed, climb rate, each of which is a linear transformation of the state, and the applied thrust.
+Figure 2.4: Longitudinal control in an aircraft. Graphic by User_A1, licensed under CC-BY-SA 4.0; it is an SVG derivative of a public domain work by G J Vigurs.
+The quantities of interest for assessing the performance of the controller are the aircraft speed, climb rate, each of which is a linear transformation of the state, and the applied thrust.
 The corresponding cost is
@@
 2.1.5 Epidemiological Models
-The COVID-19 pandemic marked an increased public interest in policy discussions on the effectiveness of certain interventions as a means of controlling the spread of infection.
-Epidemiological models forecast the spread of an infectious disease via structural models and differential equations.
-We present the discrete-time analogue of the most basic of such models – the SIR model.
-Let $S_{t}$, $I_{t}$, and $R_{t}$ denote the susceptible, infected, and recovered population sizes.
+The COVID-19 pandemic marked an increase in public interest in policy discussions on the effectiveness of certain interventions as a means of controlling infection spread.
+Epidemiological models predict the spread of an infectious disease via structural models and differential equations.
+We present a discrete-time analogue of the most basic of such models, the SIR model.
+Let $S_{t}$, $I_{t}$, and $R_{t}$ denote the population sizes of those susceptible, infected, and recovered, respectively.
 Let $u_{t}$ be the number of vaccines administered at time $t$.
@@
 $R_{t+1}$ $=R_{t}+\gamma I_{t}.$
-Here the coefficient $\beta$ measures how communicable the infection is, $\gamma$ measures how quickly an infected person recovers, and $\alpha$ represents the efficacy of the vaccine in preventing infections.
-Figure 2.5: Trajectory of a continuous-time SIR model. Source: Wikipedia.
+Here, the coefficient $\beta$ measures the level of communicability of the infection, $\gamma$ measures the rate at which an infected person recovers, and $\alpha$ represents the efficacy of the vaccine in preventing infections.
+Figure 2.5: Trajectory of a continuous-time SIR model.
 2.2 Solution Concepts for Dynamical Systems
 Before considering the control, or manipulation, of a dynamical system, we must first reason about what the goal should even be.
-In this section, we discuss various solution concepts for a dynamical system, including equilibria, stability, stabilizability, and controllability.
-Given a dynamical system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t})$, the deterministic-equivalent dynamical system, defined via the dynamics $\mathbf{x}_{t+1}=f^{\prime}(\mathbf{x}_{t},\mathbf{u}_{t})=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{0})$, models the evolution of the dynamical system in the absence of perturbations.
-Further, the autonomous dynamical system $\mathbf{x}_{t+1}=f^{\prime\prime}(\mathbf{x}_{t})=f(\mathbf{x}_{t},\mathbf{0},\mathbf{0})$ describes state transitions of the dynamical system under the absence of both control inputs and perturbations.
+In this section, we discuss various solution concepts for a dynamical system, including equilibrium, stability, stabilizability, and controllability.
+Given a dynamical system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t})$, the equivalent deterministic dynamical system, defined through the dynamics $\mathbf{x}_{t+1}=f^{\prime}(\mathbf{x}_{t},\mathbf{u}_{t})=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{0})$, models the evolution of the dynamical system in the absence of perturbations.
+In addition, the autonomous dynamical system $\mathbf{x}_{t+1}=f^{\prime\prime}(\mathbf{x}_{t})=f(\mathbf{x}_{t},\mathbf{0},\mathbf{0})$ describes the state transitions of the dynamical system in the absence of both control inputs and perturbations.
 2.2.1 Equilibrium of Dynamical Systems
-An example of an equilibrium is a ball standing at rest at the bottom of a pit.
+An example of an equilibrium is a ball sitting at rest at the bottom of a pit.
 If the ball were to be slightly moved, it would return to the bottom of the pit.
 Thus, this equilibrium is stable and attractive.
-If the ball lies at rest on the top of a mountain, pushing the ball in any direction would cause it to fall. This is an equilibrium which is unstable. A formal definition is as below.
+If the ball is at rest on the top of a mountain, pushing the ball in any direction would cause it to fall. This is an equilibrium that is unstable. A formal definition is as follows.
 Definition 2.1.
@@
 Clearly $\mathbf{0}$ is an equilibrium point as ${A}\mathbf{0}=\mathbf{0}$ for all ${A}$.
-If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vectors converges to the null vector.
-Thus $\mathbf{0}$ is a globally attractive, stable equilibrium.
-In full generality, determining if a system has an equilibrium is an intractable problem, as we show in this chapter.
+If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vector converges to the null vector.
+Thus, $\mathbf{0}$ is a globally attractive and stable equilibrium.
+In general, determining whether a system has an equilibrium is an intractable problem, as we show in this chapter.
 2.2.2 Stabilizability and Controllability
@@
 Definition 2.3.
-A Dynamical System is controllable if and only if, for every initial and target state, there exists a sequence of controls that drives the deterministic-equivalent system to the target state.
-Note that controllability clearly implies stabilizability as the zero state is one possible target state.
-In the next section, we discuss the computational difficulties that arise in determining if a system is stabilizable and/or controllable.
+A dynamical system is controllable if and only if, for every initial and target state, there exists a sequence of controls that drives the equivalent deterministic system to the target state.
+Note that controllability clearly implies stabilizability, as the zero state is one possible target state.
+In the next section, we discuss computational difficulties that arise in determining if a system is stabilizable and/or controllable.
 2.2.3 Quantitative Definitions of Stabilizability
@@
 In classical control theory this is called BIBO-stability, which stands for Bounded-Input-Bounded-Output.
-This definition can apply to any dynamical system and is an important concept in non-stochastic control.
+This definition can apply to any dynamical system and is an important concept in nonstochastic control.
 Definition 2.4.
@@
 For formal properties of this attractor set, see [Str14, Tuc99].
-This demonstrates that even studying a solution concept by itself is a formidable task.
+This shows that even studying a solution concept itself is a formidable task.
 For a formal proof of attractiveness in the continuous case, see Section 9.2 of [Str14] or the original paper of Lorenz [Lor63].
-Figure 2.6: Julia set. Souce: Wikipedia.
-Figure 2.7: Lorenz attractor. Source: Wikipedia.
+Figure 2.6: Julia set.
+Figure 2.7: Lorenz attractor. Visualization by D Schwen, licensed under CC-BY-2.5.
 2.3.1 Computational Hardness of Stabilizability
-Consider the following problem: given a dynamical system, can we determine if it stabilizable, i.e.
-for every possible starting state does there exist a choice of control inputs that drives the system to the null state?
+Consider the following problem: given a dynamical system, can we determine if it stabilizable, i.e., for every possible starting state, does there exist a choice of control inputs that drives the system to the null state?
 We now show that this decision problem is computationally hard.
@@
 The Lorenz equations were introduced to model the phenomenon of atmospheric convection [Lor63].
-For detailed information about the mathematical properties of the Lorenz attractor, see [Str14, Tuc99].
+For detailed information on the mathematical properties of the Lorenz attractor, see [Str14, Tuc99].
 For a rigorous study of Julia sets and the Mandelbrot set, see [BY08].
 Nonlinear dynamical systems admit a variety of stability criteria; we chose to define a few that are useful for further developments in the text.
-A notable omission from our discussion is that of local and global stability; a celebrated result [Lya92] proves that the existence of certain potential functions is sufficient to guarantee global stability, thereby obviating the need for a trajectory-based analysis.
-An interested reader might look to [Kha15] for an expanded discussion on how these stability criteria compare.
+A notable omission from our discussion is that of local and global stability; a celebrated result [Lya92] proves that the existence of certain potential functions is sufficient to guarantee global stability, thus obviating the need for a trajectory-based analysis.
+An interested reader might refer to [Kha15] for an expanded discussion on how these stability criteria compare.
 An excellent survey on the computational complexity of control is [BT00].
@@
 Let $\alpha,\beta,\gamma>0$.
-Does the SIR-model, as it is described in this chapter, define a stabilizable dynamical system (for an appropriate and fixed definition of the “zero state”)?
+Does the SIR-model, as is described in this chapter, define a stabilizable dynamical system (for an appropriate and fixed definition of the “zero state”)?
 What about controllability?
-We assume that only starting states with the property $S_{0}+I_{0}+R_{0}=1$ are allowed, and that $S_{t},I_{t},R_{t},u_{t}\geq 0$ has to hold at all times.
+We assume that only starting states with the property $S_{0}+I_{0}+R_{0}=1$ are allowed and that $S_{t},I_{t},R_{t},u_{t}\geq 0$ must hold at all times.
 2.4. Consider the double integrator system:
@@
 $A=\begin{bmatrix}1&\Delta\\ 0&1\end{bmatrix},\quad B=\begin{bmatrix}0\\ 1\end{bmatrix},\quad\mathbf{x}_{t}=\begin{bmatrix}x_{t}\\ v_{t}\end{bmatrix}\in\mathbb{R}^{2},\quad u_{t}\in\mathbb{R},\quad\Delta>0.$
-The goal is to ensure the system reaches the desired state $[x^{*},0]^{\top}$ as $t\to\infty$ from any initial condition.
+The goal is to ensure that the system reaches the desired state $[x^{*},0]^{\top}$ as $t\to\infty$ from any initial condition.
   1. 1.
@@
 […]
-    with $K=[k_{1}\;k_{2}]$, that places all eigenvalues of $A-BK$ strictly inside the unit circle.
-Find conditions on $\Delta,k_{1},k_{2}$ that ensure all eigenvalues of $A-BK$ are real and lie strictly between $-1$ and $1$, thus guaranteeing $\mathbf{x}_{t}\to[x^{*},0]^{\top}$.
+    with $K=[k_{1}\;k_{2}]$, which places all eigenvalues of $A-BK$ strictly within the unit circle.
+Find conditions on $\Delta,k_{1},k_{2}$ that ensure that all eigenvalues of $A-BK$ are real and lie strictly between $-1$ and $1$, thus guaranteeing $\mathbf{x}_{t}\to[x^{*},0]^{\top}$.
 Chapter 3 Markov Decision Processes
@@
 This chapter introduces a setting of equal generality as that of the control problem from the previous chapter: reinforcement learning in Markov Decision Processes (MDP).
-However, the representation of states and dynamics in a MDP is more explicit and verbose.
-In such case, we can design efficient methods aimed at optimality given the explicit state and transition representations.
-After introducing the reinforcement learning model and Markov Decision Processes, we discuss an important algorithmic technique that will be useful throughout this book: dynamic programming, and the accompanying characterization of optimal policies via the Bellman equation.
+However, the representation of states and dynamics in an MDP is more explicit and verbose.
+In such a case, we can design efficient methods aimed at optimality given the explicit state and transition representations.
+After introducing the reinforcement learning model and Markov Decision Processes, we discuss an important algorithmic technique that will be useful throughout this book: dynamic programming and the accompanying characterization of optimal policies via the Bellman equation.
 3.1 Reinforcement Learning
-Reinforcement learning is a sub-field of machine learning that formally models the setting of learning through interaction in a reactive environment.
+Reinforcement learning is a subfield of machine learning that formally models the setting of learning through interaction in a reactive environment.
 This is in contrast to learning from examples, as in supervised learning, or inference according to a probabilistic model, as in Bayesian modeling.
@@
 The environment looks at the agent’s state and hands out rewards based on a set of criteria hidden from the agent.
-Figure 3.1: Source: Wikipedia.
+Figure 3.1: The agent-environment interaction loop in reinforcement learning.
 Typically, the goal of reinforcement learning is for the agent to learn behavior that maximizes the total reward it receives from the environment.
-This methodology has led to some notable successes: machines have learned how to play Atari games, and how to beat the highest-rated human players at Go.
+This methodology has led to some notable successes: machines have learned how to play Atari games and how to beat the highest-rated human players at Go.
 A crucial assumption for the reinforcement learning model is as follows:
@@
 The reader can imagine situations where such an assumption is reasonable.
-For example, in the inverted pendulum setting described in section 2.1.3, the assumption clearly holds.
+For example, in the inverted pendulum setting described in Section 2.1.3, the assumption is clearly valid.
 It is not significant for the control problem how a certain state, describing the angle and the angular velocity, was reached, in order to optimally control the pendulum from this state onward.
@@
 For example, in the medical ventilation problem, it may very well be important how a certain state was reached, as future treatment may depend on the medical history of the patient.
-It can be argued that the Markovian assumption can always be made to work, by adding more features to the state.
-In the extreme, the state can contain the entire history, which makes the assumption true by default.
+It can be argued that the Markovian assumption can always be made to work by adding more features to the state.
+In extreme cases, the state can contain the entire history, which makes the assumption true by default.
 However, this makes the state representation inefficient.
@@
 Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel and Norvig [RN02].
-In this example, a robot starts in the bottom leftmost square, and its goal is to reach the $+1$ reward.
-The blocks $+1,-1$ represent terminal states, meaning that the game ends when the robot reaches them.
+In this example, a robot starts in the bottom left square, and its goal is to reach the $+1$ reward.
+The blocks $+1,-1$ represent terminal states, which means that the game ends when the robot reaches them.
 The block ‘X’ represents an obstacle through which the robot cannot travel.
@@
 In this example, the robot has controls that allow it to move left / right / forward / backward.
-The controls, however, are not reliable, and with $0.8$ probability the robot moves to wherever it intends.
-With the remaining $0.2$ probability the robot moves sideways, that is, with probability $0.1$ to each side.
+However, the controls are not reliable and, with $0.8$ probability, the robot moves wherever it intends to.
+With the remaining probability $0.2$, the robot moves sideways, that is, with probability $0.1$ to each side.
 Notice that the Markovian assumption clearly holds for this simple example.
@@
 A solution concept for a reinforcement learning problem is a mapping from state to action.
-This mapping is called a “policy”, and it depends on both the structure of the problem as well as on the rewards.
-For example, if the rewards at all other states are slightly negative, Bubbo wants to reach the terminal state with a persistent $+1$ reward, as given in the following figure.
+This mapping is called a “policy”, and depends both on the structure of the problem and on the rewards.
+For example, if the rewards in all other states are slightly negative, Bubbo wants to reach the terminal state with a persistent $+1$ reward, as given in the following figure.
 […]
@@
 The optimal policy changes: now the robot is more concerned with ending the game as quickly as possible rather than trying to get to the +1 reward.
-We proceed to describe a formal model for capturing this and any other reinforcement learning problem.
+We then describe a formal model to capture this and any other reinforcement learning problem.
 3.2 Markov Decision Processes
-In this chapter we describe the general and formal model for reinforcement learning: Markov decision processes.
+In this chapter, we describe the general and formal model for reinforcement learning: Markov decision processes.
 To ease the exposition, we start with a review of Markov Chains, a.k.a.
@@
 This is the first layer in abstraction before allowing for actions, as it only describes a process without exogenous inputs.
-It consists of a set of states and a transition matrix describing the probability of moving from one state to another.
+It consists of a set of states and a transition matrix that describes the probability of moving from one state to another.
 The probabilities of moving between states are the weights on the edges of the graph.
@@
     Aperiodicity. A Markov Chain is aperiodic if and only if $d(s_{i})=1$ for all states $s_{i}\in S$, where
-    $d(s_{i})=\gcd\left\{t\in\mathbb{N}_{+}\ \ |\ \ P_{ii}^{t}>0\right\},$
-    is the periodicity of state $s_{i}$, where $\gcd$ stands for “greatest common divisor”.
-The ergodic theorem asserts that every irreducible and aperiodic Markov chain has a unique stationary distribution.
+    $d(s_{i})=\gcd\left\{t\in\mathbb{N}_{+}\ \ |\ \ (P^{t})_{ii}>0\right\},$
+    is the periodicity of state $s_{i}$, where $\gcd$ stands for “greatest common divisor".
+The ergodic theorem states that every irreducible and aperiodic Markov chain has a unique stationary distribution.
 Its proof is a beautiful application of spectral graph theory.
@@
 A Markov chain is said to be reversible if there exists some distribution with respect to which the Markov chain is reversible.
-In exercise 3.6 we prove that a distribution that is reversible for a Markov chain is also stationary for this Markov chain.
+In exercise 3.6, we prove that a distribution that is reversible for a Markov chain is also stationary for this Markov chain.
 3.2.2 Markov Reward Processes
 Markov Reward Processes (MRPs) add the concept of reward to Markov Chains.
-At every state, the agent receives a reward, which we assume for simplicity to be in the range $[0,1]$.
-In terms of algorithmic consequences, no generality is lost by this assumption, and this is left as an exercise.
-The cumulative reward is a sum of discounted rewards over time, so that immediate rewards are worth more than those received in the future.
-This is consistent with the economics literature, where a fixed amount of money is worth more today than in the future, due to positive interest rates.
+In each state, the agent receives a reward, which, for simplicity, we assume to be in the range $[0,1]$.
+In terms of algorithmic consequences, no generality is lost due to this assumption, and this is left as an exercise.
+The cumulative reward is the sum of discounted rewards over time, so that the immediate rewards are worth more than those received in the future.
+This is consistent with the economics literature, where a fixed amount of money is worth more today than in the future, because of positive interest rates.
 Definition 3.3.
@@
 When the discount factor is strictly less than one, $G_{t}$ is well defined.
-In the case $\gamma=1$, an alternative definition of $G$ is presented below, assuming that a stationary distribution for the underlying Markov chain exists, denoted by $\boldsymbol{\pi}$.
+In the case $\gamma=1$, an alternative definition of $G$ is presented below, assuming that there is a stationary distribution for the underlying Markov chain, denoted by $\boldsymbol{\pi}$.
 $G:=\lim_{t\to\infty}\frac{1}{t}\sum_{i=1}^{t}R_{i}=\sum_{s\in S}\boldsymbol{\pi}(s)\hat{R}(s)$
@@
 $\mathbf{v}(s)=\mathbb{E}[G_{t}|S_{t}=s].$
-The value function tells how valuable a state is as measured by the sum of its future, discounted rewards.
+The value function tells how valuable a state is as measured by the sum of its future discounted rewards.
 One way to compute the value function is by using the Bellman equation for the Markov Reward Process, which is the following recursive equation:
@@
 $=\hat{R}(s)+\gamma\sum_{s^{\prime}}P_{s,s^{\prime}}\mathbf{v}(s^{\prime}).$
-In matrix form we can write the Bellman equation as $\mathbf{v}=\mathbf{\hat{R}}+\gamma P\mathbf{v}$, which simplifies to $\mathbf{v}=(I-\gamma P)^{-1}\mathbf{\hat{R}}$.
+In matrix form, we can write the Bellman equation as $\mathbf{v}=\mathbf{\hat{R}}+\gamma P\mathbf{v}$, which simplifies to $\mathbf{v}=(I-\gamma P)^{-1}\mathbf{\hat{R}}$.
 Given $P$ and $\mathbf{\hat{R}}$, this is a linear system of equations, and thus the value function can be solved using Gaussian elimination in cubic time in the number of states.
 3.2.3 Markov Decision Processes
-The final layer of abstraction is to allow actions, or controls in the language of control theory. This gives the full Markov Decision Process model as defined below.
+The final layer of abstraction is to allow actions or controls in the language of control theory. This gives the full Markov Decision Process model as defined below.
 Definition 3.5.
@@
     $\gamma\in[0,1)$ is the discount factor.
-The central solution concept of study in the theory of reinforcement learning is called a policy. A policy is a mapping from state to distribution over actions, i.e.
+The central solution concept of study in the theory of reinforcement learning is called a policy. A policy is a mapping from state to distribution over actions, that is,
 $\pi:S\to\Delta(A)\ ,\ \pi(a|s)=\Pr[A_{t}=a|S_{t}=s]$
-The goal of reinforcement learning is to determine a sequence of actions to maximize the long term reward.
+The goal of reinforcement learning is to determine a sequence of actions to maximize long-term reward.
 The Markovian structure of the problem suggests that this optimal value is realized by a stationary policy, i.e.
@@
 where $R_{t}$ is a random variable that denotes the reward at time $t$ that depends on the randomness in the policy as well as the transitions.
-Similar to the value function of a Markov Reward Process, we can define the value function for a Markov Decision Process as follows.
+Similarly to the value function of a Markov Reward Process, we can define the value function for a Markov Decision Process as follows.
 The value function for a state $s$ given a policy $\pi$ is defined to be the expected reward from this state onwards where the actions are chosen according to $\pi$ at each state, i.e.
@@
 $Q_{\pi}(s,a)=\mathop{\mbox{\bf E}}_{\pi,P}\left[\sum_{t=1}^{\infty}\gamma^{t-1}R_{t}|S_{1}=s,A_{1}=a\right].$
-A fundamental theorem of reinforcement learning is the following.
+A fundamental theorem of reinforcement learning is as follows.
 Theorem 3.6.
 There exists a deterministic optimal policy $\pi^{*}$, which maps every state to a single action.
-All optimal policies achieve the same optimal value $\mathbf{v}^{*}(s)$ as well as $Q^{*}(s,a)$, and this holds for every state and action.
+All optimal policies achieve the same optimal value $\mathbf{v}^{*}(s)$ as well as $Q^{*}(s,a)$, and this holds for all states and actions.
 The fact that there exists an optimal policy that is optimal regardless of the starting state in a MDP is a consequence of the Markovian nature of the transition dynamics.
@@
 The importance of this equation is paramount in algorithm design: we describe two efficient methods for reinforcement learning based on this equation later in this chapter.
-In later chapters we also make use of the Bellman equation to derive the stochastic-optimal LQR controller.
-The Bellman equation for any policy $\pi$ relates the value function for that policy to itself via one step of application for that policy. Mathematically, this amounts to the following:
+In later chapters, we also make use of the Bellman equation to derive the stochastic-optimal LQR controller.
+The Bellman equation for any policy $\pi$ relates the value function of that policy to itself through one step of application for that policy. Mathematically, this amounts to the following:
 $\mathbf{v}_{\pi}(s)=\sum_{a\in A}\pi(a|s)\left[R_{sa}+\gamma\sum_{s^{\prime}\in S}P_{ss^{\prime}}^{a}\mathbf{v}_{\pi}(s^{\prime})\right].$
@@
   1. 1.
-    $R^{\pi}$ is the induced expected reward vector when playing policy $\pi$, i.e. ${R^{\pi}}(s)=\mathop{\mbox{\bf E}}_{a\sim\pi(s)}[R_{sa}]$.
+    $R^{\pi}$ is the expected reward vector induced when playing the policy $\pi$, that is, ${R^{\pi}}(s)=\mathop{\mbox{\bf E}}_{a\sim\pi(s)}[R_{sa}]$.
   2. 2.
@@
     $P^{\pi}_{ss^{\prime}}=\sum_{a\in A}P^{a}_{ss^{\prime}}\cdot\pi(a|s).$
-The Bellman equation gains prominance when applied to the optimal policy of a Markov Decision Process.
-It is then called the Bellman optimality equation, and states that for every state and action, we have
+The Bellman equation gains prominence when applied to the optimal policy of a Markov Decision Process.
+Then it is called the Bellman optimality equation, and states that for every state and action, we have
 $\mathbf{v}^{*}(s)=\max_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}$
@@
 $\mathbf{v}^{*}(s)=\max_{a}\left\{Q^{*}(s,a)\right\},$
-where $Q^{*}$ is the $Q$ function of the optimal policy, i.e.
+where $Q^{*}$ is the $Q$ function of the optimal policy, i.e.,
 $Q^{*}(s,a)=R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\max_{a^{\prime}}\left\{Q^{*}(s^{\prime},a^{\prime})\right\}.$
-In subsequent chapters, we will also consider cases where the return of interest is either the reward accumulated in a finite number of time steps, i.e.
-$G=\sum_{t=1}^{T}R_{t}$ in the finite-time setting, or the infinite-time average reward case, i.e.
+In subsequent chapters, we will also consider cases where the return of interest is either the reward accumulated in a finite number of time steps, that is, $G=\sum_{t=1}^{T}R_{t}$ in the finite-time setting, or the infinite-time average reward case, i.e.
 $G=\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}R_{t}$.
@@
 The number of variables is $|S|$, and the number of constraints is $|S|\times|A|$, which can be solved in polynomial time in the number of states and actions.
-The feasibility of this linear program is not immediate, but will be established in the next section.
+The feasibility of this linear program is not immediate but will be established in the next section.
 3.4 The Value Iteration Algorithm
-In the previous section we established the Bellman equation and its consequence to a polynomial time solution for reinforcement learning in Markov Decision Processes.
-In this section we study a fundamental algorithm for solving Markov Decision Processes that is a more direct application of dynamic programming and the Bellman equation.
+In the previous section, we established the Bellman equation and its consequence to a polynomial-time solution for reinforcement learning in Markov Decision Processes.
+In this section, we study a fundamental algorithm for solving Markov Decision Processes that is a more direct application of dynamic programming and the Bellman equation.
 The value iteration algorithm starts from an arbitrary solution and improves it until a certain measure of optimality is reached.
@@
 Define the residual vector $\mathbf{r}_{t}(s)=\mathbf{v}^{*}(s)-\mathbf{v}_{t}(s)$.
-Then the optimal value vector $\mathbf{v}^{*}$ is unique and value iteration converges to the optimal value function and policy at a geometric rate of
+Then the optimal value vector $\mathbf{v}^{*}$ is unique and the value iteration converges to the optimal value function and policy at a geometric rate of
 $\|\mathbf{r}_{t+1}\|_{\infty}\leq\gamma^{t}\|\mathbf{r}_{1}\|_{\infty}.$
@@
 Or in vector form, $\mathbf{r}_{t+1}\leq\gamma P^{\mathbf{a}^{*}}\mathbf{r}_{t}$.
-Furthermore, $\mathbf{r}_{t}\geq 0$ is entry-wise non-negative for all $t$.
+Furthermore, $\mathbf{r}_{t}\geq 0$ is entry-wise nonnegative for all $t$.
 (See Exercise 3.6.) Therefore, it holds that
@@
 The classical text on reinforcement learning, authored by pioneers of the subject, is [SB18].
-In addition, this text extensively discusses the early ventures into the field from researchers working in artificial intelligence, psychology, and neuroscience.
-Another great reference is the lecture notes of [Sil15].
+In addition, this text extensively discusses early ventures into the field of researchers working in artificial intelligence, psychology, and neuroscience.
+Another great reference is the set of lecture notes [Sil15].
 See [AJKS19] for an excellent introduction to the theoretical aspects of reinforcement learning, such as sample complexity bounds and reductions to supervised learning.
-The reader may also look to [Ber19] for an extensive survey of computational approaches to learning in unknown Markov decision processes.
+The reader may also refer to [Ber19] for an extensive survey of computational approaches to learning in unknown Markov decision processes.
 The first author had the benefit of a discussion with [Aro20] on the proof of convergence of value iteration presented in this chapter.
@@
 It is also a good resource for an involved reader who might be interested in the completeness of the Bellman characterization and the existential concerns for the average-time setting, which we did not expand upon.
-Although tabular MDPs fail to compactly represent stochastic linear dynamical systems, function-approximation schemes [JYWJ20, DKL+21, JKA+17, KKL+20] successfully capture the stochastic LQR setting.
+Although tabular MDPs do not compactly represent stochastic linear dynamical systems, function-approximation schemes [JYWJ20, DKL+21, JKA+17, KKL+20] successfully capture the stochastic LQR setting.
 3.6 Exercises
-3.1. In this exercise we give a guided proof of the ergodic theorem. Consider a given irreducible and aperiodic Markov chain with transition matrix $P$.
+3.1. In this exercise, we give a guided proof of the ergodic theorem. Consider an irreducible and aperiodic Markov chain with a transition matrix $P$.
 A. Show that $P$ has a left eigenvalue equal to 1.
 B. Show that all eigenvalues of $P$ are at most one in magnitude.
-The irreducibility and aperiodicity of the Markov chain imply that the eigenvalue equal to one is unique.
-We are going to assume this fact, and denote by $\mathbf{v}_{1}$ the corresponding eigenvector.
+The irreducibility and aperiodicity of the Markov chain imply that the eigenvalue equal to one is unique. We are going to assume this fact and denote by $\mathbf{v}_{1}$ the corresponding eigenvector.
 C.
@@
 3.5.
-Prove that the successive iterates $\mathbf{v}_{t}$ produced by the Value Iteration algorithm are monotonically increasing, i.e.
-$\mathbf{v}_{t+1}\geq\mathbf{v}_{t}$, when initialized with $\mathbf{v}_{0}=\mathbf{0}$.
+Prove that the successive iterates $\mathbf{v}_{t}$ produced by the Value Iteration algorithm increase monotonically, that is, $\mathbf{v}_{t+1}\geq\mathbf{v}_{t}$, when initialized with $\mathbf{v}_{0}=\mathbf{0}$.
 Hence, conclude that for any $t$, $\mathbf{v}^{*}\geq\mathbf{v}_{t}$.
-3.6. Let $P$ be an aperiodic and irreducible Markov chain. Prove that $P$ is reversible if and only if for any sequence of states $s_{1},\dots s_{n}$, it holds that
+3.6. Let $P$ be an irreducible and aperiodic Markov chain. Prove that $P$ is reversible if and only if for any sequence of states $s_{1},\dots s_{n}$, it holds that
 $P_{s_{1}s_{2}}P_{s_{2}s_{3}}\dots P_{s_{n-1}s_{n}}P_{s_{n}s_{1}}=P_{s_{1}s_{n}}P_{s_{n}s_{n-1}}\dots P_{s_{2}s_{1}}.$
@@
 We have seen examples of LTI systems in Chapter 2, namely the double integrator and the aircraft dynamics.
-Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories. Source: Wikipedia.
+Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories.
+Created by XaosBits using the Mathematica packages CurvesGraphics.nb by Gianluca Gorni and DrawGraphics.nb by David and Alice Park, licensed under CC-BY-2.5.
 A more general class of linear dynamical systems is when the matrices ${A}_{t}$ and ${B}_{t}$ change over time.
@@
 Consider any dynamics that can be expressed by a smooth function $f$.
-Such a system can be approximated by a time-varying linear system to an arbitrary accuracy determined by the time step interval between samples.
+Such a system can be approximated by a time-varying linear system to an arbitrary accuracy determined by the time-step interval between samples.
 To see this, first consider the one dimensional first-order Taylor expansion of a given dynamics function $f$, which is twice differentiable, around an anchor point $(\bar{x}_{t},\bar{u}_{t})$:
@@
 $f(\mathbf{x},\mathbf{u})=f(\bar{\mathbf{x}},\bar{\mathbf{u}})+J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{x}-\bar{\mathbf{x}})+J_{\mathbf{u}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{u}-\bar{\mathbf{u}})+w_{t},$
-where $J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is the Jacobian of $f$ with respect to the variables $\bar{\mathbf{x}}$. Writing the matrices ${A}_{t},{B}_{t}$ for the Jacobians, we have
+where $J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is the Jacobian of $f$ with respect to variables $\bar{\mathbf{x}}$. Writing the matrices ${A}_{t},{B}_{t}$ for the Jacobians, we have
 $\mathbf{x}_{t+1}={A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
-This derivation explains why in practice, the class of time varying linear dynamical systems is very expressive, at least for smooth dynamical systems.
+This derivation explains why in practice the class of time-varying linear dynamical systems is very expressive, at least for smooth dynamical systems.
 This is particularly true with hidden states that can capture complicated dynamics while imposing enough structure for efficient algorithms.
-As we see next, stabilizability and controllability can be checked efficiently for linear systems, in contrast to general dynamics.
+As we shall see next, stabilizability and controllability can be checked efficiently for linear systems, in contrast to general dynamics.
 4.2 Stabilizability of Linear Systems
@@
 although much of our discussion can be generalized to LTV systems.
-Recall that a system is stabilizable if from every starting state, there is a sequence of controls that drives the system to zero.
+Recall that a system is stabilizable if, from every starting state, there is a sequence of controls that drives the system to zero.
 Determining whether a linear time-invariant system is stabilizable is determined by a particularly simple condition as follows.
-Henceforth we denote by $\rho({A})$ the spectral radius of a matrix ${A}$, that is,
+Henceforth, we denote by $\rho({A})$ the spectral radius of a matrix ${A}$, that is,
 $\rho({A})=\max\{|\lambda_{1}|,...,|\lambda_{d}|\},$
-the maximum modulus from all (possibly complex) eigenvalues. The stabilizability of a linear dynamical system can be characterized as follows.
+the maximum modulus of all (possibly complex) eigenvalues. The stabilizability of a linear dynamical system can be characterized as follows.
 Theorem 4.1.
@@
 $\rho({A}+{B}{K})<1.$
-We henceforth prove sufficiency of this condition, and leave necessity as exercise 4.6. Both directions rely on the characterization of the spectral norm as follows:
+We therefore prove the sufficiency of this condition and leave the necessity as exercise 4.6. Both directions rely on the characterization of the spectral norm as follows:
 Fact 4.2.
@@
 Theorem 4.1, sufficiency.
-Suppose there exists a ${K}$ such that $\rho({A}+{B}{K})<1$. Then choose all controls to be $\mathbf{u}_{t}={K}\mathbf{x}_{t}$. The evolution of the system without perturbation is given by
+Suppose that there exists a ${K}$ such that $\rho({A}+{B}{K})<1$. Then choose all controls to be $\mathbf{u}_{t}={K}\mathbf{x}_{t}$. The evolution of the system without perturbation is given by
 $\mathbf{x}_{t}=({A}+{B}{K})\mathbf{x}_{t-1}=...=({A}+{B}{K})^{t}\mathbf{x}_{0}.$
@@
 4.2.1 Efficient Verification of Stabilizability
-The spectral radius of a matrix can be computed to an arbitrary accuracy in polynomial time, and can be approximated by the power method.
-This implies that stability of a given system can be efficiently verified.
+The spectral radius of a matrix can be computed to an arbitrary accuracy in polynomial time and can be approximated by the power method.
+This implies that the stability of a given system can be efficiently verified.
 To efficiently check if a given system is stabilizable, we make note of the following result due to Lyapunov.
+We use the notation $A\succeq 0$ to denote that the matrix $A$ is positive semidefinite, meaning that it is symmetric and all its eigenvalues are nonnegative.
 Theorem 4.3.
@@
 To observe the latter, first note that there exists a positive constant $c<1$ such that $cP\succeq A^{\top}PA$, implying that $c^{t}P\succeq(A^{t})^{\top}PA^{t}\succeq(A^{t})^{\top}A^{t}$ via repeated applications of the previous inequality.
-Therefore, we have that $\lim_{t\to\infty}(A^{t})^{\top}A^{t}=0$, and consequently that $\lim_{t\to\infty}A^{t}=0$ since $\|A^{\top}\|\|A\|=\|A^{\top}A\|$.
+Therefore, we have $\lim_{t\to\infty}(A^{t})^{\top}A^{t}=0$, and consequently $\lim_{t\to\infty}A^{t}=0$ since $\|A^{\top}\|\|A\|=\|A^{\top}A\|$.
 For the other direction, given some matrix $A$ with $\rho(A)<1$, consider $P=\sum_{t=0}^{\infty}(A^{t})^{\top}A^{t}$.
 Since $P=I+A^{\top}PA$, $P$ satisfies the required conclusions.
-The fact that $P$ as defined exists, i.e.
-that it is finite, as long as $\rho(A)<1$ can be verified via the Jordan normal form of $A$, which in particular asserts that $\|A^{t}\|\leq O(t^{d_{x}}\rho(A)^{t})$.
+The fact that $P$ as defined exists, that is, that it is finite, as long as $\rho(A)<1$ can be verified via the Jordan normal form of $A$, which in particular asserts that $\|A^{t}\|\leq O(t^{d_{x}}\rho(A)^{t})$.
 ∎
-Using this observation, one can write a semi-definite mathematical program that checks if a given system is stabilizable, and if so produces a matrix $K$ that guarantees a spectral radius smaller than one.
+Using this observation, one can write a semidefinite mathematical program that checks if a given system is stabilizable, and if so produces a matrix $K$ that guarantees a spectral radius smaller than one.
 Theorem 4.4.
@@
 Then, ${M}\succ{0}$ if and only if ${A}\succ{0}$ and ${C}-{B}^{\top}{A}^{-1}{B}\succ{0}$.
-Note that since the spectral radius purely involves the eigenvalues of a matrix, it is invariant under transposition, and it holds that $\rho((A+BK)^{\top})=\rho(A+BL)$.
-Now using this fact and Theorem 4.3, we know that stabilizability of a linear system is equivalent to the existence of a matrices ${K}$ and ${P}$ such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)P(A+BK)^{\top}}$.
+Note that since the spectral radius purely involves the eigenvalues of a matrix, it is invariant under transposition, and we have $\rho((A+BK)^{\top})=\rho(A+BK)$.
+Now using this fact and Theorem 4.3, we know that the stabilizability of a linear system is equivalent to the existence of matrices ${K}$ and ${P}$ such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)P(A+BK)^{\top}}$.
 The above lemma allows us to reformulate this cubic constraint to a quadratic constraint.
@@
 $\begin{bmatrix}{P}&{((A+BK)P)^{\top}}\\ {(A+BK)P}&{P}\end{bmatrix}\succ{0},\qquad{P}\succeq{I}.$
-Lastly, we reparametrize this constraint as a linear semi-definite constraint as demonstrated below, by choosing ${L}={KP}$.
+Lastly, we reparameterize this constraint as a linear semidefinite constraint as demonstrated below, by choosing ${L}={KP}$.
 $\begin{bmatrix}{P}&{(AP+BL)^{\top}}\\ {AP+BL}&{P}\end{bmatrix}\succ{0},\qquad{P}\succeq{I}.$
 Notice that this is indeed a reformulation of the mathematical feasibility program, and not a relaxation, since the reparamertization ${K}={LP^{-1}}$ is invertible. ∎
-4.3 Controllability of LDS
-Recall that a system is controllable if and only if in the absence of noise and for every target state, there is a sequence of controls that drives the system state to that target.
+4.3 Controllability of Linear Dynamical Systems
+Recall that a system is controllable if and only if in the absence of noise and for every target state there is a sequence of controls that drives the system state to that target.
 Linear dynamical systems exhibit a particularly elegant characterization of controllability.
-First, we need to define an object of central importance in linear dynamical system, the Kalman matrix.
+First, we need to define an object of central importance for linear dynamical systems, the Kalman matrix.
 Definition 4.6.
@@
 ${K}_{r}({A},{B})=[{B},{AB},{A^{2}B},\ldots,{A^{r-1}B}]\in{\mathbb{R}}^{d_{x}\times rd_{u}}.$
-The controllability of a linear dynamical system can now be characterized using this matrix. Recall the rank of a matrix is the maximal number of linearly independent columns (or rows).
+The controllability of a linear dynamical system can now be characterized using this matrix. Recall that the rank of a matrix is the maximal number of linearly independent columns (or rows).
 Theorem 4.7.
@@
 Proof.
-Consider the evolution of the noiseless linear dynamical system starting from initial state $\mathbf{x}_{0}$. The state at time $T$ is given by the expression
+Consider the evolution of the noiseless linear dynamical system starting from the initial state $\mathbf{x}_{0}$. The state at time $T$ is given by the expression
 $\mathbf{x}_{T}$ $={A}\mathbf{x}_{T-1}+{B}\mathbf{u}_{T-1}$
@@
 $\mathop{\mbox{\rm span}}[{AB,\ldots,A^{t+1}B}]\subseteq\mathop{\mbox{\rm span}}[{AB,\ldots,A^{t}B}].$
-Adding ${B}$ to both matrices doesn’t affect the subset relationship.
+Adding ${B}$ to both matrices does not affect the subset relationship.
 $\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t+1}B}]\subseteq\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t}B}]$
@@
 Since the rank cannot decrease through the addition of columns, we have $r_{t+2}=r_{t+1}$.
-Since the rank can grow at most $d_{x}$ times, the equation $r_{d_{x}}=r_{T}$ holds for all $T>d_{x}$ if and only if the system is controllable, and this concludes the theorem.
+Since the rank can grow at most $d_{x}$ times, equation $r_{d_{x}}=r_{T}$ holds for all $T>d_{x}$ if and only if the system is controllable, and this concludes the theorem.
 ∎
@@
 In the rest of this text our goal is to give efficient algorithms with finite time performance bounds for online control.
-These finite time bounds naturally depend on how difficult it is for a given system to stabilize.
+These finite-time bounds naturally depend on how difficult it is for a given system to stabilize.
 Therefore, we require a quantitative definition for stabilizability, one that holds for both time-invariant and time-varying linear dynamical systems.
-4.4.1 Stabilizability of Time-Invariant LDS
+4.4.1 Stabilizability of Time-Invariant Systems
 Consider a linear time-invariant system given by the transformations $({A,B})$.
-We have already remarked that this system is called stabilizable, according to the classical, noise-free definition, if and only if there exists $K$ such that the condition $\rho({A+BK})<1$ holds.
-A stronger quantitative property can be obtained as follows, and it is in fact implied from BIBO stabilizability as defined in Definition 2.4. The proof of this fact is left as exercise 4.6.
+We have already remarked that this system is called stabilizable, according to the classical definition, if and only if there exists $K$ such that $\rho({A+BK})<1$ holds.
+A stronger quantitative property can be obtained as follows, and in fact it is implied by BIBO stabilizability as defined in Definition 2.4. The proof of this fact is left as exercise 4.6.
 Lemma 4.8.
 Consider a linear controller ${K}$ for the LTI system $({A,B})$.
-If there exists $\kappa>0$ and $\delta\in(0,1]$ such that the matrix ${A+BK}$ admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
+If there exist $\kappa>0$ and $\delta\in(0,1]$ such that the matrix ${A+BK}$ admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
 $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t},\qquad\frac{\kappa}{\delta}\leq\gamma$
@@
 $\leq\kappa\sum_{i=0}^{\infty}(1-\delta)^{i}\leq\frac{\kappa}{\delta}\leq\gamma.$
-A linear time-invariant system is thus $\gamma$-stabilizable if we can find a matrix $K$ with the properties above.
-4.4.2 Stabilizability of time-variant linear systems
+A linear time-invariant system is thus $\gamma$-stabilizable if we can find a matrix $K$ with the above properties.
+4.4.2 Stabilizability of Time-Varying Systems
 The $\gamma$-stabilizability condition, as opposed to classical stabilizability, is particularly important to generalize to time-varying systems, since the spectral radius of the product is not bounded by the product of spectral radii, see exercise 4.6 at the end of this chapter.
-We now define a sufficient condition for $\gamma$ stabilizability. Consider a time-varying linear dynamical system given by the transformations $\{{A}_{t},{B}_{t}\}$.
+We now define a sufficient condition for $\gamma$-stabilizability. Consider a time-varying linear dynamical system given by the transformations $\{{A}_{t},{B}_{t}\}$.
 Lemma 4.9.
@@
 $x_{t+1}=x_{t}+\varepsilon u_{t}.$
-Clearly, such a system is nominally controllable, as indeed is any one-dimensional system with non-zero $B$.
-However, to reach a state $x=1$ from an initial $x_{0}=0$ in a constant number of time steps requires a large control input $u_{t}$ of size at least $1/\varepsilon$.
+Clearly, such a system is nominally controllable, as is indeed any one-dimensional system with nonzero $B$.
+However, reaching a state $x=1$ from an initial $x_{0}=0$ in a constant number of time steps requires a large control input $u_{t}$ of magnitude at least $1/\varepsilon$.
 Thus, such a system may be said to be barely controllable.
-This difficulty of controllability is captured in the definition below, which requires not only that $K(A,B)$ be full row-rank, but also that such a statement is robust to small changes in $A,B$.
+This difficulty of controllability is captured in the definition below, which requires not only that $K(A,B)$ is a full row-rank, but also that such a statement is robust to small changes in $A,B$.
 Definition 4.10.
 A time-invariant linear dynamical system $(A,B)$ is $(k,\kappa)$-strongly controllable if and only if
-$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\text{ and }\|({K}_{k}(A,B))^{-1}\|\leq\kappa.$
-In the definition above, the spectral norm of the inverse of $K(A,B)$ is bounded by $\kappa$.
-For example, in the simple one-dimensional system we presented above, the matrix $K(A,B)$ is $\varepsilon$, and thus $\kappa=\frac{1}{\varepsilon}$, which indeed captures the complexity of controlling the system.
+$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\text{ and }\|({K}_{k}(A,B){K}_{k}(A,B)^{\top})^{-1}\|\leq\kappa.$
+In the above definition, in addition to the full row-rank constraint, the spectral norm of the inverse of the Gram matrix of $K(A,B)$ is bounded by $\kappa$.
+For example, in the simple one-dimensional system we presented above, with $k=1$, the matrix $K(A,B)$ is $\varepsilon$, and thus $\kappa=\frac{1}{\varepsilon^{2}}$, which indeed captures the complexity of controlling the system.
 4.5 Bibliographic Remarks
 The idea of reducing a nonlinear dynamical system to a time varying linear dynamical system had been independently discovered in disparate fields.
-In the context of control, using the linearization to characterize local stability properties of non-linear systems was pioneered by Lyapunov [Lya92], and is called the Lyapunov direct method, see e.g.
+In the context of control, using the linearization to characterize local stability properties of non-linear systems was pioneered by Lyapunov [Lya92], and is called the Lyapunov direct method; see, e.g.
 [SL91].
@@
 [Ste94, Hes18]).
-The efficient verification of stabilizability via a semi-definite program and the recovery of a stable controller form the solution of the latter appears in [Ahm16].
+The efficient verification of stabilizability via a semidefinite program and the recovery of a stable controller form the solution of the latter appears in [Ahm16].
 On the other hand, it is known that even imposition of norm constraints on a plausibly stable controller makes this problem computationally hard [BT97, BT00, BT99].
-Somewhat distinct, but related, problems of stabilizability with even a 1-dimensional linear constraint are open to the date.
+Somewhat distinct but related problems of stabilizability with even a 1-dimensional linear constraint are open to the date.
 The notion of $\gamma$-stabilizability has a precursor in the notion of strong stability introduced in [CHK+18].
-An extension of this strong stability notion to time-varying linear dynamics was established in [CKM19].
+An extension of this strong stability concept to time-varying linear dynamics was established in [CKM19].
 The definition presented in this chapter has weaker requirements than those in the listed articles, allowing us to work with and guarantee regret later with respect to a larger class of comparator policies.
@@
 Prove that the existence of a matrix $K$ such that $\rho(A+BK)<1$ is a necessary condition for a linear dynamical system to be stabilizable.
-Hint: this exercise requires the optimal solution to LQR from the next chapter.
+Hint: this exercise requires the optimal solution to LQR in the next chapter.
 4.2.
@@
 constitutes a reversible Markov chain if and only if $A$ is symmetric, i.e. $A=A^{\top}$.
-Hint: One path to this exercise is to explicitly calculate the stationary distribution of a LDS and then to verify reversibility as stated in its definition.
+Hint: One way to do this exercise is to explicitly calculate the stationary distribution of an LDS and then to verify reversibility as stated in its definition.
 Exercise 3.6 provides an alternative without the need to derive the stationary distribution.
@@
 This exercise explores the consequences of $\gamma$-BIBO stability (Definition 2.4) for linear systems.
-While we have already seen sufficient conditions to ensure BIBO stability, we ask for the necessity of such conditions.
-A. For a time-invariant linear system $(A,B)$ equipped with a linear policy $K$, prove that BIBO stability indeed implies that $\rho(A+BK)<1$. In fact, comment on how large $\rho(A+BK)$ can be.
+Although we have already seen sufficient conditions to ensure BIBO stability, we ask for the necessity of such conditions.
+A. For a time-invariant linear system $(A,B)$ equipped with a linear policy $K$, prove that BIBO stability indeed implies $\rho(A+BK)<1$. In fact, comment on how large $\rho(A+BK)$ can be.
 B.
-For a time-invariant BIBO-stable linear system, prove that there exists $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$, $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}$.
+For a time-invariant BIBO-stable linear system, prove that there exist $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$, $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}$.
 In particular, this proves that the conditions listed in Lemma 4.8 are necessary, in addition to being sufficient.
-C. Construct a time-varying linear system that is BIBO stable, yet does not have the exponential decay properties required in Lemma 4.9.
-4.6. Prove that the bound of one on the operator norm of the matrix $A$ is sufficient but not necessary condition for stabilizability.
+C. Construct a time-varying linear system that is BIBO stable but does not have the exponential decay properties required in Lemma 4.9.
+4.6. Prove that the bound of one on the operator norm of the matrix $A$ is a sufficient but not necessary condition for stabilizability.
 Chapter 5 Optimal Control of Linear Dynamical Systems
@@
 $\min_{\mathbf{u}(\mathbf{x})}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R}\mathbf{u}_{t}\right].$
-To remind the reader, the notation we use is as follows:
+To remind the reader, the notation we use is as follows.
   1. 1.
@@
     $Q\in{\mathbb{R}}^{d_{x}\times d_{x}}$, ${R}\in{\mathbb{R}}^{d_{u}\times d_{u}}$ are positive definite matrices. Thus, the cost functions are convex.
-This framework is very useful since many physical problems, such as the centrifugal governor and the ventilator control problem, can be approximated by a linear dynamical system.
-In addition, the quadratic convex cost functions are natural in many physical applications.
+This framework is very useful, since many physical problems, such as the centrifugal governor and the ventilator control problem, can be approximated by a linear dynamical system.
+In addition, quadratic convex cost functions are natural in many physical applications.
 The LQR is a special case of a Markov decision process which we have described in Chapter 3, since the transitions are stochastic.
-However, we cannot naive apply value iteration, policy iteration, or linear programming since there are infinitely many states and actions.
+However, we cannot naively apply value iteration, policy iteration, or linear programming since there are infinitely many states and actions.
 We proceed to describe a particularly elegant solution to the LQR problem that makes use of its special structure.
 5.2 Optimal Solution of the LQR
-We know from Chapter 3 that the LQR is a special case of a finite-time Markov Decision Process, and hence admits an optimal policy.
+We know from Chapter 3 that the LQR is a special case of a finite-time Markov decision process and hence admits an optimal policy.
 However, the optimal policy is in general a mapping from states to actions, both of which are infinite Euclidean spaces in our case.
-Nevertheless, it can be shown that the optimal policy for an LQR is a linear policy, i.e.
-a linear mapping from ${\mathbb{R}}^{d_{x}}$ to ${\mathbb{R}}^{d_{u}}$!
+However, it can be shown that the optimal policy for an LQR is a linear policy, that is, a linear mapping from ${\mathbb{R}}^{d_{x}}$ to ${\mathbb{R}}^{d_{u}}$!
 This astonishing fact, together with a remarkable property of the value function, is formally given in the following theorem.
-We state the theorem for a finite number of iterations, and for a linear time-invariant system.
+We state the theorem for a finite number of iterations and for a linear time-invariant system.
 Generalizing this to a time-varying linear system is left as exercise 5.6.
@@
 This implies that for any matrix $M$ which is independent of $\mathbf{w}$, we have $\mathop{\mbox{\bf E}}[\mathbf{w}^{\top}M\mathbf{w}]=\mathop{\mbox{\bf E}}[M\mathbf{w}\mathbf{w}^{\top}]=\sigma^{2}\cdot M\bullet I=\sigma^{2}{\bf Tr}(M)$.
-Since $R$ and $S_{t}$ are positive semi-definite, the expression above is convex as a function of $\mathbf{u}$. At the global minimum the gradient vanishes, and hence
+Since $R$ and $S_{t}$ are positive semidefinite, the above expression is convex as a function of $\mathbf{u}$. At the global minimum the gradient vanishes, and hence
 $\nabla_{\mathbf{u}}\Big{[}{\mathbf{x}^{\intercal}Q\mathbf{x}+\mathbf{u}^{\intercal}R\mathbf{u}+(A\mathbf{x}+B\mathbf{u})^{\intercal}S_{t}(A\mathbf{x}+B\mathbf{u})+\sigma^{2}{\bf Tr}(S_{t})}]\Big{]}=0.$
@@
 Although infinite horizon problems can never appear in the real world, in practice the horizon length may be unknown.
-Although there are techniques to iteratively update the horizon length with geometric increasing guesstimates, these are slow and impractical.
+Although there are techniques to iteratively update the horizon length with geometrically increasing guesstimates, these are slow and impractical.
 An alternative is to solve the LQR for the infinite horizon case, which results in a very elegant solution.
@@
 zero mean and have bounded variance.
-We henceforth assume that the perturbations are all zero for simplicity, adding zero-mean bounded variance perturbations is left as exercise 5.6.
+We henceforth assume that the perturbations are all zero for simplicity, and adding zero-mean bounded-variance perturbations is left as exercise 5.6.
 Without perturbations, the Bellman optimality equation becomes
@@
 $\mathbf{x}^{\top}S\mathbf{x}=\min_{\mathbf{u}}\left\{{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}+(A\mathbf{x}+B\mathbf{u})^{\top}S(A\mathbf{x}+B\mathbf{u})}\right\}.$
-Similar to the finite time case, we have the following result.
+Similarly to the finite-time case, we have the following result.
 Lemma 5.3.
@@
 Proof.
-At optimality, the gradient with respect to $\mathbf{u}$ of the equation above is zero. This implies that
+At optimality, the gradient with respect to $\mathbf{u}$ of the above equation is zero. This implies that
 ${2R\mathbf{u}+2B^{\top}SB\mathbf{u}+2B^{\top}SA\mathbf{x}=0},$
@@
 ${K=-(R+B^{\top}SB)^{-1}(B^{\top}SA)}.$
-Plugging back into the Bellman equation, we obtain
+Plugging this back into the Bellman equation, we obtain
 ${\mathbf{x}^{\top}S\mathbf{x}=\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{x}^{\top}K^{\top}RK\mathbf{x}+\mathbf{x}^{\top}(A+BK)^{\top}S(A+BK)\mathbf{x}}.$
@@
 5.3.1 Solving the DARE
-In the 1-D case, although the right side of the equation is a rational function, we can multiply both sides by ${R+B^{\top}SB}$ to form a quadratic equation in ${S}$.
-Solving the DARE therefore corresponds to finding a non-negative root of a quadratic polynomial in the one-dimensional case.
-The reason we only consider the non-negative root is since the value function is non-negative by our choice of quadratic convex costs.
+In the one-dimensional case, although the right side of the equation is a rational function, we can multiply both sides by ${R+B^{\top}SB}$ to form a quadratic equation in ${S}$.
+Solving the DARE therefore corresponds to finding a nonnegative root of a quadratic polynomial in the one-dimensional case.
+The reason we only consider the nonnegative root is that the value function is nonnegative by our choice of quadratic convex costs.
 In higher dimension, the correct generalization is somewhat more difficult to see, but using Lemma 4.5, the equation can be reformulated as the solution to the following SDP.
@@
 any possible perturbation sequence from a certain family.
-Formally, we aim to solve:
+Formally, our aim is to solve the following:
 $\begin{array}[]{c}\min_{\mathbf{u}_{1:T}(\mathbf{x})}\max_{\mathbf{w}_{1:T}}\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}{Q}\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}{R}\mathbf{u}_{t}\\ \text{ s.t.
@@
 Notice that this problem does not fall into the category of a Markov Decision Process, since $\mathbf{w}_{t}$’s are no longer stochastic.
-Therefore the typical Bellman optimality criterion does not apply.
+Therefore, the typical Bellman optimality criterion does not apply.
 Instead, the appropriate techniques to address this problem come from the theory of dynamical games.
 The above is, in fact, a two-player zero sum game.
-The first player tries to minimize the costs and the other attempts to maximize the costs, both players subject to the dynamics evolution constraint.
+The first player tries to minimize the costs, and the other attempts to maximize the costs, both players subject to the dynamics evolution constraint.
 Using techniques from dynamic games, the following can be shown. References are provided in the bibliographic section.
@@
 $\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}.$
-The proof is somewhat more involved than the LQR derivation, and is omitted from this text.
+The proof is somewhat more involved than the LQR derivation and is omitted from this text.
 5.5 Bibliographic Remarks
@@
 While seemingly a natural corollary of equivalent results for MDPs, the existence of value functions for linear control is made subtle by the continuous nature of the state space and stabilizability, or the lack thereof.
-A discussion on sufficient and necessary conditions to existence of solutions to Riccati equations can be found in [Ber07, HSK96].
-The development of $H_{\infty}$ control was driven by an observation made in the seminal paper of Doyle [Doy78] that, unlike in the fully observed case, the solution to the optimal control problem for partially observed linear systems (which we study in later chapters) is not naturally robust to miss-specification of system parameters.
-This suggests that robustness specifications constitute yet an another axis of performance for control algorithms, and may at times be at odds with the objective of optimal control.
+A discussion of sufficient and necessary conditions for the existence of solutions to Riccati equations can be found in [Ber07, HSK96].
+The development of $H_{\infty}$ control was driven by an observation made in the seminal paper of Doyle [Doy78] that, unlike in the fully observed case, the solution to the optimal control problem for partially observed linear systems (which we study in later chapters) is not naturally robust to misspecification of system parameters.
+This suggests that robustness specifications constitute yet another axis of performance for control algorithms and may at times be at odds with the objective of optimal control.
 For in-depth treatment of $H_{\infty}$ control, see [ZDG96b].
-The text [BB08] presents a succinct derivation of robust controllers entirely in the state space, as the solution to a dynamic game between the controller and an adversary, see also [Tu99] for a recent exposition.
+The text [BB08] presents a succinct derivation of robust controllers entirely in the state space, as the solution to a dynamic game between the controller and an adversary; see also [Tu18] for a recent exposition.
 Although developed in the linear setting, adaptations of the LQR controller are also widely used to control non-linear systems [Li05, LK13].
@@
 5.1. Prove Theorem 5.2 for a time-varying linear dynamical system.
-5.2. Derive the solution to the infinite horizon LQR with zero mean Gaussian noise.
-5.3. Extend the proof of Theorem 5.2 to positive semi-definite $Q,R$, rather than positive definite matrices.
+5.2. Derive the solution to the infinite-horizon LQR with zero mean Gaussian noise.
+5.3. Extend the proof of Theorem 5.2 to positive semidefinite $Q,R$, rather than positive definite matrices.
 5.4.
-Prove Theorem 5.4 as a dynamic two-player zero-sum game, where a unique saddle point solution exists.
+Prove Theorem 5.4 as a dynamic two-player zero-sum game, where a unique saddle-point solution exists.
 Hint: you can formulate this problem as a min-max optimization problem within a finite horizon.
-5.5. Prove that the function $f(\mathbf{x})=\mathbf{x}^{\top}M\mathbf{x}$ is convex if and only if $M\succeq 0$ is positive semi-definite.
+5.5. Prove that the function $f(\mathbf{x})=\mathbf{x}^{\top}M\mathbf{x}$ is convex if and only if $M\succeq 0$ is positive semidefinite.
 Part II Basics of Online Control
 Chapter 6 Regret in Control
-In the first part of this book and chapter 5 we studied linear dynamical systems with quadratic costs, namely the LQR problem.
+In the first part of this book and in chapter 5 we studied linear dynamical systems with quadratic costs, namely the LQR problem.
 We saw that in this case the optimal control takes a very special form: it is a linear function of the state.
 The Markovian property of the Markov Decision Process framework for reinforcement learning suggests that the optimal policy is always a function of the current state.
-It is characterized by the Bellman optimality equation, and can be computed in complexity proportional to the number of states and actions using, for example, the value iteration algorithm which we analyzed in a previous chapter.
+It is characterized by the Bellman optimality equation, and can be computed in complexity proportional to the number of states and actions using, for example, the value iteration algorithm, which we analyzed in a previous chapter.
 However, this changes when we deviate from the MDP framework, which may be warranted for a variety of reasons:
@@
 Rather, the learner may be required to use other signals available to her when choosing an action that is most beneficial in the long run.
-To this we add a significant consideration: computation of optimal state-based policies may be computationally intractable.
-This motivates the current chapter which considers regret analysis as a computational tractable alternative to optimal or robust control.
-We start by considering regret in the decision making framework of online convex optimization.
-We then consider several classes of control policies, as well as the relationships between them, in the context of regret analysis.
+To this we add a significant consideration: the computation of optimal state-based policies may be computationally intractable.
+This motivates the current chapter, which considers regret analysis as a computationally tractable alternative to optimal or robust control.
+We start by considering regret in the decision-making framework of online convex optimization.
+We then consider several classes of control policies and the relationships between them in the context of regret analysis.
 6.1 Online Convex Optimization
-In this section we give a brief introduction to the theory of Online Convex Optimization. The reader is referred to the bibliographic section for a comprehensive survey.
-In the decision making framework of Online Convex Optimization, a player iteratively chooses a point in a convex set $\mathcal{K}\subseteq{\mathbb{R}}^{n}$, and suffers a loss according to an adversarially chosen loss function $f_{t}:\mathcal{K}\mapsto{\mathbb{R}}$.
+In this section, we give a brief introduction to the theory of Online Convex Optimization. The reader is referred to the bibliographic section for a comprehensive survey.
+In the decision-making framework of Online Convex Optimization, a player iteratively chooses a point in a convex set $\mathcal{K}\subseteq{\mathbb{R}}^{n}$, and suffers a loss according to an adversarially chosen loss function $f_{t}:\mathcal{K}\mapsto{\mathbb{R}}$.
 Let ${\mathcal{A}}$ be an algorithm for OCO, which maps a certain game history to a decision in the decision set:
@@
 Intuitively, an algorithm performs well if its regret is sublinear as a function of $T$ (i.e.
-$\mathrm{{regret}}_{T}({\mathcal{A}})=o(T)$), since this implies that on average, the algorithm performs as well as the best fixed strategy in hindsight.
-We proceed to describe a basic algorithm for online convex optimization.
+$\mathrm{{regret}}_{T}({\mathcal{A}})=o(T)$), since this implies that, on average, the algorithm performs as well as the best fixed strategy in hindsight.
+We then describe a basic algorithm for online convex optimization.
 6.1.1 Online Gradient Descent
-Perhaps the simplest algorithm that applies to the most general setting of online convex optimization is Online Gradient Descent (OGD). Pseudo-code is given in Algorithm 3.
+Perhaps the simplest algorithm that applies to the most general setting of online convex optimization is Online Gradient Descent (OGD). The pseudocode is given in Algorithm 3.
 Algorithm 3 Online Gradient Descent
@@
 In each iteration, the algorithm takes a step from the previous point in the direction of the gradient of the previous cost.
-This step may result in a point outside of the underlying convex set.
+This step may result in a point outside the underlying convex set.
 In such cases, the algorithm projects the point back to the convex set, i.e.
 finds its closest point in the convex set.
-Despite the fact that the next cost function may be completely different than the costs observed thus far, the regret attained by the algorithm is sublinear.
+Despite the fact that the next cost function may be completely different from the costs observed thus far, the regret attained by the algorithm is sublinear.
 This is formalized in the following theorem.
@@
 The Online Gradient Descent algorithm is straightforward to implement, and updates take linear time given the gradient.
-However, the projection step may in general take significantly longer via convex optimization.
+However, in general, the projection step may take significantly longer via convex optimization.
 6.2 Regret for Control
 The most important consideration to explore beyond classical models of control, such as optimal and robust control, is computational.
-Even in linear control with convex loss functions, the optimal policy can be complex to describe, see exercises section.
-A prominent approach to deal with computational difficulty is to consider a less stringent criteria for optimality.
+Even in linear control with convex loss functions, the optimal policy can be complex to describe, see the exercises section.
+A prominent approach to deal with computational difficulty is to consider a less stringent criterion for optimality.
 Instead of the optimal policy, we can consider all policies in a certain policy class, and try to compete with the optimal policy within the class in terms of regret.
@@
 The significance of this definition is that if a simple policy class approximates another more complicated class, then it suffices to consider only the first.
-The usefulness of this definition stems from its implication for regret minimization, as per definition 1.3, as follows.
+The usefulness of this definition comes from its implications for regret minimization, according to definition 1.3, as follows.
 Lemma 6.3.
@@
 6.3 Expressivity of Control Policy Classes
-In this section we formally relate the power of different complexity classes for linear time-invariant linear dynamical systems.
-The definitions and results can be extended to time-varying linear dynamical systems, and recall from Chapter 4 that time varying linear dynamical systems were been proposed by Lyapunov as a general methodology to study nonlinear dynamics.
+In this section, we formally relate the power of different complexity classes for linear time-invariant linear dynamical systems.
+The definitions and results can be extended to time-varying linear dynamical systems, and recall from Chapter 4 that time-varying linear dynamical systems were proposed by Lyapunov as a general methodology to study nonlinear dynamics.
 For the rest of this section we assume full observation of the state and let
@@
 Schematically, the relationships we describe below are given in Figure 6.1 for fully observed systems.
-In later chapters of this text we study policies for partially observed systems.
+In later chapters of this text, we study policies for partially observed systems.
 Figure 6.1: Schematic relationship between linear Disturbance-Action Control policies and linear state-action policies of increasing diameter for linear time invariant systems.
@@
 those that choose a control signal as a (fixed) linear function of the state.
-The most important subcase of linear policies are those that are $\gamma$-stabilizing, meaning that the signal they create does not cause an unbounded state.
+The most important subcases of linear policies are those that are $\gamma$-stabilizing, meaning that the signal they create does not cause an unbounded state.
 This was formalized in Definition 2.4.
@@
 Definition 6.4 (Linear Policies).
-A linear controller $K\in{\mathbb{R}}^{d_{x}\times d_{u}}$ is a linear operator mapping state to control.
-We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers that are $\gamma$-stabilizing according to definition 2.4 and are bounded in Fronbenius norm by $\|K\|\leq\gamma$.
+A linear controller $K\in{\mathbb{R}}^{d_{x}\times d_{u}}$ is a linear operator that maps the state to the control.
+We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers that are $\gamma$-stabilizing according to definition 2.4 and are bounded in Frobenius norm by $\|K\|\leq\gamma$.
 6.3.2 Disturbance Action Controllers
-We next define the most important set of policies in online nonstochastic control, known as the disturbance action controllers, or DAC.
-The main advantages of DAC as opposed to linear controllers, are that the cost of control is convex in their parametrization, and thus allows for provably efficient optimization.
-Moreover, this set of policies is stronger than linear policies in the sense that it $\varepsilon$-approximates them, as we show below.
+We next define the most important set of policies in online nonstochastic control, known as disturbance action controllers, or DAC.
+The main advantages of DAC, as opposed to linear controllers, are that the cost of control is convex in their parameterization and thus allows for provably efficient optimization.
+Moreover, this set of policies is stronger than linear policies in the sense that it $\varepsilon$-approximates them, as we show in the following.
 Definition 6.5 (Disturbance Action Controller).
@@
 Without loss of generality, we take $M_{0}=0$, as we can add a fictitious unit coordinate to the perturbations $\{\mathbf{w}_{t}\}$, and modify the system accordingly.
-This will allow a constant shift in the control.
+This will allow for a constant shift in the control.
 Denote by $\Pi^{D}_{h,\gamma}$ the set of all disturbance action policies with $K_{1:T}$ being $\gamma$-stabilizing linear controllers for the system and
 $\sum_{i=1}^{h}\|M_{i}\|\leq\gamma.$
-The class of DAC always produces a stabilizing control signal whenever $K_{1:T}$ are sequentially stabilizing for the entire system $\{A_{t},B_{t}\}$.
+The class of DACs always produces a stabilizing control signal whenever $K_{1:T}$ are sequentially stabilizing for the entire system $\{A_{t},B_{t}\}$.
 This important fact can be seen in the following lemma.
@@
 for $\|M\|$ being the spectral norm of $M$.
-Next, since $K_{t}$’s are stabilizing, we can bound the historically distant terms as below.
-For simplicity assume $K_{t}=K$ are all equal.
+Next, since $K_{t}$’s are stabilizing, we can bound the historically distant terms as follows.
+For simplicity, assume that $K_{t}=K$ are all equal.
 Observe that the state at time $t$, following any DAC policy $\pi$, is given by
@@
 We can thus view the states $x_{t}$ as generated by a $\gamma$-stabilizing policy over noises that are of magnitude $\gamma$, and hence bounded by $\gamma^{2}$. ∎
-The above lemma is already a good indication about DAC: they generate controls that at least don’t cause a blow up. But how expressive are these policies in terms of reaching the optimal solution?
+The above lemma is already a good indication about DAC: they generate controls that at least do not cause a blow-up. But how expressive are these policies in terms of reaching the optimal solution?
 One of the main appeals of the DAC class is the following approximation lemma which shows that for time-invariant systems DAC are in fact a stronger class than linear controllers.
-Notice that the $K_{t}$ terms for DAC are only assumed to be stabilizing, and not necessarily optimal for the actual sequence.
+Notice that the $K_{t}$ terms for the DAC are only assumed to be stabilizing and not necessarily optimal for the actual sequence.
 Lemma 6.7.
@@
 $\|\mathbf{u}_{t}^{\pi_{1}}(A+BK_{0},B)-\mathbf{u}_{t}^{\pi_{2}}(A+BK_{0},B)\|\leq\varepsilon$
-when executed on some dynamical system $(A+BK_{0},B)$, then when the corresponding policies are executed on the dynamical system $(A,B)$ the iterates guarantee
-$\max\{\|\mathbf{x}_{t}^{\pi_{1}}(A,B)-\mathbf{x}_{t}^{\pi_{2}}(A,B)\|,\|\mathbf{u}_{t}^{\pi_{1}}(A,B)-\mathbf{u}_{t}^{\pi_{2}}(A,B)\|\}\leq(1+\|K_{0}\|)\varepsilon.$
+when executed on some dynamical system $(A+BK_{0},B)$, then when the corresponding policies $\pi^{\prime}_{1}$ and $\pi^{\prime}_{2}$, that is, $\pi_{1}$ and $\pi_{2}$ with $K_{0}$ superimposed, are executed on the dynamical system $(A,B)$ the iterates guarantee
+[…]
+$\|\mathbf{u}_{t}^{\pi^{\prime}_{1}}(A,B)-\mathbf{u}_{t}^{\pi^{\prime}_{2}}(A,B)\|\}\leq(1+\|K_{0}\|)\varepsilon.$
 With the aid of the above lemma, it will suffice to guarantee that the state-action sequence produced by the linear policy $K$ and an appropriately chosen DAC with the stabilizing part set to $\mathbf{0}$ on an intrinsically $\gamma$-stable system $(A,B)$ are similar enough.
@@
 $=\sum_{i=0}^{t}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}$
-$=\sum_{i=0}^{h}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}+Z_{t}$
+Now, we define $Z_{t}=\sum_{i=h+1}^{t}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}$, and continue as shown.
+$\mathbf{u}_{t+1}$ $=\sum_{i=0}^{h}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}+Z_{t}$
 $=\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}+Z_{t}$
 $=\tilde{\mathbf{u}}_{t+1}+Z_{t}$
-where $Z_{t}=\sum_{i=h+1}^{t}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}$.
-The magnitude of this residual term can be bounded, due to the existence of $\frac{\kappa}{\delta}\leq\gamma^{\prime}$ from Exercise 4.6, as
+The magnitude of this residual term $Z_{t}$ can be bounded, due to the existence of $\frac{\kappa}{\delta}\leq\gamma^{\prime}$ from Exercise 4.6, as
 $|Z_{t}|$ $\leq\left\|\sum_{i=h+1}^{t}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}\right\|$
-$\leq\sum_{i=h+1}^{t}(1-\delta)^{i+1}\left\|K\mathbf{w}_{t-i}\right\|$
+$\leq\sum_{i=h+1}^{t}\kappa(1-\delta)^{i+1}\|K\|\|\mathbf{w}_{t-i}\|$
 […]
-$=\kappa\frac{1}{\delta}e^{-\delta h}\leq\varepsilon.$
+$=\kappa^{2}\frac{1}{\delta}e^{-\delta h}\leq\varepsilon.$
 We thus conclude that $\|\mathbf{u}_{t+1}-\tilde{\mathbf{u}}_{t+1}\|\leq{\varepsilon}$.
-To make a similar observation on the corresponding state sequences, first note that
+Similarly, for the corresponding state sequences, first note that
 $\mathbf{x}_{t+1}=\sum_{i=0}^{t}A^{i}\left(\mathbf{w}_{t-i}+B\mathbf{u}_{t-i}\right).$
@@
 The reader is referred to the comprehensive texts of [CBL06, Haz16] for a detailed treatment.
-Linear state-feedback policies are a mainstay of control theory both in applications, due to their simplicity, and in theory, owing to their optimality in deterministic and stochastic settings as demonstrated in [Kal60].
-The methods for computing such policies are often based in dynamic programming.
-Often motivated by the need to impose constraints beyond optimality such as robustness, various other policy parametrizations were introduced.
-An early example is the Youla parametrization, sometimes called Q-parametrization [YJB76].
-Later on, the introduction of system level synthesis [DMW+17, ADLM19], in a similar vein, proved useful for distributed control, especially so given its emphasis on state-space representations.
-For a comparison of these and additional ones such as input-output parameterizations [FZPK19], see [ZFP+20].
-The Disturbance Action Control parametrization, introduced in [ABH+19], is similar to system level synthesis.
-Like SLS, it is a convex parametrization that is amenable to the tools of convex optimization.
+Of particular
+Linear state-feedback policies are a mainstay of control theory in both applications, due to their simplicity, and in theory, due to their optimality in deterministic and stochastic settings, as demonstrated in [Kal60].
+The methods for computing such policies are often based on dynamic programming.
+Often motivated by the need to impose constraints beyond optimality, such as robustness, various other policy parameterizations were introduced.
+An early example is the Youla parameterization, sometimes called Q-parameterization [YJB76].
+Later, the introduction of system-level synthesis (SLS) [DMW+17, ADLM19], in a similar vein, proved useful for distributed control, especially given its emphasis on state-space representations.
+For a comparison of these and additional ones such as input-output parameterizations [FZPK19], see [ZFP+20, ZZLL23].
+The Disturbance Action Control parameterization, introduced in [ABH+19], is similar to system-level synthesis.
+Like SLS, it is a convex parameterization that is amenable to the tools of convex optimization.
 However, unlike SLS, it is a more compact representation that has logarithmically many parameters in its natural representation in terms of the desired approximation guarantee against linear policies.
-This makes it better suited to machine learning.
-The study of regret in the context of control has roots going back to the work of [AYS11, AYBK14].
-These works, as well as many that followed [CHK+18, AHS19], study regret in a stochastic setting, where the perturbations are stochastic.
+This makes it better suited for machine learning.
+The study of regret in the context of control has roots in the work of [AYS11, AYBK14].
+These works, as well as many that followed [DMM+18, MTR19, CHK+18, AHS19, SF20], study regret in a stochastic setting, where the perturbations are stochastic.
 In such settings, optimal control is well defined.
 The full power of regret analysis for control, involving both adversarial perturbations, adversarial loss functions, and computational considerations, was initiated in [ABH+19].
+The notion of regret as applied to policy classes was implicit in the work of [EDKM04].
+The regret metric when applied to policies is different from the counterfactual notion of policy regret defined by [ADT12].
 6.5 Exercises
-6.1. In this exercise we prove that the optimal solution to a linear control problem can be complex to describe.
+6.1.
+Let us say that the total number of time steps $T$ is known in advance.
+Consider an implementation of online gradient descent where the step size is chosen as $\eta_{t}=D/G\sqrt{T}$, and therefore is uniform over time.
+Prove that the resultant regret is bounded by $GD\sqrt{T}$, which improves upon Theorem 6.1 by a constant factor.
+In fact, for ease of exposition, we will use this uniform step size scheme going forward.
+6.2. In this exercise, we prove that the optimal solution to a linear control problem can be complex to describe.
   1. 1.
-    Consider a linear dynamical system in one dimension, and loss functions of the form $f_{t}(\mathbf{x})=\max\{0,\mathbf{x}-c\}$, for $c_{t}\in{\mathbb{R}}$.
+    Consider a linear dynamical system in one dimension and loss functions of the form $f_{t}(\mathbf{x})=\max\{0,\mathbf{x}-c\}$, for $c_{t}\in{\mathbb{R}}$.
 Show that the optimal policy in hindsight can have as many domains as loss functions.
   2. 2.
-    Consider a similar setting in higher dimensions, and give a bound on the number of domains as a function of the number of loss functions and the dimension.
-6.2. Consider the setting of time varying linear dynamical systems, i.e.
+    Consider a similar setting in higher dimensions and give a bound on the number of domains as a function of the number of loss functions and the dimension.
+6.3. Consider the setting of time-varying linear dynamical systems, i.e.
 $\mathbf{x}_{t+1}=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
-Prove that any DAC according to definition 6.5 is stable, i.e. prove the analogue of Lemma 6.6 for time varying systems.
+Prove that any DAC according to definition 6.5 is stable, i.e., prove the analogue of Lemma 6.6 for time-varying systems.
 6.4. Prove Lemma 6.8.
 Chapter 7 Online Nonstochatic Control
-In this chapter, we explore recent advancements in machine learning that have significantly reshaped the control problem.
-These developments mark a major departure from classical control theory and, in many cases, have led to methods that not only compete with but often outperform traditional approaches.
+In this chapter, we explore recent advances in machine learning that have significantly reshaped the control problem.
+These developments mark a major departure from classical control theory and, in many cases, have led to methods that not only compete with, but often outperform traditional approaches.
 The key characteristics of these modern techniques are as follows:
@@
     Online Learning Over Precomputed Policies Unlike classical control, which relies on precomputing an optimal policy, online control adopts adaptive learning techniques to update policies dynamically.
-This aligns with the broader framework of adaptive control, where policies evolve in response to observed system behavior.
+This aligns with the broader framework of adaptive control, where policies evolve in response to observed behavior of the system.
   2. 2.
@@
     Improper Learning via Convex Relaxation A fundamental breakthrough in online nonstochastic control is the use of improper learning, where the algorithm competes against a given policy class while learning from a broader class of policies.
-By leveraging convex relaxation techniques, these methods bypass computational hardness barriers that traditionally limit optimal control, allowing for efficient algorithms with strong theoretical guarantees.
+Using convex relaxation techniques, these methods bypass computational hardness barriers that traditionally limit optimal control, and are accompanied by efficient algorithms with strong theoretical guarantees.
 These innovations redefine how we approach control problems, shifting the focus from static policy optimization to adaptive, regret-minimizing strategies that operate effectively in dynamic and adversarial environments.
 7.1 From Optimal and Robust to Online Control
-Classical control theory has been primarily shaped by two paradigms: optimal control and robust control.
+Classical control theory has been mainly shaped by two paradigms: optimal control and robust control.
 These approaches have been instrumental in various engineering applications, but they rely on specific assumptions about disturbances and system dynamics.
@@
 A key example is H-infinity control.
-While these frameworks have been successful, real-world control tasks often involve uncertainties that do not fit neatly into either a probabilistic model or a worst-case adversarial setting.
+Although these frameworks have been successful, real-world control tasks often involve uncertainties that do not fit neatly into either a probabilistic model or a worst-case adversarial setting.
 Consider the challenge of controlling a drone in an unpredictable environment:
@@
   • A worst-case robust approach may result in excessively conservative control, leading to inefficient and overly cautious maneuvers.
-These limitations motivate the need for a more adaptive approach—one that can learn and adjust dynamically based on observed disturbances rather than relying on rigid assumptions.
+These limitations motivate the need for a more adaptive approach, one that can learn and adjust dynamically based on observed disturbances rather than relying on rigid assumptions.
 This leads us to online nonstochastic control, a fundamentally different paradigm where controllers adapt in real-time without assuming a fixed model for disturbances.
@@
 Rather than relying on predefined stochastic or worst-case assumptions, online nonstochastic control introduces a new approach based on sequential decision-making and regret minimization.
-The core principles are:
+The core principles are the following.
   1. 1.
@@
 To build intuition, consider the problem of controlling an autonomous vehicle in an unknown city.
-If the vehicle had complete foresight of future traffic conditions, it could compute the optimal route in advance.
+If the vehicle had complete foresight of future traffic conditions, it could calculate the optimal route in advance.
 However, without such knowledge, the vehicle must continuously adjust its route based on real-time observations.
@@
 Unlike optimal control, which assumes disturbances follow a known probabilistic model, or robust control, which plans for the worst-case, online nonstochastic control operates in a more general and challenging setting.
-In this framework, both the disturbances and cost functions are chosen by an adversary, meaning they may follow no statistical pattern and could even be designed to hinder the controller’s decisions at every step.
-As a result, the goal is no longer to compute a fixed optimal control policy but rather to learn and adapt dynamically in response to an evolving sequence of perturbations and costs.
+In this framework, both disturbances and cost functions are chosen by an adversary, meaning that they may follow no statistically reliable pattern and could even be designed to hinder the controller’s decisions at every step.
+As a result, the goal is no longer to compute a fixed optimal control policy, but rather to learn and adapt dynamically in response to an evolving sequence of perturbations and costs.
 A fundamental difficulty in control is the computational tractability of finding the best policy.
 Classical control methods often rely on dynamic programming, which, in general, suffers from the curse of dimensionality.
-Even in simple settings, computing an optimal policy can be intractable due to the complexity of solving constrained optimization problems over trajectories.
+Even in simple settings, computing an optimal policy can be intractable due to the complexity of solving constrained optimization problems on trajectories.
 As pointed out by Rockafellar, finding an optimal control policy in arbitrary convex cost settings can be computationally prohibitive, particularly when state and control constraints are involved.
@@
 Instead of optimizing a predefined objective function, the controller seeks to minimize its total loss relative to the best policy that could have been chosen in hindsight.
-This approach bypasses the need to solve an intractable dynamic optimization problem upfront while still achieving performance guarantees over time.
+This approach bypasses the need to solve an intractable dynamic optimization problem in advance while still achieving performance guarantees over time.
 The fundamental challenge in online nonstochastic control can be summarized as follows:
@@
 where $\Pi$ is a benchmark class of policies, and $(x_{t}^{\pi},u_{t}^{\pi})$ represent the state and control sequence obtained under the best policy $\pi$ in hindsight.
-Henceforth, if $T,\Pi$ and ${\mathcal{A}}$ are clear from the context, we just refer to $\mathrm{{regret}}$ without quantifiers. The goal of an online control algorithm is to achieve sublinear regret,
+Henceforth, if $T,\Pi$ and ${\mathcal{A}}$ are clear from the context, we refer to $\mathrm{{regret}}$ without quantifiers. The goal of an online control algorithm is to achieve sublinear regret,
 $\mathrm{{regret}}_{T}=o(T),$ (7.1)
 which ensures that, as $T\to\infty$, the controller performs competitively with the best fixed policy.
-Improper Learning and Convex Relaxation
+7.2.1 Improper Learning and Convex Relaxation
 A key technique that enables online nonstochastic control algorithms to remain computationally efficient is improper learning through convex relaxation.
@@
 This approach is crucial because enforcing exact constraints on the policy space can lead to computational hardness results, as previously discussed.
-By relaxing the policy representation, for example using convex approximations or surrogate loss functions, controllers can sidestep these intractabilities while still competing effectively with the best policy in hindsight.
+By relaxing the policy representation, for example, using convex approximations or surrogate loss functions, controllers can sidestep these intractabilities while still competing effectively with the best policy in hindsight.
 In the next section, we introduce algorithmic solutions that utilize this principle to achieve low-regret online control under adversarial conditions.
@@
 After motivating the framework, defining regret, and exploring the relationships between different policy classes, we are ready to prove the main new algorithm in non-stochastic control.
-The algorithm is described in figure (4), for time varying linear dynamical systems.
+The algorithm is described in figure (4), for time-varying linear dynamical systems.
 It is a specialization of the template introduced in Algorithm 1 from Section 1.6.2.
-The main idea is to learn the parametrization of a DAC policy using known online convex optimization algorithms, namely online gradient descent.
-We henceforth prove that this parametrization is convex, allowing us to prove the main regret bound.
+The main idea is to learn the parameterization of a DAC policy using known online convex optimization algorithms, namely online gradient descent.
+We henceforth prove that this parameterization is convex, allowing us to prove the main regret bound.
 The notation we use to describe the algorithm is as follows.
-We denote by $\mathbf{x}_{t}({M}_{1:h})$ the hypothetical state arrived at by playing a DAC policy ${M}_{1:h}$ from the beginning of time.
+We denote by $\mathbf{x}_{t}({M}_{1:h})$ the hypothetical state reached by playing a DAC policy ${M}_{1:h}$ from the beginning of time.
 Note that this may not be equal to the true state $\mathbf{x}_{t}$ of the dynamical system at time $t$.
@@
 the best Disturbance Action Control policy belonging to our reference class in hindsight.
-As per the relationship between DAC and other policies, this implies vanishing regret vs.
-linear controllers as well.
+According to the relationship between DAC and other policies, this implies vanishing regret versus linear controllers as well.
 The formal statement is given as follows.
@@
   • b
-    The matrices $\{{A}_{t},{B}_{t}\}$ are $\gamma$-stabilizable, and have bounded $\ell_{2}$ norm.
+    The matrices $\{{A}_{t},{B}_{t}\}$ are $\gamma$-stabilizable and have a bounded $\ell_{2}$ norm.
 Then the GPC algorithm (4) ensures that
@@
 Proof.
-As per Theorem 6.1, the online gradient descent algorithm for convex loss functions ensures that the regret is bounded by $2GD\sqrt{T}$.
+According to Theorem 6.1, the online gradient descent algorithm for convex loss functions ensures that the regret is bounded by $2GD\sqrt{T}$.
 However, to directly apply the OGD algorithm to our setting, we need to ensure that the following conditions hold true:
   • The loss function should be a convex function in the variables ${M}_{1:h}$.
-  • We need to ensure that optimizing the loss function $\ell_{t}$ as a function of the parameters $M_{1:h}$ is similar to optimizing the cost $c_{t}$ as a function of the state and control.
-These facts are formalize in the following two lemmas, that we state here and prove at the end of this section.
+  • We must ensure that optimizing the loss function $\ell_{t}$ as a function of parameters $M_{1:h}$ is similar to optimizing the cost $c_{t}$ as a function of the state and control.
+These facts are formalize in the following two lemmas, which we state here and prove at the end of this section.
 Lemma 7.3.
@@
 7.3.1 Convexity of the Losses
-To conclude the analysis, We first show that the loss functions are convex w.r.t.
-to the variables ${M}_{1:h}$.
+To conclude the analysis, we first show that the loss functions are convex with respect to the variables ${M}_{1:h}$.
 This follows since the states and the controls are linear transformations of the variables.
@@
 $\ell_{t}({M}_{1:h})=c_{t}(\mathbf{x}_{t},\mathbf{u}_{t}).$
-Since the cost $c_{t}$ is a convex function w.r.t. its arguments, we simply need to show that $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{1:h}$. The state is given by
+Since the cost $c_{t}$ is a convex function with respect to its arguments, we simply need to show that $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{1:h}$. The state is given by
 $\mathbf{x}_{t+1}={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t}$ $={A}\mathbf{x}_{t}+{B}\left(\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i}\right)+\mathbf{w}_{t}$
@@
 The actual loss $c_{t}$ at time $t$ is not calculated on $\mathbf{x}_{t}(M_{1:h})$, but rather on the true state $\mathbf{x}_{t}$, which in turn depends on different parameters $M_{1:h}^{i}$ for various historical times $i<t$.
-Nevertheless, $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ is well approximated by $l_{t}(M^{t}_{1:h})$, as shown next.
+However, $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ is well approximated by $l_{t}(M^{t}_{1:h})$, as shown next.
 Proof of Lemma 7.4.
@@
 $\leq C\cdot\kappa\sum_{i=0}^{t}(1-\delta)^{i}\cdot i\frac{hD}{\sqrt{T}}$
-$=\frac{Ch\kappa D}{\sqrt{T}}\sum_{i=1}^{\infty}i(1-\delta)^{i-1}$ $\sum_{i=1}^{\infty}i(1-\delta)^{i}=\frac{1}{\delta^{2}}$
-$\leq\frac{C\gamma^{2}hD}{\sqrt{T}}.$
+$=\frac{Ch\kappa D}{\sqrt{T}}\sum_{i=1}^{\infty}i(1-\delta)^{i-1}$
+$\leq\frac{C\gamma^{2}hD}{\sqrt{T}},$
+where on the second last line we use $\sum_{i=1}^{\infty}i(1-\delta)^{i}=\frac{1}{\delta^{2}}$.
 The control at time $t$ depends only on the parameters at the current iteration and not on historical parameters, hence $\mathbf{u}_{t}=\mathbf{u}_{t}(M_{1:h}^{t})$.
@@
 7.4 Bibliographic Remarks
-The bulk of the material of this chapter is from [ABH+19].
+Online nonstochastic control is an extension of the framework of online convex optimization [Haz16], as it is applied to policies, and thus to loss functions with memory [AHM13].
+The bulk of the material in this chapter is from [ABH+19].
 The work of [ABH+19] defined the general online nonstochastic setting in which the disturbances and the cost functions are adversarially chosen, and where the cost functions are arbitrary convex functions.
-Follow up work by [AHS19] achieves logarithmic pseudo-regret for strongly convex, adversarially selected losses, and well-conditioned stochastic noise.
-Logarithmic regret algorithms were derived for online nonstochastic control in [Sim20, FS20].
+The follow-up work by [AHS19] achieves a logarithmic pseudo-regret for strongly convex, adversarially selected losses, and well-conditioned stochastic noise.
+Logarithmic regret algorithms for the full setting of online nonstochastic control with strongly convex costs were obtained in [Sim20, FS20, CCK20].
 Online nonstochastic control with known dynamics was further extended in several directions.
-Control with bandit feedback for the costs was considered in [GHH20].
+The task of control with bandit feedback for the costs was considered in [GHH20].
 More refined regret metrics applicable to changing dynamics were studied in [GHM20, MGSH21, ZWZ22, Sin22].
 More extensions suitable for changing environments were studied in [BW22].
+Online nonstochastic control of population dynamics was studied in [GHL+24].
 [LDL21] studies algorithms that balance regret minimization while adhering to constraints.
 An application of nonstochastic control in iterative planning in the presence of perturbations is described in [AHMS21].
-Deluca [GHM20], a control library, provides benchmarked implementations of many such algorithms.
-Online nonstochastic control is an extension of the framework of online convex optimization [Haz16] when applied to policies, and thus to loss functions with memory [AHM13].
-A similar regret notion was implicit in the work of [EDKM04].
-The regret metric when applied to policies is different from the counterfactual notion of policy regret as defined by [ADT12].
-An alternate performance metric for online control, which is also nonstochastic, is competitive ratio.
+Various open-source software packages, such as [GHM20], provide benchmarked implementations of many such algorithms.
+An alternate performance metric for online control, which is also nonstochastic, is the competitive ratio.
 This objective tracks the cost of the online controller and compares it to the best policy in hindsight.
-Control methods that minimize competitive ratio were initiated in [GW19].
+Control methods that minimize the competitive ratio were initiated in [GW19].
 This approach was extended in a series of papers [GLSW19, SLC+20, GH21].
-Thus far the success of this approach was limited to fully observed linear dynamical systems, and further to quadratic costs when achieving the optimal competitive ratio.
+Thus far, the success of this approach was limited to fully observed linear dynamical systems and further to quadratic costs when achieving the optimal competitive ratio.
 [GASH22] show that regret minimizers in the non-stochastic setting automatically achieve a near-optimal competitive ratio.
-A precursor to online nonstochastic control is a line of work, beginning with [AYBK14, CHK+18], focused on the control of stochastic systems under sequentially revealed adversarial costs functions.
+A precursor to online nonstochastic control is a line of work, beginning with [AYBK14], focused on the control of stochastic systems under sequentially revealed adversarial costs functions.
 See [CK20, CCK22] for further developments.
@@
 7.4. Consider control of a general, non-linear dynamical system. Give a version of the GPC algorithm 4 that can be applied to such a setting (without proving any performance guarantee).
-7.5. In this exercise we prove and extend Theorem 7.2 for linear time-varying systems that are stabilizable, rather than stable.
+7.5. In this exercise, we prove and extend Theorem 7.2 for linear time-varying systems that are stabilizable, rather than stable.
   1. 1.
@@
 In the previous chapters, we focused on control algorithms that assume knowledge of the system dynamics.
-However, in many real-world applications, system parameters are either unknown or change over time.
-This necessitates an approach where the controller must simultaneously learn the system model while making control decisions.
+However, in many real-world applications, system parameters are unknown or change over time.
+This necessitates an approach in which the controller must simultaneously learn the system model while making control decisions.
 This problem, known as system identification, is fundamental to adaptive control, reinforcement learning, and modern data-driven methods for control.
@@
   • Adversarial or arbitrary disturbances: The disturbances affecting the system may not follow a probabilistic model but can be chosen adversarially, leading to challenges in maintaining robustness.
-  • Balancing exploration and performance: The controller must explore sufficiently to estimate system parameters accurately while ensuring stable performance, a tradeoff that directly impacts regret minimization.
+  • Balancing exploration and performance: The controller must explore sufficiently to estimate system parameters accurately while ensuring stable performance, a trade-off that directly impacts regret minimization.
 8.2 A Nonstochastic Approach to System Identification
@@
   • Regret as a measure of learning efficiency: The effectiveness of an identification strategy is measured not in terms of statistical efficiency but in its ability to achieve low regret, ensuring that the system is identified quickly enough to enable near-optimal control.
-  • Computational tractability in identification: Classical system identification methods often involve solving nonconvex optimization problems, making them computationally demanding.
+  • Computational tractability in identification: Classical system identification methods often involve solving non-convex optimization problems, which makes them computationally demanding.
 Nonstochastic approaches frequently leverage convex relaxation techniques to sidestep intractability.
-The problem of online system identification is closely related to online convex optimization and sequential decision-making.
-Many regret-minimization techniques developed for online learning can be adapted to system identification settings.
-A key distinction, however, is that in control settings, the learned system model directly affects future decisions, creating a coupling between estimation and control that does not typically exist in standard online learning formulations.
+The problem of online system identification is closely related to online convex optimization and sequential decision making.
+Many regret minimization techniques developed for online learning can be adapted to system identification settings.
+However, a key distinction is that, in control settings, the learned system model directly affects future decisions, creating a coupling between estimation and control that does not typically exist in standard online learning formulations.
 This connection between system identification and online learning leads to several fundamental questions:
@@
 As a first step, we consider the question of whether the system matrices $A,B$ can be recovered from observations over the state, without regard to the actual cost of control.
-The method below for achieving this learns the system matrices by using a random control.
-Specifically, at each step the control $\mathbf{u}_{t}$ will be randomly selected from a Rademacher distribution, as per Algorithm 5.
+The method below to achieve this learns the system matrices by using a random control.
+Specifically, at each step the control $\mathbf{u}_{t}$ will be randomly selected from a Rademacher distribution, according to Algorithm 5.
 Algorithm 5 System identification via method of moments.
@@
   Set $\widehat{A^{j}B}=\frac{1}{T_{0}}\sum_{t=0}^{T_{0}-1}\mathbf{x}_{t(k+1)+j+1}\eta_{t(k+1)}^{\top}$, for all $j\in\{0,\dots k\}$.
-  Define $\widehat{C_{0}}=(\widehat{B},\dots\widehat{A^{k-1}B})$, $\widehat{C_{1}}=(\widehat{A^{1}B},\dots\widehat{A^{k}B})$, and return $\widehat{A},\widehat{B}$ as
+  Define $\widehat{C_{0}}=(\widehat{B},\dots\widehat{A^{k-1}B})$, $\widehat{C_{1}}=(\widehat{AB},\dots\widehat{A^{k}B})$, and return $\widehat{A},\widehat{B}$ as
 $\widehat{B}=\widehat{B},\quad\widehat{A}=\operatorname*{\arg\min}_{A}\|\widehat{C_{1}}-A\widehat{C_{0}}\|_{F}.$
@@
 $\mathop{\mbox{\bf E}}[\mathbf{x}_{t+k}\mathbf{u}_{t}^{\top}]=\mathop{\mbox{\bf E}}\left[\sum_{i=0:k}A^{i}\left(B\mathbf{u}_{t+k-i}+\mathbf{w}_{t}\right)\mathbf{u}_{t}^{\top}\right]=A^{k}B,$
-and the crucial fact that the controls are i.i.d. Thus, we can create an unbiased estimator for this matrix, and using many samples get an increasingly accurate estimate.
+and the crucial fact that the controls are i.i.d. Thus, we can create an unbiased estimator for this matrix, and using many samples, we obtain an increasingly accurate estimate.
 Proof.
@@
 $\|\widehat{C_{0}}-C_{0}\|,\|\widehat{C_{1}}-C_{1}\|\leq\sqrt{k}\varepsilon.$
-By this display, we already have $\|B-\widehat{B}\|\leq\varepsilon$. We will make use of this observation to establish the proximity of $\widehat{A}$ and $A$, whose proof is left as Exercise 8.7.
-Note that $\widehat{A}$ does not necessarily satisfy $\widehat{A}\widehat{C_{0}}=\widehat{C_{1}}$. Hence we require the following lemma.
+By this display, we already have $\|B-\widehat{B}\|\leq\varepsilon$.
+We will make use of this observation to establish the proximity of $\widehat{A}$ and $A$, using the following lemma, the proof of which is left as Exercise 8.7.
 Lemma 8.4.
-Let $X,X+\Delta X$ be the solutions to the linear system of equations $XA=B$ and $(X+\Delta X)(A+\Delta A)=B+\Delta B$. Then, as long as $\|\Delta A\|\leq\sigma_{\min}(A)$, we have
+Let $X,X+\Delta X$ be the solutions to the linear system of equations $XA=B$ and $(X+\Delta X)(A+\Delta A)=B+\Delta B$. Then, as long as $\|\Delta A\|<\sigma_{\min}(A)$, we have
 $\|\Delta X\|\leq\frac{\|\Delta B\|+\|\Delta A\|\|X\|}{\sigma_{\min}(A)-\|\Delta A\|}$
-We now use this lemma on the normal equations, i.e.
-$AC_{0}C_{0}^{\top}=C_{1}C_{0}^{\top}$ and $\widehat{A}\widehat{C_{0}}\widehat{C_{0}}^{\top}=\widehat{C_{1}}\widehat{C_{0}}^{\top}$.
+Note that $\widehat{A}$ does not necessarily satisfy $\widehat{A}\widehat{C_{0}}=\widehat{C_{1}}$.
+We now use the above lemma on the normal equations, that is, on $AC_{0}C_{0}^{\top}=C_{1}C_{0}^{\top}$ and $\widehat{A}\widehat{C_{0}}\widehat{C_{0}}^{\top}=\widehat{C_{1}}\widehat{C_{0}}^{\top}$.
 First, observe that for any $\varepsilon<1$
@@
 Therefore, using $\|A\|\leq\gamma$ and the perturbation result for linear systems, we have
-$\|A-\widehat{A}\|\leq\frac{12\gamma^{2}k}{\sigma_{\min}(A)}\varepsilon,$
-as long as $3k\kappa\varepsilon/\delta<\sigma_{\min}(C_{0})$. Finally, note that $(k,\kappa)$-strong controllability implies $\sigma_{\min}(C_{0})\geq 1/\kappa$. ∎
+$\|A-\widehat{A}\|\leq\frac{12\gamma^{2}k}{\sigma_{\min}(C_{0}C_{0}^{\top})}\varepsilon,$
+as long as $6k\gamma\varepsilon<\sigma_{\min}(C_{0}C_{0}^{\top})$. Finally, note that $(k,\kappa)$-strong controllability implies $\sigma_{\min}(C_{0}C_{0}^{\top})\geq 1/\kappa$. ∎
 8.4 From Indentification to Nonstochastic Control
@@
 The algorithms in this chapter address this by incorporating controlled perturbations in the control inputs, ensuring sufficient exploration while keeping regret low.
-However, the tradeoff results in a regret of $O(T^{2/3})$, rather than $O(\sqrt{T})$ we had in the last chapter.
+However, the trade-off results in a regret of $O(T^{2/3})$, rather than $O(\sqrt{T})$ we had in the last chapter.
     More sophisticated techniques for balancing exploration and exploitation can result in better regret bounds, some of which are surveyed in the bibliographic section.
-  • Regret Minimization as a Learning Metric Unlike traditional identification methods that aim for asymptotic convergence, online nonstochastic identification is evaluated in terms of regret, measuring performance relative to the best fixed model in hindsight.
+  • Regret Minimization as a Learning Metric Unlike traditional identification methods that aim at asymptotic convergence, online nonstochastic identification is evaluated in terms of regret, measuring the performance relative to the best fixed model in hindsight.
 8.5.1 Looking Ahead
 This part of the book considered the online nonstochastic control problem, the regret metric, and online learning techniques applied to control.
-We have thus far considered the case where the state is observable to the controller, an assumption which doesn’t always hold.
+We have thus far considered the case where the state is observable to the controller, an assumption which does not always hold.
 In the next parts of the book, we consider the setting of partial observability, where the system state is unknown, but an observed signal is available.
-We start with the problem of learning, and move onto online control.
+We start with the problem of learning and move onto online control.
 8.6 Bibliographic Remarks
 Controlling linear dynamical systems given unknown or uncertain system parameters has long been studied in adaptive control; see the text of [SBB90] for example.
-However, the results from adaptive control typically concern themselves with stability, and even when discussing implications for optimality present asymptotic bounds.
+However, the results from adaptive control typically concern themselves with stability and, even when discussing implications for optimality, present asymptotic bounds.
 In contrast, much of the work at the intersection of machine learning and control is focussed on precise non-asymptotic bounds on sample complexity or regret.
-[Fie97] which studies the discounted LQR problem under stochastic noise is the earliest example that furnishes a precise sample complexity bound.
-Similarly, the early work of [AYS11] uses an optimism-based approach to achieve an optimal regret in terms of the dependence on horizon.
+[Fie97] which studies the discounted LQR problem under stochastic noise is the earliest example that provides a precise sample complexity bound.
+Similarly, the early work of [AYS11] uses an optimism-based approach to achieve optimal regret in terms of dependence on horizon.
 This regret bound was later improved for sparse systems in [IJR12].
 [DMM+18, DMM+20] established regret and sample complexity results for the general stochastic LQR formulation with a polynomial dependence on natural system parameters.
-[FGKM18] studies the efficacy of policy-gradient-based approaches for synthesizing controllers for the LQR problem.
-The work of [MTR19] establishes that simple explore-then-commit strategies achieve the optimal dependence of regret on horizon, with [SF20] proving that such approaches are optimal in terms of their dependence on dimension of the system too.
-[CKM19] also gave a first optimal-in-horizon polynomial regret result simultaneously with the former work albeit using an optimism-based approach.
+[FGKM18] studies the efficacy of first-order optimization-based approaches in synthesizing controllers for the LQR problem.
+The work of [MTR19] establishes that simple explore-then-commit strategies achieve the optimal dependence of regret on horizon, with [SF20] proving that such approaches are optimal in terms of their dependence on dimension of the system, as well.
+[CKM19] also gave a first optimal-in-horizon polynomial regret result simultaneously with the former work, albeit using an optimism-based approach.
 The result of [PH20] gives a rate-optimal regret bound for unknown systems with convex costs.
 In terms of nonstochastic control, the method-of-moments system identification procedure and the corresponding analysis presented here appeared in [HKS20], which also gave the first end-to-end regret result for nonstochastic control starting with unknown system matrices.
-Even earlier, [SBR19] presented a regression procedure for identifying the Markov operators under nonstochastic noise.
-Both of these approaches have roots in the classical Ho-Kalman [HK66] system identification procedure, for which the first nonasymptotic sample complexity, although for stochastic noise, were furnished in [OO19].
+Even earlier, [SBR19] presented a regression procedure for identifying Markov operators under nonstochastic noise.
+Both of these approaches have roots in the classical Ho-Kalman system identification procedure [HK66], for which the first non-asymptotic sample complexity, although for stochastic noise, was furnished in [OO19].
 For strongly convex costs, logarithmic regret was established in [Sim20].
-Nonstochastic control of unknown, unstable systems without any access to prior information, also called blackbox control, was addressed in [CH21].
+Nonstochastic control of unknown, unstable systems without any access to prior information, also called black-box control, was addressed in [CH21].
 The Matrix Azuma inequality used in this chapter appears in [Tro12].
@@
 In previous chapters, we focused on control in dynamical systems, often assuming that the system state was fully observable and that system dynamics were either known or could be estimated through system identification.
-In many real-world scenarios, however, these assumptions do not hold.
+However, in many real-world scenarios, these assumptions do not hold.
 Instead, an agent may only have access to partial observations of the system and must learn to make accurate predictions without directly influencing the system’s evolution.
 This chapter focuses on learning in a dynamical system, which means predicting future system outputs based on past observations while the underlying system parameters remain unknown.
-Unlike control, where an agent actively influences system trajectories to minimize a given cost, learning in this setting is a passive task—the goal is to accurately forecast future outputs without modifying the system’s behavior.
-The ability to predict future states and outputs of an unknown system is fundamental to many domains, including time-series forecasting, financial modeling, sensor networks, autonomous systems, and most recently large language models.
+Unlike control, where an agent actively influences system trajectories to minimize a given cost, learning in this setting is a passive task; the goal is to accurately forecast future outputs without modifying the system’s behavior.
+The ability to predict future states and outputs of an unknown system is fundamental to many domains, including time series forecasting, financial modeling, sensor networks, autonomous systems, and most recently large language models.
 Moreover, learning a predictive model is often a key stepping stone toward solving more complex problems, such as control under partial observability or reinforcement learning in dynamical systems.
@@
 In this chapter, we focus on the specific task of prediction, where the objective is to estimate future outputs of a dynamical system based on past observations.
-Unlike classical system identification, which seeks to recover the system parameters explicitly, the learning approach taken here aims to optimize predictions directly.
+Unlike classical system identification, which seeks to recover the system parameters explicitly, the learning approach taken here aims to optimize the predictions directly.
 To formally define the problem, consider a sequence of observable inputs fed into a dynamical system, as illustrated in Figure 9.1.
@@
 Figure 9.1: In learning we passively observe the input sequence and attempt to predict output.
-For the remainder of this chapter, we consider a partially observable linear time invariant dynamical system described by the following equations.
-The extention of the material in this chapter to time varying linear dynamical systems is left as exercise 9.7.
+For the remainder of this chapter, we consider a partially observable linear time-invariant dynamical system described by the following equations.
+The extension of the material in this chapter to time-varying linear dynamical systems is left as exercise 9.7.
 $\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}$
@@
 Here, $\mathbf{x}_{t}$ represents the hidden system state, $\mathbf{u}_{t}$ is the observed control input, and $\mathbf{y}_{t}$ is the measured system output.
-The terms $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ account for process and observation noise, respectively.
+The terms $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ account for the process noise and the observation noise, respectively.
 A crucial challenge in this setting is that the system matrices $A,B,C$ are unknown to the learner.
-The learner’s objective is to construct an efficient prediction rule that estimates $\mathbf{y}_{t}$ based on past observations, despite the unknown system parameters.
+The learner’s objective is to construct an efficient prediction rule that estimates $\mathbf{y}_{t}$ based on past observations, despite not knowing the underlying parameters of the system.
 9.2 Online Learning of Dynamical Systems
-In an online learning setting, the learner sequentially observes data and makes predictions in real time. Formally, the process unfolds as follows:
+In an online learning setting, the learner sequentially observes the data and makes predictions in real time. Formally, the process unfolds as follows:
   1. 1.
@@
   2. 2.
-    The learner produces a prediction $\hat{\mathbf{y}}_{t}$ for the next output.
+    The learner makes a prediction $\hat{\mathbf{y}}_{t}$ for the next output.
   3. 3.
@@
 where $\hat{\mathbf{y}}^{\pi}_{t}$ denotes the prediction made at time $t$ by the best function $\pi\in\Pi$ in hindsight.
-This regret-based framework allows us to compare different classes of predictors and design learning algorithms that adapt efficiently to unknown dynamical systems.
+This regret-based framework allows us to compare different classes of predictors and design learning algorithms that efficiently adapt to unknown dynamical systems.
 9.3 Classes of Prediction Rules
@@
 This is analogous to the definitions and discussion in Chapter 6 where we discuss comparison between policy classes for control.
-However, the definitions and conclusions for prediction are simpler.
+However, the definitions and the corresponding conclusions for making predictions are simpler.
 Definition 9.1.
@@
 $\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{2})\leq\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{1})+\varepsilon T.$
-Proof of this lemma is left as an exercise to the reader at the end of this chapter. We now proceed to consider specific prediction rules.
+The proof of this lemma is left as an exercise to the reader at the end of this chapter. We now proceed to consider specific prediction rules.
 9.3.1 Linear Predictors
-The simplest prediction rule is a linear function of the observations, inputs, or both. More precisely,
+The simplest prediction rule is a linear function of observations, inputs, or both. More precisely,
 Definition 9.3 (Linear Policies).
@@
 The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $\mathbf{C}=I$ is the identity matrix.
-In this situation, the optimal prediction can be computed via the system evolution equation $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}$.
+In this situation, the optimal prediction can be calculated using the system evolution equation $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}$.
 It can be seen that this prediction rule is optimal for any zero-mean noise or perturbation model.
-However, this optimality is only true in the case of full observation of state.
-In the next chapter we consider optimal predictors for zero mean i.i.d noise in the more general partially observed case.
+However, this optimality is only true if the state is fully observable.
+In the next chapter, we consider optimal predictors for zero-mean i.i.d.
+noise in the more general partially observed case.
 For a partially observed linear dynamical system without noise we can write
@@
 $=\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t}\mathbf{x}_{0}.$
-Thus, the operator $\pi_{\{A,B,C,\mathbf{x}_{0}\}}$ takes a history of states and controls and gives the next system output accurately.
-We denote by $\prod^{\star}$ the set of all such predictors, that in the case of zero noise give a perfectly accurate prediction.
+Thus, the operator $\pi_{\{A,B,C,\mathbf{x}_{0}\}}$ takes a history of states and controls and gives the next output of the system accurately.
+We denote by $\prod^{\star}$ the set of all such predictors that in the case of zero noise give a perfectly accurate prediction.
 If the spectral radius of $A$ is larger than one, then the system is unstable and the magnitude of $\mathbf{y}_{t}$ can grow arbitrarily large, as discussed in section 4.2.
 Thus, for the rest of this section, assume that $A$ is $\gamma$-stable.
-With a decaying system, we can limit the memory of the predictor and yet maintain accuracy. Thus it makes sense to consider a more restricted class of predictors that have bounded memory, i.e.
+With a decaying system, we can limit the memory of the predictor and still maintain accuracy. Thus, it makes sense to consider a more restricted class of predictors that have bounded memory, that is,
 $\pi_{\{A,B,C,h\}}=\sum_{i=1}^{h}CA^{i-1}B\mathbf{u}_{t-i},$
-where we define all $\mathbf{u}$ vectors in negative times as zero. Denote by $\prod_{h}^{\star}$ the class of all such predictors.
-Clearly for a large $h$ this class of predictors well approximates the true observations $\mathbf{y}_{t}$ if the system is activated without any noise. We formalize what this means next.
+where as a matter of convention we say $\mathbf{u}_{t}=0$ for all negative $t$. Denote by $\prod_{h}^{\star}$ the class of all such predictors.
+Clearly, for a large $h$ this class of predictors well approximates the true observations $\mathbf{y}_{t}$ if the system is activated without any noise. We next formalize what this means.
 Lemma 9.4.
@@
 $\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|$
-$=\|\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t-1}\mathbf{x}_{0}-\sum_{i=1:h}CA^{i-1}B\mathbf{u}_{t-i}\|$
-$=\|\sum_{i=h}^{T}CA^{i}B\mathbf{u}_{t-i-1}\|$
+$=\left\|\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t-1}\mathbf{x}_{0}-\sum_{i=1:h}CA^{i-1}B\mathbf{u}_{t-i}\right\|$
+$=\left\|\sum_{i=h}^{T}CA^{i}B\mathbf{u}_{t-i-1}\right\|$
 $\leq\sum_{i=h}^{T}\kappa(1-\delta)^{i}\|C\|\|B\|\|\mathbf{u}_{t-i-1}\|$
@@
 9.4 Efficient Learning of Linear Predictors
-The previous section gives several natural classes of predictors as well some of the relationships between them.
-Notably, the class of linear dynamical predictors is optimal for certain scenarios and requires knowledge of only three matrices.
-However, the parametrization of this class is non-convex in the matrices $A,B,C$, which presents a computational challenge.
+The previous section gives several natural classes of predictors as well as some of the relationships between them.
+In particular, the class of linear dynamical predictors is optimal for certain scenarios and requires knowledge of only three matrices.
+However, the parameterization of this class is nonconvex in the matrices $A,B,C$, which presents a computational challenge.
 This challenge is circumvented by using the fact that the class $\prod^{L}$ $\varepsilon$-approximates the class $\prod^{\star}$.
-Further, the natural parametrization of $\prod^{L}$ is convex and thus can be learned using online convex optimization.
+In addition, the natural parameterization of $\prod^{L}$ is convex and thus can be learned using online convex optimization.
 The following algorithm learns the class of linear predictors in the sense that it attains a sublinear regret bound for any sequence of adversarial cost functions, as well as perturbations and observation noise.
@@
 We discussed key approaches for online learning in dynamical systems, emphasizing regret minimization as a performance metric.
-The regret framework allows for direct comparisons between learning algorithms by evaluating their cumulative prediction error against the best fixed predictor in hindsight.
+The regret framework allows direct comparisons between learning algorithms by evaluating their cumulative prediction error against the best fixed predictor in hindsight.
 A key takeaway from this chapter is that learning in LDS requires balancing expressivity and computational efficiency.
@@
 The task of predicting future observations in an online environment is often called “filtering”.
-Resting at the crossroads between statistical estimation and dynamical systems, the theory of filtering had its early beginnings in the works of Norbert Wiener [Wie49].
-For a comprehensive treatment of online learning in games and online convex optimization, see the texts [CBL06, Haz16].
+Resting at the crossroads between statistical estimation and dynamical systems, the theory of filtering had its early beginnings in the works of Norbert Wiener (see, e.g., [Wie49]).
+For a comprehensive treatment of online learning in games and online convex optimization, see [CBL06, Haz16].
 In the statistical noise setting, a natural approach for learning is to perform system identification and then use the identified system for prediction.
-Such an approach was taken by [SMT+18, SF20, SR19].
+This approach was taken by [SMT+18, SF20, SR19].
 The work of [GLS+20] extends identification-based techniques to adversarial noise and marginally stable systems.
 The class $\prod^{*}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [AHL+18].
-In that paper, it was shown how to learn partially observable, and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
+In that paper, it was shown how to learn partially observable and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
+Under probabilistic assumptions, the work of [HMR18] shows that the matrices $A,B,C,D$ of an unknown and partially observed time-invariant linear dynamical system can be learned using first-order optimization methods.
+Learning without recovery was studied in [GLS+20] even for marginally stable systems.
+More recently, tensor methods were used in [BLMY23a], and in [BLMY23b] to learn a mixture of linear dynamical systems.
 9.7 Exercises
-9.1. Prove that $\varepsilon$-approximation of prediction classes implies $\varepsilon T$ regret approximation, as formalized in Lemma 9.2.
-9.2. Prove that the class $\prod^{*}$ is optimal for fully observed linear dynamical systems with zero-mean noise.
-9.3. Show that the policy class $\prod^{\star}_{h}$ can be learned by policy class $\prod^{L}_{\tilde{h},k,\kappa}$ with $\tilde{h}=0$.
+9.1. Prove that a $\varepsilon$-approximation of prediction classes implies an additional $\varepsilon T$ regret, as formalized in the Lemma 9.2.
+9.2. Prove that the class $\prod^{*}$ is optimal for fully observed linear dynamical systems with zero mean noise.
+9.3. Show that the policy class $\prod^{\star}_{h}$ can be learned by the policy class $\prod^{L}_{\tilde{h},k,\kappa}$ with $\tilde{h}=0$.
 9.4. Give precise regret bounds for learning the class $\prod^{L}_{h,k,\kappa}$ and $\prod^{\star}_{h}$ as a function of the natural class parameters $h,k,\kappa$ and any other relevant parameter.
-9.5. Consider the square loss in algorithm 6. Give a Newton-method based algorithm based on the Online Newton Step algorithm (see [Haz16]) and write down the resulting regret bound.
+9.5. Consider the square loss in Algorithm 6. Give an algorithm based on the Online Newton Step algorithm (see [Haz16]) and write down the resulting regret bound.
 9.6.
-Write down the predictor class $\prod^{\star}$ and $\prod^{\star}_{h}$ for time varying linear dynamical systems.
-Research the notion of adaptive regret from [Haz16], and reason about conditions under which the class $\prod^{L}$ approximates $\prod^{*}$ for time varying linear dynamical systems.
+Write down the predictor class $\prod^{\star}$ and $\prod^{\star}_{h}$ for time-varying linear dynamical systems.
+Research the notion of adaptive regret from [Haz16], and reason about the conditions under which the class $\prod^{L}$ approximates $\prod^{*}$ for time-varying linear dynamical systems.
 Chapter 10 Kalman Filtering
@@
 The Kalman filter is a classical and widely used method for solving this problem, providing optimal state estimates under the assumption of linear system dynamics and Gaussian noise.
-A key feature of the Kalman filter—and a fundamental limitation—is that it requires knowledge of the system matrices, including both the state transition matrix and the observation model.
+A key feature of the Kalman filter and a fundamental limitation is that it requires knowledge of the system matrices, including both the state transition matrix and the observation model.
 These matrices encode how the system evolves over time and how observations relate to the hidden state.
-While the Kalman filter is provably optimal when these parameters are known, real-world applications often involve partially known or entirely unknown system dynamics, making explicit identification challenging.
-In contrast, online learning approaches, such as those introduced in Chapter 9, provide a model-free, open-loop alternative by directly learning optimal linear predictors from data.
+Although the Kalman filter is provably optimal when these parameters are known, real-world applications often involve partially known or entirely unknown system dynamics, making explicit identification challenging.
+In contrast, online learning approaches, such as those introduced in Chapter 9, provide a model-free open-loop alternative by directly learning optimal linear predictors from the data.
 Instead of estimating a latent state and system matrices, these methods construct forecasting models using past observations, making them more adaptive to unknown or time-varying systems.
-However, these approaches do not incorporate feedback, meaning they operate in an open-loop setting, predicting future states without correcting for new observations.
+However, these approaches do not incorporate feedback, meaning that they operate in an open-loop setting, predicting future states without correcting for new observations.
 This chapter moves beyond open-loop prediction to closed-loop state estimation, where past observations continuously refine future predictions.
@@
 Beyond classical Kalman filtering, we also introduce a closed-loop online learning method that serves as a convex relaxation of the Kalman filter.
-This approach retains the feedback mechanism of Kalman filtering while removing the need for explicit system identification, making it applicable in adversarial or uncertain environments.
+This approach retains the Kalman filtering feedback mechanism, while eliminating the need for explicit system identification, making it applicable in adversarial or uncertain environments.
 By formulating state estimation as an online convex optimization problem, this method provides provable regret guarantees and serves as a bridge between Bayesian filtering and learning-based estimation techniques.
@@
 10.1 Observable Systems
-Observability is a fundamental concept in control theory that determines whether the internal state of a system can be uniquely inferred from its outputs over time.
+Observability is a fundamental concept in control theory that determines whether the internal state of a system can be uniquely inferred from its output over time.
 In the infinite-horizon setting, prediction is meaningful only if observations remain bounded, a condition closely related to system controllability (as discussed in Section 4.3).
 A general dynamical system maps an initial state to an infinite sequence of observations.
-If this mapping is injective—meaning that each initial state corresponds to a unique sequence of observations—then the system is said to be observable.
+If this mapping is injective, meaning that each initial state corresponds to a unique sequence of observations, then the system is said to be observable.
 Formally:
@@
 This system allows only the velocity of the object to be observed, while the position remains hidden.
-It can be seen that this setting is unobservable, as the initial position cannot be determined based on the velocities and forces applied to the system alone.
+It can be seen that this setting is unobservable, as the initial position cannot be determined on the basis of the velocities and forces applied to the system alone.
 Indeed, this is a special case of the Galilean Relativity principle of Newtonian mechanics, which states that the laws of physics are the same in all inertial reference frames.
@@
 Theorem 10.2.
-A partially-observed dynamical system specified by system matrcies $(A,B,C)$ is observable if and only if $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))=d_{\mathbf{x}}$.
+A partially-observed dynamical system specified by system parameters $(A,B,C)$ is observable if and only if $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))=d_{\mathbf{x}}$.
 10.2 Optimality of the Kalman Filter
 10.2.1 The Power of Linear Predictors
-One of the important conclusions from this chapter is that the policy class of linear predictors is optimal with respect to closed loop policies. We prove this by proving two facts:
+One of the important conclusions from this chapter is that the policy class of linear predictors is optimal with respect to closed-loop policies. We prove this by proving two facts:
   1. 1.
@@
   2. 2.
-    The class of Kalman filters is optimal for closed loop policies.
-The conclusion from these two facts is that algorithm 6 can be used to learn a policy that competes with the best closed loop optimal predictor in hindsight.
+    The class of Kalman filters is optimal for closed-loop policies.
+The conclusion from these two facts is that algorithm 6 can be used to learn a policy that competes with the best closed-loop optimal predictor in hindsight.
 If the signal comes from a linear dynamical system subject to Gaussian noise, this means competing with the optimal predictor, which is a Kalman filter.
 We prove this formally at the end of this section.
-10.2.2 Closed Loop Optimal Predictors
+10.2.2 Closed-Loop Optimal Predictors
 Open-loop predictors, such as those introduced in the previous chapter, leverage system inputs and initial conditions to make predictions.
@@
 $\Sigma_{x}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[\mathbf{w}_{t}\mathbf{w}_{t}^{\top}\right],\quad\Sigma_{y}=\mathbb{E}\left[\mathbf{v}_{t}\mathbf{v}_{t}^{\top}\right].$
-So far, we have focused on predicting observations, i.e., $\mathbf{y}_{t}$.
+So far, we have focused on predicting observations, that is, $\mathbf{y}_{t}$.
 However, the Kalman filter also provides optimal state estimates $\hat{\mathbf{x}}_{t}$ given past observations $\mathbf{y}_{t-1:1}$ and controls $\mathbf{u}_{t-1:1}$.
@@
 $\mathbf{y}_{t}=C\mathbf{x}_{t}.$
-In Exercise 10.6, we ask the reader to supply a proof of this statement.
+In Exercise 10.6, we ask the reader to provide a proof of this statement.
 Based upon this observation, the following theorem is the essence of the Kalman filter, which recovers estimates for states as well as linear projection matrices for optimal prediction of the next observation.
 Theorem 10.3.
-Let the control sequence $\mathbf{u}_{t}$ be arbitrary affine functions of $\mathbf{y}_{1:t}$.
+Let the control sequence $\mathbf{u}_{t}$ be an arbitrary affine function of $\mathbf{y}_{1:t}$.
 There exists an efficiently computable sequence of linear predictors $\hat{\mathbf{x}}_{t+1}$ defined recursively as
@@
 Proof of Theorem 10.3.
-Since the estimator is recursively defined, a natural strategy is to proceed by induction. Let us say $\hat{\mathbf{x}}_{t}$ is a such a predictor. Further, let
+Since the estimator is defined recursively, a natural strategy is to proceed by induction. Let us say $\hat{\mathbf{x}}_{t}$ is such a predictor. Furthermore, let
 $\Sigma_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}.$
@@
 But, when predicting $\mathbf{x}_{t+1}$, we have more information.
-In particular, we have $\mathbf{y}_{t}$ too.
-Consider the random variables $\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}$.
+In particular, we also have $\mathbf{y}_{t}$.
+Consider random variables $\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}$.
 Can we use this covariate alone to predict $\mathbf{x}_{t+1}$?
@@
 In particular, any observable partially-observed linear dynamical system admits a solution to the Riccati equation listed above, and hence, has a steady-state Kalman filter.
-Mirroring the link between controllability and stabilizability, observability implies that $\rho(A-LC)<1$.
+Mirroring the link between controllability and stabilizability, observability implies $\rho(A-LC)<1$.
 10.2.6 Linear Predictors and the Kalman Filter
-For observable linear dynamical systems, the Kalman filter is contained in the class of linear predictors. This is fomally stated in the following Lemma, whose proof is left as Exercise 10.6.
+For observable linear dynamical systems, the Kalman filter is contained in the class of linear predictors. This is formally stated in the following lemma, whose proof is left as Exercise 10.6.
 Lemma 10.6.
@@
 In this subsection, we further restrict $\mathbf{w}_{t},\mathbf{v}_{t}$ to be i.i.d.
-draws from Gaussian distributions, and proceed to establish that the Kalman filter in such cases enjoys a stronger notion of optimality; it produces the the best estimator (under a least-squares loss) of the state given past observations among the class of all (not just linear) predictors.
+draws from Gaussian distributions, and proceed to establish that the Kalman filter in such cases enjoys a stronger notion of optimality; it produces the best estimator (under a least-squares loss) of the state given past observations among the class of all (not just linear) predictors.
 [^1] This is also called the Bayes-optimal predictor.
@@
 The setting we consider is that of Gaussian perturbations $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ and known systems matrices $(A,B,C)$.
-In particular, we consider the case when $\mathbf{w}_{t}\sim\mathcal{N}(0,\Sigma_{x})$ and $\mathbf{v}_{t}\sim\mathcal{N}(0,\Sigma_{y})$ are i.i.d.
-normally distributed draws, and $x_{0}=0$.
+In particular, we consider the case where $\mathbf{w}_{t}\sim\mathcal{N}(0,\Sigma_{x})$ and $\mathbf{v}_{t}\sim\mathcal{N}(0,\Sigma_{y})$ are i.i.d.
+Gaussian and $x_{0}=0$.
 Given the realizations $\mathbf{y}_{1},\mathbf{y}_{2}\dots,\mathbf{y}_{t}$ under some control input sequence $\mathbf{u}_{1},\mathbf{u}_{2},\dots\mathbf{u}_{t}$, we would like to compute a prediction $\hat{\mathbf{y}}_{t+1}$ that minimizes the mean square error, which is simply the expected observation as per
@@
 $\hat{\mathbf{x}}_{t+1}=(A-L_{t}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+L_{t}\mathbf{y}_{t},$
-for a sequence of matrices $L_{t}\in{\mathbb{R}}^{d_{x}\times d_{y}}$ that can be computed efficiently given $(A,B,C)$, and is independent of the realizations $\mathbf{u}_{1:T},\mathbf{y}_{1:T}$.
+for a sequence of matrices $L_{t}\in{\mathbb{R}}^{d_{x}\times d_{y}}$ that can be efficiently computed given $(A,B,C)$, and is independent of the realizations $\mathbf{u}_{1:T},\mathbf{y}_{1:T}$.
 Proof.
@@
 We also presented the filter’s steady-state behavior, particularly in the infinite-horizon setting, where it converges to a solution governed by the Riccati equation.
-Beyond its theoretical guarantees, the Kalman filter is widely used in practical applications, ranging from autonomous systems and robotics to finance and neuroscience.
+Beyond its theoretical guarantees, the Kalman filter is widely used in practical applications ranging from autonomous systems and robotics to finance and neuroscience.
 However, challenges remain, particularly in scenarios with non-Gaussian noise, time-varying dynamics, or adversarial disturbances.
-These limitations motivate extensions such as robust filtering, nonlinear filtering (e.g., the extended and unscented Kalman filters), and learning-based adaptations that integrate statistical inference with modern machine learning techniques.
+These limitations motivate extensions such as robust filtering, non-linear filtering (e.g., the extended and unscented Kalman filters) and learning-based adaptations that integrate statistical inference with modern machine learning techniques.
 The next chapter will transition from the classical filtering framework to spectral filtering, providing an alternative approach to state estimation that leverages spectral methods for efficient learning in dynamical systems.
 10.5 Bibliographic Remarks
-In 1960, Rudolph Kalman [Kal60] gave a state-space solution to the problem of filtering that was highly amenable to further extensions.
+In 1960, Rudolph Kalman gave a state-space solution to the filtering problem that was highly amenable to further extensions [Kal60].
 See [AM91] for a historical perspective on its development and aftermaths.
@@
   • Nonlinear Filtering Extensions.
-    The Kalman filter assumes linear system dynamics and Gaussian noise, limiting its effectiveness in nonlinear settings.
-The Extended Kalman Filter (see e.g.
-[UA21]) addresses this by linearizing the dynamics at each time step via a first-order Taylor series expansion.
-However, the linearization can introduce significant errors in highly nonlinear systems.
-Notable extensions using more sophisiticated sampling include the Unscented Kalman Filter [JU04], leading to better estimation accuracy in strongly nonlinear systems.
+The Kalman filter assumes that the dynamics is linear system and the noise is Gaussian, which limits its effectiveness in nonlinear settings.
+The Extended Kalman Filter (see, for example, [UA21]) addresses this issue by linearizing the dynamics at each time step via a first-order Taylor series expansion.
+However, linearization can introduce significant errors in highly non-linear systems.
+Notable extensions using more sophisticated sampling include the Unscented Kalman Filter [JU04], leading to better estimation accuracy in strongly nonlinear systems.
   • Particle Filtering (Sequential Monte Carlo Methods)
-    For highly nonlinear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative.
+    For highly non-linear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative.
 These methods approximate the posterior distribution using a set of weighted samples (particles) and employ importance sampling and resampling techniques.
@@
     Robust filtering methods seek to provide guarantees under adversarial noise.
-The $H_{\infty}$ filter [BB08] minimizes the worst-case estimation error rather than the mean-square error, making it suitable for applications with unknown disturbances.
-While the Kalman filter remains a cornerstone of state estimation, these alternative methods provide valuable extensions and improvements for nonlinear, adversarial, and data-driven settings.
-In the next chapter we will study an alternative using a technique which circumvents the need for state estimation altogether, and has stronger guarantees for sequential prediction.
+The $H_{\infty}$ filter [BB08] minimizes the worst-case estimation error rather than the mean-squared error, making it suitable for applications with unknown disturbances.
+Although the Kalman filter remains a cornerstone of state estimation, these alternative methods provide valuable extensions and improvements for nonlinear, adversarial, and data-driven settings.
+In the next chapter, we will study an alternative using a technique which circumvents the need for state estimation altogether and has stronger guarantees for sequential prediction.
 10.6 Exercises
@@
     Show that if the rank of the Kalman observability matrix is less than $d_{\mathbf{x}}$, then there are two initial states that after $d_{\mathbf{x}}$ iterations map to the same state.
-Hence the initial state is not uniquely determind.
+Hence, the initial state is not uniquely determined.
   2. 2.
-    Conversely, show that if the rank of $\hat{K}$ is full, then at any time $t$, $\mathbf{x}_{t}$ determines $\mathbf{x}_{0}$ uniquely.
+    In contrast, show that if the rank of $\hat{K}$ is full, then at any time $t$, $\mathbf{x}_{t}$ determines $\mathbf{x}_{0}$ uniquely.
 10.2. Prove Lemma 10.5, which states the following. Let $X,Y$ be possibly correlated random variables. Then
@@
 Prove this fact as formally stated in Lemma 10.6.
-Conclud that the Kalman filter can be learned using Algorithm 6.
-10.5. Prove Lemma 10.8, which is stated as follows. If a random variable pair $(\mathbf{x},\mathbf{y})$ are distributed as
+Conclude that the Kalman filter can be learned using Algorithm 6.
+10.5. Prove Lemma 10.8, which is stated below. If a random variable pair $(\mathbf{x},\mathbf{y})$ are distributed as
 $\begin{bmatrix}\mathbf{x}\\ \mathbf{y}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mu_{x}\\ \mu_{y}\end{bmatrix},\begin{bmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}\end{bmatrix}\right),$
@@
 Instead of estimating the full state and transition matrices, this approach constructs a forecasting model using past inputs and outputs.
-While this avoids the complexity of full system identification, it introduces a major limitation: the number of parameters required depends on the system’s stability constant, which can be large and unknown, making learning impractical for unstable or poorly conditioned systems.
+Although this avoids the complexity of full system identification, it introduces a major limitation: the number of parameters required depends on the stability constant of the system, which can be large and unknown, making learning impractical for unstable or poorly conditioned systems.
 In this chapter, we introduce Spectral Filtering, an alternative approach that bypasses these limitations.
-Rather than relying on state estimation or direct system identification, spectral filtering constructs predictors in a transformed basis, which does not depend on the system’s stability properties.
+Rather than relying on state estimation or direct system identification, spectral filtering constructs predictors on a transformed basis, which does not depend on the stability properties of the system.
 This method leverages spectral decompositions to efficiently extract the relevant dynamical structure, enabling more robust and scalable learning in both stable and unstable systems.
@@
 We then extend the discussion to higher-dimensional settings, demonstrating how spectral predictors enable a structured and computationally efficient representation of system dynamics.
-While this chapter primarily focuses on symmetric system matrices for clarity, the principles of spectral filtering extend to asymmetric and more general linear dynamical system settings.
+Although this chapter primarily focuses on symmetric system matrices for clarity, the principles of spectral filtering extend to asymmetric and more general linear dynamical system settings.
 These extensions require additional mathematical tools, which we reference in the bibliographic section.
 11.1 Spectral filtering in One Dimension
-In this section we give a simple introduction to the technique of spectral filtering, restricting ourselves to a single dimension.
-The restriction to a single dimension simplifies the exposition and yet is sufficient to illustrates the main ideas.
-In the next section we consider the high dimensional case.
-The single dimensional systems which we consider next are scalar systems, with $d_{u}=d_{y}=d_{x}=1$.
-We assume for simplicity that the initial state is zero $\mathbf{x}_{0}=0$.
+In this section, we give a simple introduction to the technique of spectral filtering, restricting ourselves to a single dimension.
+The restriction to a single dimension simplifies the exposition and is yet sufficient to illustrate the main ideas.
+In the next section, we consider the high-dimensional case.
+The one-dimensional systems which we consider next are scalar systems, with $d_{u}=d_{y}=d_{x}=1$.
+For simplicity, we assume that the initial state is zero $\mathbf{x}_{0}=0$.
 Recall the set of optimal predictors, the class $\Pi_{\star}$, which for scalar systems can written and simplified to,
@@
 The main idea of spectral filtering is to rewrite the vectors $\mu_{\alpha}$ in a more efficient basis, as opposed to the standard basis.
-The notion of efficiency stems from the definition of $\varepsilon$-approximation we have studied in previous chapters.
+The notion of efficiency comes from the definition of $\varepsilon$-approximation that we have studied in previous chapters.
 To define the basis, we consider the following matrix:
@@
 This theorem shows that the eigenvalues of a Hankel matrix decay very rapidly.
-A further illustration of this fact is given in Figure 11.1, which plots the eigenvalues of the matrix $Z_{T}$ for $T=30$ on a logarithmic scale.
+An additional illustration of this fact is given in Figure 11.1, which plots the eigenvalues of the matrix $Z_{T}$ for $T=30$ on a logarithmic scale.
 Figure 11.1: Eigenvalues of Hankel matrices decrease geometrically. These are the eigenvalues of the matrix $Z_{T}$, where $T=30$, plotted on a logarithmic scale.
@@
 The basis of eigenvectors of a Hankel matrix gives rise to an interesting class of predictors.
-This class of mappings do not directly apply to the history of inputs and/or observations.
+This class of mappings does not directly apply to the history of inputs and/or observations.
 Instead, we first apply a linear transformation over these vectors, and then apply a linear map.
@@
 Notice that as $h\mapsto T$, this class is equivalent to the class of all linear predictors on the inputs $\prod^{LIN}$.
-The reader may wonder at this point: Since a linear transformation composed with another linear transform remains linear, what then is the benefit here?
-The answer is that linear transformations are equivalent when we have full rank transforms.
-In our case, however, we consider low rank approximations to the transform, and the approximation properties of different subspaces can differ drastically!
+The reader may wonder at this point: Since a linear transformation composed with another linear transform remains linear, what is the benefit here?
+The answer is that linear transformations are equivalent when we have full-rank transforms.
+In our case, however, we consider low-rank approximations to the transform, and the approximation properties of different subspaces can differ drastically!
 This is the case in our setting: the spectral basis arising from the matrix $Z_{T}$ gives excellent approximations to the class of (standard) linear predictors, as we formally prove next.
@@
 Lemma 11.4.
-Let $\{\phi_{j}\}$ be the eigenvectors of Z. Then there exists constants $c_{1},c_{2}>0$ such that for all $j\in[T]$, $\alpha\in[0,1]$, we have
+Let $\{\phi_{j}\}$ be the eigenvectors of Z. Then there exist constants $c_{1},c_{2}>0$ such that for all $j\in[T]$, $\alpha\in[0,1]$, we have
 $|\phi_{j}^{\top}\mu_{\alpha}|\leq c_{1}e^{-\frac{c_{2}j}{\log T}}.$
-Using this property, we can prove Lemma 11.3, by simply shifting to the spectral basis representation as follows.
+Using this property, we can prove Lemma 11.3 by simply shifting to a representation on the spectral basis as follows.
 Proof of Lemma 11.3.
@@
 $\leq\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}+\mathbf{y}_{t-1}^{\pi_{2}}\|$$\triangle\mbox{-inequality}$
-$=\|\sum_{i=1}^{T}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}-\sum_{i=1}^{h}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}\|$
-$=\|\sum_{i=h+1}^{T}M^{\star}_{i}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-1}\|$
+$=\left\|\sum_{i=1}^{T}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}-\sum_{i=1}^{h}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}\right\|$
+$=\left\|\sum_{i=h+1}^{T}M^{\star}_{i}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-1}\right\|$
 $\leq\sum_{i=h+1}^{T}\|M_{i}^{\star}\||\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-1}|$
@@
 Consider a linear dynamical system with matrices $A,B,C$ and further assume that $A$ is symmetric.
-This implies that $A$ is diagonalizable over the real numbers as $A=U^{\intercal}DU$.
+This implies that $A$ can be diagonalized over the set of reals as $A=U^{\top}DU$.
 Decomposing over each dimension $j\in[d]$ separately, and denoting $D_{jj}=\alpha_{j}$, we can write (11.1):
@@
 11.3 Online Spectral Filtering
-The prediction class $\prod_{h}^{\lambda}$ can be learned by online gradient descent in the online convex optimization framework, as per Algorithm 3, giving rise to the following algorithm.
+The prediction class $\prod_{h}^{\lambda}$ can be learned by online gradient descent (Algorithm 3) in the online convex optimization framework, which gives rise to the following algorithm.
+Figure 11.2: The filters obtained by the eigenvectors of $Z$.
 Algorithm 7 Online Spectral Filtering
@@
 The eigenvectors of $Z$ have an interesting structure: their magnitude does not decay with time, which explains their improved approximation as captured formally in Lemma 11.3.
-Figure 11.2: The filters obtained by the eigenvectors of $Z$.
-The regret guarantees of the online gradient descent algorithm, as given in Theorem 6.1, directly implies the following corollary,
+The regret guarantees of the online gradient descent algorithm, as given in Theorem 6.1, directly imply the following corollary.
 Theorem 11.5.
@@
 In this chapter, we introduced Spectral Filtering as a novel approach to state estimation and prediction in linear dynamical systems, offering an alternative to classical methods like Kalman filtering and online learning of linear predictors.
-By leveraging spectral decompositions, we bypass the need for explicit system identification and provide a framework that is robust to system stability properties.
+Using spectral decompositions, we bypass the need for explicit system identification and provide a framework that is robust to the stability properties of the system.
 We began by outlining the core challenges of traditional state estimation approaches.
@@
 Spectral Filtering circumvents these limitations by constructing predictors in a transformed basis, allowing for efficient representation without requiring explicit system matrices.
-We demonstrated the method first in the one-dimensional case, providing intuition for how spectral predictors extract relevant dynamical structure.
+We demonstrated the method first in the one-dimensional case, providing an intuition for how spectral predictors extract relevant dynamical structure.
 This was then extended to higher-dimensional settings, where spectral filtering enables a structured and computationally efficient approach to forecasting.
@@
 From a methodological standpoint, we connected Spectral Filtering to convex optimization techniques, showing that it can be implemented via online learning in a spectral basis.
-This formulation allows for efficient sequential learning and prediction, with provable performance guarantees.
+This formulation allows for efficient sequential learning and prediction with provable performance guarantees.
 Looking ahead, Spectral Filtering opens the door for several exciting directions.
@@
 11.5 Bibliographic Remarks
-Under probabilistic assumptions, the work of [HMR18] shows that the matrices $A,B,C,D$ of an unknown and partially observed time-invariant linear dynamical system can be learned using first-order optimization methods.
-Deviating from classical recovery and into the improper prediction world, the work of [HSZ17] devised an improper learning technique using the eigenvectors of the Hankel matrix $Z$.
-The spectral properties of Hankel matrices used in this chapter are from [BT17].
+Deviating from classical recovery and into the world of improper prediction, the work of [HSZ17] devised an improper learning technique using the eigenvectors of the Hankel matrix $Z$.
+The methods we surveyed in previous chapters to learning linear dynamical systems had statistical and computational complexity that depended on the hidden dimension of the system.
+In contrast, the spectral filtering technique studied in this chapter has regret bounds and computational complexity guarantees that do not depend on the hidden dimension.
+The spectral properties of the Hankel matrices used in this chapter are from [BT17].
 The first extension of the spectral learning technique to asymmetric matrices is due to [HLS+18].
-A more recent advancement of spectral filtering to general linear dynamical systems was recently put forth in [MH25].
+A more recent advancement of spectral filtering for general linear dynamical systems was recently put forth in [MH25].
 The technique of spectral filtering has found applications in control and sequence prediction since its introduction.
-In [AHL+18], the authors leverage a spectral filtering technique to represent linear dynamical systems concisely in a linear basis, formulating the optimal control problem as a convex program.
-This method circumvents the non-convex challenges associated with explicit system identification and latent state estimation, providing provable optimality guarantees for the control signal.
-A recent line of work makes use of the spectral basis discovered in the line of work surveyed in this chapter in order to design neural architectures for sequence prediction.
-This was first suggested in the context of state space models in the spectral basis in [ASCH23], and extended to hybrid models [LND+24].
+In [AHL+18], the authors use a spectral filtering technique to concisely represent linear dynamical systems in a linear basis, formulating the optimal control problem as a convex program.
+This method circumvents the nonconvex challenges associated with explicit system identification and latent state estimation, providing provable optimality guarantees for the control signal.
+Closed-loop control using spectral filtering techniques was recently advanced in [BBDH25].
+A recent line of work makes use of the spectral basis discovered in the line of work surveyed in this chapter to design neural architectures for sequence prediction.
+This was first suggested in the context of state space models in the spectral basis in [ASCH23], and was extended to hybrid models in [LND+24].
 Later extensions explored efficient implementation [ACD+24] and the concept of length generalization in sequence prediction [MDA+24].
@@
 11.3.
-The fact that $Z_{T}$ is positive semidefinite is important to having the eigenvalues of $Z_{T}$ decay.
+The fact that $Z_{T}$ is positive semidefinite is important to ensure that the eigenvalues of $Z_{T}$ decay.
 Give an example of a family of $n\times n$ non-positive semidefinite Hankel matrices that have eigenvalues with constant absolute value.
@@
   1. (i)
-    Find a convex set $\mathcal{M}$ of tuples $(M_{1},M_{2},\dots)$’s to optimize across such that its diameter is some $\textrm{polylog}(T,1/\varepsilon)$.
+    Find a convex set $\mathcal{M}$ of tuples $(M_{1},M_{2},\dots)$’s to optimize over such that its diameter is some $\textrm{polylog}(T,1/\varepsilon)$.
 The norm we measure diameter with respect to is the sum of the Frobenius norms of the $M_{i}$’s.
   2. (ii)
-    One can further show that the norm of the gradient is at most $\textrm{polylog}(T,1/\varepsilon)$ as well.
-Use these two facts to show that with the right value of $\varepsilon$, the regret against the hypothesis class $\Pi_{\star}$ is at most $\tilde{O}(\sqrt{T})$.
+    In addition, show that the norm of the gradient is at most $\textrm{polylog}(T,1/\varepsilon)$ as well.
+Using these two facts, show that with the right value of $\varepsilon$, the regret against the hypothesis class $\Pi_{\star}$ is at most $\tilde{O}(\sqrt{T})$.
 Part IV Online Control with Partial Observation
@@
 In previous chapters, we explored control and learning techniques in fully observed linear dynamical systems, where the system state is directly accessible to the controller.
-However, in many real-world scenarios—such as robotics, economics, and signal processing—the state is only partially observed, requiring the controller to infer system dynamics from noisy or incomplete observations.
-This chapter focuses on partially observed LDS (PO-LDS), where the state is hidden, and only a linear projection of it is observed.
-For the remainder of this part of the text, we restrict ourselves to time-invariant systems, and refer to the bibliographic section for extensions.
+However, in many real-world scenarios, such as robotics, economics, and signal processing, the state is only partially observed, requiring the controller to infer system dynamics from noisy or incomplete observations.
+This chapter focuses on partially observed linear dynamical systems, where the state is hidden, and only a linear projection of it is observed.
+For the remainder of this part of the text, we restrict ourselves to time-invariant systems and refer to the bibliographic section for extensions.
 The system follows the general form:
@@
 Here, $\mathbf{x}_{t}$ represents the hidden system state, $\mathbf{u}_{t}$ the control input, $\mathbf{y}_{t}$ the observed output, and $\mathbf{w}_{t}$ the disturbance.
-Often, partially observed LDSs are defined with an additional noise term corrupting the observations.
+Often, partially observed LDSs are defined with an additional noise term that corrupts the observations.
 Exercise 12.6 asks the reader to prove that even in such circumstances we may assume the previously stated form for partially observed systems without loss of any generality.
 Unlike in the fully observed setting, where controllers can react based on the full state, partial observability introduces fundamental limitations on policy design.
-In the partial observation setting some of the control policy classes from chapter 6 are not well defined. We proceed to define policy classes and describe the relationship between them.
+In the partial observation setting, some of the control policy classes from Chapter 6 are not well defined. We then define policy classes and describe the relationship between them.
 Not all control policies designed for fully observed systems extend naturally to partially observed settings.
-In this chapter, we systematically explore different policy classes, analyzing their expressivity, computational tractability, and suitability for PO-LDS.
-  • Linear Observation-Action Controllers (LOACs): These controllers directly map observations to actions but are highly inexpressive, as they ignore past observations and fail to capture hidden state dependencies.
-  • Generalized Linear Observation-Action Controllers (GOACs): These extend LOACs by incorporating past history of observations, significantly improving expressiveness.
+In this chapter, we systematically explore different policy classes, analyzing their expressivity, computational tractability, and suitability for partially observed LDS.
+  • Linear Observation-Action Controllers (or simply Linear Controllers): These controllers directly map observations to actions, but are highly inexpressive, as they ignore past observations and fail to capture hidden state dependencies.
+  • Generalized Linear Observation-Action Controllers (or generalized linear controllers): These extend linear controllers by incorporating the history of observations, significantly improving expressiveness.
   • Linear Dynamic Controllers (LDCs): These policies model internal system dynamics, making them more powerful than simple observation-based policies.
-  • Disturbance Response Controllers (DRCs): Introduced as an alternative to state-based policies, DRCs directly act on a counterfactual signal—the “Nature’s y’s”—which captures the system’s response had no control been applied.
-We show that DRCs can approximate LDCs, making them a powerful convex relaxation.
+  • Disturbance Response Controllers (DRCs): Introduced as an alternative to state-based policies, DRCs act directly on a counterfactual signal, the “Nature’s y’s”, which captures the response of the system that result by uniformly setting the control input to zero.
+We show that DRCs can approximate LDCs, indicating that they are a highly expressive and yet convex policy class.
 We conclude by analyzing the relationships between these policy classes, showing how they fit into the broader framework of nonstochastic control with partial observation.
-This chapter lays the groundwork for online learning and regret minimization in PO-LDS, which we explore in the next chapter.
+This chapter lays the groundwork for online learning and regret minimization in partially observed LDS, which we will explore in the next chapter.
 Figure 12.1: Schematic relationship between linear Disturbance-Response Control policies and Linear Dynamical Control policies of increasing dimensions (or number of parameters) for linear time invariant systems.
@@
 A linear observation-action controller $K\in{\mathbb{R}}^{d_{u}\times d_{y}}$ is a linear operator mapping observation to control as $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}$.
-We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers bounded in Fronbenius norm by $\|K\|\leq\gamma$, that are also $\gamma$-stabilizing for given LDS.
-While this class of policies is natural to consider, it turns out to be very inexpressive.
+We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers bounded in the Fronbenius norm by $\|K\|\leq\gamma$, which are also $\gamma$-stabilizing for given LDS.
+Although this class of policies is natural to consider, it turns out to be very inexpressive.
 Even for optimal control with stochastic noise, the optimal policy cannot be approximated by a linear observation-action policy (see Exercise 12.6).
-Therefore we proceed to define more expressive and useful policies.
-However, LOACs will be useful in simplifying proofs later on.
+Therefore, we now define more expressive and useful policies.
+However, linear controllers will be useful in simplifying proofs later on.
 12.1.1 Generalized Linear Observation Action Controllers
@@
 This additional generality, of taking a history of observations, greatly increases the expressive power of this policy class.
-The formal definition is below.
+The formal definition is given below.
 Definition 12.2.
@@
 $\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}K_{i}\mathbf{y}_{t-i}.$
-Denote by $\Pi^{G}_{h,\gamma}$ be that class of all such policies satisfying $\sum_{i=0}^{h}\|K_{i}\|\leq\gamma$ that are also $\gamma$-stabilzing.
+By $\Pi^{G}_{h,\gamma}$, we denote the class of generalized linear observation-action controllers that satisfy $\sum_{i=0}^{h}\|K_{i}\|\leq\gamma$ and are also $\gamma$-stabilizing.
 Generalized observation-action policies can be seen as a special case of observation-action controllers using the technique of dimension lifting as follows.
@@
 $\ \ \tilde{C}=\begin{array}[]{|c|c|c|c|}\hline\cr C&C&\dots&C\\ \hline\cr\end{array}\ \ ,\ \ \tilde{D}=\begin{array}[]{|c|}\hline\cr I\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}.$
-The following lemma relates the original and lifted system (for a proof see exercise 12.6).
+The following lemma relates the original and lifted system (for a proof, see exercise 12.6).
 Lemma 12.3.
 For any sequence of control inputs $\mathbf{u}_{t}$, it holds that $\tilde{\mathbf{y}}_{t}=[\mathbf{y}_{t},\mathbf{y}_{t-1},\dots\mathbf{y}_{t-h}]$.
-The lemma above shows that generalized linear observation-action controllers are a special case of simple observation-action controllers such that $\mathbf{u}_{t}^{\pi}=[K_{0}\dots K_{h}]\tilde{\mathbf{y}}_{t}$ for the lifted system.
+The above lemma shows that generalized linear observation-action controllers are a special case of simple observation-action controllers such that $\mathbf{u}_{t}^{\pi}=[K_{0}\dots K_{h}]\tilde{\mathbf{y}}_{t}$ for the lifted system.
 12.2 Linear Dynamic Controllers
@@
 $\forall i\ ,\ \|A_{\pi}^{i}\|\leq\kappa(1-\delta)^{i}\ ,\ (\|B_{\pi}\|+\|C_{\pi}\|)\leq\kappa.$
-Note that LDCs are $\varepsilon$-approximated by generalized observation-action controllers, namely GOACs. The proof is left as Exercise 12.6.
+Note that LDCs are $\varepsilon$-approximated by generalized observation-action controllers. The proof is left as Exercise 12.6.
 Lemma 12.5.
@@
 The most important policy class in the context of nonstochastic control is that of Disturbance Response Controllers, or DRC.
-These policies act on a signal called Nature’s y’s, denoted $\mathbf{y}^{\mathbf{nat}}_{t}$, which is the would-be observation at time $t$ assuming all inputs to the system were zero from the beginning of time.
-To simplify presentation, we only consider linear time invariant linear dynamical systems that are stable.
-This allows us to ignore the stabilizing controller component of disturbance action control and greatly simplifies the exposition of the main ideas.
-Extension to time-varying systems is explored in the exercises.
+These policies act on a signal called Nature’s y’s, denoted $\mathbf{y}^{\mathbf{nat}}_{t}$, which is the would-be observation at time $t$ assuming that all inputs to the system were zero from the beginning of time.
+For ease of exposition, we only consider linear time invariant linear dynamical systems that are intrinsically stable.
+This allows us to ignore the stabilizing controller component of disturbance response control and greatly simplifies the presentation of the main ideas.
+The extension to time-varying systems is explored in the exercises.
 Definition 12.6 (Disturbance Response Controller).
-A disturbance response policy for stable systems $\pi_{M_{1:h}}$ is parameterized by the matrices $M_{0:h}=[{M}_{0},\ldots,{M}_{h}]$.
+A disturbance response policy for stable systems $\pi_{M_{1:h}}$ is parameterized by matrices $M_{0:h}=[{M}_{0},\ldots,{M}_{h}]$.
 It outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
@@
 where $\mathbf{y}^{\mathbf{nat}}_{t}$ is given by the system response to the zero control at time $t$, i.e.
-$\mathbf{x}^{\mathbf{nat}}_{t+1}$ $=A\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}=\sum_{i=0}^{t}A^{i}\mathbf{w}_{t-i}\ ,\ \mathbf{y}^{\mathbf{nat}}_{t+1}=C\mathbf{x}^{\mathbf{nat}}_{t+1}.$
-Denote by $\Pi^{R}_{h,\gamma}$ the set of all disturbance response policies that are $\gamma$-stabilizing 111for example, by having additional linear controllers $K_{1:T}$ that are $\gamma$-stabilizing linear controllers for the system and
+$\mathbf{x}^{\mathbf{nat}}_{t+1}$ $=A\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}=\sum_{i=0}^{t}A^{i}\mathbf{w}_{t-i}\ ,\ \mathbf{y}^{\mathbf{nat}}_{t}=C\mathbf{x}^{\mathbf{nat}}_{t}.$
+Denote by $\Pi^{R}_{h,\gamma}$ the set of all disturbance response policies that satisfy [^1] Recall that for this section, we restrict our discussion to stable systems.
+Otherwise, we would also ask the policy class to be $\gamma$-stabilizing, which can, for example, be achieved by superimposing a fixed $\gamma$-stabilizing linear dynamical controller.
 $\sum_{i=0}^{h}\|M_{i}\|\leq\gamma.$
-If the perturbations are all zero, seemingly DRC policies produce a zero control. Therefore we could add a constant offset to the control, that can be incorporated as in Definition 6.5.
+If the perturbations are all zero, seemingly DRC policies produce zero control. Therefore, we could add a constant offset to the control that can be incorporated as in Definition 6.5.
 The main advantage of this policy class, as opposed to previously considered policies in optimal control, is that they yield convex parametrization.
@@
 $\mathbf{y}^{\mathbf{nat}}_{t}$ $=\mathbf{y}_{t}-C\mathbf{z}_{t}.$
-We proceed to show that that the class of DRC approximates the class of LDC.
+We proceed to show that the class of DRC approximates the class of LDC.
 12.3.1 Expressivity of Disturbance Response Controllers
-A DRC always produces controls that ensure the system remains state-bounded. This can be shown analogously to Lemma 6.6 (see Exercise 12.6).
-While DRCs are contained inside the policy class of DACs, they are still very expressive.
+A DRC always produces controls that ensure that the system remains state-bounded. This can be shown analogously to Lemma 6.6 (see Exercise 12.6).
+Although DRCs are contained within the policy class of DACs, they are still very expressive.
 Note that a DAC policy can not be executed in a partially observed system from observations alone; even the knowledge of the entire observation sequence $\mathbf{y}_{t}$ may not uniquely determine $\mathbf{w}_{t}$.
-We have already seen that LDCs are approximately captured by GOACs in Lemma 12.5.
-It can be shown that the class of DRCs with appropriate parameters $\varepsilon$-approximates the class of GOACs for linear time-invariant systems.
-Note that this observation in conjunction with Lemma 12.5 implies that DRCs also approximate LDCs.
-To demonstrate that DRCs approximate GOACs, we instead prove that LOACs, a weaker class of policies, are $\varepsilon$-approximated by DRCs.
-Nevertheless, this observation is sufficient to conclude the stronger claim, given that GOACs can be written as LOACs on a lifted system as noted in Lemma 12.3, and simultaneously that a DRC with history $H^{\prime}$ on the lifted system can be written as a DRC with history $H^{\prime}+h$ over the original system.
+We have already seen that LDCs are approximately captured by generalized linear controllers in Lemma 12.5.
+It can be shown that the class of DRCs with appropriate parameters $\varepsilon$-approximates the class of generalized linear policies for linear time-invariant systems.
+This observation in conjunction with Lemma 12.5 implies that DRCs also approximate LDCs.
+To demonstrate that DRCs approximate generalized linear controllers, we instead prove DRCs approximate linear controllers, a weaker class of policies.
+However, this is sufficient to conclude the stronger claim.
+To see this, first recall that generalized linear controllers acting on the last $h$ states can be written as linear controllers on a lifted system with $h$ states, as noted in Lemma 12.3.
+Together with the next lemma, this implies that a DRC on the lifted system can approximate generalized linear controllers on the original system.
+However, this is not quite what we promised.
+To wrap up the argument, one can verify that a DRC with history $h^{\prime}$ on the lifted system can be written as a DRC with history $h^{\prime}+h$ on the original system.
 Lemma 12.8.
 Consider a $\gamma$-stable time-invariant linear dynamical system $(A,B,C)$.
-The class $\Pi^{R}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{L}_{\gamma}$ for parameters $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right)$, $\gamma^{\prime}=O(\gamma^{6})$.
+The class $\Pi^{R}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{L}_{\gamma}$ for the parameters $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right)$, $\gamma^{\prime}=O(\gamma^{6})$.
+Recall that for fully observed LDS, the state-action pairs produced by execution of a linear policy can be written as linear functions of the perturbations.
+This observation allowed us to prove that DACs capture linear policies for fully observed systems.
+We prove an analogue for partially observed systems: The observation-action pairs produced by the execution of linear policies can be written as linear functions of nature’s $y$’s.
+Thereafter, we argue that parameterizing the controllers to be linear in nature’s $y$’s allows us to encompass linear observation-action policies.
 Proof.
-Recall that for the fully observed case the state-action pairs produced by execution of a linear policy can be written as linear functions of the perturbations.
-It is this observation that allowed us to prove that DACs capture linear policies for fully observed systems.
-We prove an analogue for partially observed systems: The observation-action pairs produced by the execution of LOACs can be written as linear functions of nature’s $y$’s.
-Thereafter, we argue that parameterizing the controllers to be linear in nature’s $y$’s (as DRCs do) allows us to encompass LOACs.
-We introduce an augmented partially observed linear system that helps us write the observation-action pairs produced by a LOAC as a linear function of nature’s $y$’s.
-Crucially, the new system has the same nature’s $y$’s as the original system.
-Lemma 12.9.
+We introduce an augmented partially-observed linear system that helps us write the observation-action pairs produced by a linear policy as a linear function of nature’s $y$’s.
+In fact, we will shortly see that, for any sequence of control inputs, the augmented linear system produces the same sequence of obeservations as the original system.
+Thus, the new system has the same nature’s $y$’s as the original system.
 For any partially-observed system $(A,B,C)$ and perturbation sequence $\mathbf{w}_{t}$ producing observations $\mathbf{y}_{t}$, construct a lifted system $(\tilde{A},\tilde{B},\tilde{C})$ given by
 $\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t+1}\\ \Delta_{t+1}\end{bmatrix}=\begin{bmatrix}A&0\\ 0&A\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta_{t}\end{bmatrix}+\begin{bmatrix}0\\ B\end{bmatrix}\mathbf{u}_{t}+\begin{bmatrix}I\\ 0\end{bmatrix}\mathbf{w}_{t},\quad\tilde{\mathbf{y}}_{t}=\begin{bmatrix}C&C\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta_{t}\end{bmatrix}.$
-It holds uniformly for all time $t$ that $\widetilde{\mathbf{y}^{\mathbf{nat}}_{t}}={\mathbf{y}^{\mathbf{nat}}_{t}}$ and further for any sequence of control inputs $\mathbf{u}_{t}$ that $\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}$ and $\mathbf{x}_{t}=\mathbf{x}^{\mathbf{nat}}_{t}+\Delta_{t}$.
+Fix any sequence of control inputs $\mathbf{u}_{t}$.
+Let $\mathbf{x}_{t},\mathbf{y}_{t}$ be the states and observations produced by the original partially-observed system.
+The equivalents for the augmented system are $[{\mathbf{x}^{\mathbf{nat}}_{t}}^{\top},\Delta_{t}^{\top}]^{\top},\tilde{\mathbf{y}}_{t}$, as we have described above.
+Now we can verify for all time $t$ that $\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}$ and $\mathbf{x}_{t}=\mathbf{x}^{\mathbf{nat}}_{t}+\Delta_{t}$.
+Since this new system has the same input-output behavior as the original system, we might equivalently investigate how the execution of a linear controller $\pi$ affects the augmented system.
 Substituting $\mathbf{u}_{t}^{\pi}=K\mathbf{y}^{\pi}_{t}=KC(\mathbf{x}^{\mathbf{nat}}_{t}+\Delta^{\pi}_{t})$, the state-action sequence produced by a linear policy $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}^{\pi}$ can be written as
@@
 $\Delta^{\pi}_{t+1}=(A+BKC)\Delta^{\pi}_{t}+BK\mathbf{y}^{\mathbf{nat}}_{t},\quad\mathbf{y}^{\pi}_{t}=C\Delta^{\pi}_{t}+\mathbf{y}^{\mathbf{nat}}_{t},\quad\mathbf{u}^{\pi}_{t}=KC\Delta_{t}+K\mathbf{y}^{\mathbf{nat}}_{t}.$
-Now, define $M_{i}=KC(A+BKC)^{i-1}BK$ for $i>0$ and $M_{0}=K$, let $\mathbf{y}_{t},\mathbf{u}_{t}$ be the observation-action sequence produced by this DRC. We have
+Now, define $M_{i}=KC(A+BKC)^{i-1}BK$ for all positive $i\in[h^{\prime}]$ and $M_{0}=K$.
+These parameters specify our DRC.
+Let $\mathbf{y}_{t},\mathbf{u}_{t}$ be the observation-action sequence produced by this DRC.
+We have
 $\mathbf{u}^{\pi}_{t}$ $=K\mathbf{y}^{\mathbf{nat}}_{t}+KC\Delta^{\pi}_{t}$
@@
 $=\mathbf{u}_{t}+Z_{t}.$
-Given $\gamma$-stabilizability of the LOAC $K$ and $\gamma$-intrinsic stability of the of the system $(A,B,C)$, we have that for some $\kappa,\delta$ suc that $\frac{\kappa}{\delta}\leq\gamma$ that
+Given $\gamma$-stabilizability of the linear controller $K$ and $\gamma$-intrinsic stability of the of the system $(A,B,C)$, we have that for some $\kappa,\delta$ such that $\frac{\kappa}{\delta}\leq\gamma$,
 $\|Z_{t}\|$ $\leq\sum_{i>h^{\prime}}\|KC(A+BKC)^{i-1}BK\|\|\mathbf{y}^{\mathbf{nat}}_{t-i}\|$
-$\leq\sum_{i>h^{\prime}}\|KC(A+BKC)^{i-1}BK\|\left\|\sum_{j>1}CA^{j-1}B\mathbf{w}_{t-i-j}\right\|$
-$\leq\frac{\kappa^{5}}{\delta}\sum_{i>h^{\prime}}\|(A+BKC)^{i-1}\|\leq\frac{\kappa^{6}\gamma}{\delta}(1-\frac{1}{\delta})^{h^{\prime}}$
+$\leq\sum_{i>h^{\prime}}\|KC(A+BKC)^{i-1}BK\|\left(\sum_{j>1}\left\|CA^{j-1}\mathbf{w}_{t-i-j}\right\|\right)$
+$\leq\sum_{i>h^{\prime}}\|K\|^{2}\kappa(1-\delta)^{i-1}\sum_{j>1}\kappa(1-\delta)^{j}$
+$\leq\frac{\kappa^{4}}{\delta}\sum_{i>h^{\prime}}(1-\delta)^{i-1}\leq\frac{\kappa^{4}}{\delta^{2}}(1-{\delta})^{h^{\prime}}\leq\frac{\kappa^{4}}{\delta^{2}}e^{-\delta h^{\prime}}$
 Therefore, $\|\mathbf{u}_{t}-\mathbf{u}^{\pi}_{t}\|\leq\varepsilon$ by the appropriate choice of $h^{\prime}$. Once again using the intrinsic stability of the LTI $(A,B,C)$ , we have
@@
 $\leq\left\|\sum_{i=0}^{t}CA^{i}B\right\|\max_{i\leq t}\|\mathbf{u}^{\pi}_{i}-{\mathbf{u}}_{i}\|\leq\frac{\kappa}{\delta}\varepsilon.$
-The approximation in terms of observations and controls implies $\varepsilon$-approximation of policy class for any Lipschitz cost sequence, concluding the proof. ∎
+The approximation in terms of observations and controls implies $\varepsilon$-approximation of the policy class for any Lipschitz cost sequence, concluding the proof. ∎
 12.4 Summary
 In this chapter, we extended our study of control to the partially observed setting, where the controller no longer has direct access to the system state but must instead rely on noisy or incomplete observations.
-This shift fundamentally changes the nature of the control problem, requiring the design of policies that can infer and act based on limited information.
-We introduced a hierarchy of policy classes, starting with Linear Observation-Action Controllers (OACs), which directly map observations to actions, and progressing to Generalized LOACs (GOACs) and Linear Dynamic Controllers (LDCs), which incorporate history to improve decision-making.
+This shift fundamentally changes the nature of the control problem, requiring the design of policies that can infer and act on the basis of limited information.
+We introduced a hierarchy of policy classes, starting with Linear Observation-Action Controllers, which directly map observations to actions, and progressing to generalized linear policies and Linear Dynamic Controllers (LDCs), which incorporate history to improve decision-making.
 This increasing expressiveness highlighted a key limitation: memoryless policies are insufficient for optimal control, making structured state-dependent controllers necessary.
 A major contribution of this chapter was the introduction of Disturbance Response Controllers (DRCs).
-Instead of explicitly estimating the system state, DRCs act directly on a counterfactual signal—the “Nature’s y’s”—capturing what the system would have observed in the absence of control inputs.
+Instead of explicitly estimating the state of the system, DRCs act directly on a counterfactual signal, the “Nature’s y’s”, capturing what the system would have observed in the absence of control inputs.
 We established that DRCs are a convex relaxation of LDCs, making them computationally attractive while maintaining strong expressive power.
-This result provides a principled approach to designing policies that are both efficient and robust in partially observed environments.
+This result provides a principled approach to designing policies that are efficient and robust in partially observed environments.
 This chapter also set the stage for online nonstochastic control under partial observability, which we will explore in the next chapter.
@@
 12.5 Bibliographic Remarks
-The study of control in partially observed linear dynamical systems (PO-LDS) has a long history, with foundational contributions dating back to optimal estimation and control theories.
+The study of control in partially observed linear dynamical systems has a long history, with foundational contributions dating back to optimal estimation and control theories.
 This section provides an overview of key references related to the topics discussed in this chapter, including classical works, expressivity of policy classes, convex relaxations, and connections to online learning.
 The problem of control under partial observation was first rigorously formulated in the context of the Linear Quadratic Gaussian (LQG) regulator, which combines Kalman filtering with optimal control [AM79].
-The certainty equivalence principle, which asserts that optimal controllers can be designed assuming a perfect state estimate, was shown to have limitations in more general partially observed settings [Won68].
-This chapter introduced a hierarchy of policy classes, including Linear Observation-Action Controllers (OACs), Generalized OACs (GOACs), Linear Dynamic Controllers (LDCs), and Disturbance Response Controllers (DRCs).
+The certainty equivalence principle, which states that optimal controllers can be designed assuming a perfect state estimate, was shown to have limitations in more general partially-observed settings [Won68].
+This chapter introduced a hierarchy of policy classes, including linear observation-action controllers, generalized linear controllers, Linear Dynamic Controllers (LDCs), and Disturbance Response Controllers (DRCs).
 The concept of disturbance response controllers was inspired by parameterizations used in robust control, particularly the Youla-Kucera parameterization [YJB76].
-Convex relaxations have been a powerful tool for simplifying control design in PO-LDS.
+Convex relaxations have been a powerful tool for simplifying control design in partially observed LDS.
 We have surveyed convex relaxations for learning in dynamical systems in the previous part of this book.
-Recently, spectral methods for control [AHL+18] have provided an alternative approach for convex relaxation, leveraging spectral filtering to bypass direct system identification.
-The Disturbance Response Parametrization was introduced in [SSH20] as an analogue of DAC for control of partially observed systems.
+Recently, spectral methods for control [AHL+18] have provided an alternative approach to convex relaxation, utilizing spectral filtering to bypass direct system identification.
+The Disturbance Response Parametrization was introduced in [SSH20] as an analogue of DAC to control partially-observed systems.
 It can be viewed as a succinct state-space representation of the Youla parameterization [YJB76].
@@
 such that for any sequence of $\mathbf{u}_{t}$, $\mathbf{y}_{t}=\mathbf{y}^{\prime}_{t}$ holds uniformly over time.
-Additionally, verify that there exists such a construction of $\mathbf{z}_{t}$ where $\mathbf{w}^{\prime}_{t}$ are i.i.d.
+Additionally, verify that there exists a construction of $\mathbf{z}_{t}$ where $\mathbf{w}^{\prime}_{t}$ are i.i.d.
 as long as $(\mathbf{w}_{t},\mathbf{v}_{t})$ are i.i.d.
@@
 Hint: Consider the choice $\mathbf{w}^{\prime}_{t}=D^{\prime}_{t}\begin{bmatrix}\mathbf{w}_{t}\\ \mathbf{v}_{t}\end{bmatrix}$ for some $D^{\prime}_{t}$.
-Also consider using a hidden state $\mathbf{z}_{t}$ with a larger number of coordinates than $\mathbf{x}_{t}$.
+Also, consider using a hidden state $\mathbf{z}_{t}$ with a larger number of coordinates than $\mathbf{x}_{t}$.
 12.3.
@@
 Construct a partially observed linear dynamical system subject to Gaussian i.i.d.
-perturbations with quadratic costs such that simultaneously it is true that (a) there exists a GOAC policy that incurs zero (or at most constant) aggregate cost, (b) yet any LOAC policy incurs $\Omega(T)$ aggregate cost when run for $T$ time steps.
+perturbations with quadratic costs such that simultaneously it is true that (a) there exists a genearlized linear policy that incurs zero (or at most constant) aggregate cost, (b) yet any linear observation-action policy incurs $\Omega(T)$ aggregate cost when run for $T$ time steps.
 Hint: Consider a 3-dimensional hidden state so that $u_{t}$ can be chosen as a function of $y_{t-1}$ to ensure zero cost at each time step.
@@
 Unlike classical observer-based controllers, which rely on explicit state estimation, DRCs act directly on Nature’s y’s, the system’s counterfactual response in the absence of control.
-While DRCs offer a powerful alternative to state-based policies, they did not yet incorporate online learning, leaving open the question of how to adapt them dynamically in uncertain environments.
+Although DRCs offer a powerful alternative to state-based policies, we have not yet integrated them into an online learning algorithm, leaving open the question of how to adapt them dynamically in uncertain environments.
 This chapter takes the next step by introducing an online learning algorithm for optimizing DRCs in real time.
@@
   • Optimize control decisions dynamically, adapting to system behavior as new observations arrive.
-Unlike previous approaches that relied on fixed policy classes, GRC provides an adaptive control mechanism, bridging the gap between structured controllers and learning-based methods.
-A key focus of this chapter is extending policy regret minimization from the fully observed case to partial observability, introducing new challenges and solutions in learning-based control.
+Unlike previous approaches that relied on fixed policy classes, GRC provides an adaptive control mechanism that bridges the gap between structured controllers and learning-based methods.
+A key focus of this chapter is to extend policy regret minimization from the fully observed case to partial observability, introducing new challenges and solutions in learning-based control.
 To simplify the theoretical analysis, we assume system stability, allowing us to focus on the core learning problem without additional stabilization constraints.
@@
 The Gradient Response Controller algorithm is described in figure (8), for linear time-invariant dynamical systems with partial observation.
-Similar to the GPC, the main idea is to learn the parametrization of a disturbance response control policy using known online convex optimization algorithms, such as online gradient descent.
-We henceforth prove that this parametrization is convex, allowing us to prove the main regret bound.
-In algorithm (8) we denote $\mathbf{y}_{t}({M}_{1:h})$ as the observation arising by playing DRC policy ${M}_{1:h}$ from the beginning of time, and same for $\mathbf{u}_{t}({M}_{1:h})$.
+Similarly to GPC, the main idea is to learn the parameterization of a disturbance response control policy using known online convex optimization algorithms, such as online gradient descent.
+We henceforth prove that this parameterization is convex, allowing us to prove the main regret bound.
+In algorithm (8) we denote $\mathbf{y}_{t}({M}_{1:h})$ as the observation arising from playing the DRC policy ${M}_{1:h}$ from the beginning of time, and the same for $\mathbf{u}_{t}({M}_{1:h})$.
 Although it may be impossible to determine $\mathbf{w}_{t}$, especially if $C$ has fewer rows than columns, from the observations $\mathbf{y}_{t}$, $\mathbf{y}_{t}({M}_{1:h})$ can be determined from $\mathbf{y}^{\mathbf{nat}}_{t}$ alone by the application of Lemma 12.7, as
@@
 $\mathbf{y}^{\mathbf{nat}}_{t}=\mathbf{y}_{t}-C\sum_{i=0}^{t}A^{i}B\mathbf{u}_{t-i}.$
-5:     Construct loss $\ell_{t}({M}_{0:h})=c_{t}(\mathbf{y}_{t}({M}_{0:h}),\mathbf{u}_{t}({M}_{0:h}))$Update $M_{0:h}^{t+1}\leftarrow\prod_{\mathcal{K}}\left[{M}_{0:h}^{t}-\eta\nabla\ell_{t}({M}_{0:h}^{t})\right]$
-6:  end for
+5:     Construct loss $\ell_{t}({M}_{0:h})=c_{t}(\mathbf{y}_{t}({M}_{0:h}),\mathbf{u}_{t}({M}_{0:h}))$
+6:     Update $M_{0:h}^{t+1}\leftarrow\prod_{\mathcal{K}}\left[{M}_{0:h}^{t}-\eta\nabla\ell_{t}({M}_{0:h}^{t})\right]$
+7:  end for
 Algorithm 8 Gradient Response Controller(GRC)
-The GRC algorithm comes with a finite time performance guarantee: namely it guarantees vanishing worst case regret vs.
-the best disturbance response control policy in hindsight.
-As per the relationship between disturbance response control and other policies studied in section 12.3.1, this implies vanishing regret vs.
-linear dynamical controllers.
+The GRC algorithm comes with a finite-time performance guarantee: it guarantees vanishing worst-case regret versus the best disturbance response control policy in hindsight.
+Following the relationship between disturbance response control and other policies studied in Section 12.3.1, this implies vanishing regret versus linear dynamical controllers.
 The formal statement is given as follows.
@@
   • a
-    The costs $c_{t}$ are convex, bounded and have bounded gradients w.r.t. both arguments $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$.
+    The costs $c_{t}$ are convex, bounded and have bounded gradients with respect to the arguments $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$.
   • b
-    The diameter of the constrained set for the parameters $M_{0:h}$ is bounded by $D$.
+    The diameter of the constraint set for the parameters $M_{0:h}$ is bounded by $D$.
   • c
-    The matrices $\{{A},{B},C\}$ have bounded $\ell_{2}$ norm.
+    The matrices $\{{A},{B},C\}$ have bounded $\ell_{2}$ norms.
   • d
@@
 Proof.
-As per Theorem 6.1, the online gradient descent algorithm for convex loss functions ensures that the regret is bounded as a sublinear function of the number of iterations.
+According to Theorem 6.1, the online gradient descent algorithm for convex loss functions ensures that the regret is bounded as a sublinear function of the number of iterations.
 However, to directly apply the OGD algorithm to our setting, we need to make sure that the following conditions hold true:
   • The loss function should be a convex function in the variables ${M}_{0:h}$.
-  • We need to ensure that optimizing the loss function $\ell_{t}$ as a function of the parameters $M_{0:h}$ is similar to optimizing the cost $c_{t}$ as a function of the observation and control.
-First, we show that the loss function is convex w.r.t. to the variables ${M}_{0:h}$. This follows since the states and the controls are linear transformations of the variables.
+  • We must ensure that optimizing the loss function $\ell_{t}$ as a function of parameters $M_{0:h}$ is similar to optimizing the cost $c_{t}$ as a function of observation and control.
+First, we show that the loss function is convex with respect to the variables ${M}_{0:h}$. This follows since the states and the controls are linear transformations of the variables.
 Lemma 13.2.
@@
 $\ell_{t}({M}_{0:h})=c_{t}(\mathbf{y}_{t},\mathbf{u}_{t}).$
-Since the cost $c_{t}$ is a convex function w.r.t. its arguments, we simply need to show that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{0:h}$. Using Lemma 12.7,
+Since the cost $c_{t}$ is a convex function with respect to its arguments, we simply need to show that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{0:h}$. Using Lemma 12.7,
 $\mathbf{u}_{t}$ $=\sum_{i=0}^{h}M_{i}\mathbf{y}^{\mathbf{nat}}_{t-i},$
@@
 which is a linear function of the variables.
-Thus, we have shown that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ are linear transformations in ${M}_{0:h}$ and hence, the loss function is convex in ${M}_{0:h}$.
+Thus, we have shown that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ are linear transformations in ${M}_{0:h}$ and, hence, the loss function is convex in ${M}_{0:h}$.
 ∎
-Next, we use the following Lemma,
+Next, we use the following lemma.
 Lemma 13.3.
@@
 $\leq O(GDL\gamma^{3}h\sqrt{T}).$
-13.1.1 Proof of Lemma 13.3
-Proof.
-We first note that by the choice of step size by the Online Gradient Descent algorithm 3, as per Theorem 6.1, we have that $\eta=\frac{D}{G\sqrt{T}}$. Thus,
+13.1.1 Finishing up the Regret Bound
+Proof of Lemma 13.3.
+We first note that by the choice of step size by the Online Gradient Descent algorithm 3, as per Theorem 6.1, we have $\eta=\frac{D}{G\sqrt{T}}$. Thus,
 $\|M_{0:h}^{t}-M_{0:h}^{t-i}\|\leq\sum_{s=t-i+1}^{t}\|M_{0:h}^{s}-M_{0:h}^{s-1}\|\leq i\eta G=\frac{iD}{\sqrt{T}}.$
 Now, we use the fact that the iterates of OGD $M_{1:h}^{t}$ move slowly to establish that $\mathbf{y}_{t}$ and $\mathbf{y}_{t}(M^{t}_{0:h})$ are close in value.
-Before that, by definition, we have $\mathbf{u}_{t}(M^{t}_{0:h})=\mathbf{u}_{t}$ since the sequence $\mathbf{y}^{\mathbf{nat}}_{t}$ is determined independently of the choice of the policy being executed.
-Due to the $\gamma$-stability of the underlying system, we have for any $i>0$, $\|C{A}^{i}B\|\leq\kappa(1-\delta)^{i}$ for some $\kappa>0,\gamma<1$.
+Before that, by definition, we have $\mathbf{u}_{t}(M^{t}_{0:h})=\mathbf{u}_{t}$ since the sequence $\mathbf{y}^{\mathbf{nat}}_{t}$ is determined independently of the choice of the policy that is being executed.
+Due to the $\gamma$-stability of the underlying system, for any $i>0$, $\|C{A}^{i}B\|\leq\kappa(1-\delta)^{i}$ for some $\kappa>0,\gamma<1$.
 Using Lemma 12.7, we have for some $C>0$ that
@@
 Here we used the fact that $\|\mathbf{y}^{\mathbf{nat}}_{t}\|\leq\gamma$, which is left as an exercise.
-The control depends only on the parameters of the current iteration and we have that $\mathbf{u}_{t}=\mathbf{u}_{t}(M_{1:h}^{t})$.
+The control depends only on the parameters of the current iteration, and we have $\mathbf{u}_{t}=\mathbf{u}_{t}(M_{1:h}^{t})$.
 By definition, $l_{t}(M_{0:h}^{t})=c_{t}(\mathbf{y}_{t}(M_{0:h}^{t}),\mathbf{u}_{t}(M_{0:h}^{t}))$.
@@
 The GRC algorithm can be extended to time-varying dynamics as we explain in this section.
-The linear dynamical system at time $t$ is denoted by system matrices $A_{t},B_{t},C_{t}$, and we assume that they are known to the controller.
+The linear dynamical system at time $t$ is denoted by the system matrices $A_{t},B_{t},C_{t}$, and we assume that they are known to the controller.
 We further assume that these dynamical systems are all $\gamma$-stable.
@@
 The results in this chapter demonstrate that adaptive control under partial observability is possible without explicit state estimation, bridging the gap between classical observer-based controllers and modern learning-based methods.
-By leveraging regret minimization techniques, we provide a robust and flexible approach that applies even in adversarial and uncertain environments.
+Using regret minimization techniques, we provide a robust and flexible approach that applies even in adversarial and uncertain environments.
 13.4 Bibliographic Remarks
 Online nonstochastic control with partial observability was studied in [SSH20], who defined the Gradient Response Controller as well as the policy class of Disturbance Response Controllers.
-The setting was further extended to changing linear dynamics in [MGSH21], where the expressive power of different policy classes was also studied.
+The setting was further extended to time-varying linear dynamics in [MGSH21], where the expressive power of different classes of policy was also studied.
 The analogue of online nonstochastic control in the stochastic noise setting is called the Linear Quadraic Gaussian problem, and it is further limited by assuming quadratic noise.
 This classical problem in optimal control was proposed in Kalman’s work [Kal60].
-Regret bounds for controlling unknown partially observed systems were studied in [LAHA20b, LAHA20c, MTR19].
-[LAHA20a], which uses DRC policies even in context of stochastic control, obtains rate-optimal logarithmic bounds for the LQG problem.
-Control with partial information and bandit feedback with semi-adversarial noise was considered in [SNH23].
-Tight rates for quadratic control with bandit feedback, in the fully adversarial model, were obtained in [SSNH24], and more general cost functions in the bandit setting were considered in [SL24].
+Regret bounds for controlling unknown partially-observed systems were studied in [LAHA20b, LAHA20c, MTR19].
+[LAHA20a], which use DRC policies even in the context of stochastic control, obtain rate-optimal logarithmic bounds for the LQG problem.
+In [SNH23], the task of controlling a partially-observed system subject to semi-adversarial noise under bandit feedback was considered.
+The tight rates for quadratic control with bandit feedback, in the fully adversarial model, were obtained in [SSNH24], and more general cost functions in the bandit setting were considered in [SL24].
 13.5 Exercises
-13.1. In this exercise we consider the computation of nature’s y’s for time varying systems. Let
+13.1. In this exercise, we consider the computation of nature’s y’s for time-varying systems. Let
 $\mathbf{x}^{\mathbf{nat}}_{t+1}$ $=A_{t}\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}=\sum_{i=0}^{t}\left[\prod_{j=0}^{i-1}A_{t-j}\right]\mathbf{w}_{t-i}=\Phi_{t}\mathbf{w}_{1:t}.$
@@
 Prove that when a DRC policy $\pi\in\Pi^{R}_{H,\gamma}$ is executed on such a dynamical system, the resulting system is $\gamma^{2}$-stable.
-13.3. Prove that the optimal controller that minimizes the aggregate cost for a partially observed linear dynamical system assuming the perturbations are stochastic is a linear dynamical controller.
-13.4. Prove that the regret of the GRC controller for time varying linear dynamical systems is bounded by $O(\sqrt{T})$.
+13.3.
+Prove that the optimal controller that minimizes the aggregate cost for a partially observed linear dynamical system assuming that the perturbations are stochastic is a linear dynamical controller.
+13.4. Prove that the regret of the GRC controller for time-varying linear dynamical systems is bounded by $O(\sqrt{T})$.
 Hint: Follow the proof of Theorem 13.1 and modify as required.

### v6

@@
-version 0.9
+version 1.0
 [Uncaptioned image] Introduction to Online Control
-version 0.9
+version 1.0
 Elad Hazan    Karan Singh
@@
 Drew Bagnell
-Yann LeCun
-To my family: Dana, Hadar, Yoav, Oded and Deluca,
+As someone who has worked extensively on learning theory and online learning, and later applied these ideas in domains such as autonomous driving and humanoid robotics, I find this book both timely and inspiring.
+It introduces a regret-minimization framework for control that draws on the elegance and power of online learning.
+Traditional control theory often models noise either as stochastic—sometimes unrealistically optimistic—or adversarial—often overly conservative.
+This book charts a new path by asking a deeper question: while we cannot predict noise, can we perform nearly as well as if we could?
+The answer, developed here, is a novel and exciting paradigm that bridges learning theory and control, and I believe it will have a lasting impact on both research and practice.
+Shai Shalev-Shwartz
+To my family: Dana, Hadar, Yoav, Oded, and Deluca,
 —EH
@@
 In optimal control, robust control, and other control methodologies that assume stochastic noise, the goal is to perform comparably to an offline optimal strategy.
-In online nonstochastic control, both the cost functions as well as the perturbations from the assumed dynamical model are chosen by an adversary.
+In online nonstochastic control, both the cost functions and the perturbations from the assumed dynamical model are chosen by an adversary.
 Thus, the optimal policy is not defined a priori.
@@
 We remember and cherish the exact moment along with the mathematical writings in white chalk on that blackboard.
-Ever since then, the gem of an idea turned out into a complete theory which is now mainstream in machine learning and control, with numerous applications and refinements.
+Ever since then, the gem of an idea turned out into a complete theory that is now mainstream in machine learning and control, with numerous applications and refinements.
 Notable events that contributed to this theory include the following.
 Max Simchowitz, who was a graduate student then, visited Princeton during 2019-2020, where he collaborated with the group, and the theory was extended to partial observed systems.
-Paula Gradu, who was then an undergraduate, and Edgar Minasyan, while a grad student, contributed to the study of time-varying and nonlinear dynamical systems.
+Paula Gradu, who was then an undergraduate, and Edgar Minasyan, while a graduate student, contributed to the study of time-varying and nonlinear dynamical systems.
 Anirudha Majumdar joined for several important projects, notably for iterative linear control.
@@
 This book was further refined through teaching.
-Elad Hazan gratefully acknowledges the numerous contributions and insights of the students of the course computational control theory delivered at Princeton during COVID on Zoom during 2020-2021.
+Elad Hazan gratefully acknowledges the numerous contributions and insights of the students of the course computational control theory delivered at Princeton during COVID over Zoom during 2020-2021.
 The text was further enhanced by students of theoretical machine learning taught at Princeton University in the fall of 2024-2025.
-The following colleagues and students were particularly helpful, making insightful comments and in finding various typos and suggesting improvements: Gon Buzaglo, Anand Brahmbhatt, Sofiia Druchyna, and Jonathan Pillow (and many more we wish to announce very soon).
+The following colleagues and students were particularly helpful, making insightful comments, finding various typos and suggesting improvements: Gon Buzaglo, Anand Brahmbhatt, Sofiia Druchyna, Jonathan Pillow, and many more.
 Throughout the writing of this book, we engaged in extensive discussions with ChatGPT, an AI assistant.
@@
 General
-=$\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}$ definition
+$\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}$ definition
 $\operatorname*{\arg\min}\{\}$ the argument minimizing the expression in braces
@@
 $\mathbb{C}$ complex numbers
-$|A|$ determinant of matrix $A$
-${\bf Tr}(A)$ trace of matrix $A$
-$A\succeq 0$ $A$ is positive semi-definite, $\forall v\ ,\ v^{\top}Av\geq 0$
-$A\succeq B$ $A-B$ is positive semi-definite
+$|A|$ determinant of matrix AA
+${\bf Tr}(A)$ trace of matrix AA
+$\succeq 0 AA$ is positive semi-definite, $\forall v\ ,\ v^{\top}Av\geq 0$
+$A\succeq B$ ABA-B is positive semi-definite
 Control
@@
 $\mathbf{v}:S\to{\mathbb{R}}$ value function
-$Q:S\times A\to{\mathbb{R}}$ $Q$ function
+$S\times A\to{\mathbb{R}} QQ$ function
 $\pi:S\to A$ policy
@@
 $D_{p},G_{p}$ upper bound on the $p$-norm of the subgradients/diameter
-Part I Background in Control and RL
+Part I Background in Control and Reinforcement Learning
 Chapter 1 Introduction
@@
 Rather, we rethink the paradigm of control from the start.
-What can be achieved in a given dynamical system, whether known or unknown, fully observed or not, linear or non-linear, and without assumptions on the noise sequence?
+What can be achieved in a given dynamical system, whether known or unknown, fully observed or not, linear or nonlinear, and without assumptions on the noise sequence?
 This is a deep and difficult question that we fall short of answering.
@@
 This gives rise to a simple but powerful framework called nonstochastic control theory.
-The framework is a natural fit to apply techniques from online convex optimization, resulting in new, efficient, gradient-based control methods.
+The framework is a natural fit to applying techniques from online convex optimization, resulting in new, efficient, gradient-based control methods.
 1.2 The Origins of Control
-The industrial revolution was in many ways shaped by the steam engine that came to prominence in the 18th century.
+The industrial revolution was in many ways shaped by the steam engine that came to prominence in the eighteenth century.
 One of its most critical components was introduced by the Scottish inventor James Watt: the centrifugal governor, which is amongst the earliest examples of a control mechanism.
@@
 The impact of the centrifugal governor on the steam engine, and in turn on the industrial revolution, illustrates the importance of the field of control in engineering and the sciences.
-Since the 18th century, our world has changed quite a bit, and the importance of control has only grown.
+Since the eighteenth century, our world has changed quite a bit, and the importance of control has only grown.
 The challenges of controlling systems have moved from the steam engine to the design of aircraft, robotics, and autonomous vehicles.
@@
 Definition 1.2 (A generic control problem).
-Given a dynamical system with starting state $(\mathbf{x}_{1},\mathbf{y}_{1})$, iteratively construct controls $u_{t}$ to minimize a sequence of loss functions $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$, $c_{t}:{\mathbb{R}}^{d_{y}+d_{u}}\to{\mathbb{R}}$.
-This is given by the following mathematical program
+Given a dynamical system with starting state $(\mathbf{x}_{1},\mathbf{y}_{1})$, iteratively construct controls $u_{t}$ to minimize a sequence of loss functions $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$, $c_{t}\colon{\mathbb{R}}^{d_{y}+d_{u}}\to{\mathbb{R}}$.
+This is given by the following mathematical program:
 $\min_{\mathbf{u}_{1:T}}\sum_{t}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$
-$s.t.$ $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g_{t}(\mathbf{x}_{t}).$
+$s.t.\ \$$\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g_{t}(\mathbf{x}_{t}).$
 At this point, many aspects of the generic problem have not been specified. These include:
@@
 Instead of postulating a hypothesis about how our world should behave, we proceed to give an intuitive example of a control problem.
-We can then reason about the different questions and motivate the nonstochastc control setting.
+We can then reason about the different questions and motivate the nonstochastic control setting.
 1.3.1 Example: Control of a Medical Ventilator
@@
 The goal, then, is to choose a sequence of control signals $\{u_{t}\}_{t=1}^{T}$ to minimize the total cost over $T$ iterations.
-In the figure below we see an illustration of a simple controller, that we will soon describe, called the PID controller, as it performs on a ventilator and lung simulator.
+In figure 1.2 we see an illustration of a simple controller, which we will soon describe, called the PID controller, as it performs on a ventilator and lung simulator.
 Figure 1.2: Performance of the PID controller on a mechanical ventilator, from [SZG+21].
@@
 Bang-bang control is one of the simplest control strategies.
-Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a prespecified dynamic range $[x_{\min},x_{\max}]$ (which can vary over time) and where the scalar control input $u_{t}$ must at all times be within the interval $[u_{\min},u_{\max}]$, which is assumed to contain zero for simplicity.
+Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a prespecified dynamic range $x_{\min},x_{\max}]$ (which can vary over time) and where the scalar control input $u_{t}$ must at all times be within the interval $u_{\min},u_{\max}]$, which is assumed to contain zero for simplicity.
 A bang-bang controller chooses the control input at time $t$ as
@@
 A bang-bang controller saturates its inputs to extreme values in an attempt to confine the state of the system to a desirable range.
-For example, in the context of mechanical ventilation, one could choose $[p_{t}^{*}-\Delta p,p_{t}^{*}+\Delta p]$ as a reasonable region around the prescribed pressure $p^{*}_{t}$ as a desirable range for $p_{t}$.
+For example, in the context of mechanical ventilation, one could choose $p_{t}^{*}-\Delta p,p_{t}^{*}+\Delta p]$ as a reasonable region around the prescribed pressure $p^{*}_{t}$ as a desirable range for $p_{t}$.
 In such a case, a bang-bang controller would allow oxygen to enter at the highest possible rate whenever the pressure drops below $p^{*}_{t}-\Delta p$.
@@
 1.4.2 The PID Controller
-A more sophisticated control algorithm is the PID control method. Consider a control method that chooses the control to be a linear function of the observed states, i.e.
+A more sophisticated control algorithm is the PID control method. Consider a control method that chooses the control to be a linear function of the observed states, that is:
 $\mathbf{u}_{t}=\sum_{i=0}^{k}{A}_{i}\mathbf{x}_{t-i},$
@@
 However, other types of distribution can also be considered.
-As such, the control problem can be re-phrased in a more precise stochastic mathematical program,
+As such, the control problem can be rephrased in a more precise stochastic mathematical program,
 $\min_{\mathbf{u}_{1:T}}\mathop{\mbox{\bf E}}_{\mathbf{w}_{1:T}}\left[\sum_{t}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})\right]$
-$s.t.$ $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g(\mathbf{x}_{t}).$
+$s.t.\ \$$\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g(\mathbf{x}_{t}).$
 In contrast, robust control theory plans for the worst-case noise from a given set of constraints, denoted $\mathcal{K}$.
@@
 $\min_{\mathbf{u}_{1:T}}\max_{\mathbf{w}_{1:T}\in\mathcal{K}}\left[\sum_{t}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})\right]$
-$s.t.$ $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g(\mathbf{x}_{t}).$
+$s.t.\ \$$\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g(\mathbf{x}_{t}).$
 The two formulations have been debated for decades, with pros and cons for each side.
 Optimal control focuses on the average case and often yields simpler formulations for optimization, resulting in more efficient methods.
-Robust control, on the other hand, allows for more difficult noise models and some amount of model misspecification.
+Robust control, in contrast, allows for more difficult noise models and some amount of model misspecification.
 However, it is too pessimistic for many natural problems.
@@
 Planning for the worst case would mean slow and fuel-inefficient flight.
-We would like an adaptive control theory which allows us to attain the best of both worlds: an efficient and fast flight when the weather permits, and a careful drone maneuver when this is required by the conditions.
-Can we design a control theory that will allow us to take into account the specific instance perturbations and misspecifications, and give us finite time provable guarantees?
-This is the subject of online non-stochastic control!
+We would like an adaptive control theory that allows us to attain the best of both worlds: an efficient and fast flight when the weather permits, and a careful drone maneuvering when this is required by the conditions.
+Can we design a control theory that will allow us to take into account the specific instance perturbations and misspecifications, and give us finite-time provable guarantees?
+This is the subject of online nonstochastic control!
 1.6.1 Online Nonstochastic Control Theory
 This book is concerned with nonstochastic noise.
-When dealing with non-stochastic, i.e.
-arbitrary or even adversarial perturbations, the optimal policy, i.e.
-a decision making rule mapping observations to control inputs, is not clear a priori.
+When dealing with non-stochastic, that is, arbitrary or even adversarial perturbations, the optimal policy, that is, a decision making rule mapping observations to control inputs, is not clear a priori.
 Rather, an optimal policy for the observed perturbations can be determined only in hindsight.
@@
 For simplicity, we consider the case of full state observation where $\mathbf{y}_{t}=\mathbf{x}_{t}$.
-Let $\Pi=\{\pi:\mathbf{x}\to\mathbf{u}\}$ be a class of policies.
+Let $\Pi=\{\pi\colon\mathbf{x}\to\mathbf{u}\}$ be a class of policies.
 The regret of the controller with respect to $\Pi$ is defined as follows.
@@
 $\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}$ $\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}))\right),$
-where $\mathbf{u}_{t}={\mathcal{A}}(\mathbf{x}_{t})$ are the controls generated by ${\mathcal{A}}$, and $\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}$ are the counterfactual state sequence and controls under the policy $\pi$, i.e.
+where $\mathbf{u}_{t}={\mathcal{A}}(\mathbf{x}_{t})$ are the controls generated by $\mathcal{A}$}, and $\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}$ are the counterfactual state sequence and controls under the policy $\pi$, that is,
 $\mathbf{u}_{t}^{\pi}$ $=\pi(\mathbf{x}_{t}^{\pi})$
@@
     First, the best policy in hindsight out of all possible policies may be very complicated to describe and to reason about.
-This argument was made by the mathematician Tyrell Rockafellar, see the bibliographic section for more details.
+This argument was made by the mathematician Tyrell Rockafellar; see the bibliographic section for more details.
   2. 2.
@@
 The non-stochastic control problem can now be stated as finding an efficient algorithm that minimizes the worst-case regret vs.
-meaningful policies classes that we examine henceforth.
+meaningful policy classes that we examine henceforth.
 It is important to mention that the algorithm does not have to belong to the comparator set of policies!
@@
 1.6.2 A New Family of Algorithms
-The policy regret minimization viewpoint in control leads to a new approach for algorithm design.
+The policy regret minimization viewpoint in control leads to a new approach to algorithm design.
 Instead of a priori computation of the optimal controls, online convex optimization suggests modifying the controller according to the costs and dynamics, in the flavor of adaptive control, in order to achieve provable bounds on the policy regret.
@@
 2:for $t$ = $1\ldots T$ do
-3:      Use control $\mbox{Use control }\mathbf{u}_{t}=\sum_{i=1}^{h}{{M}}_{i}^{t}\mathbf{w}_{t-i}$
+3:   $\mbox{Use control }\mathbf{u}_{t}=\sum_{i=1}^{h}{{M}}_{i}^{t}\mathbf{w}_{t-i}$
 4:  Observe state $\mathbf{x}_{t+1}$, compute perturbation $\mathbf{w}_{t}=\mathbf{x}_{t+1}-f(\mathbf{x}_{t},\mathbf{u}_{t})$
@@
 Algorithm 1 Gradient Perturbation Controller(GPC) - simplified version
-The version given here assumes that the time-invariant dynamics $f$ is known.
+The version given here assumes that the time-invariant dynamics $f$ are known.
 The GPC computes a control that is a linear function of past perturbations.
@@
 The book of [Ber07] offers extensive coverage of topics from the point of view of dynamic programming.
-The texts [Ste94, ZDG96a], on the other hand, delve deep into results on linear control, the most well-understood branch of control theory.
+The texts [Ste94, ZDG96a], in contrast, delve deeply into results on linear control, the most well-understood branch of control theory.
 Rockafellar [Roc87] proposed the use of convex cost functions to model the cost in a linear dynamical system to incorporate constraints on state and control and suggested computational difficulties with this approach.
@@
 We give a simplified physical model that relates the pressure and the volume.
-This model, given in the following, can be derived by first principles from stress-strain relations, treating the left and right lungs connected by the trachea as situated in a two-balloon experiment.
+This model, given in the following, can be derived by first principles from stress-strain relations, treating the left and right lungs connected by the trachea as situated in a two-balloon experiment:
 $p_{t}=C_{0}+C_{1}v_{t}^{-1/3}+C_{2}v_{t}^{5/3}$
@@
 the real number line).
-The state $\mathbf{x}_{t}$ is 2-dimensional, with the first coordinate indicating the position of a point object and the second denoting its velocity.
+The state $\mathbf{x}_{t}$ is two dimensional, with the first coordinate indicating the position of a point object and the second denoting its velocity.
 The control algorithm decides the force $u_{t}\in{\mathbb{R}}$ that is applied to the object at each time step.
@@
 $c(\mathbf{x}_{t},u_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}x^{*}\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
-Figure 2.2: The position and velocity visualized for a double integrator.
+[…]
+Figure 2.2: Double integrator illustration showing state coordinate $x_{1}(t)$ (position), state coordinate $x_{2}(t)$ (velocity), mass $m$, and control input $u(t)$.
 2.1.3 Pendulum Swing-Up
@@
 where $\Delta$ is the time resolution of the controller.
-Note that this dynamical equation is non-linear in the state $\mathbf{x}_{t}=[\theta_{t},\dot{\theta_{t}}]^{\top}$ due to the presence of the term $\sin\theta$ in the evolution of $\dot{\theta_{t}}$.
-Figure 2.3: A pivoted massless rod connected to a point object with mass $m$.
-The point object is to subject to gravitational acceleration $g$.
-Illustration by Krishnavedala, licensed under CC-BY-SA-3.0.
+Note that this dynamical equation is nonlinear in the state $\mathbf{x}_{t}=[\theta_{t},\dot{\theta_{t}}]^{\top}$ due to the presence of the term $\sin\theta$ in the evolution of $\dot{\theta_{t}}$.
+PivotmmAngle $\theta_{t}$$m\mathbf{g}$$\mathbf{u}_{t}$$l$
+Figure 2.3: Pendulum swing-up illustration with marked angle $\theta_{t}$, gravitational force mgmg, control input torque $u_{t}$, mass $m$, and rod length $l$.
 The objective here is to get the rod to rest in a vertical upright position at an angle of $\pi$, using as little torque as possible throughout the process.
@@
 We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system, from [Boy10].
-The 2-dimensional control input $\mathbf{u}_{t}$ corresponds to the elevation angle and the thrust.
+The two dimensional control input $\mathbf{u}_{t}$ corresponds to the elevation angle and the thrust.
 The disturbances correspond to the velocity of the wind along the body axis and perpendicular to it.
 The following dynamics equation has matrices whose numerical values were computed using the dynamics equation for an actual aircraft.
-These numerical values can be determined by the mass, length, and other physical properties of the aircraft.
+These numerical values can be determined by the mass, length, and other physical properties of the aircraft:
 $\mathbf{x}_{t+1}=$ $\begin{bmatrix}-0.003&0.039&0&-0.322\\ -0.065&-0.319&7.74&0\\ 0.020&-0.101&-0.429&0\\ 0&0&1&0\end{bmatrix}\mathbf{x}_{t}+\begin{bmatrix}0.01&1\\ -0.18&-0.04\\ -1.16&0.598\\ 0&0\end{bmatrix}\mathbf{u}_{t}$
 $+\begin{bmatrix}-0.003&0.039&0&-0.322\\ -0.065&-0.319&7.74&0\end{bmatrix}\mathbf{w}_{t}$
-Figure 2.4: Longitudinal control in an aircraft. Graphic by User_A1, licensed under CC-BY-SA 4.0; it is an SVG derivative of a public domain work by G J Vigurs.
-The quantities of interest for assessing the performance of the controller are the aircraft speed, climb rate, each of which is a linear transformation of the state, and the applied thrust.
+Figure 2.4: Longitudinal control in an aircraft. Graphic by User_A1, licensed under $C$-BY-SA 4.0; it is an SVG derivative of a public domain work by G J Vigurs.
+The quantities of interest for assessing the performance of the controller are the aircraft speed, climb rate — each of which is a linear transformation of the state — and applied thrust.
 The corresponding cost is
@@
 Here, the coefficient $\beta$ measures the level of communicability of the infection, $\gamma$ measures the rate at which an infected person recovers, and $\alpha$ represents the efficacy of the vaccine in preventing infections.
-Figure 2.5: Trajectory of a continuous-time SIR model.
+Figure 2.5: Trajectory of a continuous-time SIR model with parameters $\alpha=0,\beta=0.3,\gamma=0.1$.
 2.2 Solution Concepts for Dynamical Systems
@@
 $\mathbf{x}_{t+1}={A}\mathbf{x}_{t}.$
-Clearly $\mathbf{0}$ is an equilibrium point as ${A}\mathbf{0}=\mathbf{0}$ for all ${A}$.
+Clearly $\mathbf{0}$ is an equilibrium point as ${A}\mathbf{0}=\mathbf{0}$ for all A{A}.
 If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vector converges to the null vector.
@@
 $\mathbf{x}_{t+1}^{\pi}=f(\mathbf{x}_{t}^{\pi},\pi(\mathbf{x}_{t}),\mathbf{w}_{t}).$
-Then the norm of the system’s state at every iteration is bounded by an absolute constant independent of $T$, i.e.
+Then the norm of the system’s state at every iteration is bounded by an absolute constant independent of $T$, that is,
 $\forall t\ ,\ \|\mathbf{x}_{t}^{\pi}\|\leq\gamma.$
@@
 Before formally proving the computational hardness of determining various solution concepts of dynamical systems, we start with a few intuitive and convincing examples of the complexities that arise.
-Consider the following dynamical system over the complex numbers
+Consider the following dynamical system over the complex numbers:
 $z_{t+1}=f_{c}(z_{t})=z_{t}^{2}+c.$
 The Julia set is now defined as
-$J_{c}=\{z\in C\ :\ \limsup_{t\to\infty}{|f_{c}^{t}(z)|}<\infty\},$
+$J_{c}=\{z\in C\ \colon\ \limsup_{t\to\infty}{|f_{c}^{t}(z)|}<\infty\},$
 where $f^{t}_{c}(x)=f_{c}(f_{c}(\dots f_{c}(x)))$ denotes $t$ successive applications of the function $f_{c}$.
-In other words, it is the set of all complex numbers $Z$ for which the system has a stable attractive equilibrium (which is not $\infty$).
-The boundary of the Julia set is very hard to determine, and it is known that determining if a ball around a given point belongs to the Julia set is a computationally undecidable problem, i.e.
-there is no Turing machine that decides it in finite time!
+In other words, it is the set of all complex numbers $Z$ for which the system has a stable attractive equilibrium (that is not $\infty$).
+The boundary of the Julia set is very hard to determine, and it is known that determining if a ball around a given point belongs to the Julia set is a computationally undecidable problem, that is, there is no Turing machine that can decide it in finite time!
 (See e.g.
-[BY08]) The complexity that arises near the boundary of the Julia set is depicted in figure 2.7.
-Another compelling visual evidence of the complexity of equilibrium in dynamical systems is the Lorenz attractor.
+[BY08].) The complexity that arises near the boundary of the Julia set is depicted in Figure 2.7.
+Another compelling visual illustration of the complexity of equilibrium in dynamical systems is the Lorenz attractor.
 This is a strange equilibrium resulting from simple dynamics that can be described using the following three equations over three variables, called the Lorenz equations [Lor63],
@@
 $z_{t+1}$ $=x_{t}y_{t}-\beta z_{t}.$
-The 3D trajectories of this system are attracted to a 2D manifold with two “wings” that resemble a butterfly as depicted in figure 2.7.
+The three dimensional (3D) trajectories of this system are attracted to a 2D manifold with two “wings” that resemble a butterfly, as depicted in Figure 2.7.
 For formal properties of this attractor set, see [Str14, Tuc99].
@@
 Figure 2.6: Julia set.
-Figure 2.7: Lorenz attractor. Visualization by D Schwen, licensed under CC-BY-2.5.
+Figure 2.7: Lorenz attractor. Visualization by D Schwen, licensed under $C$-BY-2.5.
 2.3.1 Computational Hardness of Stabilizability
-Consider the following problem: given a dynamical system, can we determine if it stabilizable, i.e., for every possible starting state, does there exist a choice of control inputs that drives the system to the null state?
+Consider the following problem: Given a dynamical system, can we determine if it stabilizable, that is, for every possible starting state, does there exist a choice of control inputs that drives the system to the null state?
 We now show that this decision problem is computationally hard.
@@
 decide if there exists a satisfying Boolean assignment.
-Given such a formula $F$, consider the following polynomial $p_{F}(\mathbf{u})$. For each clause $c$ in $F$, construct a degree-3 polynomial as
+Given such a formula $F$, consider the following polynomial $p_{F}(\mathbf{u})$. For each clause $c$ in $F$, construct a degree-three polynomial as
 $g_{x_{i}\lor x_{j}\lor x_{k}}(\mathbf{u})$ $\rightarrow(1-u_{1})\cdot(1-u_{2})\cdot(1-u_{3})$
@@
 $g_{\neg x_{i}\lor\neg x_{j}\lor\neg x_{k}}(\mathbf{u})$ $\rightarrow u_{1}\cdot u_{2}\cdot u_{3},$
-and the degree-6 polynomial for the entire formula as
+and the degree-six polynomial for the entire formula as
 $p_{F}(\mathbf{u})=\sum_{i=1}^{n}(u_{i}(1-u_{i}))^{2}+\sum_{j=1}^{m}(g_{c_{j}}(\mathbf{u}))^{2}.$
@@
 The text [Ted20] offers several examples and characterizations of dynamical systems useful for robotic manipulation. A more mathematically involved introduction appears in [Str14].
-[Ada20] discusses further aspects of simulations used in the early days of the COVID-19 pandemic. See [SZG+21] for examples of control algorithms used for non-invasive mechanical ventilation.
+[Ada20] discusses further aspects of simulations used in the early days of the COVID-19 pandemic. See [SZG+21] for examples of control algorithms used for noninvasive mechanical ventilation.
 The Lorenz equations were introduced to model the phenomenon of atmospheric convection [Lor63].
@@
 $A=\begin{bmatrix}1&\Delta\\ 0&1\end{bmatrix},\quad B=\begin{bmatrix}0\\ 1\end{bmatrix},\quad\mathbf{x}_{t}=\begin{bmatrix}x_{t}\\ v_{t}\end{bmatrix}\in\mathbb{R}^{2},\quad u_{t}\in\mathbb{R},\quad\Delta>0.$
-The goal is to ensure that the system reaches the desired state $[x^{*},0]^{\top}$ as $t\to\infty$ from any initial condition.
+The goal is to ensure that the system reaches the desired state $x^{*},0]^{\top}$ as $t\to\infty$ from any initial condition.
   1. 1.
@@
     Design a stabilizing state-feedback law
-[…]
-    with $K=[k_{1}\;k_{2}]$, which places all eigenvalues of $A-BK$ strictly within the unit circle.
-Find conditions on $\Delta,k_{1},k_{2}$ that ensure that all eigenvalues of $A-BK$ are real and lie strictly between $-1$ and $1$, thus guaranteeing $\mathbf{x}_{t}\to[x^{*},0]^{\top}$.
+    $u_{t}=-K(\mathbf{x}_{t}-[x^{*}\;0]^{\top}),$
+    with $K=[k_{1}\;k_{2}]$, which places all eigenvalues of ABKA-BK strictly within the unit circle.
+Find conditions on $\Delta,k_{1},k_{2}$ that ensure that all eigenvalues of ABKA-BK are real and lie strictly between 1-1 and $1$, thus guaranteeing $\mathbf{x}_{t}\to[x^{*},0]^{\top}$.
 Chapter 3 Markov Decision Processes
@@
 Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel and Norvig [RN02].
-In this example, a robot starts in the bottom left square, and its goal is to reach the $+1$ reward.
-The blocks $+1,-1$ represent terminal states, which means that the game ends when the robot reaches them.
+In this example, a robot starts in the bottom left square, and its goal is to reach the +1+1 reward.
+The blocks +1,1+1,-1 represent terminal states, which means that the game ends when the robot reaches them.
 The block ‘X’ represents an obstacle through which the robot cannot travel.
-StartX+1-1
+StartX+1+11-1
 Figure 3.2: A visual representation of Bubbo’s world.
@@
 This mapping is called a “policy”, and depends both on the structure of the problem and on the rewards.
-For example, if the rewards in all other states are slightly negative, Bubbo wants to reach the terminal state with a persistent $+1$ reward, as given in the following figure.
+For example, if the rewards in all other states are slightly negative, Bubbo wants to reach the terminal state with a persistent +1+1 reward, as given in figure 3.3.
 […]
 Figure 3.3: The optimal policy (mapping from state to action) with reward for each step taken being -0.04.
-On the other hand, if the rewards are very negative, Bubbo just wants to reach any terminal state as soon as possible, as in Figure 3.4.
+Conversely, if the rewards are very negative, Bubbo just wants to reach any terminal state as soon as possible, as in Figure 3.4.
 […]
@@
 The optimal policy changes: now the robot is more concerned with ending the game as quickly as possible rather than trying to get to the +1 reward.
-We then describe a formal model to capture this and any other reinforcement learning problem.
+We next describe a formal model to capture this and any other reinforcement learning problem.
 3.2 Markov Decision Processes
 In this chapter, we describe the general and formal model for reinforcement learning: Markov decision processes.
-To ease the exposition, we start with a review of Markov Chains, a.k.a.
-Markov Processes, then proceed to more general Markov Reward Proccesses, and conclude with the full Markov Decision Process definition.
+To ease the exposition, we start with a review of Markov Chains, also known as Markov Processes, then proceed to more general Markov Reward Proccesses, and conclude with the full Markov Decision Process definition.
 3.2.1 Markov Chains
@@
 The transition matrix is of dimension $S\times S$, where $S$ is the number of states, and in entry $i,j$ we have the probability of moving from state $i$ to state $j$.
-Figure 3.5 describes a Markov Chain corresponding to a weather process with a corresponding transition matrix $P$ as below
+Figure 3.5 describes a Markov Chain corresponding to a weather process with a corresponding transition matrix $P$ as follows:
 $P=\begin{pmatrix}0.9&0.1&0&0\\ 0.5&0&0.4&0.1\\ 0&0.5&0&0.5\\ 0&0&0&1\\ \end{pmatrix}$.
 RainSunny dayCloudySnow0.10.90.50.10.40.50.51
-Figure 3.5: A graphical description of a Markov decision process. Each state is the weather for a given day and the transition probabilities describe the chance of weather for the next day.
+Figure 3.5: A graphical description of a Markov decision process. Each state is the weather for a given day and the transition probabilities describe the chances of weather for the next day.
 Denote the distribution over states at time $t$ by $\mathbf{p}_{t}$. Then it can be seen that
@@
   2. 2.
-    Aperiodicity. A Markov Chain is aperiodic if and only if $d(s_{i})=1$ for all states $s_{i}\in S$, where
-    $d(s_{i})=\gcd\left\{t\in\mathbb{N}_{+}\ \ |\ \ (P^{t})_{ii}>0\right\},$
-    is the periodicity of state $s_{i}$, where $\gcd$ stands for “greatest common divisor".
+    Aperiodicity.
+A Markov Chain is aperiodic if and only if $d(s_{i})=1$ for all states $s_{i}\in S$, where d(s_i) = gcd{t ∈N_+ | (P^t)_ii > 0 } , is the periodicity of state $s_{i}$, where $\gcd$ stands for “greatest common divisor".
 The ergodic theorem states that every irreducible and aperiodic Markov chain has a unique stationary distribution.
@@
   3. 3.
-    $\hat{R}\in{\mathbb{R}}^{S}$ = the vector that for each state, denotes the expected reward in the next iteration as a function of the current state,
-    $\hat{R}(s)=\mathop{\mbox{\bf E}}[R_{t}|S_{t}=s].$
+    $\hat{R}\in{\mathbb{R}}^{S}$ = the vector that for each state, denotes the expected reward in the next iteration as a function of the current state, ^R(s) = E[ R_t | S_t = s] .
   4. 4.
-    $G_{t}=$ the random variable denoting the sum of the discounted rewards starting from time $t$ onwards.
-    $G_{t}=\sum_{i=0}^{\infty}R_{t+i}\gamma^{i}.$
+    $G_{t}=$ the random variable denoting the sum of the discounted rewards starting from time $t$ onwards. G_t = ∑_i=0^∞ R_t+i γ^i .
 When the discount factor is strictly less than one, $G_{t}$ is well defined.
@@
 $\mathbf{v}(s)=\mathbb{E}[G_{t}|S_{t}=s].$
-The value function tells how valuable a state is as measured by the sum of its future discounted rewards.
+The value function says how valuable a state is as measured by the sum of its future discounted rewards.
 One way to compute the value function is by using the Bellman equation for the Markov Reward Process, which is the following recursive equation:
@@
   2. 2.
-    $A$ is a set of possible actions.
+    AA is a set of possible actions.
   3. 3.
-    $P$ is a transition matrix:
-    $P_{ss^{\prime}}^{a}=\Pr[S_{t+1}=s|S_{t}=s,A_{t}=a].$
+    $P$ is a transition matrix: P_ss’^a = Pr[ S_t+1 = s | S_t = s, A_t = a].
   4. 4.
@@
 We henceforth assume w.l.o.g.
-that $R\in[0,1]^{S\times A}$, where we define
-    $R_{sa}=\sum_{s^{\prime}\in S}P_{s,s^{\prime},a}R_{s,s^{\prime},a}.$
+that $R\in[0,1]^{S\times A}$, where we define R_sa = _s’ S P_s,s’,a R_s,s’,a .
   5. 5.
@@
 $Q_{\pi}(s,a)=\mathop{\mbox{\bf E}}_{\pi,P}\left[\sum_{t=1}^{\infty}\gamma^{t-1}R_{t}|S_{1}=s,A_{1}=a\right].$
-A fundamental theorem of reinforcement learning is as follows.
+A fundamental theorem of reinforcement learning, which is constructively proven in Section 3.4, is as follows.
 Theorem 3.6.
@@
   2. 2.
-    $P^{\pi}$ is the induced transition matrix when playing policy $\pi$, i.e.
-    $P^{\pi}_{ss^{\prime}}=\sum_{a\in A}P^{a}_{ss^{\prime}}\cdot\pi(a|s).$
+    $P^{\pi}$ is the induced transition matrix when playing policy $\pi$, i.e. P^π_ss’ = ∑_a ∈A P^a_ss’ ⋅π(a|s).
 The Bellman equation gains prominence when applied to the optimal policy of a Markov Decision Process.
@@
 Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories.
-Created by XaosBits using the Mathematica packages CurvesGraphics.nb by Gianluca Gorni and DrawGraphics.nb by David and Alice Park, licensed under CC-BY-2.5.
+Created by XaosBits using the Mathematica packages CurvesGraphics.nb by Gianluca Gorni and DrawGraphics.nb by David and Alice Park, licensed under $C$-BY-2.5.
 A more general class of linear dynamical systems is when the matrices ${A}_{t}$ and ${B}_{t}$ change over time.
@@
 Determining whether a linear time-invariant system is stabilizable is determined by a particularly simple condition as follows.
-Henceforth, we denote by $\rho({A})$ the spectral radius of a matrix ${A}$, that is,
+Henceforth, we denote by $\rho({A})$ the spectral radius of a matrix A{A}, that is,
 $\rho({A})=\max\{|\lambda_{1}|,...,|\lambda_{d}|\},$
@@
 Theorem 4.1, sufficiency.
-Suppose that there exists a ${K}$ such that $\rho({A}+{B}{K})<1$. Then choose all controls to be $\mathbf{u}_{t}={K}\mathbf{x}_{t}$. The evolution of the system without perturbation is given by
+Suppose that there exists a K{K} such that $\rho({A}+{B}{K})<1$. Then choose all controls to be $\mathbf{u}_{t}={K}\mathbf{x}_{t}$. The evolution of the system without perturbation is given by
 $\mathbf{x}_{t}=({A}+{B}{K})\mathbf{x}_{t-1}=...=({A}+{B}{K})^{t}\mathbf{x}_{0}.$
@@
 For the other direction, given some matrix $A$ with $\rho(A)<1$, consider $P=\sum_{t=0}^{\infty}(A^{t})^{\top}A^{t}$.
-Since $P=I+A^{\top}PA$, $P$ satisfies the required conclusions.
+Since $I+A^{\top}PA, PP$ satisfies the required conclusions.
 The fact that $P$ as defined exists, that is, that it is finite, as long as $\rho(A)<1$ can be verified via the Jordan normal form of $A$, which in particular asserts that $\|A^{t}\|\leq O(t^{d_{x}}\rho(A)^{t})$.
@@
 Theorem 4.4.
-There exists a matrix ${K}$ such that $\rho({A}+{B}{K})<1$ if and only if there exist matrices ${P}\in\mathcal{S}^{d_{x}\times d_{x}}$ and ${L}\in{\mathbb{R}}^{d_{u}\times d_{x}}$ such that
+There exists a matrix K{K} such that $\rho({A}+{B}{K})<1$ if and only if there exist matrices ${P}\in\mathcal{S}^{d_{x}\times d_{x}}$ and ${L}\in{\mathbb{R}}^{d_{u}\times d_{x}}$ such that
 $\begin{bmatrix}{P}&({A}{P}+{B}{L})^{\top}\\ {A}{P}+{B}{L}&{P}\end{bmatrix}\succ 0,\qquad{P}\succeq{I}.$ (4.1)
-Furthermore, given such matrices ${P}$ and ${L}$ satisfying Equation 4.1, the matrix ${K}^{\prime}={L}{P}^{-1}$ satisfies $\rho({A}+{B}{K}^{\prime})<1$.
+Furthermore, given such matrices P{P} and L{L} satisfying Equation 4.1, the matrix ${K}^{\prime}={L}{P}^{-1}$ satisfies $\rho({A}+{B}{K}^{\prime})<1$.
 Proof.
@@
 Lemma 4.5.
-Define a block matrix ${M}$ as
+Define a block matrix M{M} as
 ${M}=\begin{bmatrix}{A}&{B}\\ {B}^{\top}&{C}\end{bmatrix}$. (4.2)
@@
 Note that since the spectral radius purely involves the eigenvalues of a matrix, it is invariant under transposition, and we have $\rho((A+BK)^{\top})=\rho(A+BK)$.
-Now using this fact and Theorem 4.3, we know that the stabilizability of a linear system is equivalent to the existence of matrices ${K}$ and ${P}$ such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)P(A+BK)^{\top}}$.
+Now using this fact and Theorem 4.3, we know that the stabilizability of a linear system is equivalent to the existence of matrices K{K} and P{P} such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)P(A+BK)^{\top}}$.
 The above lemma allows us to reformulate this cubic constraint to a quadratic constraint.
-The stabilizability of a system is equivalent to existence of matrices ${K}$ and ${P}$ such that
+The stabilizability of a system is equivalent to existence of matrices K{K} and P{P} such that
 $\begin{bmatrix}{P}&{((A+BK)P)^{\top}}\\ {(A+BK)P}&{P}\end{bmatrix}\succ{0},\qquad{P}\succeq{I}.$
-Lastly, we reparameterize this constraint as a linear semidefinite constraint as demonstrated below, by choosing ${L}={KP}$.
+Lastly, we reparameterize this constraint as a linear semidefinite constraint as demonstrated below, by choosing L=KP{L}={KP}.
 $\begin{bmatrix}{P}&{(AP+BL)^{\top}}\\ {AP+BL}&{P}\end{bmatrix}\succ{0},\qquad{P}\succeq{I}.$
@@
 $\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t}B}]=\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t-1}B}].$
-Since linear transformations retain containment, we can multiply by ${A}$ and get
+Since linear transformations retain containment, we can multiply by A{A} and get
 $\mathop{\mbox{\rm span}}[{AB,\ldots,A^{t+1}B}]\subseteq\mathop{\mbox{\rm span}}[{AB,\ldots,A^{t}B}].$
-Adding ${B}$ to both matrices does not affect the subset relationship.
+Adding B{B} to both matrices does not affect the subset relationship.
 $\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t+1}B}]\subseteq\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t}B}]$
@@
 4.4.1 Stabilizability of Time-Invariant Systems
-Consider a linear time-invariant system given by the transformations $({A,B})$.
+Consider a linear time-invariant system given by the transformations (A,B)({A,B}).
 We have already remarked that this system is called stabilizable, according to the classical definition, if and only if there exists $K$ such that $\rho({A+BK})<1$ holds.
@@
 Lemma 4.8.
-Consider a linear controller ${K}$ for the LTI system $({A,B})$.
-If there exist $\kappa>0$ and $\delta\in(0,1]$ such that the matrix ${A+BK}$ admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
+Consider a linear controller K{K} for the LTI system (A,B)({A,B}).
+If there exist $\kappa>0$ and $\delta\in(0,1]$ such that the matrix A+BK{A+BK} admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
 $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t},\qquad\frac{\kappa}{\delta}\leq\gamma$
@@
 $\min_{\mathbf{u}(\mathbf{x})}\lim_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}{R}\mathbf{u}_{t}\right]$
-[…]
+$\text{ s.t. }\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t},$
 where similar to the finite horizon case, we assume that the perturbations $\mathbf{w}_{t}$ are i.i.d.
@@
 5.3.1 Solving the DARE
-In the one-dimensional case, although the right side of the equation is a rational function, we can multiply both sides by ${R+B^{\top}SB}$ to form a quadratic equation in ${S}$.
+In the one-dimensional case, although the right side of the equation is a rational function, we can multiply both sides by $R+B^{\top}SB$} to form a quadratic equation in S{S}.
 Solving the DARE therefore corresponds to finding a nonnegative root of a quadratic polynomial in the one-dimensional case.
 The reason we only consider the nonnegative root is that the value function is nonnegative by our choice of quadratic convex costs.
-In higher dimension, the correct generalization is somewhat more difficult to see, but using Lemma 4.5, the equation can be reformulated as the solution to the following SDP.
+In higher dimension, the correct generalization is somewhat more difficult to see, but using Lemma 4.5, the equation can be reformulated as the solution to the following semidefinite program.
 $\min_{{X}\in\mathcal{S}^{d_{x}\times d_{x}}}\langle{X,I}\rangle$
 $\text{subject to}\begin{bmatrix}{X}&B^{T}SA\\ A^{T}SB&{R}+B^{T}SB\end{bmatrix}\succeq 0,$
-and $\text{and }S=Q+A^{T}SA-{X}.$
+$\text{and }S=Q+A^{T}SA-{X}.$
 A popular alternative is to adapt the value iteration algorithm to this setting by iteratively applying the right side of the aforementioned equation as an operator.
@@
 The text [BB08] presents a succinct derivation of robust controllers entirely in the state space, as the solution to a dynamic game between the controller and an adversary; see also [Tu18] for a recent exposition.
-Although developed in the linear setting, adaptations of the LQR controller are also widely used to control non-linear systems [Li05, LK13].
+Although developed in the linear setting, adaptations of the LQR controller are also widely used to control nonlinear systems [Li05, LK13].
 5.6 Exercises
@@
 Hence,
-$\|\mathbf{y}_{t+1}-\mathbf{x}^{\star}\|^{2}$$\leq$ $\ \|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}+\eta_{t}^{2}\|\nabla_{t}\|^{2}-2\eta_{t}\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$
-$2\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$$\leq$ $\ \frac{\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$ (6.4)
+$\|\mathbf{y}_{t+1}-\mathbf{x}^{\star}\|^{2}\ \displaystyle\leq$ $\ \|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}+\eta_{t}^{2}\|\nabla_{t}\|^{2}-2\eta_{t}\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$
+$2\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})\ \displaystyle\leq$ $\ \frac{\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$ (6.4)
 Summing (6.2) and (6.1.1) from $t=1$ to $T$, and setting $\eta_{t}=\frac{D}{G\sqrt{t}}$ (with $\frac{1}{\eta_{0}}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}0$):
@@
 $=\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi_{1}\in\Pi_{1}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{1}},\mathbf{u}_{t}^{\pi_{1}})\right)+\varepsilon T$
-Since this holds for any fixed disturbance sequence $\mathbf{w}_{1:T}$, it also holds for the maximum over all disturbance sequences $\mathbf{w}_{1:T}$ with $||\mathbf{w}_{t}||\leq 1$. ∎
+Since this holds for any fixed disturbance sequence $\mathbf{w}_{1:T}$, it also holds for the maximum over all disturbance sequences $\mathbf{w}_{1:T}$ with $\mathbf{w}_{t}||\leq 1$. ∎
 We now proceed to define and study approximation relationships between different policy classes.
@@
 $=A\mathbf{x}_{t}+B(K\mathbf{x}_{t}+\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i})+\mathbf{w}_{t}$
-$=(A+BK)\mathbf{x}_{t}+B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$ using $\text{using }M_{0}=I$
+$=(A+BK)\mathbf{x}_{t}+B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$ $\text{using }M_{0}=I$
 $=\sum_{j=0}^{t}\left((A+BK)^{j}B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i-j}\right)$
@@
 when executed on some dynamical system $(A+BK_{0},B)$, then when the corresponding policies $\pi^{\prime}_{1}$ and $\pi^{\prime}_{2}$, that is, $\pi_{1}$ and $\pi_{2}$ with $K_{0}$ superimposed, are executed on the dynamical system $(A,B)$ the iterates guarantee
-[…]
+$\max\{$ $\|\mathbf{x}_{t}^{\pi^{\prime}_{1}}(A,B)-\mathbf{x}_{t}^{\pi^{\prime}_{2}}(A,B)\|,$
 $\|\mathbf{u}_{t}^{\pi^{\prime}_{1}}(A,B)-\mathbf{u}_{t}^{\pi^{\prime}_{2}}(A,B)\|\}\leq(1+\|K_{0}\|)\varepsilon.$
@@
 $\leq\sum_{i=h+1}^{t}\kappa(1-\delta)^{i+1}\|K\|\|\mathbf{w}_{t-i}\|$
-[…]
+$\leq\kappa^{2}\sum_{i=h+1}^{\infty}(1-\delta)^{i}\leq\kappa^{2}\int_{i=h}^{\infty}e^{-\delta i}di$
 $=\kappa^{2}\frac{1}{\delta}e^{-\delta h}\leq\varepsilon.$
@@
 The reader is referred to the comprehensive texts of [CBL06, Haz16] for a detailed treatment.
-Of particular
 Linear state-feedback policies are a mainstay of control theory in both applications, due to their simplicity, and in theory, due to their optimality in deterministic and stochastic settings, as demonstrated in [Kal60].
@@
 6.4. Prove Lemma 6.8.
-Chapter 7 Online Nonstochatic Control
+Chapter 7 Online Nonstochastic Control
 In this chapter, we explore recent advances in machine learning that have significantly reshaped the control problem.
@@
 3:for $t$ = $1\ldots T$ do
-4:     Use control $\mbox{Use control }\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}{M}_{i}^{t}\mathbf{w}_{t-i}.$
+4:  $\mbox{Use control }\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}{M}_{i}^{t}\mathbf{w}_{t-i}.$
 5:  Observe state $\mathbf{x}_{t+1}$; compute noise $\mathbf{w}_{t}=\mathbf{x}_{t+1}-{A}_{t}\mathbf{x}_{t}-{B}_{t}\mathbf{u}_{t}$.
@@
  Set $\widehat{A^{j}B}=\frac{1}{T_{0}}\sum_{t=0}^{T_{0}-1}\mathbf{x}_{t(k+1)+j+1}\eta_{t(k+1)}^{\top}$, for all $j\in\{0,\dots k\}$.
-  Define $\widehat{C_{0}}=(\widehat{B},\dots\widehat{A^{k-1}B})$, $\widehat{C_{1}}=(\widehat{AB},\dots\widehat{A^{k}B})$, and return $\widehat{A},\widehat{B}$ as
+[…]
 $\widehat{B}=\widehat{B},\quad\widehat{A}=\operatorname*{\arg\min}_{A}\|\widehat{C_{1}}-A\widehat{C_{0}}\|_{F}.$
@@
     the perturbations $\mathbf{w}_{t}$ do not depend on $\mathbf{u}_{t}$,
-and for $T_{0}=\Omega\left(\frac{d^{2}k^{2}\kappa^{2}\gamma^{6}}{\varepsilon^{2}}\log\frac{1}{\Delta}\right)$, the following holds.
-Algorithm 5 outputs a pair $(\widehat{A},\widehat{B})$ that satisfies, with probability $1-\Delta$,
+[…]
 $\|\widehat{A}-A\|\leq\varepsilon,\|\widehat{B}-B\|\leq\varepsilon.$
@@
 8.4 From Indentification to Nonstochastic Control
-Having recovered approximate estimates $(\widehat{A},\widehat{B})$ of the system matrices, a natural question arises: how well does the Gradient Perturbation Algorithm 4 perform on the original system $(A,B)$ when provided with only approximate estimates of the same?
-We briefly state a result to this extent, whose proof we leave to the explorations of an interested reader.
+[…]
 Lemma 8.5.
-Given estimates of the system matrices $(\widehat{A},\widehat{B})$ such that
+[…]
 $\|\widehat{A}-A\|,\|\widehat{B}-B\|\leq\varepsilon,$
@@
 An intuitive class of prediction rules that makes use of the special structure of linear dynamical systems is that of open-loop linear dynamical predictors, which we now define.
-The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $\mathbf{C}=I$ is the identity matrix.
+The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $\mathbf{C}=I$, the identity matrix.
 In this situation, the optimal prediction can be calculated using the system evolution equation $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}$.
@@
 The work of [GLS+20] extends identification-based techniques to adversarial noise and marginally stable systems.
-The class $\prod^{*}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [AHL+18].
-In that paper, it was shown how to learn partially observable and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
+The class $\prod^{*}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [HSZ17].
+That paper showed how to learn partially observable and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
 Under probabilistic assumptions, the work of [HMR18] shows that the matrices $A,B,C,D$ of an unknown and partially observed time-invariant linear dynamical system can be learned using first-order optimization methods.
@@
 Consider the double integrator dynamical system with $\Delta=1$ and partial observability where the observation matrix is given as follows
-[…]
+$A=\begin{bmatrix}1\ 1\\ 0\ 1\end{bmatrix}\ ,\ C=[0\ 1].$
 This system allows only the velocity of the object to be observed, while the position remains hidden.
@@
 $\hat{K}(A,C)=\begin{bmatrix}C\\ CA\\ \dots\\ CA^{d_{\mathbf{x}}-1}\end{bmatrix}$.
-[…]
+where $A$ and $C$ are the system matrices that define the state transition and observation model, respectively.
+Notice that in our example, the Kalman observability matrix is given by $\hat{K}(A,C)=\begin{bmatrix}0\ 1\\ 0\ 1\end{bmatrix}$, which is degenerate.
+This is not by chance; the system is observable if and only if the observability matrix has full rank.
+We ask the reader to provide a proof of the following result in Exercise 10.6.
 Theorem 10.2.
@@
 Finally, we give a recursive relation on $\Sigma_{t}$ as follows–
-$\Sigma_{t+1}$ =$\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})^{\top}\right]$
+$\Sigma_{t+1}$ $\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})^{\top}\right]$
 $=(A-L_{t}C)\Sigma_{t}(A-L_{t}C)^{\top}+\Sigma_{x}+L_{t}\Sigma_{y}L_{t}^{\top}$
@@
 Lemma 10.6.
-Consider any partially-observed linear dynamical system $(A,B,C)$ where the Kalman filtering matrix $L$ is such that the system $A-LC$ is $\gamma$-stable.
+Consider any partially-observed linear dynamical system $(A,B,C)$ where the Kalman filtering matrix $L$ is such that the system ALCA-LC is $\gamma$-stable.
 The Kalman filtering policy is $\varepsilon$-contained in the class of linear policies $\Pi^{L}_{h,k,\kappa}$, where $h=k=\gamma\log\frac{\gamma T}{\varepsilon}$ and $\kappa=\gamma^{3}$.
@@
 From the Gaussian nature of perturbations, we observe that
-$\begin{bmatrix}\mathbf{x}_{t}|(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\\ \mathbf{y}_{t}|(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\hat{\mathbf{x}}_{t}\\ C\hat{\mathbf{x}}_{t}\end{bmatrix},\begin{bmatrix}\Sigma_{t}&\Sigma_{t}C^{\top}\\ C\Sigma_{t}&C\Sigma_{t}C^{\top}+\Sigma_{y}\end{bmatrix}\right).$
+[…]
 for some $\Sigma_{t}\in{\mathbb{R}}^{d_{\mathbf{x}}\times d_{\mathbf{x}}}$, for which we provide a recursive formula towards the end of the proof.
@@
 If a random variable pair $(\mathbf{x},\mathbf{y})$ are distributed as
-$\begin{bmatrix}\mathbf{x}\\ \mathbf{y}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mu_{x}\\ \mu_{y}\end{bmatrix},\begin{bmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}\end{bmatrix}\right),$
+[…]
 then $\mathbf{x}|\mathbf{y}\sim\mathcal{N}(\mu_{x}+\Sigma_{12}\Sigma_{22}^{-1}(\mathbf{y}-\mu_{y}),\Sigma_{11}-\Sigma_{12}\Sigma_{22}^{-1}\Sigma_{21}).$
@@
 However, challenges remain, particularly in scenarios with non-Gaussian noise, time-varying dynamics, or adversarial disturbances.
-These limitations motivate extensions such as robust filtering, non-linear filtering (e.g., the extended and unscented Kalman filters) and learning-based adaptations that integrate statistical inference with modern machine learning techniques.
+These limitations motivate extensions such as robust filtering, nonlinear filtering (e.g., the extended and unscented Kalman filters) and learning-based adaptations that integrate statistical inference with modern machine learning techniques.
 The next chapter will transition from the classical filtering framework to spectral filtering, providing an alternative approach to state estimation that leverages spectral methods for efficient learning in dynamical systems.
@@
 The Extended Kalman Filter (see, for example, [UA21]) addresses this issue by linearizing the dynamics at each time step via a first-order Taylor series expansion.
-However, linearization can introduce significant errors in highly non-linear systems.
+However, linearization can introduce significant errors in highly nonlinear systems.
 Notable extensions using more sophisticated sampling include the Unscented Kalman Filter [JU04], leading to better estimation accuracy in strongly nonlinear systems.
   • Particle Filtering (Sequential Monte Carlo Methods)
-    For highly non-linear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative.
+    For highly nonlinear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative.
 These methods approximate the posterior distribution using a set of weighted samples (particles) and employ importance sampling and resampling techniques.
@@
 10.5. Prove Lemma 10.8, which is stated below. If a random variable pair $(\mathbf{x},\mathbf{y})$ are distributed as
-$\begin{bmatrix}\mathbf{x}\\ \mathbf{y}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mu_{x}\\ \mu_{y}\end{bmatrix},\begin{bmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}\end{bmatrix}\right),$
+[…]
 then the conditional distribution of $\mathbf{x}$ given $\mathbf{y}$ can be written as $\mathbf{x}|\mathbf{y}\sim\mathcal{N}(\mu_{x}+\Sigma_{12}\Sigma_{22}^{-1}(\mathbf{y}-\mu_{y}),\Sigma_{11}-\Sigma_{12}\Sigma_{22}^{-1}\Sigma_{21}).$
@@
 To define the basis, we consider the following matrix:
-[…]
+$Z_{T}=\int_{0}^{1}\mu_{\alpha}\mu_{\alpha}^{\top}d\alpha\in{\mathbb{R}}^{T\times T}.$
 The matrix $Z_{T}$ is symmetric and is independent of $\alpha$.
@@
 An example of a Hankel matrix is the following,
-[…]
+$W_{T}$ $=\int_{0}^{1}\begin{bmatrix}1&\alpha&\alpha^{2}&&\\ \alpha&\alpha^{2}&\alpha^{3}&&\\ \alpha^{2}&\alpha^{3}&\alpha^{4}&&\\ &&&...&\\ &&&&\alpha^{2T-2}\end{bmatrix}d\alpha=\begin{bmatrix}1&\frac{1}{2}&\frac{1}{3}&&\\ \frac{1}{2}&\frac{1}{3}&\frac{1}{4}&&\\ \frac{1}{3}&\frac{1}{4}&\frac{1}{5}&&\\ &&&...&\\ &&&&\frac{1}{2T-1}\end{bmatrix}$.
 This matrix is closely related to the $Z_{T}$ matrix we have introduced before, but not equal to it.
@@
 This theorem shows that the eigenvalues of a Hankel matrix decay very rapidly.
-An additional illustration of this fact is given in Figure 11.1, which plots the eigenvalues of the matrix $Z_{T}$ for $T=30$ on a logarithmic scale.
-Figure 11.1: Eigenvalues of Hankel matrices decrease geometrically. These are the eigenvalues of the matrix $Z_{T}$, where $T=30$, plotted on a logarithmic scale.
+An additional illustration of this fact is given in Figure 11.1, which plots the eigenvalues of the matrix $Z_{T}$ for $T=100$ on a logarithmic scale.
+Figure 11.1: Eigenvalues of Hankel matrices decrease geometrically. These are the eigenvalues of the matrix $Z_{T}$, where $T=100$, plotted on a logarithmic scale.
 11.2 Spectral Predictors
@@
 Define the class of spectral linear predictors of order $k$ recursively as
-$\prod_{h}^{\lambda}=\left\{\pi_{M}\ \ \left|\ \ \pi_{M}(\mathbf{u}_{1:t},\mathbf{y}_{1:t})-\pi_{M}(\mathbf{u}_{1:t-1},\mathbf{y}_{1:t-1})=M_{0}\mathbf{u}_{t}+\sum_{i=1}^{h}M_{i}\cdot\phi_{i}(\tilde{\mathbf{u}}_{t-1})\right.\right\}.$
+[…]
 Here the operator $\phi_{i}(\mathbf{u}_{t-1})$ takes the inner product of the $i$’th eigenvector of $Z_{T}$ with $\tilde{\mathbf{u}}_{t-1}$, where $\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}$.
@@
 $\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\beta\mathbf{u}_{t}$ $=\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1})$
-[…]
-$=\beta\cdot\mu_{\alpha}^{\top}(\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top})\tilde{\mathbf{u}}_{t-1}$ basis property $\mbox{basis property }\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top}=I$
+$=\beta\cdot\mu_{\alpha}^{\top}\tilde{\mathbf{u}}_{t-1}$ (11.2)
+$=\beta\cdot\mu_{\alpha}^{\top}(\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top})\tilde{\mathbf{u}}_{t-1}$ $\mbox{basis property }\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top}=I$
 $=\sum_{i=1}^{T}M_{i}^{\star}\cdot\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}.$ $M_{i}^{\star}=\beta\phi_{i}^{\top}\mu_{\alpha}$
@@
 $\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|-\|\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{2}}\|$
-$\leq\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}+\mathbf{y}_{t-1}^{\pi_{2}}\|$$\triangle\mbox{-inequality}$
+$\leq\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}+\mathbf{y}_{t-1}^{\pi_{2}}\|$ -$\triangle\mbox{-inequality}$
 $=\left\|\sum_{i=1}^{T}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}-\sum_{i=1}^{h}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}\right\|$
@@
 $\leq\sqrt{T}\cdot\sum_{i=h+1}^{T}|M_{i}^{\star}|$ $|\mathbf{u}_{t}|\leq 1,\|\phi_{i}\|_{2}\leq 1$
-[…]
+$\leq\sqrt{T}\times O\left(\int_{h}^{\infty}e^{-\frac{i}{\log T}}di\right)$ Lemma 11.4
 $\leq\frac{\varepsilon}{T}.$ choice of $h$
@@
 First, notice that by definition of $\phi_{j}$ as the eigenvectors of $Z_{T}$, and $\sigma_{j}$ as the corresponding eigenvalues, we have
-[…]
+$\int_{\alpha=0}^{1}g(\alpha)d\alpha=\int_{\alpha=0}^{1}\big(\phi_{j}^{\top}\mu_{\alpha}\big)^{2}d\alpha=\int_{\alpha=0}^{1}\phi_{j}^{\intercal}\mu_{\alpha}\mu_{\alpha}^{\intercal}\phi_{j}d\alpha=\phi_{j}^{\intercal}Z_{T}\phi_{j}=\sigma_{j}.$
 On the other hand, we claim that the function $g(\alpha)$ is L-Lipschitz for a constant value of $L$. To see this, first notice that
@@
 Thus, the maximum value of $g(\alpha)$ is bounded by, using Theorem 11.1,
-[…]
-[…]
+$g_{\max}=\max_{\alpha}\left\{(\phi_{j}^{\top}\mu_{\alpha})^{2}\right\}$ $\leq 2\cdot 6\cdot\int_{\alpha}(\phi_{j}^{\top}\mu_{\alpha})^{2}d\alpha$
+$=12\cdot\phi_{j}^{\top}\left(\int_{\alpha}\mu_{\alpha}\mu_{\alpha}^{\top}d\alpha\right)\phi_{j}$
 $=12\sigma_{j}$
@@
 11.6 Exercises
-11.1. Generalize Lemma 11.3 to all real numbers in the range $[-1,1]$, instead of $[0,1]$.
+11.1. Generalize Lemma 11.3 to all real numbers in the range [1,1][-1,1], instead of $[0,1]$.
 11.2. Generalize Lemma 11.3 to any dimension $d_{x}>1$.
@@
 By $\Pi^{G}_{h,\gamma}$, we denote the class of generalized linear observation-action controllers that satisfy $\sum_{i=0}^{h}\|K_{i}\|\leq\gamma$ and are also $\gamma$-stabilizing.
-Generalized observation-action policies can be seen as a special case of observation-action controllers using the technique of dimension lifting as follows.
-Consider any partially-observed system $(A,B,C)$ and perturbation sequence $\mathbf{w}_{t}$.
-Consider the lifted system $(\tilde{A},\tilde{B},\tilde{C},\tilde{D})$ given by
+[…]
 $\mathbf{z}_{t+1}=\tilde{A}\mathbf{z}_{t}+\tilde{B}\mathbf{u}_{t}+\tilde{D}\mathbf{w}_{t},\quad\tilde{\mathbf{y}}_{t}=C\mathbf{z}_{t},$
@@
 Thus, the new system has the same nature’s $y$’s as the original system.
-For any partially-observed system $(A,B,C)$ and perturbation sequence $\mathbf{w}_{t}$ producing observations $\mathbf{y}_{t}$, construct a lifted system $(\tilde{A},\tilde{B},\tilde{C})$ given by
+[…]
 $\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t+1}\\ \Delta_{t+1}\end{bmatrix}=\begin{bmatrix}A&0\\ 0&A\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta_{t}\end{bmatrix}+\begin{bmatrix}0\\ B\end{bmatrix}\mathbf{u}_{t}+\begin{bmatrix}I\\ 0\end{bmatrix}\mathbf{w}_{t},\quad\tilde{\mathbf{y}}_{t}=\begin{bmatrix}C&C\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta_{t}\end{bmatrix}$.
@@
 Let $\mathbf{x}_{t},\mathbf{y}_{t}$ be the states and observations produced by the original partially-observed system.
-The equivalents for the augmented system are $[{\mathbf{x}^{\mathbf{nat}}_{t}}^{\top},\Delta_{t}^{\top}]^{\top},\tilde{\mathbf{y}}_{t}$, as we have described above.
+The equivalents for the augmented system are ${\mathbf{x}^{\mathbf{nat}}_{t}}^{\top},\Delta_{t}^{\top}]^{\top},\tilde{\mathbf{y}}_{t}$, as we have described above.
 Now we can verify for all time $t$ that $\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}$ and $\mathbf{x}_{t}=\mathbf{x}^{\mathbf{nat}}_{t}+\Delta_{t}$.
@@
 2:for $t$ = $1\ldots T$ do
-3:     Use Control $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=0}^{h}{M}_{i}^{t}\mathbf{y}^{\mathbf{nat}}_{t-i}$
+3:  $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=0}^{h}{M}_{i}^{t}\mathbf{y}^{\mathbf{nat}}_{t-i}$
 4:  Observe $\mathbf{y}_{t+1}$, compute for $t+1$:
@@
 2:for $t$ = $1\ldots T$ do
-3:     Use Control $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=0}^{h}{M}_{i}^{t}\mathbf{y}^{\mathbf{nat}}_{t-i}$
+3:  $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=0}^{h}{M}_{i}^{t}\mathbf{y}^{\mathbf{nat}}_{t-i}$
 4:  Observe $\mathbf{y}_{t+1}$, compute for $t+1$:

### v8

@@
-version 1.0
+version 1.2 $\beta$
 [Uncaptioned image] Introduction to Online Control
-version 1.0
+version 1.2 $\beta$
 Elad Hazan    Karan Singh
@@
 Given a dynamical system, the regret of an online control algorithm $\mathcal{A}$ over $T$ iterations with respect to a class of policies $\Pi$ is given by:
-$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}$ $\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}))\right),$
+$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}$ $\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right),$
 where $\mathbf{u}_{t}={\mathcal{A}}(\mathbf{x}_{t})$ are the controls generated by $\mathcal{A}$}, and $\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}$ are the counterfactual state sequence and controls under the policy $\pi$, that is,
@@
 The first mathematical analysis of such systems appeared in a paper by James Clerk Maxwell on the topic of governors [MAX68], shortly after Maxwell had published his treatise on the unification of electricity and magnetism.
-For an enlightening historical account, see [FCZI03].
+For an enlightening historical account, see [FZ03].
 For a historical account on the development of the PID controller, see [BEN93]. The theoretical and practical properties of this controller are described in detail in [ÅH95].
@@
 The text [SL91] gives a thorough introduction to adaptive control theory.
-Online nonstochastic control theory and the Gradient Perturbation Control algorithm were proposed in [ABH+19].
+Online nonstochastic control theory and the Gradient Perturbation Control algorithm were proposed in [ABH+19b].
 A flurry of novel methods and regret bounds for online nonstochastic control have appeared in recent years, many of which are surveyed in [HS21].
-Notable among them are logarithmic regret algorithms [AHS19, FS20, Sim20], nonstochastic control for unknown systems [HKS20], black-box online nonstochastic control [CH21], and nonstochastic control with partial observability [SSH20].
+Notable among them are logarithmic regret algorithms [AHS19c, FS20, SIM20], nonstochastic control for unknown systems [HKS20], black-box online nonstochastic control [CH21], and nonstochastic control with partial observability [SSH20].
 The theory is heavily based on recent developments in decision making and online learning, notably the framework of online convex optimization [HAZ16], and especially its extensions to learning with memory [AHM13].
 Chapter 2 Dynamical Systems
-In this chapter, we return to the general problem of control from Definition 1.2, that is, choosing a sequence of controls to minimize a sequence of cost functions.
-Before moving on to algorithms for control, it is important to reason about what kind of solution is even possible.
-For this reason, we consider several natural examples of dynamical systems, and reason about what constitutes desirable behavior.
-This will help us understand the kinds of observation models and control objectives that are feasible and natural in real applications.
+In this chapter, we return to the general problem of control from Definition 1.2: choosing a sequence of controls to minimize a sequence of cost functions.
+Before moving on to algorithms for control, it is useful to ask what kinds of goals are natural and what kinds of guarantees are even possible for a dynamical system.
+To that end, we consider several examples of dynamical systems and introduce a few basic solution concepts, including equilibrium, stability, stabilizability, and controllability.
+These examples also illustrate the kinds of observation models and control objectives that arise in real applications.
 2.1 Examples of Dynamical Systems
@@
 In the task of mechanical ventilation, the control algorithm dictates the amount of oxygen inflow $u_{t}$ (in $\mathrm{m^{3}/s}$) into the lungs of a patient.
-This inflow approximately determines the volume of the lungs $v_{t}$ via a simple linear relation,
+This inflow approximately determines the volume of the lungs $v_{t}$ via the simple relation
 $v_{t+1}=v_{t}+\Delta u_{t},$
@@
 Instead, the controller observes the air pressure $p_{t}$.
-We give a simplified physical model that relates the pressure and the volume.
-This model, given in the following, can be derived by first principles from stress-strain relations, treating the left and right lungs connected by the trachea as situated in a two-balloon experiment:
-$p_{t}=C_{0}+C_{1}v_{t}^{-1/3}+C_{2}v_{t}^{5/3}$
-The objective of the control algorithm is to follow an ideal pressure trajectory $p^{*}_{t}$ prescribed by the physician as closely as possible, say, as measured by squared loss $c_{t}(p_{t},u_{t})=|p_{t}-p^{*}_{t}|^{2}+|u_{t}|^{2}$, while limiting the maximum instantaneous airflow to minimize patient discomfort.
-See Figure 2.1 for a schematic representation of airflow when a lung is connected to a mechanical ventilator.
+A simplified physical model relating pressure and volume is
+$p_{t}=C_{0}+C_{1}v_{t}^{-1/3}+C_{2}v_{t}^{5/3}.$
+This relation can be derived from stress–strain considerations by modeling the left and right lungs, connected by the trachea, as in a two-balloon experiment.
+The objective of the control algorithm is to follow an ideal pressure trajectory $p_{t}^{\star}$ prescribed by the physician as closely as possible, for example via the stage cost
+$c_{t}(p_{t},u_{t})=|p_{t}-p_{t}^{\star}|^{2}+|u_{t}|^{2},$
+while penalizing large airflow rates so as to limit patient discomfort. See Figure 2.1 for a schematic representation of airflow when a lung is connected to a mechanical ventilator.
 Figure 2.1: A schematic of the respiratory circuit from [SZG+21].
@@
 The double integrator is a common example of a linear dynamical system.
-In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics on a one-dimensional trajectory (i.e.
-the real number line).
-The state $\mathbf{x}_{t}$ is two dimensional, with the first coordinate indicating the position of a point object and the second denoting its velocity.
-The control algorithm decides the force $u_{t}\in{\mathbb{R}}$ that is applied to the object at each time step.
+In its simplest form, it models a one-dimensional point object moving along the real line.
+The state $\mathbf{x}_{t}\in{\mathbb{R}}^{2}$ is two-dimensional, with the first coordinate indicating position and the second denoting velocity.
+The control algorithm chooses a scalar actuation $u_{t}\in{\mathbb{R}}$ that directly affects the velocity coordinate at each time step; after an appropriate normalization, $u_{t}$ may be interpreted as a force or acceleration input applied over one sampling interval.
 The dynamics equation can be written as
 $\mathbf{x}_{t+1}=\begin{bmatrix}1&\Delta\\ 0&1\end{bmatrix}\mathbf{x}_{t}+\begin{bmatrix}0\\ 1\end{bmatrix}u_{t}+\mathbf{w}_{t}.$
-The position of the object is linearly related to its velocity via the parameter $\Delta$, which measures the time resolution the controller operates in.
+The position is linearly related to the velocity through the sampling period $\Delta$.
 Optionally, the state is also affected by external perturbations $\mathbf{w}_{t}$ outside the influence of the controller.
-The goal here is to ensure that the point object arrives and then comes to rest at some coordinate $x^{*}\in\mathbb{R}$ as quickly as possible, while also not expending too much energy.
-The cost function is naturally expressed as
-$c(\mathbf{x}_{t},u_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}x^{*}\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
+The goal here is to ensure that the point object arrives and then comes to rest at some coordinate $x^{\star}\in\mathbb{R}$ as quickly as possible, while also not expending too much energy.
+A natural stage cost is
+$c(\mathbf{x}_{t},u_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}x^{\star}\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
 […]
-Figure 2.2: Double integrator illustration showing state coordinate $x_{1}(t)$ (position), state coordinate $x_{2}(t)$ (velocity), mass $m$, and control input $u(t)$.
+Figure 2.2: Double integrator illustration showing state coordinate $x_{1}(t)$ (position), state coordinate $x_{2}(t)$ (velocity), and control input $u(t)$.
 2.1.3 Pendulum Swing-Up
-The swing-up pendulum is one of the simplest examples of rotational kinematics.
-It corresponds to a point object of mass $m$ mounted at the end of a massless rod of length $l$.
-The other end of the rod is fixed to a pivot.
-Being vertically aligned, the system is subject to gravity with gravitational acceleration $g$.
-The controller is responsible for applying a torque $u_{t}\in\mathbb{R}$ at each time step.
-Given the vertical angle $\theta_{t}$ of the rod, the physics of this system may be summarized as
-$\theta_{t+1}=\theta_{t}+\Delta\dot{\theta}_{t},\qquad\dot{\theta}_{t+1}=\dot{\theta}_{t}+\Delta\frac{u_{t}-mg\ \sin\theta_{t}}{ml^{2}},$
-where $\Delta$ is the time resolution of the controller.
-Note that this dynamical equation is nonlinear in the state $\mathbf{x}_{t}=[\theta_{t},\dot{\theta_{t}}]^{\top}$ due to the presence of the term $\sin\theta$ in the evolution of $\dot{\theta_{t}}$.
-PivotmmAngle $\theta_{t}$$m\mathbf{g}$$\mathbf{u}_{t}$$l$
-Figure 2.3: Pendulum swing-up illustration with marked angle $\theta_{t}$, gravitational force mgmg, control input torque $u_{t}$, mass $m$, and rod length $l$.
-The objective here is to get the rod to rest in a vertical upright position at an angle of $\pi$, using as little torque as possible throughout the process.
-This may be expressed through the cost function
+The swing-up pendulum is one of the simplest examples of rotational motion.
+It consists of a point mass $m$ attached to the end of a massless rod of length $l$, with the other end fixed at a pivot.
+The system is subject to gravity with acceleration $g$, and the controller applies a torque $u_{t}\in\mathbb{R}$ at each time step.
+If $\theta_{t}$ denotes the angle of the rod measured from the downward vertical, then a discretization of the continuous-time dynamics gives
+$\theta_{t+1}$ $=\theta_{t}+\Delta\dot{\theta}_{t}\ \ ,\ \ \ \dot{\theta}_{t+1}=\dot{\theta}_{t}+\Delta\left(\frac{u_{t}}{ml^{2}}-\frac{g}{l}\sin\theta_{t}\right).$
+Thus the state is $\mathbf{x}_{t}=[\theta_{t},\dot{\theta}_{t}]^{\top}$, and the dynamics are nonlinear because of the term $\sin\theta_{t}$.
+PivotmmAngle $\theta_{t}$$m\mathbf{g}$$u_{t}$$l$
+Figure 2.3: Pendulum swing-up illustration with marked angle $\theta_{t}$, gravitational force mgmg, control torque $u_{t}$, mass $m$, and rod length $l$.
+The objective here is to bring the rod to rest in the upright position $\theta=\pi$, using as little torque as possible. A simple stage cost is
 $c(\mathbf{x}_{t},u_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}\pi\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
-2.1.4 Aircraft Dynamics
-A final example of a dynamical system we consider is the longitudinal (vertical tilt) motion of an aircraft that is regulated by automated controllers.
-We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system, from [Boy10].
-The two dimensional control input $\mathbf{u}_{t}$ corresponds to the elevation angle and the thrust.
-The disturbances correspond to the velocity of the wind along the body axis and perpendicular to it.
-The following dynamics equation has matrices whose numerical values were computed using the dynamics equation for an actual aircraft.
-These numerical values can be determined by the mass, length, and other physical properties of the aircraft:
-$\mathbf{x}_{t+1}=$ $\begin{bmatrix}-0.003&0.039&0&-0.322\\ -0.065&-0.319&7.74&0\\ 0.020&-0.101&-0.429&0\\ 0&0&1&0\end{bmatrix}\mathbf{x}_{t}+\begin{bmatrix}0.01&1\\ -0.18&-0.04\\ -1.16&0.598\\ 0&0\end{bmatrix}\mathbf{u}_{t}$
-$+\begin{bmatrix}-0.003&0.039&0&-0.322\\ -0.065&-0.319&7.74&0\end{bmatrix}\mathbf{w}_{t}$
-Figure 2.4: Longitudinal control in an aircraft. Graphic by User_A1, licensed under $C$-BY-SA 4.0; it is an SVG derivative of a public domain work by G J Vigurs.
-The quantities of interest for assessing the performance of the controller are the aircraft speed, climb rate — each of which is a linear transformation of the state — and applied thrust.
-The corresponding cost is
-$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\begin{bmatrix}1&0&0&0\\ 0&-1&0&7.74\end{bmatrix}\mathbf{x}_{t}-\mathbf{x}_{t}^{*}\right\|^{2}+\left\|\begin{bmatrix}0&1\end{bmatrix}\mathbf{u}_{t}\right\|^{2}.$
+When interpreting this objective globally, the angular error should be understood modulo $2\pi$.
+2.1.4 Autoregressive Language Models
+A modern example of a dynamical system is an autoregressive language model.
+Let $\mathcal{V}$ be a finite vocabulary, and let $y_{t}\in\mathcal{V}$ denote the token observed at time $t$.
+The model maintains a hidden state $h_{t}$ that summarizes the preceding context.
+For a recurrent neural network, one may take $h_{t}\in\mathbb{R}^{d}$ to be a fixed-dimensional latent vector.
+For a transformer, it is more natural to regard $h_{t}$ as the model’s internal cache or prefix representation associated with the sequence $y_{1:t}$; this state is typically very high-dimensional and is not directly observed.
+A simple state-space description is
+$h_{t+1}=f_{\theta}(h_{t},e(y_{t})),$
+where $e(y_{t})$ is the embedding of token $y_{t}$ and $f_{\theta}$ is a learned nonlinear update map. From the hidden state, the model produces a probability distribution over the next token,
+$p_{t+1}=g_{\theta}(h_{t+1})\in\Delta(\mathcal{V}),$
+where $\Delta(\mathcal{V})$ denotes the probability simplex over the vocabulary.
+The objective in next-token prediction is to assign high probability to the realized next token $y_{t+1}$. A natural stage cost is therefore the log loss
+$c_{t}(h_{t},y_{t+1})=-\log p_{t+1}(y_{t+1}).$
+Equivalently, one may view training as minimizing the sum of these stage costs over time.
+This example illustrates several features that are common in modern machine learning systems.
+The hidden state is extremely large and latent, the observations are discrete symbols rather than physical measurements, and the dynamics are highly nonlinear.
+At generation time, the model recursively predicts a distribution over the next token, chooses or samples a token from that distribution, appends it to the context, and repeats the process.
 2.1.5 Epidemiological Models
-The COVID-19 pandemic marked an increase in public interest in policy discussions on the effectiveness of certain interventions as a means of controlling infection spread.
-Epidemiological models predict the spread of an infectious disease via structural models and differential equations.
-We present a discrete-time analogue of the most basic of such models, the SIR model.
-Let $S_{t}$, $I_{t}$, and $R_{t}$ denote the population sizes of those susceptible, infected, and recovered, respectively.
-Let $u_{t}$ be the number of vaccines administered at time $t$.
-The dynamics of the basic SIR model are given by
-$S_{t+1}$ $=S_{t}-\beta S_{t}I_{t}-\alpha u_{t}$
-$I_{t+1}$ $=I_{t}+\beta S_{t}I_{t}-\gamma I_{t}$
-$R_{t+1}$ $=R_{t}+\gamma I_{t}.$
-Here, the coefficient $\beta$ measures the level of communicability of the infection, $\gamma$ measures the rate at which an infected person recovers, and $\alpha$ represents the efficacy of the vaccine in preventing infections.
-Figure 2.5: Trajectory of a continuous-time SIR model with parameters $\alpha=0,\beta=0.3,\gamma=0.1$.
+The COVID-19 pandemic renewed public interest in policy interventions for controlling the spread of infectious diseases.
+A basic compartmental model is the SIR model.
+Let $S_{t}$, $I_{t}$, and $R_{t}$ denote the fractions of the population that are susceptible, infected, and removed (recovered or immune), respectively, so that $S_{t}+I_{t}+R_{t}=1$.
+Let $u_{t}\geq 0$ denote the vaccination effort at time $t$, and let $\alpha\in[0,1]$ denote vaccine efficacy.
+A discrete-time controlled SIR model is
+$S_{t+1}$ $=S_{t}-\beta S_{t}I_{t}-\alpha u_{t},$
+$I_{t+1}$ $=I_{t}+\beta S_{t}I_{t}-\gamma I_{t},$
+$R_{t+1}$ $=R_{t}+\gamma I_{t}+\alpha u_{t}.$
+Here, $\beta$ measures transmissibility, $\gamma$ is the recovery rate, and $\alpha u_{t}$ is the fraction of susceptible individuals effectively moved into the removed/immune compartment by vaccination.
+Under these dynamics, the identity $S_{t}+I_{t}+R_{t}=1$ is preserved whenever it holds initially.
+A natural control objective is to keep infections small while also limiting intervention cost, for example through the stage cost
+$c_{t}(S_{t},I_{t},R_{t},u_{t})=I_{t}^{2}+\lambda u_{t}^{2},$
+for some tradeoff parameter $\lambda>0$, subject to $u_{t}\geq 0$ and $\alpha u_{t}\leq S_{t}$.
+Figure 2.4: Typical trajectory of a continuous-time SIR model without vaccination, with parameters $\beta=0.3$ and $\gamma=0.1$.
+These examples span hidden-state and fully observed systems, linear and nonlinear dynamics, and a variety of natural control objectives.
+We next abstract away from the application details and formalize several general solution concepts.
 2.2 Solution Concepts for Dynamical Systems
-Before considering the control, or manipulation, of a dynamical system, we must first reason about what the goal should even be.
-In this section, we discuss various solution concepts for a dynamical system, including equilibrium, stability, stabilizability, and controllability.
-Given a dynamical system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t})$, the equivalent deterministic dynamical system, defined through the dynamics $\mathbf{x}_{t+1}=f^{\prime}(\mathbf{x}_{t},\mathbf{u}_{t})=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{0})$, models the evolution of the dynamical system in the absence of perturbations.
-In addition, the autonomous dynamical system $\mathbf{x}_{t+1}=f^{\prime\prime}(\mathbf{x}_{t})=f(\mathbf{x}_{t},\mathbf{0},\mathbf{0})$ describes the state transitions of the dynamical system in the absence of both control inputs and perturbations.
+Before considering how to control a dynamical system, we should decide what success means.
+In this section, we discuss several solution concepts, including equilibrium, stability, stabilizability, and controllability.
+Given a dynamical system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t})$, its deterministic counterpart
+$\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{0})$
+models the evolution of the system in the absence of perturbations. The autonomous system
+$\mathbf{x}_{t+1}=f(\mathbf{x}_{t})=f(\mathbf{x}_{t},\mathbf{0},\mathbf{0})$
+describes the state transitions in the absence of both control inputs and perturbations. In both cases, we use the same symbol $f$ when the omitted arguments are understood to be zero.
 2.2.1 Equilibrium of Dynamical Systems
-An example of an equilibrium is a ball sitting at rest at the bottom of a pit.
-If the ball were to be slightly moved, it would return to the bottom of the pit.
-Thus, this equilibrium is stable and attractive.
-If the ball is at rest on the top of a mountain, pushing the ball in any direction would cause it to fall. This is an equilibrium that is unstable. A formal definition is as follows.
+An example of an equilibrium is a ball sitting at rest at the bottom of a pit. If the ball is moved slightly, it returns to the bottom of the pit. Thus this equilibrium is both stable and attractive.
+If a ball is at rest on the top of a mountain, then pushing it in any direction causes it to roll away. This is an equilibrium that is unstable. We now formalize these notions.
 Definition 2.1.
-A point $\mathbf{x}$ is an equilibrium point of an autonomous dynamical system $f$ if and only if $\mathbf{x}=f(\mathbf{x})$.
-An equilibrium point $\mathbf{x}$ is stable and attractive if and only if $\mathbf{x}=\lim_{t\rightarrow\infty}f^{t}(\mathbf{x}_{0})$ for every initial state $\mathbf{x}_{0}$ in some neighborhood of $\mathbf{x}$.
-If the neighborhood includes the entire set of states, we say that the equilibrium is globally attractive.
+A point $\mathbf{x}_{e}$ is an equilibrium point of an autonomous dynamical system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t})$ if and only if $f(\mathbf{x}_{e})=\mathbf{x}_{e}$.
+The equilibrium $\mathbf{x}_{e}$ is stable if for every $\varepsilon>0$ there exists $\delta>0$ such that
+$\|\mathbf{x}_{0}-\mathbf{x}_{e}\|<\delta\quad\Longrightarrow\quad\|\mathbf{x}_{t}-\mathbf{x}_{e}\|<\varepsilon\qquad\text{for all }t\geq 0.$
+The equilibrium $\mathbf{x}_{e}$ is attractive if there exists $\delta>0$ such that
+$\|\mathbf{x}_{0}-\mathbf{x}_{e}\|<\delta\quad\Longrightarrow\quad\lim_{t\to\infty}\mathbf{x}_{t}=\mathbf{x}_{e}.$
+The equilibrium $\mathbf{x}_{e}$ is asymptotically stable if it is both stable and attractive.
+If the attraction property holds for every initial state in the state space, we say that the equilibrium is globally attractive; if it is also stable, we say that it is globally asymptotically stable.
 Another example is given by a linear transformation in Euclidean space. Consider the system
 $\mathbf{x}_{t+1}=A\mathbf{x}_{t}.$
-Clearly $\mathbf{0}$ is an equilibrium point as ${A}\mathbf{0}=\mathbf{0}$ for all A{A}.
-If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vector converges to the null vector.
-Thus, $\mathbf{0}$ is a globally attractive and stable equilibrium.
-In general, determining whether a system has an equilibrium is an intractable problem, as we show in this chapter.
+Clearly $\mathbf{0}$ is an equilibrium point, since $A\mathbf{0}=\mathbf{0}$ for every matrix $A$.
+If every eigenvalue of $A$ has modulus strictly less than one, then $A^{t}\mathbf{x}_{0}\to\mathbf{0}$ for every initial state $\mathbf{x}_{0}$.
+Thus, $\mathbf{0}$ is a globally asymptotically stable equilibrium.
+In general, determining or even describing equilibrium sets and long-run behavior can be difficult, as the examples later in this chapter illustrate.
 2.2.2 Stabilizability and Controllability
-Two fundamental properties of dynamical systems are stabilizability and controllability, as defined next.
+Two fundamental properties of dynamical systems are stabilizability and controllability.
 Definition 2.2.
-A dynamical system is stabilizable if and only if, for every initial state, there exists a sequence of controls that drives the deterministic-equivalent system to the zero state.
-That is, $\lim_{t\rightarrow\infty}{\mathbf{x}_{t}}=\mathbf{0}$.
+A dynamical system is asymptotically stabilizable if there exists a control policy $\pi$ such that, for every initial state, the deterministic closed-loop trajectory satisfies
+$\lim_{t\to\infty}\mathbf{x}_{t}^{\pi}=\mathbf{0}.$
 Definition 2.3.
-A dynamical system is controllable if and only if, for every initial and target state, there exists a sequence of controls that drives the equivalent deterministic system to the target state.
-Note that controllability clearly implies stabilizability, as the zero state is one possible target state.
-In the next section, we discuss computational difficulties that arise in determining if a system is stabilizable and/or controllable.
+A dynamical system is controllable if, for every initial state $\mathbf{x}_{0}$ and every target state $\mathbf{x}^{\star}$, there exist a finite horizon $T$ and controls $\mathbf{u}_{0},\ldots,\mathbf{u}_{T-1}$ such that the deterministic counterpart of the system satisfies $\mathbf{x}_{T}=\mathbf{x}^{\star}$.
+Controllability implies finite-time reachability of the zero state.
+If the zero state is an equilibrium that can be maintained once reached, then controllability also implies asymptotic stabilizability.
+Some texts reserve the word stabilizable for a weaker algebraic property of linear systems. Here we use the more direct trajectory-based notion above, since it is convenient for our purposes.
+In the next subsection, we introduce a more quantitative robust-boundedness notion that is useful in nonstochastic control.
 2.2.3 Quantitative Definitions of Stabilizability
-The quantified notion of stabilizability we introduce below is more refined than driving the system to zero.
-We require the ability to drive the system to zero under adversarial bounded noise in the dynamics.
-In classical control theory this is called BIBO-stability, which stands for bounded-input-bounded-output.
-This definition can apply to any dynamical system and is an important concept in nonstochastic control.
+The notion below is stronger in one sense and weaker in another: it allows adversarial bounded disturbances, but it asks only that the state remain uniformly bounded rather than converge to zero.
+For linear systems, related ideas include robust invariance and input-to-state stability.
 Definition 2.4.
-A control policy $\pi$ for a dynamical system $f$ is $\gamma$-stabilizing if and only if the following holds.
-For every starting state $\|\mathbf{x}_{0}\|\leq 1$ and every sequence of adversarially chosen bounded perturbations $\mathbf{w}_{1:T}$ such that $\|\mathbf{w}_{t}\|\leq 1$ for all $t\in[T]$, let $\mathbf{x}_{t}^{\pi}$ be the state obtained by using policy $\pi$,
-$\mathbf{x}_{t+1}^{\pi}=f(\mathbf{x}_{t}^{\pi},\pi(\mathbf{x}_{t}),\mathbf{w}_{t}).$
-Then the norm of the system’s state at every iteration is bounded by an absolute constant independent of $T$, that is,
-$\forall t\ ,\ \|\mathbf{x}_{t}^{\pi}\|\leq\gamma.$
+A control policy $\pi$ for a dynamical system $f$ is $\gamma$-stabilizing if the following holds.
+For every initial state satisfying $\|\mathbf{x}_{0}^{\pi}\|\leq 1$ and every disturbance sequence $\mathbf{w}_{0},\ldots,\mathbf{w}_{T-1}$ such that $\|\mathbf{w}_{t}\|\leq 1$ for all $t\in\{0,\ldots,T-1\}$, let $\mathbf{x}_{t}^{\pi}$ be the trajectory generated by policy $\pi$:
+$\mathbf{x}_{t+1}^{\pi}=f(\mathbf{x}_{t}^{\pi},\pi(\mathbf{x}_{t}^{\pi}),\mathbf{w}_{t}).$
+Then
+$\|\mathbf{x}_{t}^{\pi}\|\leq\gamma\qquad\text{for all }t=0,1,\ldots,T.$
 A dynamical system is said to be $\gamma$-stabilizable if and only if it admits a $\gamma$-stabilizing control policy.
-2.3 Intractability of Equilibrium, Stabilizability, and Controllability
-Before formally proving the computational hardness of determining various solution concepts of dynamical systems, we start with a few intuitive and convincing examples of the complexities that arise.
-Consider the following dynamical system over the complex numbers:
+2.3 Complexity of Dynamical Systems and Hardness of Stabilizability
+Before formally proving the computational hardness of stabilizability, we begin with two examples that illustrate how complicated invariant sets and attractors can become even for simple dynamical rules.
+Consider the complex quadratic map
 $z_{t+1}=f_{c}(z_{t})=z_{t}^{2}+c.$
-The Julia set is now defined as
-$J_{c}=\{z\in C\ \colon\ \limsup_{t\to\infty}{|f_{c}^{t}(z)|}<\infty\},$
-where $f^{t}_{c}(x)=f_{c}(f_{c}(\dots f_{c}(x)))$ denotes $t$ successive applications of the function $f_{c}$.
-In other words, it is the set of all complex numbers $Z$ for which the system has a stable attractive equilibrium (that is not $\infty$).
-The boundary of the Julia set is very hard to determine, and it is known that determining if a ball around a given point belongs to the Julia set is a computationally undecidable problem, that is, there is no Turing machine that can decide it in finite time!
-(See e.g.
-[BY08].) The complexity that arises near the boundary of the Julia set is depicted in Figure 2.7.
-Another compelling visual illustration of the complexity of equilibrium in dynamical systems is the Lorenz attractor.
-This is a strange equilibrium resulting from simple dynamics that can be described using the following three equations over three variables, called the Lorenz equations [Lor63],
-$x_{t+1}$ $=\sigma(y_{t}-x_{t})$
-$y_{t+1}$ $=x_{t}(\rho-z_{t})-y_{t}$
-$z_{t+1}$ $=x_{t}y_{t}-\beta z_{t}.$
-The three dimensional (3D) trajectories of this system are attracted to a 2D manifold with two “wings” that resemble a butterfly, as depicted in Figure 2.7.
-For formal properties of this attractor set, see [Str14, Tuc99].
-This shows that even studying a solution concept itself is a formidable task.
-For a formal proof of attractiveness in the continuous case, see section 9.2 of [Str14] or the original paper of Lorenz [Lor63].
-Figure 2.6: Julia set.
+Its filled Julia set is
+$K_{c}=\left\{z\in\mathbb{C}\ \colon\ \sup_{t\geq 0}|f_{c}^{t}(z)|<\infty\right\},$
+where $f_{c}^{t}(z)$ denotes $t$ successive applications of the function $f_{c}$.
+The Julia set is the boundary $J_{c}=\partial K_{c}$.
+Depending on the parameter $c$, these sets can have extremely intricate geometry, and there are strong negative results on the computability of Julia sets for certain values of $c$; see, for example, [BY08].
+A related parameter-space object is the Mandelbrot set, shown in Figure 2.6.
+Another compelling visual illustration of complexity in dynamical systems is the Lorenz system, governed by the continuous-time equations
+$\dot{x}$ $=\sigma(y-x),$
+$\dot{y}$ $=x(\rho-z)-y,$
+$\dot{z}$ $=xy-\beta z.$
+For classical parameter values, trajectories are attracted to a complicated invariant set known as the Lorenz attractor, depicted in Figure 2.6.
+For background on this attractor, see [STR14, TUC99, LOR63].
+These examples show that even describing the long-run behavior of a seemingly simple dynamical system can be a formidable task.
+Figure 2.5: Mandelbrot set.
 Figure 2.6: Lorenz attractor. Visualization by D Schwen, licensed under $C$-BY-2.5.
 2.3.1 Computational Hardness of Stabilizability
-Consider the following problem: Given a dynamical system, can we determine if it stabilizable, that is, for every possible starting state, does there exist a choice of control inputs that drives the system to the null state?
+Consider the following decision problem: given a dynamical system, can we determine whether it is asymptotically stabilizable, that is, whether there exists a control policy that drives the system to the zero state from every initial condition?
 We now show that this decision problem is computationally hard.
 Theorem 2.5.
-For a family of dynamical systems described as polynomials with integer coefficients, determining the stabilizability of any member is NP-hard.
-We show that this problem can be reduced to the 3SAT problem, a well-known NP-hard problem. This implies that determining the stabilizability of a dynamical system is NP-hard in general.
+For the family of discrete-time polynomial dynamical systems with integer coefficients, deciding asymptotic stabilizability is NP-hard.
 Proof.
-We first claim that checking if a given polynomial has a zero is NP-hard.
-This is achieved by a reduction from 3SAT.
-Recall the 3SAT problem: Given a Boolean formula in the conjunctive normal form over $n$ variables and with $m$ clauses where each clause is restricted to three literals, such as
-$F(x_{1:n})=(x_{1}\lor\neg x_{2}\lor x_{3})\land(\neg x_{1}\lor\neg x_{2}\lor x_{3})\land(x_{1}\lor x_{3}\lor x_{4}),$
-decide if there exists a satisfying Boolean assignment.
-Given such a formula $F$, consider the following polynomial $p_{F}(\mathbf{u})$. For each clause $c$ in $F$, construct a degree-three polynomial as
-$g_{x_{i}\lor x_{j}\lor x_{k}}(\mathbf{u})$ $\rightarrow(1-u_{1})\cdot(1-u_{2})\cdot(1-u_{3})$
-$g_{\neg x_{i}\lor x_{j}\lor x_{k}}(\mathbf{u})$ $\rightarrow u_{1}\cdot(1-u_{2})\cdot(1-u_{3})$
-$g_{\neg x_{i}\lor\neg x_{j}\lor x_{k}}(\mathbf{u})$ $\rightarrow u_{1}\cdot u_{2}\cdot(1-u_{3})$
-$g_{\neg x_{i}\lor\neg x_{j}\lor\neg x_{k}}(\mathbf{u})$ $\rightarrow u_{1}\cdot u_{2}\cdot u_{3},$
-and the degree-six polynomial for the entire formula as
-$p_{F}(\mathbf{u})=\sum_{i=1}^{n}(u_{i}(1-u_{i}))^{2}+\sum_{j=1}^{m}(g_{c_{j}}(\mathbf{u}))^{2}.$
-For a disjunctive clause to be false, all its literals must evaluate to false. Therefore, given any satisfying Boolean assignment $x_{1:n}$ for the formula $F$, it holds that $p_{F}(x_{1:n})=0$.
-Similarly, given any $\mathbf{u}\in{\mathbb{R}}^{n}$ such that $p_{F}(\mathbf{u})=0$, it must be that $\mathbf{u}_{i}\in\{0,1\}$ for all $i$.
-Furthermore, since $p_{F}$ is a sum of squares, the Boolean assignment $x_{1:n}=\mathbf{u}$ must satisfy every clause in $F$.
-This shows the hardness of determining the existence of a zero for polynomials.
-Now, consider the scalar-state dynamical equation given by
-$x_{t+1}=(1-\varepsilon)x_{t}+p_{F}(\mathbf{u}_{t}),$
-with $p$ being the polynomial family we just constructed and $\varepsilon>0$ an arbitrarily small constant.
-For the system to reach the null state from an arbitrary start state, we would need to verify if there is an assignment of $\mathbf{u}$ such that $p_{F}(\mathbf{u})=0$, which as we just proved is NP-hard.
-∎
-The proof establishes that even for a dynamical system that is a polynomial of degree six, checking for stabilizability is NP-hard.
-Exercise 2.5 asks the reader to strengthen this result and show that checking the stabilizability of dynamical systems describable as polynomials of degree four is also NP-hard.
+We reduce from 3SAT. Let
+$F(x_{1:n})=c_{1}\land c_{2}\land\cdots\land c_{m}$
+be a Boolean formula in conjunctive normal form with three literals per clause. Introduce real variables $u_{1},\ldots,u_{n}$. For each literal $\ell$, define
+[…]
+For a clause $c=\ell_{1}\lor\ell_{2}\lor\ell_{3}$, let
+$g_{c}(\mathbf{u})=r(\ell_{1})\,r(\ell_{2})\,r(\ell_{3}).$
+On Boolean assignments $\mathbf{u}\in\{0,1\}^{n}$, the quantity $g_{c}(\mathbf{u})$ equals $1$ exactly when the clause $c$ is false and 0 otherwise.
+Now define the polynomial
+$p_{F}(\mathbf{u})=\sum_{i=1}^{n}\bigl(u_{i}(1-u_{i})\bigr)^{2}+\sum_{j=1}^{m}g_{c_{j}}(\mathbf{u})^{2}.$
+Then $p_{F}(\mathbf{u})=0$ if and only if $\mathbf{u}\in\{0,1\}^{n}$ encodes a satisfying assignment of $F$.
+Indeed, the first sum vanishes if and only if each $u_{i}\in\{0,1\}$, and the second sum vanishes if and only if no clause is false.
+Hence deciding whether a degree-six polynomial with integer coefficients has a real zero is NP-hard.
+Consider now the scalar dynamical system
+$x_{t+1}=p_{F}(\mathbf{u}_{t}).$
+If $F$ is satisfiable, choose a satisfying assignment $\mathbf{u}^{\star}$ and apply the constant control $\mathbf{u}_{t}\equiv\mathbf{u}^{\star}$.
+Then $p_{F}(\mathbf{u}^{\star})=0$, so $x_{t}=0$ for all $t\geq 1$, regardless of the initial state.
+Therefore the system is asymptotically stabilizable.
+Conversely, suppose $F$ is unsatisfiable. Then $p_{F}(\mathbf{u})>0$ for all $\mathbf{u}\in{\mathbb{R}}^{n}$. Since $p_{F}$ is continuous and
+$\sum_{i=1}^{n}\bigl(u_{i}(1-u_{i})\bigr)^{2}\to\infty\qquad\text{as }\|\mathbf{u}\|\to\infty,$
+the function $p_{F}$ attains a strictly positive global minimum $m>0$. Hence, for every control sequence,
+$x_{t+1}=p_{F}(\mathbf{u}_{t})\geq m\qquad\text{for all }t\geq 0.$
+So the state cannot converge to 0. Therefore the system is asymptotically stabilizable if and only if $F$ is satisfiable, proving NP-hardness. ∎
+The proof already shows hardness for an especially simple family of polynomial control systems.
+Exercise 2.5 asks the reader to strengthen this result and show that deciding asymptotic stabilizability for polynomial systems of degree four is also NP-hard.
 2.4 Bibliographic Remarks
@@
 For detailed information on the mathematical properties of the Lorenz attractor, see [STR14, TUC99].
-For a rigorous study of Julia sets and the Mandelbrot set, see [BY08].
-Nonlinear dynamical systems admit a variety of stability criteria; we chose to define a few that are useful for further developments in the text.
-A notable omission from our discussion is that of local and global stability; a celebrated result [Lya92] proves that the existence of certain potential functions is sufficient to guarantee global stability, thus obviating the need for a trajectory-based analysis.
-An interested reader might refer to [Kha15] for an expanded discussion on how these stability criteria compare.
+For discussion of computability questions for Julia sets and related sets, see [BY08].
+Nonlinear dynamical systems admit many stability notions beyond those introduced here, including Lyapunov stability, local and global asymptotic stability, robust invariance, and input-to-state stability.
+Lyapunov’s direct method shows that suitable potential functions can certify stability without requiring explicit trajectory formulas; see [LYA92, KHA15] for further discussion.
 An excellent survey on the computational complexity of control is [BT00].
@@
 The NP-hardness theorem in this chapter appeared in [AHM16].
-The computational complexity of mathematical programming – both constrained and unconstrained – is commented upon in [Haz19].
+The computational complexity of mathematical programming—both constrained and unconstrained—is commented upon in [HAZ19].
 2.5 Exercises
 2.1.
-Strengthen Theorem 2.5 to show that checking whether a dynamical system that is a polynomial of degree four is stabilizable is NP-hard.
+Strengthen Theorem 2.5 to show that deciding whether a dynamical system given by a polynomial of degree four is asymptotically stabilizable is NP-hard.
 Hint: Reduce from the NP-hard problem of MAX-2-SAT instead of 3-SAT.
 2.2. Let $A\in\mathcal{R}^{d\times d}$ be a diagonalizable matrix. Consider the following dynamical system:
-$\mathbf{x}_{t+1}=\frac{A\mathbf{x}_{t}}{\|\mathbf{x}_{t}\|}.$
-What are the stable and attractive equilibrium points $\mathbf{x}\in\mathcal{R}^{d}$ of the system? Are they also globally attractive? Your answer may depend on $A$.
+$\mathbf{x}_{t+1}=\begin{cases}\dfrac{A\mathbf{x}_{t}}{\|\mathbf{x}_{t}\|},&\mathbf{x}_{t}\neq\mathbf{0},\\[3.87498pt] \mathbf{0},&\mathbf{x}_{t}=\mathbf{0}.\end{cases}$
+What are the equilibrium points $\mathbf{x}\in\mathcal{R}^{d}$ of the system?
+Which of them are stable, attractive, or asymptotically stable?
+Are any of them globally attractive?
+Your answer may depend on $A$.
 2.3.
 Let $\alpha,\beta,\gamma>0$.
-Does the SIR-model, as is described in this chapter, define a stabilizable dynamical system (for an appropriate and fixed definition of the “zero state”)?
+Does the SIR model, as described in this chapter, define an asymptotically stabilizable dynamical system for a natural fixed choice of target state, such as a disease-free equilibrium?
 What about controllability?
-We assume that only starting states with the property $S_{0}+I_{0}+R_{0}=1$ are allowed and that $S_{t},I_{t},R_{t},u_{t}\geq 0$ must hold at all times.
+We assume that only starting states with the property $S_{0}+I_{0}+R_{0}=1$ are allowed and that $S_{t},I_{t},R_{t},u_{t}\geq 0$ and $\alpha u_{t}\leq S_{t}$ hold at all times.
 2.4. Consider the double integrator system:
@@
 $A=\begin{bmatrix}1&\Delta\\ 0&1\end{bmatrix},\quad B=\begin{bmatrix}0\\ 1\end{bmatrix},\quad\mathbf{x}_{t}=\begin{bmatrix}x_{t}\\ v_{t}\end{bmatrix}\in\mathbb{R}^{2},\quad u_{t}\in\mathbb{R},\quad\Delta>0.$
+Assume throughout this exercise that $\mathbf{w}_{t}\equiv\mathbf{0}$.
 The goal is to ensure that the system reaches the desired state $x^{*},0]^{\top}$ as $t\to\infty$ from any initial condition.
@@
 The previous chapters described and motivated the problem of control in dynamical systems.
-We have seen that in full generality, even simple questions about dynamical systems such as stabilizability and controllability are computationally hard.
-In this chapter, we consider a different representation of the control problem for which computationally efficient methods exist.
-This chapter introduces a setting of equal generality as that of the control problem from the previous chapter: reinforcement learning in Markov Decision Processes (MDP).
-However, the representation of states and dynamics in an MDP is more explicit and verbose.
-In such a case, we can design efficient methods aimed at optimality given the explicit state and transition representations.
-After introducing the reinforcement learning model and Markov Decision Processes, we discuss an important algorithmic technique that will be useful throughout this book: dynamic programming and the accompanying characterization of optimal policies via the Bellman equation.
+We have seen that, in full generality, even simple questions about dynamical systems such as stabilizability and controllability can be computationally hard.
+In this chapter, we study a more structured model for sequential decision making for which computationally efficient methods are available.
+This chapter focuses on finite-state, finite-action, discounted, fully observed Markov decision processes.
+In this setting, the state space and transition dynamics are explicitly represented, and this extra structure makes efficient planning algorithms possible.
+In addition, throughout this chapter we assume that the transition probabilities and rewards are known.
+When these quantities are unknown and must be inferred from interaction with the environment, one arrives at reinforcement learning in the narrower sense.
+After introducing the model of Markov decision processes, we discuss dynamic programming and the accompanying characterization of optimal policies via the Bellman equation.
+We then present two algorithmic consequences: a linear-programming formulation and the value-iteration algorithm.
 3.1 Reinforcement Learning
-Reinforcement learning is a subfield of machine learning that formally models the setting of learning through interaction in a reactive environment.
-This is in contrast to learning from examples, as in supervised learning, or inference according to a probabilistic model, as in Bayesian modeling.
+Reinforcement learning is a subfield of machine learning that studies learning through interaction with a reactive environment.
+This is in contrast to learning from labeled examples, as in supervised learning, or inference in a probabilistic model, as in Bayesian statistics.
 In reinforcement learning, we have an agent and an environment.
-The agent observes its position (or state) in the environment and takes actions that transition it to a new state.
-The environment looks at the agent’s state and hands out rewards based on a set of criteria hidden from the agent.
-Figure 3.1: The agent-environment interaction loop in reinforcement learning.
-Typically, the goal of reinforcement learning is for the agent to learn behavior that maximizes the total reward it receives from the environment.
-This methodology has led to some notable successes: machines have learned how to play Atari games and how to beat the highest-rated human players at Go.
-A crucial assumption for the reinforcement learning model is as follows:
-The Markovian assumption: Future states, observations, and rewards are independent of the past given the current state.
+The agent observes its current state in the environment and takes an action; the environment then transitions to a new state and emits a reward.
+Figure 3.1: The agent–environment interaction loop in reinforcement learning.
+Typically, the goal is to learn behavior that maximizes the total reward accumulated over time.
+This viewpoint has led to some notable successes: machines have learned how to play Atari games and how to defeat the strongest human players at Go.
+A crucial modeling assumption is the following.
+The Markovian assumption: Conditional on the current state and current action, the distribution of the next state and immediate reward is independent of the past.
 The reader can imagine situations where such an assumption is reasonable.
-For example, in the inverted pendulum setting described in Section 2.1.3, the assumption is clearly valid.
-It is not significant for the control problem how a certain state, describing the angle and the angular velocity, was reached, in order to optimally control the pendulum from this state onward.
-Other problems may call into question this assumption.
-For example, in the medical ventilation problem, it may very well be important how a certain state was reached, as future treatment may depend on the medical history of the patient.
-It can be argued that the Markovian assumption can always be made to work by adding more features to the state.
-In extreme cases, the state can contain the entire history, which makes the assumption true by default.
-However, this makes the state representation inefficient.
+For example, in the inverted pendulum setting from Section 2.1.3, the assumption is natural: once the current angle and angular velocity are known, the future evolution under a chosen torque does not depend on how that state was reached.
+Other problems call this assumption into question.
+In the medical ventilation problem, for instance, future treatment may depend on the medical history of the patient and not only on the present physiological variables.
+One possible remedy is to expand the state description so that it includes enough information to make the process Markovian.
+In the extreme, the state can contain the entire history, in which case the Markovian assumption becomes true by construction.
+Of course, such a representation may be computationally inefficient.
+In this chapter we restrict attention to the fully observed setting, in which the state itself is available to the agent.
+When only partial observations are available, the appropriate formal model is a partially observed Markov decision process (POMDP), which is beyond our present scope.
 3.1.1 A Classical Example: Bubbo’s World
-Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel and Norvig [RN02].
-In this example, a robot starts in the bottom left square, and its goal is to reach the +1+1 reward.
-The blocks +1,1+1,-1 represent terminal states, which means that the game ends when the robot reaches them.
+Before diving into the formalism, we start with a classical example from the text of Russel and Norvig [RN02].
+In this example, a robot starts in the bottom-left square, and its goal is to reach the terminal state with reward +1+1.
+The square marked 1-1 is also terminal, but undesirable.
 The block ‘X’ represents an obstacle through which the robot cannot travel.
@@
 Figure 3.2: A visual representation of Bubbo’s world.
-In this example, the robot has controls that allow it to move left / right / forward / backward.
-However, the controls are not reliable and, with $0.8$ probability, the robot moves wherever it intends to.
-With the remaining probability $0.2$, the robot moves sideways, that is, with probability $0.1$ to each side.
-Notice that the Markovian assumption clearly holds for this simple example.
-What is an optimal solution to this example?
-A solution concept for a reinforcement learning problem is a mapping from state to action.
-This mapping is called a “policy”, and depends both on the structure of the problem and on the rewards.
-For example, if the rewards in all other states are slightly negative, Bubbo wants to reach the terminal state with a persistent +1+1 reward, as given in figure 3.3.
+The robot can choose to move left, right, up, or down.
+However, the controls are noisy.
+With probability $0.8$, the robot moves in the intended direction; with the remaining probability $0.2$, it moves sideways, with probability $0.1$ to each side.
+Notice that the Markovian assumption clearly holds in this example.
+Once the robot’s current square and chosen action are specified, the distribution of the next square and the immediate reward no longer depend on the earlier history.
+A solution to this problem is a mapping from states to actions.
+Such a mapping is called a policy.
+The best policy depends not only on the transition structure but also on how rewards are assigned.
+To connect the example to the discounted infinite-horizon formalism used later in the chapter, it is convenient to model terminal states as absorbing states: entering the upper terminal state yields reward +1+1, entering the lower terminal state yields reward 1-1, and thereafter the process remains in that terminal state with reward 0.
+For example, if the reward for each non-terminal step is mildly negative, say 0.04-0.04, then Bubbo prefers to steer toward the +1+1 terminal while avoiding the risk of falling into the 1-1 terminal, as in Figure 3.3.
 […]
-Figure 3.3: The optimal policy (mapping from state to action) with reward for each step taken being -0.04.
-Conversely, if the rewards are very negative, Bubbo just wants to reach any terminal state as soon as possible, as in Figure 3.4.
+Figure 3.3: An optimal policy when the reward for each non-terminal step is 0.04-0.04.
+Conversely, if the step cost is very negative, then Bubbo prefers to terminate the episode as quickly as possible, even at the risk of ending in the 1-1 terminal.
+This leads to a different optimal policy, as in Figure 3.4.
 […]
-Figure 3.4: Here, we change the reward of each step to be -2.
-The optimal policy changes: now the robot is more concerned with ending the game as quickly as possible rather than trying to get to the +1 reward.
-We next describe a formal model to capture this and any other reinforcement learning problem.
+Figure 3.4: If the reward of each non-terminal step is changed to 2-2, the optimal policy changes: the robot is now more concerned with ending the game quickly than with reaching the +1+1 terminal.
+We next describe a formal model that captures this and many other reinforcement-learning problems.
 3.2 Markov Decision Processes
-In this chapter, we describe the general and formal model for reinforcement learning: Markov decision processes.
-To ease the exposition, we start with a review of Markov Chains, also known as Markov Processes, then proceed to more general Markov Reward Proccesses, and conclude with the full Markov Decision Process definition.
+We now describe the mathematical model used in this chapter: finite discounted Markov decision processes.
+To ease the exposition, we begin with Markov chains, proceed to Markov reward processes, and then give the full Markov decision process definition.
 3.2.1 Markov Chains
-A Markov Chain (or Markov Process) is a model of a memoryless stochastic process as a weighted directed graph.
-This is the first layer in abstraction before allowing for actions, as it only describes a process without exogenous inputs.
-It consists of a set of states and a transition matrix that describes the probability of moving from one state to another.
-The probabilities of moving between states are the weights on the edges of the graph.
-The transition matrix is of dimension $S\times S$, where $S$ is the number of states, and in entry $i,j$ we have the probability of moving from state $i$ to state $j$.
-Figure 3.5 describes a Markov Chain corresponding to a weather process with a corresponding transition matrix $P$ as follows:
+A Markov chain is a stochastic process that evolves without exogenous actions.
+It consists of a finite set of states together with a transition matrix that specifies the probability of moving from one state to another.
+Formally, let $S$ be a finite set of states.
+A Markov chain on $S$ is specified by a row-stochastic matrix $P\in[0,1]^{S\times S}$, where the entry $P_{ij}$ is the probability of moving from state $i$ to state $j$ in one step.
+Figure 3.5 describes a Markov chain corresponding to a simple weather process with transition matrix
 $P=\begin{pmatrix}0.9&0.1&0&0\\ 0.5&0&0.4&0.1\\ 0&0.5&0&0.5\\ 0&0&0&1\\ \end{pmatrix}$.
 RainSunny dayCloudySnow0.10.90.50.10.40.50.51
-Figure 3.5: A graphical description of a Markov decision process. Each state is the weather for a given day and the transition probabilities describe the chances of weather for the next day.
-Denote the distribution over states at time $t$ by $\mathbf{p}_{t}$. Then it can be seen that
+Figure 3.5: A graphical description of a Markov chain. Each state is the weather on a given day, and the transition probabilities describe the distribution of the next day’s weather.
+Denote the distribution over states at time $t$ by $\mathbf{p}_{t}$. Then
 $\mathbf{p}_{t+1}=\mathbf{p}_{t}P=\mathbf{p}_{0}P^{t+1},$
-where $\mathbf{p}_{0}$ is the initial state distribution represented as a vector.
-A central object in the study of Markov Chains is stationary distributions.
+where $\mathbf{p}_{0}$ is the initial state distribution.
+A central object in the study of Markov chains is a stationary distribution.
 Definition 3.1.
-A stationary distribution of a Markov Process is a limiting probability distribution
-$\boldsymbol{\pi}=\lim_{t\to\infty}\mathbf{p}_{0}P^{t}.$
-A stationary distribution satisfies
+A distribution $\boldsymbol{\pi}$} over the state space $S$ is a stationary distribution of a Markov chain with transition matrix $P$ if and only if
 ${\boldsymbol{\pi}}={\boldsymbol{\pi}}P.$
-A stationary distribution does not always exist nor is it always unique.
-Existence, uniqueness, and convergence to stationarity are the subject of Ergodic theory, a rich and well-studied branch of mathematics.
-For our discussion, we are interested in sufficient conditions for stationary distributions to exist.
-They are:
+For finite Markov chains, a stationary distribution always exists. It need not be unique, however, and convergence from an arbitrary initial distribution need not occur without further assumptions.
+Two standard structural properties are the following.
   1. 1.
-    Irreducibility. A Markov Chain is irreducible if and only if $\forall i,j\;\exists t\;s.t.\;P_{ij}^{t}>0$. In other words, every state is reachable from every other state.
+    Irreducibility. A Markov chain is irreducible if and only if
+    $\forall i,j\in S\;\exists t\geq 0\text{ such that }(P^{t})_{ij}>0.$
+    In words, every state is reachable from every other state.
   2. 2.
-    Aperiodicity.
-A Markov Chain is aperiodic if and only if $d(s_{i})=1$ for all states $s_{i}\in S$, where d(s_i) = gcd{t ∈N_+ | (P^t)_ii > 0 } , is the periodicity of state $s_{i}$, where $\gcd$ stands for “greatest common divisor".
-The ergodic theorem states that every irreducible and aperiodic Markov chain has a unique stationary distribution.
-Its proof is a beautiful application of spectral graph theory.
-Exercise 3.6 asks the reader for a guided proof of the ergodic theorem.
+    Aperiodicity. A Markov chain is aperiodic if and only if every state has period one. The period of a state $s_{i}$ is defined by
+    $d(s_{i})=\gcd\left\{t\in\mathbb{N}_{+}\ \middle|\ (P^{t})_{ii}>0\right\}.$
+The following finite-state ergodic theorem summarizes the role of these properties.
+Theorem 3.2 (Ergodic theorem for finite Markov chains).
+Let $P$ be a finite Markov chain.
+  1. 1.
+    $P$ has at least one stationary distribution.
+  2. 2.
+    If $P$ is irreducible, then the stationary distribution is unique.
+  3. 3.
+    If $P$ is irreducible and aperiodic, then for every initial distribution $\mathbf{p}_{0}$,
+    $\lim_{t\to\infty}\mathbf{p}_{0}P^{t}={\boldsymbol{\pi}},$
+    where $\boldsymbol{\pi}$} is the unique stationary distribution.
+Its proof is a beautiful application of spectral graph theory. Exercise 3.6 guides the reader through one standard proof of the convergence statement.
 AB11
-Figure 3.6: In this very simple Markov Chain, the state alternates every time step, so there is no unique stationary distribution. This Markov Chain has periodicity of two.
+Figure 3.6: This two-state Markov chain is irreducible but periodic with period $2$.
+It has a unique stationary distribution, namely ${\boldsymbol{\pi}}=(\frac{1}{2},\frac{1}{2})$, but the distribution of the chain does not converge to $\boldsymbol{\pi}$} from every initial state because the chain alternates deterministically between the two states.
 120.50.50.50.5
-Figure 3.7: Adding a self-loop to every node makes the periodicity to be one, and this chain is also irreducible.
-The stationary distribution exists and is unique, it is equal to $\boldsymbol{\pi}=(\frac{1}{2},\frac{1}{2})$.
-Reversible Markov Chains
-An interesting special case of Markov chains is the reversible Markov chain.
+Figure 3.7: Adding a self-loop to every node makes the period equal to one.
+This chain is irreducible and aperiodic, so its stationary distribution exists, is unique, and equals ${\boldsymbol{\pi}}=(\frac{1}{2},\frac{1}{2})$.
+An important special case is the reversible Markov chain.
 Definition 3.3.
-A Markov Chain $P$ is said to be reversible with respect to distribution $\pi$ if and only if the following hold for every two states $i,j$:
+A Markov chain $P$ is said to be reversible with respect to a distribution $\boldsymbol{\pi}$} if and only if, for every two states $i,j$,
 ${\boldsymbol{\pi}}_{i}P_{ij}={\boldsymbol{\pi}}_{j}P_{ji}.$
-A Markov chain is said to be reversible if there exists some distribution with respect to which the Markov chain is reversible.
-In exercise 3.6, we prove that a distribution that is reversible for a Markov chain is also stationary for this Markov chain.
+A Markov chain is said to be reversible if there exists some distribution with respect to which it is reversible.
+Exercise 3.6 asks the reader to prove that any distribution satisfying the reversibility equations is automatically stationary.
 3.2.2 Markov Reward Processes
-Markov Reward Processes (MRPs) add the concept of reward to Markov Chains.
-In each state, the agent receives a reward, which, for simplicity, we assume to be in the range $[0,1]$.
-In terms of algorithmic consequences, no generality is lost due to this assumption, and this is left as an exercise.
-The cumulative reward is the sum of discounted rewards over time, so that the immediate rewards are worth more than those received in the future.
-This is consistent with the economics literature, where a fixed amount of money is worth more today than in the future, because of positive interest rates.
+Markov reward processes add rewards to Markov chains.
+At each time step, the process transitions between states and emits a reward.
+For simplicity, we assume rewards lie in the range $[0,1]$.
+This normalization loses no essential generality and keeps the formulas uncluttered.
 Definition 3.4.
 A Markov reward process is a tuple $(S,P,R,\gamma)$, where:
-  • $S$ is a set of states that can be finite or infinite.
-  • $P\in{\mathbb{R}}^{S\times S}$ is the transition matrix.
-  • $R\in[0,1]^{S\times S}$ is the reward function, where $R(s,s^{\prime})$ is the reward for moving from state $s$ to $s^{\prime}$.
-  • $\gamma\in[0,1]$ is the discount factor.
-We henceforth consider instantiations of Markov Reward Processes, and use the following notation:
+  • $S$ is a finite set of states.
+  • $P\in[0,1]^{S\times S}$ is a row-stochastic transition matrix.
+  • $R\in[0,1]^{S\times S}$ is a reward function, where $R(s,s^{\prime})$ is the reward obtained when moving from state $s$ to state $s^{\prime}$.
+  • $\gamma\in[0,1)$ is the discount factor.
+We henceforth use the following notation.
   1. 1.
-    We denote by $S_{t}\in S$ the random state of the process at time $t$.
+    $S_{t}\in S$ denotes the random state of the process at time $t$.
   2. 2.
-    $R_{t}$ = the random reward received at time $t$, starting from state $S_{t}$ and moving to $S_{t+1}$
+    $R_{t}$ denotes the random reward obtained when transitioning from $S_{t}$ to $S_{t+1}$.
   3. 3.
-    $\hat{R}\in{\mathbb{R}}^{S}$ = the vector that for each state, denotes the expected reward in the next iteration as a function of the current state, ^R(s) = E[ R_t | S_t = s] .
+    $\hat{R}\in{\mathbb{R}}^{S}$ denotes the vector of expected one-step rewards:
+    $\hat{R}(s)=\mathop{\mbox{\bf E}}[R_{t}\mid S_{t}=s].$
   4. 4.
-    $G_{t}=$ the random variable denoting the sum of the discounted rewards starting from time $t$ onwards. G_t = ∑_i=0^∞ R_t+i γ^i .
-When the discount factor is strictly less than one, $G_{t}$ is well defined.
-In the case $\gamma=1$, an alternative definition of $G$ is presented below, assuming that there is a stationary distribution for the underlying Markov chain, denoted by $\boldsymbol{\pi}$.
-$G:=\lim_{t\to\infty}\frac{1}{t}\sum_{i=1}^{t}R_{i}=\sum_{s\in S}\boldsymbol{\pi}(s)\hat{R}(s)$
-One of the most important definitions in reinforcement learning is the value function:
+    $G_{t}$ denotes the discounted return from time $t$ onward:
+    $G_{t}=\sum_{i=0}^{\infty}\gamma^{i}R_{t+i}.$
+Since $\gamma<1$ and rewards are bounded, the random variable $G_{t}$ is well defined.
+A natural question is: what quantity should summarize the future behavior of a Markov reward process?
+For an ordinary Markov chain, one often studies long-run state frequencies via stationary distributions.
+Once rewards and discounting are introduced, however, the central quantity becomes the expected discounted return from a given starting state.
+This dependence on the starting state is unavoidable: for example, when $\gamma=0$, the return reduces to the immediate reward, which can vary from state to state.
+This motivates the central object in the theory of Markov reward processes: the value function.
+Because the process is time-homogeneous, the conditional expectation $\mathop{\mbox{\bf E}}[G_{t}\mid S_{t}=s]$ depends only on the state $s$, and not on the particular time index $t$.
 Definition 3.5.
-The value function maps states to their expected infinite-horizon reward,
-$\mathbf{v}(s)=\mathbb{E}[G_{t}|S_{t}=s].$
-The value function says how valuable a state is as measured by the sum of its future discounted rewards.
-One way to compute the value function is by using the Bellman equation for the Markov Reward Process, which is the following recursive equation:
-$\mathbf{v}(s)=\hat{R}(s)+\gamma\sum_{s^{\prime}}P_{s,s^{\prime}}\mathbf{v}(s^{\prime}).$
-To see that this equation holds we can write,
-$\mathbf{v}(s)$ $=\mathbb{E}[G_{t}|S_{t}=s]$
-$=\mathbb{E}[\sum_{i=0}^{\infty}\gamma^{i}R_{t+i}|S_{t}=s]$
-$=\mathbb{E}[R_{t}+\gamma\sum_{i=0}^{\infty}\gamma^{i}R_{t+1+i}|S_{t}=s]$
-$=\mathbb{E}[R_{t}+\gamma G_{t+1}|S_{t}=s]$
-$=\mathbb{E}[R_{t}+\gamma\mathbf{v}(S_{t+1})|S_{t}=s]$
-$=\hat{R}(s)+\gamma\sum_{s^{\prime}}P_{s,s^{\prime}}\mathbf{v}(s^{\prime}).$
-In matrix form, we can write the Bellman equation as $\mathbf{v}=\mathbf{\hat{R}}+\gamma P\mathbf{v}$, which simplifies to $\mathbf{v}=(I-\gamma P)^{-1}\mathbf{\hat{R}}$.
-Given $P$ and $\mathbf{\hat{R}}$, this is a linear system of equations, and thus the value function can be solved using Gaussian elimination in cubic time in the number of states.
+The value function of a Markov reward process maps states to expected discounted return:
+$\mathbf{v}(s)=\mathbb{E}[G_{t}\mid S_{t}=s].$
+The value function measures how desirable each state is in terms of expected future reward. It satisfies the Bellman equation for a Markov reward process:
+$\mathbf{v}(s)=\hat{R}(s)+\gamma\sum_{s^{\prime}\in S}P_{s,s^{\prime}}\mathbf{v}(s^{\prime}).$
+To verify this identity, write
+$\mathbf{v}(s)$ $=\mathbb{E}[G_{t}\mid S_{t}=s]$
+$=\mathbb{E}\left[\sum_{i=0}^{\infty}\gamma^{i}R_{t+i}\;\middle|\;S_{t}=s\right]$
+$=\mathbb{E}\left[R_{t}+\gamma\sum_{i=0}^{\infty}\gamma^{i}R_{t+1+i}\;\middle|\;S_{t}=s\right]$
+$=\mathbb{E}[R_{t}+\gamma G_{t+1}\mid S_{t}=s]$
+$=\mathbb{E}[R_{t}+\gamma\mathbf{v}(S_{t+1})\mid S_{t}=s]$
+$=\hat{R}(s)+\gamma\sum_{s^{\prime}\in S}P_{s,s^{\prime}}\mathbf{v}(s^{\prime}).$
+In matrix form, the Bellman equation is
+$\mathbf{v}=\hat{R}+\gamma P\mathbf{v}.$
+Equivalently, $\mathbf{v}$ is the unique solution to the linear system
+$(I-\gamma P)\mathbf{v}=\hat{R}$.
+Since $\gamma<1$ and $P$ is a stochastic matrix, the matrix $I-\gamma P$ is invertible, and this system has a unique solution.
+Thus, for a known Markov reward process, the value function is the solution to a linear system and can be computed in polynomial time in the number of states.
 3.2.3 Markov Decision Processes
-The final layer of abstraction is to allow actions or controls in the language of control theory. This gives the full Markov Decision Process model as defined below.
+The final layer of abstraction is to allow actions, or controls in the language of control theory. This gives the full Markov decision process model.
 Definition 3.6.
@@
   1. 1.
-    $S$ is a set of states.
+    $S$ is a finite set of states.
   2. 2.
-    AA is a set of possible actions.
+    AA is a finite set of possible actions.
   3. 3.
-    $P$ is a transition matrix: P_ss’^a = Pr[ S_t+1 = s | S_t = s, A_t = a].
+    $P$ is a transition kernel, written as
+    $P^{a}_{ss^{\prime}}=\Pr[S_{t+1}=s^{\prime}\mid S_{t}=s,A_{t}=a].$
+    For each action $a\in A$, the matrix $P^{a}$ is row-stochastic.
   4. 4.
-    $R$ is a reward function that gives each transition and action a reward $R_{s,s^{\prime},a}$ which is a reward for moving from $s$ to $s^{\prime}$ using action $a$.
-We henceforth assume w.l.o.g.
-that $R\in[0,1]^{S\times A}$, where we define R_sa = _s’ S P_s,s’,a R_s,s’,a .
+    $R$ is a reward function on state–action–next-state triples, where $R_{ss^{\prime}a}$ is the reward for moving from $s$ to $s^{\prime}$ using action $a$. We also write
+    $R_{sa}=\sum_{s^{\prime}\in S}P^{a}_{ss^{\prime}}R_{ss^{\prime}a}$
+    for the corresponding expected immediate reward. As above, we assume rewards lie in $[0,1]$.
   5. 5.
     $\gamma\in[0,1)$ is the discount factor.
-The central solution concept of study in the theory of reinforcement learning is called a policy. A policy is a mapping from state to distribution over actions, that is,
-$\pi:S\to\Delta(A)\ ,\ \pi(a|s)=\Pr[A_{t}=a|S_{t}=s]$
-The goal of reinforcement learning is to determine a sequence of actions to maximize long-term reward.
-The Markovian structure of the problem suggests that this optimal value is realized by a stationary policy, i.e.
-there exists a fixed map from states to actions that maximizes the expected rewards.
-Thus, the goal of reinforcement learning is to find a policy which maximizes the expected reward starting from a certain starting state, i.e.
-$\max_{\pi\in S\to\Delta(A)}\mathop{\mbox{\bf E}}\left\{\sum_{t=1}^{\infty}\gamma^{t-1}R_{t}|a_{t}\sim\pi(s_{t})\ ,\ S_{1}=s_{1}\right\},$
-where $R_{t}$ is a random variable that denotes the reward at time $t$ that depends on the randomness in the policy as well as the transitions.
-Similarly to the value function of a Markov Reward Process, we can define the value function for a Markov Decision Process as follows.
-The value function for a state $s$ given a policy $\pi$ is defined to be the expected reward from this state onwards where the actions are chosen according to $\pi$ at each state, i.e.
-$\mathbf{v}_{\pi}(s)=\mathop{\mbox{\bf E}}_{\pi,P}\left[\sum_{t=1}^{\infty}\gamma^{t-1}R_{t}|S_{1}=s\right].$
-Similarly, the $Q$ function is defined over state-action pairs given a policy $\pi$ to be the expected cumulative reward gained when starting from some state $s$ and action $a$ and thereafter choosing actions according to $\pi$ at each future state, i.e.
-$Q_{\pi}(s,a)=\mathop{\mbox{\bf E}}_{\pi,P}\left[\sum_{t=1}^{\infty}\gamma^{t-1}R_{t}|S_{1}=s,A_{1}=a\right].$
-A fundamental theorem of reinforcement learning, which is constructively proven in Section 3.4, is as follows.
+The central solution concept is a policy.
+Definition 3.7.
+A policy is a rule for choosing actions as the process evolves.
+A policy may in principle depend on the entire history.
+A stationary policy is a policy that depends only on the current state, and is therefore a mapping
+$\pi:S\to\Delta(A),\qquad\pi(a\mid s)=\Pr[A_{t}=a\mid S_{t}=s].$
+If each $\pi(\cdot\mid s)$ is a point mass, the policy is called deterministic.
+In this chapter we consider infinite-horizon discounted return. Starting from a state $s$, the value of a policy $\pi$ is
+$\mathbf{v}_{\pi}(s)=\mathop{\mbox{\bf E}}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{t}\;\middle|\;S_{0}=s\right].$
+The optimal value function is defined pointwise by
+$\mathbf{v}^{*}(s)=\sup_{\pi}\mathbf{v}_{\pi}(s).$
+A fundamental theorem of discounted MDPs is the following.
 Theorem 3.8.
-There exists a deterministic optimal policy $\pi^{*}$, which maps every state to a single action.
-All optimal policies achieve the same optimal value $\mathbf{v}^{*}(s)$ as well as $Q^{*}(s,a)$, and this holds for all states and actions.
-The fact that there exists an optimal policy that is optimal regardless of the starting state in a MDP is a consequence of the Markovian nature of the transition dynamics.
+For every finite discounted Markov decision process, there exists a deterministic stationary optimal policy $\pi^{*}$. Moreover, every optimal policy has the same value function $\mathbf{v}^{*}$.
+The existence of an optimal stationary policy is a consequence of the Markovian structure of the problem and the Bellman optimality equation, which we now discuss.
 3.3 The Bellman Equation
-A fundamental property of Markov Decision Processes is that they satisfy the Bellman equation for dynamic programming.
-The importance of this equation is paramount in algorithm design: we describe two efficient methods for reinforcement learning based on this equation later in this chapter.
-In later chapters, we also make use of the Bellman equation to derive the stochastic-optimal LQR controller.
-The Bellman equation for any policy $\pi$ relates the value function of that policy to itself through one step of application for that policy. Mathematically, this amounts to the following:
-$\mathbf{v}_{\pi}(s)=\sum_{a\in A}\pi(a|s)\left[R_{sa}+\gamma\sum_{s^{\prime}\in S}P_{ss^{\prime}}^{a}\mathbf{v}_{\pi}(s^{\prime})\right].$
-It is sometimes convenient to use matrix-vector notation, where we can write:
+A fundamental property of Markov decision processes is that they satisfy a recursive optimality principle. This principle is expressed by the Bellman equation and is the basis for dynamic programming.
+For a fixed stationary policy $\pi$, the Bellman equation takes the form
+$\mathbf{v}_{\pi}(s)=\sum_{a\in A}\pi(a\mid s)\left[R_{sa}+\gamma\sum_{s^{\prime}\in S}P_{ss^{\prime}}^{a}\mathbf{v}_{\pi}(s^{\prime})\right].$
+In matrix–vector notation this can be written as
 $\mathbf{v}_{\pi}=\mathbf{R^{\pi}}+\gamma P^{\pi}\mathbf{v}_{\pi},$
-where we use the notation
+where
   1. 1.
-    $R^{\pi}$ is the expected reward vector induced when playing the policy $\pi$, that is, ${R^{\pi}}(s)=\mathop{\mbox{\bf E}}_{a\sim\pi(s)}[R_{sa}]$.
+    $R^{\pi}$ is the expected reward vector induced by $\pi$, defined by
+    $R^{\pi}(s)=\mathop{\mbox{\bf E}}_{a\sim\pi(\cdot\mid s)}[R_{sa}],$
   2. 2.
-    $P^{\pi}$ is the induced transition matrix when playing policy $\pi$, i.e. P^π_ss’ = ∑_a ∈A P^a_ss’ ⋅π(a|s).
-The Bellman equation gains prominence when applied to the optimal policy of a Markov Decision Process.
-Then it is called the Bellman optimality equation, and states that for every state and action, we have
-$\mathbf{v}^{*}(s)=\max_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}$
-Equivalently, this equation can be written in terms of the $Q$ function as
-$\mathbf{v}^{*}(s)=\max_{a}\left\{Q^{*}(s,a)\right\},$
-where $Q^{*}$ is the $Q$ function of the optimal policy, i.e.,
-$Q^{*}(s,a)=R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\max_{a^{\prime}}\left\{Q^{*}(s^{\prime},a^{\prime})\right\}.$
-In subsequent chapters, we will also consider cases where the return of interest is either the reward accumulated in a finite number of time steps, that is, $G=\sum_{t=1}^{T}R_{t}$ in the finite-time setting, or the infinite-time average reward case, i.e.
-$G=\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}R_{t}$.
-In the finite-time setting, both optimal policies and the optimal value function are non-stationary – they depend on the time step in addition to the current state.
-In this setting, the corresponding Bellman optimality equation can be written as
-$\mathbf{v}_{t}^{*}(s)=\max_{a}\left\{R_{sa}+\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}_{t+1}(s^{\prime})\right\}$
-starting with $\mathbf{v}_{T}(s)=0$ for all states.
-In the average-reward case, the Bellman optimality equation offers the following characterization: $\mathbf{v}^{*}:S\to{\mathbb{R}}$ is the optimal value function if there exists some $\lambda\in{\mathbb{R}}$ such that
-$\lambda+\mathbf{v}^{*}(s)=\max_{a}\left\{R_{sa}+\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}.$
-Here the scalar $\lambda$ corresponds to the optimal average long-term reward, and $\mathbf{v}^{*}(s)$ reflects the transient advantage of being at state $s$.
-3.3.1 Reinforcement Learning as Linear Programming
-One of the immediate conclusions from the Bellman optimality equation is an efficient algorithm for reinforcement learning in Markov Decision Processes via linear programming.
-Consider the problem of computing the value function corresponding to an optimal policy.
-The optimal value function gives the optimal policy, and deriving this implication explicitly is left as an exercise.
-According to the Bellman optimality equation, the optimal value function satisfies:
-$\mathbf{v}^{*}(s)=\max_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\},$
-or in vector form,
-$\mathbf{v}^{*}=\max_{\mathbf{a}\in A^{|S|}}\left\{\mathbf{R^{\mathbf{a}}}+\gamma P^{\mathbf{a}}\mathbf{v}^{*}\right\},$
-where the maximization over actions takes place independently for each state. Equivalently, we can write the following mathematical program:
-$\min_{\mathbf{v}^{*}}\mathbf{1}^{\top}\mathbf{v}^{*}$
-subject to $\mathbf{v}^{*}(s)\geq R_{sa}+\gamma\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})$ $\forall(s,a)\in S\times A.$
-Notice that this is a linear program!
-The number of variables is $|S|$, and the number of constraints is $|S|\times|A|$, which can be solved in polynomial time in the number of states and actions.
-The feasibility of this linear program is not immediate but will be established in the next section.
+    $P^{\pi}$ is the transition matrix induced by $\pi$, defined by
+    $P^{\pi}_{ss^{\prime}}=\sum_{a\in A}P^{a}_{ss^{\prime}}\,\pi(a\mid s).$
+Since $\gamma<1$, the matrix $I-\gamma P^{\pi}$ is invertible, and therefore the Bellman equation for a fixed policy has a unique solution.
+The Bellman equation becomes especially powerful when applied to the optimal value function.
+$\mathbf{v}^{*}(s)=\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}$
+This is the Bellman optimality equation. It says that the optimal value of a state is obtained by choosing the action with the largest one-step reward plus discounted continuation value.
+The proof is a one-step decomposition.
+Fix a state $s$.
+For any policy, the value obtained from $s$ is bounded above by the best possible choice of first action plus the optimal continuation value thereafter, so
+$\mathbf{v}^{*}(s)\leq\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}.$
+Conversely, choose an action $a^{*}$ achieving the maximum on the right-hand side.
+For each successor state $s^{\prime}$, choose a policy that is $\varepsilon$-optimal from $s^{\prime}$.
+By first taking action $a^{*}$ and then following the chosen continuation policy, one achieves a value at least
+$R_{sa^{*}}+\gamma\sum_{s^{\prime}\in S}P^{a^{*}}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime}),$
+and the reverse inequality follows.
+The Bellman optimality equation immediately implies the existence of a deterministic optimal policy. Indeed, define a policy $\pi^{*}$ by choosing at each state any maximizing action,
+$\pi^{*}(s)\in\arg\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}.$
+Then $\mathbf{v}^{*}$ satisfies
+$\mathbf{v}^{*}(s)=R_{s\pi^{*}(s)}+\gamma\sum_{s^{\prime}\in S}P^{\pi^{*}(s)}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\qquad\text{for all }s\in S,$
+which is exactly the Bellman equation for the fixed policy $\pi^{*}$.
+By uniqueness of the solution of the Bellman equation for a fixed policy, we obtain $\mathbf{v}_{\pi^{*}}=\mathbf{v}^{*}$, proving that $\pi^{*}$ is optimal.
+This proves Theorem 3.8.
+3.3.1 Linear Programming Formulation
+The Bellman optimality equation characterizes $v^{*}$ as a fixed point. An equivalent characterization views $v^{*}$ as the smallest function satisfying the Bellman inequalities.
+Consider the following linear program:
+$\min_{\mathbf{v}\in\mathbb{R}^{|S|}}\quad$ $\sum_{s\in S}\mathbf{v}(s)$
+s.t. $\mathbf{v}(s)\geq r(s,a)+\gamma\sum_{s^{\prime}\in S}P(s^{\prime}\mid s,a)\,\mathbf{v}(s^{\prime})\quad\forall s\in S,\ a\in A.$
+This program has $|S|$ variables and |S||A||S|\,|A| constraints, so it can be solved in polynomial time in the numbers of states and actions.
+The next theorem shows that its unique optimum is precisely the optimal value function.
+Theorem 3.9.
+The unique optimal solution of the above linear program is the optimal value function $\mathbf{v}^{*}$.
+Proof.
+First, $\mathbf{v}^{*}$ is feasible, since it satisfies the Bellman optimality equation and therefore all Bellman inequalities.
+Now let $\mathbf{v}$ be any feasible solution. Then for all $s$,
+$\mathbf{v}(s)\geq\max_{a}\left\{r(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}\mid s,a)\,\mathbf{v}(s^{\prime})\right\}=(T\mathbf{v})(s),$
+where $T$ is the Bellman optimality operator. Thus $\mathbf{v}\geq T\mathbf{v}$. Iterating,
+$\mathbf{v}\geq T\mathbf{v}\geq T^{2}\mathbf{v}\geq\cdots.$
+Since $T^{k}\mathbf{v}\to\mathbf{v}^{*}$ as $k\to\infty$, it follows that $\mathbf{v}\geq\mathbf{v}^{*}$ componentwise.
+Therefore $\mathbf{v}^{*}$ is the smallest feasible vector, and hence minimizes any strictly increasing objective such as $\sum_{s}\mathbf{v}(s)$. Uniqueness follows. ∎
 3.4 The Value Iteration Algorithm
-In the previous section, we established the Bellman equation and its consequence to a polynomial-time solution for reinforcement learning in Markov Decision Processes.
-In this section, we study a fundamental algorithm for solving Markov Decision Processes that is a more direct application of dynamic programming and the Bellman equation.
-The value iteration algorithm starts from an arbitrary solution and improves it until a certain measure of optimality is reached.
-A formal description is given in Algorithm 2.
-The performance guarantee of Value Iteration, as well as its analysis, is particularly elegant and does not require expertise in mathematical optimization theory.
+The Bellman optimality equation also leads to one of the most fundamental algorithms for discounted MDPs: value iteration.
+Starting from an initial estimate of the value function, the algorithm repeatedly applies the Bellman optimality update.
 Algorithm 2 Value Iteration
 Input: MDP
- Set $\mathbf{v}_{0}(s)=0$ to all states
-for $t=1$ to T do
+ Set $\mathbf{v}_{0}(s)=0$ for all states $s$
+for $t=0$ to T1T-1 do
   Update for all states,
-$\mathbf{v}_{t+1}(s)=\max_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P_{ss^{\prime}}^{a}\mathbf{v}_{t}(s^{\prime})\right\}$
+$\mathbf{v}_{t+1}(s)=\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P_{ss^{\prime}}^{a}\,\mathbf{v}_{t}(s^{\prime})\right\}$
 end for
-return $\pi_{T}$ which is derived from the last value function as
-$\pi_{T}(s)=\operatorname*{\arg\max}_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P_{ss^{\prime}}^{a}\mathbf{v}_{T}(s^{\prime})\right\}$
-Beyond being simple and providing a finite-time convergence rate, the analysis also establishes feasibility of the optimal solution, as given in the following theorem.
+ Define a greedy policy with respect to $\mathbf{v}_{T}$ by
+$\pi_{T}(s)\in\arg\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P_{ss^{\prime}}^{a}\,\mathbf{v}_{T}(s^{\prime})\right\}$
+return $\mathbf{v}_{T}$ and $\pi_{T}$
+The next theorem gives the basic convergence guarantee.
 Theorem 3.10.
-Define the residual vector $\mathbf{r}_{t}(s)=\mathbf{v}^{*}(s)-\mathbf{v}_{t}(s)$.
-Then the optimal value vector $\mathbf{v}^{*}$ is unique and the value iteration converges to the optimal value function and policy at a geometric rate of
-$\|\mathbf{r}_{t+1}\|_{\infty}\leq\gamma^{t}\|\mathbf{r}_{1}\|_{\infty}.$
+Let $(\mathbf{v}_{t})_{t\geq 0}$ be the sequence produced by Value Iteration with initialization $\mathbf{v}_{0}=\mathbf{0}$. Then:
+  1. 1.
+    $\mathbf{0}\leq\mathbf{v}_{t}\leq\mathbf{v}_{t+1}\leq\mathbf{v}^{*}$ entry-wise for every $t\geq 0$.
+  2. 2.
+    If we define the residual vector $\mathbf{r}_{t}=\mathbf{v}^{*}-\mathbf{v}_{t}$, then
+    $\|\mathbf{r}_{t+1}\|_{\infty}\leq\gamma\|\mathbf{r}_{t}\|_{\infty}\qquad\text{for every }t\geq 0.$
+    Consequently,
+    $\|\mathbf{v}^{*}-\mathbf{v}_{t}\|_{\infty}\leq\frac{\gamma^{t}}{1-\gamma}.$
+In particular, $\mathbf{v}_{t}\to\mathbf{v}^{*}$ geometrically fast.
 Proof.
+Monotonicity and the upper bound are left as an exercise. We now prove the geometric decay of the residual.
 For every state we have by choice of the algorithm that
@@
 $\|\mathbf{r}_{t+1}\|_{\infty}\leq\gamma^{t}\|\mathbf{r}_{1}\|_{\infty}.$
+Finally, because rewards lie in $[0,1]$, the total discounted reward is at most $1+\gamma+\gamma^{2}+\cdots=1/(1-\gamma)$, so
+$\|\mathbf{r}_{0}\|_{\infty}=\|\mathbf{v}^{*}\|_{\infty}\leq\frac{1}{1-\gamma}.$
+Combining the two inequalities yields
+$\|\mathbf{v}^{*}-\mathbf{v}_{t}\|_{\infty}=\|\mathbf{r}_{t}\|_{\infty}\leq\frac{\gamma^{t}}{1-\gamma}.$
 3.5 Bibliographic Remarks
-The section on reinforcement learning in the text [RN02] introduces basic concepts, along with commentaries on related topics, such as the origins of the scalar reward model from utility theory.
-The classical text on reinforcement learning, authored by pioneers of the subject, is [SB18].
-In addition, this text extensively discusses early ventures into the field of researchers working in artificial intelligence, psychology, and neuroscience.
-Another great reference is the set of lecture notes [Sil15].
-See [AJKS19] for an excellent introduction to the theoretical aspects of reinforcement learning, such as sample complexity bounds and reductions to supervised learning.
-The reader may also refer to [Ber19] for an extensive survey of computational approaches to learning in unknown Markov decision processes.
-The first author had the benefit of a discussion with [Aro20] on the proof of convergence of value iteration presented in this chapter.
-Dynamic programming approaches, in addition to being widely used in algorithm design [CLRS22], are at the heart of Markov decision processes.
-This view is emphasized in [Ber07], which discusses the sufficiency and necessity of the Bellman optimality conditions at length.
-It is also a good resource for an involved reader who might be interested in the completeness of the Bellman characterization and the existential concerns for the average-time setting, which we did not expand upon.
-Although tabular MDPs do not compactly represent stochastic linear dynamical systems, function-approximation schemes [JYWJ20, DKL+21, JKA+17, KKL+20] successfully capture the stochastic LQR setting.
+The section on reinforcement learning in the text [RN02] introduces basic concepts and places them in the broader context of artificial intelligence.
+The classical textbook on reinforcement learning is [SB18], which develops the subject from both the algorithmic and conceptual perspectives.
+Another excellent reference is the set of lecture notes [SIL15].
+See [AJK+19a] for an introduction to the theoretical aspects of reinforcement learning, including sample-complexity guarantees and reductions to supervised learning.
+The reader may also refer to [BER19] for a broad survey of computational approaches to learning in unknown Markov decision processes.
+Dynamic programming is at the heart of Markov decision processes.
+This viewpoint is emphasized in [BER07], which discusses the Bellman equations and related algorithmic ideas at length.
+It is also a useful source for the finite-horizon and average-reward settings, which we have only mentioned briefly here.
+Although tabular MDPs do not compactly represent stochastic linear dynamical systems, function-approximation schemes [JYW+20, DKL+21, JKA+17, KKL+20] can successfully capture problems such as stochastic LQR.
 3.6 Exercises
-3.1. In this exercise, we give a guided proof of the ergodic theorem. Consider an irreducible and aperiodic Markov chain with a transition matrix $P$.
-A. Show that $P$ has a left eigenvalue equal to 1.
-B. Show that all eigenvalues of $P$ are at most one in magnitude.
-The irreducibility and aperiodicity of the Markov chain imply that the eigenvalue equal to one is unique. We are going to assume this fact and denote by $\mathbf{v}_{1}$ the corresponding eigenvector.
-C.
-Consider an abirtrary distribution $\mathbf{p}_{0}$, and the distribution after $t$ iterations of the Markov chain $\mathbf{p}_{t}=\mathbf{p}_{0}P^{t}$.
-Prove that $\lim_{t\to\infty}\mathbf{p}_{t}=\mathbf{v}_{1}.$
-3.2. Prove the geometric convergence of the value iteration algorithm for rewards that can be arbitrary real numbers, as opposed to in the range $[0,1]$.
-3.3. Consider the derivation of an optimal policy for a Markov Decision Process given its value function as described in algorithm 2. Prove that the resulting policy is indeed optimal.
-3.4. Let Markov chain $P$ be reversible with respect to stationary distribution $\boldsymbol{\pi}$}, i.e.
+3.1. In this exercise, we give a guided proof of the convergence part of the ergodic theorem. Consider a finite irreducible and aperiodic Markov chain with transition matrix $P$.
+A. Show that the all-ones vector is a right eigenvector of $P$ with eigenvalue $1$. Conclude that $1$ is also an eigenvalue of $P^{\top}$.
+B. Show that all eigenvalues of $P$ have magnitude at most one.
+The irreducibility and aperiodicity of the Markov chain imply that the eigenvalue $1$ of $P^{\top}$ is simple and that all other eigenvalues have magnitude strictly smaller than one.
+Assume this fact, and let $\mathbf{v}_{1}$ denote the corresponding left eigenvector normalized to sum to one.
+C. Consider an arbitrary initial distribution $\mathbf{p}_{0}$, and the distribution after $t$ iterations of the Markov chain, $\mathbf{p}_{t}=\mathbf{p}_{0}P^{t}$. Prove that
+$\lim_{t\to\infty}\mathbf{p}_{t}=\mathbf{v}_{1}.$
+3.2. Prove the geometric convergence of the value-iteration algorithm for rewards that can be arbitrary bounded real numbers, as opposed to lying in the range $[0,1]$.
+3.3. Let $\mathbf{v}^{*}$ be the optimal value function of a discounted Markov decision process, and define a greedy policy by
+$\pi(s)\in\arg\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}.$
+Prove that $\pi$ is optimal.
+3.4. Let Markov chain $P$ be reversible with respect to a distribution $\boldsymbol{\pi}$}, i.e.
 $\forall i,j\ ,\ {\boldsymbol{\pi}}_{i}P_{ij}={\boldsymbol{\pi}}_{j}P_{ji}.$
 Prove that $\boldsymbol{\pi}$} is stationary for $P$.
-3.5.
-Prove that the successive iterates $\mathbf{v}_{t}$ produced by the Value Iteration algorithm increase monotonically, that is, $\mathbf{v}_{t+1}\geq\mathbf{v}_{t}$, when initialized with $\mathbf{v}_{0}=\mathbf{0}$.
-Hence, conclude that for any $t$, $\mathbf{v}^{*}\geq\mathbf{v}_{t}$.
-3.6. Let $P$ be an irreducible and aperiodic Markov chain. Prove that $P$ is reversible if and only if for any sequence of states $s_{1},\dots s_{n}$, it holds that
-$P_{s_{1}s_{2}}P_{s_{2}s_{3}}\dots P_{s_{n-1}s_{n}}P_{s_{n}s_{1}}=P_{s_{1}s_{n}}P_{s_{n}s_{n-1}}\dots P_{s_{2}s_{1}}.$
+3.5. Suppose Value Iteration is initialized at an arbitrary vector $\mathbf{v}_{0}$ satisfying
+$\mathbf{v}_{0}(s)\leq\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P^{a}_{ss^{\prime}}\mathbf{v}_{0}(s^{\prime})\right\}\qquad\text{for all }s\in S.$
+Prove that the successive iterates $\mathbf{v}_{t}$ increase monotonically, that is, $\mathbf{v}_{t+1}\geq\mathbf{v}_{t}$ for all $t$, and that $\mathbf{v}_{t}\leq\mathbf{v}^{\star}$.
 Chapter 4 Linear Dynamical Systems
 In Chapter 2, we observed that even simple questions about general dynamical systems are intractable.
-This motivates the restriction to a special class of dynamical systems for which we can design efficient algorithms with provable guarantees.
+This motivates restricting attention to a special class of dynamical systems for which we can design efficient algorithms with provable guarantees.
 Our treatment in this chapter is restricted to linear dynamical systems (LDS): dynamical systems that evolve linearly according to the equation
 $\mathbf{x}_{t+1}={A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
-If the matrices ${A}_{t}={A}$ and ${B}_{t}={B}$ are fixed, we say that the dynamical system is linear time-invariant (LTI).
-We have seen examples of LTI systems in Chapter 2, namely the double integrator and the aircraft dynamics.
+If the matrices ${A}_{t}={A}$ and ${B}_{t}={B}$ are fixed, we say that the dynamical system is linear time-invariant (LTI). We have already seen examples of LTI systems in Chapter 2.
 Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories.
 Created by XaosBits using the Mathematica packages CurvesGraphics.nb by Gianluca Gorni and DrawGraphics.nb by David and Alice Park, licensed under $C$-BY-2.5.
-A more general class of linear dynamical systems is when the matrices ${A}_{t}$ and ${B}_{t}$ change over time.
+A more general class of linear dynamical systems arises when the matrices ${A}_{t}$ and ${B}_{t}$ change over time.
 These are called linear time-varying (LTV) systems.
@@
 4.1 Approximating General Dynamics via LTV Systems
-Consider any dynamics that can be expressed by a smooth function $f$.
-Such a system can be approximated by a time-varying linear system to an arbitrary accuracy determined by the time-step interval between samples.
-To see this, first consider the one dimensional first-order Taylor expansion of a given dynamics function $f$, which is twice differentiable, around an anchor point $(\bar{x}_{t},\bar{u}_{t})$:
-$x_{t+1}$ $=f(x_{t},u_{t})+w_{t}$
-$=f(\bar{x}_{t},\bar{u}_{t})+\underbrace{\frac{\partial}{\partial x}f(x,u)\bigg|_{\bar{x}_{t},\bar{u}_{t}}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}A_{t}}(x_{t}-\bar{x}_{t})+\underbrace{\frac{\partial}{\partial u}f(x,u)\bigg|_{\bar{x}_{t},\bar{u}_{t}}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}B_{t}}(u_{t}-\bar{u}_{t})+w_{t},$
-where $w_{t}$ is the approximation error, or remainder, which by the mean value theorem behaves as $w_{t}\sim(x_{t}-\bar{x}_{t})^{2}+(u-\bar{u}_{t})^{2}$. Thus, this simplifies to
-$x_{t+1}=A_{t}x_{t}+B_{t}u_{t}+w_{t}.$
-An analogous approximation holds for the higher-dimensional case. The Taylor expansion is
-$f(\mathbf{x},\mathbf{u})=f(\bar{\mathbf{x}},\bar{\mathbf{u}})+J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{x}-\bar{\mathbf{x}})+J_{\mathbf{u}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{u}-\bar{\mathbf{u}})+w_{t},$
-where $J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is the Jacobian of $f$ with respect to variables $\bar{\mathbf{x}}$. Writing the matrices ${A}_{t},{B}_{t}$ for the Jacobians, we have
-$\mathbf{x}_{t+1}={A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
-This derivation explains why in practice the class of time-varying linear dynamical systems is very expressive, at least for smooth dynamical systems.
-This is particularly true with hidden states that can capture complicated dynamics while imposing enough structure for efficient algorithms.
+Consider deterministic dynamics of the form
+$\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t}),$
+where $f$ is twice continuously differentiable. Fix a nominal state-control trajectory $(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})$ satisfying
+$\bar{\mathbf{x}}_{t+1}=f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t}),$
+and define the deviation variables
+$\delta\mathbf{x}_{t}=\mathbf{x}_{t}-\bar{\mathbf{x}}_{t},\qquad\delta\mathbf{u}_{t}=\mathbf{u}_{t}-\bar{\mathbf{u}}_{t}.$
+A first-order Taylor expansion around $(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})$ gives
+$f(\mathbf{x}_{t},\mathbf{u}_{t})=f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})+J_{\mathbf{x}}f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})(\mathbf{x}_{t}-\bar{\mathbf{x}}_{t})+J_{\mathbf{u}}f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})(\mathbf{u}_{t}-\bar{\mathbf{u}}_{t})+r_{t},$
+where $J_{\mathbf{x}}f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})$ and $J_{\mathbf{u}}f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})$ are the Jacobians of $f$ with respect to $\mathbf{x}$ and $\mathbf{u}$, evaluated at $(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})$, and where the remainder satisfies
+$\|r_{t}\|=O\!\left(\|\delta\mathbf{x}_{t}\|^{2}+\|\delta\mathbf{u}_{t}\|^{2}\right)$
+locally around the nominal trajectory. Writing
+$A_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}J_{\mathbf{x}}f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t}),\qquad B_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}J_{\mathbf{u}}f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t}),$
+and subtracting the nominal dynamics yields the deviation dynamics
+$\delta\mathbf{x}_{t+1}=A_{t}\delta\mathbf{x}_{t}+B_{t}\delta\mathbf{u}_{t}+r_{t}.$
+Thus, near a nominal trajectory, a smooth nonlinear dynamical system is well approximated by a time-varying linear system in the deviation variables.
+If the original nonlinear system also has additive perturbations, then those perturbations simply enter additively in the linearized deviation dynamics as well.
+This derivation explains why in practice the class of time-varying linear dynamical systems is very expressive, at least locally for smooth dynamical systems.
+In chapter 13 we see another reduction to LDS from nonlinear dynamics.
 As we shall see next, stabilizability and controllability can be checked efficiently for linear systems, in contrast to general dynamics.
@@
 Verifying stability of a system, and more so stabilizability, was shown to be intractable in our discussion of Chapter 2.
-For linear dynamical systems, however, the picture is quite different: stability and stabilizability can be efficiently verified!
-Furthermore, the theoretical analysis of this fact is interesting and elegant.
+For linear dynamical systems, however, the picture is quite different: stability and stabilizability can be efficiently verified.
+Furthermore, the theoretical analysis of this fact is elegant.
 For simplicity, in the rest of this section we consider time-invariant linear dynamical systems of the form
@@
 although much of our discussion can be generalized to LTV systems.
-Recall that a system is stabilizable if, from every starting state, there is a sequence of controls that drives the system to zero.
-Determining whether a linear time-invariant system is stabilizable is determined by a particularly simple condition as follows.
+Recall that a system is stabilizable if, in the absence of perturbations, from every starting state there exists a control policy that drives the state to zero.
+Determining whether a linear time-invariant system is stabilizable admits a particularly simple characterization.
 Henceforth, we denote by $\rho({A})$ the spectral radius of a matrix A{A}, that is,
-$\rho({A})=\max\{|\lambda_{1}|,...,|\lambda_{d}|\},$
-the maximum modulus of all (possibly complex) eigenvalues. The stabilizability of a linear dynamical system can be characterized as follows.
+$\rho({A})=\max\{|\lambda_{1}|,\ldots,|\lambda_{d}|\},$
+the maximum modulus of all (possibly complex) eigenvalues. The stabilizability of a linear dynamical system can then be characterized as follows.
 Theorem 4.1.
@@
 $\rho({A}+{B}{K})<1.$
-We therefore prove the sufficiency of this condition and leave the necessity as exercise 4.6. Both directions rely on the characterization of the spectral norm as follows:
+We prove the sufficiency of this condition and leave the necessity as Exercise 4.6. Both directions rely on the following characterization of the spectral radius.
 Fact 4.2.
-$\rho({A})<1$ if and only if $\lim_{t\mapsto\infty}{A}^{t}=0$. Furthermore, if $\rho({A})>1$, then for any norm, $\lim_{t\mapsto\infty}\|{A}^{t}\|=\infty$.
-The proof of this fact is left as exercise 4.6. Based on this fact, we can now prove the sufficiency part of Theorem 4.1.
+$\rho({A})<1$ if and only if $\lim_{t\to\infty}{A}^{t}=0$. Furthermore, if $\rho({A})>1$, then for any norm,
+$\lim_{t\to\infty}\|{A}^{t}\|=\infty.$
+The proof of this fact is left as Exercise 4.6. Based on it, we can now prove the sufficiency part of Theorem 4.1.
 Theorem 4.1, sufficiency.
-Suppose that there exists a K{K} such that $\rho({A}+{B}{K})<1$. Then choose all controls to be $\mathbf{u}_{t}={K}\mathbf{x}_{t}$. The evolution of the system without perturbation is given by
-$\mathbf{x}_{t}=({A}+{B}{K})\mathbf{x}_{t-1}=...=({A}+{B}{K})^{t}\mathbf{x}_{0}.$
-Thus, using the previous fact concerning the spectral radius, we have
-$\lim_{t\to\infty}\|\mathbf{x}_{t}\|=\lim_{t\to\infty}\|({A}+{B}{K})^{t}x_{0}\|=\|0\mathbf{x}_{0}\|=0.$
+Suppose that there exists a matrix K{K} such that $\rho({A}+{B}{K})<1$.
+Choose the linear state-feedback controller $\mathbf{u}_{t}={K}\mathbf{x}_{t}$.
+In the noiseless system, the closed-loop dynamics are
+$\mathbf{x}_{t+1}=({A}+{B}{K})\mathbf{x}_{t},$
+and therefore
+$\mathbf{x}_{t}=({A}+{B}{K})^{t}\mathbf{x}_{0}.$
+By the fact above, $(A+BK)^{t}\to 0$ as $t\to\infty$. Hence
+$\lim_{t\to\infty}\|\mathbf{x}_{t}\|=\lim_{t\to\infty}\|({A}+{B}{K})^{t}\mathbf{x}_{0}\|=0.$
+Thus the system is stabilizable. ∎
 4.2.1 Efficient Verification of Stabilizability
-The spectral radius of a matrix can be computed to an arbitrary accuracy in polynomial time and can be approximated by the power method.
-This implies that the stability of a given system can be efficiently verified.
-To efficiently check if a given system is stabilizable, we make note of the following result due to Lyapunov.
-We use the notation $A\succeq 0$ to denote that the matrix $A$ is positive semidefinite, meaning that it is symmetric and all its eigenvalues are nonnegative.
+Stabilizability admits a convenient convex characterization.
+We use the following standard form of the discrete-time Lyapunov theorem: a matrix $M$ satisfies $\rho(M)<1$ if and only if there exists a matrix $P\succ 0$ such that
+$MPM^{\top}\prec P.$
+By scaling $P$, we may additionally impose $P\succeq I$.
+Applying this to the closed-loop matrix $A+BK$ gives the following linear matrix inequality.
 Theorem 4.3.
-$\rho({A})<1$ holds if and only if there exists a matrix ${P}\in{\mathbb{R}}^{d_{x}\times d_{x}}$ such that ${P}\succeq{I}$ and ${P}\succ{A}^{\top}{P}{A}$.
+There exists a matrix $K$ such that $\rho(A+BK)<1$ if and only if there exist matrices $P\in\mathcal{S}^{d_{x}\times d_{x}}$ and $L\in{\mathbb{R}}^{d_{u}\times d_{x}}$ such that
+$\begin{bmatrix}P&({A}{P}+{B}{L})^{\top}\\ {A}{P}+{B}{L}&P\end{bmatrix}\succ 0,\qquad P\succeq I.$
+Moreover, any feasible pair $(P,L)$ yields a stabilizing controller $K=LP^{-1}$.
 Proof.
-Say there exists a matrix $P$ such that $P\succeq I$ and $P\succ A^{\top}PA$, then to show that $\rho(A)<1$, it suffices to prove that $\lim_{t\to\infty}A^{t}=0$.
-To observe the latter, first note that there exists a positive constant $c<1$ such that $cP\succeq A^{\top}PA$, implying that $c^{t}P\succeq(A^{t})^{\top}PA^{t}\succeq(A^{t})^{\top}A^{t}$ via repeated applications of the previous inequality.
-Therefore, we have $\lim_{t\to\infty}(A^{t})^{\top}A^{t}=0$, and consequently $\lim_{t\to\infty}A^{t}=0$ since $\|A^{\top}\|\|A\|=\|A^{\top}A\|$.
-For the other direction, given some matrix $A$ with $\rho(A)<1$, consider $P=\sum_{t=0}^{\infty}(A^{t})^{\top}A^{t}$.
-Since $I+A^{\top}PA, PP$ satisfies the required conclusions.
-The fact that $P$ as defined exists, that is, that it is finite, as long as $\rho(A)<1$ can be verified via the Jordan normal form of $A$, which in particular asserts that $\|A^{t}\|\leq O(t^{d_{x}}\rho(A)^{t})$.
-∎
-Using this observation, one can write a semidefinite mathematical program that checks if a given system is stabilizable, and if so produces a matrix $K$ that guarantees a spectral radius smaller than one.
-Theorem 4.4.
-There exists a matrix K{K} such that $\rho({A}+{B}{K})<1$ if and only if there exist matrices ${P}\in\mathcal{S}^{d_{x}\times d_{x}}$ and ${L}\in{\mathbb{R}}^{d_{u}\times d_{x}}$ such that
-$\begin{bmatrix}{P}&({A}{P}+{B}{L})^{\top}\\ {A}{P}+{B}{L}&{P}\end{bmatrix}\succ 0,\qquad{P}\succeq{I}.$ (4.1)
-Furthermore, given such matrices P{P} and L{L} satisfying Equation 4.1, the matrix ${K}^{\prime}={L}{P}^{-1}$ satisfies $\rho({A}+{B}{K}^{\prime})<1$.
-Proof.
-We begin by observing the following result, using the Schur complement, regarding the semi-definiteness of block matrices.
-Lemma 4.5.
-Define a block matrix M{M} as
-${M}=\begin{bmatrix}{A}&{B}\\ {B}^{\top}&{C}\end{bmatrix}$. (4.2)
-Then, ${M}\succ{0}$ if and only if ${A}\succ{0}$ and ${C}-{B}^{\top}{A}^{-1}{B}\succ{0}$.
-Note that since the spectral radius purely involves the eigenvalues of a matrix, it is invariant under transposition, and we have $\rho((A+BK)^{\top})=\rho(A+BK)$.
-Now using this fact and Theorem 4.3, we know that the stabilizability of a linear system is equivalent to the existence of matrices K{K} and P{P} such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)P(A+BK)^{\top}}$.
-The above lemma allows us to reformulate this cubic constraint to a quadratic constraint.
-The stabilizability of a system is equivalent to existence of matrices K{K} and P{P} such that
-$\begin{bmatrix}{P}&{((A+BK)P)^{\top}}\\ {(A+BK)P}&{P}\end{bmatrix}\succ{0},\qquad{P}\succeq{I}.$
-Lastly, we reparameterize this constraint as a linear semidefinite constraint as demonstrated below, by choosing L=KP{L}={KP}.
-$\begin{bmatrix}{P}&{(AP+BL)^{\top}}\\ {AP+BL}&{P}\end{bmatrix}\succ{0},\qquad{P}\succeq{I}.$
-Notice that this is indeed a reformulation of the mathematical feasibility program, and not a relaxation, since the reparamertization ${K}={LP^{-1}}$ is invertible. ∎
+Suppose first that $\rho(A+BK)<1$. By the discrete-time Lyapunov theorem, there exists $P\succ 0$ such that
+$(A+BK)P(A+BK)^{\top}\prec P.$
+After scaling, we may assume $P\succeq I$. By the Schur complement, this is equivalent to
+$\begin{bmatrix}P&((A+BK)P)^{\top}\\ (A+BK)P&P\end{bmatrix}\succ 0.$
+Setting $L=KP$ gives the displayed condition.
+Conversely, suppose $P$ and $L$ satisfy the displayed condition, and define $K=LP^{-1}$. Another application of the Schur complement yields
+$(A+BK)P(A+BK)^{\top}\prec P.$
+Hence $\rho(A+BK)<1$ by the discrete-time Lyapunov theorem. ∎
+Thus stabilizability can be checked efficiently via semidefinite feasibility, and any feasible solution directly produces a stabilizing controller.
 4.3 Controllability of Linear Dynamical Systems
-Recall that a system is controllable if and only if in the absence of noise and for every target state there is a sequence of controls that drives the system state to that target.
+Recall that a system is controllable if and only if, in the absence of noise and for every target state, there is a sequence of controls that drives the system state to that target.
 Linear dynamical systems exhibit a particularly elegant characterization of controllability.
-First, we need to define an object of central importance for linear dynamical systems, the Kalman matrix.
+First, we define an object of central importance for linear dynamical systems: the Kalman controllability matrix.
 Definition 4.4.
@@
 Proof.
-Consider the evolution of the noiseless linear dynamical system starting from the initial state $\mathbf{x}_{0}$. The state at time $T$ is given by the expression
+Consider the evolution of the noiseless linear dynamical system starting from the initial state $\mathbf{x}_{0}$. The state at time $T$ is given by
 $\mathbf{x}_{T}$ $={A}\mathbf{x}_{T-1}+{B}\mathbf{u}_{T-1}$
@@
 $\mathbf{x}_{T}-{A}^{T}\mathbf{x}_{0}=[{B},{A}{B},{A^{2}}{B},\ldots,{A^{T-1}}{B}]\begin{bmatrix}\mathbf{u}_{T-1}\\ \vdots\\ \mathbf{u}_{0}\end{bmatrix}$.
-By the fundamental theory of linear algebra, this equation has a solution if and only if ${K}_{T}({A,B})$ has full rank.
-That is, for every initial state $\mathbf{x}_{0}$ and final state $\mathbf{x}_{T}$, there exists a set of control vectors $\{\mathbf{u}_{t}\}$ that satisfy the equation if and only if ${K}_{T}({A,B})$ has full rank.
-The rank of the Kalman matrix is upper bounded by the number of rows of ${K}_{T}$, which is $d_{x}$.
-Thus, the system is controllable if and only if the rank of ${K}_{T}$ is $d_{x}$ for all $T\geq d_{x}$.
-Denote by $r_{t}=\mathop{\mbox{\rm rank}}({K}_{t})$ the rank of the Kalman matrix after $t$ iterations.
-Since the rank can never decrease with $t$, it remains to show that the system is controllable if and only if $r_{t}=r_{d_{x}}$ for all $t\geq d_{x}$.
-We claim that once $r_{t}=r_{t+1}$, the rank never increases again by increasing $t$. Assuming $r_{t}=r_{t+1}$, we have
-$\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t}B}]=\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t-1}B}].$
-Since linear transformations retain containment, we can multiply by A{A} and get
-$\mathop{\mbox{\rm span}}[{AB,\ldots,A^{t+1}B}]\subseteq\mathop{\mbox{\rm span}}[{AB,\ldots,A^{t}B}].$
-Adding B{B} to both matrices does not affect the subset relationship.
-$\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t+1}B}]\subseteq\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t}B}]$
-Therefore, $r_{t+2}\leq r_{t+1}$.
-Since the rank cannot decrease through the addition of columns, we have $r_{t+2}=r_{t+1}$.
-Since the rank can grow at most $d_{x}$ times, equation $r_{d_{x}}=r_{T}$ holds for all $T>d_{x}$ if and only if the system is controllable, and this concludes the theorem.
-∎
+Therefore, for every initial state $\mathbf{x}_{0}$ and target state $\mathbf{x}_{T}$, there exists a control sequence $\mathbf{u}_{0},\ldots,\mathbf{u}_{T-1}$ that satisfies this equation if and only if the column span of ${K}_{T}(A,B)$ is all of ${\mathbb{R}}^{d_{x}}$, i.e., if and only if ${K}_{T}(A,B)$ has rank $d_{x}$.
+Denote by $r_{t}=\mathop{\mbox{\rm rank}}({K}_{t})$ the rank of the Kalman matrix after $t$ steps.
+Since adding columns cannot decrease rank, the sequence $(r_{t})_{t\geq 1}$ is nondecreasing.
+We claim that once $r_{t}=r_{t+1}$, the rank can never increase again.
+Indeed, if $r_{t}=r_{t+1}$, then
+$A^{t}B\in\mathop{\mbox{\rm span}}[B,AB,\ldots,A^{t-1}B].$
+Multiplying by $A$ gives
+$A^{t+1}B\in\mathop{\mbox{\rm span}}[AB,A^{2}B,\ldots,A^{t}B]\subseteq\mathop{\mbox{\rm span}}[B,AB,\ldots,A^{t}B].$
+Hence adding the next block $A^{t+1}B$ does not increase the rank, so $r_{t+2}=r_{t+1}$. By induction, once the rank stops increasing, it never increases again.
+Since the rank can increase at most $d_{x}$ times, it follows that $r_{t}=r_{d_{x}}$ for all $t\geq d_{x}$. Therefore, the system is controllable if and only if
+$\mathop{\mbox{\rm rank}}K_{d_{x}}(A,B)=d_{x}.$
 4.4 Quantitative Definitions
-In the rest of this text our goal is to give efficient algorithms with finite time performance bounds for online control.
-These finite-time bounds naturally depend on how difficult it is for a given system to stabilize.
-Therefore, we require a quantitative definition for stabilizability, one that holds for both time-invariant and time-varying linear dynamical systems.
+In the rest of this text our goal is to give efficient algorithms with finite-time performance bounds for online control.
+These bounds naturally depend on how difficult it is to stabilize and control a given system.
+Since the main focus of the sequel is on linear time-invariant systems, we now introduce quantitative notions of stabilizability and controllability for that setting.
 4.4.1 Stabilizability of Time-Invariant Systems
-Consider a linear time-invariant system given by the transformations (A,B)({A,B}).
-We have already remarked that this system is called stabilizable, according to the classical definition, if and only if there exists $K$ such that $\rho({A+BK})<1$ holds.
+Consider a linear time-invariant system given by the pair (A,B)({A},{B}).
+We have already remarked that this system is classically stabilizable if and only if there exists a matrix $K$ such that $\rho(A+BK)<1$.
 A stronger quantitative property can be obtained as follows, and in fact it is implied by BIBO stabilizability as defined in Definition 2.4. The proof of this fact is left as Exercise 4.6.
 Lemma 4.6.
-Consider a linear controller K{K} for the LTI system (A,B)({A,B}).
-If there exist $\kappa>0$ and $\delta\in(0,1]$ such that the matrix A+BK{A+BK} admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
-$\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t},\qquad\frac{\kappa}{\delta}\leq\gamma$
+Consider a linear controller K{K} for the LTI system (A,B)({A},{B}). If there exist $\kappa>0$ and $\delta\in(0,1]$ such that
+$\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}\qquad\text{for all }t\geq 0,$
+and moreover $\kappa/\delta\leq\gamma$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
 Proof.
-Let $\tilde{K}=A+BK$}. By unrolling the dynamics equation we have that
-$\mathbf{x}_{t+1}$ $={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t}=\tilde{K}\mathbf{x}_{t}+\mathbf{w}_{t},$
-$=\sum_{i=0}^{t-1}\tilde{{K}}^{i}\mathbf{w}_{t-i}+\tilde{{K}}^{t+1}\mathbf{x}_{0}.$
-Thus, as long as $\|\mathbf{w}_{t}\|,\|\mathbf{x}_{0}\|\leq 1$, we have
-$\|\mathbf{x}_{t+1}\|$ $\leq\sum_{i=0}^{t}\|\tilde{K}^{i}\|\|\mathbf{w}_{t-i}\|+\|\tilde{K}^{t+1}\|\|\mathbf{x}_{0}\|,$
-$\leq\kappa\sum_{i=0}^{\infty}(1-\delta)^{i}\leq\frac{\kappa}{\delta}\leq\gamma.$
+Let $\tilde{K}=A+BK$}. Under the linear controller $u_{t}=Kx_{t}$, the dynamics are
+$\mathbf{x}_{t+1}=\tilde{K}\mathbf{x}_{t}+\mathbf{w}_{t}.$
+Unrolling the recursion gives
+$\mathbf{x}_{t+1}=\tilde{K}^{t+1}\mathbf{x}_{0}+\sum_{s=0}^{t}\tilde{K}^{t-s}\mathbf{w}_{s}.$
+Define an augmented disturbance sequence by $v_{-1}=\mathbf{x}_{0}$ and $v_{s}=\mathbf{w}_{s}$ for $s\geq 0$. Then
+$\mathbf{x}_{t+1}=\sum_{s=-1}^{t}\tilde{K}^{\,t-s}v_{s}.$
+Therefore, whenever $\|\mathbf{x}_{0}\|\leq 1$ and $\|\mathbf{w}_{s}\|\leq 1$ for all $s$,
+$\|\mathbf{x}_{t+1}\|$ $\leq\sum_{s=-1}^{t}\left\|\tilde{K}^{\,t-s}\right\|\|v_{s}\|$
+$\leq\kappa\sum_{i=0}^{t+1}(1-\delta)^{i}$
+$\leq\frac{\kappa}{\delta}\leq\gamma.$
+Thus $K$ is $\gamma$-stabilizing. ∎
 A linear time-invariant system is thus $\gamma$-stabilizable if we can find a matrix $K$ with the above properties.
-4.4.2 Stabilizability of Time-Varying Systems
-The $\gamma$-stabilizability condition, as opposed to classical stabilizability, is particularly important to generalize to time-varying systems, since the spectral radius of the product is not bounded by the product of spectral radii, see exercise 4.6 at the end of this chapter.
-We now define a sufficient condition for $\gamma$-stabilizability. Consider a time-varying linear dynamical system given by the transformations $\{{A}_{t},{B}_{t}\}$.
-Lemma 4.9.
-Consider a sequence of linear controllers ${K}_{1:T}$ for the time-varying linear dynamical system $\{{A}_{1:T},{B}_{1:T}\}$.
-If the matrices $\tilde{K}_{t}=A_{t}+B_{t}K_{t}$} satisfy the following inequality, then this linear controller sequence is $\gamma$-stabilizing.
-$\forall t_{1},t_{2}\ .\ \left\|\prod_{t=t_{1}}^{t_{2}}\tilde{K}_{t}\right\|\leq\kappa(1-\delta)^{t_{1}-t_{2}},\ \frac{\kappa}{\delta}\leq\gamma.$
-Proof.
-Notice that under the condition of the lemma, by unrolling the dynamics equation we have that
-$\|\mathbf{x}_{t+1}\|$ $=\|{A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}\|=\left\|\tilde{{K}}_{t}\mathbf{x}_{t}+\mathbf{w}_{t}\right\|,$
-$=\left\|\sum_{i=0}^{t-1}(\prod_{j=1}^{i}\tilde{{K}}_{t-j})\mathbf{w}_{t-i}+(\prod_{j=1}^{t}\tilde{{K}}_{t-j})\mathbf{x}_{0}\right\|,$
-$\leq\kappa\sum_{i=0}^{\infty}e^{-\delta i}\leq\frac{\kappa}{\delta}\leq\gamma.$
 4.4.2 Controllability of Time-Invariant Systems
@@
 Clearly, such a system is nominally controllable, as is indeed any one-dimensional system with nonzero $B$.
-However, reaching a state $x=1$ from an initial $x_{0}=0$ in a constant number of time steps requires a large control input $u_{t}$ of magnitude at least $1/\varepsilon$.
+However, reaching a state $x=1$ from an initial state $x_{0}=0$ in a constant number of time steps requires a large control input $u_{t}$ of magnitude at least $1/\varepsilon$.
 Thus, such a system may be said to be barely controllable.
-This difficulty of controllability is captured in the definition below, which requires not only that $K(A,B)$ is a full row-rank, but also that such a statement is robust to small changes in $A,B$.
+This difficulty is captured by the definition below, which requires not only that $K_{k}(A,B)$ has full row rank, but also that the corresponding Gram matrix be well conditioned.
 Definition 4.7.
 A time-invariant linear dynamical system $(A,B)$ is $(k,\kappa)$-strongly controllable if and only if
-$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\text{ and }\|({K}_{k}(A,B){K}_{k}(A,B)^{\top})^{-1}\|\leq\kappa.$
-In the above definition, in addition to the full row-rank constraint, the spectral norm of the inverse of the Gram matrix of $K(A,B)$ is bounded by $\kappa$.
-For example, in the simple one-dimensional system we presented above, with $k=1$, the matrix $K(A,B)$ is $\varepsilon$, and thus $\kappa=\frac{1}{\varepsilon^{2}}$, which indeed captures the complexity of controlling the system.
+$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\qquad\left\|({K}_{k}(A,B){K}_{k}(A,B)^{\top})^{-1}\right\|\leq\kappa.$
+In the above definition, in addition to the full-row-rank constraint, the spectral norm of the inverse of the Gram matrix of $K_{k}(A,B)$ is bounded by $\kappa$.
+For example, in the simple one-dimensional system above with $k=1$, the matrix $K_{1}(A,B)$ is just $\varepsilon$, and thus $\kappa=1/\varepsilon^{2}$, which indeed captures the complexity of controlling the system.
 4.5 Bibliographic Remarks
-The idea of reducing a nonlinear dynamical system to a time varying linear dynamical system had been independently discovered in disparate fields.
-In the context of control, using the linearization to characterize local stability properties of non-linear systems was pioneered by Lyapunov [Lya92], and is called the Lyapunov direct method; see, e.g.
-[SL91].
-The asymptotic notions of stability, stabilizability and controllability discussed in this chapter constitute a core part of theory of linear control, and are found in most texts on the subject (e.g.
+The idea of reducing a nonlinear dynamical system to a time-varying linear dynamical system has appeared independently in several fields.
+In the context of control, using linearization to characterize local stability properties of nonlinear systems was pioneered by Lyapunov [LYA92], and is called the Lyapunov direct method; see, e.g., [SL91].
+The asymptotic notions of stability, stabilizability, and controllability discussed in this chapter constitute a core part of linear control theory, and are found in most texts on the subject (e.g.
 [STE94, HES18]).
-The efficient verification of stabilizability via a semidefinite program and the recovery of a stable controller form the solution of the latter appears in [Ahm16].
-On the other hand, it is known that even imposition of norm constraints on a plausibly stable controller makes this problem computationally hard [BT97, BT00, BT99].
-Somewhat distinct but related problems of stabilizability with even a 1-dimensional linear constraint are open to the date.
+The efficient verification of stabilizability via a semidefinite program, together with the recovery of a stabilizing controller from a solution to that program, appears in [AHM16].
+On the other hand, it is known that even imposing norm constraints on a plausibly stable controller makes this problem computationally hard [BT97, BT00, BT99].
+Somewhat distinct but related problems of stabilizability with even a 1-dimensional linear constraint remain open to this day.
 The notion of $\gamma$-stabilizability has a precursor in the notion of strong stability introduced in [CHK+18].
 An extension of this strong stability concept to time-varying linear dynamics was established in [CKM19].
-The definition presented in this chapter has weaker requirements than those in the listed articles, allowing us to work with and guarantee regret later with respect to a larger class of comparator policies.
+The definition presented in this chapter has weaker requirements than those in the cited articles, allowing us to work with and guarantee regret later with respect to a larger class of comparator policies.
 4.6 Exercises
@@
 Hint: this exercise requires the optimal solution to LQR in the next chapter.
-4.2.
-Prove that $\rho(A)<1$ if and only if $\lim_{t\mapsto\infty}A^{t}=0$.
-Furthermore, if $\rho(A)>1$, then for any norm, $\lim_{t\mapsto\infty}\|A^{t}\|=\infty$.
-Finally, prove that $\rho(A)=\lim_{t\to\infty}\|A^{t}\|^{1/t}$.
+4.2. Prove that
+$\rho(A)<1\quad\text{if and only if}\quad\lim_{t\to\infty}A^{t}=0.$
+Furthermore, if $\rho(A)>1$, then for any norm,
+$\lim_{t\to\infty}\|A^{t}\|=\infty.$
+Finally, prove that
+$\rho(A)=\lim_{t\to\infty}\|A^{t}\|^{1/t}.$
 4.3. Prove that the spectral radius of a product of matrices is not necessarily bounded by the product of the spectral radii of the individual matrices. That is, show that in general,
-$\rho(\prod_{i=1}^{n}A_{i})\nleq\prod_{i=1}^{n}\rho(A_{i}).$
-4.4. Prove that a stochastic linear dynamical system with Gaussian perturbations
+$\rho\!\left(\prod_{i=1}^{n}A_{i}\right)\nleq\prod_{i=1}^{n}\rho(A_{i}).$
+4.4. Assume $\rho(A)<1$. Prove that the stochastic linear dynamical system with Gaussian perturbations
 $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+\mathbf{w}_{t},\qquad\mathbf{w}_{t}\sim\mathcal{N}(0,I),$
 constitutes a reversible Markov chain if and only if $A$ is symmetric, i.e. $A=A^{\top}$.
-Hint: One way to do this exercise is to explicitly calculate the stationary distribution of an LDS and then to verify reversibility as stated in its definition.
-Exercise 3.6 provides an alternative without the need to derive the stationary distribution.
+Hint: One way to do this exercise is to explicitly calculate the stationary distribution of the chain and then verify detailed balance directly.
 4.5.
@@
 Although we have already seen sufficient conditions to ensure BIBO stability, we ask for the necessity of such conditions.
-A. For a time-invariant linear system $(A,B)$ equipped with a linear policy $K$, prove that BIBO stability indeed implies $\rho(A+BK)<1$. In fact, comment on how large $\rho(A+BK)$ can be.
-B.
-For a time-invariant BIBO-stable linear system, prove that there exist $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$, $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}$.
+A. For a time-invariant linear system $(A,B)$ equipped with a linear policy $K$, prove that BIBO stability implies $\rho(A+BK)<1$. In fact, comment on how large $\rho(A+BK)$ can be.
+B. For a time-invariant BIBO-stable linear system, prove that there exist $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$,
+$\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}.$
 In particular, this proves that the conditions listed in Lemma 4.6 are necessary, in addition to being sufficient.
-C. Construct a time-varying linear system that is BIBO stable but does not have the exponential decay properties required in Lemma 4.9.
-4.6. Prove that the bound of one on the operator norm of the matrix $A$ is a sufficient but not necessary condition for stabilizability.
+4.6. Prove that the strict condition $\|A\|<1$ on the operator norm of $A$ is sufficient, but not necessary, for stabilizability.
 Chapter 5 Optimal Control of Linear Dynamical Systems
-In the chapters thus far, we have studied more basic notions than control, namely solution concepts of dynamical systems such as stabilizability and controllability.
+In the chapters thus far, we have studied more basic notions than optimal control, namely structural properties of dynamical systems such as stabilizability and controllability.
 Even for these more basic problems, we have seen that they are computationally intractable in general.
 This motivated us to look into simpler dynamics, namely linear dynamical systems, for which we can at least answer basic questions about stabilizability and controllability.
-In this chapter, we consider how we can control linear dynamical systems optimally.
+In this chapter, we consider how to control linear dynamical systems optimally.
 The answer is positive in a strong sense.
 An archetypal problem in optimal control theory is the control of linear dynamical systems with quadratic cost functions.
-This setting is called the Linear Quadratic Regulator which we define and study in this chapter.
+This setting is called the Linear Quadratic Regulator, which we define and study in this chapter.
 The resulting efficient and practical algorithm is one of the most widely used subroutines in control.
@@
 The linear quadratic regulator (LQR) problem is that of optimal control of a linear dynamical system with known and fixed quadratic costs.
-For simplicity, we mostly consider the time-invariant version, where the linear dynamical system is fixed, and the costs are also fixed quadratic, as given formally below:
+For simplicity, we mostly consider the time-invariant version, where the dynamics and the costs do not vary with time.
 Definition 5.1.
-The (time-invariant) Linear Quadratic Regulator (LQR) input is a linear dynamical system that evolves according to
-$\mathbf{x}_{t+1}={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t},$
-where $\mathbf{w}_{t}\sim\mathcal{N}(0,\sigma^{2})$ is i.i.d. Gaussian, and convex quadratic cost functions are specified by matrices $Q$ and $R$.
-The goal is to generate a sequence of controls $\mathbf{u}_{t}$ that minimizes the expected cost, taken over the randomness of the perturbations, given by quadratic loss functions,
-$\min_{\mathbf{u}(\mathbf{x})}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R}\mathbf{u}_{t}\right].$
+The finite-horizon (time-invariant) Linear Quadratic Regulator (LQR) problem consists of a linear dynamical system evolving according to
+$\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t},\qquad t=1,\dots,T,$
+where the perturbations $\mathbf{w}_{t}\in{\mathbb{R}}^{d_{x}}$ are i.i.d. with
+$\mathbb{E}[\mathbf{w}_{t}]=0,\qquad\mathbb{E}[\mathbf{w}_{t}\mathbf{w}_{t}^{\top}]=\Sigma.$
+Given positive definite matrices $Q\in{\mathbb{R}}^{d_{x}\times d_{x}}$ and $R\in{\mathbb{R}}^{d_{u}\times d_{u}}$, the goal is to choose policies $\pi_{1},\dots,\pi_{T}$, where $\mathbf{u}_{t}=\pi_{t}(\mathbf{x}_{t})$, that minimize the expected quadratic cost
+$\min_{\pi_{1},\dots,\pi_{T}}\mathbb{E}\!\left[\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}R\mathbf{u}_{t}\right].$
 To remind the reader, the notation we use is as follows.
@@
   3. 3.
-    $\mathbf{w}_{t}\in\mathbb{R}^{d_{w}}$ is the perturbation sequence drawn i.i.d.
-from normal distributions, or any other zero-mean distribution as long as the draws are independent of states or controls.
+    $\mathbf{w}_{t}\in{\mathbb{R}}^{d_{x}}$ is the perturbation sequence, assumed i.i.d.
+with zero mean and covariance $\Sigma$.
+A Gaussian disturbance is a common special case, but Gaussianity is not essential for the derivation below.
   4. 4.
-    $A\in{\mathbb{R}}^{d_{x}\times d_{x}}$, $B\in{\mathbb{R}}^{d_{x}\times d_{u}}$ are the system matrices.
+    $A\in{\mathbb{R}}^{d_{x}\times d_{x}}$ and $B\in{\mathbb{R}}^{d_{x}\times d_{u}}$ are the system matrices.
   5. 5.
-    $Q\in{\mathbb{R}}^{d_{x}\times d_{x}}$, ${R}\in{\mathbb{R}}^{d_{u}\times d_{u}}$ are positive definite matrices. Thus, the cost functions are convex.
+    $Q\in{\mathbb{R}}^{d_{x}\times d_{x}}$ and $R\in{\mathbb{R}}^{d_{u}\times d_{u}}$ are positive definite matrices. Thus, the stage costs are strictly convex in the control.
 This framework is very useful, since many physical problems, such as the centrifugal governor and the ventilator control problem, can be approximated by a linear dynamical system.
 In addition, quadratic convex cost functions are natural in many physical applications.
-The LQR is a special case of a Markov decision process which we have described in Chapter 3, since the transitions are stochastic.
+The LQR is a special case of a Markov decision process, which we described in Chapter 3, since the transitions are stochastic.
 However, we cannot naively apply value iteration, policy iteration, or linear programming since there are infinitely many states and actions.
@@
 5.2 Optimal Solution of the LQR
-We know from Chapter 3 that the LQR is a special case of a finite-time Markov decision process and hence admits an optimal policy.
+We know from Chapter 3 that the LQR is a special case of a finite-horizon Markov decision process and hence admits an optimal policy.
 However, the optimal policy is in general a mapping from states to actions, both of which are infinite Euclidean spaces in our case.
-However, it can be shown that the optimal policy for an LQR is a linear policy, that is, a linear mapping from ${\mathbb{R}}^{d_{x}}$ to ${\mathbb{R}}^{d_{u}}$!
-This astonishing fact, together with a remarkable property of the value function, is formally given in the following theorem.
-We state the theorem for a finite number of iterations and for a linear time-invariant system.
+It turns out that the optimal policy for an LQR is a linear policy, that is, a linear mapping from ${\mathbb{R}}^{d_{x}}$ to ${\mathbb{R}}^{d_{u}}$.
+This remarkable fact, together with a corresponding quadratic form for the value function, is given in the following theorem.
+We state the theorem for a finite horizon and a linear time-invariant system.
 Generalizing this to a time-varying linear system is left as Exercise 5.6.
 Theorem 5.2.
-For the LQR problem with finite $T>0$ iterations, there exist positive semi-definite matrices $S_{t}$, ${K}_{t}$ and scalars $c_{t}$, such that the optimal value function and optimal policy satisfy:
-$\mathbf{v}_{t}^{*}(\mathbf{x})=\mathbf{x}^{\top}{S}_{t}\mathbf{x}+c_{t},\quad\mathbf{u}_{t}^{*}(\mathbf{x})={K}_{t}\mathbf{x}.$
+For the finite-horizon LQR problem, there exist symmetric matrices $S_{t}\succeq 0$, gains $K_{t}\in{\mathbb{R}}^{d_{u}\times d_{x}}$, and scalars $c_{t}\in{\mathbb{R}}$ such that, for every stage $t$,
+$\mathbf{v}_{t}^{*}(\mathbf{x})=\mathbf{x}^{\top}S_{t}\mathbf{x}+c_{t},\qquad\mathbf{u}_{t}^{*}(\mathbf{x})=K_{t}\mathbf{x}.$
 Proof.
-Recall that the Bellman Optimality equation for a general MDP is written as
-$\mathbf{v}_{t}^{*}(s)=\max_{a}\left\{R_{sa}+\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}_{t+1}^{*}(s^{\prime})\right\}.$
-We henceforth write the special case applicable to LQR, and prove the theorem by backward induction on $t$.
+For $t=1,\dots,T$, let $\mathbf{v}_{t}^{*}(\mathbf{x})$ denote the optimal cost-to-go from stage $t$ onward when $\mathbf{x}_{t}=\mathbf{x}$, and set $\mathbf{v}_{T+1}^{*}(\mathbf{x})\equiv 0$.
+We prove the claim by backward induction on $t$.
 Base Case: For $t=T$, we have
-$\mathbf{v}^{*}_{T}(\mathbf{x})=\min_{\mathbf{u}}\big\{\mathbf{x}^{\intercal}Q\mathbf{x}+\mathbf{u}^{\intercal}R\mathbf{u}\big\}.$
-Since ${R}\succ 0$, a minimum occurs at $\mathbf{u}=0$. Hence, ${S}_{T}={Q}\succ{0}$, ${K}_{T}={0}$ and $c_{T}=0$.
-Inductive Step: Assume for some $t>1$ that $\mathbf{u}_{t}^{*}(\mathbf{x})={K}_{t}\mathbf{x}$ and $\mathbf{v}_{t}^{*}(\mathbf{x})=\mathbf{x}^{\intercal}{S}_{t}\mathbf{x}+c_{t}$. Then we have
-$\mathbf{v}^{*}_{t-1}(\mathbf{x})$ $=\min_{\mathbf{u}}\big\{\mathbf{x}^{\intercal}{Q}\mathbf{x}+\mathbf{u}^{\intercal}{R}\mathbf{u}+\mathop{\mbox{\bf E}}_{\mathbf{w}}\mathbf{v}^{*}_{t}({A}\mathbf{x}+{B}\mathbf{u}+\mathbf{w})\big\}$
-$=\min_{\mathbf{u}}\big\{\mathbf{x}^{\intercal}{Q}\mathbf{x}+\mathbf{u}^{\intercal}{R}\mathbf{u}+\mathop{\mbox{\bf E}}_{\mathbf{w}}\left[({A}\mathbf{x}+{B}\mathbf{u}+\mathbf{w})^{\intercal}S_{t}({A}\mathbf{x}+{B}\mathbf{u}+\mathbf{w})\right]\big\}+c_{t}$
-$=\min_{\mathbf{u}}\big\{\mathbf{x}^{\intercal}{Q}\mathbf{x}+\mathbf{u}^{\intercal}{R}\mathbf{u}+({A}\mathbf{x}+{B}\mathbf{u})^{\intercal}{S}_{t}({A}\mathbf{x}+{B}\mathbf{u})+\mathop{\mbox{\bf E}}_{\mathbf{w}}[\mathbf{w}^{\top}S_{t}\mathbf{w}]\big\}+c_{t},$
-where in the last equality we used the fact that $\mathop{\mbox{\bf E}}[\mathbf{w}]=0$, and that the perturbation is independent of the state and control.
-We henceforth denote the second moment of the noise by $\mathop{\mbox{\bf E}}[\mathbf{w}\mathbf{w}^{\top}]=\sigma^{2}I$.
-This implies that for any matrix $M$ which is independent of $\mathbf{w}$, we have $\mathop{\mbox{\bf E}}[\mathbf{w}^{\top}M\mathbf{w}]=\mathop{\mbox{\bf E}}[M\mathbf{w}\mathbf{w}^{\top}]=\sigma^{2}\cdot M\bullet I=\sigma^{2}{\bf Tr}(M)$.
-Since $R$ and $S_{t}$ are positive semidefinite, the above expression is convex as a function of $\mathbf{u}$. At the global minimum the gradient vanishes, and hence
-$\nabla_{\mathbf{u}}\Big[{\mathbf{x}^{\intercal}Q\mathbf{x}+\mathbf{u}^{\intercal}R\mathbf{u}+(A\mathbf{x}+B\mathbf{u})^{\intercal}S_{t}(A\mathbf{x}+B\mathbf{u})+\sigma^{2}{\bf Tr}(S_{t})}]\Big]=0.$
-Since $Q,R,S_{t}$} are all positive semi-definite and symmetric, this gives us
-$2R\mathbf{u}+2B^{\intercal}S_{t}A\mathbf{x}+2B^{\intercal}S_{t}B\mathbf{u}=0\implies(R+B^{\intercal}S_{t}B)\mathbf{u}=-B^{\intercal}S_{t}A\mathbf{x}$}.
-Hence, we have that
-$\mathbf{u}_{t-1}(\mathbf{x})=K_{t-1}\mathbf{x},\ \ \ K_{t-1}=-(R+B^{\intercal}S_{t}B)^{-1}(B^{\intercal}S_{t}A)$}.
-Notice that the matrix $R+B^{\top}S_{t}B$ is full rank since $R\succ 0$. Plugging this in the expression for the value function, we get
-$\mathbf{v}_{t-1}^{*}(\mathbf{x})$
-$={\mathbf{x}^{\intercal}Q\mathbf{x}+\mathbf{x}^{\intercal}K_{t-1}^{\intercal}RK_{t-1}\mathbf{x}+\mathbf{x}^{\intercal}(A-BK_{t-1})^{\intercal}S_{t}(A-BK_{t-1})\mathbf{x}+\sigma^{2}{\bf Tr}(S_{t})}+c_{t}$
-$=\mathbf{x}^{\intercal}{S}_{t-1}\mathbf{x}+c_{t-1}$
-where $S_{t-1}=Q+K_{t-1}^{\intercal}RK_{t-1}+(A-BK_{t-1})^{\intercal}S_{t}(A-BK_{t-1})$} is positive definite, and $c_{t-1}=c_{t}+\sigma^{2}{\bf Tr}(S_{t})$.
+$\mathbf{v}_{T}^{*}(\mathbf{x})=\min_{\mathbf{u}}\left\{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}\right\}.$
+Since $R\succ 0$, the unique minimum occurs at $\mathbf{u}=0$. Hence
+$S_{T}=Q,\qquad K_{T}=0,\qquad c_{T}=0.$
+Inductive Step: Assume for some $t>1$ that
+$\mathbf{v}_{t}^{*}(\mathbf{x})=\mathbf{x}^{\top}S_{t}\mathbf{x}+c_{t}$
+for a symmetric matrix $S_{t}\succeq 0$. Then
+$\mathbf{v}_{t-1}^{*}(\mathbf{x})$ $=\min_{\mathbf{u}}\left\{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}+\mathbb{E}_{\mathbf{w}}\!\left[\mathbf{v}_{t}^{*}(A\mathbf{x}+B\mathbf{u}+\mathbf{w})\right]\right\}$
+$=\min_{\mathbf{u}}\left\{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}+\mathbb{E}_{\mathbf{w}}\!\left[(A\mathbf{x}+B\mathbf{u}+\mathbf{w})^{\top}S_{t}(A\mathbf{x}+B\mathbf{u}+\mathbf{w})\right]\right\}+c_{t}.$
+Using $\mathbb{E}[\mathbf{w}]=0$ and $\mathbb{E}[\mathbf{w}\mathbf{w}^{\top}]=\Sigma$, we obtain
+$\mathbb{E}_{\mathbf{w}}\!\left[(A\mathbf{x}+B\mathbf{u}+\mathbf{w})^{\top}S_{t}(A\mathbf{x}+B\mathbf{u}+\mathbf{w})\right]=(A\mathbf{x}+B\mathbf{u})^{\top}S_{t}(A\mathbf{x}+B\mathbf{u})+{\bf Tr}(S_{t}\Sigma).$
+Therefore,
+$\mathbf{v}_{t-1}^{*}(\mathbf{x})=\min_{\mathbf{u}}\left\{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}+(A\mathbf{x}+B\mathbf{u})^{\top}S_{t}(A\mathbf{x}+B\mathbf{u})\right\}+c_{t}+{\bf Tr}(S_{t}\Sigma).$
+Since $R\succ 0$ and $S_{t}\succeq 0$, the expression inside the braces is a strictly convex quadratic function of $\mathbf{u}$. Its gradient is
+$2(R+B^{\top}S_{t}B)\mathbf{u}+2B^{\top}S_{t}A\mathbf{x}.$
+Setting the gradient to zero yields
+$(R+B^{\top}S_{t}B)\mathbf{u}=-B^{\top}S_{t}A\mathbf{x}.$
+Since $R\succ 0$, the matrix $R+B^{\top}S_{t}B$ is positive definite and therefore invertible. Hence the optimal control is
+$\mathbf{u}_{t-1}^{*}(\mathbf{x})=K_{t-1}\mathbf{x},\qquad K_{t-1}=-(R+B^{\top}S_{t}B)^{-1}B^{\top}S_{t}A.$
+Substituting this back into the value function gives
+$\mathbf{v}_{t-1}^{*}(\mathbf{x})$ $=\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{x}^{\top}K_{t-1}^{\top}RK_{t-1}\mathbf{x}$
+$\qquad\qquad+\mathbf{x}^{\top}(A+BK_{t-1})^{\top}S_{t}(A+BK_{t-1})\mathbf{x}+c_{t}+{\bf Tr}(S_{t}\Sigma)$
+$=\mathbf{x}^{\top}S_{t-1}\mathbf{x}+c_{t-1},$
+where
+$S_{t-1}=Q+K_{t-1}^{\top}RK_{t-1}+(A+BK_{t-1})^{\top}S_{t}(A+BK_{t-1}),$
+and
+$c_{t-1}=c_{t}+{\bf Tr}(S_{t}\Sigma).$
+Since each term in the definition of $S_{t-1}$ is positive semidefinite, we have $S_{t-1}\succeq 0$. This completes the induction. ∎
 5.3 Infinite Horizon LQR
-Although infinite horizon problems can never appear in the real world, in practice the horizon length may be unknown.
-Although there are techniques to iteratively update the horizon length with geometrically increasing guesstimates, these are slow and impractical.
-An alternative is to solve the LQR for the infinite horizon case, which results in a very elegant solution.
-Not only that, this turns out to be a practical solution to which the finite-horizon solution converges quickly.
-The infinite horizon LQR problem can be written as follows:
-$\min_{\mathbf{u}(\mathbf{x})}\lim_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}{R}\mathbf{u}_{t}\right]$
-$\text{ s.t. }\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t},$
-where similar to the finite horizon case, we assume that the perturbations $\mathbf{w}_{t}$ are i.i.d.
-zero mean and have bounded variance.
-We henceforth assume that the perturbations are all zero for simplicity, and adding zero-mean bounded-variance perturbations is left as exercise 5.6.
-Without perturbations, the Bellman optimality equation becomes
+Although real control tasks always have finite duration, in practice the horizon length may be very large or unknown.
+This motivates the infinite-horizon formulation, which leads to an especially elegant stationary controller.
+In this section we consider the deterministic infinite-horizon total-cost problem
+$\min_{\pi}\quad$ $\sum_{t=1}^{\infty}\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}R\mathbf{u}_{t}$
+s.t. $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t},$
+where $\mathbf{u}_{t}=\pi(\mathbf{x}_{t})$. We assume throughout that $Q\succ 0$, $R\succ 0$, and that the pair $(A,B)$ is stabilizable.
+Persistent additive noise is more naturally treated with an average-cost formulation rather than a total-cost one; see Exercise 5.6.
+Under the above assumptions, the optimal policy is stationary and linear, and the value function has quadratic form. Writing the Bellman optimality equation for the total cost, we obtain
 $\mathbf{v}^{*}(\mathbf{x})=\min_{\mathbf{u}}\left\{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}+\mathbf{v}^{*}(A\mathbf{x}+B\mathbf{u})\right\}.$
-Assuming that the value function is quadratic, we have:
+Assuming that the value function is quadratic, say $\mathbf{v}^{*}(\mathbf{x})=\mathbf{x}^{\top}S\mathbf{x}$, we get
 $\mathbf{x}^{\top}S\mathbf{x}=\min_{\mathbf{u}}\left\{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}+(A\mathbf{x}+B\mathbf{u})^{\top}S(A\mathbf{x}+B\mathbf{u})\right\}.$
-Similarly to the finite-time case, we have the following result.
-Lemma 5.3.
-The optimal policy is given by a transformation $\mathbf{u}^{*}(\mathbf{x})={K}\mathbf{x}$.
-Proof.
-At optimality, the gradient with respect to $\mathbf{u}$ of the above equation is zero. This implies that
-$2R\mathbf{u}+2B^{\top}SB\mathbf{u}+2B^{\top}SA\mathbf{x}=0$},
-which implies
-$K=-(R+B^{\top}SB)^{-1}(B^{\top}SA)$}.
-Plugging this back into the Bellman equation, we obtain
+Since the objective is strictly convex in $\mathbf{u}$, the optimum is characterized by a vanishing gradient:
+$2R\mathbf{u}+2B^{\top}SB\mathbf{u}+2B^{\top}SA\mathbf{x}=0.$
+Hence the optimal policy is linear and given by
+$\mathbf{u}^{*}(\mathbf{x})=K\mathbf{x},\qquad K=-(R+B^{\top}SB)^{-1}B^{\top}SA.$
+Plugging this back into the Bellman equation yields
 $\mathbf{x}^{\top}S\mathbf{x}=\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{x}^{\top}K^{\top}RK\mathbf{x}+\mathbf{x}^{\top}(A+BK)^{\top}S(A+BK)\mathbf{x}.$
-Since this equality holds for all $\mathbf{x},$, we get
+Since this equality holds for all $\mathbf{x}$, we obtain
 $S=Q+K^{\top}RK+(A+BK)^{\top}S(A+BK).$
-Simplifying this expression we obtain
-$S=Q+A^{T}SA-A^{T}SB\left({R}+B^{T}SB\right)^{-1}B^{T}SA.$
-This equation is called the Discrete Time Algebraic Riccati Equation (DARE) and is one of the most important equations in control theory.
+Simplifying this expression gives
+$S=Q+A^{\top}SA-A^{\top}SB\left(R+B^{\top}SB\right)^{-1}B^{\top}SA.$
+This equation is called the Discrete-Time Algebraic Riccati Equation (DARE) and is one of the most important equations in control theory.
+Its stabilizing solution determines the optimal infinite-horizon controller.
 5.3.1 Solving the DARE
-In the one-dimensional case, although the right side of the equation is a rational function, we can multiply both sides by $R+B^{\top}SB$} to form a quadratic equation in S{S}.
-Solving the DARE therefore corresponds to finding a nonnegative root of a quadratic polynomial in the one-dimensional case.
-The reason we only consider the nonnegative root is that the value function is nonnegative by our choice of quadratic convex costs.
-In higher dimension, the correct generalization is somewhat more difficult to see, but using Lemma 4.5, the equation can be reformulated as the solution to the following semidefinite program.
-$\min_{{X}\in\mathcal{S}^{d_{x}\times d_{x}}}\langle{X,I}\rangle$
-$\text{subject to}\begin{bmatrix}{X}&B^{T}SA\\ A^{T}SB&{R}+B^{T}SB\end{bmatrix}\succeq 0,$
-$\text{and }S=Q+A^{T}SA-{X}.$
-A popular alternative is to adapt the value iteration algorithm to this setting by iteratively applying the right side of the aforementioned equation as an operator.
-Concretely, starting from ${S}_{0}\succeq 0$, we choose
-$S_{t+1}=Q+A^{T}S_{t}A-A^{T}S_{t}B\left({R}+B^{T}S_{t}B\right)^{-1}B^{T}S_{t}A.$
+In the one-dimensional case, although the right-hand side of the equation is a rational function, we can multiply both sides by $R+B^{\top}SB$ to form a quadratic equation in $S$.
+Solving the DARE therefore corresponds to finding a nonnegative root of a quadratic polynomial.
+We only consider the nonnegative root because the value function is nonnegative by our choice of quadratic convex costs.
+In higher dimensions, a convenient convex formulation is obtained from the associated Riccati inequality. By the Schur complement, the matrix inequality
+$\begin{bmatrix}Q+A^{\top}SA-S&A^{\top}SB\\ B^{\top}SA&R+B^{\top}SB\end{bmatrix}\succeq 0$
+is equivalent to
+$S\preceq Q+A^{\top}SA-A^{\top}SB\left(R+B^{\top}SB\right)^{-1}B^{\top}SA.$
+Under the standard stabilizability assumptions, the maximal positive semidefinite solution of this Riccati inequality is precisely the stabilizing solution of the DARE.
+Therefore, one can recover $S$ via the semidefinite program
+$\max_{S\in\mathcal{S}^{d_{x}\times d_{x}}}\langle S,I\rangle$
+subject to
+$\begin{bmatrix}Q+A^{\top}SA-S&A^{\top}SB\\ B^{\top}SA&R+B^{\top}SB\end{bmatrix}\succeq 0,\qquad S\succeq 0.$
+A popular alternative is the Riccati iteration: starting from some $S_{0}\succeq 0$, define
+$S_{t+1}=Q+A^{\top}S_{t}A-A^{\top}S_{t}B\left(R+B^{\top}S_{t}B\right)^{-1}B^{\top}S_{t}A.$
+Under the same assumptions, this sequence converges to the stabilizing solution of the DARE.
 5.4 Robust Control
@@
 is a strong one, and numerous attempts have been made over the years to generalize the noise model.
-One of the most important and prominent generalizations is that of robust control, also known as $\text{H}_{\infty}$-control.
-The goal of $\text{H}_{\infty}$-control is to come up with a more robust solution, by considering the optimal solution vs.
-any possible perturbation sequence from a certain family.
-Formally, our aim is to solve the following:
-$\begin{array}[]{c}\min_{\mathbf{u}_{1:T}(\mathbf{x})}\max_{\mathbf{w}_{1:T}}\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}{Q}\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}{R}\mathbf{u}_{t}\\ \text{ s.t.
-}\mathbf{x}_{t+1}=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t},\end{array}$
-where we use the exact same notation as in previous sections.
-Notice that this problem does not fall into the category of a Markov Decision Process, since $\mathbf{w}_{t}$’s are no longer stochastic.
-Therefore, the typical Bellman optimality criterion does not apply.
-Instead, the appropriate techniques to address this problem come from the theory of dynamical games.
-The above is, in fact, a two-player zero sum game.
-The first player tries to minimize the costs, and the other attempts to maximize the costs, both players subject to the dynamics evolution constraint.
-Using techniques from dynamic games, the following can be shown. References are provided in the bibliographic section.
+One of the most important and prominent generalizations is robust control, also known as $H_{\infty}$-control.
+The goal of $H_{\infty}$-control is to come up with a more robust solution by comparing the controller against an adversary that chooses the disturbance sequence.
+A standard finite-horizon formulation fixes a robustness parameter $\gamma>0$ and considers
+$\begin{array}[]{cl}\min_{\pi_{1},\dots,\pi_{T}}\max_{\psi_{1},\dots,\psi_{T}}&\displaystyle\sum_{t=1}^{T}\left(\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}R\mathbf{u}_{t}-\gamma^{2}\mathbf{w}_{t}^{\top}\mathbf{w}_{t}\right)\\[10.00002pt] \text{s.t.}&\mathbf{x}_{t+1}=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t},\\ &\mathbf{u}_{t}=\pi_{t}(\mathbf{x}_{t}),\qquad\mathbf{w}_{t}=\psi_{t}(\mathbf{x}_{t}).\end{array}$
+The parameter $\gamma$ controls the tradeoff between nominal performance and robustness.
+Notice that this problem does not fall into the category of a Markov decision process, since the disturbances are no longer stochastic.
+Instead, the appropriate techniques come from the theory of dynamic games.
+The above problem is a two-player zero-sum game: the controller tries to minimize the cost, while the adversary tries to maximize it, both subject to the same dynamical constraint.
+Using techniques from dynamic games, one can show the following. References are provided in the bibliographic section.
 Theorem 5.3.
-For the $\text{H}_{\infty}$ control problem with quadratic convex loss functions, the optimal solution is a linear policy of the state given by
-$\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}.$
+Under the standard solvability conditions for finite-horizon $H_{\infty}$ control, the above dynamic game admits a saddle point in linear state-feedback strategies.
+In particular, there exist matrices $K_{t}$ and $L_{t}$ such that
+$\mathbf{u}_{t}=K_{t}\mathbf{x}_{t},\qquad\mathbf{w}_{t}=L_{t}\mathbf{x}_{t}.$
 The proof is somewhat more involved than the LQR derivation and is omitted from this text.
@@
 The LQR problem and its solution were introduced in the seminal work of Kalman [KAL60], who also introduced the eponymous filtering procedure for linear estimation.
-While seemingly a natural corollary of equivalent results for MDPs, the existence of value functions for linear control is made subtle by the continuous nature of the state space and stabilizability, or the lack thereof.
-A discussion of sufficient and necessary conditions for the existence of solutions to Riccati equations can be found in [Ber07, HSK96].
-The development of $H_{\infty}$ control was driven by an observation made in the seminal paper of Doyle [Doy78] that, unlike in the fully observed case, the solution to the optimal control problem for partially observed linear systems (which we study in later chapters) is not naturally robust to misspecification of system parameters.
+While the finite-horizon derivation above follows the spirit of dynamic programming for Markov decision processes, the infinite-horizon theory is subtler because of the continuous nature of the state space and the role of stabilizability.
+A discussion of sufficient and necessary conditions for the existence of stabilizing solutions to Riccati equations can be found in [BER07, HSK96].
+The development of $H_{\infty}$ control was driven in part by the observation, made in the seminal paper of Doyle [DOY78], that optimal controllers need not be robust to misspecification of system parameters.
 This suggests that robustness specifications constitute yet another axis of performance for control algorithms and may at times be at odds with the objective of optimal control.
@@
 5.1. Prove Theorem 5.2 for a time-varying linear dynamical system.
-5.2. Derive the solution to the infinite-horizon LQR with zero mean Gaussian noise.
-5.3. Extend the proof of Theorem 5.2 to positive semidefinite $Q,R$, rather than positive definite matrices.
-5.4.
-Prove Theorem 5.4 as a dynamic two-player zero-sum game, where a unique saddle-point solution exists.
-Hint: you can formulate this problem as a min-max optimization problem within a finite horizon.
-5.5. Prove that the function $f(\mathbf{x})=\mathbf{x}^{\top}M\mathbf{x}$ is convex if and only if $M\succeq 0$ is positive semidefinite.
+5.2. Formulate the infinite-horizon LQR with i.i.d. zero-mean Gaussian noise as an average-cost control problem, and derive the corresponding Bellman equation and optimal linear policy.
+5.3. Extend the proof of Theorem 5.2 to the case $Q\succeq 0$ and $R\succ 0$.
+5.4. Prove Theorem 5.3 by formulating the robust control problem above as a finite-horizon dynamic two-player zero-sum game.
+5.5. Prove that the function $f(\mathbf{x})=\mathbf{x}^{\top}M\mathbf{x}$ is convex if and only if the symmetric part $\frac{1}{2}(M+M^{\top})$ is positive semidefinite.
 Part II Basics of Online Control
@@
 Definition 6.4 (Linear Policies).
-A linear controller $K\in{\mathbb{R}}^{d_{x}\times d_{u}}$ is a linear operator that maps the state to the control.
+A linear controller $K\in{\mathbb{R}}^{d_{u}\times d_{x}}$ is a linear operator that maps the state to the control.
 We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers that are $\gamma$-stabilizing according to definition 2.4 and are bounded in Frobenius norm by $\|K\|\leq\gamma$.
@@
 Lemma 6.6.
-Any DAC policy in $\Pi_{h,\gamma}^{D}$ is $\gamma^{2}$-stabilizing.
+Any DAC policy in $\Pi_{h,\gamma}^{D}$ is $2\gamma^{2}$-stabilizing.
 Proof.
-For convenience, denote by $M_{0}=I$, and the fictitious noise term $\tilde{\mathbf{w}}_{t}=\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$.
-First, due to the bound on $M_{1:h}$ and $\mathbf{w}_{t}$, we have that
-$\|\tilde{\mathbf{w}}_{t}\|=\|\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}\|\leq\sum_{i=0}^{h}\|M_{i}\|\|\mathbf{w}_{t-i}\|\leq\gamma,$
+For convenience, we define a fictitious noise term $\tilde{\mathbf{w}}_{t}=B\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i}$. First, due to the bound on $M_{1:h}$ and $\mathbf{w}_{t}$, we have that
+$\|\tilde{\mathbf{w}}_{t}\|=\left\|\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i}\right\|\leq\sum_{i=1}^{h}\|M_{i}\|\|\mathbf{w}_{t-i}\|\leq\gamma,$
 for $\|M\|$ being the spectral norm of $M$.
@@
 $\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}$
-$=A\mathbf{x}_{t}+B(K\mathbf{x}_{t}+\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i})+\mathbf{w}_{t}$
-$=(A+BK)\mathbf{x}_{t}+B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$ $\text{using }M_{0}=I$
-$=\sum_{j=0}^{t}\left((A+BK)^{j}B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i-j}\right)$
-$=\sum_{j=0}^{t-1}\left((A+BK)^{j}B\tilde{\mathbf{w}}_{t-j}\right).$
-We can thus view the states $x_{t}$ as generated by a $\gamma$-stabilizing policy over noises that are of magnitude $\gamma$, and hence bounded by $\gamma^{2}$. ∎
+$=A\mathbf{x}_{t}+B\left(K\mathbf{x}_{t}+\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i}\right)+\mathbf{w}_{t}$
+$=(A+BK)\mathbf{x}_{t}+B\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i}+\mathbf{w}_{t}$
+$=\sum_{j=0}^{t}\left((A+BK)^{j}\left(B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i-j}+\mathbf{w}_{t-j}\right)\right)$
+$=\sum_{j=0}^{t-1}\left((A+BK)^{j}(B\tilde{\mathbf{w}}_{t-j}+\mathbf{w}_{t-j})\right).$
+We can thus view the states $\mathbf{x}_{t}$ as generated by a $\gamma$-stabilizing policy over noises that are of magnitude $2\gamma$, and hence bounded by $2\gamma^{2}$. ∎
 The above lemma is already a good indication about DAC: they generate controls that at least do not cause a blow-up. But how expressive are these policies in terms of reaching the optimal solution?
@@
 $\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|$ $=\left\|\sum_{i=1}^{t}A^{i-1}B(\mathbf{u}_{t-i}-\tilde{\mathbf{u}}_{t-i})\right\|$
-$\leq\left\|\sum_{i=0}^{t}A^{i}B\right\|\max_{i\leq t}\|\mathbf{u}_{t}-\tilde{\mathbf{u}}_{i}\|\leq\frac{\kappa}{\delta}\varepsilon,$
+$\leq\left\|\sum_{i=0}^{t}A^{i}B\right\|\max_{i\leq t}\|\mathbf{u}_{i}-\tilde{\mathbf{u}}_{i}\|\leq\frac{\kappa}{\delta}\varepsilon,$
 by using the $\gamma$-stability of $(A,B)$. The approximation in terms of states and controls implies $\varepsilon$-approximation of policy class for any Lipschitz cost sequence, concluding the proof.
@@
 Regret minimization, online learning, and online convex optimization have a rich history and have been studied for decades in the machine learning and game theory literature.
-The reader is referred to the comprehensive texts of [CBL06, Haz16] for a detailed treatment.
+The reader is referred to the comprehensive texts of [CL06, HAZ16] for a detailed treatment.
 Linear state-feedback policies are a mainstay of control theory in both applications, due to their simplicity, and in theory, due to their optimality in deterministic and stochastic settings, as demonstrated in [KAL60].
@@
 An early example is the Youla parameterization, sometimes called Q-parameterization [YJB76].
-Later, the introduction of system-level synthesis (SLS) [DMW+17, ADLM19], in a similar vein, proved useful for distributed control, especially given its emphasis on state-space representations.
-For a comparison of these and additional ones such as input-output parameterizations [FZPK19], see [ZFP+20, ZZLL23].
-The Disturbance Action Control parameterization, introduced in [ABH+19], is similar to system-level synthesis.
+Later, the introduction of system-level synthesis (SLS) [DMW+17, ADL+19], in a similar vein, proved useful for distributed control, especially given its emphasis on state-space representations.
+For a comparison of these and additional ones such as input-output parameterizations [FZP+19], see [ZFP+20, ZZL+23].
+The Disturbance Action Control parameterization, introduced in [ABH+19b], is similar to system-level synthesis.
 Like SLS, it is a convex parameterization that is amenable to the tools of convex optimization.
@@
 This makes it better suited for machine learning.
-The study of regret in the context of control has roots in the work of [AYS11, AYBK14].
-These works, as well as many that followed [DMM+18, MTR19, CHK+18, AHS19, SF20], study regret in a stochastic setting, where the perturbations are stochastic.
+The study of regret in the context of control has roots in the work of [AS11, ABK14].
+These works, as well as many that followed [DMM+18, MTR19, CHK+18, AHS19c, SF20], study regret in a stochastic setting, where the perturbations are stochastic.
 In such settings, optimal control is well defined.
-The full power of regret analysis for control, involving both adversarial perturbations, adversarial loss functions, and computational considerations, was initiated in [ABH+19].
-The notion of regret as applied to policy classes was implicit in the work of [EDKM04].
+The full power of regret analysis for control, involving both adversarial perturbations, adversarial loss functions, and computational considerations, was initiated in [ABH+19b].
+The notion of regret as applied to policy classes was implicit in the work of [EKM04].
 The regret metric when applied to policies is different from the counterfactual notion of policy regret defined by [ADT12].
@@
 Online nonstochastic control is an extension of the framework of online convex optimization [HAZ16], as it is applied to policies, and thus to loss functions with memory [AHM13].
-The bulk of the material in this chapter is from [ABH+19].
-The work of [ABH+19] defined the general online nonstochastic setting in which the disturbances and the cost functions are adversarially chosen, and where the cost functions are arbitrary convex functions.
-The follow-up work by [AHS19] achieves a logarithmic pseudo-regret for strongly convex, adversarially selected losses, and well-conditioned stochastic noise.
+The bulk of the material in this chapter is from [ABH+19b].
+The work of [ABH+19b] defined the general online nonstochastic setting in which the disturbances and the cost functions are adversarially chosen, and where the cost functions are arbitrary convex functions.
+The follow-up work by [AHS19c] achieves a logarithmic pseudo-regret for strongly convex, adversarially selected losses, and well-conditioned stochastic noise.
 Logarithmic regret algorithms for the full setting of online nonstochastic control with strongly convex costs were obtained in [SIM20, FS20, CCK20].
 Online nonstochastic control with known dynamics was further extended in several directions.
-The task of control with bandit feedback for the costs was considered in [GHH20].
-More refined regret metrics applicable to changing dynamics were studied in [GHM20, MGSH21, ZWZ22, Sin22].
+The task of control with bandit feedback for the costs was considered in [GHH20a].
+More refined regret metrics applicable to changing dynamics were studied in [GHM20b, MGS+21, ZWZ22, SIN22].
 More extensions suitable for changing environments were studied in [BW22].
@@
 [LDL21] studies algorithms that balance regret minimization while adhering to constraints.
-An application of nonstochastic control in iterative planning in the presence of perturbations is described in [AHMS21].
-Various open-source software packages, such as [GHM20], provide benchmarked implementations of many such algorithms.
+An application of nonstochastic control in iterative planning in the presence of perturbations is described in [AHM+21].
+Various open-source software packages, such as [GHM20b], provide benchmarked implementations of many such algorithms.
 An alternate performance metric for online control, which is also nonstochastic, is the competitive ratio.
@@
 Control methods that minimize the competitive ratio were initiated in [GW19].
-This approach was extended in a series of papers [GLSW19, SLC+20, GH21].
+This approach was extended in a series of papers [GLS+19, SLC+20, GH21].
 Thus far, the success of this approach was limited to fully observed linear dynamical systems and further to quadratic costs when achieving the optimal competitive ratio.
-[GASH22] show that regret minimizers in the non-stochastic setting automatically achieve a near-optimal competitive ratio.
-A precursor to online nonstochastic control is a line of work, beginning with [AYBK14], focused on the control of stochastic systems under sequentially revealed adversarial costs functions.
+[GAS+22] show that regret minimizers in the non-stochastic setting automatically achieve a near-optimal competitive ratio.
+A precursor to online nonstochastic control is a line of work, beginning with [ABK14], focused on the control of stochastic systems under sequentially revealed adversarial costs functions.
 See [CK20, CCK22] for further developments.
@@
 [FIE97] which studies the discounted LQR problem under stochastic noise is the earliest example that provides a precise sample complexity bound.
-Similarly, the early work of [AYS11] uses an optimism-based approach to achieve optimal regret in terms of dependence on horizon.
+Similarly, the early work of [AS11] uses an optimism-based approach to achieve optimal regret in terms of dependence on horizon.
 This regret bound was later improved for sparse systems in [IJR12].
 [DMM+18, DMM+20] established regret and sample complexity results for the general stochastic LQR formulation with a polynomial dependence on natural system parameters.
-[FGKM18] studies the efficacy of first-order optimization-based approaches in synthesizing controllers for the LQR problem.
+[FGK+18] studies the efficacy of first-order optimization-based approaches in synthesizing controllers for the LQR problem.
 The work of [MTR19] establishes that simple explore-then-commit strategies achieve the optimal dependence of regret on horizon, with [SF20] proving that such approaches are optimal in terms of their dependence on dimension of the system, as well.
@@
 For the remainder of this chapter, we consider a partially observable linear time-invariant dynamical system described by the following equations.
-The extension of the material in this chapter to time-varying linear dynamical systems is left as exercise 9.7.
-$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}$
+$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t},$
 $\mathbf{y}_{t}$ $=C\mathbf{x}_{t}+\mathbf{v}_{t}.$
@@
 9.2 Online Learning of Dynamical Systems
-In an online learning setting, the learner sequentially observes the data and makes predictions in real time. Formally, the process unfolds as follows:
-  1. 1.
-    At each time step $t$, the learner observes the control input $\mathbf{u}_{t-1}$.
-  2. 2.
-    The learner makes a prediction $\hat{\mathbf{y}}_{t}$ for the next output.
-  3. 3.
-    The true output $\mathbf{y}_{t}$ is revealed.
-  4. 4.
-    The learner incurs a loss based on the difference between $\mathbf{y}_{t}$ and $\hat{\mathbf{y}}_{t}$, using a predefined loss function $\ell(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})$, such as squared error:
-    $\ell(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})=\|\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\|^{2}.$
-The goal is to minimize the total loss over time.
-However, in adversarial or unpredictable settings, direct loss minimization may be ill-posed.
-Instead, we consider a regret-based formulation, where the learner’s performance is measured relative to the best possible predictor from a predefined class of functions.
-More precisely, let $\Pi$ be a family of prediction rules mapping past observations to predictions. The regret of an online learning algorithm $A$ with respect to $\Pi$ is given by:
-$\text{Regret}_{T}(A,\Pi)=\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\Pi}\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}^{\pi}_{t}\|^{2}$
-where $\hat{\mathbf{y}}^{\pi}_{t}$ denotes the prediction made at time $t$ by the best function $\pi\in\Pi$ in hindsight.
-This regret-based framework allows us to compare different classes of predictors and design learning algorithms that efficiently adapt to unknown dynamical systems.
+In the problem of online learning in a linear dynamical system, sequentially for $t=1,\ldots,T$, the online learning algorithm:
+  • observes input $\mathbf{u}_{t-1}\in{\mathbb{R}}^{d_{u}}$,
+  • predicts output $\hat{\mathbf{y}}_{t}\in{\mathbb{R}}^{d_{y}}$,
+  • observes true system output $\mathbf{y}_{t}$,
+  • suffers loss according to a pre-specified loss function, $\ell_{t}(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})$, e.g.
+    $\ell_{t}(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})=\|\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\|^{2}.$
+The goal of the online learner is to minimize total loss over all game iterations.
+If the observations ${\mathbf{y}}_{t}$ are arbitrary, this is an ill-defined goal, and rather than minimizing loss, we consider a comparative analogue.
+More precisely, we consider a family of prediction rules from available information to prediction, and consider the metric of regret with respect to the best of these rules in hindsight.
+Formally, let $\Pi$ be a family of mappings from $\{\mathbf{u}_{\tau},\mathbf{y}_{\tau}\}_{\tau=1}^{t-1}$ to a prediction $\hat{\mathbf{y}}_{t}$, i.e.
+$\pi\in\Pi\ ,\ \pi(\mathbf{u}_{1:t-1},\mathbf{y}_{1:t-1})=\hat{\mathbf{y}}_{t}^{\pi}\in{\mathbb{R}}^{d_{y}}.$
+Let $\mathcal{A}$} be a prediction algorithm that iteratively produces predictions $\hat{\mathbf{y}}_{t}$. Then its regret with respect to $\Pi$ is defined as
+$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\sum_{t=1}^{T}\ell_{t}(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}\ell_{t}(\hat{\mathbf{y}}^{\pi}_{t},\mathbf{y}_{t}).$
+We proceed to define common prediction classes in the context of linear dynamical systems.
 9.3 Classes of Prediction Rules
@@
 Definition 9.1.
-We say that a class of predictors $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ if the following holds.
-For every sequence of $T$ inputs and outputs $\{\mathbf{u}_{1:T},\mathbf{y}_{1:T}\}$, and every $\pi_{2}\in\Pi_{2}$, there exists $\pi_{1}\in\Pi_{1}$ such that,
-$\sum_{t=1}^{T}\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|\leq T\varepsilon,$
-where $\mathbf{y}_{t}^{\pi_{1}}=\pi_{1}(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\ ,\ \mathbf{y}_{t}^{\pi_{2}}=\pi_{2}(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})$
+We say that a class of predictors $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ over horizon $T$ if the following holds.
+For every sequence of inputs and outputs $\{\mathbf{u}_{1:T},\mathbf{y}_{1:T}\}$ and every $\pi_{2}\in\Pi_{2}$, there exists $\pi_{1}\in\Pi_{1}$ such that
+$\sum_{t=1}^{T}\|\hat{\mathbf{y}}_{t}^{\pi_{1}}-\hat{\mathbf{y}}_{t}^{\pi_{2}}\|\leq T\varepsilon,$
+where
+$\hat{\mathbf{y}}_{t}^{\pi_{1}}=\pi_{1}(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\ ,\ \hat{\mathbf{y}}_{t}^{\pi_{2}}=\pi_{2}(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}).$
 The significance of this definition is that if a simple predictor class approximates another more complicated class, then it suffices to consider only the first.
@@
 Lemma 9.2.
-Let predictor class $\Pi_{1}$ $\varepsilon$-approximate class $\Pi_{2}$. Then for any prediction algorithm $\mathcal{A}$} we have
-$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{2})\leq\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{1})+\varepsilon T.$
+Assume that each loss function $\ell_{t}(\hat{\mathbf{y}},\mathbf{y}_{t})$ is $G$-Lipschitz in its first argument on the relevant prediction domain.
+If predictor class $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$, then for any prediction algorithm $\mathcal{A}$} we have
+$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{2})\leq\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{1})+G\varepsilon T.$
 The proof of this lemma is left as an exercise to the reader at the end of this chapter. We now proceed to consider specific prediction rules.
 9.3.1 Linear Predictors
-The simplest prediction rule is a linear function of observations, inputs, or both. More precisely,
-Definition 9.3 (Linear Policies).
-A linear controller parametrized by matrices $\{M_{1,i},M_{2,j}\}\in{\mathbb{R}}^{h\times d_{y}\times d_{y}+k\times d_{y}\times d_{u}}$ is a linear operator mapping state to control as
+The simplest prediction rule is a linear function of past observations and past inputs.
+Definition 9.3 (Linear Predictors).
+A linear predictor parameterized by matrices
+$M_{1,i}\in{\mathbb{R}}^{d_{y}\times d_{y}}\ \text{ for }\ i=1,\ldots,h,\qquad M_{2,j}\in{\mathbb{R}}^{d_{y}\times d_{u}}\ \text{ for }\ j=1,\ldots,k,$
+maps past outputs and inputs to a prediction according to
 $\hat{\mathbf{y}}_{t}=\sum_{i=1}^{h}M_{1,i}\mathbf{y}_{t-i}+\sum_{j=1}^{k}M_{2,j}\mathbf{u}_{t-j}.$
-We denote by $\Pi_{h,k,\kappa}^{L}$ the set of all linear controllers bounded as $\sum_{i=1}^{h}\|M_{1,i}\|+\sum_{j=1}^{k}\|M_{2,j}\|\leq\kappa$.
+We denote by $\Pi_{h,k,\kappa}^{L}$ the set of all such predictors satisfying
+$\sum_{i=1}^{h}\|M_{1,i}\|+\sum_{j=1}^{k}\|M_{2,j}\|\leq\kappa.$
+This class is computationally attractive because it is parameterized by a convex set of matrices.
+At first sight, however, it may appear too restrictive to model the predictions generated by a linear dynamical system with hidden state.
+The next subsection explains why this concern is misleading.
 9.3.2 Linear Dynamical Predictors
-An intuitive class of prediction rules that makes use of the special structure of linear dynamical systems is that of open-loop linear dynamical predictors, which we now define.
-The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $\mathbf{C}=I$, the identity matrix.
-In this situation, the optimal prediction can be calculated using the system evolution equation $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}$.
+An intuitive class of prediction rules that makes use of the special structure of linear dynamical systems is that of open-loop linear dynamical predictors.
+The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $C=I$, the identity matrix.
+In this situation, the optimal prediction can be calculated by unrolling the system evolution equation $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}$.
 It can be seen that this prediction rule is optimal for any zero-mean noise or perturbation model.
@@
 $\mathbf{y}_{t}^{\pi}=\pi_{\{A,B,C,\mathbf{x}_{0}\}}(t)$ $=C\mathbf{x}_{t}$
-$=CA\mathbf{x}_{t-1}+CB\mathbf{u}_{t-1}$
-$\vdots$
-$=\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t}\mathbf{x}_{0}.$
-Thus, the operator $\pi_{\{A,B,C,\mathbf{x}_{0}\}}$ takes a history of states and controls and gives the next output of the system accurately.
-We denote by $\prod^{\star}$ the set of all such predictors that in the case of zero noise give a perfectly accurate prediction.
-If the spectral radius of $A$ is larger than one, then the system is unstable and the magnitude of $\mathbf{y}_{t}$ can grow arbitrarily large, as discussed in section 4.2.
-Thus, for the rest of this section, assume that $A$ is $\gamma$-stable.
-With a decaying system, we can limit the memory of the predictor and still maintain accuracy. Thus, it makes sense to consider a more restricted class of predictors that have bounded memory, that is,
-$\pi_{\{A,B,C,h\}}=\sum_{i=1}^{h}CA^{i-1}B\mathbf{u}_{t-i},$
-where as a matter of convention we say $\mathbf{u}_{t}=0$ for all negative $t$. Denote by $\prod_{h}^{\star}$ the class of all such predictors.
-Clearly, for a large $h$ this class of predictors well approximates the true observations $\mathbf{y}_{t}$ if the system is activated without any noise. We next formalize what this means.
+$=CA^{t}\mathbf{x}_{0}+\sum_{i=1}^{t}CA^{i-1}B\mathbf{u}_{t-i}.$
+Thus, the operator $\pi_{\{A,B,C,\mathbf{x}_{0}\}}$ takes a history of inputs and gives the next output of the system accurately.
+We denote by $\Pi^{\star}$ the set of all such predictors that in the case of zero noise give a perfectly accurate prediction.
+The class $\Pi^{\star}$ is a very natural comparator class for the noiseless open-loop prediction problem.
+It is important to note, however, that in the general partially observed noisy setting the optimal predictor need not belong to this class, since it may depend on past outputs in a nontrivial way.
+We return to this more general question in the next chapter.
+For the present chapter, $\Pi^{\star}$ serves as a clean structural benchmark whose induced predictors we can approximate or represent by convex classes.
+9.4 Convex Relaxations of Linear Dynamical Predictors
+In the previous sections we have defined what it means to learn in a dynamical system, and described two natural classes of predictors.
+The class of linear dynamical predictors is most natural, as it exactly captures the noiseless evolution of a linear dynamical system.
+However, it is nonconvex in its parametrization.
+In this section we describe convex relaxation of this class that approximate it extremely well and allow for efficient learning.
+9.4.1 Convex Relaxation of Stable Systems
+The first relaxation we discuss is that which assumes additional structure from the dynamics, namely stability. For a noiseless system, define the truncated open-loop predictor
+$\hat{\mathbf{y}}_{t}=\pi_{\{A,B,C,h\}}(t)=\sum_{i=1}^{h}CA^{i-1}B\mathbf{u}_{t-i},$
+where by convention $\mathbf{u}_{s}=0$ for all $s\leq 0$. Denote by $\Pi_{h}^{\star}$ the class of all such truncated predictors.
+When the system is stable, this truncation yields a simple approximation to the class of open-loop linear dynamical predictors.
+To state this in the language of Definition 9.1, we restrict attention to a stable subclass with uniformly bounded parameters and zero initial state.
 Lemma 9.4.
-Suppose that $\|B\|+\|C\|\leq\gamma$, and $\|\mathbf{u}_{t}\|\leq U$ for all $t\in[T]$.
-Then the class $\prod_{h}^{\star}$ for $h=\gamma\log\frac{T\gamma^{3}U}{\varepsilon}$, $\varepsilon$-approximates the class $\prod^{\star}$.
+Fix $M\geq 1$, and define
+$\Pi^{\star}_{\mathrm{stab}}(M)\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\left\{\pi_{\{A,B,C\}}\in\Pi^{\star}:\|A^{t}\|\leq Me^{-t/M}\ \text{for all }t\geq 0,\quad\|B\|+\|C\|\leq M\right\}.$
+Assume that $\mathbf{x}_{0}=0$ and that
+$\max_{t\in[T]}\|\mathbf{u}_{t}\|\leq M.$
+Then $\Pi_{h}^{\star}$ $\varepsilon$-approximates $\Pi^{\star}_{\mathrm{stab}}(M)$ over horizon $T$ provided
+$h\geq M\log\!\left(\frac{2M^{5}}{\varepsilon}\right).$
 Proof.
-Consider any sequence and any predictor from $\prod^{\star}$, say $\pi_{1}=\pi_{A,B,C,\mathbf{x}_{0}}$.
-Consider the predictor $\pi_{2}=\pi_{A,B,C,h}\in\prod_{h}^{\star}$ for the same system matrices.
-Then we have
-$\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|$
-$=\left\|\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t-1}\mathbf{x}_{0}-\sum_{i=1:h}CA^{i-1}B\mathbf{u}_{t-i}\right\|$
-$=\left\|\sum_{i=h}^{T}CA^{i}B\mathbf{u}_{t-i-1}\right\|$
-$\leq\sum_{i=h}^{T}\kappa(1-\delta)^{i}\|C\|\|B\|\|\mathbf{u}_{t-i-1}\|$
-$\leq\frac{\gamma^{2}\kappa U}{\delta}(1-\delta)^{h}\leq\frac{\gamma^{2}\kappa U}{\delta}e^{-\delta h},$
-for some $\frac{\kappa}{\delta}\leq\gamma$. Using the value of $H$, and summing up over all $T$ iterations, we have
-$\sum_{t}\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|$ $\leq\sum_{t}\frac{\gamma^{2}\kappa U}{\delta}e^{-\delta h}=\varepsilon.$
+Fix any predictor $\pi_{\{A,B,C\}}\in\Pi^{\star}_{\mathrm{stab}}(M)$, and let $\pi_{\{A,B,C,h\}}\in\Pi_{h}^{\star}$ be its truncated predictor. For each $t\in[T]$,
+$\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C\}}}-\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C,h\}}}$ $=\sum_{i=h+1}^{t}CA^{i-1}B\mathbf{u}_{t-i}.$
+Hence
+$\left\|\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C\}}}-\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C,h\}}}\right\|$ $\leq\sum_{i=h+1}^{t}\|C\|\,\|A^{i-1}\|\,\|B\|\,\|\mathbf{u}_{t-i}\|$
+$\leq M^{4}\sum_{i=h+1}^{t}e^{-(i-1)/M}.$
+Since $M\geq 1$, we have $1-e^{-1/M}\geq\frac{1}{2M}$, and therefore
+$\sum_{i=h+1}^{t}e^{-(i-1)/M}\leq\sum_{i=h+1}^{\infty}e^{-(i-1)/M}\leq\frac{e^{-h/M}}{1-e^{-1/M}}\leq 2Me^{-h/M}.$
+Thus
+$\left\|\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C\}}}-\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C,h\}}}\right\|\leq 2M^{5}e^{-h/M}.$
+Summing over $t=1,\dots,T$ yields
+$\sum_{t=1}^{T}\left\|\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C\}}}-\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C,h\}}}\right\|\leq 2TM^{5}e^{-h/M}.$
+Choosing
+$h\geq M\log\!\left(\frac{2M^{5}}{\varepsilon}\right)$
+makes the right-hand side at most $T\varepsilon$, proving the claim. ∎
+9.4.2 Exact Representation via Cayley Hamilton
+In contrast to the previous approximation result, which is dimension-free but requires stability, the Cayley–Hamilton argument below gives an exact finite-memory representation for arbitrary noiseless finite-dimensional systems, at the price of dependence on the hidden-state dimension $d_{x}$.
+For a given matrix $A$ that describes the state evolution of a linear dynamical system with hidden-state dimension $d_{x}$, let
+$p_{A}(\lambda)=\lambda^{d_{x}}+a_{1}\lambda^{d_{x}-1}+\cdots+a_{d_{x}}$
+be the characteristic polynomial of $A$, and define $a_{0}=1$. The following theorem is the key structural observation of this chapter.
+Theorem 9.5.
+Consider the noiseless linear dynamical system
+$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t},$
+$\mathbf{y}_{t}$ $=C\mathbf{x}_{t},$
+with characteristic polynomial $p_{A}(\lambda)$. Then there exist matrices
+$M_{1,i}\in{\mathbb{R}}^{d_{y}\times d_{y}},\qquad M_{2,j}\in{\mathbb{R}}^{d_{y}\times d_{u}},\qquad i,j=1,\ldots,d_{x},$
+such that for every $t>d_{x}$,
+$\mathbf{y}_{t}=\sum_{i=1}^{d_{x}}M_{1,i}\mathbf{y}_{t-i}+\sum_{j=1}^{d_{x}}M_{2,j}\mathbf{u}_{t-j}.$
+In particular, one may take
+$M_{1,i}=-a_{i}I_{d_{y}},\qquad M_{2,j}=C\left(\sum_{i=0}^{j-1}a_{i}A^{j-1-i}\right)B.$
+Proof.
+Write $d=d_{x}$. By unrolling the state dynamics for $d$ steps, we obtain
+$\mathbf{x}_{t}=A^{d}\mathbf{x}_{t-d}+\sum_{j=1}^{d}A^{j-1}B\mathbf{u}_{t-j}.$ (9.1)
+Similarly, for each $i=1,\ldots,d$,
+$\mathbf{x}_{t-i}=A^{d-i}\mathbf{x}_{t-d}+\sum_{j=i+1}^{d}A^{j-i-1}B\mathbf{u}_{t-j},$
+so
+$A^{d-i}\mathbf{x}_{t-d}=\mathbf{x}_{t-i}-\sum_{j=i+1}^{d}A^{j-i-1}B\mathbf{u}_{t-j}.$
+By the Cayley–Hamilton theorem, $A^{d}+a_{1}A^{d-1}+\cdots+a_{d}I=0$, hence $A^{d}=-\sum_{i=1}^{d}a_{i}A^{d-i}$. Substituting this into (9.1), and then using the previous identity, gives
+$\mathbf{x}_{t}$ $=-\sum_{i=1}^{d}a_{i}A^{d-i}\mathbf{x}_{t-d}+\sum_{j=1}^{d}A^{j-1}B\mathbf{u}_{t-j}$
+$=-\sum_{i=1}^{d}a_{i}\mathbf{x}_{t-i}+\sum_{i=1}^{d}a_{i}\sum_{j=i+1}^{d}A^{j-i-1}B\mathbf{u}_{t-j}+\sum_{j=1}^{d}A^{j-1}B\mathbf{u}_{t-j}$
+$=-\sum_{i=1}^{d}a_{i}\mathbf{x}_{t-i}+\sum_{j=1}^{d}\left(A^{j-1}+\sum_{i=1}^{j-1}a_{i}A^{j-1-i}\right)B\mathbf{u}_{t-j}$
+$=-\sum_{i=1}^{d}a_{i}\mathbf{x}_{t-i}+\sum_{j=1}^{d}\left(\sum_{i=0}^{j-1}a_{i}A^{j-1-i}\right)B\mathbf{u}_{t-j}.$
+Multiplying both sides by $C$ yields
+$\mathbf{y}_{t}=-\sum_{i=1}^{d}a_{i}\mathbf{y}_{t-i}+\sum_{j=1}^{d}C\left(\sum_{i=0}^{j-1}a_{i}A^{j-1-i}\right)B\mathbf{u}_{t-j},$
+which is exactly the claimed recurrence. ∎
+The theorem shows that every noiseless $d_{x}$-dimensional linear dynamical system induces an exact linear predictor of memory at most $d_{x}$.
+Thus, if an upper bound on the hidden-state dimension is known, then it suffices to compete with a class of linear predictors of that memory length.
+Corollary 9.6.
+Let $\pi_{\{A,B,C,\mathbf{x}_{0}\}}\in\Pi^{\star}$ be any noiseless linear dynamical predictor with hidden-state dimension at most $d_{x}$, and let $\{\mathbf{y}_{t}\}_{t=1}^{T}$ be the output sequence it generates under an arbitrary input sequence $\{\mathbf{u}_{t}\}_{t=1}^{T}$.
+Then there exists $\kappa<\infty$ and a predictor $\pi\in\Pi^{L}_{d_{x},d_{x},\kappa}$ such that
+$\hat{\mathbf{y}}_{t}^{\pi}=\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C,\mathbf{x}_{0}\}}}=\mathbf{y}_{t}\qquad\text{for all }t>d_{x}.$
+Consequently, along every realizable trajectory of a noiseless $d_{x}$-dimensional linear dynamical system, the class $\Pi^{L}_{d_{x},d_{x},\kappa}$ exactly represents the induced linear dynamical predictor after a warm-up of $d_{x}$ rounds.
+Proof.
+Theorem 9.5 explicitly constructs matrices $M_{1,1:d_{x}}$ and $M_{2,1:d_{x}}$ that realize the predictor as a linear predictor of memory $d_{x}$ for all times $t>d_{x}$.
+Choosing $\kappa$ larger than the corresponding sum of matrix norms yields the claim.
+∎
+This exact finite-memory representation is one of the main reasons linear predictors are useful.
+It says that, in the noiseless setting, learning a linear predictor is not merely a heuristic surrogate for learning an unknown linear dynamical system: after an initial transient of length at most $d_{x}$, it is expressive enough to represent the predictor exactly.
+In the noisy setting, the same argument gives the same linear recurrence plus an additive disturbance term that depends on recent process and observation noises.
+Thus, linear predictors remain a natural benchmark even in the stochastic case, though they are no longer exact.
 9.5 Efficient Learning of Linear Predictors
 The previous section gives several natural classes of predictors as well as some of the relationships between them.
-In particular, the class of linear dynamical predictors is optimal for certain scenarios and requires knowledge of only three matrices.
-However, the parameterization of this class is nonconvex in the matrices $A,B,C$, which presents a computational challenge.
-This challenge is circumvented by using the fact that the class $\prod^{L}$ $\varepsilon$-approximates the class $\prod^{\star}$.
-In addition, the natural parameterization of $\prod^{L}$ is convex and thus can be learned using online convex optimization.
-The following algorithm learns the class of linear predictors in the sense that it attains a sublinear regret bound for any sequence of adversarial cost functions, as well as perturbations and observation noise.
-Algorithm 6 Learning $\prod^{L}$ by Online Gradient Descent
-1: Input: $M^{1}=M^{1}_{1,1:h},M^{1}_{2,1:k}$, convex constraints set $\mathcal{K}\subseteq{\mathbb{R}}^{h\times d_{y}+k\times d_{u}}$
+In particular, the class of linear dynamical predictors is natural but nonconvex in the matrices $A,B,C$.
+The Cayley–Hamilton theorem circumvents this challenge: every noiseless $d_{x}$-dimensional LDS induces a predictor in the convex class $\Pi^{L}_{d_{x},d_{x},\kappa}$ for an appropriate $\kappa$.
+The following algorithm learns the class of linear predictors in the sense that it attains a sublinear regret bound for any sequence of convex loss functions.
+Algorithm 6 Learning $\Pi^{L}$ by Online Gradient Descent
+1: Input: initial predictor matrices $M^{1}=(M^{1}_{1,1:h},M^{1}_{2,1:k})\in\mathcal{K}$, convex constraint set $\mathcal{K}$
 2:for $t=1$ to $T$ do
-3:  Predict $\hat{\mathbf{y}}_{t}(M^{t})=\sum_{i=1}^{h}M^{t}_{i}\mathbf{u}_{t-i}+\sum_{j=1}^{k}M^{t}_{j}\mathbf{y}_{t-j}$ and observe true $\mathbf{y}_{t}$.
-4:  Suffer loss $\ell_{t}(\hat{\mathbf{y}}_{t}(M^{t}),\mathbf{y}_{t})$, and update:
-$M^{t+1}=M^{t}-\eta_{t}\nabla\ell_{t}(M^{t})$
-$M^{t+1}=\mathop{\Pi}_{\mathcal{K}}(M^{t+1})$
+3:  Predict
+$\hat{\mathbf{y}}_{t}(M^{t})=\sum_{i=1}^{h}M^{t}_{1,i}\mathbf{y}_{t-i}+\sum_{j=1}^{k}M^{t}_{2,j}\mathbf{u}_{t-j},$
+and observe true $\mathbf{y}_{t}$.
+4:  Suffer loss $\ell_{t}(\hat{\mathbf{y}}_{t}(M^{t}),\mathbf{y}_{t})$ and update
+$M^{t+\frac{1}{2}}=M^{t}-\eta_{t}\nabla\ell_{t}(M^{t})$
+$M^{t+1}=\mathop{\Pi}_{\mathcal{K}}(M^{t+\frac{1}{2}})$
 5:end for
+Here $\ell_{t}(M)$ is shorthand for the loss incurred by predictor $M$ at time $t$, namely $\ell_{t}(\hat{\mathbf{y}}_{t}(M),\mathbf{y}_{t})$.
 This algorithm is an instance of the Online Gradient Descent algorithm, and Theorem 6.1 directly gives the following bound for it.
-We denote by $D,G$ the diameter of the set $\mathcal{K}$ as in Algorithm 6, and by $G$ an upper bound on the norm of the gradients of the loss functions $\ell_{t}$.
+We denote by $D$ the diameter of the set $\mathcal{K}$ and by $G$ an upper bound on the norm of the gradients of the loss functions $\ell_{t}(M)$ over $\mathcal{K}$.
 Corollary 9.7.
-For choice of step sizes $\eta_{t}=\frac{D}{G\sqrt{t}}$, and $\mathcal{A}$} being Algorithm 6, we have that
-$\mathrm{{regret}}({\mathcal{A}})=\sum_{t}\ell(\hat{\mathbf{y}}_{t}^{\mathcal{A}},\mathbf{y}_{t})-\min_{\pi\in\prod^{L}}\ell(\hat{\mathbf{y}}_{t}^{\pi},\mathbf{y}_{t})\leq 2GD\sqrt{T}.$
+For the choice of step sizes $\eta_{t}=\frac{D}{G\sqrt{t}}$, and $\mathcal{A}$} being Algorithm 6, we have
+$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi^{L}_{h,k,\kappa})\leq 2GD\sqrt{T}.$
+Combining Corollary 9.6 with Corollary 9.7, we conclude that if the observations are generated by an unknown noiseless linear dynamical system of hidden-state dimension at most $d_{x}$, then learning $\Pi^{L}_{d_{x},d_{x},\kappa}$ suffices to obtain sublinear regret with respect to the best linear dynamical predictor in hindsight, up to the lower-order cost of the first $d_{x}$ initialization rounds.
 9.6 Summary
@@
 We began by formulating the problem in terms of partially observable LDS, highlighting the challenges posed by unknown system matrices and noisy observations.
-We discussed key approaches for online learning in dynamical systems, emphasizing regret minimization as a performance metric.
-The regret framework allows direct comparisons between learning algorithms by evaluating their cumulative prediction error against the best fixed predictor in hindsight.
-A key takeaway from this chapter is that learning in LDS requires balancing expressivity and computational efficiency.
-Simple linear predictors may be computationally efficient but lack expressivity, whereas more sophisticated dynamical predictors require careful algorithmic design to remain tractable.
-The regret analysis provided insights into how online learning algorithms adapt to unknown system parameters while ensuring competitive performance.
-This discussion sets the stage for the next chapter on Kalman filtering, where we transition from open-loop predictors to closed-loop estimation techniques that integrate feedback from observations to refine state estimates.
+We then introduced several natural classes of predictors and explained how to compare them through regret minimization.
+The key technical observation of the chapter is that every noiseless finite-dimensional linear dynamical system induces an exact finite-memory linear predictor.
+This follows from the Cayley–Hamilton theorem, and it explains why simple linear predictors are far more expressive than they might initially appear.
+Rather than identifying the unknown system matrices explicitly, we can compete directly with the induced input-output predictor.
+Finally, we showed that linear predictors can be learned efficiently by online convex optimization, using Online Gradient Descent to obtain sublinear regret.
 9.7 Bibliographic Remarks
@@
 Resting at the crossroads between statistical estimation and dynamical systems, the theory of filtering had its early beginnings in the works of Norbert Wiener (see, e.g., [WIE49]).
-For a comprehensive treatment of online learning in games and online convex optimization, see [CBL06, Haz16].
+Learning linear predictors with finite history is a classical consequence of the Cayley–Hamilton theorem and realization theory; see, e.g., [KAI80, CHE99].
+For a comprehensive treatment of online learning in games and online convex optimization, see [CL06, HAZ16].
 In the statistical noise setting, a natural approach for learning is to perform system identification and then use the identified system for prediction.
@@
 The work of [GLS+20] extends identification-based techniques to adversarial noise and marginally stable systems.
-The class $\prod^{*}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [HSZ17].
+The class $\Pi^{\star}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [HSZ17].
 That paper showed how to learn partially observable and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
@@
 Learning without recovery was studied in [GLS+20] even for marginally stable systems.
-More recently, tensor methods were used in [BLMY23a], and in [BLMY23b] to learn a mixture of linear dynamical systems.
+More recently, tensor methods were used in [BLM+23a], and in [BLM+23b] to learn a mixture of linear dynamical systems.
+The idea that a finite-dimensional state-space model induces a finite-order input-output recurrence is classical in linear systems theory.
+In the context of this chapter, the Cayley–Hamilton theorem provides a particularly clean explanation for why linear predictors can exactly represent noiseless finite-dimensional LDS.
 9.8 Exercises
-9.1. Prove that a $\varepsilon$-approximation of prediction classes implies an additional $\varepsilon T$ regret, as formalized in the Lemma 9.2.
-9.2. Prove that the class $\prod^{*}$ is optimal for fully observed linear dynamical systems with zero mean noise.
-9.3. Show that the policy class $\prod^{\star}_{h}$ can be learned by the policy class $\prod^{L}_{\tilde{h},k,\kappa}$ with $\tilde{h}=0$.
-9.4. Give precise regret bounds for learning the class $\prod^{L}_{h,k,\kappa}$ and $\prod^{\star}_{h}$ as a function of the natural class parameters $h,k,\kappa$ and any other relevant parameter.
-9.5. Consider the square loss in Algorithm 6. Give an algorithm based on the Online Newton Step algorithm (see [Haz16]) and write down the resulting regret bound.
-9.6.
-Write down the predictor class $\prod^{\star}$ and $\prod^{\star}_{h}$ for time-varying linear dynamical systems.
-Research the notion of adaptive regret from [Haz16], and reason about the conditions under which the class $\prod^{L}$ approximates $\prod^{*}$ for time-varying linear dynamical systems.
+9.1. Prove that Lemma 9.2 holds.
+9.2.
+Assume the system is fully observed, so that $C=I$, and that the process noise has zero conditional mean given the past.
+Show that the open-loop predictor obtained by unrolling the dynamics minimizes the expected squared prediction error among all predictors measurable with respect to the past inputs.
+9.3. Show that every truncated open-loop predictor of the form
+$\hat{\mathbf{y}}_{t}=\sum_{i=1}^{h}H_{i}\mathbf{u}_{t-i}$
+is a special case of a linear predictor in the class $\Pi^{L}_{0,h,\kappa}$ for an appropriate choice of $\kappa$.
+9.4.
+Assume that $\|\mathbf{u}_{t}\|\leq U$, $\|\mathbf{y}_{t}\|\leq Y$ for all $t$, and that the parameter space defining $\Pi^{L}_{h,k,\kappa}$ is endowed with a norm under which its diameter is at most $D$.
+Derive an explicit upper bound on $\mathrm{{regret}}_{T}({\mathcal{A}},\Pi^{L}_{h,k,\kappa})$ for Online Gradient Descent in terms of $h,k,\kappa,U,Y,D$, and $T$.
+9.5.
+Assume the square loss, and suppose moreover that predictions, outputs, and the parameter domain are uniformly bounded.
+Give an algorithm based on Online Newton Step for learning $\Pi^{L}_{h,k,\kappa}$, and derive the corresponding regret bound.
 Chapter 10 Kalman Filtering
-State estimation plays a central role in control and prediction tasks, where an agent seeks to infer latent states from noisy observations.
-The Kalman filter is a classical and widely used method for solving this problem, providing optimal state estimates under the assumption of linear system dynamics and Gaussian noise.
-A key feature of the Kalman filter and a fundamental limitation is that it requires knowledge of the system matrices, including both the state transition matrix and the observation model.
-These matrices encode how the system evolves over time and how observations relate to the hidden state.
-Although the Kalman filter is provably optimal when these parameters are known, real-world applications often involve partially known or entirely unknown system dynamics, making explicit identification challenging.
-In contrast, online learning approaches, such as those introduced in Chapter 9, provide a model-free open-loop alternative by directly learning optimal linear predictors from the data.
-Instead of estimating a latent state and system matrices, these methods construct forecasting models using past observations, making them more adaptive to unknown or time-varying systems.
-However, these approaches do not incorporate feedback, meaning that they operate in an open-loop setting, predicting future states without correcting for new observations.
-This chapter moves beyond open-loop prediction to closed-loop state estimation, where past observations continuously refine future predictions.
-The Kalman filter accomplishes this by recursively updating estimates based on incoming measurements, achieving Bayes-optimal performance under Gaussian noise.
-This structured feedback mechanism allows for more accurate and robust estimation compared to open-loop methods, particularly in dynamic and uncertain environments.
-Beyond classical Kalman filtering, we also introduce a closed-loop online learning method that serves as a convex relaxation of the Kalman filter.
-This approach retains the Kalman filtering feedback mechanism, while eliminating the need for explicit system identification, making it applicable in adversarial or uncertain environments.
-By formulating state estimation as an online convex optimization problem, this method provides provable regret guarantees and serves as a bridge between Bayesian filtering and learning-based estimation techniques.
-We provide a rigorous introduction to Kalman filtering, including its derivation, optimality properties, and computational aspects.
-We also discuss its limitations and contrast it with learning-based alternatives, setting the stage for spectral filtering techniques in the next chapter, which further generalize state estimation in unknown dynamical systems.
+In the previous chapter we defined what it means to learn in dynamical systems, and established relationship between prediction classes.
+Most notably, we showed that despite nonconvexity of the class of linear dynamical predictors, they can be learned exactly using the convex relaxation of linear predictors.
+In this chapter we show another approach to learn the class of linear dynamical predictors.
+This approach is more limited in terms of the generality where it applies, it is based on generative data and statistical assumptions about the noise.
+However, it allows for more than just prediction: Kalman filtering allows state estimation which is very useful by its own right in various applications.
 10.1 Observable Systems
-Observability is a fundamental concept in control theory that determines whether the internal state of a system can be uniquely inferred from its output over time.
-In the infinite-horizon setting, prediction is meaningful only if observations remain bounded, a condition closely related to system controllability (as discussed in Section 4.3).
-A general dynamical system maps an initial state to an infinite sequence of observations.
-If this mapping is injective, meaning that each initial state corresponds to a unique sequence of observations, then the system is said to be observable.
-Formally:
+Observability asks whether the hidden state of a partially observed dynamical system can be recovered from its observations.
+This is the basic structural condition that makes state estimation meaningful: if two different hidden states can produce the same observation history, then no estimator—Kalman filter or otherwise—can determine the state uniquely.
+This notion is dual to controllability, but conceptually different.
+Controllability asks whether one can drive the state by choosing controls; observability asks whether one can infer the state from what is observed.
+For linear systems, observability admits a clean rank characterization through the Kalman observability matrix.
 Definition 10.1.
-A partially observed dynamical system is observable if and only if every sequence of observations uniquely determines the initial hidden state.
-A simple example of an unobservable system is given as the following variant of the double integrator from section 2.1.2.
-Consider the double integrator dynamical system with $\Delta=1$ and partial observability where the observation matrix is given as follows
-$A=\begin{bmatrix}1\ 1\\ 0\ 1\end{bmatrix}\ ,\ C=[0\ 1].$
+Consider the noiseless partially observed linear system
+$\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t},\qquad\mathbf{y}_{t}=C\mathbf{x}_{t},$
+with known input sequence $\{\mathbf{u}_{t}\}_{t\geq 0}$.
+We say that the pair $(A,C)$ is observable if for every two initial states $\mathbf{x}_{0},\mathbf{x}_{0}^{\prime}$, equality of the resulting observation sequences under the same inputs implies $\mathbf{x}_{0}=\mathbf{x}_{0}^{\prime}$.
+Equivalently, the observation history uniquely determines the initial state.
+A simple example of an unobservable system is the following variant of the double integrator from Section 2.1.2.
+Consider the double integrator dynamical system with $\Delta=1$ and partial observability where the observation matrix is given by
+$A=\begin{bmatrix}1&1\\ 0&1\end{bmatrix},\qquad C=\begin{bmatrix}0&1\end{bmatrix}$.
 This system allows only the velocity of the object to be observed, while the position remains hidden.
@@
 Indeed, this is a special case of the Galilean relativity principle of Newtonian mechanics, which states that the laws of physics are the same in all inertial reference frames.
-For linear systems, observability can be verified using a rank condition on the Kalman observability matrix, given by:
-$\hat{K}(A,C)=\begin{bmatrix}C\\ CA\\ \dots\\ CA^{d_{\mathbf{x}}-1}\end{bmatrix}$.
-where $A$ and $C$ are the system matrices that define the state transition and observation model, respectively.
-Notice that in our example, the Kalman observability matrix is given by $\hat{K}(A,C)=\begin{bmatrix}0\ 1\\ 0\ 1\end{bmatrix}$, which is degenerate.
-This is not by chance; the system is observable if and only if the observability matrix has full rank.
+For linear systems, observability can be verified using a rank condition on the Kalman observability matrix, given by
+$\hat{K}(A,C)\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\begin{bmatrix}C\\ CA\\ \dots\\ CA^{d_{\mathbf{x}}-1}\end{bmatrix}$.
+where $A$ and $C$ are the system matrices that define the state transition and observation model, respectively. Notice that in our example, the Kalman observability matrix is given by
+$\hat{K}(A,C)=\begin{bmatrix}0&1\\ 0&1\end{bmatrix}$,
+which is degenerate.
+This is not by chance: the system is observable if and only if the observability matrix has full rank.
 We ask the reader to provide a proof of the following result in Exercise 10.6.
 Theorem 10.2.
-A partially-observed dynamical system specified by system parameters $(A,B,C)$ is observable if and only if $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))=d_{\mathbf{x}}$.
-10.2 Optimality of the Kalman Filter
-10.2.1 The Power of Linear Predictors
-One of the important conclusions from this chapter is that the policy class of linear predictors is optimal with respect to closed-loop policies. We prove this by proving two facts:
-  1. 1.
-    The class of linear policies $\varepsilon$-approximates the class of Kalman filters, in the sense of Definition 9.1.
-  2. 2.
-    The class of Kalman filters is optimal for closed-loop policies.
-The conclusion from these two facts is that algorithm 6 can be used to learn a policy that competes with the best closed-loop optimal predictor in hindsight.
-If the signal comes from a linear dynamical system subject to Gaussian noise, this means competing with the optimal predictor, which is a Kalman filter.
-We prove this formally at the end of this section.
-10.2.2 Closed-Loop Optimal Predictors
-Open-loop predictors, such as those introduced in the previous chapter, leverage system inputs and initial conditions to make predictions.
-However, they do not incorporate feedback from observations, making them susceptible to modeling errors and disturbances.
-Closed-loop prediction rules address this limitation by continuously adjusting estimates based on new observations.
-The most well-known example of such a predictor is the Kalman filter, which produces the best linear predictor in a least-squares sense under the assumption that process and observation noise are Gaussian with known covariances.
-Formally, let $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ denote the process and observation noise, assumed to be zero-mean and i.i.d., with covariance matrices:
-$\Sigma_{x}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[\mathbf{w}_{t}\mathbf{w}_{t}^{\top}\right],\quad\Sigma_{y}=\mathbb{E}\left[\mathbf{v}_{t}\mathbf{v}_{t}^{\top}\right].$
-So far, we have focused on predicting observations, that is, $\mathbf{y}_{t}$.
-However, the Kalman filter also provides optimal state estimates $\hat{\mathbf{x}}_{t}$ given past observations $\mathbf{y}_{t-1:1}$ and controls $\mathbf{u}_{t-1:1}$.
-This is motivated by the fact that the least-squares optimal predictor of observations can be expressed as:
-$\mathbf{y}_{t}=C\mathbf{x}_{t}.$
-In Exercise 10.6, we ask the reader to provide a proof of this statement.
-Based upon this observation, the following theorem is the essence of the Kalman filter, which recovers estimates for states as well as linear projection matrices for optimal prediction of the next observation.
+A partially-observed dynamical system with system matrices $(A,C)$ is observable if and only if $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))=d_{\mathbf{x}}$.
+10.2 The Kalman Filter
+Chapter 9 focused on open-loop predictors, which map past observations and controls to a forecast of the next output.
+Such predictors do not maintain or update an explicit estimate of the hidden state.
+The Kalman filter is the canonical closed-loop predictor for linear systems with zero-mean process and observation noise.
+It recursively combines the system dynamics with the latest measurement, producing the best linear least-squares estimate of the hidden state when the system matrices and noise covariances are known.
+Formally, let $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ denote the process and observation noise, assumed to be zero-mean and i.i.d., with covariance matrices
+$\Sigma_{x}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}[\mathbf{w}_{t}\mathbf{w}_{t}^{\top}],\qquad\Sigma_{y}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}[\mathbf{v}_{t}\mathbf{v}_{t}^{\top}].$
+Although our ultimate goal is to predict observations, it is convenient to first estimate the hidden state.
+Once we have a least-squares estimate $\hat{\mathbf{x}}_{t}$ of $\mathbf{x}_{t}$ given the past, the corresponding least-squares prediction of the observation is $\hat{\mathbf{y}}_{t}=C\hat{\mathbf{x}}_{t}$; see Exercise 10.6.
+The following theorem is the essence of the Kalman filter. It constructs a recursive state estimator that is optimal among linear estimators based on past observations and controls.
 Theorem 10.3.
 Let the control sequence $\mathbf{u}_{t}$ be an arbitrary affine function of $\mathbf{y}_{1:t}$.
-There exists an efficiently computable sequence of linear predictors $\hat{\mathbf{x}}_{t+1}$ defined recursively as
+There exists an efficiently computable sequence of linear state estimates $\hat{\mathbf{x}}_{t+1}$ defined recursively as
 $\hat{\mathbf{x}}_{t+1}=(A-L_{t}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+L_{t}\mathbf{y}_{t},$
 such that
-$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{L}}\mathbb{E}\|\mathbf{x}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
+$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L}\mathbb{E}\|\mathbf{x}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2},$
+where the minimum ranges over all linear maps from $(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})$ to $\mathbb{R}^{d_{\mathbf{x}}}$.
 To prove this theorem, we require some basic tools from linear regression which we state next.
@@
 $\mathbb{E}\|X-A^{*}Y-B^{*}Z\|^{2}\leq\min_{A,B}\mathbb{E}\|X-AY-BZ\|^{2}.$
-In the remaining proof, we use this observation to construct an optimal least-squares estimator $\hat{\mathbf{x}}_{t+1}$ for $\mathbf{x}_{t+1}$ given $\hat{\mathbf{x}}_{t}$.
-In particular, we know by induction hypothesis that
-$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{L}}\mathbb{E}\|\mathbf{x}_{t}-L\left[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}\right]\|^{2}.$
-But, is $\hat{\mathbf{x}}_{t}$ any good for predicting $\mathbf{x}_{t+1}$? To answer this, we note the following structural characterization of least-squares (Exercise 10.6).
+In the remaining proof, we use this observation to construct an optimal least-squares estimator $\hat{\mathbf{x}}_{t+1}$ for $\mathbf{x}_{t+1}$ from the available information.
+In particular, by the induction hypothesis,
+$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L}\mathbb{E}\|\mathbf{x}_{t}-L\left[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}\right]\|^{2},$
+where the minimum ranges over all linear maps from $(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})$ to $\mathbb{R}^{d_{\mathbf{x}}}$.
+But is $\hat{\mathbf{x}}_{t}$ any good for predicting $\mathbf{x}_{t+1}$?
+To answer this, we note the following structural characterization of least-squares (Exercise 10.6).
 Lemma 10.5.
@@
 10.2.2 Analysis
-We are now ready to prove the main property of the Kalman filter below.
+We are now ready to prove the main property of the Kalman filter below. Throughout this section, assume that $\Sigma_{y}\succ 0$.
 Proof of Theorem 10.3.
 Since the estimator is defined recursively, a natural strategy is to proceed by induction. Let us say $\hat{\mathbf{x}}_{t}$ is such a predictor. Furthermore, let
-$\Sigma_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}.$
-Since for any $s<t$
+$\Sigma_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}\right].$
+Since for any $s<t$,
 $\mathbb{E}\left[(\mathbf{x}_{t+1}-B\mathbf{u}_{t})\mathbf{y}_{s}^{\top}\right]=\mathbb{E}\left[(A\mathbf{x}_{t}+\mathbf{w}_{t})\mathbf{y}_{s}^{\top}\right]=A\mathbb{E}\left[\mathbf{x}_{t}\mathbf{y}_{s}^{\top}\right],$
 using the first part of Lemma 10.5, we have that
-$\mathbb{E}\|\mathbf{x}_{t+1}-A\hat{\mathbf{x}}_{t}-B\mathbf{u}_{t}\|^{2}\leq\min_{L\in\Pi^{L}}\mathbb{E}\|\mathbf{x}_{t+1}-B\mathbf{u}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
+$\mathbb{E}\|\mathbf{x}_{t+1}-A\hat{\mathbf{x}}_{t}-B\mathbf{u}_{t}\|^{2}\leq\min_{L}\mathbb{E}\|\mathbf{x}_{t+1}-B\mathbf{u}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2},$
+where the minimum ranges over all linear maps from $(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})$ to $\mathbb{R}^{d_{\mathbf{x}}}$.
 But, when predicting $\mathbf{x}_{t+1}$, we have more information.
 In particular, we also have $\mathbf{y}_{t}$.
-Consider random variables $\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}$.
+Consider the random variable $\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}$.
 Can we use this covariate alone to predict $\mathbf{x}_{t+1}$?
 Note that
-$\mathbb{E}\left[(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})^{\top}\right]$ $=\Sigma_{v}+C\mathbb{E}\left[(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}\right]C^{\top}$
-$=\Sigma_{v}+C\Sigma_{t}C^{\top},$
+$\mathbb{E}\left[(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})^{\top}\right]$ $=\Sigma_{y}+C\mathbb{E}\left[(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}\right]C^{\top}$
+$=\Sigma_{y}+C\Sigma_{t}C^{\top},$
 $\mathbb{E}\left[(\mathbf{x}_{t+1}-B\mathbf{u}_{t})(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})^{\top}\right]$ $=\mathbb{E}\left[(A\mathbf{x}_{t}+\mathbf{w}_{t})(C(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})+\mathbf{v}_{t})^{\top}\right]$
@@
 where the equality indicated in braces follows from the second part of Lemma 10.5, by noting that $\hat{\mathbf{x}}_{t}$ is a linear function of $\mathbf{y}_{1:t-1}$.
-In the preceding paragraphs, we have constructed two least-squares predictors for $\hat{\mathbf{x}}_{t+1}$. Observe that for any $s<t$
+In the preceding paragraphs, we have constructed two least-squares predictors for ${\mathbf{x}}_{t+1}$. Observe that for any $s<t$,
 $\mathbb{E}(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})\mathbf{y}_{s}^{\top}=C\mathbb{E}(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})\mathbf{y}_{s}^{\top}=0,$
@@
 To establish that such predictors can be computed efficiently, note that
-$\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1}=(A-L_{t}C)(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})+\mathbf{w}_{t}+L_{t}\mathbf{v}_{t}.$
-Finally, we give a recursive relation on $\Sigma_{t}$ as follows–
+$\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1}=(A-L_{t}C)(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})+\mathbf{w}_{t}-L_{t}\mathbf{v}_{t}.$
+Finally, we give a recursive relation on $\Sigma_{t}$ as follows:
 $\Sigma_{t+1}$ $\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})^{\top}\right]$
 $=(A-L_{t}C)\Sigma_{t}(A-L_{t}C)^{\top}+\Sigma_{x}+L_{t}\Sigma_{y}L_{t}^{\top}$
-$=A\Sigma_{t}A^{\top}+\Sigma_{x}+L_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})L_{t}^{\top}-2A\Sigma_{t}C^{\top}L_{t}^{\top}$
-$=A\Sigma_{t}A^{\top}-A\Sigma C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}C\Sigma A^{\top}+\Sigma_{x}.$
+$=A\Sigma_{t}A^{\top}+\Sigma_{x}+L_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})L_{t}^{\top}-A\Sigma_{t}C^{\top}L_{t}^{\top}-L_{t}C\Sigma_{t}A^{\top}$
+$=A\Sigma_{t}A^{\top}-A\Sigma_{t}C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}C\Sigma_{t}A^{\top}+\Sigma_{x}.$
 10.2.3 Infinite Horizon Kalman Filtering
-In the infinite-horizon setting, the steady-state Kalman filter can be described by the following linear dynamical system
+In the infinite-horizon setting, when the covariance recursion converges, the Kalman filter admits a steady-state description with a fixed gain matrix $L$.
+The resulting estimator is itself a linear dynamical system driven by the control inputs and the incoming observations:
 $\hat{\mathbf{x}}_{t+1}$ $=(A-LC)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+L\mathbf{y}_{t},$
 $\hat{\mathbf{y}}_{t}$ $=C\hat{\mathbf{x}}_{t},$
-where $L=A\Sigma(C\Sigma C^{\top}+\Sigma_{y})^{-1}$ and $\Sigma$ satisfies the Riccati equation
+where $L=A\Sigma C^{\top}(C\Sigma C^{\top}+\Sigma_{y})^{-1}$ and $\Sigma$ satisfies the Riccati equation
 $\Sigma=A\Sigma A^{\top}-A\Sigma C^{\top}(C\Sigma C^{\top}+\Sigma_{y})^{-1}C\Sigma A^{\top}+\Sigma_{x}.$
-However, when does a solution to this Riccati equation exist?
-A sufficient condition for this is the observability of the underlying dynamical system.
-In particular, any observable partially-observed linear dynamical system admits a solution to the Riccati equation listed above, and hence, has a steady-state Kalman filter.
-Mirroring the link between controllability and stabilizability, observability implies $\rho(A-LC)<1$.
-10.2.6 Linear Predictors and the Kalman Filter
-For observable linear dynamical systems, the Kalman filter is contained in the class of linear predictors. This is formally stated in the following lemma, whose proof is left as Exercise 10.6.
+The existence of a stabilizing solution to this Riccati equation is a classical question.
+Under standard assumptions—typically detectability of $(A,C)$ together with an appropriate stabilizability condition for the state-noise model—the Riccati equation admits a stabilizing solution.
+In that case the steady-state Kalman gain $L$ is well defined and the closed-loop matrix ALCA-LC is stable.
+We do not pursue the sharp conditions here.
+10.2.4 Linear Predictors Approximate the Kalman Filter
+At first sight, the steady-state Kalman filter appears to lie outside the class of predictors studied in Chapter 9 because it maintains an internal state estimate and uses feedback from the newest observation.
+The next lemma gives a relationship between the class of predictors.
+Once the closed-loop matrix ALCA-LC is stable, the Kalman recursion can be unrolled and truncated, yielding a finite-memory linear predictor with exponentially small approximation error.
+This is the closed-loop analogue of the truncation phenomenon from Chapter 9.
 Lemma 10.6.
-Consider any partially-observed linear dynamical system $(A,B,C)$ where the Kalman filtering matrix $L$ is such that the system ALCA-LC is $\gamma$-stable.
-The Kalman filtering policy is $\varepsilon$-contained in the class of linear policies $\Pi^{L}_{h,k,\kappa}$, where $h=k=\gamma\log\frac{\gamma T}{\varepsilon}$ and $\kappa=\gamma^{3}$.
+Fix $M\geq 1$. Consider the family of steady-state Kalman filters
+$\hat{\mathbf{x}}_{t+1}=F\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+L\mathbf{y}_{t},\qquad\hat{\mathbf{y}}_{t}=C\hat{\mathbf{x}}_{t},$
+with $\hat{\mathbf{x}}_{1}=0$, where F=ALCF=A-LC satisfies
+$\|F^{s}\|\leq Me^{-s/M}\qquad\text{for all }s\geq 0,$
+and $\|B\|+\|L\|+\|C\|\leq M$, as wel as $\max_{t\in[T]}\|\mathbf{u}_{t}\|\leq M,\qquad\max_{t\in[T]}\|\mathbf{y}_{t}\|\leq M$.
+Then the class $\Pi^{L}_{h,h,\kappa}$ $\varepsilon$-approximates this family over horizon $T$ provided
+$h\geq M\log\!\left(\frac{4M^{5}}{\varepsilon}\right),\qquad\kappa\geq 4M^{4}.$
 10.3 Bayes-Optimality under Gaussian Noise
-In this subsection, we further restrict $\mathbf{w}_{t},\mathbf{v}_{t}$ to be i.i.d.
-draws from Gaussian distributions, and proceed to establish that the Kalman filter in such cases enjoys a stronger notion of optimality; it produces the best estimator (under a least-squares loss) of the state given past observations among the class of all (not just linear) predictors.
+The previous section established the Kalman recursion as the optimal linear closed-loop estimator.
+Under Gaussian noise, this linear optimality strengthens to full Bayes optimality.
 [^1] This is also called the Bayes-optimal predictor.
 See bibliographic materials for more details and literature.
-.
-The setting we consider is that of Gaussian perturbations $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ and known systems matrices $(A,B,C)$.
-In particular, we consider the case where $\mathbf{w}_{t}\sim\mathcal{N}(0,\Sigma_{x})$ and $\mathbf{v}_{t}\sim\mathcal{N}(0,\Sigma_{y})$ are i.i.d.
-Gaussian and $x_{0}=0$.
-Given the realizations $\mathbf{y}_{1},\mathbf{y}_{2}\dots,\mathbf{y}_{t}$ under some control input sequence $\mathbf{u}_{1},\mathbf{u}_{2},\dots\mathbf{u}_{t}$, we would like to compute a prediction $\hat{\mathbf{y}}_{t+1}$ that minimizes the mean square error, which is simply the expected observation as per
-$\hat{\mathbf{y}}_{t+1}=\mathbb{E}[\mathbf{y}_{t+1}|\mathbf{y}_{1:t},\mathbf{u}_{1:t}].$
-Since $\hat{\mathbf{y}}_{t+1}=C\hat{\mathbf{x}}_{t+1}$, we will recursively compute $\hat{\mathbf{x}}_{t+1}$ where
-$\hat{\mathbf{x}}_{t+1}=\mathbb{E}[\mathbf{x}_{t+1}|\mathbf{y}_{1:t},\mathbf{u}_{1:t}].$
+In this section we assume that $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ are i.i.d.
+Gaussian and show that the Kalman filter computes the conditional expectation of the hidden state, and hence the mean-square optimal predictor of the next observation.
+Concretely, assume that $\mathbf{w}_{t}\sim\mathcal{N}(0,\Sigma_{x})$ and $\mathbf{v}_{t}\sim\mathcal{N}(0,\Sigma_{y})$ are i.i.d.
+Gaussian and that $\mathbf{x}_{0}=0$.
+Given realizations $\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{t}$ under a control sequence $\mathbf{u}_{1},\mathbf{u}_{2},\dots,\mathbf{u}_{t}$, we would like to compute
+$\hat{\mathbf{y}}_{t+1}=\mathbb{E}[\mathbf{y}_{t+1}\mid\mathbf{y}_{1:t},\mathbf{u}_{1:t}].$
+Since $\hat{\mathbf{y}}_{t+1}=C\hat{\mathbf{x}}_{t+1}$, it suffices to recursively compute
+$\hat{\mathbf{x}}_{t+1}=\mathbb{E}[\mathbf{x}_{t+1}\mid\mathbf{y}_{1:t},\mathbf{u}_{1:t}].$
 Theorem 10.7.
@@
 $\hat{\mathbf{x}}_{t+1}=(A-L_{t}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+L_{t}\mathbf{y}_{t},$
-for a sequence of matrices $L_{t}\in{\mathbb{R}}^{d_{x}\times d_{y}}$ that can be efficiently computed given $(A,B,C)$, and is independent of the realizations $\mathbf{u}_{1:T},\mathbf{y}_{1:T}$.
+for a sequence of matrices $L_{t}\in\mathbb{R}^{d_{\mathbf{x}}\times d_{y}}$ that can be efficiently computed given $(A,B,C)$, and is independent of the realizations $\mathbf{u}_{1:T},\mathbf{y}_{1:T}$.
 Proof.
@@
 Therefore, we have that
-$\mathbf{x}_{t}|(\mathbf{y}_{1:t},\mathbf{u}_{1:t-1})$ $\sim\mathcal{N}(\hat{\mathbf{x}}_{t}+\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}),\Sigma^{\prime}_{t}),$
+$\mathbf{x}_{t}|(\mathbf{y}_{1:t},\mathbf{u}_{1:t-1})$ $\sim\mathcal{N}\Big(\hat{\mathbf{x}}_{t}+\Sigma_{t}C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}),\Sigma^{\prime}_{t}\Big),$
 $\Sigma^{\prime}_{t}$ $=\Sigma_{t}-\Sigma_{t}C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}C\Sigma_{t}.$
@@
 $\hat{\mathbf{x}}_{t+1}$ $=A\mathbb{E}[\mathbf{x}_{t}|\mathbf{y}_{1:t},\mathbf{u}_{1:t-1}]+B\mathbf{u}_{t}$
-$=(A-\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}\mathbf{y}_{t}.$
+$=(A-\underbrace{A\Sigma_{t}C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}\mathbf{y}_{t}.$
 It remains to obtain a recursive relation on $\Sigma_{t}$. To this extent, from the state evolution of a linear dynamical system, we have that
@@
 10.4 Conclusion
-In this chapter, we explored the Kalman filter as a fundamental tool for state estimation in dynamical systems, particularly in the presence of noise and uncertainty.
-We established its optimality in the least-squares sense and examined its connections to online convex optimization, reinforcing its relevance in both classical control and modern learning-based paradigms.
-Key insights from this chapter include the role of observability in ensuring accurate state estimation, the derivation of the Kalman filter update equations, and the recursive structure that allows for efficient real-time implementation.
-We also presented the filter’s steady-state behavior, particularly in the infinite-horizon setting, where it converges to a solution governed by the Riccati equation.
-Beyond its theoretical guarantees, the Kalman filter is widely used in practical applications ranging from autonomous systems and robotics to finance and neuroscience.
-However, challenges remain, particularly in scenarios with non-Gaussian noise, time-varying dynamics, or adversarial disturbances.
-These limitations motivate extensions such as robust filtering, nonlinear filtering (e.g., the extended and unscented Kalman filters) and learning-based adaptations that integrate statistical inference with modern machine learning techniques.
-The next chapter will transition from the classical filtering framework to spectral filtering, providing an alternative approach to state estimation that leverages spectral methods for efficient learning in dynamical systems.
+This chapter extends the story of Chapter 9 from open-loop prediction to closed-loop state estimation.
+The main conceptual point is that feedback changes the benchmark but not the convex relaxation: the Kalman filter is the natural optimal closed-loop predictor, yet stable Kalman filters are still approximable by finite-memory linear predictors.
+We first introduced observability as the structural condition that makes state estimation possible.
+We then derived the Kalman recursion and its steady-state form, and showed that under Gaussian noise it is Bayes-optimal.
+Finally, we connected this classical estimator back to the learning perspective of the previous chapter by showing that finite-memory linear predictors can approximate stable Kalman filters.
+Consequently, the online learning machinery developed in Chapter 9 continues to apply in partially observed stochastic systems.
+The next chapter moves from Kalman filtering to spectral filtering.
+There the emphasis shifts again from explicit state estimation to improper learning of predictors, but the central theme remains the same: exploit linear structure while avoiding direct nonconvex identification of the underlying dynamical system.
 10.5 Bibliographic Remarks
 In 1960, Rudolph Kalman gave a state-space solution to the filtering problem that was highly amenable to further extensions [KAL60].
-See [AM91] for a historical perspective on its development and aftermaths.
+See [AM91] for a historical perspective on its development and aftermath.
 For unknown systems, in the stochastic case, one may estimate the system matrices, as we have discussed before, before applying optimal filtering.
-Improper learning approaches to this problem were studied in [KMTM19].
+Improper learning approaches to this problem were studied in [KMT+19].
 State estimation is a fundamental problem in control and signal processing, with numerous approaches beyond the classical Kalman filter. Alternative methods include:
   • Nonlinear Filtering Extensions.
-The Kalman filter assumes that the dynamics is linear system and the noise is Gaussian, which limits its effectiveness in nonlinear settings.
+The Kalman filter assumes that the dynamics is linear and the noise is Gaussian, which limits its effectiveness in nonlinear settings.
 The Extended Kalman Filter (see, for example, [UA21]) addresses this issue by linearizing the dynamics at each time step via a first-order Taylor series expansion.
@@
 Notable extensions using more sophisticated sampling include the Unscented Kalman Filter [JU04], leading to better estimation accuracy in strongly nonlinear systems.
-  • Particle Filtering (Sequential Monte Carlo Methods)
+  • Particle Filtering (Sequential Monte Carlo Methods).
 For highly nonlinear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative.
 These methods approximate the posterior distribution using a set of weighted samples (particles) and employ importance sampling and resampling techniques.
-  • Adversarial & Robust Filtering
+  • Adversarial & Robust Filtering.
 Robust filtering methods seek to provide guarantees under adversarial noise.
 The $H_{\infty}$ filter [BB08] minimizes the worst-case estimation error rather than the mean-squared error, making it suitable for applications with unknown disturbances.
-Although the Kalman filter remains a cornerstone of state estimation, these alternative methods provide valuable extensions and improvements for nonlinear, adversarial, and data-driven settings.
-In the next chapter, we will study an alternative using a technique which circumvents the need for state estimation altogether and has stronger guarantees for sequential prediction.
+In the next chapter, we will study an alternative using a technique which circumvents the need for state estimation altogether and has dimension-free guarantees for certain cases of sequential prediction.
 10.6 Exercises
-10.1. In this exercise we prove Theorem 10.2 in two parts:
+10.1. In this exercise we prove Theorem 10.2.
   1. 1.
-    Show that if the rank of the Kalman observability matrix is less than $d_{\mathbf{x}}$, then there are two initial states that after $d_{\mathbf{x}}$ iterations map to the same state.
-Hence, the initial state is not uniquely determined.
+    Suppose $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))<d_{\mathbf{x}}$.
+Show that there exists a nonzero vector $v\in\ker(\hat{K}(A,C))$.
+Conclude that the two initial states $\mathbf{x}_{0}$ and $\mathbf{x}_{0}+v$ generate the same observation sequence under the same input sequence, and hence the system is not observable.
   2. 2.
-    In contrast, show that if the rank of $\hat{K}$ is full, then at any time $t$, $\mathbf{x}_{t}$ determines $\mathbf{x}_{0}$ uniquely.
+    Suppose $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))=d_{\mathbf{x}}$. Show that the map
+    $\mathbf{x}_{0}\mapsto\begin{bmatrix}C\mathbf{x}_{0}\\ CA\mathbf{x}_{0}\\ \vdots\\ CA^{d_{\mathbf{x}}-1}\mathbf{x}_{0}\end{bmatrix}$
+    is injective. Conclude that, given the input sequence, the first $d_{\mathbf{x}}$ observations uniquely determine the initial state.
 10.2. Prove Lemma 10.5, which states the following. Let $X,Y$ be possibly correlated random variables. Then
@@
 10.4.
-In this exercise we conclude that the Kalman filter can be improperly learned by the class of linear predictors.
-Prove this fact as formally stated in Lemma 10.6.
-Conclude that the Kalman filter can be learned using Algorithm 6.
+Prove Lemma 10.6 by unrolling the steady-state Kalman recursion and truncating the resulting linear predictor.
+Conclude that, under the assumptions of the lemma, the family of stable steady-state Kalman filters can be learned improperly using Algorithm 6.
 10.5. Prove Lemma 10.8, which is stated below. If a random variable pair $(\mathbf{x},\mathbf{y})$ are distributed as
@@
 Chapter 11 Spectral Filtering
-In the previous chapter, we explored two major approaches for state estimation and prediction in unknown linear dynamical systems:
+In the previous chapters, we explored two major approaches for state estimation and prediction in unknown linear dynamical systems:
   1. 1.
@@
 Instead of estimating the full state and transition matrices, this approach constructs a forecasting model using past inputs and outputs.
-Although this avoids the complexity of full system identification, it introduces a major limitation: the number of parameters required depends on the stability constant of the system, which can be large and unknown, making learning impractical for unstable or poorly conditioned systems.
+Although this avoids the complexity of full system identification, it introduces a major limitation: the number of parameters required depends on the hidden dimension of the system, or the stability constant of the system, which can be large and unknown, making learning impractical for high-dimensional or poorly conditioned systems.
 In this chapter, we introduce Spectral Filtering, an alternative approach that bypasses these limitations.
@@
 For simplicity, we assume that the initial state is zero $\mathbf{x}_{0}=0$.
-Recall the set of optimal predictors, the class $\Pi_{\star}$, which for scalar systems can written and simplified to,
-$\mathbf{y}_{t}$ $=\sum_{i=1}^{t}CA^{i-1}B\mathbf{u}_{t-i}=\beta\sum_{i=1}^{t}\alpha^{i-1}\mathbf{u}_{t-i},$
-Where $\beta=CB\in{\mathbb{R}}$. Thus, we can write the difference between two consecutive iterates as
-$\mathbf{y}_{t}-\mathbf{y}_{t-1}$ $=\beta(\sum_{i=1}^{t}\alpha^{i-1}\mathbf{u}_{t-i}-\sum_{i=1}^{t-1}\alpha^{i-1}\mathbf{u}_{t-1-i})$
-$=\beta(\mathbf{u}_{t-1}+\sum_{i=1:t-1}(\alpha^{i}-\alpha^{i-1})\mathbf{u}_{t-1-i})$
-$=\beta\mathbf{u}_{t-1}+\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1}),$ (11.1)
-which is a geometric series involving $\alpha$ multiplying the past controls. We used the operator notation to denote the operator
-$\mu_{\alpha}(\mathbf{u}_{t-1:1})=\sum_{i=1:t-1}(\alpha-1)\alpha^{i-1}\mathbf{u}_{t-1-i}=\mu_{\alpha}^{\top}\tilde{\mathbf{u}}_{t-1}.$ (11.2)
-Above we used the vector notation for
-$\mu_{\alpha}=(\alpha-1)\begin{bmatrix}1&\alpha&\alpha^{2}&\ldots&\alpha^{T-1}\end{bmatrix}\in{\mathbb{R}}^{T},$
+Recall the set of optimal predictors, the class $\Pi_{\star}$, which for scalar systems can be written and simplified to,
+$\mathbf{y}_{t}$ $=\sum_{i=1}^{t-1}CA^{i-1}B\mathbf{u}_{t-i}=\beta\sum_{i=1}^{t-1}\alpha^{i-1}\mathbf{u}_{t-i}=\beta\cdot\tilde{\mathbf{u}}_{t}^{\top}\mu_{\alpha},$ (11.1)
+where $\beta=CB\in{\mathbb{R}}$ is a scalar and we used the vector notation for
+$\mu_{\alpha}=\begin{bmatrix}1&\alpha&\alpha^{2}&\ldots&\alpha^{T-1}\end{bmatrix}^{\top}\in{\mathbb{R}}^{T},$
 and we define $\tilde{\mathbf{u}}_{t}$ to be the padded vector $\mathbf{u}_{t-1:1}$:
-$\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}.$
+$\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0&\ldots&0\end{bmatrix}^{\top}\in{\mathbb{R}}^{T}.$
 The main idea of spectral filtering is to rewrite the vectors $\mu_{\alpha}$ in a more efficient basis, as opposed to the standard basis.
@@
 The matrix $Z_{T}$ is symmetric and is independent of $\alpha$.
-Let $\phi_{1},...,\phi_{k},...\in{\mathbb{R}}^{T}$ be the eigenvectors of $Z_{T}$ ordered by the magnitude of their corresponding eigenvalues.
+Let $\phi_{1},...,\phi_{h},...\in{\mathbb{R}}^{T}$ be the eigenvectors of $Z_{T}$ ordered by the magnitude of their corresponding eigenvalues.
 Our basis for representing $\mu_{\alpha}$ is spanned by the eigenvectors of $Z_{T}$, and we proceed to study their properties.
+Figure 11.1: The filters obtained by the eigenvectors of $Z_{T}$.
 11.1.1 The Magic of Hankel Matrices
-The matrix $Z_{T}$ is a special case of a Hankel Matrix, which is a square Hermitian matrix with each ascending skew-diagonal from left to right having the same value.
-An example of a Hankel matrix is the following,
-$W_{T}$ $=\int_{0}^{1}\begin{bmatrix}1&\alpha&\alpha^{2}&&\\ \alpha&\alpha^{2}&\alpha^{3}&&\\ \alpha^{2}&\alpha^{3}&\alpha^{4}&&\\ &&&...&\\ &&&&\alpha^{2T-2}\end{bmatrix}d\alpha=\begin{bmatrix}1&\frac{1}{2}&\frac{1}{3}&&\\ \frac{1}{2}&\frac{1}{3}&\frac{1}{4}&&\\ \frac{1}{3}&\frac{1}{4}&\frac{1}{5}&&\\ &&&...&\\ &&&&\frac{1}{2T-1}\end{bmatrix}$.
-This matrix is closely related to the $Z_{T}$ matrix we have introduced before, but not equal to it.
+The matrix $Z_{T}$ is called the Hilbert matrix, and it is a special case of a Hankel Matrix, which is a square Hermitian matrix with each ascending skew-diagonal from left to right having the same value.
+To spell out $Z_{T}$ more precisely,
+$Z_{T}$ $=\int_{0}^{1}\begin{bmatrix}1&\alpha&\alpha^{2}&&\\ \alpha&\alpha^{2}&\alpha^{3}&&\\ \alpha^{2}&\alpha^{3}&\alpha^{4}&&\\ &&&...&\\ &&&&\alpha^{2T-2}\end{bmatrix}d\alpha=\begin{bmatrix}1&\frac{1}{2}&\frac{1}{3}&&\\ \frac{1}{2}&\frac{1}{3}&\frac{1}{4}&&\\ \frac{1}{3}&\frac{1}{4}&\frac{1}{5}&&\\ &&&...&\\ &&&&\frac{1}{2T-1}\end{bmatrix}$.
 The properties of Hankel matrices have been studied extensively, and the following theorem captures one of their crucial properties.
 Theorem 11.1.
-Let $\sigma_{k}$ be the $k$-th largest singular value of a Hankel matrix $M\in{\mathbb{R}}^{T\times T}$, then there exists constants $c_{1},c_{2}$ such that :
-$\sigma_{k}(M)\leq c_{1}\cdot e^{-\frac{c_{2}k}{\log T}}$
+Let $\sigma_{k}$ be the $k$-th largest singular value of the Hilbert matrix $Z_{T}\in{\mathbb{R}}^{T\times T}$, then for $k,T>1$:
+$\sigma_{k}(Z_{T})\leq 4\pi^{2}\cdot e^{-\frac{k}{\log T}}$
 The proof of this theorem is beyond our scope, and references to details are given in the bibliographic material at the end of this chapter.
@@
 Definition 11.2.
-Define the class of spectral linear predictors of order $k$ recursively as
+Define the class of spectral linear predictors of order $h$ as
 […]
-Here the operator $\phi_{i}(\mathbf{u}_{t-1})$ takes the inner product of the $i$’th eigenvector of $Z_{T}$ with $\tilde{\mathbf{u}}_{t-1}$, where $\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0\ldots&0\end{bmatrix}\in{\mathbb{R}}^{T}$.
-Notice that as $h\mapsto T$, this class is equivalent to the class of all linear predictors on the inputs $\prod^{LIN}$.
+Here $\tilde{\mathbf{u}}_{t}^{\top}\phi_{i}$ evaluates the inner product of the $i$’th eigenvector of $Z_{T}$ with the padded history vector $\tilde{\mathbf{u}}_{t}$, where $\tilde{\mathbf{u}}_{t}=\begin{bmatrix}\mathbf{u}_{t-1}&\ldots&\mathbf{u}_{1}&0&\ldots&0\end{bmatrix}^{\top}\in{\mathbb{R}}^{T}$.
+Notice that as $h\to T$, this class is equivalent to the class of all linear predictors on the inputs $\Pi^{LIN}$.
 The reader may wonder at this point: Since a linear transformation composed with another linear transform remains linear, what is the benefit here?
@@
 This is the case in our setting: the spectral basis arising from the matrix $Z_{T}$ gives excellent approximations to the class of (standard) linear predictors, as we formally prove next.
-The following lemma is analogous to Lemma 9.4, with the significant difference that the number of parameters $h$ does not depend on the stabilizability parameter of the system $\gamma$ at all!
+The following Theorem is analogous to Lemma 9.4, with the significant difference that the number of parameters $h$ does not depend on the stabilizability parameter of the system $\gamma$ at all!
+Theorem 11.3.
+Suppose that $\mathbf{x}_{0}=0$ and $|u_{t}|\leq 1$ for all $t$. Then the class $\Pi_{h}$ for $h=O\!\left(\log^{2}\frac{T}{\varepsilon}\right)$ $\varepsilon$-approximates the class $\Pi_{\star}$.
+The proof of this lemma relies on the following crucial property of the spectral basis.
 Lemma 11.4.
-Suppose that $\mathbf{x}_{0}=0$, then the class $\prod_{h}^{\lambda}$ for $h=O(\log^{2}\frac{T}{\varepsilon})$, $\varepsilon$-approximates the class $\prod_{\star}$.
-The proof of this lemma relies on the following crucial property of the spectral basis, whose proof we defer to later.
-Lemma 11.4.
-Let $\{\phi_{j}\}$ be the eigenvectors of Z. Then there exist constants $c_{1},c_{2}>0$ such that for all $j\in[T]$, $\alpha\in[0,1]$, we have
-$|\phi_{j}^{\top}\mu_{\alpha}|\leq c_{1}e^{-\frac{c_{2}j}{\log T}}.$
-Using this property, we can prove Lemma 11.3 by simply shifting to a representation on the spectral basis as follows.
+Let $\{\phi_{j}\}$ be the orthonormal eigenvectors of $Z_{T}$, with corresponding eigenvalues $\sigma_{j}$. Then for all $j\in[T]$, $j,T>1$, and $\alpha\in[0,1]$, we have
+$|\phi_{j}^{\top}\mu_{\alpha}|\;\leq\;(4T^{2}\sigma_{j})^{1/4}.$
+In particular, using Theorem 11.1,
+$|\phi_{j}^{\top}\mu_{\alpha}|=O\!\left(T^{1/2}e^{-\frac{j}{4\log T}}\right).$
+Using this property, we can prove Theorem 11.3 by shifting to a representation on the spectral basis as follows.
+Proof of Theorem 11.3.
+Consider any sequence and any predictor from $\Pi_{\star}$, say $\pi_{1}=\pi_{A,B,C,0}$. Following equation (11.1), we can write
+$\mathbf{y}_{t}^{\pi_{1}}$ $=\beta\cdot\tilde{\mathbf{u}}_{t}^{\top}\mu_{\alpha}$
+$=\beta\cdot\tilde{\mathbf{u}}_{t}^{\top}\left(\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top}\right)\mu_{\alpha}\qquad\mbox{(basis property }\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top}=I)$
+$=\sum_{i=1}^{T}M_{i}^{\star}\big(\tilde{\mathbf{u}}_{t}^{\top}\phi_{i}\big),\qquad M_{i}^{\star}=\beta(\phi_{i}^{\top}\mu_{\alpha}).$
+By Lemma 11.4,
+$|M_{i}^{\star}|$ $=|\beta|\,|\phi_{i}^{\top}\mu_{\alpha}|\leq|\beta|\,(4T^{2}\sigma_{i})^{1/4}$
+$=O\!\left(|\beta|\,T^{1/2}e^{-\frac{i}{4\log T}}\right).$ (11.2)
+Consider the predictor $\pi_{2}\in\Pi_{h}$ that keeps only the first $h$ coefficients. Then we have
+$|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}|$ $=\left|\sum_{i=1}^{T}M_{i}^{\star}\big(\tilde{\mathbf{u}}_{t}^{\top}\phi_{i}\big)-\sum_{i=1}^{h}M_{i}^{\star}\big(\tilde{\mathbf{u}}_{t}^{\top}\phi_{i}\big)\right|$
+$=\left|\sum_{i=h+1}^{T}M_{i}^{\star}\big(\tilde{\mathbf{u}}_{t}^{\top}\phi_{i}\big)\right|$
+$\leq\sum_{i=h+1}^{T}|M_{i}^{\star}|\,\big|\tilde{\mathbf{u}}_{t}^{\top}\phi_{i}\big|$
+$\leq\sum_{i=h+1}^{T}|M_{i}^{\star}|\,\|\phi_{i}\|_{2}\|\tilde{\mathbf{u}}_{t}\|_{2}$ (Cauchy–Schwarz)
+$\leq\sqrt{T}\,\sum_{i=h+1}^{T}|M_{i}^{\star}|$ $\|\phi_{i}\|_{2}=1,\;|u_{t}|\leq 1$
+$\leq\sqrt{T}\,|\beta|\,T^{1/2}\sum_{i=h+1}^{\infty}e^{-\frac{i}{4\log T}}$ (11.2)
+The geometric tail can be strictly bounded via an integral:
+$\sum_{i=h+1}^{\infty}e^{-\frac{i}{4\log T}}\leq\int_{h}^{\infty}e^{-\frac{x}{4\log T}}dx=4\log T\cdot e^{-\frac{h}{4\log T}}.$
+Thus, the overall error is bounded by:
+$|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}|\leq|\beta|\,T\big(4\log T\big)e^{-\frac{h}{4\log T}}.$
+Choosing
+$h\geq 4\log T\left(\log T+\log(4\log T)+\log\frac{|\beta|}{\varepsilon}\right)$
+ensures the per-round error is at most $\varepsilon$, completing the proof. ∎
 Proof of Lemma 11.4.
-Consider any sequence and any predictor from $\prod_{\star}$, say $\pi_{1}=\pi_{A,B,C,0}$. Following equation (11.1), we can write the optimal predictor in the spectral basis:
-$\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\beta\mathbf{u}_{t}$ $=\beta\cdot\mu_{\alpha}(\mathbf{u}_{t-1:1})$
-$=\beta\cdot\mu_{\alpha}^{\top}\tilde{\mathbf{u}}_{t-1}$ (11.2)
-$=\beta\cdot\mu_{\alpha}^{\top}(\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top})\tilde{\mathbf{u}}_{t-1}$ $\mbox{basis property }\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top}=I$
-$=\sum_{i=1}^{T}M_{i}^{\star}\cdot\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}.$ $M_{i}^{\star}=\beta\phi_{i}^{\top}\mu_{\alpha}$
-By Lemma 11.4 we have that
-$M_{i}^{\star}=\beta\phi_{i}^{\top}\mu_{\alpha}=O(e^{-\frac{i}{\log T}}).$
-We proceed by an inductive claim on $\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|$.
-Consider the predictor $\pi_{2}\in\prod_{h}^{\lambda}$ for the same system matrices $M^{\star}$.
-Then we have
-$\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|-\|\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{2}}\|$
-$\leq\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}+\mathbf{y}_{t-1}^{\pi_{2}}\|$ -$\triangle\mbox{-inequality}$
-$=\left\|\sum_{i=1}^{T}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}-\sum_{i=1}^{h}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}\right\|$
-$=\left\|\sum_{i=h+1}^{T}M^{\star}_{i}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-1}\right\|$
-$\leq\sum_{i=h+1}^{T}\|M_{i}^{\star}\||\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-1}|$
-$\leq\sum_{i=h+1}^{T}\|M_{i}^{\star}\|\|\phi_{i}\|_{2}\|\tilde{\mathbf{u}}_{t-1}\|_{2}$ Cauchy-Schwartz
-$\leq\sqrt{T}\cdot\sum_{i=h+1}^{T}|M_{i}^{\star}|$ $|\mathbf{u}_{t}|\leq 1,\|\phi_{i}\|_{2}\leq 1$
-$\leq\sqrt{T}\times O\left(\int_{h}^{\infty}e^{-\frac{i}{\log T}}di\right)$ Lemma 11.4
-$\leq\frac{\varepsilon}{T}.$ choice of $h$
-Thus, since same argument exactly applies to $\|\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{2}}\|$, we have that
-$\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|\leq\sum_{t=1}^{T}\frac{\varepsilon}{T}\leq\varepsilon.$
-It remains to prove Lemma 11.4, which is the main technical lemma of this section.
-Lemma 11.4.
-Consider the scalar function $g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}$ over the interval $[0,1]$.
-First, notice that by definition of $\phi_{j}$ as the eigenvectors of $Z_{T}$, and $\sigma_{j}$ as the corresponding eigenvalues, we have
-$\int_{\alpha=0}^{1}g(\alpha)d\alpha=\int_{\alpha=0}^{1}\big(\phi_{j}^{\top}\mu_{\alpha}\big)^{2}d\alpha=\int_{\alpha=0}^{1}\phi_{j}^{\intercal}\mu_{\alpha}\mu_{\alpha}^{\intercal}\phi_{j}d\alpha=\phi_{j}^{\intercal}Z_{T}\phi_{j}=\sigma_{j}.$
-On the other hand, we claim that the function $g(\alpha)$ is L-Lipschitz for a constant value of $L$. To see this, first notice that
-$\|\mu_{\alpha}\|_{2}^{2}=\sum_{i=0}^{T-1}(1-\alpha)^{2}\alpha^{2i}\leq\frac{(1-\alpha)^{2}}{1-\alpha^{2}}=\frac{1-\alpha}{1+\alpha}\leq 1-\alpha,$
-implying that $g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}\leq\|\phi_{j}\|^{2}\|\mu_{\alpha}\|^{2}\leq 1-\alpha$. Next, we have
-$\left\|\frac{\partial}{\partial\alpha}\mu_{\alpha}\right\|_{2}^{2}$ $=\sum_{t=0}^{T-1}((1-\alpha)i\alpha^{i-1}-\alpha^{i})^{2}$
-$\leq 2(1-\alpha)^{2}\sum_{t=0}^{T-1}i^{2}\alpha^{2i-2}+2\sum_{t=1}^{T}\alpha^{2i}$ $(a+b)^{2}\leq 2(a^{2}+b^{2})$
-$\leq 2(1-\alpha)^{2}\left(\frac{1}{(1-\alpha^{2})^{2}}+\frac{\alpha^{2}}{(1-\alpha^{2})^{3}}\right)+\frac{2}{1-\alpha^{2}}$
-$\leq\frac{6}{1-\alpha},$
-where we use the identity $\sum_{i=0}^{\infty}i^{2}\beta^{i-1}=\frac{1}{(1-\beta)^{2}}+\frac{\beta}{(1-\beta)^{3}}$ with the substitution $\beta=\alpha^{2}$.
-Using this, we obtain an upper bound on the Lipschitz constant of $g(\alpha)$ as
-$|g^{\prime}{(\alpha)}|$ $=\left\lvert\frac{\partial}{\partial{\alpha}}(\mu_{\alpha}^{\top}\phi_{j})^{2}\right\rvert=2|\mu_{\alpha}^{\top}\phi_{j}|\left\lvert\frac{\partial}{\partial{\alpha}}(\mu_{\alpha}^{\top}\phi_{j})\right\rvert$
-$\leq 2\|\mu_{\alpha}\|_{2}\|\phi_{j}\|_{2}\left\lvert\frac{\partial}{\partial{\alpha}}(\mu_{\alpha}^{\top}\phi_{j})\right\rvert$ Cauchy-Schwarz
-$\leq 2\|\mu_{\alpha}\|_{2}\|\phi_{j}\|^{2}_{2}\left\|\frac{\partial}{\partial{\alpha}}\mu_{\alpha}\right\|_{2}$ Cauchy-Schwarz
-$\leq 2\sqrt{1-\alpha}\sqrt{\frac{6}{1-\alpha}}\leq 6.$ $\|\phi_{i}\|_{2}\leq 1$
-Let $g_{\max}=\max_{\alpha}g(\alpha)$ be the maximal value that $g$ attains from zero to one. Any non-negative L-Lipschitz function that reaches $g_{\max}$ and has the area of $A$, satisfies:
-$A\geq\frac{g_{\max}^{2}}{2L}\ \ \Rightarrow\ \ g_{\max}^{2}\leq{2LA},$
-since the minimal area is realized by a triangle.
-Thus, the maximum value of $g(\alpha)$ is bounded by, using Theorem 11.1,
-$g_{\max}=\max_{\alpha}\left\{(\phi_{j}^{\top}\mu_{\alpha})^{2}\right\}$ $\leq 2\cdot 6\cdot\int_{\alpha}(\phi_{j}^{\top}\mu_{\alpha})^{2}d\alpha$
-$=12\cdot\phi_{j}^{\top}\left(\int_{\alpha}\mu_{\alpha}\mu_{\alpha}^{\top}d\alpha\right)\phi_{j}$
-$=12\sigma_{j}$
-$\leq 12c_{1}e^{-\frac{c_{2}j}{\log T}}$ Theorem 11.1
+Consider the scalar function
+$g(\alpha)=(\mu_{\alpha}^{\top}\phi_{j})^{2}$
+over the interval $[0,1]$.
+First,
+$\int_{0}^{1}g(\alpha)\,d\alpha$ $=\int_{0}^{1}(\phi_{j}^{\top}\mu_{\alpha})^{2}d\alpha$
+$=\phi_{j}^{\intercal}\left(\int_{0}^{1}\mu_{\alpha}\mu_{\alpha}^{\intercal}d\alpha\right)\phi_{j}$
+$=\phi_{j}^{\intercal}Z_{T}\phi_{j}=\sigma_{j}.$
+Next, we bound the Lipschitz constant of $g$.
+We have
+$\|\mu_{\alpha}\|_{2}^{2}=\sum_{t=0}^{T-1}\alpha^{2t}\leq T,$
+and
+$\left\|\frac{\partial}{\partial\alpha}\mu_{\alpha}\right\|_{2}^{2}=\sum_{t=0}^{T-1}(t\alpha^{t-1})^{2}\leq T^{3}.$
+Thus
+$|g^{\prime}(\alpha)|$ $=2|\mu_{\alpha}^{\top}\phi_{j}|\left|\frac{\partial}{\partial\alpha}(\mu_{\alpha}^{\top}\phi_{j})\right|$
+$\leq 2\|\mu_{\alpha}\|_{2}\|\phi_{j}\|_{2}\left\|\frac{\partial}{\partial\alpha}\mu_{\alpha}\right\|_{2}$
+$\leq 2\sqrt{T}\sqrt{T^{3}}=2T^{2}.$
+Hence $g$ is $L$-Lipschitz with $L\leq 2T^{2}$.
+Let $g_{\max}=\max_{\alpha\in[0,1]}g(\alpha)$.
+Among all nonnegative $L$-Lipschitz functions attaining value $g_{\max}$, the minimal area is achieved by a triangle.
+Because the half-width of this triangle’s base is at most $g_{\max}/L\leq T/(2T^{2})=1/(2T)\leq 1/2$, the triangle cannot be truncated by both boundaries of the interval $[0,1]$ simultaneously.
+Thus, the area strictly within $[0,1]$ is at least half the total area of the unbounded triangle, yielding
+$\int_{0}^{1}g(\alpha)\,d\alpha\geq\frac{g_{\max}^{2}}{2L}.$
+Therefore,
+$g_{\max}^{2}\leq 2L\sigma_{j}\leq 4T^{2}\sigma_{j}.$
+Taking fourth roots,
+$\max_{\alpha}|\phi_{j}^{\top}\mu_{\alpha}|\leq(4T^{2}\sigma_{j})^{1/4}.$
+Using Theorem 11.1,
+$\sigma_{j}\leq 4\pi^{2}e^{-\frac{j}{\log T}},$
+which yields
+$|\phi_{j}^{\top}\mu_{\alpha}|=O\!\left(T^{1/2}e^{-\frac{j}{4\log T}}\right).$
 11.2.1 Generalization to High Dimension
-Consider a linear dynamical system with matrices $A,B,C$ and further assume that $A$ is symmetric.
+Consider a linear dynamical system with matrices $A,B,C$.
+The expression $\tilde{\mathbf{u}}_{t}$ is now a block matrix of dimensions $T\times d_{u}$, and thus $\tilde{\mathbf{u}}_{t}^{\top}\mu_{\alpha}$ is a vector of dimension $d_{u}$.
+We further assume that $A$ is symmetric.
 This implies that $A$ can be diagonalized over the set of reals as $A=U^{\top}DU$.
-Decomposing over each dimension $j\in[d]$ separately, and denoting $D_{jj}=\alpha_{j}$, we can write (11.1):
-$\mathbf{y}_{t}-\mathbf{y}_{t-1}-CB\mathbf{u}_{t-1}$
-$=\sum_{i=1}^{t-1}CU^{\top}(D^{i-1}-D^{i-2})UB\mathbf{u}_{t-1-i}$
-$=\sum_{i=1}^{t-1}CU^{\top}(\sum_{j=1}^{d_{x}}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})e_{j}e_{j}^{\top})UB\mathbf{u}_{t-i-1}$
-$=\sum_{j=1}^{d_{x}}c_{j}b_{j}^{\top}\sum_{i=1}^{t-1}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})\mathbf{u}_{t-1-i}$ $c_{j},b_{j}\leftarrow j^{th}\mbox{ row/col of }CU^{\top},UB$
-$=\sum_{j=0}^{d_{x}}\left(M^{\star}_{j}\mathbf{u}_{t}+{M}^{\star}_{j}\mu_{\alpha_{j}}(\mathbf{u}_{1:t-1})\right)$
-$=\tilde{M}^{\star}\mathbf{u}_{t}+\tilde{M}^{\star}\mu_{\alpha_{j}}(\mathbf{u}_{1:t-1})$
-Thus, we can apply the same approximation techniques from the previous section, and derive an analogous conclusion to Lemma 11.3 for dimension higher than one.
+Decomposing over each dimension $j\in[d_{x}]$ separately, and denoting $D_{jj}=\alpha_{j}$, we can write (11.1):
+$\mathbf{y}_{t}=\sum_{i=1}^{t-1}CU^{\top}D^{i-1}UB\mathbf{u}_{t-i}$
+$=\sum_{i=1}^{t-1}CU^{\top}\left(\sum_{j=1}^{d_{x}}\alpha^{i-1}_{j}e_{j}e_{j}^{\top}\right)UB\mathbf{u}_{t-i}$
+$=\sum_{j=1}^{d_{x}}c_{j}b_{j}^{\top}\sum_{i=1}^{t-1}\alpha^{i-1}_{j}\mathbf{u}_{t-i}$ $c_{j},b_{j}\leftarrow j^{th}\mbox{ col of }CU^{\top},UB$
+$=\sum_{j=1}^{d_{x}}M_{j}\big(\tilde{\mathbf{u}}_{t}^{\top}\mu_{\alpha_{j}}\big)$ $\text{where }M_{j}:=c_{j}b_{j}^{\top}\in{\mathbb{R}}^{d_{y}\times d_{u}}.$
+Thus, we can apply the same approximation techniques from the previous section, and derive an analogous conclusion to Theorem 11.3 for dimension higher than one.
 Crucially, the number of parameters required does not depend on the hidden dimension of the original system.
@@
 11.3 Online Spectral Filtering
-The prediction class $\prod_{h}^{\lambda}$ can be learned by online gradient descent (Algorithm 3) in the online convex optimization framework, which gives rise to the following algorithm.
-Figure 11.2: The filters obtained by the eigenvectors of $Z$.
-Algorithm 7 Online Spectral Filtering
+The prediction class $\Pi_{h}$ can be learned by online gradient descent in the online convex optimization framework, spelled out in Algorithm 7.
+The regret guarantees of the resulting algorithm, as given in Theorem 6.1, directly imply the following.
+Theorem 11.5.
+Let $D$ be the diameter of the parameter set $\mathcal{K}$, and $G$ an upper bound on the norm of the gradients for the loss functions $f_{t}$ as in Algorithm 7.
+For choice of step sizes $\eta_{t}=\frac{D}{G\sqrt{t}}$, we have that
+$\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\Pi_{h}}\sum_{t=1}^{T}\|\mathbf{y}_{t}-\mathbf{y}_{t}^{\pi}\|^{2}\leq 2GD\sqrt{T}.$
+Algorithm 7 Online Spectral Filtering with OGD
 1: Input: $M^{1}_{1:h}$, convex constraints set $\mathcal{K}\subseteq{\mathbb{R}}^{h\times d_{y}\times d_{u}}$, horizon $T$
-2: Compute $h$ top eigenvectors of $Z_{T}$, denoted $\phi_{1},...,\phi_{h}$.
+2: Compute $h$ top eigenvectors of $Z_{T}$, denoted $\phi_{1},\dots,\phi_{h}$.
 3:for $t=1$ to $T$ do
-4:  Predict $\hat{\mathbf{y}}_{t}=M_{0}^{t}\mathbf{u}_{t}+\sum_{j=1}^{h}M^{t}_{j}\phi_{j}(\tilde{\mathbf{u}}_{t-i})$ , as per definition 11.2, and observe true $\mathbf{y}_{t}$.
-5:  Define cost function $f_{t}(M)=\|\mathbf{y}_{t}-\tilde{\mathbf{y}}_{t}\|^{2}$.
+4:  Predict $\hat{\mathbf{y}}_{t}=\sum_{j=1}^{h}M^{t}_{j}\big(\tilde{\mathbf{u}}_{t}^{\top}\phi_{j}\big)$ and observe true $\mathbf{y}_{t}$.
+5:  Define loss $f_{t}(M)=\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}$.
 6:  Update and project:
@@
 7:end for
-The eigenvectors of $Z$ have an interesting structure: their magnitude does not decay with time, which explains their improved approximation as captured formally in Lemma 11.3.
-The regret guarantees of the online gradient descent algorithm, as given in Theorem 6.1, directly imply the following corollary.
-Theorem 11.5.
-Let $D$ be the diameter of the parameter set $\mathcal{K}$, and $G$ an upper bound on the norm of the gradients for the loss functions $f_{t}$ as in Algorithm 7.
-For choice of step sizes $\eta_{t}=\frac{D}{G\sqrt{t}}$, we have that
-$\sum_{t}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\prod_{h}^{\lambda}}\|\mathbf{y}_{t}-\mathbf{y}_{t}^{\pi}\|^{2}\leq 2GD\sqrt{T}.$
+However, the gradients of the loss can depend polynomially on $T$, since the feature vectors can have norms that grow with the horizon.
+This renders the regret guarantee of OGD uninformative in our setting.
+To address this issue, we turn instead to the Vovk-Azoury-Warmuth (VAW) forecaster from Appendix A.
+In the spectral filtering setting, the feature vector is known before the prediction is made, so the VAW protocol is a natural fit.
+For simplicity of exposition, we present the scalar setting ($d_{y}=d_{u}=1$).
+The resulting regret depends multiplicatively only on the label bound $Y$, while the dependence on the feature magnitudes appears only through a logarithmic determinant term.
+Algorithm 8 Spectral Filtering with VAW
+1:Input: Filter size $h$, $\lambda>0$, horizon $T$.
+2:Precompute: Top $h$ eigenvectors of $Z_{T}$, stored in $\Phi_{h}\in{\mathbb{R}}^{T\times h}$.
+3:Initialize: $A_{0}=\lambda I_{h}$, $v_{0}=0\in{\mathbb{R}}^{h}$.
+4:for $t=1,2,\dots,T$ do
+5:  $\tilde{\mathbf{u}}_{t}=[u_{t-1},\dots,u_{1},0,\dots,0]^{\top}$.
+6:  $h_{t}=\Phi_{h}^{\top}\tilde{\mathbf{u}}_{t}$.
+7:  Curvature update: $A_{t}=A_{t-1}+h_{t}h_{t}^{\top}$.
+8:  Predict $\hat{y}_{t}=h_{t}^{\top}A_{t}^{-1}v_{t-1}$ and observe $y_{t}$.
+9:  Update target vector: $v_{t}=v_{t-1}+y_{t}h_{t}$.
+10:end for
+Corollary 11.6 (VAW regret for spectral filtering).
+Assume $|y_{t}|\leq Y$ for all $t$, and let $\Theta\subseteq{\mathbb{R}}^{h}$ satisfy $\|\theta\|_{2}\leq B$ for all $\theta\in\Theta$. Then
+$\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}-\min_{\theta\in\Theta}\sum_{t=1}^{T}(y_{t}-\theta^{\top}h_{t})^{2}\;\leq\;\lambda B^{2}+Y^{2}\log\det\!\left(I_{h}+\frac{1}{\lambda}\sum_{t=1}^{T}h_{t}h_{t}^{\top}\right).$
+If in addition $|u_{s}|\leq 1$ for all $s$, and setting $\lambda=1/B^{2}$, gives
+$\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}-\min_{\theta\in\Theta}\sum_{t=1}^{T}(y_{t}-\theta^{\top}h_{t})^{2}\;\leq\;1+hY^{2}\log\!\left(1+B^{2}T^{2}\right).$
+Indeed, applying the VAW regret bound from Appendix A with $a_{t}=h_{t}$ and $b_{t}=y_{t}$ yields
+$\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}-\sum_{t=1}^{T}(y_{t}-\theta^{\top}h_{t})^{2}\;\leq\;\lambda\|\theta\|_{2}^{2}+Y^{2}\log\det\!\left(I_{h}+\frac{1}{\lambda}\sum_{t=1}^{T}h_{t}h_{t}^{\top}\right)$
+for every $\theta\in{\mathbb{R}}^{h}$. Taking $\theta$ to be the best predictor in $\Theta$ gives the corollary.
+Since $\Phi_{h}$ has orthonormal columns, $\|\Phi_{h}^{\top}\|_{2}=1$.
+If $|u_{s}|\leq 1$, then $\|\tilde{\mathbf{u}}_{t}\|_{2}\leq\sqrt{T}$ and hence $\|h_{t}\|_{2}\leq\sqrt{T}$.
+Therefore $\|h_{t}h_{t}^{\top}\|_{2}\leq T$, so
+$\left\|\sum_{t=1}^{T}h_{t}h_{t}^{\top}\right\|_{2}\leq T^{2},$
+which implies
+$\log\det\!\left(I_{h}+\frac{1}{\lambda}\sum_{t=1}^{T}h_{t}h_{t}^{\top}\right)\leq h\log\!\left(1+\frac{T^{2}}{\lambda}\right).$
+Even if the feature magnitudes grow polynomially with $T$, their contribution remains only logarithmic.
+Thus VAW avoids the large multiplicative constants that arise in first-order methods, and the leading multiplicative dependence is only on the label magnitude $Y$.
 11.4 Conclusion
@@
 Extensions to asymmetric system matrices, as referenced in the bibliographic section, provide a pathway toward broader applicability.
-Moreover, integrating spectral methods with control strategies may lead to novel approaches for nonstochastic control with partial observations, which we explore in the next chapter.
+Moreover, integrating spectral methods with control strategies may lead to novel approaches for nonstochastic control with partial observations, which we explore in the next part of this text.
 11.5 Bibliographic Remarks
-Deviating from classical recovery and into the world of improper prediction, the work of [HSZ17] devised an improper learning technique using the eigenvectors of the Hankel matrix $Z$.
-The methods we surveyed in previous chapters to learning linear dynamical systems had statistical and computational complexity that depended on the hidden dimension of the system.
+Deviating from classical recovery and into the world of improper prediction, the work of [HSZ17] devised an improper learning technique using the eigenvectors of the Hankel moment matrix $Z$.
+The methods we surveyed in previous chapters for learning linear dynamical systems had statistical and computational complexity that depended on the hidden dimension.
 In contrast, the spectral filtering technique studied in this chapter has regret bounds and computational complexity guarantees that do not depend on the hidden dimension.
-The spectral properties of the Hankel matrices used in this chapter are from [BT17].
-The first extension of the spectral learning technique to asymmetric matrices is due to [HLS+18].
-A more recent advancement of spectral filtering for general linear dynamical systems was recently put forth in [MH25].
+The spectral decay properties of the Hankel (Hilbert) matrices used in this chapter can be derived from the displacement-structure bounds of [BT17], and are closely connected to the classical approximation theory of Hankel operators; see, e.g., [PEL03] for a comprehensive treatment.
+Connections to realization and model-reduction theory trace back to the Ho–Kalman construction [HK66] and to balanced truncation and related approximation methods [ANT05].
+Spectral filtering can also be interpreted through the lens of operator approximation: it approximates the input–output map via projection onto leading modes of a compact Hankel operator.
+This viewpoint is closely related to Koopman/operator-theoretic approaches and data-driven spectral methods such as Dynamic Mode Decomposition [TRL+14].
+The first extension of spectral filtering beyond symmetric settings to general (asymmetric) linear dynamical systems is due to [HLS+18].
+More recent dimension-free advances for learning asymmetric linear dynamical systems were developed in [MH25].
 The technique of spectral filtering has found applications in control and sequence prediction since its introduction.
-In [AHL+18], the authors use a spectral filtering technique to concisely represent linear dynamical systems in a linear basis, formulating the optimal control problem as a convex program.
-This method circumvents the nonconvex challenges associated with explicit system identification and latent state estimation, providing provable optimality guarantees for the control signal.
-Closed-loop control using spectral filtering techniques was recently advanced in [BBDH25].
-A recent line of work makes use of the spectral basis discovered in the line of work surveyed in this chapter to design neural architectures for sequence prediction.
-This was first suggested in the context of state space models in the spectral basis in [ASCH23], and was extended to hybrid models in [LND+24].
-Later extensions explored efficient implementation [ACD+24] and the concept of length generalization in sequence prediction [MDA+24].
+In [AHL+18], the authors use a spectral filtering representation to cast control of unknown linear dynamical systems as a convex program.
+Closed-loop control using spectral filtering techniques was recently advanced in [BBD+25].
+A recent line of work makes use of spectral bases to design neural architectures for sequence prediction.
+This was first suggested in the context of spectral state space models [ASC+23], and was extended to hybrid models such as FlashSTU [LND+24].
+Later extensions explored efficient implementation [ACD+24] and the concept of length generalization [MDA+24].
+Structured state-space models such as S4 [GGR22] also exploit diagonalizations of linear recurrences, though derived from continuous-time parameterizations rather than Hankel moment structure.
+At a conceptual level, spectral filtering replaces explicit state estimation by operator approximation: instead of reconstructing latent states, it approximates the input–output map in a near-optimal spectral basis.
+This perspective unifies ideas from realization theory, compact operator approximation, and modern sequence modeling.
 11.6 Exercises
-11.1. Generalize Lemma 11.3 to all real numbers in the range [1,1][-1,1], instead of $[0,1]$.
-11.2. Generalize Lemma 11.3 to any dimension $d_{x}>1$.
+11.1. Generalize Theorem 11.3 to all real numbers in the range [1,1][-1,1], instead of $[0,1]$.
+11.2. Generalize Theorem 11.3 to any dimension $d_{x}>1$.
 11.3.
@@
 Give an example of a family of $n\times n$ non-positive semidefinite Hankel matrices that have eigenvalues with constant absolute value.
-11.4.
-The view of Theorem 13.5 is a bit misleading, as it compares the iterates of Online Gradient Descent with the hypothesis class $\Pi_{h}^{\lambda}$ and uses big-Oh notation to obscure the diameter and Lipschitz constants of the predictors.
-Assume that $d_{x}=1$.
-  1. (i)
-    Find a convex set $\mathcal{M}$ of tuples $(M_{1},M_{2},\dots)$’s to optimize over such that its diameter is some $\textrm{polylog}(T,1/\varepsilon)$.
-The norm we measure diameter with respect to is the sum of the Frobenius norms of the $M_{i}$’s.
-  2. (ii)
-    In addition, show that the norm of the gradient is at most $\textrm{polylog}(T,1/\varepsilon)$ as well.
-Using these two facts, show that with the right value of $\varepsilon$, the regret against the hypothesis class $\Pi_{\star}$ is at most $\tilde{O}(\sqrt{T})$.
+11.4. The trace of a matrix is the sum of its eigenvalues, which measures the “total energy” of the system.
+  1. 1.
+    Using the integral definition $Z_{T}=\int_{0}^{1}\mu_{\alpha}\mu_{\alpha}^{\top}d\alpha$, show that the $(i,j)$-th entry of $Z_{T}$ (where indices $i,j$ run from 0 to T1T-1) is given by:
+    $(Z_{T})_{i,j}=\frac{1}{i+j+1}.$
+  2. 2.
+    Compute the trace of $Z_{T}$ as a summation.
+  3. 3.
+    Show that $\text{Tr}(Z_{T})\approx\frac{1}{2}\ln T$.
+Note: This contrasts with the operator norm $\|Z_{T}\|_{2}$, which is bounded by $\pi$.
+The fact that the trace grows logarithmically while the norm is bounded implies that the vast majority of eigenvalues must be extremely small, reinforcing the motivation for spectral filtering.
+11.5. Consider the simplest non-trivial case where the horizon is $T=2$.
+  1. 1.
+    Compute the matrix $Z_{2}$ explicitly using the formula derived in the previous exercise.
+  2. 2.
+    Calculate the two eigenvalues of $Z_{2}$.
+  3. 3.
+    Calculate the ratio $\lambda_{1}/\lambda_{2}$. What does this ratio tell you about how much information is captured by the first spectral feature compared to the second?
+11.6.
+In spectral filtering, we learn a parameter vector $\theta\in{\mathbb{R}}^{h}$ such that our prediction is $\hat{y}_{t}=\theta^{\top}(\Phi_{h}^{\top}\tilde{\mathbf{u}}_{t})$, where $\Phi_{h}\in{\mathbb{R}}^{T\times h}$ contains the top $h$ eigenvectors of $Z_{T}$.
+Standard linear predictors (FIR filters) use a weight vector $\mathbf{w}\in{\mathbb{R}}^{T}$ such that $\hat{y}_{t}=\mathbf{w}^{\top}\tilde{\mathbf{u}}_{t}$.
+  1. 1.
+    Find the expression for the equivalent FIR weights $\mathbf{w}_{\text{spec}}$ implied by the learned spectral parameters $\theta$.
+  2. 2.
+    If $h=T$ (we use all eigenvectors), can we represent any arbitrary FIR filter $\mathbf{w}$ using spectral parameters $\theta$? Why or why not?
+Chapter 12 Luenberger Observers
+In Chapter 11 we developed spectral filtering for learning linear dynamical systems (LDS) without dependence on the hidden state dimension.
+However, those guarantees relied implicitly on favorable spectral structure.
+In particular, systems with symmetric or real spectrum are significantly easier to learn than fully asymmetric systems with complex eigenvalues near the unit circle.
+In this chapter we address the learning of asymmetric linear dynamical systems.
+We begin by showing that without spectral restrictions, learning can be information-theoretically hard.
+We then introduce the Luenberger observer as a mechanism for spectrally reshaping the observer error dynamics.
+This leads to a regret bound governed by a new quantity: the Luenberger complexity $Q_{\star}$.
+12.1 Lower Bounds for Learning Asymmetric LDS
+We first show that without spectral restrictions, learning an LDS may require a number of samples proportional to the hidden dimension.
+Theorem 12.1.
+For any online learning algorithm $\mathcal{A}$, there exists a linear dynamical system with hidden state dimension $d_{h}$ such that the expected cumulative squared loss of $\mathcal{A}$ over $T\geq d_{h}$ rounds is at least $\Omega(d_{h})$.
+Proof.
+We construct an autonomous LDS for which prediction is equivalent to predicting independent coin flips.
+Consider the system:
+$x_{t+1}$ $=Ax_{t},$
+$y_{t}$ $=Cx_{t},$
+where $A$ is the $d_{h}\times d_{h}$ cyclic permutation matrix
+$A=\begin{pmatrix}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\\ 1&0&0&\cdots&0\end{pmatrix}$
+and $C=(1,0,\dots,0)$.
+Let the initial state $x_{0}\in\{-1,1\}^{d_{h}}$ be drawn uniformly at random, with independent Rademacher entries.
+Then $x_{t}=A^{t}x_{0}$, and for $t=0,\dots,d_{h}-1$,
+$y_{t}=(x_{0})_{t+1}.$
+Thus each observation reveals a fresh independent coordinate of $x_{0}$.
+Fix any realization of the learner’s internal randomness. Conditional on this realization, the prediction at time $t$ has the form
+$\hat{y}_{t}=f_{t}(y_{0},\dots,y_{t-1}),$
+which depends only on $(x_{0})_{1},\dots,(x_{0})_{t}$.
+But $y_{t}=(x_{0})_{t+1}$ is independent of these values, hence
+$\mathbb{E}[y_{t}\mid y_{0},\dots,y_{t-1}]=0.$
+Therefore,
+$\mathbb{E}[(y_{t}-\hat{y}_{t})^{2}]$ $=\mathbb{E}[y_{t}^{2}-2y_{t}\hat{y}_{t}+\hat{y}_{t}^{2}]$
+$=1+\mathbb{E}[\hat{y}_{t}^{2}]$
+$\geq 1.$
+Summing over $t=0,\dots,d_{h}-1$ yields expected cumulative loss at least $d_{h}$. ∎
+This lower bound shows that without restricting spectral structure, learning asymmetric systems may require a number of samples proportional to the hidden dimension.
+We now show that this hardness can be overcome when the system admits an observer representation that reshapes the error dynamics.
+12.2 Pole Placement
+In this section we prove a fundamental and elegant result from classical control theory: the eigenstructure of the observer error dynamics can be modified completely under the observability assumption.
+The key structural condition that allows this radical shift in eigenstructure is observability, as defined earlier in the book. Recall that $(A,C)$ is observable if the observability matrix
+$\mathcal{O}_{d_{h}}=\begin{pmatrix}C\\ CA\\ \vdots\\ CA^{d_{h}-1}\end{pmatrix}$
+has full column rank. Equivalently, the only state $x$ satisfying $CA^{k}x=0$ for all $k\geq 0$ is $x=0$. Observability ensures that the hidden state can be completely reconstructed from the outputs.
+The following result shows that this same condition allows for the arbitrary assignment of the spectrum of the observer error dynamics.
+Theorem 12.2 (Observer Pole Placement).
+Let $(A,C)$ be observable.
+Then for any self-conjugate set of $d_{h}$ complex numbers $\Lambda$, there exists an observer gain matrix $L\in\mathbb{R}^{d_{h}\times d_{y}}$ such that the spectrum (the set of eigenvalues) of ALCA-LC is exactly $\Lambda$.
+Proof.
+For clarity we provide a constructive, self-contained proof for the single-output case ($d_{y}=1$), where $C$ is a $1\times d_{h}$ row vector and $L$ is a $d_{h}\times 1$ column vector.
+Let the characteristic polynomial of the unobserved dynamics $A$ be
+$p(z)=\det(zI-A)=z^{d_{h}}+a_{d_{h}-1}z^{d_{h}-1}+\dots+a_{1}z+a_{0}.$
+Because the pair $(A,C)$ is observable, there exists an invertible change-of-basis matrix $T$ such that in the new coordinates $x=T\bar{x}$, the system matrices $\bar{A}=T^{-1}AT$ and $\bar{C}=CT$ are in Observable Canonical Form:
+$\bar{A}=\begin{pmatrix}0&0&\dots&0&-a_{0}\\ 1&0&\dots&0&-a_{1}\\ 0&1&\dots&0&-a_{2}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dots&1&-a_{d_{h}-1}\end{pmatrix},\quad\bar{C}=\begin{pmatrix}0&0&\dots&0&1\end{pmatrix}$.
+Let the desired target spectrum $\Lambda$ be defined by the roots of a monic polynomial:
+$q(z)=z^{d_{h}}+\tilde{a}_{d_{h}-1}z^{d_{h}-1}+\dots+\tilde{a}_{1}z+\tilde{a}_{0}.$
+We wish to find an observer gain in the transformed coordinates, $\bar{L}=(l_{0},l_{1},\dots,l_{d_{h}-1})^{\top}$, such that $\bar{A}-\bar{L}\bar{C}$ has characteristic polynomial $q(z)$.
+Observe the structure of the outer product $\bar{L}\bar{C}$: it is a $d_{h}\times d_{h}$ matrix of all zeros, except for its last column which equals exactly $\bar{L}$.
+Therefore, subtracting it from $\bar{A}$ simply modifies the last column:
+$\bar{A}-\bar{L}\bar{C}=\begin{pmatrix}0&0&\dots&0&-(a_{0}+l_{0})\\ 1&0&\dots&0&-(a_{1}+l_{1})\\ 0&1&\dots&0&-(a_{2}+l_{2})\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dots&1&-(a_{d_{h}-1}+l_{d_{h}-1})\end{pmatrix}$.
+Because this resulting matrix is also in canonical form, its characteristic polynomial can be read directly from its last column:
+$\det(zI-(\bar{A}-\bar{L}\bar{C}))=z^{d_{h}}+(a_{d_{h}-1}+l_{d_{h}-1})z^{d_{h}-1}+\dots+(a_{0}+l_{0}).$
+To assign the spectrum $\Lambda$, we simply match the coefficients to $q(z)$ by choosing $l_{i}=\tilde{a}_{i}-a_{i}$ for all $i=0,\dots,d_{h}-1$.
+Finally, we transform the gain back to the original coordinates by setting $L=T\bar{L}$.
+The closed-loop matrix is $A-LC=T(\bar{A}-\bar{L}\bar{C})T^{-1}$, which is similar to $\bar{A}-\bar{L}\bar{C}$ and therefore possesses the exact assigned spectrum $\Lambda$.
+∎
+Constructing the observable canonical form.
+In the single-output case ($d_{y}=1$), the observability matrix
+$\mathcal{O}_{d_{h}}(A,C)=\begin{pmatrix}C\\ CA\\ \vdots\\ CA^{d_{h}-1}\end{pmatrix}$
+is a $d_{h}\times d_{h}$ matrix.
+Since $(A,C)$ is observable, it has full rank and is therefore invertible.
+Likewise, the observability matrix of the observable companion pair $(A_{\mathrm{obs}},C_{\mathrm{obs}})$ is invertible.
+Because under the coordinate change $x=T\bar{x}$ we have
+$\mathcal{O}_{d_{h}}(\bar{A},\bar{C})=\mathcal{O}_{d_{h}}(A,C)\,T,$
+we may compute the similarity transform from
+$\mathcal{O}_{d_{h}}(A,C)\,T=\mathcal{O}_{d_{h}}(A_{\mathrm{obs}},C_{\mathrm{obs}}),\qquad\text{so}\qquad T=\mathcal{O}_{d_{h}}(A,C)^{-1}\mathcal{O}_{d_{h}}(A_{\mathrm{obs}},C_{\mathrm{obs}}).$
+Output coordinates and dynamical interpretation.
+Still in the single-output case, there is a second coordinate change, closely related to observability, that makes the output dynamics transparent. Define
+$z_{t}:=\mathcal{O}_{d_{h}}(A,C)x_{t}=\begin{pmatrix}y_{t}\\ y_{t+1}\\ \vdots\\ y_{t+d_{h}-1}\end{pmatrix}$.
+Because $(A,C)$ is observable, $\mathcal{O}_{d_{h}}(A,C)$ is invertible, so this is a valid state coordinate. In these coordinates the dynamics become
+$z_{t+1}=\begin{pmatrix}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \vdots&&\ddots&\ddots&\vdots\\ 0&0&\cdots&0&1\\ -a_{0}&-a_{1}&\cdots&-a_{d_{h}-2}&-a_{d_{h}-1}\end{pmatrix}z_{t},\qquad y_{t}=e_{1}^{\top}z_{t}.$
+The last row is exactly the Cayley–Hamilton relation, so unrolling the recursion yields
+$y_{t+d_{h}}+a_{d_{h}-1}y_{t+d_{h}-1}+\cdots+a_{1}y_{t+1}+a_{0}y_{t}=0.$
+Thus an observable $d_{h}$-dimensional autonomous linear system may be viewed as a degree-$d_{h}$ autoregressive model: in suitable coordinates, the state is encoded by a block of $d_{h}$ consecutive outputs, and the dynamics are a shift together with a linear recurrence.
+From this perspective, pole placement simply changes the coefficients of that recurrence, and hence its roots.
+We proceed to show how pole placement can be used to create the Luenberger observer system: a powerful technique in learning of dynamical systems.
+12.3 The Luenberger Observer
+Consider the linear dynamical system
+$x_{t+1}=Ax_{t}+Bu_{t},\qquad y_{t}=Cx_{t}.$
+Definition 12.3 (Luenberger Observer).
+Given a gain matrix $L\in\mathbb{R}^{d_{h}\times d_{y}}$, the Luenberger observer is the dynamical system
+$\tilde{x}_{t+1}=A\tilde{x}_{t}+Bu_{t}+L(y_{t}-\tilde{y}_{t}),\qquad\tilde{y}_{t}=C\tilde{x}_{t}.$
+Define the estimation error
+$e_{t}:=x_{t}-\tilde{x}_{t}.$
+Subtracting the observer dynamics from the true dynamics yields
+$e_{t+1}$ $=x_{t+1}-\tilde{x}_{t+1}$
+$=(Ax_{t}+Bu_{t})-(A\tilde{x}_{t}+Bu_{t}+L(y_{t}-\tilde{y}_{t}))$
+$=(A-LC)e_{t}.$
+Thus the observer error evolves autonomously under the matrix ALCA-LC, entirely independent of the control inputs $u_{t}$.
+By Theorem 12.2, if $(A,C)$ is observable then $L$ can be chosen so that the spectrum of ALCA-LC matches any desired target spectrum.
+We define the resulting closed-loop matrix as $\tilde{A}:=A-LC$.
+12.3.1 Luenberger Complexity
+Although pole placement guarantees that we may assign the eigenvalues of $\tilde{A}$ arbitrarily, the transient behavior of the observer depends not only on the spectrum, but also on the conditioning of the eigenbasis.
+We henceforth suppose that the chosen closed-loop matrix $\tilde{A}=A-LC$ is diagonalizable.
+For any diagonalization $\tilde{A}=HDH^{-1}$, the factor $\|H\|\,\|H^{-1}\|$ measures the conditioning of the eigenbasis.
+Definition 12.4 (Luenberger Complexity).
+For a strictly stable, diagonalizable matrix $\tilde{A}$, let $\gamma=1-\rho(\tilde{A})>0$ be its stability margin. Define
+$\kappa_{\mathrm{diag}}(\tilde{A})=\min\{\|H\|\,\|H^{-1}\|:\tilde{A}=HDH^{-1},\ D\text{ diagonal}\},$
+and its Luenberger complexity by
+[…]
+Thus Q(A~)Q($\tilde${A}) measures the transient amplification and time scale required before contraction dominates. By definition, $Q(\tilde{A})\geq 1/\gamma$.
+Theorem 12.5 (Error Contraction and Burn-In).
+Let $(A,C)$ be observable, choose $L$, and let $\tilde{A}=A-LC$. Assume $\tilde{A}$ is strictly stable and diagonalizable with stability margin $\gamma=1-\rho(\tilde{A})>0$. Then
+$\|e_{t}\|\leq\kappa_{\mathrm{diag}}(\tilde{A})\,(1-\gamma)^{t}\,\|e_{0}\|.$
+Consequently, for any $\varepsilon>0$, to guarantee $\|e_{t}\|\leq\varepsilon$ it suffices that
+$t\geq Q(\tilde{A})\log\frac{1+\|e_{0}\|}{\varepsilon}.$
+Proof.
+[…]
+$\tilde{A}^{t}=HD^{t}H^{-1}.$
+Thus
+$\|\tilde{A}^{t}\|\leq\|H\|\,\|H^{-1}\|\,\|D^{t}\|\leq\kappa_{\mathrm{diag}}(\tilde{A})\,(1-\gamma)^{t}.$
+Multiplying by $\|e_{0}\|$ yields the first inequality. The burn-in bound follows by solving
+$\kappa_{\mathrm{diag}}(\tilde{A})(1-\gamma)^{t}\|e_{0}\|\leq\varepsilon,$
+and using the standard inequality $\log(1-\gamma)\leq-\gamma$. ∎
+Pole placement ensures existence of a stable observer.
+However, the effective convergence speed depends on both the spectral gap $\gamma$ and the conditioning of the eigenbasis (which is constrained because we can only assign the eigenvalues, not the exact matrix $\tilde{A}$).
+The Luenberger complexity Q(A~)Q($\tilde${A}) captures precisely the scale that governs this burn-in time.
+12.3.2 One Dimensional Example
+Consider the simple one dimensional (and thus fully observable) linear dynamical system
+$x_{t+1}=2x_{t},\qquad y_{t}=x_{t}.$
+Define
+$\tilde{x}_{t+1}=2\tilde{x}_{t}+L(x_{t}-\tilde{x}_{t}).$
+Then
+$e_{t+1}=(2-L)e_{t}.$
+Choosing $L=2.5$ yields
+$e_{t+1}=-0.5e_{t},$
+so the error contracts exponentially even though the system diverges.
+12.4 Reduction to Stable Learning
+We now exploit the reshaping principle as a general reduction for online learning.
+Section 12.3 showed that if $(A,C)$ is observable, we may choose an observer gain $L$ so that the error dynamics evolve under a strictly stable closed-loop matrix $\tilde{A}=A-LC$.
+The effective transient of this observer is governed by the Luenberger complexity Q(A~)Q($\tilde${A}).
+Crucially, this transforms the problem of predicting an unstable, partially-observed system with infinite memory into the problem of predicting the outputs of a strictly stable, well-behaved system.
+12.4.1 A Learnable Spectral Region
+To guarantee that the observer forgets its initial estimation error $e_{0}$, we must ensure that the closed-loop matrix $\tilde{A}$ is strictly stable.
+Fix a desired stability margin $\gamma\in(0,1)$ and define the target spectral region simply as the strictly stable complex disk:
+$\Sigma_{\gamma}:=\mathbb{D}_{1-\gamma}=\{z\in\mathbb{C}:|z|\leq 1-\gamma\}.$
+We enforce the margin $\gamma$ entirely to guarantee rapid error dissipation (burn-in).
+Marginally stable modes (e.g., $|\lambda|=1$) are strictly excluded.
+Because the Luenberger observer is a closed-loop system, an eigenvalue of $1$ would cause the unforced initial estimation error $e_{t}=\tilde{A}^{t}e_{0}$ to persist indefinitely as a constant offset.
+Because online learning algorithms build predictive features solely from forced inputs and outputs, they cannot accurately predict this unforced hidden offset, resulting in linear regret.
+All eigenvalues of the original system $A$ — in particular unstable modes or complex eigenvalues near the unit circle — constitute the difficult modes; we aim to move them strictly into the stable region $\Sigma_{\gamma}$ via the observer.
+12.4.2 The Luenberger Program
+Pole placement guarantees feasibility under observability, but not necessarily good conditioning.
+Even if there exists an $L$ such that the spectrum satisfies $\sigma(A-LC)\subseteq\Sigma_{\gamma}$, the resulting observer may have a poorly conditioned eigenbasis, leading to a large Luenberger complexity.
+We therefore define an optimization program to find the best-conditioned observer among all feasible pole placements for an arbitrary stable target set $\Sigma$.
+Definition 12.6 (Luenberger Program).
+Fix an arbitrary strictly stable spectral constraint set $\Sigma\subseteq\mathbb{D}_{1-\varepsilon}$ for some $\varepsilon>0$. Define the optimal Luenberger complexity as:
+$Q_{\star}(A,C;\Sigma):=\min_{L}\;Q(A-LC)\quad\text{s.t.}\quad\sigma(A-LC)\subseteq\Sigma,$
+with the convention $Q_{\star}(A,C;\Sigma)=+\infty$ if infeasible.
+For our specific stable disk $\Sigma=\Sigma_{\gamma}$, we write $Q_{\star}^{\gamma}$.
+We call $Q_{\star}$ the Optimal Luenberger Complexity.
+12.4.3 Regret Reduction
+Assume $Q_{\star}(A,C;\Sigma_{\gamma})<\infty$ and let $L^{\star}$ attain the minimum in the Luenberger program. Define the optimal observer matrix
+[…]
+By recursively expanding the observer state, we can express the true system output $y_{t}$ entirely as a function of the past observations $y_{t-i}$ and controls $u_{t-i}$, plus an unforced error term:
+$y_{t}=\underbrace{C\tilde{A}^{t}\tilde{x}_{0}+\sum_{i=1}^{t}C\tilde{A}^{i-1}(Bu_{t-i}+L^{\star}y_{t-i})}_{=:\bar{y}_{t}}+C\tilde{A}^{t}e_{0}.$
+The sequence $\bar{y}_{t}$ represents a strictly stable mapping from past data to the current output.
+Because $\tilde{A}$ is strictly stable, the sequence of coefficient matrices $C\tilde{A}^{i-1}B$ and $C\tilde{A}^{i-1}L^{\star}$ decay to zero exponentially fast.
+This provides a powerful, general reduction for online learning in dynamical systems: we have reduced the problem to learning a stable transfer function.
+Because the coefficients decay exponentially fast, the mapping can be effectively approximated using a finite memory length $m=O(Q_{\star}\log T)$.
+We may now employ any suitable online sequence prediction algorithm to learn this mapping.
+For example:
+  • Linear Regression (FIR Filters): We can parameterize the predictor as a standard finite impulse response filter of length $m$ and learn the coefficients using Online Gradient Descent.
+  • Spectral Filtering & ONS: We can project the observations onto the highly efficient Hankel basis (which requires only $O(\log^{2}T)$ features for real modes) and use the Online Newton Step to achieve logarithmic regret.
+Regardless of the chosen learning algorithm, the total regret against the optimal observer is bounded by the algorithm’s inherent regret for stable sequence prediction, plus an initial, bounded penalty from the Luenberger burn-in period.
+The true observation $y_{t}$ differs from the idealized stable prediction $\bar{y}_{t}$ only by the unforced observer error $C\tilde{A}^{t}e_{0}$.
+By Theorem 12.5, the total squared error introduced by the observer’s transient burn-in over all time is bounded by a constant independent of $T$:
+$E_{\text{burn-in}}:=\sum_{t=1}^{T}\|y_{t}-\bar{y}_{t}\|^{2}$ $\leq\|C\|^{2}\kappa_{\mathrm{diag}}(\tilde{A})^{2}\|e_{0}\|^{2}\sum_{t=1}^{\infty}(1-\gamma)^{2t}$
+$\leq\frac{\|C\|^{2}\kappa_{\mathrm{diag}}(\tilde{A})^{2}\|e_{0}\|^{2}}{2\gamma}.$
+[…]
+Let $\mathcal{F}_{\gamma}$ denote any class of stable linear predictors with margin $\gamma$ that contains the stable surrogate sequence $\bar{y}_{t}$ defined above.
+Theorem 12.7 (Regret Reduction).
+Let $\mathcal{A}$ be an online learning algorithm that guarantees $\mathrm{Regret}_{\mathcal{A}}(T)$ against the class $\mathcal{F}_{\gamma}$.
+If we apply $\mathcal{A}$ to the sequence of observations from the partially observed system $(A,C)$, then
+$\sum_{t=1}^{T}\|y_{t}-\hat{y}_{t}\|^{2}\leq\inf_{f\in\mathcal{F}_{\gamma}}\sum_{t=1}^{T}\|y_{t}-f_{t}\|^{2}+\mathrm{Regret}_{\mathcal{A}}(T)\leq\mathrm{Regret}_{\mathcal{A}}(T)+E_{\mathrm{burn-in}}(L^{\star}).$
+This reduction completely separates the control problem (using observability to reshape the error dynamics and minimize $Q_{\star}$) from the learning problem (running an algorithm on a stable system).
+The regret against the true underlying (potentially unstable and asymmetric) system is therefore governed by the regret of learning a stable system, plus a constant burn-in penalty.
+12.5 Summary
+This chapter shows that partial observation fundamentally changes the learning problem for linear dynamical systems.
+Without additional structure, one cannot hope for dimension-free guarantees in general: the lower bound demonstrates that a learner may be forced to uncover the hidden state one coordinate at a time, incurring loss of order $d_{h}$ before the system becomes predictable.
+Thus the hidden dimension is a genuine barrier if we work directly with the original dynamics.
+The Luenberger observer provides a way around this obstacle. When $(A,C)$ is observable, we may introduce an output-injection matrix $L$ and study the estimation error dynamics
+$e_{t+1}=(A-LC)e_{t}.$
+Pole placement shows that the spectrum of this closed-loop matrix can be moved into any prescribed stable region.
+This does not by itself solve the learning problem, because fast asymptotic decay is not enough: one must also control transient growth.
+That is the role of the Luenberger complexity Q(A~)Q($\tilde${A}), and of the optimized quantity $Q_{\star}$, which measure how quickly the observer forgets its initial error and how badly the non-normal geometry can amplify that error before contraction dominates.
+Once a well-conditioned observer is chosen, the original partially observed system is converted into a stable input-output prediction problem, up to a burn-in term coming from the decaying initial estimation error.
+The chapter therefore separates into two clean pieces.
+Control theory supplies the observer that reshapes the error dynamics into a stable form, and online learning supplies an algorithm for the resulting stable sequence-prediction task.
+In this way, the effective difficulty of learning is governed not directly by the hidden dimension $d_{h}$, but by the best attainable observer geometry encoded by $Q_{\star}$.
+12.6 Bibliographic Remarks
+The fundamental hardness of learning linear dynamical systems without spectral assumptions, as discussed in the beginning of this chapter, contrasts with the symmetric or stable cases.
+While [HMR18] showed that gradient descent efficiently learns stable systems, the lower bounds for general asymmetric systems typically rely on constructions where the system dynamics obscure the initial state for long periods, similar to the hardness results in [SMT+18].
+The information theoretic lower bound described here is due to [HM26].
+The theoretical foundation for reconstructing internal states from outputs lies in the work of [LUE64], who introduced the concept of the state observer.
+The specific architecture discussed in this chapter is widely known as the Luenberger Observer.
+[LUE71] provides a comprehensive introduction to this theory.
+The powerful result that observability allows for arbitrary assignment of the closed-loop spectrum (Theorem 12.2) is known as the Pole Placement Theorem, originally established by [WON67].
+While classical control theory focuses on asymptotic stability, this chapter emphasizes the transient behavior of the error dynamics, which is crucial for finite-time learning guarantees.
+The study of transient growth in stable but non-normal matrices (where eigenvectors are ill-conditioned) is famously treated in the text on pseudospectra by [TE05].
+The concept of Luenberger Complexity ($Q_{\star}$) serves as a quantitative measure of this transient behavior specifically optimized for the learning setting.
+It was introduced in [DBH25].
+The reduction framework presented in this chapter, which utilizes the Luenberger observer to transform an unstable or asymmetric learning problem into a stable prediction task, is based on the work of [DBH25].
+This work showed that dimension-free regret bounds are achievable for general linear systems provided one can find a well-conditioned observer, shifting the complexity burden from the state dimension $d_{h}$ to the spectral conditioning of the dynamics.
+12.7 Exercises
+12.1.
+This exercise illustrates the "peaking phenomenon," where placing eigenvalues far into the stable region can cause the error norm to grow transiently before decaying (i.e., high Luenberger complexity).
+Consider the two-dimensional system with $A=\begin{bmatrix}1&1\\ 0&1\end{bmatrix}$ and $C=\begin{bmatrix}1&0\end{bmatrix}$.
+  1. 1.
+    Verify that the system is observable.
+  2. 2.
+    Find the observer gain vector $L=[l_{1},l_{2}]^{\top}$ such that the closed-loop matrix $\tilde{A}=A-LC$ has both eigenvalues equal to some $\lambda\in[0,1)$.
+  3. 3.
+    Show that as we demand faster convergence (i.e., as $\lambda\to 0$), the norm of the gain $\|L\|$ tends to infinity.
+  4. 4.
+    Calculate the matrix $\tilde{A}$ explicitly for $\lambda=0$. Compute $\|\tilde{A}e_{0}\|$ for an initial error $e_{0}=[0,1]^{\top}$. What happens to the error magnitude in the first step?
+12.2. A deadbeat observer is one where the closed-loop poles are all placed at the origin (0).
+  1. 1.
+    Let $\tilde{A}=A-LC$. Prove that if all eigenvalues of $\tilde{A}$ are 0, then $\tilde{A}$ is nilpotent, meaning $\tilde{A}^{d_{h}}=0$.
+  2. 2.
+    Conclude that for a deadbeat observer, the estimation error $e_{t}$ becomes exactly zero for all $t\geq d_{h}$, regardless of the initial error.
+  3. 3.
+    Why might we not always choose a deadbeat observer in practice, particularly in the presence of observation noise $\mathbf{v}_{t}$? (Hint: Consider the magnitude of $L$ derived in Exercise 12.7).
+12.3.
+In Theorem 12.1, we used a cyclic permutation matrix to prove a lower bound on learning.
+Let $A$ be the $d_{h}\times d_{h}$ cyclic permutation matrix where $A_{i,i+1}=1$ for $1\leq i<d_{h}$, $A_{d_{h},1}=1$, and 0 elsewhere.
+Let $C=[1,0,\dots,0]$.
+  1. 1.
+    Write down the observability matrix $\mathcal{O}_{d_{h}}$ for this pair $(A,C)$.
+  2. 2.
+    Prove that this system is fully observable.
+  3. 3.
+    Since it is observable, Theorem 12.2 implies we can stabilize it. Find a gain vector $L$ such that ALCA-LC is nilpotent.
+12.4.
+Prove that the set of stable matrices is not convex.
+specifically, give two matrices $A_{1},A_{2}\in\mathbb{R}^{2\times 2}$ such that $\rho(A_{1})<1$ and $\rho(A_{2})<1$, but $\rho(\frac{1}{2}A_{1}+\frac{1}{2}A_{2})>1$.
+Note: This geometric fact implies that the Luenberger Program (Definition 12.6), which optimizes over stable matrices, involves a non-convex constraint set.
+Chapter 13 Learning in Nonlinear Dynamical Systems
+In the previous chapters, we developed a comprehensive theory for learning and predicting in linear dynamical systems.
+We moved from symmetric systems in Chapter 11 to general, asymmetric, and potentially unstable linear systems in Chapter 12 via the Luenberger observer framework.
+However, the physical world is rarely strictly linear. In this chapter, we address the fundamental problem of learning nonlinear dynamical systems. Formally, we consider a system of the form
+$\mathbf{x}_{t+1}$ $=f(\mathbf{x}_{t})$ (13.1)
+$\mathbf{y}_{t}$ $=h(\mathbf{x}_{t}),$
+where $f:\mathcal{X}\to\mathcal{X}$ is the nonlinear transition function and $h:\mathcal{X}\to{\mathbb{R}}^{d_{y}}$ is the observation function.
+The state $\mathbf{x}_{t}$ is hidden, and the functions $f$ and $h$ are unknown.
+For clarity, we treat the autonomous case in this chapter; the controlled extension $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})$ is significantly more complex.
+Historically, learning such systems involved system identification: attempting to approximate the function $f$ using neural networks or kernels.
+In this chapter, we take a different approach rooted in improper learning.
+Rather than trying to recover the nonlinearity explicitly, we compete with a linear surrogate that predicts well.
+More precisely, we will show two things.
+First, every bounded nonlinear system with non-expansive dynamics admits a finite-dimensional approximate linear representation over any fixed horizon via discretization.
+Second, if such a lifted linear approximation admits a stable, well-conditioned observer, then the learning guarantees from Chapter 12 allow us to compete with it.
+Thus, nonlinear prediction can be reduced to linear prediction, provided the corresponding lifted system has moderate observer complexity.
+A passive prediction viewpoint.
+A useful way to think about the learning problem in this chapter is as passive sequential prediction.
+Here the learner does not act on the system and does not choose any control input.
+Instead, it only observes a sequence generated by hidden nonlinear dynamics and attempts to predict the next observation.
+Example 1: Lorenz dynamics.
+A canonical example is the Lorenz system, given in continuous time by
+$\dot{x}^{(1)}(\tau)$ $=\sigma\bigl(x^{(2)}(\tau)-x^{(1)}(\tau)\bigr),$
+$\dot{x}^{(2)}(\tau)$ $=x^{(1)}(\tau)\bigl(\rho-x^{(3)}(\tau)\bigr)-x^{(2)}(\tau),$
+$\dot{x}^{(3)}(\tau)$ $=x^{(1)}(\tau)x^{(2)}(\tau)-\beta x^{(3)}(\tau),$
+where $\sigma,\rho,\beta>0$ are fixed parameters. Writing
+$\mathbf{x}(\tau)=\begin{bmatrix}x^{(1)}(\tau)\\ x^{(2)}(\tau)\\ x^{(3)}(\tau)\end{bmatrix}\in{\mathbb{R}}^{3},$
+this system may be viewed abstractly as
+$\dot{}\mathbf{x}(\tau)=F(\mathbf{x}(\tau)).$
+Fix a sampling interval $\Delta>0$, and define the discrete-time state
+$\mathbf{x}_{t}=\mathbf{x}(t\Delta).$
+Then the sampled trajectory evolves according to an autonomous nonlinear dynamical system
+$\mathbf{x}_{t+1}=f(\mathbf{x}_{t}),$
+where $f$ is the time-$\Delta$ flow map of the differential equation. If we observe only a scalar signal
+$\mathbf{y}_{t}=h(\mathbf{x}_{t}),$
+for example one coordinate of the state, then the learning problem is to predict the next observation $\mathbf{y}_{t}$ from the past observations $\mathbf{y}_{1},\ldots,\mathbf{y}_{t-1}$.
+Thus a nonlinear physical system gives rise naturally to the passive prediction problem studied in this chapter.
+Example 2: large language models.
+A large language model can be viewed in the same way.
+Let $\mathbf{y}_{t}\in V$ be the token emitted at time $t$, where $V$ is a finite vocabulary.
+The model maintains a hidden state $\mathbf{x}_{t}$ that summarizes its internal activations together with the current context, cache, and recent tokens.
+The evolution of this hidden state is governed by the nonlinear dynamics of the underlying deep neural network, and the emitted token is a function of the current state.
+Abstractly, this may be written as
+$\mathbf{x}_{t+1}=f(\mathbf{x}_{t}),\qquad\mathbf{y}_{t}=h(\mathbf{x}_{t}),$
+where $f$ is induced by one forward pass of the model together with the decoding rule.
+In the deterministic case, for example under greedy decoding, this fits exactly into the autonomous framework studied in this chapter.
+If the decoding rule is randomized, for example when the next token is sampled from a truncated distribution such as top-$k$ or nucleus sampling, then the dynamics become stochastic rather than deterministic.
+In that case one obtains a stochastic analogue of the same framework; see Exercise 13.5.
+In the sequel we return to the Euclidean observation model $\mathbf{y}_{t}\in{\mathbb{R}}^{d_{y}}$.
+The language-model example is included as motivation; one may embed discrete tokens in this framework, for example, via one-hot or probability-vector representations.
+13.1 Global Linear Approximation via Discretization
+The core mathematical insight enabling linear learning of nonlinear dynamics is that nonlinear dynamics in a finite-dimensional space can be lifted to linear dynamics in a higher, potentially infinite-dimensional space.
+This is often associated with the Koopman operator, which we discuss in Section 13.2.
+We begin with a concrete finite-horizon construction based on discretization.
+13.1.1 The $\varepsilon$-Net Construction
+We make the following standard assumptions about the regularity of the system:
+  1. 1.
+    Boundedness: The state space $\mathcal{X}$ is contained in a ball of radius $R$.
+  2. 2.
+    Lipschitz continuity: The dynamics $f$ and observation map $h$ are $1$-Lipschitz. That is,
+    $\|f(x)-f(z)\|\leq\|x-z\|,\qquad\|h(x)-h(z)\|\leq\|x-z\|$
+    for all $x,z\in\mathcal{X}$.
+Let $\mathcal{S}$ be an $\varepsilon$-net of the state space $\mathcal{X}$. This is a finite set of points
+$\mathcal{S}=\{s^{(1)},\dots,s^{(N)}\}$
+such that for every $x\in\mathcal{X}$, there exists some $s\in\mathcal{S}$ with $\|x-s\|\leq\varepsilon$. Its cardinality satisfies
+$N=O\!\left((R/\varepsilon)^{d_{x}}\right).$
+We define a lifted state space of dimension $N$, where the lifted state $\mathbf{z}_{t}\in{\mathbb{R}}^{N}$ is a one-hot vector encoding the current approximate grid location.
+Let $\pi(x)$ denote a nearest grid point to $x$ in $\mathcal{S}$, with ties broken arbitrarily:
+$\pi(x)\in\operatorname*{\arg\min}_{s\in\mathcal{S}}\|x-s\|.$
+We now define a linear transition matrix $A^{\prime}\in{\mathbb{R}}^{N\times N}$ and observation matrix $C^{\prime}\in{\mathbb{R}}^{d_{y}\times N}$ as follows:
+  1. 1.
+    Dynamics ($A^{\prime}$): If the system is at grid point $s^{(j)}$, the next state is approximated by the grid point nearest to $f(s^{(j)})$. Thus
+    $A^{\prime}_{ij}=1\quad\text{if}\quad s^{(i)}=\pi(f(s^{(j)})),\qquad A^{\prime}_{ij}=0\quad\text{otherwise.}$
+  2. 2.
+    Observations ($C^{\prime}$): The $j$-th column of $C^{\prime}$ is the observation vector at the grid point $s^{(j)}$:
+    $C^{\prime}_{:,j}=h(s^{(j)}).$
+This defines a linear system
+$\mathbf{z}_{t+1}=A^{\prime}\mathbf{z}_{t},\qquad\mathbf{y}^{\prime}_{t}=C^{\prime}\mathbf{z}_{t}.$
+If $s^{(j)}=\pi(\mathbf{x}_{0})$ and we initialize $\mathbf{z}_{0}=\mathbf{e}_{j}$, then $\mathbf{z}_{t}$ remains a one-hot vector for all $t$.
+13.1.2 Approximation Guarantee
+We now prove that this high-dimensional linear system remains close to the true nonlinear trajectory over a finite horizon.
+Theorem 13.1 (Finite-Horizon Linear Approximation via Discretization).
+Let $\mathbf{y}_{0},\dots,\mathbf{y}_{T-1}$ be generated by the nonlinear system (13.1) under the boundedness and $1$-Lipschitz assumptions above.
+Let $A^{\prime},C^{\prime}$ be the matrices constructed from an $\varepsilon$-net $\mathcal{S}$, and let $\mathbf{y}^{\prime}_{t}$ be the output of the lifted linear system initialized at the one-hot vector corresponding to $\pi(\mathbf{x}_{0})$.
+Then for every $t=0,1,\dots,T-1$,
+$\|\mathbf{y}_{t}-\mathbf{y}^{\prime}_{t}\|\leq(t+1)\varepsilon.$
+Consequently,
+$\sum_{t=0}^{T-1}\|\mathbf{y}_{t}-\mathbf{y}^{\prime}_{t}\|^{2}\leq\varepsilon^{2}\sum_{t=0}^{T-1}(t+1)^{2}=\varepsilon^{2}\frac{T(T+1)(2T+1)}{6}\leq T^{3}\varepsilon^{2}.$
+Proof.
+Let $\mathbf{x}_{0},\dots,\mathbf{x}_{T}$ be the true trajectory of the nonlinear system. Let $s_{t}$ be the grid point occupied by the lifted linear system at time $t$. Then
+$s_{0}=\pi(\mathbf{x}_{0}),\qquad s_{t+1}=\pi(f(s_{t})).$
+Define the tracking error
+$\xi_{t}:=\|s_{t}-\mathbf{x}_{t}\|.$
+We analyze its evolution:
+$\xi_{t+1}$ $=\|s_{t+1}-\mathbf{x}_{t+1}\|$
+$=\|\pi(f(s_{t}))-f(\mathbf{x}_{t})\|.$
+By the triangle inequality,
+$\xi_{t+1}\leq\|\pi(f(s_{t}))-f(s_{t})\|+\|f(s_{t})-f(\mathbf{x}_{t})\|.$
+The first term is at most $\varepsilon$ by the definition of the $\varepsilon$-net and nearest-neighbor projection.
+The second term is at most $\|s_{t}-\mathbf{x}_{t}\|=\xi_{t}$ by the $1$-Lipschitz property of $f$.
+Hence
+$\xi_{t+1}\leq\xi_{t}+\varepsilon.$
+Since
+$\xi_{0}=\|\pi(\mathbf{x}_{0})-\mathbf{x}_{0}\|\leq\varepsilon,$
+an induction gives
+$\xi_{t}\leq(t+1)\varepsilon\qquad\text{for all }t\geq 0.$
+Now $\mathbf{y}_{t}=h(\mathbf{x}_{t})$ and $\mathbf{y}^{\prime}_{t}=h(s_{t})$, so by the $1$-Lipschitz property of $h$,
+$\|\mathbf{y}_{t}-\mathbf{y}^{\prime}_{t}\|=\|h(\mathbf{x}_{t})-h(s_{t})\|\leq\|\mathbf{x}_{t}-s_{t}\|=\xi_{t}\leq(t+1)\varepsilon.$
+This proves the pointwise bound. Summing the squared bound over $t=0,\dots,T-1$ yields
+$\sum_{t=0}^{T-1}\|\mathbf{y}_{t}-\mathbf{y}^{\prime}_{t}\|^{2}\leq\varepsilon^{2}\sum_{t=0}^{T-1}(t+1)^{2}=\varepsilon^{2}\frac{T(T+1)(2T+1)}{6}\leq T^{3}\varepsilon^{2}.$
+Thus, by choosing $\varepsilon\asymp T^{-3/2}$, the cumulative squared approximation error is $O(1)$ over the horizon $T$. The price is that the lifted dimension
+$N=O\!\left((R/\varepsilon)^{d_{x}}\right)$
+may be exponentially large in the original state dimension $d_{x}$.
+13.2 The Koopman Operator
+The discretization argument above gives a finite-dimensional linear approximation over a fixed horizon. The Koopman perspective explains why linear structure is present more fundamentally.
+The Koopman operator $\mathcal{K}$ is an infinite-dimensional linear operator acting on scalar observable functions $g:\mathcal{X}\to{\mathbb{R}}$ by
+$(\mathcal{K}g)(x)=g(f(x)).$
+Even if $f$ is nonlinear, the operator $\mathcal{K}$ is linear.
+Thus every nonlinear dynamical system admits an exact linear representation at the level of observables, but in general this representation lives in an infinite-dimensional function space.
+The discretization in Section 13.1 should be viewed as a concrete finite-state approximation to this principle: by tracking which cell of an $\varepsilon$-net the system occupies, we obtain a finite-dimensional lifted linear system.
+For the purposes of prediction, our perspective is pragmatic.
+We do not need to recover the full infinite-dimensional Koopman operator.
+We only need to know that a finite-dimensional linear approximation exists and that this approximation can be learned competitively by the online algorithms developed earlier.
+13.3 Improper Learning of Nonlinear Dynamics
+We can now combine the existence result of Theorem 13.1 with the learning algorithms from Chapter 12.
+The key point is existential: the lifted linear system constructed above is used only as a surrogate comparator in the analysis.
+The learner never constructs the grid $\mathcal{S}$, the matrices $(A^{\prime},C^{\prime})$, or an observer gain $L$ explicitly.
+The lifted linear system can have enormous hidden dimension
+$N=O\!\left((R/\varepsilon)^{d_{x}}\right),$
+yet the regret bounds from Chapter 12 do not depend directly on that hidden dimension.
+This makes improper learning possible: we never manipulate the lifted state $\mathbf{z}_{t}$ explicitly, but instead compete with the predictions of the lifted linear surrogate on the observed sequence $\mathbf{y}_{t}$.
+Dimension-independence alone, however, is not enough.
+To apply the Luenberger reduction from Chapter 12, the lifted system must admit a stable, well-conditioned observer.
+This leads to the following complexity measure.
+Definition 13.2 (Nonlinear Luenberger Complexity).
+Fix $\varepsilon>0$ and a stability margin $\gamma\in(0,1)$. Define
+$Q^{\star}_{\gamma}(f,h;\varepsilon):=\inf_{\begin{subarray}{c}\mathcal{S}\ \text{an }\varepsilon\text{-net of }\mathcal{X}\\ \pi\ \text{a nearest-neighbor map}\end{subarray}}\;\inf_{\begin{subarray}{c}L\in{\mathbb{R}}^{N\times d_{y}}\\ A^{\prime}-LC^{\prime}\text{ diagonalizable}\\ \sigma(A^{\prime}-LC^{\prime})\subseteq D_{1-\gamma}\end{subarray}}Q(A^{\prime}-LC^{\prime}),$
+where $(A^{\prime},C^{\prime})$ are the lifted matrices induced by $(\mathcal{S},\pi)$ as in Section 13.1. If no feasible choice exists, we set
+$Q^{\star}_{\gamma}(f,h;\varepsilon)=+\infty.$
+This quantity asks whether some $\varepsilon$-resolution linearization of the nonlinear system can be stabilized by a Luenberger observer with moderate transient behavior.
+Feasibility is not automatic: unlike the setting of Chapter 12, the discretized pair $(A^{\prime},C^{\prime})$ need not be observable, and the nonlinear Luenberger program may therefore have value $\infty$.
+Because the setting is autonomous, the relevant comparator class now consists of stable linear predictors based only on past outputs.
+We now combine the discretization error with the linear regret bound.
+Theorem 13.3 (Improper Learning Guarantee).
+Fix $\varepsilon>0$ and $\gamma\in(0,1)$.
+Let $(A^{\prime},C^{\prime})$ be any lifted linear system obtained from an $\varepsilon$-net construction in Section 13.1, and suppose there exists a gain matrix $L\in{\mathbb{R}}^{N\times d_{y}}$ such that $A^{\prime}-LC^{\prime}$ is diagonalizable and
+$\sigma(A^{\prime}-LC^{\prime})\subseteq D_{1-\gamma}.$
+Let $\bar{\mathbf{y}}_{t}$ denote the stable surrogate sequence obtained by applying the Luenberger reduction of Chapter 12 to the lifted linear system $(A^{\prime},C^{\prime})$, and let $E_{\mathrm{burn\mbox{-}in}}(A^{\prime},C^{\prime},L)$ denote the corresponding burn-in term.
+If $\mathcal{A}$ is any online learning algorithm with regret $\mathrm{Regret}_{\mathcal{A}}(T)$ against the class of stable linear predictors with margin $\gamma$, then its predictions $\hat{\mathbf{y}}_{t}$ on the nonlinear sequence satisfy
+$\sum_{t=0}^{T-1}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}\leq\mathrm{Regret}_{\mathcal{A}}(T)+2T^{3}\varepsilon^{2}+2E_{\mathrm{burn\mbox{-}in}}(A^{\prime},C^{\prime},L).$
+Proof.
+Let $\mathbf{y}^{\prime}_{t}$ denote the output of the lifted linear system $(A^{\prime},C^{\prime})$ constructed from the $\varepsilon$-net. By Theorem 13.1,
+$\sum_{t=0}^{T-1}\|\mathbf{y}_{t}-\mathbf{y}^{\prime}_{t}\|^{2}\leq T^{3}\varepsilon^{2}.$
+Next, applying the reduction from Chapter 12 to the lifted linear system $(A^{\prime},C^{\prime})$ yields a stable surrogate sequence $\bar{\mathbf{y}}_{t}$ such that
+$\sum_{t=0}^{T-1}\|\mathbf{y}^{\prime}_{t}-\bar{\mathbf{y}}_{t}\|^{2}\leq E_{\mathrm{burn\mbox{-}in}}(A^{\prime},C^{\prime},L).$
+Since $\bar{\mathbf{y}}_{t}$ belongs to the comparator class of stable linear predictors with margin $\gamma$, the regret guarantee of $\mathcal{A}$ gives
+$\sum_{t=0}^{T-1}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}\leq\sum_{t=0}^{T-1}\|\mathbf{y}_{t}-\bar{\mathbf{y}}_{t}\|^{2}+\mathrm{Regret}_{\mathcal{A}}(T).$
+Finally, using
+$\|a+b\|^{2}\leq 2\|a\|^{2}+2\|b\|^{2},$
+with
+$a=\mathbf{y}_{t}-\mathbf{y}^{\prime}_{t},\qquad b=\mathbf{y}^{\prime}_{t}-\bar{\mathbf{y}}_{t},$
+we obtain
+$\sum_{t=0}^{T-1}\|\mathbf{y}_{t}-\bar{\mathbf{y}}_{t}\|^{2}$ $\leq 2\sum_{t=0}^{T-1}\|\mathbf{y}_{t}-\mathbf{y}^{\prime}_{t}\|^{2}+2\sum_{t=0}^{T-1}\|\mathbf{y}^{\prime}_{t}-\bar{\mathbf{y}}_{t}\|^{2}$
+$\leq 2T^{3}\varepsilon^{2}+2E_{\mathrm{burn\mbox{-}in}}(A^{\prime},C^{\prime},L).$
+Substituting this into the regret bound proves the claim. ∎
+In particular, choosing $\varepsilon\asymp T^{-3/2}$ makes the discretization term $O(1)$.
+Therefore, instantiating $\mathcal{A}$ with any stable-predictor learner from Chapter 12 yields sublinear regret whenever $Q^{\star}_{\gamma}(f,h;\varepsilon)$ and the burn-in term are moderate.
+The lifted dimension $N$ does not appear explicitly in the guarantee; instead, the difficulty is pushed into the observer complexity and burn-in term.
+13.4 Bibliographic Remarks
+The operator-theoretic viewpoint on nonlinear dynamics goes back to the work of [KOO31], who observed that nonlinear state evolution induces a linear evolution on observables.
+Modern Koopman theory, especially as it relates to model reduction, spectral analysis, and control, was further developed in [MEZ05, MEZ13].
+Useful modern overviews include the review of [OR21], the survey of [BBK+22], and the volume [MMS20].
+The finite-state discretization used in Section 13.1 is also closely related to the older literature on transfer operators and set-oriented numerics.
+In particular, our $\varepsilon$-net construction is more naturally viewed as an Ulam-type finite approximation than as a standard EDMD construction: one partitions the state space, represents the dynamics by a finite matrix on those cells, and then studies the resulting linear surrogate.
+See [ULA60], [FRO99], and the overview [FRO01] for this line of work.
+On the data-driven side, [SCH10] popularized Dynamic Mode Decomposition (DMD) in fluid mechanics, and [TRL+14] clarified its operator-theoretic interpretation.
+Extended Dynamic Mode Decomposition (EDMD), which replaces the identity observable by a user-chosen dictionary, was introduced by [WKR15].
+Related developments include generator-based methods such as gEDMD [KNP+20] and time-delay liftings [KKB+20].
+These methods aim to approximate Koopman eigenfunctions or invariant subspaces directly from data.
+A separate strand of the literature studies when nonlinear systems admit exact or nearly exact finite-dimensional Koopman representations.
+Finite-dimensional invariant subspaces for control were explored by [BBP+16], while learned Koopman coordinates and invariant subspaces were developed by [TKY17].
+This distinction is conceptually important for the present chapter: exact finite-dimensional Koopman closures are special, whereas the discretization argument here gives a universal but approximate lifting.
+For systems with inputs and control, operator-theoretic liftings lead naturally to linear or bilinear predictor models.
+Particularly relevant references are [PBK18] on Koopman operators with inputs and control, and [KM18] on linear predictors for nonlinear controlled systems and model predictive control.
+These works are natural points of comparison for the autonomous presentation adopted in this chapter.
+Finally, the traditional alternative to the improper-learning viewpoint taken here is nonlinear system identification, which attempts to estimate the nonlinear transition map or latent state-space model directly from data.
+Good entry points to the classical literature are the perspective article of [LJU10] and the state-space identification treatment of [SWN11].
+More recently, there has been significant progress in providing non-asymptotic, finite-sample guarantees for learning nonlinear dynamical systems directly from a single trajectory.
+For instance, [FSR20] establish optimal sample complexity bounds for learning generalized nonlinear dynamics without relying on strict mixing assumptions.
+Similarly, [SO22] develop noise-sensitive, uniform convergence guarantees for learning nonlinear systems via gradient descent.
+The specific improper-learning reduction developed in this chapter, which combines a lifted linear approximation with the Luenberger framework from Chapter 12, is based on [DBH25].
+13.5 Exercises
+13.1.
+Consider the proof of Theorem 13.1.
+How does the dimension $N$ of the approximating linear system scale with the dimension of the true state $d_{x}$?
+Why does this curse of dimensionality not appear explicitly in the regret bounds derived from Chapter 12?
+13.2. Theorem 13.1 assumed the dynamics $f$ are $1$-Lipschitz. Suppose instead that $f$ is $L$-Lipschitz with $L>1$.
+  1. 1.
+    Derive the recurrence satisfied by the tracking error $\xi_{t}$.
+  2. 2.
+    Show that $\xi_{t}$ can grow exponentially with $t$.
+  3. 3.
+    What does this imply about the grid scale $\varepsilon$ required to maintain low approximation error over a horizon $T$?
+13.3. Assume $x_{0}>0$, and consider the nonlinear system
+$x_{t+1}=x_{t}^{2},\qquad y_{t}=x_{t}.$
+  1. 1.
+    Is this system linear in the state variable $x$?
+  2. 2.
+    Show that the observable
+    $g(x)=\log x$
+    evolves linearly along the dynamics, in the sense that
+    $g(x_{t+1})=2g(x_{t}).$
+  3. 3.
+    Explain why this gives a one-dimensional Koopman-style lifting of the dynamics on the positive half-line.
+13.4.
+Consider the large-language-model example in the introduction to this chapter.
+Assume that the hidden state $\mathbf{x}_{t}$ contains the current internal activations together with the prefix, cache, and other information needed to generate the next token.
+  1. 1.
+    Formulate deterministic decoding as an autonomous nonlinear dynamical system of the form
+    $\mathbf{x}_{t+1}=f(\mathbf{x}_{t}),\qquad\mathbf{y}_{t}=h(\mathbf{x}_{t}),$
+    where $\mathbf{y}_{t}\in V$ is the emitted token.
+  2. 2.
+    Now consider randomized decoding, where the next token is sampled from a distribution determined by the current hidden state. A natural formulation is
+    $\mathbf{y}_{t}\sim p(\cdot\mid\mathbf{x}_{t}),\qquad\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{y}_{t}).$
+    Explain why this gives a stochastic nonlinear dynamical system, and rewrite it equivalently either in noise-driven form
+    $\mathbf{x}_{t+1}=F(\mathbf{x}_{t},\xi_{t}),\qquad\mathbf{y}_{t}=H(\mathbf{x}_{t},\xi_{t}),$
+    or in Markov-kernel form.
+  3. 3.
+    Suppose that, after an appropriate discretization or lifting, the relevant hidden-state space becomes finite.
+Explain why the evolution of the resulting distribution over lifted states is linear and governed by a stochastic matrix, rather than by a zero-one transition matrix as in the deterministic setting.
 Part IV Online Control with Partial Observation
-Chapter 12 Policy Classes for Partially Observed Systems
+Chapter 14 Policy Classes for Partially Observed Systems
 In previous chapters, we explored control and learning techniques in fully observed linear dynamical systems, where the system state is directly accessible to the controller.
@@
 Often, partially observed LDSs are defined with an additional noise term that corrupts the observations.
-Exercise 12.6 asks the reader to prove that even in such circumstances we may assume the previously stated form for partially observed systems without loss of any generality.
+Exercise 14.6 asks the reader to prove that even in such circumstances we may assume the previously stated form for partially observed systems without loss of any generality.
 Unlike in the fully observed setting, where controllers can react based on the full state, partial observability introduces fundamental limitations on policy design.
@@
 Although this class of policies is natural to consider, it turns out to be very inexpressive.
-Even for optimal control with stochastic noise, the optimal policy cannot be approximated by a linear observation-action policy (see Exercise 12.6).
+Even for optimal control with stochastic noise, the optimal policy cannot be approximated by a linear observation-action policy (see Exercise 14.6).
 Therefore, we now define more expressive and useful policies.
@@
 $\ \ \tilde{C}=\begin{array}[]{|c|c|c|c|}\hline\cr C&C&\dots&C\\ \hline\cr\end{array}\ \ ,\ \ \tilde{D}=\begin{array}[]{|c|}\hline\cr I\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}.$
-The following lemma relates the original and lifted system (for a proof, see exercise 12.6).
+The following lemma relates the original and lifted system (for a proof, see exercise 14.6).
 Lemma 14.3.
@@
 $\forall i\ ,\ \|A_{\pi}^{i}\|\leq\kappa(1-\delta)^{i}\ ,\ (\|B_{\pi}\|+\|C_{\pi}\|)\leq\kappa.$
-Note that LDCs are $\varepsilon$-approximated by generalized observation-action controllers. The proof is left as Exercise 12.6.
+Note that LDCs are $\varepsilon$-approximated by generalized observation-action controllers. The proof is left as Exercise 14.6.
 Lemma 14.5.
@@
 Thus, this policy class allows for provably efficient optimization and algorithms that we explore in the next chapter.
-We conclude with the following important lemma on an expression that is important to compute nature’s y’s. The proof of this lemma follows directly from the definition and is left as Exercise 12.6.
+We conclude with the following important lemma on an expression that is important to compute nature’s y’s. The proof of this lemma follows directly from the definition and is left as Exercise 14.6.
 Lemma 14.7.
@@
 14.3.1 Expressivity of Disturbance Response Controllers
-A DRC always produces controls that ensure that the system remains state-bounded. This can be shown analogously to Lemma 6.6 (see Exercise 12.6).
+A DRC always produces controls that ensure that the system remains state-bounded. This can be shown analogously to Lemma 6.6 (see Exercise 14.6).
 Although DRCs are contained within the policy class of DACs, they are still very expressive.
@@
 It can be viewed as a succinct state-space representation of the Youla parameterization [YJB76].
-For time-varying dynamics, the power of various policy classes was contrasted in [MGSH21].
+For time-varying dynamics, the power of various policy classes was contrasted in [MGS+21].
 14.6 Exercises
@@
 Also, consider using a hidden state $\mathbf{z}_{t}$ with a larger number of coordinates than $\mathbf{x}_{t}$.
-12.3.
+14.3.
   1. 1.
@@
 14.5. For any $\gamma$-stable partially observed system, prove that any DRC policy in $\Pi_{H,\gamma}^{D}$ is $\gamma^{2}$-stabilizing.
-12.6.
+14.6.
 Construct a partially observed linear dynamical system subject to Gaussian i.i.d.
@@
 Hint: Consider a 3-dimensional hidden state so that $u_{t}$ can be chosen as a function of $y_{t-1}$ to ensure zero cost at each time step.
-Chapter 13 Online Nonstochastic Control with Partial Observation
+Chapter 15 Online Nonstochastic Control with Partial Observation
 In the previous chapter, we introduced Disturbance Response Controllers (DRCs) as a structured class of policies for controlling partially observed linear dynamical systems (PO-LDS).
@@
 Online nonstochastic control with partial observability was studied in [SSH20], who defined the Gradient Response Controller as well as the policy class of Disturbance Response Controllers.
-The setting was further extended to time-varying linear dynamics in [MGSH21], where the expressive power of different classes of policy was also studied.
+The setting was further extended to time-varying linear dynamics in [MGS+21], where the expressive power of different classes of policy was also studied.
 The analogue of online nonstochastic control in the stochastic noise setting is called the Linear Quadraic Gaussian problem, and it is further limited by assuming quadratic noise.
 This classical problem in optimal control was proposed in Kalman’s work [KAL60].
-Regret bounds for controlling unknown partially-observed systems were studied in [LAHA20b, LAHA20c, MTR19].
-[LAHA20a], which use DRC policies even in the context of stochastic control, obtain rate-optimal logarithmic bounds for the LQG problem.
+Regret bounds for controlling unknown partially-observed systems were studied in [LAH+20b, LAH+20c, MTR19].
+[LAH+20a], which use DRC policies even in the context of stochastic control, obtain rate-optimal logarithmic bounds for the LQG problem.
 In [SNH23], the task of controlling a partially-observed system subject to semi-adversarial noise under bandit feedback was considered.
-The tight rates for quadratic control with bandit feedback, in the fully adversarial model, were obtained in [SSNH24], and more general cost functions in the bandit setting were considered in [SL24].
+The tight rates for quadratic control with bandit feedback, in the fully adversarial model, were obtained in [SSN+24], and more general cost functions in the bandit setting were considered in [SL24].
 15.5 Exercises
@@
 $\mathbf{y}^{\mathbf{nat}}_{t}$ $=\mathbf{y}_{t}-C_{t}\mathbf{z}_{t}.$
-13.2.
+15.2.
 Consider a partially observed linear dynamical system that is $\gamma$-stable in the absence of control inputs.
 Prove that when a DRC policy $\pi\in\Pi^{R}_{H,\gamma}$ is executed on such a dynamical system, the resulting system is $\gamma^{2}$-stable.
-13.3.
+15.3.
 Prove that the optimal controller that minimizes the aggregate cost for a partially observed linear dynamical system assuming that the perturbations are stochastic is a linear dynamical controller.
@@
 Hint: Follow the proof of Theorem 15.1 and modify as required.
+Appendix A Second Order Online Learning with the Square Loss
+This appendix presents second order methods for the special case of online convex optimization with the square loss.
+The first algorithm is a specialization of the Online Newton Step (ONS) algorithm for the square loss.
+ONS operates strictly within the standard Online Convex Optimization (OCO) protocol, meaning no feature vector is revealed to the learner prior to making a decision.
+In addition, ONS gracefully handles arbitrary constrained convex sets for the learned parameters.
+The proof follows the same potential-based argument as in [HAZ16, Theorem 4.5], but exploits the exact quadratic structure of square loss.
+The second method we describe is the classical VAW algorithm, which does not strictly fall within the framework of online convex optimization due to two limitations.
+First, it requires knowing the feature vector before prediction, and secondly, it does not allow for constraints.
+However, in many control scenarios studied in this text, these limitations are acceptable, and in return it offers superior regret bounds in some respects.
+A.1 Setting
+We consider the standard Online Convex Optimization (OCO) protocol. At each round $t=1,\dots,T$:
+  • The learner chooses a parameter $\mathbf{x}_{t}\in\mathcal{K}\subseteq{\mathbb{R}}^{d}$, where $\mathcal{K}$ is a closed convex set with diameter $D$.
+  • The environment reveals a feature vector $\mathbf{a}_{t}\in{\mathbb{R}}^{d}$ such that $\|\mathbf{a}_{t}\|\leq R$, and a true label $b_{t}\in{\mathbb{R}}$, defining the loss function $\ell_{t}(\mathbf{x})=\frac{1}{2}(\mathbf{x}^{\top}\mathbf{a}_{t}-b_{t})^{2}$.
+  • The learner suffers loss $\ell_{t}(\mathbf{x}_{t})$. Assume that the losses suffered by the learner are bounded by a constant $C>0$:
+    $\max_{t=1,\dots,T}\ell_{t}(\mathbf{x}_{t})=\max_{t}\frac{1}{2}(\mathbf{x}_{t}^{\top}\mathbf{a}_{t}-b_{t})^{2}\leq C.$
+Define regret against $\mathbf{x}^{\star}\in\mathcal{K}$:
+$\mathrm{{regret}}_{T}(\mathbf{x}^{\star})=\sum_{t=1}^{T}\ell_{t}(\mathbf{x}_{t})-\ell_{t}(\mathbf{x}^{\star}).$
+A.2 Algorithm: ONS for Square Loss
+Algorithm 11 ONS for Square Loss
+1: Input: regularization $\lambda>0$
+2: Initialize $A_{0}=\lambda I_{d}$, $\mathbf{x}_{1}\in\mathcal{K}$
+3:for $t=1$ to $T$ do
+4:  Play $\mathbf{x}_{t}$
+5:  Observe $\mathbf{a}_{t},b_{t}$ and suffer loss $\ell_{t}(\mathbf{x}_{t})=\frac{1}{2}(\mathbf{x}_{t}^{\top}\mathbf{a}_{t}-b_{t})^{2}$
+6:  Compute gradient $\nabla_{t}=\nabla\ell_{t}(\mathbf{x}_{t})=(\mathbf{x}_{t}^{\top}\mathbf{a}_{t}-b_{t})\mathbf{a}_{t}$
+7:  Update curvature matrix:
+$A_{t}=A_{t-1}+\mathbf{a}_{t}\mathbf{a}_{t}^{\top}$
+8:  Update parameter:
+$\mathbf{x}_{t+1}=\mathop{\Pi}_{\mathcal{K}}^{A_{t}}\left(\mathbf{x}_{t}-A_{t}^{-1}\nabla_{t}\right)$
+9:end for
+A.3 Regret Bound
+Theorem A.1 (Logarithmic regret for square loss).
+For Algorithm 11, assuming losses are bounded by $C$ and $\mathcal{K}$ has diameter $D$, for all $\mathbf{x}^{\star}\in\mathcal{K}$,
+$\mathrm{{regret}}_{T}(\mathbf{x}^{\star})\leq\frac{\lambda}{2}D^{2}+{C}\log\det\!\left(I_{d}+\frac{1}{\lambda}\sum_{t=1}^{T}\mathbf{a}_{t}\mathbf{a}_{t}^{\top}\right).$
+In particular, setting $\lambda=1/D^{2}$, and $d,C,T,R,D>2$ we get
+$\mathrm{{regret}}_{T}(\mathbf{x}^{\star})\leq\frac{\lambda}{2}D^{2}+{dC}\log\!\left(1+\frac{TR^{2}}{\lambda}\right)\leq 4dC\log\!\left(TRD\right).$
+Proof.
+Fix $\mathbf{x}^{\star}\in\mathcal{K}$. Let $r_{t}=\mathbf{x}_{t}^{\top}\mathbf{a}_{t}-b_{t}$ denote the prediction residual. For square loss we have the exact identity
+$\ell_{t}(\mathbf{x}_{t})-\ell_{t}(\mathbf{x}^{\star})=\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})-\frac{1}{2}((\mathbf{x}_{t}-\mathbf{x}^{\star})^{\top}\mathbf{a}_{t})^{2}.$
+Indeed, expanding the difference in losses yields
+$\ell_{t}(\mathbf{x}_{t})-\ell_{t}(\mathbf{x}^{\star})$ $=\frac{1}{2}\left(r_{t}^{2}-(r_{t}-(\mathbf{x}_{t}-\mathbf{x}^{\star})^{\top}\mathbf{a}_{t})^{2}\right)$
+$=r_{t}\,(\mathbf{x}_{t}-\mathbf{x}^{\star})^{\top}\mathbf{a}_{t}-\frac{1}{2}((\mathbf{x}_{t}-\mathbf{x}^{\star})^{\top}\mathbf{a}_{t})^{2}$
+$=\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})-\frac{1}{2}((\mathbf{x}_{t}-\mathbf{x}^{\star})^{\top}\mathbf{a}_{t})^{2}.$
+We define the potential $\Phi_{t}=\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|_{A_{t-1}}^{2}$. By the generalized Pythagorean theorem for projection onto a convex set in the $A_{t}$-norm,
+$\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|_{A_{t}}^{2}\leq\|\mathbf{x}_{t}-A_{t}^{-1}\nabla_{t}-\mathbf{x}^{\star}\|_{A_{t}}^{2}.$
+Expanding the right-hand side gives
+$\|\mathbf{x}_{t}-A_{t}^{-1}\nabla_{t}-\mathbf{x}^{\star}\|_{A_{t}}^{2}=\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|_{A_{t}}^{2}-2\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})+\nabla_{t}^{\top}A_{t}^{-1}\nabla_{t}.$
+Rearranging this inequality to bound the linear term yields
+$\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})\leq\frac{1}{2}\left(\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|_{A_{t}}^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|_{A_{t}}^{2}\right)+\frac{1}{2}\nabla_{t}^{\top}A_{t}^{-1}\nabla_{t}.$
+Since the curvature matrix is updated via $A_{t}=A_{t-1}+\mathbf{a}_{t}\mathbf{a}_{t}^{\top}$, we can relate the norms as
+$\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|_{A_{t}}^{2}=\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|_{A_{t-1}}^{2}+((\mathbf{x}_{t}-\mathbf{x}^{\star})^{\top}\mathbf{a}_{t})^{2}.$
+Plugging this expansion into our previous inequality and moving the resulting quadratic term to the left-hand side gives
+$\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})-\frac{1}{2}((\mathbf{x}_{t}-\mathbf{x}^{\star})^{\top}\mathbf{a}_{t})^{2}\leq\frac{1}{2}\left(\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|_{A_{t-1}}^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|_{A_{t}}^{2}\right)+\frac{1}{2}\nabla_{t}^{\top}A_{t}^{-1}\nabla_{t}.$
+By the exact identity established above, the left-hand side is exactly the instantaneous regret $\ell_{t}(\mathbf{x}_{t})-\ell_{t}(\mathbf{x}^{\star})$.
+Summing over $t=1$ to $T$ telescopes the potential terms:
+$\mathrm{{regret}}_{T}(\mathbf{x}^{\star})\leq\frac{1}{2}\|\mathbf{x}_{1}-\mathbf{x}^{\star}\|_{A_{0}}^{2}+\frac{1}{2}\sum_{t=1}^{T}\nabla_{t}^{\top}A_{t}^{-1}\nabla_{t}.$
+Since $A_{0}=\lambda I_{d}$, we have $\frac{1}{2}\|\mathbf{x}_{1}-\mathbf{x}^{\star}\|_{A_{0}}^{2}=\frac{\lambda}{2}\|\mathbf{x}^{\star}-\mathbf{x}_{1}\|_{2}^{2}\leq\frac{\lambda}{2}D^{2}$, where the last inequality uses the assumption that $\mathcal{K}$ has diameter $D$.
+Using $\nabla_{t}=(\mathbf{x}_{t}^{\top}\mathbf{a}_{t}-b_{t})\mathbf{a}_{t}$, we can rewrite the gradient term as $\nabla_{t}^{\top}A_{t}^{-1}\nabla_{t}=2\ell_{t}(\mathbf{x}_{t})(\mathbf{a}_{t}^{\top}A_{t}^{-1}\mathbf{a}_{t})$.
+By our assumption that the losses are bounded by $C$, we have $\ell_{t}(\mathbf{x}_{t})\leq C$, and thus
+$\sum_{t=1}^{T}\nabla_{t}^{\top}A_{t}^{-1}\nabla_{t}\leq 2C\sum_{t=1}^{T}\mathbf{a}_{t}^{\top}A_{t}^{-1}\mathbf{a}_{t}.$
+By the matrix determinant lemma, $A_{t}=A_{t-1}+\mathbf{a}_{t}\mathbf{a}_{t}^{\top}$, and the vector harmonic series lemma from [HAZ16] Theorem 4.5, we have
+$\sum_{t}\mathbf{a}_{t}^{\top}A_{t}^{-1}\mathbf{a}_{t}=\sum_{t}A_{t}^{-1}\bullet(A_{t}-A_{t-1})\leq\log\frac{\det(A_{T})}{\det(A_{0})}.$
+Thus, the regret is bounded by
+$\mathrm{{regret}}_{T}(\mathbf{x}^{\star})\leq\frac{\lambda}{2}D^{2}+{C}\log\frac{\det(A_{T})}{\det(A_{0})}.$
+To explicitly bound the log-determinant ratio, we bound the determinant of $A_{T}$ using its trace.
+Since $A_{T}=\lambda I_{d}+\sum_{t=1}^{T}\mathbf{a}_{t}\mathbf{a}_{t}^{\top}$, the trace is linearly bounded:
+$\text{Tr}(A_{T})=\text{Tr}(\lambda I_{d})+\sum_{t=1}^{T}\|\mathbf{a}_{t}\|_{2}^{2}\leq\lambda d+TR^{2}.$
+By the AM-GM inequality, the determinant of a positive definite matrix is bounded by the average of its eigenvalues raised to the dimension $d$, which gives:
+$\det(A_{T})\leq\left(\frac{\text{Tr}(A_{T})}{d}\right)^{d}\leq\left(\frac{\lambda d+TR^{2}}{d}\right)^{d}\leq(\lambda+TR^{2})^{d}.$
+Since $A_{0}=\lambda I_{d}$, we have $\det(A_{0})=\lambda^{d}$. Taking the logarithm of the ratio yields:
+$\log\frac{\det(A_{T})}{\det(A_{0})}\leq\log\frac{(\lambda+TR^{2})^{d}}{\lambda^{d}}=d\log\left(1+\frac{TR^{2}}{\lambda}\right).$
+Substituting this upper bound into our regret inequality yields the final stated bound.
+A.4 The VAW Algorithm for Online Least Squares
+We explicitly compare this approach to the well-known Vovk-Azoury-Warmuth (VAW) forecaster [VOV01, AW01] to highlight two critical distinctions.
+First, ONS operates strictly within the standard Online Convex Optimization (OCO) protocol, meaning no feature vector is revealed to the learner prior to making a decision.
+Second, ONS gracefully handles arbitrary constrained convex sets $\mathcal{K}$, whereas VAW is natively an unconstrained algorithm.
+The projection step $\mathop{\Pi}_{\mathcal{K}}^{A_{t}}$ in Algorithm 11 plays a crucial mathematical role in keeping the algorithm stable within the strict OCO protocol.
+To see why, consider the unprojected version of the algorithm where $\mathcal{K}={\mathbb{R}}^{d}$.
+The parameter update simplifies to:
+$\mathbf{x}_{t+1}=\mathbf{x}_{t}-A_{t}^{-1}\mathbf{a}_{t}(\mathbf{x}_{t}^{\top}\mathbf{a}_{t}-b_{t}).$
+Multiplying both sides by $A_{t}=A_{t-1}+\mathbf{a}_{t}\mathbf{a}_{t}^{\top}$ gives:
+$A_{t}\mathbf{x}_{t+1}=(A_{t-1}+\mathbf{a}_{t}\mathbf{a}_{t}^{\top})\mathbf{x}_{t}-\mathbf{a}_{t}\mathbf{a}_{t}^{\top}\mathbf{x}_{t}+b_{t}\mathbf{a}_{t}=A_{t-1}\mathbf{x}_{t}+b_{t}\mathbf{a}_{t}.$
+By induction (and assuming $A_{0}\mathbf{x}_{1}=0$), we obtain:
+$\mathbf{x}_{t+1}=A_{t}^{-1}\sum_{s=1}^{t}b_{s}\mathbf{a}_{s}.$
+The vulnerability of unprojected ONS is that, because the decision $\mathbf{x}_{t}$ must be made before seeing $\mathbf{a}_{t}$, the resulting prediction $\mathbf{x}_{t}^{\top}\mathbf{a}_{t}$ can be overly aggressive.
+A.4.1 The VAW Forecaster
+The Vovk-Azoury-Warmuth (VAW) forecaster achieves logarithmic regret without requiring projections, but it must relax the strict OCO protocol to do so.
+In the VAW framework, the unlabelled feature $\mathbf{a}_{t}$ is revealed before the learner predicts.
+Algorithm 12 Vovk-Azoury-Warmuth (VAW) Forecaster
+1: Input: regularization $\lambda>0$
+2: Initialize $A_{0}=\lambda I_{d}$, target vector $\mathbf{v}_{0}=\mathbf{0}\in{\mathbb{R}}^{d}$
+3:for $t=1$ to $T$ do
+4:  Observe feature vector $\mathbf{a}_{t}$
+5:  Update curvature matrix: $A_{t}=A_{t-1}+\mathbf{a}_{t}\mathbf{a}_{t}^{\top}$
+6:  Predict $\hat{b}_{t}=\mathbf{a}_{t}^{\top}A_{t}^{-1}\mathbf{v}_{t-1}$
+7:  Observe true label $b_{t}$ and suffer loss $\ell_{t}(\hat{b}_{t})=\frac{1}{2}(\hat{b}_{t}-b_{t})^{2}$
+8:  Update target vector: $\mathbf{v}_{t}=\mathbf{v}_{t-1}+b_{t}\mathbf{a}_{t}$
+9:end for
+Because $\mathbf{a}_{t}$ is known in advance, VAW exploits this by updating the precision matrix with the current feature before committing to a prediction.
+By the Sherman-Morrison formula, $A_{t}^{-1}\mathbf{a}_{t}=\frac{A_{t-1}^{-1}\mathbf{a}_{t}}{1+\mathbf{a}_{t}^{\top}A_{t-1}^{-1}\mathbf{a}_{t}}$.
+Consequently, the VAW prediction can be written in terms of the unprojected ONS prediction (where $\mathbf{x}_{t}=A_{t-1}^{-1}\mathbf{v}_{t-1}$):
+$\hat{b}_{t}^{\text{VAW}}=\frac{\mathbf{x}_{t}^{\top}\mathbf{a}_{t}}{1+\mathbf{a}_{t}^{\top}A_{t-1}^{-1}\mathbf{a}_{t}}.$
+This demonstrates that VAW fundamentally shrinks the standard unprojected Ridge Regression prediction by the factor $(1+\mathbf{a}_{t}^{\top}A_{t-1}^{-1}\mathbf{a}_{t})$.
+Whenever the model is highly uncertain in the direction of $\mathbf{a}_{t}$, the denominator grows and shrinks the prediction.
+This implicit regularization automatically bounds the loss relative to the true labels, entirely bypassing the need for worst-case domain or magnitude bounds.
+Theorem A.2 (Regret of the VAW Forecaster).
+For Algorithm 12, assume the true labels are bounded such that $|b_{t}|\leq Y$ for all $t=1,\dots,T$. For any benchmark $\mathbf{x}^{\star}\in{\mathbb{R}}^{d}$, the regret is bounded by:
+$\sum_{t=1}^{T}\frac{1}{2}(\hat{b}_{t}-b_{t})^{2}-\sum_{t=1}^{T}\frac{1}{2}(\mathbf{x}^{\star}{}^{\top}\mathbf{a}_{t}-b_{t})^{2}\leq\frac{\lambda}{2}\|\mathbf{x}^{\star}\|_{2}^{2}+\frac{Y^{2}}{2}\log\det\!\left(I_{d}+\frac{1}{\lambda}\sum_{t=1}^{T}\mathbf{a}_{t}\mathbf{a}_{t}^{\top}\right).$
+In particular, if $\|\mathbf{a}_{t}\|_{2}\leq R$, and setting $\lambda=1/\|\mathbf{x}^{\star}\|^{2}$, with $d,Y,T,R>2$ we get:
+$\mathrm{{regret}}_{T}(\mathbf{x}^{\star})\leq\frac{\lambda}{2}\|\mathbf{x}^{\star}\|_{2}^{2}+\frac{dY^{2}}{2}\log\!\left(1+\frac{TR^{2}}{\lambda}\right)\leq 4dY^{2}\log(TR\|\mathbf{x}^{*}\|).$
+Comparing Theorem A.1 to Theorem A.2 highlights a distinct advantage of VAW: the regret bound depends on $Y^{2}$, which is determined strictly by the maximum magnitude of the observed labels.
+In contrast, the ONS bound depends on $C$, which scales with both the diameter of the decision set $\mathcal{K}$ and the feature norm $R$.
+However, this elegant shrinkage in VAW relies on the domain being unconstrained (${\mathbb{R}}^{d}$) and requires violating the strict zero-lookahead OCO timeline, making projected ONS the necessary algorithm when dealing with strict constraints or true zero-lookahead settings.
+A.4.2 VAW with Known Additive Offsets
+We now consider a shifted prediction setting in which, at each round $t$, the learner observes both the feature vector $\mathbf{a}_{t}\in{\mathbb{R}}^{d}$ and a known scalar offset $c_{t}\in{\mathbb{R}}$ before making a prediction.
+The learner then predicts $\hat{z}_{t}\in{\mathbb{R}}$, after which the environment reveals the true label $z_{t}\in{\mathbb{R}}$, and the learner suffers square loss
+$\ell_{t}(\hat{z}_{t})=\frac{1}{2}(\hat{z}_{t}-z_{t})^{2}.$
+We compare against benchmarks of the form $\mathbf{x}^{\star}{}^{\top}\mathbf{a}_{t}+c_{t}$, i.e. predictors that are allowed to use the same known offset:
+$\mathrm{{regret}}_{T}^{\mathrm{off}}(\mathbf{x}^{\star})=\sum_{t=1}^{T}\frac{1}{2}(\hat{z}_{t}-z_{t})^{2}-\sum_{t=1}^{T}\frac{1}{2}(\mathbf{x}^{\star}{}^{\top}\mathbf{a}_{t}+c_{t}-z_{t})^{2}.$
+The key point is that, since $c_{t}$ is known prior to prediction, one can simply subtract it from the labels and run the usual VAW forecaster on the centered targets.
+Algorithm 13 VAW Forecaster with Known Additive Offsets
+1: Input: regularization $\lambda>0$
+2: Initialize $A_{0}=\lambda I_{d}$, target vector $\mathbf{v}_{0}=\mathbf{0}\in{\mathbb{R}}^{d}$
+3:for $t=1$ to $T$ do
+4:  Observe feature vector $\mathbf{a}_{t}$ and known offset $c_{t}$
+5:  Update curvature matrix:
+$A_{t}=A_{t-1}+\mathbf{a}_{t}\mathbf{a}_{t}^{\top}$
+6:  Predict the centered label:
+$\hat{y}_{t}=\mathbf{a}_{t}^{\top}A_{t}^{-1}\mathbf{v}_{t-1}$
+7:  Output the shifted prediction:
+$\hat{z}_{t}=\hat{y}_{t}+c_{t}$
+8:  Observe true label $z_{t}$ and suffer loss
+$\ell_{t}(\hat{z}_{t})=\frac{1}{2}(\hat{z}_{t}-z_{t})^{2}$
+9:  Update target vector using the centered label:
+$\mathbf{v}_{t}=\mathbf{v}_{t-1}+(z_{t}-c_{t})\mathbf{a}_{t}$
+10:end for
+Theorem A.3 (VAW regret with known additive offsets).
+For Algorithm 13, assume the centered labels are bounded:
+$|z_{t}-c_{t}|\leq Y\qquad\text{for all }t=1,\dots,T.$
+Then for any benchmark $\mathbf{x}^{\star}\in{\mathbb{R}}^{d}$,
+$\sum_{t=1}^{T}\frac{1}{2}(\hat{z}_{t}-z_{t})^{2}-\sum_{t=1}^{T}\frac{1}{2}(\mathbf{x}^{\star}{}^{\top}\mathbf{a}_{t}+c_{t}-z_{t})^{2}\leq\frac{\lambda}{2}\|\mathbf{x}^{\star}\|_{2}^{2}+\frac{Y^{2}}{2}\log\det\!\left(I_{d}+\frac{1}{\lambda}\sum_{t=1}^{T}\mathbf{a}_{t}\mathbf{a}_{t}^{\top}\right).$
+In particular, the regret bound depends only on the centered labels $z_{t}-c_{t}$, and has no dependence on the magnitude of the offsets $c_{t}$.
+If furthermore $\|\mathbf{a}_{t}\|_{2}\leq R$ for all $t$, then
+$\sum_{t=1}^{T}\frac{1}{2}(\hat{z}_{t}-z_{t})^{2}-\sum_{t=1}^{T}\frac{1}{2}(\mathbf{x}^{\star}{}^{\top}\mathbf{a}_{t}+c_{t}-z_{t})^{2}\leq\frac{\lambda}{2}\|\mathbf{x}^{\star}\|_{2}^{2}+\frac{dY^{2}}{2}\log\!\left(1+\frac{TR^{2}}{\lambda}\right).$
+Proof.
+Define the centered labels and centered predictions by
+$y_{t}:=z_{t}-c_{t},\qquad\hat{y}_{t}:=\hat{z}_{t}-c_{t}.$
+By construction of Algorithm 13, we have
+$\hat{y}_{t}=\mathbf{a}_{t}^{\top}A_{t}^{-1}\mathbf{v}_{t-1},\qquad\mathbf{v}_{t}=\mathbf{v}_{t-1}+y_{t}\mathbf{a}_{t}.$
+Therefore, when written in terms of $(\mathbf{a}_{t},y_{t})$, Algorithm 13 is exactly the standard VAW forecaster of Algorithm 12 applied to the centered labels $y_{t}$.
+Since $|y_{t}|=|z_{t}-c_{t}|\leq Y$, Theorem A.2 implies that for every $\mathbf{x}^{\star}\in{\mathbb{R}}^{d}$,
+$\sum_{t=1}^{T}\frac{1}{2}(\hat{y}_{t}-y_{t})^{2}-\sum_{t=1}^{T}\frac{1}{2}(\mathbf{x}^{\star}{}^{\top}\mathbf{a}_{t}-y_{t})^{2}\leq\frac{\lambda}{2}\|\mathbf{x}^{\star}\|_{2}^{2}+\frac{Y^{2}}{2}\log\det\!\left(I_{d}+\frac{1}{\lambda}\sum_{t=1}^{T}\mathbf{a}_{t}\mathbf{a}_{t}^{\top}\right).$
+It remains to relate this centered regret to the original shifted problem. By definition of $y_{t}$ and $\hat{y}_{t}$,
+$\hat{z}_{t}=\hat{y}_{t}+c_{t},\qquad z_{t}=y_{t}+c_{t}.$
+Hence the square loss is exactly translation invariant:
+$\frac{1}{2}(\hat{z}_{t}-z_{t})^{2}=\frac{1}{2}\big((\hat{y}_{t}+c_{t})-(y_{t}+c_{t})\big)^{2}=\frac{1}{2}(\hat{y}_{t}-y_{t})^{2}.$
+Likewise, for any benchmark $\mathbf{x}^{\star}$,
+$\frac{1}{2}(\mathbf{x}^{\star}{}^{\top}\mathbf{a}_{t}+c_{t}-z_{t})^{2}=\frac{1}{2}\big(\mathbf{x}^{\star}{}^{\top}\mathbf{a}_{t}+c_{t}-(y_{t}+c_{t})\big)^{2}=\frac{1}{2}(\mathbf{x}^{\star}{}^{\top}\mathbf{a}_{t}-y_{t})^{2}.$
+Substituting these identities into the previous bound yields
+$\sum_{t=1}^{T}\frac{1}{2}(\hat{z}_{t}-z_{t})^{2}-\sum_{t=1}^{T}\frac{1}{2}(\mathbf{x}^{\star}{}^{\top}\mathbf{a}_{t}+c_{t}-z_{t})^{2}\leq\frac{\lambda}{2}\|\mathbf{x}^{\star}\|_{2}^{2}+\frac{Y^{2}}{2}\log\det\!\left(I_{d}+\frac{1}{\lambda}\sum_{t=1}^{T}\mathbf{a}_{t}\mathbf{a}_{t}^{\top}\right),$
+which proves the first claim.
+The simplified bound under $\|\mathbf{a}_{t}\|_{2}\leq R$ follows exactly as in Theorem A.2, by bounding the log-determinant term via
+$\log\det\!\left(I_{d}+\frac{1}{\lambda}\sum_{t=1}^{T}\mathbf{a}_{t}\mathbf{a}_{t}^{\top}\right)\leq d\log\!\left(1+\frac{TR^{2}}{\lambda}\right).$
+Thus the regret bound is completely independent of the magnitude of the offsets $c_{t}$; only the centered labels $z_{t}-c_{t}$ matter. ∎